A Complete Bibliography of Publications in Statistics and Probability Letters: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.06

Title word cross-reference

(−∞, +∞) [Sim09]. (7, 35, 15, 3, 5) [MGSB08]. (n, f, k) [CKLX06]. (p − 1) × 1 [Muk09]. (r, p) [CB07]. 1 [Asc09, TP01]. 1/2 [IKP08]. 12 [MS04a]. 16 [But05]. 2 [DS00b, Ete07, Joh02, KF08, ZAC03]. 2^{n_{1}} × 3^{n_{2}} [Lia01]. 2 × 2 [SS08c]. 64 [MT04]. [0, 1]^{d} [AR02]. _{2} [GV02a]. A [CGK06]. α [Bra02a, MM09, OK06, TG02, XYC09, ZL07b]. AR(p) [ZL09]. ARCH(p) [Che08b]. ARMA(p, q) [LH08b]. B [Den05, Shi04c]. β [BL09b, LS05a, LWD07]. C [Raj09, Sim07]. C_{p} [Kou03]. C_{pm} [Wri00]. χ^{2} [Hor04, Jia01b]. D [DMB02, Fan02b, Fan03, HP07, Hil00, Jen04, KS08b, Pro03, XNF07, AK09, CFR06, LHLC04]. D[0, 1] [Bez06, Pak08b]. δ [BB07b, LK07]. δ^{2} [KSG08]. E [But08, FRSWT08, MW00, SS03]. E(f_{NOD}) [KM05]. E(X^{k} · Y^{ℓ}) = E(X^{k}) · E(Y^{ℓ}) [BS06d]. ϵ [GVS00, Gri03, SV02]. F [Cac01, Sen01, vE00]. G [Fot05, Zha08, CP00, HHC09, Jia06, Jia09b, Wan09b, Zha09a]. G_{2} [But04]. Γ [WZL00]. Γ(n + 1, 1) [AJ05]. GARCH(p, q) [BHK04, LS05a]. H(n)

[YWWY08]. INAR(p) [SS06c]. K [ASG04, ARB09, DR04, DLB05, Dem08, GBQ03, Kli06, KH08a, Lal08, LBMD05, Lou03, MPV01, Mir05, Por03, Sas08, Ver07]. k_{T} [Lab08]. k ≥ 3 [CKS06b]. k ∈ (0, 1) [Won09]. l [JF00]. L^{1} [LLOS09, dlHRB05]. L^{2} [dlHRB05]. L^{p} [PS09c, PS09b]. L_{1} [BCG01, BBR05, BMR07, Cho05b, Esa04, Gir04, LJ01, Ole01, Fan02a]. L_{2} [CFCGM08, Lie01, NP00]. L_{p} [HZ03, HZ04b, QH09]. L_{q} [ML09a]. ⟨n, f, k⟩ [CKLX06]. M [BC01, BKL01, COS03, CT04, Ciu09, DW00, Fur02, GV06, KP02, OS05, TS09b, Tom09, RW00, SB00]. L^{p} [SZ05]. R^{d} [Ilt00, Fot05]. R^{n} [Nan08]. Z^{2} [Vol06]. D′(R^{⌈}) [MF05b]. L_{∞} [LOS02, LOS03]. R^{∈} [Bar05]. μ_{t+s} = μ_{s} ∗ T_{s}μ_{t} [SS01c]. N [CH09a, CM01, Mut08, AJ00, BA08, DG09, Kev07, NT06a, NP08a, SB00]. n − k + 1 [BA08]. OA(4r, r^{1}2^{p}, 2) [KyS05]. P [Raq03, Ara01, Bis00, BT02, CV01, CH07, Dah06, GVMSN04, HW07, HCSV00, Hu09, Khr01, KM02b, Kub08, Li06, Li09a, LC05, MC09, QH09, Ros09a, TT07a, TT07b, Wan07a, XW08, YYL08]. ϕ [AHV06, CHV09, Mal01]. p × 1 [Muk09]. Q [LPY04, CY02b, Li07, LV07, ZM06]. Q_{1} [MY09]. Q_{ϕ} [MY09]. R [FS08b, Ome08, Cha05, SW00b]. R = P (Y < X) [KR09]. r^{2} [Fit06]. ρ [CG08b, JH08]. ρ∗ [JD08c]. ρ– d [Shi04b]. s [AHL08]. s^{n−m} [AZ04a]. σ [CMNP07, Yan08]. sup [Gai07]. t [AT08c, BvdV08, BvdV09, BC05, DK02, GL08, GL00, HS00, IKIS09, Jon02b, KM03a, KM09a, Nas06, Tar05, Wat01]. τ [KNB00, MSW07]. θ [Hör03]. ρ̃ [WJ08]. U [AW06, BW08, DS09, DR02, DR15, HKM04, HZ06, JSV01, JGMGPM03, JYZ08, Kla09, Mav08, Nas09, Neu04, Pyc07, SS02, SW00a, XNF07, YJ01, ZS08]. V [FM00, JGMGPM03]. VAR(p) [TR02]. Z [CWTB09]. Z_{p}^{k} [Asc09]. Z_{n} [GII08].

-1 [Sch03]. -aberration [But04]. -adic [Khr01]. -algebras [Yan08]. -approach [LLOS09]. -ary [AJ00, Mut08]. -based [CY02b]. -bit [MT04]. -Capacities [CB07]. -contaminated [GVS00, SV02]. -convergence [QH09]. -deficiency [LOS03]. -demimartingales [CH09a]. -dependent [RW00, TP01]. -dimensional [CFR06, LHLC04]. -distribution [DK02]. -distributions [BvdV08, BvdV09]. -efficiency [Jen04]. -error [Esa04, Gir04, LOS02]. -estimates [CT04]. -Estimation [Fur02, BC01, Ciu09, DW00]. -estimator [COS03]. -estimators [GV06, KP02, Ome08, TS09b, Tom09]. -exchangeable [Ete07]. -expectation [Jia06, Jia09b]. -expectations [HHC09, Zha09a]. -factor [ZAC03]. -factors [KH08a, YWWY08]. -fields [CMNP07]. -functions [Li07]. -indeterminate [OS05]. -level [AHL08, DS00b]. -martingales [CP00, Wan09b]. -median [BCG01]. -method [Hör03]. -minimax [WZL00]. -mixing [BL09b, Bra02a, CG08b, CHV09, JH08, JD08c, LS05a, Mal01, Shi04b, WJ08, XYC09]. -mixingale [ML09a]. -model [ASG04]. -nearest [Lal08]. -norm [Fan02a, Raq03]. -norms [HZ03, HZ04b]. -occupation [Lab08]. -optimal [Ara01, But08, CGK06, DMB02, Fan02b, Fan03, Hil00, KS08b, MW00].

-optimality [FRSWT08, SS03, XNF07]. -optimum [HP07, Pro03]. -Paul [Kev07]. -power [BB07b]. -processes [AW06, Neu04, SW00a]. -quantiles [BKL01, JSV01, ZS08]. -rate [BMR07]. -records [ARB09]. -resolvable [OK06, KM02a]. -risk [PS09b, PS09c]. -run [But05, MS04a]. -sample [Sim07]. -scan [SW00b]. -semi-selfdecomposable [Raj09]. -semimartingales [Kub08]. -sequences [JF00, Joh02]. -shock [LK07]. -spacings [Mir05]. -squared [FS08b]. -stable [BT02, MM09]. -state [LV07]. -statistic [GL00, Kla09]. -statistics [DR15, BW08, DS09, DR02, HKM04, HZ06, JYZ08, Mav08, Nas09, Pyc07, SS02, YJ01]. -step [Ver07]. -subgaussian [AHV06]. -Subgeometric [FM00]. -symmetric [MPV01]. -test [GL08, Jia01b, Tar05, Cac01, CFCGM08]. -tests [Lie01, Sen01]. -time [Che08b]. -topology [MSW07]. -trees [CM01]. -tuple [Lou03]. -type [NP00, XNF07]. -uniformly [QH09]. -upcrossings [Gri03]. -value [Li09a, LC05, TT07a, XW08]. -valued [Den05, Shi04c]. -values [Bis00, Dah06, GVMSN04, KM02b, Li06, MC09, TT07b, Wan07a, Hu09]. -variation [CH07, YYL08]. -variations [Ros09a]. -Winsorized [TG02].

1 [Zel06]. 186/125 [KMvE00].

27 [Bro11].

31 [HV00]. 35 [Lin00]. 38 [HV00].

43 [Pal01b]. 46 [Oeh01]. 47 [dE00b].

51 [FT01a, Rou01a, vEZ02]. 57 [DR15].

63 [KM09a]. 68 [Duc10].

76 [Ano06o]. 77 [GV08, PK08, Yu09b]. 78 [BvdV09, Gat09, Hal09, Kim13, SF12]. 79 [HWYZ09a, PS09b].

A-optimal [HT05]. a.s [MF05b]. Aalen [DS01b, Kra04]. Abelian [CF00]. aberration [AZ04b, But04, HZ09a, LCLZ06, QC04, YZL07]. abilities [CCDG02]. above [FBVW00]. absence [HZ04a]. Absolute [YH08, AM02b, AP04, Dar01, Haa09, KL03c, MR09, MS09b, NS06b, SM09, YZL08]. absolutely [AVG08, LN03]. absorption [GLX07, GJZ06]. Acceleration [KSG08]. acceptance [AAHJ04]. accumulated [IMS06]. accumulation [GDF01]. Accuracy [Bis01, Bha08, Nov03]. Accurate [WN09b, WN09c]. ACD [Fer04]. acetylation [YPD09]. Acknowledgement [Bro11, HWYZ09a]. action [CKS04]. actions [CS00a]. active [VZ03]. Adaptive [Eni06, LdC07, Tri04, Aic09, BM08, BR04, BM01b, But00, DS01c, GP00b, HZ09b, JNR09, KSW06a, KSW06b, MPS06, PP01, YW07]. Addendum [HV00]. Adding [Tot04]. addition [EKM02]. additive [Che00a, CA00, Clé01, GT04, HS08a, Hat06, Ilt00, Kra04, MZ05, MY08, MN04, MRP04, PW08, RAS08b, SZ08a, Yeo07, YZ07b, Zap03]. additive-multiplicative [Kra04]. additivity [DV08, Jia06]. adic [Khr01]. adjusted [CM06, NS09, PD09, SN09]. adjustment [Coo02, FC08, Luc00, Wan02b]. admissible [JNS02]. advert [Ano01a]. affected [BNO03, Nik08]. affine [Asc09, Zaj09]. after [And05, BM08, CF01]. against [dBC06, Dub02, GJ02, KD04, Kla00, Nag09, Sen09, SGK09]. Age [IM05, YY07a, Ahm04, FH06, JT08]. age-dependent [JT08]. age-period-cohort [FH06]. ageing [BI02, LCJ06]. aggregated [CC08b, Sta05]. aging [BGHP04, CW05, NBA07, ST01b]. agreeing [GV02a]. ahead [GTH06]. AIDS [NPC05]. Aitken [KSG08]. Akaike [SC08]. al [BS04]. Algebraic [FS08a, DMB02, Ser02]. algebras [Yan08]. algorithm [Che09, DKM02, EZA04, FM00, GL03, Gut00, Hor04, Kar02, MOM07, SXL05]. Algorithmic [WH07]. algorithms [BSS01, Bet00, CLF06, Dev09, HP07, KSG08, Lel08, Pro03, SBF00]. aliasing [Man05a]. alignments [MGW02]. Allocating [DDM06]. allocation [Aic09, BM08, BS08b, BR04, HK08, KZS08, Ruk07b, TF05, VZ03]. allocations [LH08a]. allowing [Fuk05]. Almost [BW06a, CL06b, CL08b, Che02a, CG02c, HKM04, Lel08, Mon06, NZ01, Shi04b, Sta02, TP09, WZ06a, Bez08, Dud03, Dud08, Fah00, Gon08, HH03b, Hör06, Jam03, KKH08, KRV05, LW08a, LS01, LT00b, Mat05, RAS08b, SM06b, WL08a, GAW08]. Almost-sure [Che02a]. alone [Jan08]. along [VT00]. alpha [YAT09]. alpha-integrable [YAT09]. alternating [CP06a]. Alternative [BL07, ÖW00, Ars08, Ban08, FGMS08, MW00, SK06a, SS01b, Sen01, Sen09, SGK09]. alternatives [CP08, dBC06, Dub02, GJ02, Kla00, Voc06]. always [KN05]. American [CI09]. among [GS08b, NRS05, TV07]. amount [BLS08]. analogs [Oli01]. analyses [Li02, Sch02]. Analysis [Min09, AS07, AH01, AHS07, BR01, Bou06, Cho05b, DMP01b, DP00, DWC04, Esh04, FD05, FH06, GS03b, GLV00, Gut00, HCW07, Hug01, Iac01, IKAT09, IMS06, Jow06, IV00, LRM00, LL02, LB03, Mon08, Nki05, Omo07, Poo06, Ser00, Shu03, Sun00, SZ02b, TG02, Tri08b, WH07, YK03, Zan03]. Analytic [Bos09]. Analytical [MS02a, DW05]. analytics [AFS^{+}07]. Analyzing [ML03, ST09, KN00]. ancestor [MN09]. Andersen [LL09a, Ren09a, YH09]. Anderson [Man05b, Pyc03, TO03]. anisotropic [Hue03]. annealing [OW03]. announcement [Ano07t]. ANOCOVA [BdC00]. ANOVA [CKM03, Hut03, SN04, Wul08, XW08]. anticipative [MS08c]. any [MB08, PS07a]. appearance [HA03]. applicability [FJS03]. Application [Mas06a, Rao02b, AY02, AV01a, AAV02, BL09a, BE08, CKM03, CL08c, Deb09, Fot04, HKV04, HKS00, HNRV00, Jia01b, JO08, KS06, Kim00a, Lef08b, LB03, LC05, MY04, MMB07, PBH09, RW05, Rod00, Rou01a, Rou01b, YJ01, ZQ06, Zhe01, dlPZ02]. Applications [Jou07, Ahm02, Aue04, BBZ08, BS01a, BHH08, BHHM04, BR06c, Bro09, BP08, DaY08, Duc04, Duc10, FJ04, Hut00, KO01, KPI09, Kat09, Kle02, KvD05, LN09, Li03a, LS00c, LaYY03, MS08a, MPY03, Otm09, Pou08, Rom02, Sha07, Shi04c, SKPP03, SKS09, XYC09, YK00, YK03, YSY08, Zha07a, ZNS08]. applied [PPRW08]. applying [HCW07]. apportionment [SD06]. approach [BM03c, CLR08, CW01, CH09b, DF02, EKM02, EK03a, FW08b, FSX02, FSC03, GLML09, Gua07, Irl04, JF00, Kim01, KSH07, KN00, LLOS09, Lal08, LCZ09, Li09a, LH08b, Man05a, Mor05, PR09a, sS09c, SM06a, Zho09, dL05a, dlHRB05]. approaches [MGB05, Pro08, SK01]. approaching [dL05b]. Approximate [Cor08, HK09, MS08c, RC03, Rek09, Seg05, Bis06, CW07, GS08a, Gir09, Wu08b, YK05]. approximately [LLS06]. Approximating [GM00, JZ04]. Approximation [BH06, Bou06, CKS06a, Duc09, AVV09, Are03, Aue04, BJR00, Bar05, BM03c, Bis01, Bor00, BS01b, Cai00, hCW09, Car04, CLF06, DR01b, DG06b, FV08, GJ02, Gat08, Gat09, HH02b, HK03b, Hou05, KT03, Kol00a, LSEB00, Lia09b, LW03, LS08b, Mak08b, Mon08, NFL09, Nov03, PW00, Roo01, SG06, Sta00, SD01, UWF06, WSM09, Zha01a, hZbG08]. Approximations [CG02b, CTY02, HKS00, ADR05, DR01a, FF09, FR06, HMR08, KvD05, Nas09, SZ09b, SW00b, Vin08, XWS05]. arbitrage [Gap04, Stu00]. arbitrarily [Cat01]. arbitrary [Joh02, LY03, Nik08, Won09, YY08a]. arc [NO00]. arc-sine [NO00]. ARCH [Duc10, Sar00, AT08b, DG07, Duc04, HL02, HW01e, LJ01]. Archimedean [CS08b, WO08]. arcsine [SZ09a]. areas [Jur03]. Argmax [Fer09]. Argmax-stable [Fer09]. ARIMA [DDN08]. arising [Gut00, KS00, Lu06, RS06a, Zel06]. ARMA [BS09, FG04, GP07, KMS08, Lee00a, LS08a, Yi05, dJP04]. array [DDM06, Kan07, SSY02]. arrays [AAV02, BW07, BDM08, CFM01, DS00b, Dey05, EKK06, FM04, Has05, Has06b, HU06a, HCSV00, HSV98, HV00, IW04, KK06b, KVH06, Kuc04b, Kuc07, Kuc09, KyS05, Man05a, ML09a, ML09b, PLZ03, PZL04a, PZL04b, QH09, Sue03, SVH05, SHV05, Sun08a, Tóm05]. arrival [Bro08b, Led08b]. artificial [BL07]. ary [AJ00, Mut08]. Asian [CL06c]. aspect [Hua02]. assay [Zel06]. Assessing [CW01, JJ09, Kim03, MU01, Yat02, BV07, IWWJ04]. assessment [SW04]. assigned [Efr08]. assignment [Mid08]. assisted [WW01]. associated [BR02b, BS01b, hCW09, Che08a, CH09a, DR02, DR05, DR15, Fer03, Fie03, GRT07, HO08, HWYZ09a, HWYZ09b, HX02, Hua03b, HZ06, Hug01, JT08, KKH08, KK08a, KKC08, Kla09, Kuc09, LQ07, LYZ08, Li09b, Lia00a, Lin08c, Liu07c, LGLDG00, Mas02c, Mir07, Oli05, Rao02c, Sha08b, Sun07c, Sun08b, Tud08, Wan04, WT04, WCS07, WL08a, Yan03a, Zah00, ZW01]. Association [DD06, Li03a, CP05a, Esh04, GQR06, LW08a, Rou00, TV07]. assumption [CHV09, Coe08, Kim03, SH09a, SC08, Wan08a]. assumptions [HR08a, Hal09, HZ04a, Hut03, KK08b, xLxZ04, RM08, YAT09]. asymmetric [AT08a, DG07, GS09, Haa08, Hua02, JK07, JR08, KvD05, Nag03, WB05]. asymmetrical [Hua09]. asymmetrically [Lai01]. asymmetry [Bos07a, UJ09]. Asymptotic [AHEA04, AT08b, AR09b, Asc09, BS08a, BEA09a, BPvZ04, BW08, dCB00a, CM04a, Che04b, CYJL01a, CY02b, FJA09, FM05, FH06, GS07, GP00a, GN08, GVMSN04, GV02b, Has01, HCV05, HZ06, HP09, Jar03, KSW06a, IV00, KK04, KKAK08, KH01, KKH02, Lad07, Mir07, MC07b, MW06, Nag03, Nag09, Pen02a, PP01, Pro09b, Psa09a, Rou00, SL09a, SN09, SS00b, SKJ07, SS06c, VvdMD08, Wan02d, WL08b, WSM09, WW08, YC04, Bar06, BR01, BB02b, Bia02, BM05, CC08a, CBR00, DG03, DS02b, FS02, FE08, Fuj07, GQR06, Gir04, GPC09, HO09, Hür04, HW01e, Ign09, IKIS09, KMS08, LLOS09, LOS02, Leó01, LS06b, LYZ08, LQ04, MN09, McG06, NHO03, Nic08b, OH05, Per09, Psa02, Sch01a, Sco03, Sgi01a, Sgi01b, SS06b, Ten01, Ten03, TR02, UWF06, Var08]. asymptotic [Vou06, Woe03, XTW06, Yan03a, Yi05]. Asymptotic-power [IV00]. Asymptotically [GP06, HK08, GH00, Jar08, PP06]. Asymptotics [ARN08, Che08a, Deg08, Gel04, Has09, HZ03, HZ04b, Hua03a, Kos09a, MP01b, RW05, Sha08a, She08, She09, WLG01, AS01, BGM03, CLR02, DS01c, FL08, Fro03, Gan00, Ilt00, KP00, LL06, LTY08, NT06a, PP09, WZ03, YX06, ZY08]. atom [BR06a]. atom-on-demand [BR06a]. attracting [Has06b]. attraction [Has05, MR08b, Sch01b, Seg01, Shm06]. attractiveness [PG09]. Aumann [FM03a]. Author [Ano01a, Ano01b, Ano01c, Ano01e, Ano01f, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e]. auto [Gri02]. auto-Poisson [Gri02]. autocorrelated [Bha05a]. autocorrelation [Bha05a, CC08b, Gui01, Haa08, MR00, Yi05]. autocorrelations [Deb07, GH00]. autocovariance [LH08b]. autocovariances [Kav08]. autocovariation [Gal01]. automata [LS00d, Lou05]. autonomous [WD08]. autoregression [BRRZ03, HC09, Mar01b, Vou06, YZ08]. autoregressions [AG09b, CS05a]. Autoregressive [JTS08, LJ06, AT08a, AM02b, BD05a, BLS04, Bat05, Bos09, BR06c, CS08c, Cho05a, Coo03, Coo04, Cor08, Fit06, Gal01, Gal09, Gui01, Haf06, HZ03, HZ04b, HBK06, KJ08, LVL09, LN02, Mar01c, MB06, Nag09, NBW07, PS06a, PMVL09, PW06, Psa00, RP02, Ris05, SSR08, SN04, Sha06b, Shu09, SFS03, TMN01, VvdMD08, Vou08a, WIL01, WB05, Wor01, YWS09, ZCCH09, ZB08, ZB05]. auxiliary [Hen00, JD08b]. Availability [BS00, SS01a, CK01, Lee04a, Mi06, SL06, VT00]. available [ST03a]. average [AAV02, AK09, Bru09, CHV09, DY04, sDXlyY05, FSX02, Gal09, HS03a, HM06, KK08b, Kin01, xLxZ04, PLLJ06, Pre00, Ros07, SL06, Sun09b, WLG01, YX06]. Averaged [ZL07a, SBF01]. averages [Ete07, HK03a, KH08b, Sch01a, SH09b]. averaging [Cha02]. avoiding [CLM03].

Backward [NO09, WH09, AH06b, GJZ06, Hyn07, Jia08, Jia05, LX05, Lin09, LR02, RH07, Sit02, WX09]. backward-forward [AH06b]. badly [Nik08]. Bahadur [Coe08, Lin08c, MS09b, Pal01a, Sun06b]. Balakrishnan [YGK08]. balance [CGK06]. Balanced [ADJ06, ADJ07, AJ00, FM03b, JMP06, LK06, LLS06, MGSB08, NG02, OMJ^{+}04, TF05]. balanced-type [JMP06]. ball [DLL00, Naz09, Shm06]. Banach [AAV02, AM02b, CV01, Den03, Den07, HCSV00, HNRV00, HCV01, QH09, Rao02b, Tóm05]. band [Luc00]. bandit [Wan00b, Wan02d]. bands [Bur00, Lu06, MZ02, MZ06]. Bandwidth [APS09b, AP02b, BBR05, Gua07, HH03a, Ten01, WCC07]. bandwidths [AF01, HW01a]. barrier [LX05, YH09, YH08]. barriers [KK04, WY08a, ZJ09a]. Bartlett [BC05, LWT08]. based [AB05, AB08a, AF09, AW06, BA03, BM03a, BK08a, Bar00, BBR07, BM03c, BRSL08, Ber08, BF00, CP08, CQ03, CY02b, DS00a, DZ02, DGZ02, DN09, DOKN07, EZA04, EHMM08, ED02, FS01, FM03a, Gao03, GH00, GQR06, GdL00, GKM03, GV06, GBQ03, GJP01, Haf06, HS00, JG04, JO08, Jur00a, KH01, Küh01, LS00b, Li02, LZC05, LPY04, LA07, LLC06, Lin04, LT08, LSL00, LS08b, LS08c, Mac09, Mad08, Man05a, MOM07, MU04, MS07b, MN03, Miz00, Moj00, Moj01b, NZY09, Ome08, Osm01, PW08, Pen02b, Pyc07, RS07, RB08b, Ruk07a, Sch02, SH09a, SO00, Sim07, Sub09, TG02, Tia09, VL03, Vou08b, Wan08a, WN09b, WN09c, WP00, XY07, Yeo07, YPD09, Yoo09, dlHRB05]. bases [BHZ00]. bathtub [Che00b]. Baum [LS04a, Sto08, Sto09]. Baxter [IKP08]. Bayes [BGM03, Bis01, CM04a, CS00c, FC03, HH06, HY01, HK08, KE05, Kou03, KZ08d, LG03, Lia04, Lia09d, MP01a, MLT00, SL00, TT03, WV08, WC03, ZWY05]. Bayes-type [KZ08d]. Bayesian [Cer07, CK01, CMNP07, CK05a, CS00a, CW03, DFK03, DG06b, Fer00, FGMW09, GM03a, GVGP08, GLV00, Hua02, KW07a, KM03b, KY00, Lau06, LRM00, LZ02a, Lia07, Lia02a, LT08, LL03, MS02a, MNOP03, MGB05, NOP00, NOP01, ODO07, Omo07, OC04, PL07, RA08, Tia09, TM09, YM01, YW04b, dlHRB05]. be [Din00b, KN01a, Kla09, PS07a]. beamlet [Huo05]. Becker [ASS02, Hil00]. before [BS00, YZ01]. Behavior [OH07, Asc09, BB02b, Cri04, CFH02, GQR06, JKP01, Lad07, LK07, LQR09, Nic08b, PP06, Pin02, PS07b, RAS08a, VvdMD08, Yi05]. behaviors [Liu06]. Behaviour [Sen01, Sen08a, BM05, CC08a, CHV09, GMM00, Khr01, Psa02, Sgi01a, Sgi01b, Ten01]. Behrens [Drt08, KY04, TT07a]. Bell [SKS09]. below [Poz04]. bend [CLR02]. Benford [EL03]. Benjamini [GR08]. Bent [CLR02]. Bent-cable [CLR02]. Bernoulli [JJQ08, LP03, Mav07, Sun07a, Xia08, YZ02]. Bernstein [BC06, KW07a]. Berry [Bos04b, Bos03b, UWF06]. Besov [Rad07, Ros09a, Zha01c]. Bessel [CL07, DO07, Dev02, FJ04, Lef08a, Man05b, SLF09, YZ04a, YL05, YY08b]. Bessel-squared [CL07]. Best [MM09, Ryc02, AR03a, cICP09, IS01a, KL00b, Mor02, Ose09a, Por02b, Por02a, Por03, RL01, Shi01, TW06]. Beta [ALP08, BGN08, Cha04, Mar01c, OL03, SP00, Ris05, WZ09b]. Beta-hypergeometric [ALP08]. Bethe [YL00]. Better [Kab01, AAK00, Pre00, TS09b]. better-than-average [Pre00]. Betting [Sin07, BL03]. between [Bar06, CH04a, CQX08, Che08c, Cho08b, DV01, DL04, Efr03, Esh04, FF08, GP07, GVMSN04, GBQ03, Gor09, HY04a, HHC09, Huo05, Hür04, JM08, KK00a, KS07, KvD05, Kou06, Kow00, KZS08, KL06, Lia02a, LA07, Mac09, MB08, MC07a, MZG08, MY09, NC09a, NOP00, Ren09a, Roz09a, San05, SBF01, SZ06, TKO03, Van08, WBG04, Wen03, ZD04, Zuo01, dFF02]. beyond [Spo07]. BGAR [Ris05]. bi [BW06b]. bi-Poisson [BW06b]. Bias [Mid08, PL09a, GPC09, HF04b, IKK05, MBY00, MP08a, Nai01, OR01, dHC06]. bias-corrected [GPC09]. bias-robustness [MBY00]. biased [BLV04, sS09b, ST09, dUÁ03]. Biases [CP05b, CB01]. Bickel [BD05a, Che04a, LN02, Lee06]. bifractional [LN09]. bifurcating [ZB05]. bilateral [HWLZ09, KT08, Xie08, Xie09]. bilinear [BO02, BhRH04, Bib05, BL09b, CmHP02, DvdAW08, Gha00, LT00a, PS06b]. bimodal [LL08]. Binary [INA07, CHL08, FS06, KN00, Llo08, Nat01, SBF01, Sim00, TM09, Tso00, YKB07]. Bingham [KW07b]. Binned [MU04, KK00c]. Binomial [Car04, AA08, AR05, BH02, Cai00, Cek02, Fur07, HW07, KC03, LB02, LW08b, MSH03, ST04a, VU07, Wei09, WW08, Yu09a, ZR05]. binomials [PR09b]. biography [Ano07u]. biological [CHL01]. biomedical [Jan08]. bipolar [FG06]. Birnbaum [GOPB09]. Birth [vDZ05, Ath08, CM08b, DGNR08, GDF01]. Birth-death [vDZ05, CM08b, DGNR08]. bisexual [GMM00, MdPR07]. bit [MT04]. Bivariate [AE08, Has04, MRK08, Bar02a, BV02, BH02, Bry00, CP01, DB09, FR01, FR02, Fin03b, FE08, Han08, Has06a, Haz00, HM09, Hür03, Hut00, JL07, Jon02b, Kos02, KM09b, LX00a, LT06, LB03, MS02a, Mod03, Nav08, NQMRLÚF01, OL03, PS02, PS09a, Ris06, RLÚF04, Sch01b, SV02, SZ02a, SZ08b, VT08, WM03, Wan00a, Wan08a, WO08, dUÁMM08]. Blackwell [DW05]. block [ASG04, Che08c, Efr00, Ger08, KM02a, MGSB08, MB08, NG02, Sia00, SKPP03, Sim07, SS03, Udd08]. blocked [AH06a, CCDG02, IB09]. blocks [ASS02, ASG04, BB07c, EK03a, Jac08, MGSB08]. Blockwise [Bra02a, LZ01, Psa06]. Blowing [BB04]. blu [YOH00]. BLUPs [Mor02]. BLUS [Vou08b]. Board [Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06p, Ano06q, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h]. Board [Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. bond [Gap04]. bonds [CL07]. Bonferroni [CGNW01, Gor07, PAJ05]. Bonferroni-type [CGNW01].

Bootstrap [AFMGJG05, JTQ09, MP08a, MS07b, PP05, PL07, Psa00, APR03, BV07, BC01, Bur00, Cho05a, CAM03, Csö03, HKS00, HK03b, HH00b, JSV01, JSV02, JGMGdlV06, KH07, KS05, LLS06, Paw00, Psa03, Psa06, SC08, VT00, Yat02, YC06b, ZvdH03]. bootstrapped [HCV05, SB00, Sha01]. Bootstrapping [JSV05, JGMGPM03, SB06, LY05, LX00b, Mor02, NP00]. Borel [Cha08, HWLZ09, Lia09d, Lim06, Pet02, Pet04, Xie08, Yan08]. borne [GS08a]. bound [BH05, Dar04, FvE01, HK03a, KL03b, LS00c, MK08c, The07, UWF06, Wit05, Xia08]. boundaries [Gir04, GJ08, LWW09, Par01]. Boundary [Ren01, Abu02, BH05, CN03, FBVW00, HM09, HPR02, KZ08a, PK04, RMAA01, Spo07]. boundary-crossing [CN03]. bounded [Clé01, Den03, Dro06, JNS02, KMvE00, Le08, LL09c, MP02, MP06, Mar02, Moj01a, PG09, Pen02b, Poz04, Pro04, San05, Sto08, Xie09, dSF08, vE00, vEZ01, vEZ02]. Bounding [DY04, AP04]. Bounds [BCK01, Bis00, Dow09, Fer04, Mül01, BCLBMR00, hCW09, Che02a, CI01, DR03, DR04, DR00, GO01a, GMM02, HN08, HP08, Ima08, Jon00, KTZ03, LL09b, Llo05, Nyr05, OR01, Pap01, PW00, Raq03, SC05, Sha04, Sou01, Wor01, YH09]. box [DOKN07, CYJL01a, FM06, HB04, Liu06, Liu07a, Sar00]. Bradley [SZ06]. Branching [LO09, CLR06, CLR08, DFX05, Dev07, GMM00, GMdP04, GMM04, He09, JT08, KO08, LZ06, Ma09, MPY03, MMY09, MdPR07, MS02b, RAS08a, SY04, Tal08, YY07a, YP09]. break [Sen01, Sen07, Sen08a, Sen09]. Breakdown [Che00c, MM02, CT04, CFH02, HH00b, Oli01, Zuo01]. breaks [MR00]. bridge [BHHM04, BH05, Deh07, GL04, Lev00, YZ04b, Hob07]. bridge/motion [BH05]. bridges [Man05b, Wit05, YY08b]. brief [Ano07u]. Brittle [KBH07]. Brownian [DF11, Pou08, Abu02, Alb08, BJT03, BNO03, BTW01, BHHM04, BH05, BGT04, CH07, Car04, CI01, Deh07, DMMY00, Dun06, DLL00, DF02, DvZ05, EN03, Eng04b, ESOO09, GW05, GG02, GL04, GK04a, GK04b, HY04a, Hob07, HL01, Hoo02, HNS08, IKP08, Ist06, Kin01, LN09, Lev00, LZ07, LLW09, Mak09, MNN09, McG06, MMV01, Nan08, Nie04, NS06b, Ole08, SL00, SZ05, SZ09b, Tak04, Tot04, Tud08, WL08b, Wit05, YYL08, YZ04b, ZBZ09]. Brunk [HCW08a]. Bruss [Sza07]. BSDE [El 08]. BSDEs [FH08, Jia09a, Otm09, Roz04, ZZ08]. budget [VT00]. Buehler [Kab01, KL03a]. Building [UJ09]. built [MP08a]. built-in [MP08a]. Burkholder [Nik08, Ren08]. Burman [MS04a].

C [DWC04]. cable [CLR02]. calculated [Had03]. Calculation [KL00c, PB09, Son08]. calculations [YKB07]. calculus [YY07b]. Calibration [KSH07, GP05b, Lia02b, Tsa01, Tsa04]. callbacks [PNJ08]. Cambanis [Ano06o, CA06a, CA06b]. Can [Din00b, PS07a, PTA03, KN01a]. cannot [Kla09]. Canonical [DP03, Esh04, FR01, JFZ05, Mon08, Nki05, Pai08]. Cantelli [Pet04, Cha08, Che08b, HWLZ09, Moj01b, Pet02, Xie08]. Cantrell [CV09]. capability [Wri00]. Capacities [CB07, Cha09]. capacity [AZ04b, AK07, LV05, LV07]. capital [Mil07, Ole01]. capture [Hug01, Hug07, ZvdH03]. capture-recapture [Hug01, Hug07, ZvdH03]. cardioid [APS09a]. care [Lou00]. Carlo [BBP09, BSS01, EZA04, JD08a, Kos00, Lia09b, Mir01]. carrier [GS08a]. carrier-borne [GS08a]. case [AFMGJG05, BB09b, BJR00, CD01a, CMNP07, CM08a, Faz03, Fuj07, GLV00, He08, KY06, Lia02b, LL08, OE08a, Por02a, Sgi01b, TP01, XL06, ZWWC07]. case-control [GLV00]. cases [Aic09, Che04b]. casting [HC09]. catalytic [He09]. catastrophes [DGNR08]. catch [CP05a, PDM01]. Categorical [ZB09c, Bis04, Bro09, CKS04, Shi01, VL03, ZB09d]. categories [Bro08b]. Cauchy [AB00, BJB06, BM01a, HC07, MM02, Pen08, WL08b]. causal [Ima08, Lec08]. causality [Bib05]. caveats [Sin07]. Cayley [LW03]. cells [SS08c]. cellular [LS00d, Lou05]. censored [BCK01, BM03a, BK08a, Bet00, BBG05, Cra04, DS00a, Dub02, FJA09, FS01, Fer00, HcW08b, IP03, KZ08d, LOS02, Mad02, MNZ08, PBZ08, Pon06, Raq03, SSB09, sS09b, sS09c, SZ00, Sun06a, TG02, WH05, XL09, ZG01, dUÁMM08]. Censoring [GO03, HK09, AAB04, BCC02, BE08, BV08, BCK06, DB09, HY00, KL08a, KD03, KD04, Kos02, KY06, Lau06, LLC06, MN03, PB09, PD09, Rab00, SB08, WO08, Zhe01]. censorship [DV08, GS02, OS06, QZ07, SH06, SB08]. Center [YK03]. Center-similar [YK03]. centered [AVV09, Deh07]. centering [HSX05, JS02]. Central [Bir08, CM03, DR02, DR15, HP05, KMM00, PMJ08, Cha09, Gra09b, Gut06, Hör06, Hor04, Shu09, Yu09c, BHH08, Bla07, BdB09, CL08b, CC02, CPD00, Dud03, Dud08, FU08, GA00, JY08, JH02, JH08, JR07, KB01, KK08a, LW08a, Mal01, Neu04, RW00, RS03b, SW06, Sou01, WL08a, Web06, vZ00]. centrality [LZD09]. centre [PG09]. certain [AL06, AM02a, BCLBMR00, DDM06, Dud08, FS00, LS05a, LK04, MF05b, Nai01, Shi09, SS03, SS07d, Sta02, dlCV01, vW05]. Cesàro [YAT09]. Chain [GR07, BSS01, ES09, FSX02, FSC03, JF00, KH07, Lef08a, Mam02, MSW01, PG09, Son08, Spr01, Vag03]. chains [Ath08, Bou09, CP06a, Cha06b, Che00a, CQX08, Clé01, CW08b, CD00, CD01b, GTH06, GO02, Ilt00, JY08, JB03, Kos00, Leó01, MY08, Mir01, Nic08b, Pal00, Pas06, PPSW04, Poz08, RAS08b, Spr03, Sta05, SP00, YL00, Yan03b, Yao00]. Chamberlain [BBP09]. chance [KKAK08]. change [AH01, AS02, AH04, BR01, BHHM04, BZ06, Che09, Coo06, DFK03, Dup06, Ewa08, FJ01, FJ07, Fot09, Gom00, HW05, Jar03, JTQ09, Küh01, Lan08, LN04, LS08a, Nin05, PK04, WZW03, ZL09, ZLQW07]. change-point [AS02, BR01, BHHM04, Che09, DFK03, Dup06, FJ01, Fot09, Gom00, HW05, Nin05, WZW03, ZL09, ZLQW07]. changes [AGJ04, Bro08b, FG04, HR09, HKS07, Hus03]. changing [AS02]. Chaos [LT00a]. chaoses [Hei01]. character [CD05, Su05]. characterisation [FS08c]. characteristic [DK02, FvE01, He08, MU04, TSJ09, Zha07b]. characteristics [Bos03a]. Characterization [CG08a, FH07, HBA05, KKRV01, Kra00, LPY04, LG09, Por02a, RC00, AF09, AZS08, BA03, Cer08, CFM01, EFL02, FR02, LL04, MM07, Nov07, SS02, SZ09a, WC04, Yan00b, Zhe01]. Characterization-based [LPY04]. Characterizations [AR09c, BGHP04, LH01, MS08b, Nav08, VT08, AV01b, AE00, BM07, CL09, CH04b, EWI04, KC08]. characterizes [LBSM00]. Characterizing [KGC09, Lil04]. Charlier [KSO05]. charts [FSC03]. Chebyshev [BL02, CMS05]. Chebyshev-type [CMS05]. check [LSL00]. checking [CHD08, Duc04, Duc10, GR01b, SSY02, Wan08a]. checks [SQMK06]. chemical [CGK06]. Chernoff [The07, WZ08, Xia08]. Chi [BBP09, Gra09b, LZD09, WBG04, Wit02]. chi-square [Gra09b, WBG04, Wit02]. Chi-Squared [BBP09, LZD09]. choice [DH06, Kou03, Lee04a, Por02b, Por02a]. Cholesky [SS05b]. Choosing [MS04a, HSX05, Por03]. Choquet [HHC09]. Chover [PQ03]. Chover-type [PQ03]. Chung [Che00a, JKK05, ZL07b]. CIR [Guo08, Hyn07]. circle [DMB02, Jon08, KS08a, Pyc07]. circulant [BM02]. circular [LK06, MPT09, Pal00, UJ09]. claim [BLS08, Has04, MZG08]. claims [CTY02, KZ08c]. class [Ars05, BK08b, Bea09b, BJB06, Bis00, Bra05, CL06a, CÖ02, CP01, CFUOV00, CD00, DW02, DDP08, DR00, DGJ09, EsO08, FR01, Fin03b, FB02, GE07, Gor07, HV06, Has06a, He09, Jam03, Jia08, KW00, KA07, KM06, KS08a, Kla00, KJ08, Lan06, Led08b, LLJ00, Li04, LCZ09, Lin09, LY03, LW03, MPA07, MGSB08, Moj01a, MdPR07, Nai01, Nan09, NQMRLÚF01, ÖK07, PL09b, PMA08, PDM01, RLÚF04, Sal02, SBF01, SSA09, SS03, SS07d, Tar01, Udd02, zW06, Wan08b, WHSL08, WH02, WX08, WH07, XW09, Yin08, ZF06, dHC06, vW05]. classes [AAK00, Ahm04, BS06c, Bla08, JT08, LS06a, Lia01, OS05, Raj09, SP00, ST04b, Vin08]. classical [BW06b, DR08, Fot09, GN08, KL08b, LWD07, ZZG06, ZBZ09]. Classification [SS08c, DS02b, Duc09, FT04, KKM^{+}08, LA07, Lin04, MOM07, Moj00, Shi01, WC03, Wul08, ZF06]. classified [KSG08]. Classifier [Moj01a]. Clayton [Wan08a]. clear [AZ04a, AH06a, LCLZ06, YB07]. clinical [MPS06, YW07]. clipping [Ste05]. closed [CHL08, FNA09, ZM06, dFF02]. closed-form [CHL08]. closeness [AB09, BIKM09]. closest [MN09]. Closure [Li04, Gel09]. CLS [FM05]. CLT [Kan07, LS01, MS03a, Mas06a, Mir05, Moo08, Poz04, XW09]. Cluster [MOM07, DS01c, DWC04, Jia01a, SBF01, Tso00, Zan03]. Cluster-based [MOM07]. cluster-specific [SBF01]. clustered [Hug04, Jow06, KG06, Mid08]. Clustering [IR08, GBQ03, KNB00, Shu03, YZ02]. clusters [Seg05]. clutter [JK08]. co [Bar02a]. co-ordinate-wise [Bar02a]. Cochran [LF07]. coding [LC05]. Codispersion [RV08]. Coefficient [HL02, AG09b, BM06, El 08, ET07, FV08, Fra06, Fuj00, GAW08, Had03, HWL08, HWZ09b, HZ09b, HZ09c, HBK06, JO08, LSEB00, LR02, Man08, Mar01c, Nag09, NS09, Nov04, PP00, Rou07, Roz04, Sch01b, SV02, SW07b, TPA08, Tso00, YZ06, Yu08b, ZL07a, ZX09, ZY04]. coefficients [AH06b, BM07, Bha05a, FS08a, GJZ06, Jia08, Jia09a, Kon08, Lee03, Lin09, NO09, RV08, ST03a, Sit02, WH01, Wan08c, WH09, Yu07b, ZWWC07].

Cohen [CP05b]. coherence [BKM04]. coherent [BS08b, NRS05]. cohort [FH06]. cointegration [Coo06]. cokriging [MS02a]. collapse [GK04a]. collected [Li02]. collections [Ale01]. Collision [Nak08, Wen03, CWZ08]. Colombeau [ÇA01]. colorings [Jon02a]. column [BB08a, Udd02]. columns [DDM06, MS04a]. combination [cICP09, Hou05, JB09, JL09a, JL09b, RN00, UI00]. combinations [BK07, MK08b, Wit02]. combined [Moj00, Wei08]. Combining [KM02b, CCDG02, Lee04b, YZ07a]. comes [Mor02]. coming [AM09]. comment [Ali03, Ros09b]. Comments [Bar06, GS03a, GMZ07, HWLZ09, Lia09c, BV02, Mil07]. common [FV09, Hut02, MN09, ZP09]. comonotonic [Jia06]. comonotonicity [LR08c]. compact [ETP08]. compactly [CV01]. compactness [Aur08b]. comparative [GP05b]. Comparing [Mor05, Cai00, CCC02, CFCGM08, HT05, Li09a, Lia09a, Nic08a]. Comparison [FP07, Fig07, LR02, Ome08, PF07, Sch00, WX09, AA08, Als01, Che08c, Dro06, GN08, GBQ03, Gør02, Jia05, KS05, NW05, Naz09, Ose09b, QSP01, Voc06, Yue01, ZD04, ZB09b]. comparisons [Bar01, BS01a, CH04a, HW01c, HW01d, HZ05, Kak08, LH08a, LS00d, MSH03, NRS05, Sha07]. compatibility [BCT08]. Compensator [RB06]. competing [FE08]. complement [OE08b]. complements [Bos04b]. Complete [JD08c, KK08b, xLxZ04, Lia00a, Sun07b, Yu09c, AAV02, BK08a, BW07, CMNP07, Csö03, Den05, FRSWT08, Ger08, HSV98, HV00, HX02, KL03b, KVH06, Kuc04b, Kuc07, Kuc09, LL06, MGSB08, Pas06, SS09a, SS07d, SVH05, Sun08a, Sun09b, WZ06b, YX06, ZY08]. Completely [Hel09, Ger08]. completeness [IPP04, Irl04]. completion [CP06b]. complex [GLX07]. complex-valued [GLX07]. complexity [Dai04a]. Component [dCB00b, Alt08, Cho05b, Sun00, VZ03]. components [ASS02, BA08, BS08b, BF00, Bou06, dCB00a, Leh05, Lo05a, Nat01, NRS05, PMM09, SS08a, Wul08]. Compound [Mas02b, Cha05, FSC03, Gat08, Gat09, HH02b, LW08b, Mas08, Nov03, Psa09b, TSJZ01, YZL08, YZ09, YH08]. compounds [BRVM03]. computable [Dai04a, KL00a]. Computation [LC05, MNZ08]. computational [BCC02]. compute [CL08c, FP08]. Computing [Pap06, Paw00, Fot04, FSC03]. concave [KKM^{+}08, KL06, MR08b, QS05]. concavity [CXH09]. Concentration [HS05, JR09, Ole01, Zan03, AFS^{+}07, AHV06, Fai01, WJ05]. concept [Hag09]. concerning [Eng04a, HCV01, HWYZ09a, Nic08b]. concomitant [EB03]. Concomitants [WN09a, AA07, VT08]. Condition [Jen04, AL06, CLF06, CPD00, JD08a, Jia09b, MSW07, Mas06a, Röl08, Zha01b]. Conditional [Din00c, IB09, LB03, MB08, Mas06b, Nic02, Abu02, AT08b, BHZ00, BCG01, BV08, CV01, CGNW01, CG02b, Cri04, CP05c, CDRV01, Deb09, FS06, FJ04, Fot04, GM06, Haa09, IP03, KD04, Kol00a, Kow00, LLOS09, Lil04, Liu00, NOP00, OS06, STKF08, SS05b, WP00, XL08, YK05]. conditionally [Dav04, HKLB07, Ose09a]. conditionals [ACSGV00]. conditioned [GL04, MN09]. conditioning [BNO03, MB08, Sin07]. Conditions [CW08b, JK08, Ker08, BHGRV08, BCT08, Cha08, CI09, CD01b, HY04b, Hua03a, KP00, Kuc04b, Lel08, LWW09, dR06, RMAA01, Ter03, Wan05a]. cone [Log03, OS02, Sha03]. cones [AW04, Del03]. Confidence [AB04, BM09, Bos03a, FK08, HJ06, KLL08, Lud09, Tar05, AB05, AB08a, ARB09, Bar02a, Bur00, CQ03, Din00b, DG06b, GL08, HS00, HH02c, IF08, KL00a, KL00b, Kab01, KL03a, Kab08, KG09, LB02, Lu06, MZ02, MZ06, MP08a, Ome08, Ruk07a, Sha03, SW01, Tri04, VT00, Wan02a]. confident [AVV09]. configuration [HP08]. confined [Pad01]. conformability [BCEO01]. confounded [EK03a]. conjecture [BK02, Eng04a, Hür03, LO05b, PN09, Ros09b]. conjugate [BB09a, LZ02a, PR09a]. conjunction [WF00]. connected [dCB00a, CLO09]. connectedness [QO09]. connection [CH04a, GP07, MC07a, NC09a, Ren09a, SZ06]. connections [JM08, Lia02a]. connectivity [AR02]. consecutive [Cui02]. consecutive-k-out-of-n [Cui02]. consensus [BL03]. Conservative [Ruk07a]. conservativeness [Dah06]. Consistency [BBR07, BHGRV08, Che02b, DSBA09, FS01, Had03, JGMGdlV06, LLOS09, LYZ08, Lia07, Shu04, Tia09, XTW06, BH03, Ber03, BRRZ03, Blo07, FI06, GS02, HZ04a, JSV01, KF08, Kos02, KP02, Lab08, LL05, LC04, Mas02c, Neu07, STM01, WC04, ZCCH09]. Consistent [BBG05, Bha08, RS00, ZK06, Mas08, NS09, QF02]. consistently [CZ07, Liu07c, SL08]. constancy [BHK04, HL02]. constant [AR03a, IS01a, KZ08c, KW07b, LB02, Ose09a, RL01, Yi05, YH08]. constants [CBR00, DK08, II08, Roo01]. Constrained [PK07b, PK08, DG08, FC03, KT03, MSW01, SVB07, Yao00]. Constraints [FV09, AT08c, FI06, MLT00, Sha03]. construct [Jar08]. constructed [KS04]. Constructing [cICP09, Kon08, LB02, SS09a, Sha03]. Construction [GKM03, GSKP09, KM05, Sue03, YZL07, BRSL08, ÇA01, FQ03, Jac08, KM08, LLZ04, LLS06, Mas08, PZL04b, SS07d, Udd02, Won09]. constructions [AJ00, PW06]. constructive [FW08a]. contact [Hue03]. containing [AZ04a, Tri08a, YB07, ZZ09]. contaminated [GVS00, SV02]. contamination [Pal01a]. Contents [Ano07a, Ano01g]. context [FI07]. contiguous [HB01]. contingency [KA07, ZB09c]. continued [ALP08]. Continuity [FR06, BK01a, CC01a, Dar01, Jon00, Kab09, KL03c, Lou00, Lou01, NS06b, Zap03]. Continuous [CP06a, Alb08, AH06b, Ara01, AVG08, Bar02a, BR02b, Cha02, Dub03, El 08, Ery07, EF08, HV06, HR08b, IPP04, Jia09a, LN03, LS06a, LR02, LW08b, MS05, MNX09, PTA03, Ren08, Roz04, TS01, ZvdH03, dL05a, vZ00]. continuous-parameter [LS06a]. continuous-time [BR02b, LW08b, PTA03]. continuous-valued [Ery07, EF08]. contoured [BG09]. contours [GQT07, GV08, Kim00a]. contractions [OAAA05]. contracts [RS07]. contrast [Bis06, Gut00, PLLJ06]. control [FSX02, FSC03, GMdP04, GLV00, HT05, Lec08, Lep03, LdC07, MN01a, PP01, Sit02, ZBZ09]. controlled [GMdP04, HL01, Maz00, MR03, MdPR07, Tan09]. controlling [GSTS07, KW02]. controls [Oeh00, Oeh01]. Convergence [AD01, BCAC01, BGK07, Bre06, CS08b, DZ02, ETP08, HU06a, HCV01, Li09b, Mac09, Mim08, Sap08, Sim01, Tóm05, Waa04, AAV02, BB09a, Bar06, BJT03, BW07, Bis06, BMR07, But00, Cek02, Cha06b, CLF06, CC02, CP05c, Csö03, Dai00, DE06, Den05, DM05, EZA04, EP02, FR02, Fot09, FvE01, Fuj08, GW05, GRT07, Gui01, GJ01, HW05, HY00, HSV98, HV00, HRV08, Hür04, IF08, JD08c, KV02, KSW06b, Kav08, Kim00a, KK00b, KKH08, KK08b, KVH06, Kub08, Kuc04b, Kuc07, Kuc09, Led08b, Lel08, xLxZ04, Lia00a, Lia04, LL06, LLW09, Maj05, Mas02a, MF05b, Mei07, Mon06, OS06, Pak08b, PS09b, PS09c, QH09, RS08, SW06, SM09, Sha01, SWM09, Shi04b, SM06b, SVH05, Sun07b, Sun08a, Sun09b, TP09, Ver07, WZ06a, WZ06b, Wu08a, XL08, YX06, Zap08]. convergence [ZW01, ZY08, vEG08]. convergences [Arc00, HX02]. convergent [KO04]. Converse [Jia05, LS06b]. Converses [Fro05]. convex [AW04, BR08, BK07, Har04, Lil04, Llo02]. convexity [Alz06, CM04b, DLS00]. convexly [BS01a]. convolution [Fur07, MR08a, NT06a]. convolutions [MK08a, MSZ08, Psa09b, ZB09a]. Cook [KLP01]. coordinate [FWW01]. copies [Mat03, NZ00]. copula [ARN08, BW05, WO08]. copulas [CS08b, DFS08, FJS03, Hür03, LP09, NQMRLÚF01, RLÚF03, RLÚF04, WH02]. correct [SLC09]. Corrected [CFUOV00, FCN02, Buz09, GPC09, KZ08a]. Correcting [Coo02]. Correction [FT01a, Oeh01, Pal01b, vEZ02, LWT08, MP08a, PL09a, SXLX07]. Corrections [BX00, BC05, FR06, Hal09]. Correlated [MNX09, ADJ07, Bis04, Cac01, CG02a, CG08a, CHN04, FS06, FM04, JJQ08, KN00, LXZS09, Omo03, Özk08, PMM09, PM01, PMA08, SBF01, SN09, SS00b, Udd08, YKB07]. correlation [AA07, BB05, Bon03, CW08a, CWTB09, DMP01a, ET07, FD05, Fuj00, Gua07, HWL08, Huo05, JO08, LB03, MT07b, Mon08, NS09, Nov04, PPRW08, RL08, Sas08, SV02, Sim00, SW07b, Tso00, YW06, Yu08b]. correlations [BR06c, NS06a]. correspondence [SBF01]. corresponding [KKRV01, NMO08]. Corrigendum [BvdV09, DR15, Duc10, DF11, KM09a, Kim13]. corrupted [MNN09]. cost [Luc00, ZZ09]. costs [Pou08]. count [Hag09, HM06, Jow06, JS09, Pas06, ZJ09b]. countable [ST01a]. Counter [NT03]. Counter-intuitive [NT03]. counterparts [Als01]. Counting [Nic08a, Bro08a, Fie03, RB06]. counts [And00, BMR00, Omo03, Tan09, Wei08]. coupled [DHS00]. couples [MdPR07]. coupling [Lou05, Pel02]. coupon [CL07]. course [PI09]. covariables [dUÁ03]. Covariance [LPS09, Ruk03, WZ09b, Zaj09, Bar05, Bea09b, BD05b, CG02a, Cor04b, DH09, GM08, HS07a, HRV08, Hut00, LK04, LZ05, Lyh05, Ma03, MB08, Mas06a, MU01, NHO03, Pan04, Pap06, PMZP07, Pri01, Rom02, SG06, Sch02, SS05b, YW09, YC04, YOH00]. covariance-matrix-based [Sch02]. covariate [Aic09, NS09, SN09, WW01]. covariate-adjusted [SN09]. covariates [MSW09, SDR01, SZ08a, Wu08b, ZvdH03]. Cover [Ano03t, DW02, PDM01, WX08, XW09]. coverage [AVV09, Din00b, KL00a, OH03]. covering [Jon08, KTZ03]. Cowan [Sza07]. Cox [CYJL01a, FM06, HS08b, HB04, Kra04, LWT08, Liu06, Liu07a, LXZS09, MS08b, Sar00, SZ06, TD08, Waa04]. Cramér [Jan03, PP09, CI01]. Cramér-Rao [Jan03, CI01]. Crash [RS07]. Cressie [Bra02a]. criteria [But01, DWZ05, GP07, LCLZ06]. criterion [Ano06o, Ath09, CLR06, CA06a, CA06b, GM03a, Haf06, KL03c, MW00, Nin05, ÖK07, QF02, SC08, Zap03, ZQ06]. Critical [MMY09, CS08a, HBK06, KL00c, Sgi01b]. Cross [KTM08, Lia09a, BR06c, CW08a, GD04, GDM06, Gua07, HT05, Li02, SS09a, SS07d, Wal02]. cross-correlation [CW08a]. cross-correlations [BR06c]. Cross-spectral [KTM08]. Cross-validation [Lia09a, Gua07, Wal02]. Crossing [ZP09, Abu02, BM03b, BH05, CCC02, CN03, FP07, Gir09]. crossings [GG00, ST03b, TS01]. crossover [BM03d, KG09]. Csörgo [Fro05]. Csörgo-Révész [Fro05]. Cumulative [Mar06, CM08a, Wri00, dFF02]. cumulatives [Sch00]. cure [LTS01, PZ08]. Current [ARB09, AB05, AF09, Rab00]. curvature [AA01]. Curve [PAJ05, Llo02, Pan04]. curved [DS02b]. curves [DZ02, Gør02, HH03a]. curvilinear [Mak08a]. cusum [WN09b, WN09c]. cut [DKM02]. Cutoff [DN09, GH05a]. cutpoints [HF04a]. cuts [BLP03]. cyclically [DFX05]. cylinder [Yan08].

D [KF08]. Dalal [BK02]. Darling [BW06a, CLO09, DO07, Man05b, Pyc03, TO03]. Darling-Erdos [BW06a]. Darmois [KW00]. Data [Gao03, GJP01, LL02, Rom09, AP02a, AL00, AS06, BS06b, BK08a, BL08, BL09a, BBR07, BHGRV08, Bet00, BBG05, Bha05a, Bro09, CP08, Cai03, CQ03, Che04b, Cho05b, CKS04, DSBA09, DS00a, DKM02, Dub02, FS01, FS06, Fer00, GP00a, GL03, GMdRV02, GLV00, HM06, Han08, Haz00, HS07a, Hug04, Hug07, HcW08b, IP03, IB09, Ing00, IWWJ04, IR08, JS07, JS02, Jow06, JS09, KL00a, Kab08, KK00c, KG06, KN00, KZ08d, KY00, KSG08, LOS02, LS00b, Li02, dL01, Mad02, MNZ08, Mar02, Mei07, MN03, Moj01b, PI09, PBZ08, PPRW08, PMZP07, PD09, Pro08, QBZ09, Rab00, Rei08, Rod00, RA08, RL08, SS05a, SK06a, SBF01, SSB09, sS00a, sS09b, sS09c, Shi09, Sim00, Son09, SXL05, SS08c, Sun00, SZ00, SZ01, SZ02b, Sun06a, ST09, TG02]. data [Ten01, Tri08b, Tso00, VL03, VV00, WH05, WZ05, Wan08a, WP00, YZ05, Yu07a, Yu09b, ZZ06, ZZ09, ZG01, ZJ09b]. data-dependent [Ten01]. data-driven [Ing00]. data-invariance [JS02]. datasets [Rei08]. Davis [Nik08, Ren08]. DDA [DR03]. dead [Luc00]. deadband [LdC07]. death [CM08b, DGNR08, Ima08, PD09, vDZ05]. debit [HWZ09a, YZG08]. decay [Hue03]. decision [CS00a, Ili00, KKM^{+}08]. decisions [Mas08]. decomposability [CS00b, Koz05a, Koz05b, iS01b]. decomposable [AM02a, Koz05a, Koz05b]. Decomposition [Cas07, KNB00, Deb09, LN09, LT00a, PLZ03, SS05b]. Deconvolution [Cat01, QF03, DH06, DGJ09, Mei09, Pen02b, PS09b, PS09c, RS00, vEG08]. decoupled [Kla09]. Decoupling [dlPZ02]. decreasing [AY02, Aur08a, CLM03, Fai01]. defective [Psa09a]. deficiency [LOS03]. deficit [CTY02, XZJ08]. defined [Asc09, FG06, Fig07]. defining [MC09]. definite [YOH00]. definiteness [SvPtK08]. Definition [AVdPS02, Deb09, FW08a, Mil07, Pai08, SS01b]. deformation [GP00c, PS00]. Degeneracy [BC03]. Degenerate [OE08a, BW08, JGMGPM03, JYZ08, Kur08]. degree [ASG04, BM06, Gup09, Hua09, Wat01]. degree-of-freedom [Wat01]. degrees [Jon02b, Wit02]. Delay [AH04, BHY09, GR06, WZ09a]. delayed [GW09, GK04b, Luc00, Wan00b]. delays [HR08a, Hal09]. deletion [EKM02, HK09]. delta [Ber08]. delta-sequence [Ber08]. demand [BR06a]. demeaning [Kuz05]. demimartingales [Chr00, CH09a, Wan04]. denormalized [Kla09]. densities [ACSGV00, BSS01, BW05, CFCGM08, KT08, MR08b, Nag03, Pen02a, Sha04]. Density [Pen02b, AGG02, AF01, AR09b, BC06, BYB08, BBR05, Ber08, BBG05, Bia02, BGN08, Bre08, But00, CD01a, CKS06a, Che08a, Che02a, Che02b, DZ01, Dow09, FJA09, For00, FI06, GS02, Gir09, Haz00, HM09, HW01a, HZ03, HZ04b, KSW06a, KSW06b, KZ08a, KE05, KF09, KK00c, Lab08, LCPY04, LN04, LOS02, Lia09a, LS08c, Mas02c, Mei07, Mim08, Miz00, Mna08b, Nai01, Neg01, NP00, Osm01, Pel05, RS00, Rao02a, Rou00, Rou01a, Rou01b, SW06, SWM09, Shi04a, SY07, SXLX07, SZ01, TLRvdV08, Ten01, Ten03, Var08, VV00, WCC07, YY08b, vE01]. Dependence [FD09, KK05, AA07, AP02b, BL08, BvdV08, BvdV09, BV02, Cha08, Col08, CM08b, CPD00, Duc04, Duc10, DFS08, FR01, Fal05, Fra06, FF08, Fro06, GAW08, HY04b, JL07, KK08b, Kow00, xLxZ04, Lou01, Mar01a, MF05a, MZG08, MZ04, OH05, Pen01b, dR06, Sch01b, Sun06b, VL03, WH02, YAT09]. Dependency [CL09]. dependent [AZB07, AVG08, BC06, BYB08, BCG01, Bib05, BV08, CZ07, CL08b, CY02a, CM03, Coe08, DL06, DS09, Dub08, Dud03, FJA09, FMS04, FGMW09, Fur02, bGhWhZhW07, GMM02, GMdRV02, GMM04, HW05, HRV08, JT08, Jon02b, KW00, KB01, KRV05, KM02b, Kuc05, Kuc07, KH08b, KS00, LX00a, LZ01, Liu09, Mar02, Mat05, Muk09, Mül01, NRS05, Pap01, Pel02, PD09, QL06, Rab00, RW00, Ryc08, San05, Sar00, SDR01, Sek08, SL08, SK06b, SZ01, Ten01, TP01, WLG01, Wan02b, WN09a, Wan09a, YV06, YZ02]. depending [Lee04a, MdPR07]. depth [Gao03, HR08b, Kim00a, KH01, Rom09]. derivative [Bir08, Eni06, Ewa08, HC02, LX01]. derivatives [AGG02, BR06b, KW07b, TS09a, Ten03]. derive [CDMLDK01]. derived [Cer07]. description [Lee00b]. design [Aic09, Ath09, BR04, But01, CGK06, CL08a, Che07b, DWZ05, EK03b, FY00, FLQ03, Gai07, GM03a, GRT07, HP07, Hil00, HW01b, Ioa04, Jar03, KM08, KK06b, MS04a, MPS06, Nia09, NG02, PM01, Pro03, Pro04, Pro09b, RH02, SLC09, Sim07, XNF07, Yue01, ZAC03]. design-adaptive [Aic09]. designs [AJ00, ASS02, ASG04, ADJ06, ADJ07, AZ04b, AH06a, AHL08, AK07, BD01b, BM06, BM03d, But04, But05, But08, CWL08, DMB02, EKM02, EK03a, EK04, EKS05, Fan02b, FQ03, Fan03, FM03b, FRSWT08, GKM03, Geo07, Ger08, GD04, GDM06, GSKP09, Han03, HC04, JL07, Jou07, KM02a, IV00, KK00b, KS08b, KS04, KM05, LL02, LLZ04, Lia00b, Lia01, MGSB08, MS07a, MW00, OW03, OMJ^{+}04, OK06, Ozt07, PM06, PM08, PP06, QC04, QO09, SB06, SS09a, Sia00, SKPP03, SS03, SS07d, Udd02, Udd08, YZL07, YB07, Yue02, YW04b, ZNS08, AZ04a]. detect [BHHM04, GLML09]. detecting [NW05, YKB07]. Detection [AGJ04, WIL01, And05, AH04, Fuk05, Gom00, HKS07, Hus03, JK08, Pan04, SS08b]. determinants [MPA07, RW05]. Determination [Le 03, DMB02, HW01b, Ibr08, LB02]. determine [GVS00]. determined [CH01]. determining [MMB07]. deterministic [BN01, CL08c, DZ02, HY01, Kuz05]. development [SC08]. developments [HY04b]. Deviation [LS00c, Mór09, BLS08, Com07, sDXlyY05, DdHL03, HK03a, Ing00, LCZ09, LS01, Lin08b, LW03, MSW07, Mac09, Man08, MS09b, Nag03, SW00a, SM09, zW06, Wan08b, Wit05]. deviations [Arc02, Aur08b, Aur08a, BS06a, BJ00, Ber08, Bis08, CZ07, Dar04, Eng04b, Esa04, Gao01, GJ09, HO08, HCV05, Ilt00, Jou07, Li03b, Lin08a, Liu07c, Liu09, LV05, Mac08, Osm01, Pou08, Roz09b, SL08, Shi02, SD09, TSJZ01, Wor01, XL06]. DF [Rod00]. DFRA [DR03]. DHF [dBC06]. diagnosing [ET05, Wan07b]. Diagnostic [CHD08, cICP09, Duc04, Duc10, EHMM08]. diagnostics [DW00, GPC05, HK09, Jen01, Jia01b, KPK02, PMVL09, TWW00, XWL09]. diagonal [BGK07, Hei01, MB08]. diagram [TT08]. diallel [GD04, GDM06, HT05, Li02, SS09a, SS07d]. diallels [CCDG02]. diameter [Tar01]. dichotomization [WG03]. dichotomous [MPS06]. Dickey [Coo02, Sen01, Sen07, Sen08a]. Difference [XY07, CHD08, CC08c, KG09, KL06, PZL04b, SH06, vEZ01, vEZ02]. Difference-based [XY07]. differences [AB08a, HCW08a, NW05, Nie04]. different [Ale01, Arc00, BNO03, Dav05, HY04a, Jon02b, Lee04b, ZL07a]. differentiability [Sch01a]. differential [BHY09, Bez08, DaY08, FT01a, FT01b, GJZ06, HR08a, Hal09, HNS08, Ign09, Jia08, Jia05, LX05, Lin09, LT00b, LR02, LL09c, MN04, MS08c, Nic08c, NO09, RH07, RMAA01, SL09a, ST01a, Sit02, WD08, WH09, WH07, Xu08]. differing [Spo07]. difficulties [Hug01]. diffusion [AL00, CKS06a, CN03, Dow09, GR06, Gas09, Gat08, Gat09, Lee06, LL04, LXZS09, Man08, Maz00, Nak04, Neg01, Shi09, Stu00, Tak04, VA07, WY08a, Xi04, Zha01c, Zha07a]. diffusion-type [GR06]. diffusion-wave [AL00]. diffusions [DFX05, Gri03, He09, LL05, MR03]. digraph [CP05a]. digraphs [PDM01]. dilution [Zel06]. Dimension [YC06a, BV07, BD05b, Szu02, WY08b, Wen07, Xia09, YC04, Yoo09]. dimensional [Bri00, CG02b, Che04b, CFR06, DW02, GLX07, Jia09a, LHLC04, NS06b, OE08b, Sun00, Tak04, UWF06, WY08a, WX08, XW09, Zha01c]. dimensionality [CY02b, Nki05]. Dimensions [KO03, SY04, HH00a, KO08, LX08, RR01]. direct [AGG02, AB07, Van08, Ano01g, Ano01h]. directions [BX00]. Dirichlet [BDM08, EGM06, FH07, FW08a, HSX05, Jur00b, KGC09, Let02, Lic09, TJ06, WM03]. disaggregated [Sco03]. disclosure [Rei08]. discontinuity [HC02]. discontinuous [Ara05, Jia08, KKP00, NO09, Yu07a, Yu09b]. Discordant [Pan04]. discounted [LW08b, Ren09a, Sun05, ZZG06]. discovery [GSTS07, GR08, KW02, Oeh00, Oeh01]. discrepancy [CQ08, ZNS08, ZRQ09]. Discrete [BS09, OC04, Sza01, TS01, AS01, BR02b, BL02, CG08a, CQ08, CQX08, CS00b, DLB05, FGMS08, GW08, KL00a, Kab08, Lia09c, Liu03, LBMD05, MP06, MR08a, Mór09, PSS00, PN09, PTA03, SSA09, VJ08, WH08, YZL09]. Discrete-valued [BS09]. discretely [Bis06, Lon09]. discreteness [Jam03]. discretized [Waa04]. discriminant [Che04b, PP00, Poo06, RM05, Shu03]. discrimination [DL04, GP07]. discussion [GVGP08]. disease [YZ02]. disk [LR08a]. disk-percolation [LR08a]. Dispersion [RS03a, BB07a, BLV04, BRSL08, CF06, ML03, Rom09, TWW00, XTW06]. dispersive [KK00a, RS03a]. Distance [San05, AFMGJG05, BBR06, BW05, DR00, Hil01, Hil03, KLP01, KL03c, KH08a, Li00, SV00, SLF06, SR04, vEG08]. distances [DV01, EZA04, dlHRB05]. distinguishability [Erm00]. distortion [Coo02, CV04]. distortions [MS08a]. distributed [CP06a, FJ01, Gup09, KKAK08, Kru08, Lai01, NC09b, Sch03, SH09b, SvPtK08, XL06, YT08]. Distribution [AB05, Cha05, CM01, Ery08, KP02, MK08b, NQMRLÚF01, Pap01, RLÚF03, AHEA04, APS09a, ADL05, AJ05, ARB09, AM09, AFMGJG05, Ali03, Alz06, AR09b, AB07, AVG08, AB00, Ars08, Ars09, BC06, BK02, BM03a, BK08a, Ban08, BHY09, Bar02a, BB04, BSCN09, BB02a, BM04, BTW01, BH02, BG09, BM02, Bos04a, BR06c, Bra02b, BP08, Cek02, CM08a, Che00b, CC01b, CCC02, Che02b, CW03, CS00b, DB09, Dem08, DR00, DK02, DC01, ETP08, Fal05, FY00, FS01, FH07, Faz03, Fer04, FG06, FNA09, FV08, FM06, FvE01, FSC03, Fuj07, FS08c, Fur08, GH00, GTH06, GX08, GT04, GOPB09, Gon08, GMdP04, Gra09b, GR00, Han08, HSX05, Has05, Has06b, HR08b, HW07, Hen00, HA03, Hou05, HH06, HC07, HS05, IKIS09, JBB08, JB09]. distribution [Jar03, JK07, JF00, Joh02, JR08, Jon02b, JZ04, Kab09, Kar02, Khr01, KM03a, KM03b, Kim08, KM09a, Kim13, KSO05, KM09b, KL06, KW07b, KR09, Lan08, Le 03, Le 06, LSSG08, LWD07, LL09a, LZD09, LB03, Lim06, LL03, LBMD05, Lou00, Lou01, Lou03, LQ04, Lu06, MN09, MP06, Mar06, MN03, Mim08, Mod03, Mor05, MSW09, Mus00, NPC05, NQMRLÚF03, OH05, Oht02, OR01, OL03, PSS00, PK07a, Pen01a, PS02, PR08, PDM01, RC00, Roz09a, SS08b, Sch01b, SZ09a, Sco03, Sek08, Sen07, SSB09, Ser00, Sgi01a, Sgi01b, sS09c, SK06b, ST04a, SS06c, Sta00, Ste05, SBH03, SZ02a, SW00b, SF12, Sza01, SR04, TR02, UWF06, VU07, Vou06, WZL00, WO08, WM06, Wit02, Wri00, YGK08, YSN01, YZ01, Yan00b, Yeo07, Yi05, YC04, ZPK08, Zel06, Zha08]. distribution [ZJ09a, dFF02, dHC06, dUÁMM08, dCD09, vE00]. Distribution-free [AB05, KP02, ARB09, GH00, OR01, Ser00]. Distributional [TT07a, Sta02]. Distributions [BMW06, MiSW00, Sch03, AE08, AAK00, AY02, Ahm04, AZS08, Alb08, AV01b, AVdPS02, Ars05, ALP08, BB09b, BvdV08, BvdV09, BCLBMR00, Bar00, Bar01, Bar02b, BS06c, Bar09, BR06a, BNO03, BJB06, BH06, BS01a, BR02a, Bla08, BLS05, BW08, Bos03a, Bos07a, Bos09, BL02, Bro09, Bry00, Cai00, ÇA01, Cha04, CH04b, CKS04, CY02b, DLB05, Dev09, DL04, DZ04, DMRA07, Fan02a, Fin03b, Fot04, Fot05, Fra06, FV09, Gel09, GH05a, GLC09, GQT07, GV08, Gri08, GV02b, Hag09, HKV04, HV06, Has06a, Has09, Hoo02, Hor04, HW01d, HC06, IKIS09, JM08, JT08, Jur07, KW07a, KS08a, KMvE00, KKRV01, Kle02, KN01b, KFW02, KvD05, Kow00, Kum02, LX00a, LN03, Lan06, LT06, LS02a, Let02, LLJ00, Li03a, LL04, Lia09d, LH01]. distributions [LL08, Lop08, LBMD05, dL05b, LRW09, MM07, MR09, MB08, Mas08, MR08a, Mi06, Mna08a, Mór09, MR08b, NK03, Nad05, NT06a, NP08a, OS05, PF07, Pel00, Pin02, Psa09a, Psa09b, Pui08, Qi03, Raj09, Ryc02, Seg05, SV02, Shi02, SSA09, Spo07, SP00, ST04b, Umb06, Umb08, UJ09, VT08, WBG04, Wan05a, WJ05, WA01, XL08, YY07a, YK03, Yeh02, Yu09c, Yu09a, ZN05]. distributions-positive [Mór09]. disturbances [Lai01, Wat01]. Divergence [EHMM08, BBR07, DN09, Haf06, Yoh08]. Divergence-based [EHMM08]. divergences [Tom09]. dividend [LL09a, YZ09, YH09, YH08, YZG08]. divisibility [BK08c, DMRA07, KN08]. divisible [Bar00, EsO08, Jur07, KK06a, KM09b, MPA07]. division [IMS06]. do [KL00b, Yan08]. does [BS06d]. domain [AM07, DF02, Has05, Mak08a, MR08b, Sch01b, Seg01, Shm06, Wan05a]. domains [BB07b, FY00, KZS08, SLF09]. dominance [Nam03, Oht01]. dominated [Jia09b, MP01a]. domination [CP05a, Has02, PDM01, Zha09a, dlPZ02]. Donati [Hob07]. Donkey [Let02]. Donoho [DH09, Ger02]. dose [MMB07, TSGG02]. Double [ASAK00, Aic09, AG09a, Cha04, Din00b, GS07, MY04, NC09a, QH09, XZJ08, ZZ08]. double-covariate [Aic09]. Double-ranked [ASAK00]. double-sided [XZJ08]. doubling [HZ09a]. Doubly [YW04a, FS01, Fer00, Lin09, Mad02, NO09, sS09c]. down [GS08b]. downcrossing [CP00]. drift [ESOO09, GK04b, HH02a, LT00a, McG06, Pou08, Shi09, SL00, Stu00, Tan07, Xi04]. drifted [LS05b]. drifting [LdC07]. drifts [HY04a]. driven [BN08, CL07, El 08, HNS08, Ing00, NS06b, Otm09, Pre09, RH07, YYL08, ZB09d, ZB08]. driving [Shu09]. drop [MOM07]. dual [BEA09a, CKLX06]. Duals [KPI09]. due [FD09, Hob07]. Dufresne [HY04a]. dummy [Pro08]. Dunnett [KL00c]. duration [HWZ09a]. durations [CP06a]. DVRL [AZS08]. Dynamic [AES05, LLL00, Dav04, Lyh05]. Dynamical [Jon08, FF08, Hil03, KS00]. dynamics [CM08b, LV07, Luc00]. Dyson [BCT08]. each [Muk09]. EB [Gut00]. EC2SLS [BL09a]. EC3SLS [BL09a]. ECM [KSG08]. ecosystems [PL09b]. edge [Dai05, DZ01, JY02, PRB99, Pal01b]. edge-reinforced [Dai05]. edge-transitive [PRB99, Pal01b]. edges [CBR00, Kol00a]. Edgeworth [FMS04, Mar01b, MY04, SB00, Sze08a]. Editor [Joh07]. Editorial [Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06p, Ano06q, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h]. Editorial [Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. Effect [Aic09, And05, BB08a, BB07c, Bru09, LTY08, Nic08a, SS07a, WH01, YZ07a]. effective [TSGG02]. Effects [EK03a, AJ00, ADJ06, ADJ07, AZ04a, AH06a, BL08, BBP09, CKM03, cC01c, ET05, FM03b, Hut03, Ima08, Jac08, Jia01a, Jow06, KGC09, LB03, Luc00, MLT00, Nat01, NG02, Omo03, Van08, Wu08b, ZF06]. Efficiency [Mir01, QSP01, AT08b, BES01, Bos07b, EK03b, Jen04, KL03a, KM02a, SS00b, SKJ07, SK01]. Efficient [BM04, CCG09, DG06b, HZ09b, Mai08, Mil07, Omo07, PPSW04, AM07, AH06a, BCC02, Che09, GP06, GDM06, GSKP09, HT05, Hut02, HB01, MRP04, Neg01, PS02, PS09a, Ver07, YZ07b]. efficiently [BM06]. Ehrenfest [BMW06]. eigenstructure [PB03]. eigenvalue [Asc09]. elemental [OH07]. elements [AAV02, AHV06, CV01, HCSV00, HNRV00, HCV01, QH09, Sun08a, Sun09a, Tóm05]. Elfving [Kra09]. ellipsoidal [ST03b, TS01]. elliptic [LL04, Ren01, RSY02]. Elliptical [FJS03, Has05, BB09b, CDRV01, GQT07, GV08, Has06b, Has06c, Has09, KM03b, KN01b, Lan06, LT06, LP09, Lop08, PMVL09, Sch02]. elliptically [BG09, Nad05]. ellipticity [Gag02]. Ellis [Nyr05]. EM-algorithm [Gut00]. Emperors [KBH07]. Empirical [ADR05, AAB04, BGM03, CQ03, CC08c, DR01a, DS02a, HMR08, HWZ09b, HH06, HZ09c, JH02, JYZ08, KE05, KZ08b, Lia09d, LZZ08, QZ07, QLL09, SH06, WC03, YZ06, Zha06, ZLQW07, Ahm02, Arc02, Bar07, BdB09, Bor00, Bra02a, CM08a, Dar04, DE06, Din00a, Din03, DOKN07, Esa04, Faz03, Fer09, HO08, KV02, Kou03, KL06, LS02a, LG03, Lia04, LZ01, MZ02, MU04, MS07b, Paw00, PPRW08, QBZ09, RS08, Ser02, Sha01, Shi02, Tsa01, Tsa04, WV08, Wu04, YZ07a, ZWY05, ZY08]. empirical-type [CM08a]. endogenous [Lec08]. endpoint [GT04]. enlarged [Otm09]. entire [Sim06]. entrance [LZC05]. entropies [AF09, ZN05]. Entropy [ET07, AE00, CG02a, Cha06b, Mon07, Son00, Yu09c]. entry [DS00a]. enumeration [AB07]. environment [dCB00b, Dev07, FS09]. environmental [FD09]. environments [KO08, SK06a, SY04, Tak04, YK00, YP09]. epidemic [GS08a, HYL04, Kyp09]. epoch [QL06]. equal [Yan08, YOH00]. equality [Bha05a, TW06, Wul08, YK00]. equation [EGM06, HKV04, HNS08, LT08, SS01c, SK01, ZAC03, Zha09b]. Equations [MN08, Ago06, AL00, AR03b, BHY09, BM03c, Bez08, Cai03, cC01c, CA00, Drt08, DaY08, FT01a, FT01b, GJZ06, HR08a, Hal09, Ign09, Jia08, Jia05, KL08b, Kur08, LT00a, LX05, Lin09, LT00b, LR02, LL09c, MW09, MN04, MS08c, Nic08c, NO09, Psa09a, Ren01, RSY02, RH07, SL09a, ST01a, Ser02, Sit02, Tso00, WD08, Wan08c, WH09, WH07, Xu08, Yan00a]. equicontinuity [Ros06]. Equidistant [KS08b, Jar03]. equifrequency [Bur02]. equilibrium [DHS00, GTH06, LL04, NP08a, XWS05]. Equivalence [OE08a, Cho08b, DWZ05, Efr03, NOP01, Pro08, Pro04, WBG04]. Equivalent [KKY09, Guo08]. equivariant [GO01b]. Erdos [KO01, BW06a]. ergodic [Cha06b, CH01, DZ02, Deb09, GO02, Gne03, HH00a, JY08, Neg01, SP00, Tal08]. Ergodicity [CS05a, PG09, BRR01, CW08b, FM00, LS00a, Li07, LJ01, ZCLH01]. Erlang [Leo09, MSZ08, Sun05]. Erratum [Ano06o, Bos04b, Gat09, GV08, Lin00, PS09b, PK08, Rou01a, SF12, Yu09b]. error [AFS^{+}07, AR05, AG08, Baí03, BM01a, Bir08, BRR01, CM04a, Che02a, Che02b, Dav05, DG08, DGZ02, DS02a, Duc09, Esa04, Fuk05, Gir04, Gor09, HH02c, JSV02, JNS02, Kos00, LTC04, LOS02, Liu07b, MSW09, Oht02, ODO07, ÖK07, PF07, QF03, Ten01, Vou08a, XY07, vE01]. errors [Bha05a, Bia02, BdC00, CGK06, CHD08, CS08c, CHN04, EP02, FRG06, GP05a, GT04, HWL08, HWZ09b, HKLB07, Li00, LF07, Mor05, NIK02, Özk08, PMVL09, PM01, PMA08, SS07a, SN09, SS00b, SS06b, SFS03, Son09, SZ08a, Udd08, Wan02b, Wan06, ZWWC07]. errors-in-variables [FRG06, HWL08, HWZ09b, Li00, Son09]. es0971 [Ano01j]. Esscher [Mon07]. Esseen [Bos04b, Rou01a, Bos03b, Rao02a, Rou01b, UWF06]. Esseen-type [Rou01a, Rao02a, Rou01b]. Establishing [HH02a]. Estermann [Lau05]. Estimability [KC08]. estimate [BM06, Bru09, Dub08, Fot09, GP00b, McG06, Rou00, She09, TG02, TR02, She08]. estimated [Bru09, JSV05, JGMGPM03, KN01a, MS07b, SXLX07]. estimates [AJ05, ADR05, BBR05, BB03, BB02b, BK01b, CT04, CB01, Cor04b, DK08, FI06, GP06, Gir04, Gui01, Gut00, Had03, Hen00, Lee04b, LR08b, MBY00, Pro09b, RS00, ST03a, TMN01, WS04, XTW06, Yat02, Yoh08]. Estimating [AT08a, CBR00, DS00a, Dav05, DMP01a, FNA09, GT04, Hut00, KK01, Lia09c, LK02, MN08, MSW09, Pen01a, SDR01, VJ08, WB05, vEZ01, vEZ02, Ago06, AS06, AS02, cC01c, Che09, CA00, Dro06, HH03a, KF08, LRM00, LZD09, LT08, OR01, Par01, SK01, Tar01, TP03, Tso00, Yan00a]. Estimation [AAC00, BCG01, Bur02, Dup06, ESOO09, FR01, Fan02a, FG04, FB02, Fur02, GP05b, HC02, KR09, LVL09, Li00, Lim06, Liu07b, Llo02, MW09, Mar02, Mod03, NIK02, Omo03, OS02, PSS00, gPB02, PK04, PNJ08, RM03, SSB09, Shi04a, SS05b, WZW03, ZWWC07, dCD09, AGG02, Ahm02, AZ04b, AFMGJG05, AW04, AFS^{+}07, AK07, Ars09, ALOS08, BC06, BC00b, BCAC01, BR08, BP02, BK08a, BES01, BBZ08, BV02, BYB08, Ber08, BM01a, BM04, BBG05, Bha08, BW05, BD01b, BC03, BGN08, BC01, BL02, Bre08, BM01b, But00, Cai01, Cai03, CCG09, CD01a, CO06, CF01, CYJL01b, Ciu09, CAM03, CHL01, CS05b, CFUOV00, CDRV01, CHN04, DL06, Dar03, DG08, De 08, Deg08, DKM02, DG03, Din00c, DR01b, DZ01, DW00, DG06b, Efr00, Efr08, Eni06, EP02]. estimation [EK03b, FS01, FRG06, FJ01, Gai07, GS02, GSGMFB03, Gas09, GdL00, GD04, Gre09, Gro03, Haz00, HM09, HZ04a, Hil01, Hil03, HS07a, Hua02, HF04b, HH06, HZ09b, HYL04, Hug04, Hug07, HLM06, Hut03, HY01, HBK06, HKLB07, HK08, HC09, Ili00, Jan08, JL07, JR08, JMP06, KKP00, Kar02, KE05, KKH02, KK00c, Kos00, KH08a, Krä07, KMV09, Kyp09, LB07, Lee03, LCPY04, LZ02a, Lia07, Lia09a, Lia09d, LS00c, LC04, Mad02, Mad08, Mai08, MS09a, MZ05, MP06, Mar01c, Mas02b, Mas02c, Mei07, MN01a, Mir05, Miz00, MP08b, MZ04, MN08, Neg01, Nki05, OL02, PR09a, Pel05, Pen01b, PS02, Per02, Pon06, QSP01, dR06, RS06a, RS07, RC03, RM08, SVB07, SS07a, ST03a, SN04, sS09c, SV02, Shi09, SFS03, SV00, SXL05, SZ01, SW07b]. estimation [SZ08b, TMH02, TLRvdV08, Tri04, Tso00, UI00, VT00, Vou08a, Wal02, WZL00, WP08, Wan09a, WC03, WCC07, XY07, YZ08, YW09, YC06b, YZ07b, YV06, YWS09, ZPK08, Zha01b, ZL07a, ZY04, ZB05, dHC06, vE00]. estimations [Gua07]. estimator [AVV09, Baí03, BHGRV08, BH03, Bis01, COS03, Che00c, CP06b, CFH02, DB09, DH09, FJ07, Fuj07, Fuj00, Gao01, GS07, GO01b, Ger02, GPC09, Han05, HW05, HY00, IP03, JNS02, JKP01, KL08a, KSW06a, KSW06b, KZ08a, Kos02, Küh01, KZ08d, Lab08, LLOS09, LOS02, LOS03, Li00, LYZ08, Lon09, Man08, MP02, Mid08, Mim08, MRP04, MB06, Nam03, NS09, NC07, NT03, OH05, Oht01, Oht02, Osm01, OS06, ÖK07, Özk08, QF03, SS05a, SWM09, SKJ07, Shu04, SS06c, SKS09, Sub09, Tia09, Vou06, Wan00a, WZ06a, XL06, Yan03a, Yu07b, YW04a, ZWY05, ZR06, ZCCH09, dUÁMM08, vE01]. estimators [AM07, AF01, AT08b, BL09a, BB04, BBR06, BBR07, Ber03, Bia02, Bir08, Bis01, Bis06, Blo07, BR06b, CP05b, Che08a, Che02a, Che02b, DFK03, DS01b, Din00b, FJA09, FM05, FC03, FH06, GQR06, GJ08, GV06, Gne03, GO03, GN08, GRT07, HW01a, HZ03, HZ04b, Hut02, HPR02, Jan03, JB05, KH01, KSH07, KP02, LS06a, LS08c, LL05, MM02, MSW01, Nai01, NO06, NBK08, Neu07, NMO08, Ome08, ÖK07, Pap06, PP06, PPSW04, PS09a, PK07b, PK08, SB06, SW06, Son00, Ten01, Ten03, TW06, TS09b, Tom09, Var08, Ver07, VV00, WW01, Wan02a, WV08, Wul08, YSkPkL02, YZ05, YOH00].

Euclidean [CBR00, JY02, LZ01]. Euler [FT01a, FT01b, ST01a, Sun07a, TSJ09, WH07]. Eulerian [Bro08a]. evaluating [LB03]. Evaluation [EKS05, KTZ03]. evaluations [Ryc02, Ryc09]. events [Dub08, Tan09]. Every [AL06]. everywhere [Zha01c]. evidence [SZ02a]. evolution [Bez08, Sta05, Zha09b]. evolving [Bri00]. EWMA [LWW09, VvdMD08]. Exact [LOS02, MR03, OM06, Oht02, Sgi01a, Ste06, SBH03, TT03, Wit02, YKB07, AVG08, Dev09, Gal09, Kab08, KMS08, Llo08, Lou00, Lou03, LF07, Lu06, Lyh05, MW00, OW03, Sgi01b]. exactness [LSEB00]. example [Eng04a, Kan07, Ros06]. Examples [CH01, Sch01b, Muk09]. exceedances [Ast01, BE09]. exceptional [Ber01]. excess [LS04b, SD06]. exchange [RMLD07]. exchangeability [Röl08, Ter03]. Exchangeable [GST00, Tso00, ETC00, Ery08, Ete07, GM00, HP00, IW04, JH02, MSW07, MS04b, ST01b]. excited [DDP08]. exciting [Wan05b]. exclusion [JS07]. excursion [TSJ09]. Excursions [FBVW00, TS01, YY08b]. exist [KL00b]. Existence [AH06b, BM03a, Bez08, Wan08c, AB08b, Ale08, CS05a, CKR02, Guo08, HR08a, Hal09, Roz04, Sha04, SS07d, Zha09b, ZZ08]. existing [Tot04]. Exit [Mak09, BN08, CY02a, Kin01, LL04, Mak08a, SLF09]. exiting [Vol06]. expansion [Deh07, DS02b, DvZ05, FJA09, Ist06, IKIS09, MY04, SB00, vE01]. expansions [FMS04, GR00, KK04, KKAK08, Mar01b, Pen02a, Sze08a, WW08]. expectation [Eis08, FM03a, Jia06, Jia09b, Pap01, STKF08, TS01, Won09]. expectations [Cri04, CP05c, DR03, DR04, Din00c, GO01a, Gha00, HHC09, RMLD07, Ryc09, TB09, Zha09a]. Expected [Mah04, ST03b, SL09b, Duc09, HF04a, HK03a, Kol00b, LW08b, Ren09a, Sun05, dlCV01, dlCC05b]. experimental [Jou07, WK07, ZNS08]. Experimentation [WK07]. experiments [FY00, Geo07, Ger08, HW01b, HT05, Hug01, Ima08, Li02, Mid08, PM01, STM01, Tia09]. Explicit [DR00, Nad08, Won09, Xu08, DvZ05, GC03, Hoo02, SC05, Wit05, ZCL08]. exploding [Stu00]. explosive [HKLB07, Nag09]. exponent [MP08b]. Exponential [BLB08, CH09a, LS05b, MN01b, SM09, WD08, AA01, ADL05, AM09, Ann07, AR03b, BLP03, BLV04, BSCN09, Bar02b, Bri00, Cha01, CTY02, CW03, CF06, DS02b, EL03, Esh04, FC08, GM03b, HY04a, HH06, JR08, JNS02, Kag02, KT03, KKAK08, KBH07, Kim00b, Kok01, KK06a, KM09b, KK01, KMV09, Lia09c, LT00b, LBSM00, Mad02, Mad08, Mas06b, NT06a, Nic02, NZY09, Oli05, Pal01a, PS06b, Pom01, RB06, SSB09, Sun07c, TG02, TT08, VJ08, YT08, ZB09a]. exponential-geometric [ADL05]. exponential-Poisson [BSCN09]. exponentiality [AZS08, BH08, GJ02, Kla00]. exponentially [Aur08a, Cer07, CP06a, FJ01, FvE01, Gup09]. exponentials [Won09]. exposure [GLV00]. expression [IR08, Wit00]. Expressions [WM06, ZN05, Bos09, Nad08]. Extended [Che08b, Kum02, PW06, Pri01, CW03, SK01]. extending [HCW08a, KL08b]. Extension [Ery07, APS09a, ADL05, CG02a, Fah00, GG02, GOPB09, HY04a, Hör06, KW00, KO01, LSSG08, PL07, SSR08, VB08, dL05a]. Extensions [CB07, KH07, Pet08, Pyc03, SS07b]. extensively [Leh05]. Extinction [YP09, Gho03, HR09, MN09]. extraction [BH09]. extrapolation [DS02a]. Extremal [CP01, LP03, FP08, Fra06, KN01b, MF05a, PMM09]. Extreme [Hai03, Ryc08, Sim06, BEA09a, BV02, CH09b, DG03, FR02, Fal05, FF08, GJ08, GN08, He08, Hür03, HLM06, Jur00a, Pou08, Sch01b, Zha07b, dHC06, dlCV01, dlCC05b]. Extremes [Fer03, HS03a, MPY03, BDG07, DdHL03, FS09, HM06, Kol00b, LRW09, PW08, Seg05]. extremum [Krä07].

F [Cui02]. faces [QS05]. facilities [SL06]. factor [AZ04a, MP01a, SD09, YB07, ZAC03]. factorial [AM05, AZ04b, AH06a, AHL08, But08, EKM02, EK03a, Ger08, Lia01, QC04, YZL07, YW04b]. factorials [Man05a, ML03]. Factorization [GL03, HBA05, Zhe01]. factors [DDM06, DS00b, FD09, KH08a, MPS06, Mon08, SL00, YWWY08]. fails [Kan07]. failure [BLS05, Che00b, DKM02, FS01, FE08, FD09, GJ02, LL08, Llo05, MK08c, Nav08]. failures [AB07, OL02]. fake [Ole08]. false [GSTS07, GR08, KW02, Oeh00, Oeh01]. familial [SS04]. familial-longitudinal [SS04]. families [Ann07, BLP03, BLB08, Bri00, DR03, DR04, HW01d, Kag02, Kok01, KK06a, Mas02a, Mas06b, Pom01]. family [AA01, BPR01, BHH08, CF06, FC08, GQT07, GV08, JNS02, KF08, LX00a, Lia09c, LBSM00, Mut08, Nic02, NZY09, Pal01a, Pui08, VJ08, Zhe01, ZJ09b]. familywise [Gor09]. Fast [LLS06, Fai01]. Fatou [KO04]. Fatou-type [KO04]. FBSDE [FW08b]. FBSDEs [Yin08]. Feasible [FGMS08, SO00]. feature [FWW01]. Feller [Bra02b, CD00, CD01b]. fence [JNR09]. few [Pat09, QS05]. FGM [Has01]. field [CW08a, DKM02, Duc09, Hu08, MF05b, MC07b, Per09, YL00]. fields [Ast01, BM07, BHZ00, Bia02, BS01b, hCW09, CWTB09, CMNP07, Che08a, Del03, FP08, GS09, KKC08, LCPY04, LSEB00, LS06a, LW03, Mal01, MS03a, MP01b, NT06b, PS00, Sha08b, Sha04, TSJ09, WL08a, Wan09a, Xia09, ZW01]. filter [BK01a, GC03]. filtered [Lef08b]. filtering [Efr03, FF09, Gri02, KL08b, MY04, Piz09]. filters [CP05c, JD08b]. filtration [Otm09]. filtrations [Zap03]. finance [Otm09]. financial [CS06, Irl04, Nic08c, RS07, TV07]. financing [Sal02]. Finding [JK08, DS00b, MW00]. Finetti [BKM04, GM00]. Finite [CS06, VV00, AHEA04, ADR05, Bos04a, Bri00, CT04, CWZ08, CQX08, CV04, Din00b, ES09, Had02, Hil02, HWL08, KSW06a, KSW06b, KZ08c, KS05, Leó01, LR07, Mir01, Nic08b, PP09, Pro09b, Sim06, Spr01, Spr03, WD08, Yat02, Zuo01]. finite-dimensional [Bri00]. finite-order [HWL08]. finite-sample [CV04]. finite-time [KZ08c]. finiteness [GM03b]. Finitizing [GLC09]. First [BLS04, Kok01, Ros09a, ZB08, ADJ06, ADJ07, Als01, CY02a, DY04, Dit07, Gir09, HA03, HZ03, HZ04b, HZ08, HBK06, JKZ09, LC00, LN02, LT00a, LWD07, PP09, PMVL09, Pen08, PS06b, SLF09, TS09a, WY08a]. first-crossing-time [Gir09]. First-order [BLS04, ZB08, HZ03, HZ04b, HBK06, LN02, PMVL09, PS06b]. first-passage [Dit07]. Fisher [AA07, BS06b, Bra02b, CWTB09, Drt08, HBA05, KMS08, KY04, PK07a, PBZ08, PB09, SBH03, TT07a, WH05, WJ05, Zhe01, ZG01]. fit [BN00, Ban08, BH08, BM06, Bur00, CS08c, Fie03, FG06, GV02b, HcW08b, JG04, JSV05, KS05, KY06, Kra04, LS00b, LP09, Mak08b, MU04, MS07b, Son09, YW07]. fits [HC02]. fitted [DP03]. Fitting [BT04, ZM06, Gal01, Gør02, LSL00]. Fixed [Ioa04, BBP09, GRT07, IF08, SB06]. fixed-width [IF08]. FKG [YK00]. flats [Lia00b]. Fleming [dSF06a, dSF06b, dSF08]. flexible [CL09]. flights [CLO09, Gre09]. floating [BRR01]. floating-point [BRR01]. Fluctuation [LZ06, CQX08]. Fluctuations [HL01, MW06, NZ00]. folded [PR09b]. foldover [AHL08, But04, FLQ03]. follow [sS00a]. follow-up [sS00a]. following [RH02]. food [VT00]. force [KZ08c, YZ01, YZL08]. Forecasting [Dav04, TAP05, MW05]. forests [PI09]. form [BD01b, CHL08, Com07, GA00, Ole01, Sze08a, dJP04]. formalism [Mab08]. formation [BLB08, Li04]. forming [BHZ00]. forms [BNO03, BGK07, GHL01, Gha00, HC06, KM03a, KM09a, Kra02, LH01, Lop01, PMJ05, PMJ08, Ruk09]. formula [Ano06o, BK04, CA06a, CA06b, DO07, Gha00, Gir09, Leo09, SY07, Tan09]. formulae [GvLM04, TS01, TT03]. Formulation [Hor04]. Fortet [Zah00]. Forward [Hyn07, AH06b, GJZ06, Kin08, WX09]. Forward-backward [Hyn07, GJZ06]. Foster [JD08a]. Founding [Joh07]. four [ASS02, CKS04, PK07a]. four-parameter [PK07a]. Fourier [AD01, XNF07]. fractal [RMAA01]. fractals [Din00a]. Fractional [Nie04, AM05, AZ04b, AH06a, AHL08, AL00, BJT03, Bis08, BGT04, CH07, CI01, DDN08, Dun06, DF11, DF02, DvZ05, EN03, Eni06, ESOO09, Gap04, HNS08, IKP08, Ist06, Jon02c, Man05a, MNN09, ML03, MF05b, MMV01, MN04, NS06b, PB03, QC04, RMAA01, SZ05, SZ09b, Yan04, YYL08, YZL07, ZBZ09]. fractional-order [RMAA01]. fractions [ALP08]. Fragment [GH05a]. fragmentations [GH05a]. Frailty [AH02, BN00, FE08, GQR06, MRK08, PZ08]. framework [HHC09, OC04, RB08b]. Frechet [BT02, HN08]. free [AB05, ARB09, Ban08, GH00, Hu08, Jur07, KM06, KP02, OR01, Ser00]. free-infinitely [Jur07]. freedom [Jon02b, Wat01, Wit02]. frequencies [NIK02]. frequency [DZ01, DF02, Shu03, Wan05a]. frequentist [CM04a, CM08a]. Front [Ano03t]. full [Geo07]. Fuller [Coo02, Sen01, Sen07, Sen08a]. function [AE00, BC06, BC00b, BA03, Bla08, Bou06, Che00b, Che02b, CL08c, CDRV01, DB09, DH09, Deg08, DK02, FS01, Faz03, Fer04, FvE01, Gal01, Gal09, GS02, Ger02, GMdP04, Gua07, Hen00, HBA05, HPR02, IP03, Jar08, JO08, JR08, JMP06, Kos09a, Kos09b, KL06, KS00, Lau05, LPS09, LN04, Lia07, Lil04, Liu03, LW08b, LBSM00, Lud09, Mak08b, MU04, MK08c, MU01, Mod03, MT07b, MSW09, NBA07, Nav08, Neu07, NS02, ODO07, Ose09a, PK04, PP00, Pyc07, RS00, RS06a, RM05, Ros06, Rou00, Roz09a, RS08, SDR01, SS01b, Sgi09, sS09c, SV06, SY07, SZ01, TP01, WY08b, Wri00, WZW03, Yi05, Yu07b, Zah00, ZZG06, Zhe01, dFF02, dUÁMM08]. Functional [Sto03, WH01, APV06, APS09b, AG08, ALOS08, BB09b, Bat05, BHGRV08, BHH08, Bez08, BD01b, BF00, CC01a, CB07, CL08b, Coe08, Com07, DMB02, DR01b, GW05, GK09, HR08a, Hal09, HKV04, Kos09b, LLOS09, Lal08, LT00b, MS03a, Moo08, MB06, PS09b, PS09c, Poz04, Rom02, Tal08, VW08, YYL08, ZL07a]. functional-algebraic [DMB02]. functional-coefficient [ZL07a]. functionals [Che00a, Clé01, HY04a, Jan03, KKM^{+}08, Maj05, MY08, PPSW04, PS02, RAS08b]. functions [AF01, Bar02a, Bar05, BM03b, BBG05, BR06b, CM08a, CCC02, Che07a, Che02a, Che04b, CÖ02, CDMLDK01, Dev09, DR00, Fai01, Fin06, For00, GDF01, He08, HC02, JFZ05, JB05, KKP00, KS04, Lan08, LQ07, Li07, Lin04, Lop08, LK02, Ma03, Mar06, MK08b, Mas06a, MZ02, MZ06, Mir05, Mna08b, Moj01a, Mor05, MRS06, NQMRLÚF01, NQMRLÚF03, NP08b, NC09b, NZY09, Rao02a, Ren09a, RLÚF03, RMLD07, Rou01a, Rou01b, Sas08, Sch00, SG06, Sen08b, SH06, Sim06, Sou01, Sza01, TG02, Ter03, Ter08, TP03, WZ09b, Woz09, Yeo07, YWS09, ZCLH01, Zha07b, ZP09, dHC06]. Further [FL02, HY04b, PZL04a, Piz09]. future [AM09, DS00a, RB08a]. futures [LB03]. fuzzy [AR09a, JKK05, LS08c].

gain [FMS00, Fro03]. gains [CS07, SL09b]. Galton [Dai05, Gho03, GMM00, Kin08, Ma09, MN09, MS02b, Mut08, SV00]. gambler [ES09, Lef08a]. gambles [Bea07]. gambling [Poz08]. game [Sza07]. games [CS07]. Gamma [Cer08, SF12, Alz06, AL08, BDM08, Cha04, CH04b, Fur08, GQR06, Jur00b, MN03, NC09a, Rou07, ST04b, Yu09c, ZPK08, KT08, PL07, Ris05]. gap [IMS06, dUÁMM08]. gapped [MGW02]. GARCH [AB08b, BHH08, Haa08, Haa09, HB04, LS08a, Liu06, Liu07a, Mim08, PBH09, Wan06, BB07b]. GARMA [KH08a]. Gärtner [Nyr05]. Gärtner-Ellis [Nyr05]. gas [AFS^{+}07]. Gauss [IPP04]. Gaussian [AZB07, ARN08, BB09a, BJT03, BSS01, BB05, BC03, CC08b, CL06b, CH04b, Coe08, CG02c, DZB04, Del03, Duc09, Dud08, FF09, For00, Gri08, HH02b, Hei01, Hu08, HZ08, Kab09, Kar02, KK04, KN08, Le08, LS02a, LSSG08, LHLC04, MB08, Mar01b, MY04, Mei08, MV01, MS04b, MC07b, Nas06, Nin05, OS05, PS07a, San05, Sou01, Vit08, Vou06, Waa04, Wor01, Xia09, Yu09c, ZJ09b]. Gaussian-generalized [FF09]. Gaussianity [HCW07]. GCV [AD01]. Gebelein [Nov04]. Gegenbauer [Sun07a]. GEM [YSN01]. Gene [PI09, IR08]. General [BSS01, CLR06, DaY08, BC00b, BM03a, BV07, Bib05, Bla08, BM03d, CD01a, CP00, CL08a, Che02a, DMP01b, Deb09, FM04, FS08b, Gor07, GS08b, He08, HYL04, Ilt00, JS07, JY02, Ker08, Kyp09, Lin08a, LH08b, LS00d, Mai08, NC09b, OC04, PP09, PL09a, RW00, RB08b, SM06a, TW06, WX08, XNF07, XWS05, YSY08, ZAC03, dlCC05a]. Generalised [BC01, ZJ09b]. Generalization [PR07, BSCN09, Bea09b, EG01, Lan06, Nic08b, NFL09, Pet04, SSR08, ST04a, YGK08]. generalizations [Man05b, TO03, VU07]. Generalized [DZ01, HC07, HKLB07, Jur03, Kev07, LTC04, Lef08b, MN08, TP03, TJ09, Vou08a, Yeo07, ZS08, AV01a, AZB07, AGJ04, AES05, AR09c, BEA09a, Bou09, BK08c, CW08a, Cer08, CM04a, CXH09, CH04b, CH09b, CB01, Cor04a, Cor04b, CKR02, DZ02, DN09, DW00, EM08, ET07, EK04, FW08a, FF09, For00, FV09, FB02, GP00a, GOPB09, Gor09, HZ05, JBB08, Jia01b, JFZ05, JL09a, KM08, KK05, KSO05, KMR00, Kum02, LLOS09, LJ06, Li09a, Li09c, LC05, Liu07b, LRW09, MK08a, Mir07, Mon08, PB03, PZL04a, QC04, RAS08a, RH02, Sas08, Son08, ST04b, SK01, TT07a, TJ06, Tso00, TT07b, WBG04, WP08, WV08, XL09, XW08, Yi05, ZR06, ZB09d, ZZ08, dCD09]. generated [BCD09, Din00a, KB01, KKH08, KK08a, KKC08, Moo08, NK03, OMJ^{+}04, OK06, PMM09, Zap03]. Generating [Bis04, Che07a, CÖ02, FS06, MRS06]. generation [Fal05, FY00, MN09]. generator [SC05]. Geometric [AM02b, Hue03, ADL05, AK09, BRVM03, BTW01, BK01b, BRR01, Dem08, DMMY00, Eis08, EK03a, Fot09, Gup09, GK04b, HH02a, HA03, KY09, Kim01, LS00a, LJ06, LJ01, Mak09, MPY03, PW00, Psa09b, Ruk07b, Wan02d, ZB09a, Pou08]. geometrically [Hua03b, Sto03]. geometry [CW08a]. geostatistics [MU01]. Gerber [ZZG06]. Gibbs [Cer08, FS00, MP00, MS03a, NBW07, Omo07]. Gibbsian [NT06b]. GIC [IWWJ04]. GIG [FJ04]. Gini [JG04]. Girsanov [MV01, SY07]. Gittins [Son08]. given [Dub03, GMM02, Hut02, Hut03, Leh05, MB08, Mül01, Ost01, WC04]. giving [D’A06]. Glivenko [Che08b, Moj01b]. Global [DGJ09, JJ09, Kos02, MS03b]. GLSE [SO00]. GMM [JS09]. Gnedenko [Gut06]. good [KL00a, TS09b]. Goodness [FG06, KS05, KY06, Kra04, LP09, YW07, BN00, Ban08, BH08, CS08c, Din00b, Fie03, GV02b, HcW08b, JG04, JSV05, LS00b, MU04, MS07b]. Goodness-of-fit [FG06, KS05, KY06, Kra04, LP09, YW07, BN00, BH08, CS08c, GV02b, JG04, JSV05, LS00b, MU04, MS07b]. governed [Mac09]. GQL [JS09]. GR [TMN01, TR02]. GR-estimate [TR02]. GR-estimates [TMN01]. grade [Kow00]. gradient [Mon06, PR09a]. gradually [AS02]. graduation [GJP01]. graph [AR02, GBQ03, Gup09, KL03b, LO09]. Graphical [Voc06, GL03, LLL00, Mei08]. graphs [FD05, GR07, HF04a, JY02, KM06, LR08a, LS02a, PRB99, Pal01b, YWWY08]. Green [BV08, KD04, Pyc07, ZR06]. Grossman [KN05]. group [Aic09, GT08, HW07, KC03, KKRV01, Zap03]. group-valued [Zap03]. grouped [BvdV08, BvdV09, BBR07, Mei07, dL05a]. grouping [MN03]. groups [CF00, HC04, Hil02]. growing [BM05, Hua09]. Growth [Ath08, And05, Ano02f, Eng04b, KRV05, Pan04, WHSL08]. guarantee [HH06]. Gundy [Nik08, Ren08].

Haar [Gir04]. Hadamard [AK07, PZL04a]. Hájek [HWYZ09a, Rao02c, Sun08b, HWYZ09b, RB08b]. Halton [RS06b]. Hammersley [RS06b]. hard [Che08c, Dro06]. harmonic [Kin08, Pui08, Sen08b]. harmonics [BM07]. harness [BW06b]. Harris [CLR06, LL05]. hash [CM01]. Hastings [FM00, SBF00]. Hausdorff [LFdVA05, Mna08a, Mna08b, Szu02]. Hausman [BL07]. having [Muk09, Por02b]. Hawkes [MT07b]. haystack [JK08]. hazard [AE00, BGHP04, BBG05, DG06b, Dup06, EP02, GVMSN04, HBA05, Lau06, dR06, WZW03, ZP09, Zhe01]. Hazards [sS09b, AH02, CM07, DS01b, Di 00, GE07, Hat06, SZ06, Yu07a, Yu09b, ZZ06]. Heavy [CFR06, Ibr08, AKK^{+}02, BB07b, CS06, HS03a, HK03b, Kle02, Lin08a, Liu09, MY06, NIK02, Pen01a, Psa09a, SD09, TSJZ01, WT04, Wan06, WCS07, ZJ09b]. heavy-tailed [CS06, HS03a, HK03b, Kle02, Lin08a, MY06, Psa09a, SD09, TSJZ01, Wan06, WCS07, ZJ09b]. Heavy-tailedness [Ibr08]. heights [Ale08]. Heisenberg [D’A06]. Heisenberg-type [D’A06]. Hellinger [Hil01, Hil03, SV00]. heredity [GD04]. Hermite [Wil05, Wit00]. Hessian [BX00]. heterogeneity [GL00, Han08, ST01b]. heterogeneous [Hug01, Lim05, WK07, ZB09a]. Heteroscedastic [CPG07, BB02b, CWL08, Dav04, Efr03, HKLB07, Kak08, Özk08, WIL01, YW04a]. heteroscedasticity [XW08, YC05, YC06b]. heteroskedastic [BC05]. heteroskedasticity [FCN02]. Heyde [SL09b]. hidden [Alt08, Cho08b, EK03b, LLL00, Spr01, Spr03]. Hierarchical [LQ07, Lia00b]. High [AVV09, Oli01, Ast01, BR06b, Che04b, DLS00, FS02, LX08, OH03, RR01, UWF06, YC06a]. high-level [Ast01]. Higher [Bro08a, AB08b, BGM03, Mab08, SQMK06, SD01, TSJ09]. higher-order [AB08b, SD01]. highest [BM06]. Hilbert [Bos04b, FT01a, Bos03b, FT01b, Kim01, KKH08, KK08a, LT08, Liu08, LL09c, Sit02]. Hilbertian [Gui01]. Hill [GPC09]. Hinkley [FJ07]. Histogram [LS08c, BMR07]. histograms [Bur02]. histone [YPD09]. hit [Bél00]. hit-and-run [Bél00]. hitting [Che07a, HF04a, JKZ09, Pal09, dFF02]. HIV [Zel06]. HIV-1 [Zel06]. HNBUE [Kla00, Pel00]. HNBUE-type [Pel00]. Hochberg [GR08]. Hoeffding [Bea09b, Bou09, GO02]. Hölder [LLW09, PMA08]. holding [Hil02]. Holgate [SZ02a]. Holm [GSTS07, GS08b]. Holm-type [GSTS07]. Homogeneity [XWL09, Cer02, NZY09, SGK09, Ste06]. homogeneous [Dow09, Hu06b, Hu09, Jon08, Led08b, dL01, NBK08, Yan03b]. homology [Lou03]. homoscedasticity [Cac01]. Hopfield [LV05]. horizon [GJZ06, Yin08]. household [VT00]. Huang [Ano06o, CA06a, CA06b]. hunting [PI09]. Hurst [IKP08]. Hüsler [Has05, Has06b]. Hutchinson [Hür03]. hybrid [PBZ08, PB09]. hydrology [Lef08b]. hyperbolic [DG07, Mah04]. hypercube [BC06]. hypergeometric [ALP08, HS05, Kum02, RC00]. hypersphere [FG06, Fig07]. hypotheses [De 08, LGH03, Tor06, TT07b, YKB07]. hypothesis [AY02, Erm00, GVS00, KKRV01, Li06, LX08, PP05, Sen08a].

i.i.d [Ast01, Aur08a, Cai02, LQR09, PS07b, Roz09b, Sun01, Sun09a]. identical [MW09, Mi06]. identically [Kru08, NC09b, XL06, YT08]. Identifiability [LTS01, PZ08, HH02c, STKF08, vW05]. Identification [GP00c, MNN09, Ros07, CHL08, LX01, SLC09, TSGG02]. Identifying [Poo06, AH06a]. Identities [Ruk09, CO00, GG02, Pri01]. identity [BG09, Hob07, Kat09, Lac08, NS02, Ruk03, dlPZ02]. if [dE00a, dE00b]. IFC [Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano06p, Ano06q]. IFR [BGHP04, Cui02]. ignorable [CS05b]. II [DG06a, PRB99, Pal01b, AAB04, BCK01, BM03a, Fer00, LF07, Raq03, Zhe01]. iid [LS04a, Eis08, Muk09]. III [Has09]. ill [DSBA09]. ill-posed [DSBA09]. illness [PD09]. illness-death [PD09]. image [KK00b, NH07]. images [Mas02a]. immediately [YZ01]. immigration [GMM00, HL01, LZ06, Ma09, Pre09, RAS08a]. impact [TV07]. imperfect [BS00, BC07, CK01]. Implementing [ZvdH03]. implications [KNB00, Lop02]. implies [MB08]. imply [BS06d]. importance [Bha08, dCB00b, Che05, NH07]. important [DDM06]. imprecise [GLV00]. improve [Har04]. Improved [GO01a, HH03a, Lee04b, Mad08, MSW01, SZ08b, VL03, Wan02a, WJ05, YW09, CAM03, CF06]. improvement [AT08c, KGC09]. Improvements [HP07, Zha01b, KZ08a]. Improving [AAHJ04, MBY00, PS09a]. impulsive [SL09a]. imputation [AS06, GL03, MSC08, Rei08, TMH02]. imputations [Sub09]. imputed [AS06, Moj01b, Rei08, SS05a]. inactive [ZB09b]. Inagaki [RB08b]. INAR [Wei08]. Inclusion [JS07]. income [CY02a]. Incomplete [Ger08, OMJ^{+}04, OK06, AP02a, CP08, Haz00, MGSB08, Mon07, NG02, Udd02, XL09]. Inconsistency [Cho05a, Nag09, NBK08]. incorporating [ODO07]. increased [Coo04]. Increasing [Lou05, AY02, Che00b, FMS00, GJ02, JF00, Joh02, Par01]. increments [AM02a, Bal02, Din03, GW05, Kab09, LLW09, MiSW00, Wan01, Wan02c]. incubation [NPC05]. indefinite [SXLX07]. Independence [BM08, Ano06o, Bar06, BS06d, CS08a, CA06a, CA06b, FF08, Hür04, IB09, JS02, Kou06, KW02, LX08, MB08, NOP00, Sch08, SZ02a, TKO03, WP00]. independent [AM02a, Bal02, BW07, Den05, DG09, Hou05, Kan07, Khr01, Kos09a, Kra02, Kru08, Lim05, MiSW00, MK08a, Mat03, Mav07, NZ00, NZ01, NC09b, RW05, Sha06a, Sim01, Sor09, Spo07, SS05b, Sun08a, TB09, XL06, Yos09]. independently [YT08]. indeterminate [OS05]. Index [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano01c, Ano01e, Ano01f, Ano01d, Ano03b, Ano03c, Ano03d, Ano03e, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano01b, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano03a, BV02, BM04, DG06a, FP08, GP00b, GO03, GN08, HLM06, IKP08, JG04, Lou01, Pen01b, Son08, Wri00, YPD09]. indexed [Bal01, Bal02, Bor00, Cas07, CM03, HCSV00, IS03, Kra00, MMY09, Yan03b, Yos09]. indexing [Ale01]. indicators [SB08]. indices [Li09b]. indirect [AGG02, Van08]. Indistinguishability [LN03]. individual [VT00]. individuals [HYL04]. induced [MS08a, PD09, SSA09]. inductive [Zan08]. Inequalities [CK05b, Gor09, MMV01, SZ05, Spo07, dlCC05b, Ahm02, AZS08, BB08b, Bea09b, BJ00, BB05, CGNW01, Chr00, CH09a, Clé01, CP05c, CM04b, CMS05, FJ04, GLT08, Har08, HX02, JPL09, Kra02, Mav07, ML09a, MN01b, Mus00, Nan08, Nik08, Pri01, Ren08, Ren09b, RB06, Sha06a, Shi04c, TS09a, Wan00a, Wan04, Wan09b, WZ09b, Yan04, YT08, Yu08a, Zan08, Zha09a, dlPZ02]. inequality [AY02, AR03a, BR08, Bha05b, Bos03b, Bos04b, Bou09, CP00, EG01, Gel04, GII08, GO02, Goe09, HWYZ09a, HWLZ09, HWYZ09b, IS01a, II08, IKP08, Jan03, LS02b, Li03b, LGLDG00, Mav08, Mer05, NS06a, Oli05, Ose09a, Rao02a, Rao02b, Rao02c, RL01, Rou01a, Rou01b, SM09, Sim08, Sim09, Sun07c, Sun08b, WZ08, Xie08, Xie09, XYC09, YZ04a, dlCC05a]. inertia [RM05]. infection [Hue03]. Inference [FJ07, GM08, SZ08a, ZL09, AV01a, Aic09, BPvZ04, CCC02, CC08c, CS00c, DOKN07, Fer00, GTH06, HW01e, HP09, KW07a, Kra04, MLT00, OC04, PAJ05, QZ07, QBZ09, Rek09, SD01, Wu04, YK05]. Inferences [MN03, JS09]. inferential [Lop02]. inferiority [JL09b]. infill [LTY08]. infimum [DK05]. Infinite [BK08c, KN08, BR06c, dCB00a, DMRA07, GJZ06, HH00a, HP09, JTQ09, Kru08, Mal01, Moo08, OE08b, Pen01b, Ros07, Sgi09, Sim01, WZ09a, YZ08, Yin08, ZL09]. infinite-dimensional [OE08b]. infinitely [Bar00, EsO08, Jur07, KK06a, KM09b, MPA07]. Influence [HCW07, KPK02, PMVL09, TWW00, CW01, DH09, Ger02, HK09, SS01b, SW04, TG02]. influential [JJ09, Poo06]. Information [Kle03, LLC06, Nin05, AA07, AES05, BS06b, BBZ08, Bra02b, Deb09, FK08, Hen00, HBA05, HN08, HH02c, KG09, KKM^{+}08, KMS08, NH07, PK07a, PBZ08, PB09, SC08, SBH03, WH05, WC04, XL09, Zhe01, ZG01]. informative [ST09]. inhomogeneous [CW08b, DaY08, Iac01]. initial [HYL04, Kyp09]. Inner [Tud08]. innovation [Sen07, Sen09]. innovations [AT08a, BYB08, DL06, DDN08, HS03a, LS08a, Mim08, WLG01, WB05]. inquiry [FL02]. insight [Lia02b]. inspected [SS01a]. inspection [FWW01, HW01c]. instability [Ker08, KG06]. instrumental [Hol08]. Insurance [ZBZ09, CS06, Has04]. insurer [Zho09]. integer [DvdAW08, FR06, HS03a, Wan08b, ZB08]. integer-valued [DvdAW08, FR06, HS03a, ZB08]. integrability [AL06, CV01, GMZ07, HCV01, MNZ05, Yao01]. integrable [YAT09]. integral [BH08, DMMY00, Gap08, GR01a, Jur07, MM07, NQMRLÚF01, TLRvdV08, YYL08]. integrals [Ano06o, BTW01, BdB09, Bre06, CA06a, CA06b, DO07, DLL00, HY00, Kub08, Küc04a, MMV01, SZ05, SZ09b, WM06, Yan04, YL05]. integrands [Tud08]. Integrated [BM01a, Bir08, DS01b, Mac08, Mak09, SKJ07, Ten01, Ten03, ZL07b, vE01]. integration [BB07a, Lia09b, MRP04]. integro [Bez08]. integro-differential [Bez08]. intensities [Fie03]. intensity [BZ06, ETC00, HS08b, NBA07]. intensity-marked [HS08b]. interaction [LWW09]. interactions [AZ04a, DDM06, Ma03, YB07]. interarrival [MY06]. intercept [Nia09]. interest [CL07, bGhWhZhW07, HWZ09a, KZ08c, LWD07, WH08, YZ01, YZL08, YH09, YZG08, ZWWC07]. interference [BK04, FM04, FRSWT08, KKAK08]. Interior [RSY02, Tot04]. intermediate [NO00, Seg01]. intermittent [MW05]. interpolation [GP05a]. interpretation [Ath09, Hör03, LL03, Man05b, NH07]. interpreted [PS07a]. Interval [LGLDG00, SW07b, BH04, CWL08, CYJL01b, DZB04, ETC00, FS01, GD04, HY01, IMS06, KY06, LB02, Sch03, ZCL08]. interval-censored [FS01]. Interval-valued [LGLDG00]. intervals [AB04, AB05, AB08a, ARB09, AM09, APR03, BGM03, BM09, CQ03, Din00b, DG06b, FK08, GL08, HS00, HH02c, IF08, Kab08, KG09, KLL08, Lud09, MP08a, MZG08, Ome08, RB08a, Ruk07a, Sha03, SW01, Tar05, Tri04, VT00, Wan02a, WN09b, WN09c]. Intrinsic [Kim00b, IKAT09, Vit08]. intrinsically [Sas08]. intuitive [NT03]. Invariance [Cha02, LC00, DR08, JS02, KKC08, Küh01, Lan08, MNOP03, NOP00, NT06b, RAS08b, Sha08b]. Invariant [CD00, dSF06a, BLV04, CW07, Cho05b, CD01b, HW07, JNS02, Mad02, Neg01]. inventory [LB03]. Inverse [BGN08, FM06, TD08, dlPGM08, Bou09, CL06a, CH04b, DB09, FF09, GP01, HV06, Kar02, KKY09, LSSG08, NC07, OS05, sS09c, Szk02, TT03, TL04, WSM09, WW08, Yu09c, ZJ09b, HK09]. inverse-probability-weighted [sS09c]. inversion [VA07]. inverted [AL08, Wit02]. invertible [BW06b, GL00, SW07a, ZM06]. Inverting [Are03, LH08b]. investigation [Piz09]. investment [KN05, Zho09]. involving [Li02, NS06a, PMJ08, ST04b, YKB07, Zan08]. IPCW [HK09]. irregular [PP06]. ISE [FJA09]. isomorphic [EKK06, Had02]. Isoperimetric [HP08, Nan08, Zan08]. Isoperimetric-type [Nan08]. isotonic [Hu06b, Hu09]. isotropic [BM07, JM08]. isotropy [PS00]. Itô [Bre06, Kur08, LT00a]. items [EG09]. Iterated [LQR09, YL05, Cai02, Den03, Den07, FU08, Hu08, Hua03b, Jar08, Kon02, KS00, Li09b, MY08, MS03b, Nan08, Nan09, PQ03, QF02, RMLD07, Sto03, Su05, WZ03, XL09, Yan04, YZ05, Zah00, ZY08]. iterations [GT08]. iterative [GAW08, Oht01, Sta00]. IV [YB07].

Jackknife [SS05a]. jackknifed [Özk08]. Jackknifing [DW00, YSkPkL02]. Jacobi [Dev09, Sza01]. James [AVV09]. Jeffreys [Nat01]. Jensen [Mer05]. Jirina [Li09c]. Johansen [DS01b]. Johnson [Ano07u, Cha06a]. joint [AB07, BK02, Bre06, Dud08, GQR06, HS07a, SK06b]. jointly [SH09b]. journals [Jan08]. judge [Din00b]. jump [AAHJ04, CL07, GJ01, LX01, Man08, Nak04, Shm06, dSF06a, dSF06b, dSF08]. jump-diffusion [Man08, Nak04]. jump-type [dSF06a, dSF06b, dSF08]. Jumps [Wei09, CL06c, Gap08, Gir09, Kur08, PL09b, Sit02, WIL01, WH07, XZJ08, Zha07a, Zha09b]. juxtaposition [EKK06, Sue03].

Kalman [Piz09]. Kantorovich [BBR06]. Kaplan [JSV02, LOS03, NT03, Sub09]. kappa [PK07a]. Karcher [KS08a]. Karhunen [Deh07, Ist06]. Karlin [Ste01]. Katz [LS04a, Sto08, Sto09]. Kendall [KNB00, NQMRLÚF03, WO08]. Kernel [BF00, CD01a, KKP00, MZ05, Pel05, AF01, Ahm02, BBR05, Blo07, BGN08, Che08a, DGZ02, DH06, EP02, FJA09, FI06, GS02, GJ08, HW01a, Hua01, JB05, KSW06a, KSW06b, KZ08a, KE05, KP02, LLOS09, LOS02, LC04, Mim08, Miz00, Moj00, NK03, Osm01, OS06, QF03, Rou00, SW06, SWM09, Shi02, Ten01, Ten03, Wal02, WW01, WCC07, Zha01b, vE01, vEG08]. Kernel-based [BF00, DGZ02, Moj00]. kernel-type [Shi02]. kernels [Cat01, HM09, Wan08c]. Kesten [Zha01a]. Khinchin [Kra02]. Khmaladze [BS04, KS05]. killing [vDZ05]. Kimeldorf [GV02a]. kind [LL06]. Kingman [Cer07]. known [Bru09, PS02]. Koch [Led08a]. Kolmogorov [BHHM04, BKM04, Dai04a, Dur03, Mav07, Mav08]. Kolmogorov-type [Dur03]. Kotz [Ars05, BK08c, DZ04, Has09]. Kotz-type [Ars05, DZ04]. Koziol [BV08, KD04, ZR06]. Kozubowski [BK08c]. Kuels [Oeh01, Oeh00]. Kulikova [Goe09]. Kullback [DOKN07, Haf06, Yoh08]. kurtosis [Haa09, KMvE00, Mar07, PP00].

l [GP00b]. Lack [BM06, Son09, AR09c]. Lack-of-fit [Son09]. Lack-of-fit-efficiently [BM06]. LAD [SW04]. ladder [Ale08]. lag [BL08, Wan02c]. Laha [Yan00b]. Lai [Hür03]. Lambda [UWF06]. lamplighter [GT08]. Lange [ET05]. Langevin [SBF00]. Laplace [Bar00, BK08c, HH02a, JK07, JTS08, KvD05, KN08, LCZ09, NC09a, NFL09, Roz09a, Sha07, TJ09]. Laplacian [Pyc07]. Large [Arc02, BJ00, Ber01, Ber08, Bis08, CV09, CC01b, CZ07, Chr00, Dar04, Eng04b, Esa04, Gut06, HO08, HU06a, JKK05, KG09, Kru08, Kuc05, LL05, MSW07, Mac08, Man08, Ole01, Osm01, Shi02, SM06a, SD09, TSJZ01, Ter03, Ter08, Tóm05, zW06, Wor01, YL00, YSY08, YJ01, dUÁ03, BS06a, BLS08, Bez06, BD05b, BM06, CDMLDK01, Com07, DdHL03, FQ03, FT08, GLS08, HCW08a, HC09, Ilt00, Ing00, Jou07, KN01a, Li03b, LS01, Lin08a, Lin08b, Liu07c, Liu09, Mac09, ML09a, ML09b, Nag03, QH09, Sch08, SW00a, SL08, Shu09, WHSL08, Wit05, YAT09]. Large-sample [KG09]. larger [dE00a, dE00b]. largest [CS07]. Lasso [Mei08]. last [HZ08, LL04, MOM07]. Latent [FD05, Bro09, Lyh05, vW05]. Lattice [Duk07, KTM08, Mar06, MK08b, Tot04, WP00]. Laugesen [PN09]. Law [Chr00, Gut06, Kru08, Kuc05, QF02, SM06a, Ter03, zW06, YL00, YSY08, YJ01, Ano02f, Cai02, Che00a, EL03, EGM06, FU08, GK09, HCSV00, HNRV00, HCW08a, Hu08, Hua03b, Kon02, Li09b, MY08, MS03b, NO00, Nov07, Pui08, QH09, Shu09, Sto03, Su05, Sun06a, Sun09a, WHSL08, YY08b, YZ05, ZY08, Ber01, HU06a, JKK05, LL05, Ole01, Ter08, Tóm05]. Laws [Bez06, CV09, GLS08, Nan09, XL09, YAT09, AL00, BC00a, BDG07, Bre06, Bri00, CP01, CDMLDK01, CH01, DS09, Den03, Den07, FMS00, Fro05, JY02, KN08, LS04a, MC07a, ML09a, ML09b, PQ03, SS06a, Sim09, SSA09, Sun01, SHV05, WZ03]. layer [YZL08, YZ09]. lazy [Hil02]. Least [HBK06, Lon09, SN04, ZB05, AT08b, BB07a, BE08, Blo07, GR01b, Gua07, Gue07, HLM06, HKLB07, JKP01, SKJ07, Shu04, TW06, Vou08a, YSkPkL02]. Least-squares [SN04, ZB05, GR01b, Gua07, JKP01, SKJ07]. Lee [ZNS08, ZRQ09]. Leffler [KM09b]. left [BP08, HcW08b, KL08a, Pon06, SB08, Wan08a]. left-truncated [HcW08b, Pon06, Wan08a]. Leibler [DOKN07, Yoh08]. Lemma [HWLZ09, Xie08, Bea09b, Cha08, KO04, Pet02, Lan06, Pet04]. Lenglart [Zha09a]. length [BLV04, CBR00, FSC03, IMS06, JT08, Lee00b, PR08, sS09b, Vag03, MS09a]. length-biased [BLV04, sS09b]. lengths [FSX02, KBH07]. less [IKP08, dHC06]. Lett [Ano06o, BvdV09, DR15, Duc10, dE00b, FT01a, GV08, HV00, HWYZ09a, KM09a, Kim13, Lin00, Oeh01, Pal01b, PS09b, PK08, Rou01a, SF12, Yu09b, vEZ02]. Letters [Bro11, Gat09, Hal09, Joh07]. level [AHL08, Ast01, BCAC01, Baí03, But04, But08, CL08a, DS00b, DC01, EKM02, EK04, FP07, GKM03, GSKP09, KK06b, KS04, KM05, LLZ04, Naz09, RM08, ST03b, TS01, YB07, YPD09]. level-crossing [FP07]. level-crossings [ST03b, TS01]. levels [EKS05, GPC09]. Levy [DR00, BN08, Bra05, El 08, FL08, Jur03, KK06a, Küc04a, Lon09, Otm09, PP09, RH07, Shm06, SV06, Zho09]. Lévy-driven [BN08]. Life [LK07, AAK00, AY02, Ahm04, AE00, Bar01, BS06c, BM03b, Bla08, JT08, LLJ00, NBA06, QZ07, SSB09, Sin07]. lifelengths [BA08]. lifetime [Che00b, DL04, Fin06, HH06, KY00, Mi06, PD09, YY07a]. lifetimes [GST00]. lightly [BD01a]. Likelihood [CF01, GTH06, Li02, Lo05a, RL08, ZB09a, Ago02, Ago06, Ann07, Ars09, BK08a, Bar07, BH03, BB03, BC03, Bis01, CL06a, CP05b, CC01b, CQ03, CC08c, CW07, CS05b, CFUOV00, CB01, Cor04b, CFH02, CF06, Dar03, Drt08, DG06b, EZA04, FS01, FCN02, FC08, FJ01, Fuj07, Fuj08, Gal09, Gao01, GS07, Gom00, HS07a, HWZ09b, HZ09c, JYZ08, Kar02, KZ08b, KN00, Kyp09, LCPY04, LNS01, LZ01, LZ02b, LZZ08, Llo05, LZ05, MK08a, MZ02, MN01a, MZ04, PR09a, Paw00, QZ07, QLL09, QBZ09, SH06, So03, Ste06, SD01, Szk02, Tia09, Tsa01, Ver07, WJ05, WS04, WX00, Wu04, XTW06, XL06, YZ07a, YZ06, YW04a, Zha06, ZR06, ZLQW07, dL05a]. Likelihood-based [Li02, CQ03, EZA04]. Likelihood-look-ahead [GTH06]. likelihoods [CM06, CM08a]. LIL [CB07, ZL07b]. Limit [BDG07, BHH08, Bla07, BMR00, BdB09, CL08b, CC02, CPD00, DS06, DK05, Dud03, Dud08, Faz03, FU08, GA00, GAW08, GMM00, HS03b, HPR02, JJQ08, JY08, JH02, JH08, JR07, Khr01, KB01, KK08a, Kul06, KMM00, Lau05, LW08a, Li09c, Mal01, Mat03, MS05, Neu04, NC09b, PMM09, Qi03, RW00, RS03b, SW06, Son00, Sou01, WL08a, Web06, vZ00, AZB07, AKK^{+}02, BR02a, Bir08, Blo02, Cha09, CP06b, CM03, CP01, CG02c, DR02, DR15, Efr03, Fah00, FH08, GS03a, GW05, Gon08, Gut06, HY00, HP05, HKM04, Hör06, KL00a, KL06, LZ06, Lin08b, Liu00, LY03, LaYY03, Liu03, Ma09, MY04, Mat05, NZ01, PMJ08, Sha04, Shi04c, Shu09, Sta02, Wan00a, Wan01, Woz09, WJ08, Yan03b, YY08a, Yu09c, ZCL08]. limitation [Rei08].

Limitations [MW05, FJS03]. limited [BM09, Wan02b]. Limiting [BM02, CHV09, GMdP04, Kab09, Mi06, Pin02, PS07b, RAS08a, SL06, Zha08, PMJ05, Zel06]. limits [CL06b, Eng04a, GV06, GK09, GT08, KL00b, Kab01, KL03a, Kos09b, Kra00]. line [And05, Huo05, SS08b, Zha01b]. Linear [HM09, IPP04, Abu02, AV01a, Als01, AP02b, APV06, AGJ04, AG08, BN00, BC00a, BYB08, Bos03b, Bos04b, CS05a, cICP09, CLR06, CC08c, CYJL01a, CYJL01b, CN03, CS05b, CB01, Cor04a, Cor04b, CHN04, CPG07, DG08, Det00, DP03, DaY08, Efr00, ET07, FS06, FMS04, Fer04, FvE01, GSGMFB03, GP00a, GC03, Geo07, Gør02, GP00c, HKS07, HWL08, HWZ09b, HZ09b, HZ09c, Hug07, JB09, JS07, Jar03, Jia01b, JFZ05, JL09a, JL09b, JR07, KB01, KKH08, KK08a, KZ08b, KKC08, Kle03, KS04, KC08, Kul06, KSG08, KGC09, LTC04, Lia02b, LH01, LC05, Liu07b, Lop08, MW09, Maj05, MZ05, MS08a, MK08b, MP01b, MS07a, MM09, Moo08, Mor05, NS09, Nki05, Özk08, PLGM06, PMVL09, PP00, PS02, PTA03, QLL09, RM05, Riv04, RH02, RN00, SB06, SW06]. linear [SW07a, Ser02, ST03a, She09, TW06, UI00, WP08, WN09b, Wit02, WC04, XTW06, XL09, Yeo05, YSkPkL02, YZ05, YC05, YZ06, YC06b, Yu07b, ZR05, ZWY05, ZF06, ZK06, ZX09, ZY04]. linearity [BA03, HB01, PP01]. linearly [dCB00a, Cas07, KB01]. LINEX [WZL00, Hua02]. Link [Kow00, FF08, JFZ05, KvD05, Kou06, Mac09, Roz09a, dFF02]. linkage [LA07]. Linnik [LJ06]. Lipschitz [Dub03, Lin09, San05, Wan08c, WH09]. Lipschitzian [Sit02]. LLN [GS09]. load [JL07]. Local [COS03, GSGMFB03, IS03, LCPY04, MPT09, SG06, Woe03, WX00, AKK^{+}02, Bar05, Blo07, Blo02, CS05a, Cha02, CW01, CA00, CO00, ETC00, GW09, GL04, Gør02, HC02, KLP01, Lin08b, LSL00, Mei09, MGW02, NS09, NT06b, ODO07, PP01, Sap08, Seg01, SV06, Sto03, Wan01, WZ03, WL08b, Wit05, YY07b, Zha01c, vZ00]. Localized [FT08]. locally [Pal01a, TN09]. locating [GM03a]. location [Che00c, CT04, Fan02a, FS02, Ger02, KH01, ML03, Per09, Sor09, Voc06, YS00]. location-independent [Sor09]. locations [Hu06b]. loci [ZCL08]. Loève [Deh07, Ist06]. Log [AL08, CXH09, Naz09, CS05b, FJ04, Han08, KG06, KSG08, Li09a, MR08b, So03, Waa04, YZ07a]. log-concave [MR08b]. Log-concavity [CXH09]. log-gamma [AL08]. Log-level [Naz09]. log-likelihood [So03]. log-linear [CS05b, KSG08]. log-negative [AL08]. Log-normal [AL08, Han08, Li09a]. log-rank [KG06, YZ07a]. logarithm [Cai02, Den03, Den07, FU08, Hu08, Hua03b, Kon02, Li09b, LQR09, MY08, MS03b, Nan09, PQ03, QF02, Shu09, Sto03, Su05, Sun09a, WZ03, XL09, YZ05, ZY08]. logarithmic [Cha02, Gan00]. Logconcavity [Cra04]. logistic [AV01a, Buz09, CAM03, CFH02, Dai00, GS07, Had03, Poo06, QF02, WH01]. logit [Bro09, Rek09]. logit-stable [Bro09]. loglikelihood [Tsa04]. lognormal [ARN08, BM09, WS04]. logrank [GE07]. Logspline [KK00c]. Long [LV03, BHGRV08, Cho05b, FMS04, HW05, KLL08, KH08b, LSEB00, MV01, PR09a, Pen01b, RCdCLN09, Wan09a, YV06, ZZ06]. long-memory [LSEB00, MV01]. long-range [HW05, Pen01b, YV06]. long-range-dependent [FMS04]. long-term [RCdCLN09, ZZ06]. longest [CFCL03, Mus00, Vag03]. longitudinal [QBZ09, QSP01, SS04, ST09, WH01, YZ07b]. look [CQ08, Che05, GTH06, Wan01]. Lorentz [JPL09]. Lorenz [DS01a, DZ02, KKY09, Koc06]. loss [BC00b, DV01, GMM02, Hua02, JR08, JNS02, JMP06, LLC06, Lin04, ODO07, Ste06, SZ08b, WZL00, YW09, Yu07b]. losses [Fro03, dlHRB05]. Lotka [WZ09a]. Lower [BCLBMR00, Mir05, BH05, CK05b, Dar04, Jon00, KRV05, LG03, MK08c, SC05, vE00]. lower-bounded [vE00]. loyal [BB04]. LRT [ZCL08]. ls [YOH00, PP06]. LSE [LSEB00]. LTD [Col08]. Lukacs [Yan00b]. lumpability [JB03]. Lyapunov [DS01a, JD08a]. Lynden [SKS09].

M [CHN04]. M-estimation [CHN04]. MA [GP00b, SXLX07, ZM06]. machines [FWW01]. MacQueen [SL09b]. made [Muk09]. magnitudes [He08, Zha07b]. main [AZ04a, BB08a, BB07c, Jac08]. maintained [BS00, CK05a]. majorant [KL06]. majorants [QS05]. majority [Sch07]. Malliavin [Ewa08]. Mallows [AD01, BK02]. Mandelbrot [Szu02]. manifolds [Pel05]. many [DS00b, KN01a, Leh05]. map [Hai03]. mapping [ZCL08]. mappings [Mas02a, Zap08]. maps [Dai00]. Marcinkiewicz [BC00a, DS09, GS09, Kra02, RL01]. margin [Lin04]. margin-based [Lin04]. Marginal [Haz00, Sek08, Alb08, Bri00, CH01, KH07, MRP04, Pel00, PS09a, SDR01, YC06a]. marginals [Dub03, Jon02b, JTS08, KM09b, LJ06, Mar01c, Mül01, PPSW04, PS02, PS09a]. marked [Fer09, HS08b]. markers [cICP09]. market [FW08b, INA07, LB03, Mon07, Zho09]. markets [Gap04, Irl04, Mil07]. Markov [Alt08, Ath08, Bal01, BSS01, BN01, Bou09, CG08a, Cek02, CP06a, CC08b, Cha06b, Che00a, CLR06, CLR08, CQX08, CL08c, Cho08b, Clé01, CmHP02, CW08b, CD00, CD01b, DBS07, DS01b, Dav04, EG01, ES09, EFL02, FP07, FSX02, FSC03, GTH06, GO02, Haa08, IPP04, Ilt00, JY08, JB03, JF00, KH07, Kos00, LLL00, Led08b, Lef08a, Leó01, Lim06, LW03, Liu07a, Lou01, Mam02, MY08, Mir01, MSW01, Nic08b, NBW07, OL02, PG09, Pal00, gPB02, Pas06, PPSW04, Pon06, Poz08, Psa02, RAS08b, Sen08b, SK06b, Son08, Spr01, Spr03, Sta05, SP00, Vag03, YL00, Yan03b, YZL09, Yao01]. Markov-additive [Ilt00]. Markov-dependent [SK06b]. Markov-switching [Dav04, Haa08, Liu07a]. Markovian [BHY09, FG04, Gap04, KKAK08, MNN09, Pet08]. Marshall [BR08]. Martin [Hob07]. Martingale [MY09, MV01, SKS09, Alb08, Ara01, Ara05, CHD08, CC08c, Guo08, HCSV00, HCW08a, JPL09, LS02b, Lep03, Li03b, LBSM00, Nie04, Ose09a, SS07c, Yu08a, Zho09]. martingale-differences [Nie04]. martingales [AR03a, Cha02, CP00, GLX07, IM08, LS00c, MN01b, Rao02b, Ren08, Ren09b, Wan09b, Yos09, dlPZ02, vZ00]. masked [KY00]. Masking [ZF06]. mass [JO08, Roz09b, RS08]. match [HH03b]. Matched [Lac04, Llo08]. Matching [CM08a, CKM03, CFCL03, CM06, MSC08]. mathematical [GLML09, PL09b]. mating [MdPR07]. matrices [AK07, CG02a, CC08a, EM09, FH07, GM08, GKM03, Gha00, Kle03, KMS08, MPA07, PMJ08, PLZ03, ST06, Zaj09]. matricial [Duc04, Duc10]. matrix [Asc09, BD05b, BG09, Bra02b, BP08, Cor04b, DaY08, Fal05, Fan02a, LH08b, LS08b, LZ05, Lyh05, MB08, Mon06, NHO03, ÖK07, PZL04b, Pap06, PK07a, Sch02, SvPtK08, SXLX07, SS05b, SZ08b, Tri08b, WZ08, WZ09b, YW09]. matrix-valued [DaY08]. matrix-variate [Tri08b]. Matsumoto [Hob07, Kou06]. Matusita [AFMGJG05]. Max [SS06a, CL06b, Fah00, Has05, Has06a, HS00, MR08b]. max-domain [Has05, MR08b]. max-limit [Fah00]. max-limits [CL06b]. Max-semi-selfdecomposable [SS06a]. max-stable [Has06a]. Maxima [BDM08, WT04, AKK^{+}02, Bar01, Dud08, HH01, Has02, HH05, Has06b, HS03b, Mat03, NZ00, Ose09b, Sco03, TB09]. Maximal [Chr00, HZ09a, MR08a, ML09a, Wan04, Wan09b, Yan04, Bha05b, FMS00, Fro03, HK03a, Har08, HX02, IMS06, Kle03, XYC09, Yu08b]. Maximized [So03]. maximizing [EK03b]. Maximum [AK07, Ars09, CG02a, FJ01, GZ06, HS07a, KR04, Leó01, MN01a, Mla09, MZ04, Per02, AHEA04, AZ04b, Ano02f, AK09, AVG08, BK08a, BH03, BB03, BC03, Bis01, Car04, CL06a, CP05b, CNX06, CG08b, CS05b, CFUOV00, CB01, Cor04b, CFH02, CG02c, Dar03, Del03, Eis08, FS01, Fuj07, Gao01, GS07, Has04, Kar02, KC03, Kyp09, LC05, Mar02, MMB07, PR09a, Per09, dR06, RC00, SLF09, Sta00, Szk02, Tia09, WS04, XTW06, XL06, ZR06]. maximum-likelihood [CFUOV00]. maxiset [Che08c]. Maxisets [Riv04]. McKay [FV08]. MCMC [AAHJ04]. Mean [DGZ02, SR04, AA01, AY02, AP04, AE00, BD01a, Bar07, BM03b, CM04a, Coo02, DR01a, Deg08, Deh07, Dro06, FK08, Fin06, GM08, HO08, HS07a, HCV05, HMR08, Hut02, KY09, KS08a, Kim00a, KvD05, Lil04, LZZ08, dL01, Mak08a, MP02, Mar01c, NBA06, NC09a, Oli01, ÖK07, Pen01a, PR08, QF03, QZ07, RM03, Sen09, SSB09, WY08b, Yan00a, ZL09, vE01]. mean-centered [Deh07]. mean-covariance [HS07a]. means [BCK01, Csö03, GMM02, HK03a, HCV05, Hu06b, Hu09, JGMGdlV06, KV02, Kim00b, Kru08, LTC04, Li09a, NW05, Ruk07a, UI00, vEZ01, vEZ02]. measurable [LFdVA05]. Measure [HR09, Ara01, Ara05, Dar04, DL04, DN09, Irl04, Kin08, Lan08, Lou00, Mon07, MS02b, SS07c, Zan03]. measure-theoretic [Irl04]. measurement [BdC00, Dav05, DG08, DS02a, Fuk05, GP05b, HH02c, Liu07b, SS07a, SZ08a, Wan02b, XY07]. measurements [DS02a, GP00a, GLV00, RL08, WH01]. measures [Arc00, AES05, BP06, BJ00, BT02, Bos07a, BR02b, CF00, CLM03, CD00, CD01b, Dar01, Duc04, Duc10, Esa04, FS00, Gri03, GM00, Guo08, GG00, HH02a, HHC09, Hel09, HS07a, Jam03, Jia09b, KL03c, KK06a, Kow00, Mas02a, OC04, Pak08b, Pro04, Rad07, RS03a, Sha01, VL03, Zaj09, Zap08, dSF06a]. Measuring [BCEO01, MF05a, ET07, FWW01]. Median [She08, AJ05, Alz06, BIKM09, BCG01, CDRV01, DGZ02, Lia01, MS09b, NHO03, Ozt07, SB00, SM09, Ste05, Wan07b]. median-resolution [Lia01]. medians [GM06, Hut02, Mer05]. medical [ZZ09]. medium [HS03b]. Meier [JSV02, LOS03, NT03, Sub09]. Memory [Wan09a, AR09c, BHGRV08, INA07, KLL08, LV03, LSEB00, MV01, PR09a, WD08]. Merwe [KY04]. Message [MS09a]. Method [Poz08, AH06a, AB07, AS06, BCC02, BV07, BM03c, BBG05, CL08c, CAM03, CDMLDK01, CM04b, Del03, DG03, Gal01, HK09, Hör03, Kou03, KSG08, Lep03, LX05, LT08, PPRW08, Röl08, SSY02, Sha03, TL04, Tor06, Tsa01, Tsa04, Udd02, Wu08b, YSY08, Zan08]. methodology [Ago02, KS05]. methods [AE08, AD01, BCD09, Che08c, FG06, GKM03, GJP01, HH03a, HH00b, Lia02a, MR00, QSP01, SD06, WH01, YC04]. metric [San05]. metrically [Kim00a]. metrics [Arc00]. Metropolis [FM00, SBF00, Yao00]. middle [MNZ08]. middle-censored [MNZ08]. minification [Ris06]. minihypers [AM05]. minima [Bar01, DG06a, HBA05]. Minimal [BKK09, CMNP07, Coe08, CBR00, DKM02, GBQ03, KL03b, Mon07, RM08, SS07c]. minimand [PL07]. Minimax [But05, Dro06, Huo05, Jon00, MP06, PM08, PS09b, PS09c, SV02, TLRvdV08, WZL00, vE00, CWL08, Che07b, Eni06, JNS02, MP02, OW03, Pro04, ST03a, Sha08a, SZ08b, Yu07b]. Minimization [Nak04, KKM^{+}08]. Minimum [BBP09, But04, GO01b, KH08a, Luc00, SV00, ZQ06, AZ04b, BBR06, BBR07, BW05, Bis06, Bon03, Hil00, HZ09a, HF04b, Lee00b, Li00, LCLZ06, Mar02, QC04, SLF09, TSGG02, YZL07, MS09a]. minority [Sch07]. minus [KMvE00]. mis [Coo06]. mis-specified [Coo06]. misclassification [Che04b]. Mises [WJ05]. mismeasured [CHL08]. Missing [BR04, Pro08, Tri08b, CLR02, GP00a, GL03, KPI09, MP08b, Moj01b, QLL09, Rei08, SXL05, SB08, WW01, WP00, Wu08b]. misspecification [Jow06]. misspecified [DFK03, Hat06]. Mitigating [SS07a]. Mittag [KM09b]. Mixed [Jia01a, PW08, AFMGJG05, Alt08, CL08a, DS00b, EN03, FM03b, GN08, GSKP09, KM05, Lau06, LLZ04, Lia01, Man05a, MN03, PZL04b, ST03a, SC08, Sue03, SD01, Wu08b]. Mixed-effects [Jia01a, Wu08b]. mixed-level [CL08a, GSKP09, KM05, LLZ04]. mixing [BL09b, Bra02a, CG08b, CHV09, Gan00, JH08, JD08c, KV02, Kon08, LN04, LS05a, Mal01, Mas08, Moo08, Shi04b, WJ08, XYC09]. mixingale [ML09a]. Mixture [GvLM04, Sha06b, AHEA04, ASS02, Ars05, Bos09, DP00, GST00, HSX05, Hor04, KSW06a, KSW06b, LL08, Lo05a, NZY09, PZ08, Ste05, Wan05b, WZ06a, WX00]. Mixtures [Bar02b, BLS05, AA08, BS01a, BC03, CS05a, CC01b, CMS05, EFL02, FY00, FJ04, Fot04, FV09, FS08c, Gri08, Kim08, Kim13, Lim05, MSH03, Pal01a, PM06, PM08, gPB02, STKF08, Vin08, YS00]. ML [RC03]. MLE [DGJ09, XL09]. MLEs [BM03a]. MM [SB06]. MM-estimators [SB06]. MMLEs [TS09b]. modal [Gut00]. Modality [BLS05, FI06]. mode [GS02, HZ04a, MK08c, RS00, SWM09]. Model [Jow06, PBH09, SQMK06, Yue02, ZP02, ASG04, Ago02, AT08c, BGM03, BB03, BRRZ03, BN01, BM05, BM03d, BV08, BK08c, CC08a, CK01, CFM01, CK05a, CP05b, CWL08, CL09, CL07, Ciu09, CFH02, DS00a, DG08, DMP01b, Di 00, DG07, Dup06, Efr03, EHMM08, FJA09, FWW01, FS06, FRG06, FF09, FM03b, FM04, FRSWT08, Fin03a, Fit06, For00, FM05, GP07, GW09, GR06, Gap08, GO01b, GQR06, Gri02, HM06, Han03, HWZ09a, HN02, HWL08, HZ09b, HYL04, Hyn07, IPP04, IM05, JS07, JK07, Jia01b, JL07, KD04, KF08, KC08, Kra04, Kyp09, LTC04, LB07, LR08a, Led08a, Lee03, Leo09, LZC05, LWD07, LK07, LWT08, LYZ08, LL09a, Lia02a, LC05, LSL00, LW08b, LV05, LJ01, LXZS09, LX01, Ma03, Mam02, MZ05, Mar01c, Mar02, MY04, Mei08]. model [MZG08, MGW02, MS02a, Mon07, MGB05, Nak04, Nam03, NO06, NBA06, NC09a, Nia09, Nin05, OM06, Omo07, Orm00, Özk08, PW08, Pan04, Pap06, PL09a, PZ08, PS06b, Per02, PS09b, PS09c, PD09, Pre09, QF02, Rek09, Ren09a, Sar00, SLC09, SSR08, Sek08, Ser00, Shu04, So03, Son09, ST01b, Ste05, SL09b, SZ06, SZ00, SZ02b, SS05b, Sun06a, SZ08a, SD01, TSJZ01, TJ06, TW06, WZ09a, Wan05b, Wan07b, WP08, Wan08a, WC03, WIL01, WC04, WX00, XL09, XZJ08, XWS05, YZL08, YZ09, YZL09, YH09, YPD09, Yu07a, Yu09b, YH08, YZG08, ZWY05, ZZG06, ZR06, ZBZ09, ZK06, dJP04, dL05a, dlHRB05]. model-based [PW08]. Model-robust [Yue02]. modeled [LP09]. Modeling [HC09, Nic08c, DMP01a, Haf06, KF09, KY00, Lau06, ZZ09]. Modelling [Han08, HS08b, PMZP07, ZJ09b, MRP04, UD09]. models [AHEA04, AH02, AV01a, ASS02, AT08a, AB08b, AG09a, Alt08, AT08b, ANV06, AP02b, AGJ04, Ara01, BN00, BE09, BL08, BL09a, BLV04, BPvZ04, BC05, BBZ08, BM09, BHK04, BB02b, BO02, BhRH04, Bib05, BHHM04, Bis08, BMR00, BdC00, Bos09, Cer07, CHD08, CC08b, CKM03, CD05, cC01c, CM07, CC08c, CHL08, CS08c, CYJL01a, CYJL01b, CÖ02, Cho08b, CH09b, CA00, CS05b, Con05, CFUOV00, CB01, Cor04a, Cor04b, CDRV01, CHN04, CF06, CPG07, CV08, DL06, DS01b, Dav05, DMB02, DS02a, DGJ09, DP03, DvdAW08, Duc04, Duc10, Esh04, ET07, EWI04, FR01, FR02, Fan02b, FGMS08, FC08, FE08, FG04, FS08b, Fuj08, FB02, GPC05, GP07, Gal01, GP06, bGhWhZhW07, GP00a, GL03, Geo07, GST00, GP05b, GAW08, GR07, GH05b, GV02b, HL02, HS03a]. models [HSX05, Hat06, Hil01, HS07a, Hol08, HZ03, HZ04b, HWZ09b, HZ09c, Hug07, HH02c, INA07, Jia01a, JFZ05, JL09a, KSW06a, KSW06b, KA07, KK00b, KPK02, KZ08b, Kle03, KMR00, KTZ03, KSG08, KGC09, LLL00, Lau06, Lee00a, LN02, LS05a, LS08a, Li00, LTS01, Li03a, Lia00b, Liu07b, Lyh05, MP01a, Mai08, MRK08, MS07a, Mor05, MSW01, Nag09, Nat01, NBK08, OS03, PW08, PP05, PLGM06, PMVL09, PS09a, Pro09b, PK07b, PK08, Psa00, QLL09, QSP01, RC03, RA08, RCdCLN09, RH02, RMAA01, SBF01, Sco03, SC08, Sha06b, SS00b, Shi04a, SKJ07, SS04, SX08, SB08, SD09, TWW00, TPA08, Wan02b, WZ06a, WO08, WS04, WH08, Wei08, WC04, WB05, WP00, WZW03, Wu08b, Wul08, XTW06, XNF07, XWL09, XY07, YZ08, Yeo05, Yeo07, YSkPkL02, YZ05, YC05, YZ06, YC06b, YZ07b, YWS09, ZM06, ZL07a]. models [ZP09, ZZ06, ZX09, ZB09c, ZB09d, ZY04, ZvdH03, vW05]. Moderate [sDXlyY05, Gao01, GJ09, XL06, Arc02, LV05, Wor01]. modes [ACSGV00, Le 03, Le 06]. modifications [MS09b, SM09]. Modified [CS08a, Coo03, JSV01, KY04, LK06, Wan07a, FCN02, JJ09, JSV02, JSV05, Nat01]. moduli [Jon00]. modulus [CC01a, Kab09]. Moment [AZS08, Clé01, GLT08, LL09b, Sha06a, Shi04c, Sou01, Sun07a, Wan00a, WP08, AY02, Ahm02, CM04b, DG06a, FL08, GN08, HRV08, HB04, KK08b, xLxZ04, Lin97, Lin00, LL06, LG09, Mna08a, Mna08b, MRS06, NZY09, Pen02a, Pen08, RM05, Sgi09, Sim09, Tar01, Umb06, Umb08, XYC09, Yoo09, YX06]. moment-based [NZY09, Yoo09]. moment-type [CM04b]. Moments [DWC04, GHL01, Gra09b, KM03a, KM09a, WY08a, AB08b, Ale08, ANV06, BCC02, BM09, Bla07, Bro08a, CS05a, CG08b, CTY02, CKR02, DMMY00, FNA09, Fot04, FN07, GP01, GM03b, II08, JZ04, KN01a, KO04, KK04, KKAK08, Lad07, Mar06, MMV01, Nad08, OM06, Raq03, Ruk09, Sim08, Sze08b, TSJ09, Wil05, WSM09, WW08, YC06a]. momentum [Coo03, Coo04, PS06a]. momentum-threshold [Coo04]. Monitoring [Bro08b]. monotone [AH06b, ABG07, BK02, KO04, KD04, Li07, Log03, LK02, Mas02b, Neu07, Pal08, PLLJ06, PCTA00, Por02a, WP00, ZCLH01]. Monotonicity [CW05, Llo05, dlCV01, CQX08, Dur03, Mei09, NP08b, SV06, Yu09c, ZB09b]. Monte [BBP09, BSS01, EZA04, JD08a, Kos00, Lia09b, Mir01]. Moore [LO05b]. Moran [IM05]. Morpurgo [PN09]. mosaic [MPV01]. motion [Abu02, Alb08, BNO03, BTW01, BH05, BGT04, CH07, Car04, Dun06, DF11, DLL00, DF02, EN03, Eng04b, ESOO09, GW05, GG02, GK04a, GK04b, HL01, HNS08, IKP08, LN09, LZ07, LLW09, Mak09, MNN09, McG06, MMV01, Nan08, Nie04, NS06b, SL00, SZ05, SZ09b, Tud08, YYL08, ZBZ09, Pou08]. motion-type [IKP08]. motions [BT02, CI01, DMMY00, Ist06, Mac09, Ole08, SP00]. Mourier [Zah00]. moving [AAV02, CHV09, sDXlyY05, Gal09, HS03a, HM06, KK08b, Kin08, KH08b, xLxZ04, NZ00, Ros07, Sun09b, WLG01, YX06]. moving-average [xLxZ04, YX06]. mrl [Fro06]. MSE [Oht01, WV08]. MST [LS02a]. MTAR [LS00a]. multi [Ciu09, DGZ02, Ery08, Fuj08, GW09, HC09, Joh02, KS06, KKC08, KS04, SK06b, Sia00, Sun00, Tak04, Vag03, YZL08, YZ09, ZF06]. multi-casting [HC09]. multi-class [ZF06]. multi-delayed [GW09]. multi-dimensional [Sun00, Tak04]. multi-layer [YZL08, YZ09]. multi-level [KS04]. multi-parameter [KKC08]. multi-phase [Ciu09, Fuj08]. multi-server [KS06]. multi-sets [Joh02]. multi-stage [DGZ02]. multi-state [Ery08, SK06b, Vag03]. multi-way [Sia00]. multidimensional [BJR00, BDG07, DR00, EM08, GK04a, Li09b, Stu00]. multidimensionally [CM03]. Multifold [Fai01]. Multifractal [FS00, AL08, BM01b, Kin08, Lud09, Mab08]. multifractional [LZ07]. multigroup [PI09]. multilevel [GR07]. multinomial [CKS06b, GV02b, Le 03, Le 06, Rek09, Ruk03]. multinomials [Lie01]. multinormality [LPY04]. multiparameter [CM08a, CA00, MS08b]. multiphase [MS03b, Min09]. Multiple [ACSGV00, CP08, Drt08, HH05, Kak08, KFW02, Ano06o, BD01a, BdB09, Bre06, CA06a, CA06b, Gor07, GS08b, Gor09, JL09a, JL09b, KL00c, Lia09a, LC05, MK08c, NW05, PLLJ06, Rei08, She09, Sub09, TMH02, TT07b]. multiplicative [Kra04, UD09]. multiresponse [Yue02]. Multisample [Miz00]. Multiscale [MPV01, LK04]. multistage [DS00a, Ten03]. Multistratum [AZ04b]. Multitype [GMM04, MN09]. multivalued [Xu08]. Multivariate [AP02a, Bur00, CL09, HW01c, LRW09, Mas02c, NP08a, PLGM06, Pyc03, SS07b, ST01b, AT08c, Ars05, Ars08, Ars09, BB09b, BT04, BRSL08, Bha05a, Blo02, BG09, BKL01, Bro07, Bro11, CG02a, CG08a, CC08a, CHD08, Dub02, ETC00, Fal05, Fra06, FD05, Fur08, GM08, GP00a, GR01a, Ger02, GVGP08, Gri08, HH01, HH05, Has06b, HS07b, Hu09, HC06, Hut02, Kak08, Kim00a, Kos00, KY04, Li03a, LGH03, LL03, Liu03, Log03, LJ01, MS07a, MB08, Mar07, MSW09, Nov03, PR09a, PL09a, PS07a, RS06b, Rao02a, RLÚF03, Rom09, RL08, Ser00, Sha07, SF12, TR02, Umb08, Voc06, WH02, Wit00, WCC07, YK03, Yeh02, ZS08, ZN05, dFF02, vZ00, RC00]. mutagrams [MW06]. mutation [dSF08]. MV [Udd08]. MV-optimal [Udd08]. My [Joh07].

N [Hal09, Cui02]. N. [Ros09b]. Nadaraya [Cai01]. Nardzewski [HNRV00]. Narvarte [Che00c]. natural [BLP03, Kok01, KK06a, LBSM00, Mas06b, NZY09, Pom01]. Navier [BM03c]. NBU [Li04]. NBUC [LLJ00]. ND [KZ08c]. near [BB07a, HH01, HS03b, QL06, YJ01, Per02]. near-epoch [QL06]. near-integration [BB07a]. near-maxima [HS03b]. nearest [GBQ03, Lal08]. nearly [BPvZ04, FQ03, HP09, Sal02]. Necessary [CD01b, LWW09, CLF06, Jia09b, MSW07]. needle [JK08]. Negative [LCJ06, AA08, AL08, DMMY00, EG09, Fur07, HWZ09a, HW07, HY04b, JZ04, Le 06, Mac08, Mór09, Pen08, PS06b, Roz09b, Ruk09, ST04a, ST01b, Tan07, VU07, WW08, Yu09a]. negatively [AZB07, hCW09, CZ07, CH09a, GRT07, HX02, Hua03b, HZ06, KKH08, KRV05, Kuc09, Li09b, Lia00a, Lin08c, Liu07c, Sha08b, SL08, WT04, WCS07, WL08a, Yan03a, ZW01]. neighbor [FM03b, KM08, Lal08]. neighbors [GBQ03]. Nel [KY04]. Nelson [DS01b]. Nested [Raj09, BB08a]. nets [DR01b]. Network [dlPGM08, Din00c, Hua09]. networks [DHS00, LV07, STKF08]. neural [LV07, STKF08]. neutral [BHY09]. Neveu [KO01]. Newman [Oeh00, Oeh01]. Newton [Ver07]. Neyman [Ing00]. no [Pal01a, ZWWC07]. noise [DZ02, GvLM04, KF08, MSC08, MN01a, MN04, RC03, SS07c, Wan05a]. noises [Lon09]. Non [AR03b, EKK06, PW00, Alb08, AP02a, And05, BW08, BC01, CLR06, CS05b, CD01b, DS01b, DR08, ET05, EKS05, FS08a, FS06, Fuj07, GC03, Gra09b, Hor04, Hou05, HCW07, IM08, JL09b, JYZ08, KP02, Led08b, LZD09, LWW09, Lia02a, Lin09, dL01, MP01a, MN09, Mak08b, Mar07, Mi06, MPS06, NBK08, PS06b, PS00, Pro03, Ren09a, Roz09b, Sch00, Shi09, Sit02, TP09, Tri04, WD08, Wan08c, WH09, ZWWC07]. non-autonomous [WD08]. non-Bayesian [Lia02a]. non-Brownian [Alb08]. non-central [Gra09b, Hor04]. non-centrality [LZD09]. non-classical [DR08]. non-degenerate [BW08, JYZ08]. non-dichotomous [MPS06]. non-discounted [Ren09a]. non-dominated [MP01a]. Non-exponential [AR03b]. non-extinction [MN09]. non-Gaussianity [HCW07]. non-homogeneous [Led08b, dL01, NBK08]. non-identical [Mi06]. non-ignorable [CS05b]. non-independent [Hou05]. non-inferiority [JL09b]. non-interaction [LWW09]. Non-isomorphic [EKK06]. non-linear [CLR06, FS06, GC03]. non-Lipschitz [Lin09, Wan08c, WH09]. non-Lipschitzian [Sit02]. non-Markov [DS01b]. non-negative [PS06b, Roz09b]. non-normal [ZWWC07]. non-normality [ET05, Mar07]. non-optimal [Pro03]. non-orthogonal [EKS05]. non-parametric [AP02a, And05, KP02, Mak08b, Sch00, Tri04]. non-recurrent [Shi09]. non-regular [BC01, Fuj07, IM08]. non-response [CS05b]. non-singular [CD01b]. non-stationary [PS00, TP09]. non-symmetric [FS08a]. Non-uniform [PW00]. Nonanticipative [Lep03]. Noncanonical [OE08b]. noncentral [SvPtK08]. Nonclassical [BK04]. Noncommutative [Sto09]. nonhomogeneity [CGK06]. nonhomogeneous [LLL00, LRM00]. nonidentical [Xia08]. nonidentically [San05]. nonignorably [Wu08b]. noninformative [SL00]. Nonlinear [HB01, Als01, BK01a, CHD08, CS08c, Ciu09, CFUOV00, CF06, CV08, FC08, GPC05, GP00a, HO09, Hil03, HH02c, KKH02, Lee00a, LS05a, Li00, LJ01, Pro09b, Ren01, RSY02, SL09a, Shu04, SFS03, TWW00, TP03, WN09c, Wu08b, Xi04, XWL09, YH09]. nonlinearity [Wan07b]. Nonnegative [YOH00, Bha05b, GP01, IS01a, KRV05, Roz09a, zW06, Wan08b, WSM09, Xie09]. Nonnegative-definite [YOH00]. nonnormality [GL00]. nonoptimal [HP07]. nonorthogonal [Han05]. Nonparametric [AB08a, Bre08, Cai03, CO06, CCS06, DL06, Efr08, Gør02, Hut03, LSL00, NG02, dR06, RM08, SK06a, VT00, ALOS08, BCAC01, BD01b, COS03, Cer07, CL06a, Che02b, CF01, Con05, DV08, Deg08, Det00, Erm00, FS01, FGMW09, GP06, GR01b, GMdRV02, HZ04a, HZ09c, Hug04, Jan03, KW07a, KE05, Kul01, LN04, LZ02a, Mak08c, MSC08, MP08a, MN01a, Mor05, MSW09, Nai01, OC04, PF07, RS00, RS06b, Son00, TM09, Voc06, Wan07b, YZ07b, Yue01, YW04b, ZD04]. nonproportional [GE07]. Nonregular [YB07, BCLBMR00]. nonstandard [Arc02, Jia01b]. Nonstationary [Ma03, dBC06, CL06b, Cho05a, HP09]. Nontransitivity [GE07]. nontrivial [AL06]. nonuniform [Mal01]. nonvanishing [BGK07]. NOP [Lou00]. norm [Fan02a, Gai07, KK00b, LLW09, Mim08, Ose09b, Ren09b, YZ04b, Yu08a, Raq03]. Normal [BS01b, PZL04b, Ruk07b, Ser00, Wil05, AL08, ACSGV00, BM03a, BGM03, BM09, Bis01, BLS05, Cac01, Cai00, hCW09, CC01b, Dro06, FK08, FBVW00, FNA09, FJ04, Fot04, FM06, GM08, GO01b, GV02a, GHL01, Han08, HC06, Hut03, IF08, JBB08, JB09, JTS08, Kar02, Kim08, Kim13, LSEB00, Li09a, LB03, LL03, LL08, Lo05a, Lop01, LRW09, MP02, NK03, Nov07, PL09a, PR08, RR01, RA08, Ruk09, Seg01, SV02, SS05b, SW07b, SZ08b, Sza01, TJ09, Wan02a, WBG04, WM06, WC04, WX00, Wul08, XWL09, YGK08, Yan00b, ZWWC07, vEZ01, vEZ02]. normal-agreeing [GV02a]. normal-inverse [Kar02]. normal-Laplace [JTS08, TJ09]. normal-symmetric [HC06]. normalising [KW07b]. normality [AW06, DG03, ET05, GP00a, Gir04, HO09, HZ06, Hut02, KSW06a, LLOS09, LYZ08, Mar07, Mir07, Rou00, Ten03, Woe03, XTW06, Yan03a]. Normalizations [ED02]. normalized [AHEA04, Cer08, Den03, Den07, JH08, JR07, Kos09a, Kul06, MC07a, Wan02c]. normally [SvPtK08]. normed [CG08b]. normings [Bis01]. norms [HZ03, HZ04b, SZ05, Zah00]. North [EM09]. Note [DvdAW08, Ann07, Ano03u, AS02, Aur08a, BD05a, BC00b, BS06a, BD01a, BS04, BL09a, BHZ00, BLP03, BLV04, BE08, BTW01, BW05, BO02, Bib05, BD05b, Blo02, BI02, Bor00, BKL01, BRR01, BZ06, Bru09, BPC09, Buz09, Cer07, CC08b, Che06, CM07, CHL08, CWZ08, CI09, CmHP02, Dai03, DG08, DK08, DGNR08, DG03, DK02, Dud03, ETC00, Ewa08, FQ03, FLQ03, Fuj07, GP05a, Gan00, GLX07, GL08, GJ08, GPC09, GMdRV02, Guo08, HS08a, HM06, Han03, HC04, Hil02, HW01a, HSV98, HV00, HH00b, HZ05, HF04b, Hug01, HYL04, Hug04, IM08, Ili00, Jac08, Jia06, JD08b, JB05, Jur00b, KM03b, Kim08, Kim13, Kol00b, Koz05a, Koz05b, Kyp09, Lac08, Lan08, Lec08, LCLZ06, Li07, Lin04, LT00b, LO05b, Lop08, LBMD05, LQ04, MS08a, Maz00]. note [Mei08, MRS06, NW05, NRS05, Neu07, Pak08b, Pal00, PLZ03, Pet02, Pro08, Psa09b, Pui08, Rod00, Röl08, RN00, Sco03, SH09a, SWM09, SW01, Ste01, STKF08, SLF06, Sun00, Sun07c, Sun08a, Sun08b, Sun09b, Szk02, TSJ09, Tar01, TN09, Vol06, WW01, WBG04, Wan06, WL08a, Wan08a, Wen07, YY08a, YZ08, YV06, Yu07a, Yu09b, ZR06, Zha07b, ZCCH09, ZRQ09, dSF06b]. Notes [Ago06, Shi09]. notion [BGHP04, NOP01]. notions [BI02]. novel [Has06a]. Novikov [Stu00]. NPMLE [MNZ08]. nuclear [Mas06a]. Nugget [LTY08]. nuisance [KZ08b, LY05]. null [FvE01, GVS00, GVGP08, Sen08a]. null-distribution [FvE01]. Number [HA03, BB04, dCB00a, CP05a, CL08a, DHS00, Eis09, FQ03, GLS08, HH01, HS03b, HYL04, Küh01, Kyp09, LRM00, Lo05a, Mah04, MdPR07, OMM09, Por02a, Pou08, PDM01, Rei08, Sch08, ST03b]. numbers [Bez06, CDMLDK01, GST00, GLS08, HCW08a, Jen04, ML09a, ML09b, QH09, Shu09, WHSL08, YAT09, Ber01, CV09, Chr00, Gut06, HU06a, JKK05, Kru08, Kuc05, LL05, Ole01, SM06a, Ter03, Ter08, Tóm05, zW06, YL00, YSY08, YJ01]. Nummelin [Clé01]. NWU [Ryc02].

O [ZJ09a]. objects [Por02a, Por03]. Observation [ZB09d, GT04, Lee04a, ST09, Tri08b, ZB08]. Observation-driven [ZB09d, ZB08]. observations [ADJ07, AF01, BHH08, BCG01, CQ03, DS09, Duc09, FM04, GS03b, HK03b, JJ09, JTQ09, MNN09, MP08b, Muk09, Pon06, Poo06]. observed [Bis06, CP06a, Din03, LS04b, Lon09]. obstacles [ZZ08]. obtain [AB07]. obtained [EKK06, OW03, PLZ03, PZL04a, Yat02]. obtaining [BL07]. occupancy [Mir07, Nak08, Zel06]. occupation [BGT04, DS00a, DS01b, GST00, Hoo02, Lab08, Tal08, YY08b]. occurrence [Cha05, OL02, Poz08]. occurrences [HA03]. odd [II08, Wit02]. odds [CM07, Kim03, KD04]. offspring [GMM04]. often [QS05]. Oja [NHO03, She08]. OLS [Vou06, YPD09]. OLS-based [YPD09]. OLSE [SS00b]. omitted [Nam03, Oht02]. On-line [And05, SS08b]. once [TT08, Dai05]. once-reinforced [TT08]. one [BM04, BB02b, CW07, DW02, Fin03b, GLX07, Gut00, HS00, Hut03, Jia09a, KK00a, Kok01, LX05, LGH03, LWW09, Lin08b, Log03, NS06b, Pal01a, Por03, Wan02a, Wan07a, WY08a, WX08, XW09, Zha01c, vEZ01, vEZ02]. one-dimensional [DW02, GLX07, Jia09a, NS06b, WY08a, WX08, XW09, Zha01c]. one-parameter [Pal01a]. one-sample [KK00a]. one-sided [BM04, CW07, HS00, LGH03, LWW09, Lin08b, Log03, Wan02a, Wan07a]. one-step [BB02b]. one-way [Gut00, Hut03]. only [dE00a, dE00b]. Operator [ST06, Cho08a, Ewa08, LVL09, MB06, Sze08a, Yu08a]. operator-valued [Yu08a]. operators [Mas06a, Rom02, dSF08]. opposite [HY04a]. Optimal [ASS02, AM05, AF01, AHL08, BB08a, BBR05, Bia02, BD01b, BC07, BM03d, BB07c, CWL08, CL08a, GD04, II08, KS04, LG03, MS07a, Mei07, MT07a, PM01, SXLX07, TF05, Tom09, WG03, WZ05, Yoh08, Zho09, Ara01, Ara05, Ath09, BS08b, BM06, BZ06, BCK06, But08, CGK06, CP06b, CI09, DH06, DMB02, Fan02b, FLQ03, Fan03, FT04, GR06, GP00b, GPC09, Han03, HC04, Hil00, HW01a, HN08, HT05, HK08, Jac08, JL09a, JL09b, Jou07, KS08b, KN05, KM05, Lee04a, Lia04, LR08c, MW00, MN01a, Mül01, Nia09, Pal01a, Por03, Pro03, SS09a, Udd02, Udd08, VZ03, ZAC03]. Optimality [CCDG02, DvZ05, FM03b, FM04, GS08b, But01, DWZ05, FRSWT08, GR08, Mar01a, STM01, SS03, XNF07]. optimization [Che04a, Yao00]. Optimum [PM06, GM03a, HP07, PM08, Pro03, Pro04]. option [Gag02, Gap08, XWS05]. optional [FI07]. options [CL06c, RS07]. Order [AR09a, CLM03, Pak08a, ADJ06, ADJ07, AM09, AVV09, AB08b, AS01, BK02, BCK01, BCC02, BA03, BS08a, BB08b, BCD09, BLB08, Bar06, BEA09a, BLS04, BGM03, BPR01, BGHP04, BB03, BK07, BRVM03, BR06b, BR06c, Bro08a, CV01, CP05b, CXH09, CW01, CS00c, CKS06b, CB01, Cor04b, CKR02, Cra04, DP03, DS02b, EB03, FM03a, Hei01, HZ03, HZ04b, HZ05, HWL08, Hür04, Hut00, HBK06, JB09, Jon02c, KO01, KK00a, KK05, Kle02, KF08, Koc06, LN02, LQ07, LB03, LNS01, Log03, Lop08, Mab08, MLT00, MR09, MK08a, Nad08, NP08a, Nic08c, OR01, OAAA05, Ozt07, Pap01, PMVL09, PS06b, Raq03, RP02, Ris05, Ros09a, RMAA01, Ruk07b, Ryc08, Ryc09, Sas08, Seg01, Sha07, Sim06, Sor09, Sta02, SQMK06, SD01, VT08, WN09a]. order [Woz09, YW06, ZP02, ZB08, dlCV01, dlCC05b]. order-preserving [BB03]. Ordered [YSN01, Ban08, BS01a, CP08, Cas07, Cho08b, CKS04, CKS06b, Dub02, DC01, FT04, Ili00, IB09, IS03, KW07a, Li02, NW05, SGK09]. Ordering [AA08, BB07a, AF09, Fin06, Fro06, GV02a, KK00a, Kle02, Koc06, LT06, LH01, PR09b, WH02, Zha07a, ZB09a]. orderings [BRSL08, Col08, GW08, KK05, KS07, RS03a]. orders [Bar00, Bar02b, BS06c, BRR02, Cha01, DLS00, DZ04, LNS01, MR09]. ordinal [Bis04, KA07]. ordinary [BB07a, BZ06, QF03, SKJ07, TW06]. ordinate [Bar02a, BRF06]. origin [Fan02b, GL04, Roz09b]. Orlicz [Ren09b, Yu08a]. Orlicz-norm [Ren09b]. Ornstein [Bis01, Bis06, Bis08, BN08, Dit07, EM08, GJ09, KS08b, Liu08, Lon09]. Orthogonal [ASG04, Lia00b, PLLJ06, ASS02, CFM01, DDM06, DS00b, Dey05, EM09, EKS05, EKK06, FM04, Geo07, KK06b, KyS05, Lia01, OE08b, PLZ03, PZL04a, PZL04b, SSY02, Sue03]. Orthogonality [XNF07, Dar01]. orthogonally [CCDG02]. oscillations [Din00a, GK04b]. Ostrowski [dlCC05a]. Ostrowski-type [dlCC05a]. other [BJT03, Li02, ZWWC07]. otherwise [CKS04]. outcomes [Jan08, YKB07]. outlier [Pan04]. Outliers [Bat05, BD01a, HS08a]. outlyingness [DH09]. over-relaxation [BSS01]. overdispersed [CB01]. overdispersion [Orm00]. overshoot [Spo07, Tan07].

p [Lee00a, Wei08]. packets [LVL09]. packing [Xia09]. pair [Gua07, LWW09, MT07b, Yan00a]. paired [GS03b]. pairs [Kow00, Li02, Llo08]. Pairwise [dL05a, Kru08, KN00]. Pakes [LJ06]. panel [BL08, BL09a, JS09, PPRW08, SZ02b, YZ05]. Papakonstantinou [APS09a]. paper [BS04, GMZ07, HWLZ09]. paradox [Kev07]. parallel [Li04, LH08a, Lia00b, LV07]. parallel-flats [Lia00b]. Parameter [BBZ08, AM09, Aic09, Ars09, Aur08b, BP02, BES01, BJR00, BJT03, BHK04, Che00b, Fal05, Fan02a, FGMS08, FV09, Gas09, GQR06, HW07, HKS07, HW01b, HH06, JBB08, JGMGPM03, JR08, JNS02, KKC08, Kle03, KK06b, Kou03, KR09, LS08a, LSSG08, LS06a, LZD09, Lia09c, MW09, MP06, MN03, MZ04, MN01b, OS02, Pal01a, PK07a, Pui08, RA08, SXL05, SX08, VJ08, WZL00, Wan07a, Wan09a, Woe03, YSN01, ZPK08, vE00]. parameterization [PL09a]. parameterizations [But01]. parameters [AM07, AT08a, AA07, BK08a, BLS04, BM01a, FR01, FW08b, Fit06, FNA09, Ili00, IF08, JS07, JSV05, KZ08b, KS08b, KC08, LY05, LG03, LdC07, MW09, MS07b, NBK08, PK07b, PK08, WJ05, WB05, Yu07b, ZWY05]. Parametric [Gre09, GLV00, LY05, AP02a, And05, Hol08, HWL08, KP02, Lau06, Mak08b, PPSW04, Pen01b, PS09a, Pon06, Sch00, Tri04, ZvdH03]. Pareto [Bra02b, CL09, LR08c, dCD09]. part [AVV09, Han03, Nam03, NO06, SS08b]. Partial [Ign09, Yoo09, AP02b, APV06, BBZ08, BE08, BJ00, CC08c, CG08b, Deb07, DMB02, DK05, FD05, GH00, GL03, GDM06, Gon08, GK09, HCSV00, Kos09b, LT00b, LaYY03, Mav07, Mla09, MC07b, Poz04, QL06, WD08, Wan02c, WA01, WC04, Yi05, ZR06]. partial-sums [WA01]. partially [Din03, FT04, GM00, HWL08, HWZ09b, HZ09b, HZ09c, IS03, IW04, KSG08, Liu07b, MZ05, PLGM06, YSkPkL02, YZ05, YC05, YZ06, YC06b, ZX09, ZY04]. Particle [FF09, JD08b]. particular [HM06, KKH02]. partition [LS08c]. partitioning [Wal02]. partitions [ETC00, Esa04]. parts [HZ09c, Mac08]. Pascal [MSZ08]. passage [Als01, DY04, Dit07, Kok01, LC00, PP09, WY08a]. past [DL04]. pasting [CI09]. Path [BK01a, GK04a, LHLC04, Cha06b, DKM02, Sek08, Yos09]. path-dependent [Sek08]. pathological [KP00]. paths [dSF08]. pattern [Cha05, HA03, MR00]. patterns [AB07, CP05a, HS08b, Poz08, Wan05a].

Paul [Kev07]. payments [LL09a, YZG08]. peaks [FR01]. peaks-over-threshold [FR01]. Pearson [Ago06, Cor04a]. pebbled [CM01]. Penalization [LX05]. Penalized [ABG07, DG06b, Pal08, Bar07, GS07, Gue07, YW04a]. penalty [Hua01, LW08b, Ren09a, Sun05, ZZG06]. percentiles [HJ06]. percolation [LR08a, MPV01]. perfect [BS00, HH03b, SS01a]. performance [dBC06, Mak08c, NO06, VV00]. period [FH06, KG09]. Periodic [AG09b, AB08b, AG09a, Bez08, BB05, BhRH04, BL09b, SN04, Sha06b]. periodically [BLS04, CG02a, CG08a, SS01a]. Periodogram [BRF06, ED02]. periodogram-based [ED02]. Permutation [AH01, CW07, JL09b]. Permutations [Lev00, JF00, Joh02]. perpetual [CI09]. Perpetuities [BR01]. Persistence [Haa09]. Persistent [PBH09]. Persistent-threshold-GARCH [PBH09]. perspective [Dav05]. Perturbation [Rom02, ZBZ09]. perturbed [Gir09, LC00, LXZS09, YZ09]. Perturbing [PL07]. Petersburg [CS07]. Phase [TT08, Ciu09, Fuj08, Li03a, NH07]. phase-type [Li03a]. pHd [CY02b]. phenomenon [AHV06]. Pickands [DK08]. piece [BN01]. piece-wise [BN01]. piecewise [Abu02, CL08c, PMA08]. piecewise-linear [Abu02]. Pitman [AB09, BIKM09, IK04]. pivots [Cor08]. place [KS00]. place-dependent [KS00]. placebo [DWZ05]. placebo-treatment [DWZ05]. Plackett [MS04a]. planar [Gre09]. planned [FI07]. plans [AM05, AHL08, BB08a, BB07c, Jac08]. players [Eis09]. plot [AZ04b, OMJ^{+}04, OK06, YZL07]. plots [LPY04]. plotting [ET05]. PLS [CW01]. plug [Baí03]. plug-in [Baí03]. plus [Jan08]. PMSE [Nam03, NO06]. Podgórski [BK08c]. point [AH01, AS02, BR01, BHHM04, BM05, BRR01, CM04a, CR08, Che00c, CCC02, Che09, CM08b, DFK03, DE06, Din00b, Dup06, FJ01, FJ07, Fot09, GM03a, Gir04, Gom00, GVS00, GVGP08, GvLM04, Gri03, HW05, Has04, HS08b, KTM08, LZ02a, LWD07, MNOP03, Nin05, NOP00, PMM09, PK04, WZW03, ZL09, ZLQW07, Zuo01]. point-lattice [KTM08]. points [And05, AR02, BB09b, CT04, CFR06, Gup09, HP07, HH01, Küh01, LX01, Mat03, MM02, MS02b, NZ01, Pou08, Pro03, RC00, Tot04, Wan01, YS00, Yue01]. Pointwise [SW06, BB09a, But00, HK08, KZ08d, Lab08]. Poisson [Gat09, AR05, BK08b, BJR00, BSCN09, Bor00, BZ06, Bro08b, BW06b, Cek02, Cer07, CP01, DFK03, FM05, Gat08, Gir04, Gra09a, Gri02, HH02b, Jon08, JZ04, KT03, Led08b, Lef08b, LRM00, LT00a, MS08b, MGW02, MSH03, MT07a, MN08, NBK08, Nia09, Nov03, Roo01, SZ02a, SD01, SR04, Szk02, Vin08, VW08, WZL00, WS04, YZL08, YZ09, Yu09a, YH08, Zha09b, ZJ09b]. Poisson-binomial [AR05]. Poisson-inverse [ZJ09b]. Poisson-lognormal [WS04]. Pole [EM09]. policies [HW01c]. Polish [ETP08, Yan08]. Pólya [Gri08, KW07a, KMR00]. polygon [DZ01]. polyhedral [AW04]. polynomial [BM06, Bry00, CWL08, Fan02b, Fan03, GR01b, Hil00, Hua03a, HC02, KLP01, LSL00, Mah04, MPT09, SS00b, Sun07a]. polynomials [BC06, Bro08a, Duk07, FS08a, KW07a, Mar06, MK08b, Sen08b, Sun07a, Wil05, Wit00]. population [AAC00, AB09, BC00b, BIKM09, Bar07, CHL01, Din00b, GMM04, Hug07, IM05, KK01, LRM00, Lia09c, Mai08, Nic08a, RM03, SBF01, VJ08]. population-averaged [SBF01]. populations [ADR05, Bha05a, Bos04a, Cer02, CKS06b, DS02b, GM08, Kak08, KMV09, Li09a, LR07, NZ01, Ryc09]. portfolio [FW08b]. Portmanteau [BP06]. posed [DSBA09]. position [AK09]. positive [AVV09, Che00a, Col08, DMMY00, DFS08, GO03, IKK05, LX00a, MRK08, Mór09, Nam03, NO06, Roz09b, SvPtK08, Sze08b, Wit02]. positive-part [AVV09, Nam03, NO06]. positively [KB01, Mat05, Sha08b]. possession [FL02]. possible [BZ06, CCC02, CQ03, MR03]. possibly [FMS04, GMM02, HP09, Pap01]. Posterior [HW07, Pom01, BB09a, CKS06a, CM08a, Dah06, GVMSN04, RA08]. posteriors [Bas00, GL03, Waa04]. potentials [LL04]. Potts [LV07]. power [BB07b, BHHM04, CW03, Con05, Coo04, CBR00, DG07, ET07, FM06, GL08, GLC09, Haa08, HN02, IV00, KBH07, KvD05, Mak08b, Sen09, SZ02a, Tar05, TLRvdV08, YKB07, ZD04]. power-normal [FM06]. power-weighted [CBR00]. powerful [KW02, Llo08]. powers [MR08a, OS05]. practical [Kim03]. pre [Oht01, Oht02]. pre-test [Oht01, Oht02]. Precise [BLS08, LL06, Liu07c, Liu09, SL08, WZ03, YX06, ZY08, LS06b, Lin08a]. precision [SZ08b]. preconditioned [PR09a]. Predictable [IW04, Ose09b]. predicting [CKM03, YPD09]. Prediction [AM09, CYJL01b, Dun06, DF11, HP00, HH02c, JB09, Kli06, RB08a, SW07a, APR03, APS09b, AG08, BGM03, Bos07b, GP05a, GV06, KPI09, KM03b, LR07, MM09, NS02]. predictive [BR02a, Bos04a, Bos09, CCS06, Cor08, Dah06, ET07, MGB05, dlHRB05]. predictor [DGZ02, YPD09]. predictors [CM04a, JL09a, MPT09, RN00, WY08b]. Preface [KAS07]. premium [ZZG06]. preorder [DBS07]. prescribed [Fal05, Kon08]. presence [AHS07, Coo06, FS08b, GT04, Hen00, Hol08, JK08, KZ08b, NIK02, Psa02, SDR01, Sen09]. present [BD01a, dUÁ03]. Preservation [BS06c, Bla08, JT08, LNS01, BRR02]. preserved [BLB08]. preserving [BB03, BRVM03, Mar01a]. prespecified [CWL08]. price [XWS05]. prices [Gag02]. Pricing [CL07]. Principal [BB09b, Leh05, BX00, BF00, Bou06, Cho05b, Pen08, Sun00, Van08, WL08b, YS00]. principle [DR08, KKC08, Küh01, Lee00b, LS01, Lin08a, Man08, Naz09, NT06b, RAS08b, Sha08b]. principles [Cha02, sDXlyY05, LC00, MSW07]. prior [BB09a, FK08, GVS00, HW07, KG09, LZ02a, Nat01, ST03a]. priority [Bro11, HWYZ09a]. priors [Cer07, CKM03, GVS00, Har04, Kim00b, SL00]. Prob [Ano06o]. Probab [BvdV09, DR15, Duc10, dE00b, FT01a, GV08, HV00, HWYZ09a, KM09a, Kim13, Lin00, Oeh01, Pal01b, PS09b, PK08, Rou01a, SF12, Yu09b, vEZ02]. Probabilistic [CLR08, IMS06, AG09a, AVdPS02, BM03c, GM06, Hör03, Lee00a, LS00d, Lou05]. Probabilities [Sch07, Ber08, BH05, BCEO01, BL03, BK04, CK05b, Che04b, DS00a, DS01b, DLL00, GLX07, GVMSN04, Hil02, Hou05, Hue03, Hug01, KS00, Liu00, MR08a, Mut08, Naz09, OM06, Osm01, PP09, RR01, Shm06, TV07, WH08, YH09]. Probability [Bro11, CKM03, CM06, FN07, Hal09, HK09, Joh07, SLC09, Vol06, AGG02, AVV09, AD01, Arc00, Bea07, BKM04, BK04, CF00, CFCL03, CS06, Clé01, Con05, CD00, CD01b, DB09, Dar01, De 08, Dub08, FS06, Fer04, GW09, Gat08, Gat09, GR01a, Gho03, GVS00, Had02, HR08b, HN08, HCV05, Ign09, Jia09b, KL03c, KZ08c, KZ08d, Kum02, LS00c, LB02, LXZS09, Mak09, Mas02c, Mna08b, Nak08, NQMRLÚF01, OC04, Pak08b, PNJ08, RS00, RS06a, Rao02a, RC00, Rou00, Rou01a, Rou01b, SM09, sS09c, Wan00a, Wen03, hZbG08]. Problem [EM09, BV07, BR06a, BHHM04, Bon03, But00, Cha06a, CI09, CKS04, DSBA09, Dai04a, DW02, DWZ05, Drt08, ES09, GR06, GVGP08, HV06, HH03b, He08, IK04, JKZ09, Kra09, KY04, Lee04a, Lef08a, LZZ08, Log03, LG09, Mna08a, Mna08b, Nak08, Pal08, Por02a, Ren01, RSY02, Sza07, Szk02, TT07a, WX08, XW09, ZLQW07]. problems [AB07, BC01, CL06a, DFK03, bGhWhZhW07, KPI09, IV00, KK00a, Lin97, Lin00, LR07, Mei09, Mir07, Nag03, Por02b, RS00, Ruk07b, SVB07, SXL05, YZL09, dlPGM08]. procedure [GSTS07, GVS00, GBQ03, Gor07, GS08b, Gor09, GR08, HY01, JNR09, JR08, KL00c, KW02, LB07, Sta00]. procedures [CH04a, CS00c, Gor07, GS08b, Lia09a, LLC06, MSC08, NW05, Riv04, TSGG02]. process [AAB04, AS02, BMW06, BJR00, BB07b, Bis01, Bis06, BdB09, BN01, BM05, BZ06, Bro08b, BW06b, CP06a, CR08, CC01a, CL08c, CY02a, CmHP02, DFK03, Dar04, DZB04, Deg08, DD06, Din00a, Din03, Dit07, El 08, FW08a, Fer04, FGMW09, Fuj08, Gal09, GS08a, GJ09, Gas09, Gat08, Gat09, Gho03, Gir09, GMM00, Guo08, Haa08, Hai03, HSX05, HP05, Hue03, Iac01, JSV02, JR07, KTM08, KB01, KKH08, KK08a, KS08b, KMS08, Koz05a, Koz05b, KL06, Led08b, LS00a, Lef08a, LRM00, LZC05, LdC07, Lic09, LHLC04, LH08b, Liu06, Liu07a, Man08, MB06, Nag09, NBK08, Otm09, PMM09, PS06a, PS07a, Pet08, Pon06, Pyc03, RAS08a, RP02, Ris05, Ros07, ST03b, Shu09, SV06, SLF09, Sun05, Szu02, Tal08, TJ09, WM03, Wan00b, Wor01, Wri00, Yao01, Yi05, Zha01c, ZL07b]. process [ZY08, ZJ09a, ZL09]. processes [Ahm02, AAV02, AM02a, AM02b, AL08, Arc02, AW06, Aur08b, BH09, BK08b, Bal01, Bal02, BJT03, BLS04, BYB08, BL09b, BS09, Bor00, BN08, Bos03b, Bos04b, BR02b, Bra05, Bra02a, Bro08a, CKS06a, CG02a, CG08a, Cer08, CLR06, CLR08, CHV09, CN03, Cho05a, CL07, CM08b, Coo03, Cor08, CH01, DBS07, Dav04, DE06, Deb07, DGNR08, sDXlyY05, Dow09, DM05, Dub03, Dun06, DF11, EM08, EFL02, EsO08, FH08, FMS04, Fer09, FP07, Fie03, FL08, GA00, GLT08, Gir04, Gne03, GLML09, GMdP04, GMM04, GDF01, GvLM04, Gui01, GJ01, Haa09, Has04, HS08b, Hoo02, HKS00, HZ08, HB01, HW01e, HB04, HBK06, HKLB07, HP09, IKP08, IKAT09, IS03, JT08, JTS08, Kab09, KV02, KPI09, KKY09, KS06, KB01, KK08b, KKC08, KO08, Kon08, KH08a, KJ08, Küc04a]. processes [Kul06, LPS09, Le08, Led08b, LN04, Lee06, Lef08b, LJ06, LS06a, LS05b, LX00b, LL04, xLxZ04, Li09c, Lim06, LZ01, Liu08, Lon09, LK04, Ma09, Mac08, Mac09, MiSW00, Mas02c, MS08b, MMB07, MV01, MPY03, MY06, MMY09, Mla09, MM09, MdPR07, MT07b, Mon06, MC07b, Moo08, Nan09, Neg01, Neu04, NBW07, OL02, OE08a, PR09a, PB03, PP09, PMM09, gPB02, PBH09, Pat09, PTA03, PW06, RH07, RB06, Ris06, RB08b, ST06, Sas08, SS06a, SW06, SW07a, SS07c, Sen08b, SW00a, SY04, Shi09, Shm06, SS06c, SV00, Ste06, Stu00, Sun09b, Szk02, Tak04, TS01, TD08, VvdMD08, VW08, VA07, Waa04, WLG01, Wan02d, Wei09, Woe03, YY07a, YZ04a, YL05, YV06, YP09, YX06, Zap03, ZCCH09, ZB08, ZB05, dSF06a, dSF06b, dSF08, vDZ05]. Product [Ano06o, CA06a, CA06b, DG09, CP06b, DMP01b, GX08, Gon08, HY00, Man05a, Nad05, Ole01, PZL04a, Tud08, Wan00a]. product-limit [HY00, Wan00a]. product-sum [DMP01b]. Products [Cha04, AL08, CMNP07, Gha00, LW08a, LQ04, MC07a, Qi03, RW05, Sha06a, Sim01]. Profile [LZ02b, FCN02, IR08]. progenitor [MdPR07]. prognostic [MPS06]. programs [Sha08a]. progressive [AAB04, BCK01, BCC02, BCK06, Raq03]. progressively [BM03a, Cra04]. Prohorov [Arc00, DR00]. Prohorov-type [Arc00]. projected [PR08]. Projection [Dey05, AK07, EK03b, Hen00, HCW07, KH01, Ose09b, PLZ03, ZQ06]. Projections [Del03, DOKN07, MBY00, Var08]. projective [EK03a]. projectivity [EK04]. Prokhorov [Goe09, HCW08a]. Prolate [LR08b]. Proof [JL07, DW05, GS02, HY04a, Hob07, Jam03, JY08]. propagation [AFS^{+}07]. propensity [Bru09]. Proper [LT08]. Properties [BhRH04, CS00c, Jen01, Lai01, NBA06, NBA07, Ahm04, AG09a, Ara05, AVdPS02, Aur08b, BS08a, BM03b, BRVM03, BO02, But04, CW05, CC08b, CR08, CL09, CV08, DGZ02, Dey05, EK03b, FM05, FH06, FN07, GS07, Gel09, GM06, Had02, Hat06, HRV08, KL00a, Kab08, KTM08, KH01, KKH02, Lee00a, LP03, LHLC04, LZ07, Liu03, LL08, LK04, dL05b, MSZ08, MC07b, Nag03, NT03, Pro09b, SN09, Sia00, Stu00, TT07a, VT08, Wan01, Wan02d, WO08, WL08b, Yan03b, Yeh02, dlCV01]. property [Alz06, AR09c, BK02, Bal01, BB03, CF00, CM06, CWZ08, Cui02, GV02a, GLML09, Haa09, KY09, Kou06, LS05a, LCJ06, Lil04, LH01, Liu00, LBSM00, MS08b, Spr01, VA07]. Proportional [ZZ06, AH02, BRR02, CM07, Di 00, Kim03, Lau06, NBA06, SD06, Sch07, SZ06, Yu07a, Yu09b]. proportionality [TW06]. proportions [CCS06, KC03, Nic08a, SH09b]. proposal [PMZP07]. proposals [AAHJ04, SBF00]. Propriety [RA08, Nat01]. protocols [ÖW00]. provider [Lou01]. proving [CM04b]. proximity [CP05a]. proxy [NO06]. Pseudo [ST03a, JM08, LTC04]. pseudo-generalized [LTC04]. pseudo-isotropic [JM08]. Pseudo-minimax [ST03a]. Publisher [Ano03u, Ano07t]. pure [Ath08, GDF01]. pursuit [HCW07].

q [Lee00a]. QAL [PD09]. QME [Lai01]. QML [Pap06]. quadrant [KB01, LX00a, RR01]. quadrant-dependent [KB01]. Quadratic [HC06, LR07, Lop01, BGK07, Bia02, BW06b, CO06, CAM03, EP02, GP07, GA00, GHL01, GP00c, KM03a, KM09a, KS04, Kra02, Li02, LBSM00, NZY09, PMJ05, Pom01, Ruk09, Yu07b]. quadratics [PMJ08]. qualitative [Had02]. quality [FSX02, PD09, Yat02]. quantification [LGLDG00]. quantile [AB05, AB08a, AAB04, Che02a, DG03, Hen00, IP03, KKH02, LLOS09, NP08b, OH05, OS06, Ryc02, SS07a, SKPP03, Sim07, SX08, YZ08, YV06, YM01]. quantiles [AB04, AB09, Bar02a, BKL01, Coe08, GM06, Ioa04, JSV01, JSV02, Jur00a, KK01, Lic09, Lin08c, Mai08, OR01, PCTA00, SW01, Sun06b, ZS08]. quantisation [MT07a]. quantitative [ZCL08, Zuo01]. quantization [SS06b]. quantum [BK04]. Quasi [CH07, Hut02, TL04, Ann07, BH03, BS01b, KA07, LZ02b, Szk02, TP01, Zha01c]. quasi-associated [BS01b]. quasi-everywhere [Zha01c]. quasi-likelihood [Ann07, LZ02b]. quasi-maximum [BH03, Szk02]. Quasi-medians [Hut02]. Quasi-residuals [TL04]. quasi-score [TP01]. Quasi-sure [CH07]. quasi-symmetry [KA07]. quasilikelihood [Per02]. quasirandom [Yue01]. questions [EG09]. queueing [BBZ08, Ker08, Leo09]. queues [KS06, MS03b, Min09]. Quine [GS03a].

R. [Ano07u]. r.c.l.l [LX05]. Racetrack [BL03]. radial [GJ09]. radioactive [AFS^{+}07]. radius [AA01]. Radix [GR00]. Radon [LFdVA05]. Raikov [Gut06]. Rand [DWC04]. Random [BK07, Din00a, Jur07, Kon02, KS00, Lic09, NS02, Pad01, PRB99, Pal01b, Pel00, QS05, Sha04, SP00, TPA08, AL06, AAV02, AR09a, AG09b, AA08, AZB07, And00, AHV06, AK09, AVG08, ALP08, Asc09, AKK^{+}02, ARN08, Ast01, Aur08a, BC06, BE09, BM07, BL08, BHZ00, BH04, Bar01, BJ00, BW07, Bia02, Bis01, Bis04, BR02b, dCB00b, BS01b, BPC09, BP08, CV01, Cai02, CW05, hCW09, CLO09, ÇA01, CW08a, CWTB09, CC08a, CP05a, Cha01, CKM03, Che06, cC01c, Che07a, CZ07, Che08a, CG08b, CWZ08, Che07b, CM03, CH09a, CM01, Ciu09, CDMLDK01, CFR06, D’A06, DSBA09, Dai00, DR01a, Dai03, Dai04a, Dai04b, Dai05, DHS00, DZ02, DO07, DV08, DZB04, Deg08, DDP08, Den05, DV01, DG09, Dev02, Dev07, DR02, DR15, DK05, Duc09]. random [Dud03, Duk07, EM09, ET05, Eis08, ETP08, EL03, EGM06, Ete07, FS00, FS08a, FW08b, FM03a, FP08, FBVW00, FT08, FJ01, FS09, Fur07, Fur02, Gai07, GS02, GA00, GW09, GP01, GHL01, Gha00, GH05a, GLX07, GW08, Gir09, GAW08, GX08, GMdP04, GT08, Gre09, GMZ07, GRT07, Gup09, GS09, HH00a, Has01, HH02b, Has06c, HY00, Hel09, HO08, Hil02, HN02, HCSV00, HNRV00, HCV01, HY04b, HRV08, HMR08, HWYZ09a, HWYZ09b, HX02, Hua03b, HZ06, Hut03, HBK06, IS01a, II08, Jam03, JJQ08, Jia01a, JH02, JF00, Joh02, Jon02a, JKK05, JD08c, Jur00b, KW00, KO04, KPI09, KM06, KK04, KKAK08, Khr01, KM03a, KKH08, KK08a, KM09a, Kin08, KD03, KRV05, KO03, Kon09, Kra02, Kra09, Kru08, Kuc05, Kuc07, Kuc09, KGC09, Lac08]. random [Lad07, LS04a, LC00, LS02a, Lee03, LCPY04, Lee04a, LP03, LSEB00, LS06a, LCJ06, LCZ09, Li09b, LQR09, Lia00a, LH01, Lin08a, Liu00, LY03, LaYY03, LW03, Liu03, Liu07c, LL09b, Liu09, Lop01, LGLDG00, Lud09, Mac09, MPA07, Mah04, MNZ05, Maj05, Mal01, MK08a, MP01b, Mar01c, Mas06a, Mat05, Mav07, MF05b, MS05, MNX09, ML09b, MSZ08, Mid08, MSH03, MM09, MS04b, MC07b, Moo08, Muk09, Mül01, Nag09, Nas06, Nat01, Nia09, NC09b, Nik08, NG02, Nyr05, OMM09, Omo03, OAAA05, OC04, Ost01, OS06, Pal09, PMJ05, PMJ08, PL07, Pel02, Pen02a, PL09b, Per09, PS00, PS07a, PS07b, Por02a, Poz04, PDM01, QZ07, QH09, RM03, RMLD07, Rom09, RW00, Roz09a, Roz09b, Ruk07b, San05, Sch03, Sha07, Sha06a, Sha08b, SL08, Shi04b, Shi04c, Sim01]. random [SH09b, SLF09, Sta00, Sto03, SLF06, Sun01, SHV05, Sun07b, Sun07c, Sun08a, Sun08b, Sun09a, TSJ09, TT08, TP09, TSJZ01, Tan07, TKO03, Ter03, Ter08, TB09, Tóm05, TM09, Vol06, WH01, Wan04, WT04, zW06, WCS07, Wan08b, WHSL08, Wan09a, WZ09b, WL08b, Wen03, Wit02, WJ08, WSM09, Xi04, Xia08, Xia09, Xie09, YWWY08, YK00, YT08, YZ02, Yue01, Zan08, Zha01a, ZB09a, hZbG08, dFF02, dlPGM08]. random-effect-coefficients [WH01]. Randomization [AS07, JKZ09, OM06, ZR05]. randomized [Ger08, Ima08, Ozt07, SKPP03, Sim07, Tan09]. randomly [CNX06, DS00a, HCSV00, HL01, HCV01, Kra00, Lel08, MMY09, Moj01b, SS05a, SL08, XL09, ZG01, hZbG08]. range [AB04, AHS07, FMS04, HW05, HP05, KH08b, Pen01b, Wan09a, YV06]. ranges [Hut03]. Rank [cC01c, GH00, CP08, DR05, EB03, FvE01, Gao03, GQR06, HZ09a, Hus03, KG06, LLC06, NMO08, SvPtK08, VL03, YZ07a, ZR05, ZS08]. Rank-based [GH00, CP08, GQR06, LLC06, VL03]. ranked [ASAK00, BES01, MZ04, ÖW00, RM03, WZ05]. ranking [CH04a]. Rao [Pan04, CI01, Jan03]. Rasch [BB03]. Rate [Kim00a, Sha01, Ver07, AAV02, AAHJ04, BB09a, Bar06, BGHP04, BH03, BBG05, BLS05, BMR07, But00, Cha06b, Che00b, CHL01, DGJ09, Duc09, Eng04b, FR02, Fot09, FvE01, GW05, GJ02, GSTS07, GVMSN04, GRT07, GR08, Hür04, IF08, KRV05, KW02, LG03, Lia04, LL08, LLW09, Nav08, Oeh00, Oeh01, OL02, OS06, SWM09, SM06a, WHSL08, YPD09, ZZG06]. Rates [Bis06, Blo07, Csö03, Gui01, HW05, HZ04a, KV02, KSW06b, KK00b, Mas02a, Arc02, Ath08, BCAC01, Bro08b, EP02, FE08, bGhWhZhW07, Gap04, Gor09, HU06a, JSV01, Li09b, Llo05, Mac09, Mas02c, Mei07, PS09b, PS09c, Sap08, SW06, Tóm05, WH08]. Ratio [MN08, ADR05, AS06, BS08a, Bar07, CC01b, CW07, CF06, DS06, FC08, Fuj08, HJ06, KF09, Lad07, LNS01, Lo05a, LZ05, Mad08, MK08a, RL08, Ste06, Tsa01, Tsa04, Wan02a, WJ05, WX00, YZ04a, ZB09a, ZLQW07]. ratios [AR09b, Gom00, MZ02, Pui08, Su05]. Rayleigh [Fer00, MK08a]. rays [Hu06b]. RCA [Aue04]. Read [Bra02a]. real [sDXlyY05, Mah04]. Realistic [GH05b]. rebirth [GK04a]. recapture [Hug01, Hug07, ZvdH03]. Reconstruction [Mna08a, Mna08b, NH07]. record [AB04, AB08a, AB09, Ali03, AA07, BS06b, BB02a, DR04, Eis09, GO01a, HN02, HBA05, KS07, NZ01, OAAA05, SBH03, ZB09b]. records [AB05, ARB09, AM09, AF09, BS06a, DLB05, DS06, Dem08, GLS08, KR04, Kli06, LBMD05, Mad08, RB08a]. Recurrence [Gha00, Tak04, SH09a, Sze08b]. recurrent [Che06, Che00a, Shi09, Sze08b, Tan09]. recursion [Asc09, KSO05]. Recursive [Kuz05, Coo02, Son08, SLF06, WZ06a]. redescending [CT04].

Redistribution [Bet00]. reduced [JGMGdlV06, Nai01]. Reducing [PS00]. reduction [DP03, MP00, WY08b, Wen07, YC04, YC06a, Yoo09]. redundancy [LH08a, VZ03]. Refined [LS04a]. refinement [YK05]. refinements [Xia08]. Reflected [GL04, RH07, LX05, Tak04, ZJ09a, ZZ08]. Reflecting [Kel06, LWW09]. reflection [Pro09a]. regarding [Dai00, Dai04b]. regeneration [JY08]. regenerative [Bha08]. regime [BhRH04, FG04, SH09a]. region [Bar02a, JL09c, ZAC03, ZM06]. regions [FGMS08, HJ06]. Regression [Che07b, Lee00b, SZ02b, AV01a, ADR05, Ago02, Aic09, AP02b, APV06, ABG07, AG08, AT08c, ALOS08, BA03, BV07, BC05, BGM03, BE08, BHGRV08, BB02b, BD01b, BHHM04, BMR00, Blo07, BR06b, Buz09, Cai01, CFM01, CO06, CHL08, CW01, Che02b, CÖ02, CAM03, CFUOV00, CFH02, DV08, DMB02, Dup06, Dur03, Efr03, Efr08, Fan02b, Fan03, FS08b, FGMW09, Fuj08, GP06, GS07, GSGMFB03, GP00a, Geo07, GMdRV02, Gør02, GRT07, Had03, HO09, HKS07, HS07b, Hu06b, Hu09, Hua03a, HZ09b, HC02, Ioa04, IWWJ04, JJ09, Jar03, Jen01, JR09, JL09c, Jur00a, KKP00, KL08a, KLP01, KKH02, KPK02, KP00, KS05, Kra04, Kul01, LTC04, Lal08, Lee03, Lee04b, LSEB00, LWT08, LYZ08, Lia02a, LC04, LX01, MW09, Mai08, MS08a, MBY00, MPT09, MP08a, Mid08, MM02]. regression [MN08, MSW09, Nam03, NO06, NP08b, NC07, Nia09, Oht01, Oht02, OH03, OH07, OS03, Özk08, Pal08, PP05, PF07, PK04, PL09a, PCTA00, PMA08, QF02, QLL09, QSP01, RS06b, RC03, SB06, SK06a, Sch00, SN09, ST03a, She09, sS09b, SS00b, STM01, Shu04, SFS03, Son09, SX08, SW04, TL04, TM09, Wal02, WH01, WW01, Wan07b, WY08b, XNF07, XWL09, XY07, Yan03a, Yeo05, YC04, YC06a, YSkPkL02, YC05, YZ06, YM01, Yu07b, YWS09, YW04a, YW04b, ZD04, ZF06, ZL07a, ZY04, ZWWC07]. regression-scale [MM02]. regressions [BL07, Bry00, KLL08, LX00b, WV08]. regressor [ALOS08, CHL08]. regressors [LLOS09, Nam03, SKJ07]. regret [Dro06]. regular [AHL08, BC01, Clé01, Fuj07, Has06c, IM08, Jon02a, MP08b]. regularity [AM02b, CLR06, Rad07, Zha01c]. regularization [AD01, DSBA09]. regularly [CNX06]. reimputation [SS05a]. Reinforced [DM05, Dai03, Dai04b, Dai05, MMB07, TT08]. Reiss [Has05, Has06b]. rejection [Che05]. related [BBZ08, Dev09, Dun06, DF11, Geo07, GT08, HP00, Has04, HP08, HB01, KC08, Lef08a, PAJ05, SS06a, SS07b, SV06, Yeh02]. relating [DO07]. relation [BGT04, CQX08, DBS07, HY04a]. relations [CLM03, D’A06, Man05a, Van08]. relationship [GR01b, HHC09, NOP00, Sor09, Sza01]. relationships [CLO09, Det00, GVMSN04, Umb06, Umb08, Zuo01]. relative [AR05, Fin06, KD03, Li06, MN09, Sch08]. relatively [Hut02]. relaxation [BSS01]. release [BC07]. relevant [Nam03]. Reliability [KMV09, BS01a, dCB00a, CL08c, CKLX06, GG00, KTZ03, KY00, Led08a, MK08c, NBK08, SSB09, Wan05b]. reliable [JK08]. remainder [Mir05]. remaining [BA08]. Remark [Dar03, Vou06, RR01, YZ04b]. Remarks [Hal09, HH02b, Has02, Roz09a, YY08b, Pat09, Pel02, YY07b]. remediation [CS00a]. REML [RC03, Wul08]. Removing [Pro03, HP07, MR00]. Ren [Hal09]. renewal [AAK00, BI02, Bro08a, CP06a, DW05, GW09, HP05, KS06, MS08b, MY06, Psa09a, Sap08, Sgi09, Sze08b, TSJZ01, TD08]. renewals [OMM09]. renormalization [Pak08a]. Rènyi [HWYZ09a, HWYZ09b, KO01, Rao02c, Sun08b, ZN05]. Rènyi-type [HWYZ09a, HWYZ09b, Rao02c]. repair [BKK09, BS00, BC07, CK01, HW01c, Mi06, SS01a, SL06]. repairs [BS00]. Repeated [Gra09a, GP00a, HS07a, MGSB08, RL08, WG03, WM06, XY07]. repeated-measures [HS07a]. replacement [BS00]. replicate [DS02a, GP05b]. replication [CD05]. reported [Lop02]. Reporting [Jan08]. representation [Bar09, Coe08, Cri04, IP03, Lin08c, OE08b, RB08b, SD06, Sch07, Sun06b, Sun07a, Vit08, ZCL08]. representations [Jur07, MM07, MS09b, PTA03, SKS09, TN09]. Reproducibility [De 08]. reproductive [TWW00, XTW06]. Resampling [dL01, PL07, Yat02]. Rescaled [AHS07, Bor00]. reserve [CY02a]. reserve-dependent [CY02a]. Residual [AE00, AY02, AJ00, ADJ06, ADJ07, BA08, BM03b, BR06c, Fin06, KS07, NBA06, Pal08, QZ07, Sin07, YY07a, YAT09]. Residuals [CV08, Ago06, Bro07, Bro11, Cor04a, KY09, TL04, Vou08b]. Resistance [KO08]. Resistant [Gne03, RS07]. resolution [Lia01, YB07]. resolvable [KM02a, OMJ^{+}04, OK06]. Resolving [KG06]. respect [BdB09, EsO08, Khr01, KKRV01, Kub08, Küc04a, MSW07, MMV01, SZ05, SZ09b, YL05]. response [But01, CS05b, DP00, GM03a, HF04b, MPS06, PNJ08, YW07, Yue01]. responses [BT04, BR04, QLL09, Wan00b, Wan02d]. Restricted [Gro03, CS00c, DR04, FY00, HW01d, OS02, Ozt07, Piz09, Voc06, WZL00, Wan07a, ZAC03]. restriction [ZP02]. restrictions [BBP09, DP03, HRV08]. result [Dai00, Dit07, DR01b, GS08a, Nic08b]. resultant [PR08]. results [Abu02, AF09, BCD09, BLS08, CM04a, CN03, Cho08a, CP05c, Dai04b, DS01c, Di 00, Din03, DK05, DGJ09, DF02, FL02, FC03, Has01, Jia09a, LLJ00, LS06b, Lop02, MS02a, Mór09, PZL04a, Psa09a, QC04, SM09, Var08, XL08, dUÁ03]. retarded [Liu08]. retrial [Ker08]. retrieval [LV07]. returns [FJ04]. reversals [JR09]. reverse [LBSM00]. reversed [AR09c, Di 00]. reversibility [CQX08]. reversible [AAHJ04]. reversion [The07]. Révész [Fro05]. revisited [KMM00, Rou07]. Revisiting [BB08b]. reward [DBS07]. rho [SS07b]. Rice [Del03]. rich [EsO08]. Ridge [DP00, Gro03, JJ09, JR09, Özk08]. Riemannian [GO01b, KL03c, Pel05]. right [BPR01, DB09, DS00a, HcW08b, KL08a, Lau06, Pon06, SH06, sS09b, SB08, Zap03]. right-censored [DS00a, HcW08b, Pon06, sS09b]. risk [CY02a, Con05, DS02b, Efr08, Fro06, bGhWhZhW07, GW09, GO01b, Har04, HHC09, HWZ09a, KK00b, KD03, Lep03, LZC05, LWD07, LW08b, LXZS09, LR08c, MZG08, Nak04, PS09b, PS09c, Ren09a, Sun05, SZ08a, WH08, XZJ08, YZ09, YZL09, YH09, YH08, YZG08, ZZG06, ZBZ09]. riskier [Sor09]. risks [CS06, FE08, GMM02, Li06, TV07]. RNG [MT04]. Robust [AV01a, Ago02, BR06b, CDRV01, FRG06, GdL00, GV06, OH03, QBZ09, SFS03, ALOS08, Ber03, BM04, CCG09, HW01b, Hut02, KK06b, NC07, RS07, Tom09, Yoh08, Yue02, ZCCH09]. robustness [Bos04a, CW03, MBY00, ODO07, WS04]. ROC [HH03a, JL09c, Llo02]. role [GST00]. Roman [KBH07]. root [CK05a, Coo03, CV04, Coo04, ED02, Fuk05, HK03b, LX00b, Mak08c, Nag09, Psa02, Sen07, Sen08a, Sen09, Vou06, Vou08b, Wan06]. roots [BRF06, MR00, PS06a, Psa00, SO00]. Rosalsky [CV09]. Rosenblatt [BD05a, Che04a, LN02, Lee06]. Rosenthal [IS01a, II08]. roughness [Hua01]. roughness-penalty [Hua01]. roundoff [BRR01]. row [BB08a, Udd02]. row-column [BB08a, Udd02]. rowwise [Kuc07, Kuc09, Sun08a]. RTI [Col08]. Ruin [bGhWhZhW07, GW09, WH08, YZL09, CS06, CTY02, CY02a, Con05, ES09, Gat08, Gat09, HZ08, KZ08c, Lef08a, LXZS09, Mac09, Sun05, TV07, XZJ08, YZ01, YZL08, YH09, YH08]. rule [CS00a, HY01, IWWJ04, KBH07, Moj00, NO06, Pre00]. rules [Bis00, FSC03, HP00, HK08, KF08]. run [Bél00, But05, CFCL03, FSX02, FSC03, MS04a, Mus00, Vag03]. Runs [EF08, AK07, CL08a, Ery07, Ery08, FQ03, FMS00, SK06b]. Ryan [ET05]. Ryll [HNRV00].

Sackrowitz [CP05b]. Saddlepoint [Kol00a, XWS05, AE08, ADR05, Are03, DR01a, FR06, GJ02, Gat08, Gat09, HMR08, PPRW08]. Safe [KF09]. same [CKS04]. Sample [KZS08, LB02, Tan09, dSF08, AK09, BD01a, BIKM09, Cha06b, CC01b, CT04, Coe08, CV04, EB03, GJ08, GL00, Had03, Haf06, HCV05, HC09, JGMGdlV06, KG09, KN01a, KY09, KC03, KK00a, Kol00a, KS05, Lin08c, LZZ08, Mar07, Mat03, MP08b, Mla09, Neu04, NC09a, Omo07, Ozt07, PW08, PS07a, SS02, Sch08, SSB09, SKPP03, Sim07, Son00, Sun06b, Tsa01, VV00, YSN01, Zuo01, dUÁ03]. sample-path [Cha06b]. sampled [CR08, HS03a, Shi09]. sampler [MP00, Omo07]. samplers [JD08a]. samples [BM03a, Bis04, BW08, Bre08, FN07, HH05, LYZ08, NMO08, RM03, Ryc08, WZ05, WN09a, Yan03a, Yat02]. Sampling [LS08b, ASAK00, BLV04, BES01, Bél00, Bha08, Bos04a, cICP09, Che05, DR01a, DS01c, Din00b, FWW01, FI07, Fit06, FS08b, Gra09a, HP00, HMR08, IW04, KSH07, KZS08, Lac04, LK06, Li02, MZ04, ÖW00, ST09, TF05, Wat01, Zha01b, dUÁ03]. Sampson [GV02a]. Satisfactory [SSY02]. satisfies [AL06]. saturated [EKS05, QO09]. Saunders [GOPB09]. scalar [AR03b, BT04]. Scale [Ars05, BP02, CCG09, CW07, CMS05, FJ04, FS08c, JNS02, KM06, Kim08, Kim13, Mad02, Mad08, MN03, MM02, Miz00, Ste06, TG02, Woe03, vE00]. scale-invariant [JNS02]. scale-parameter [JNS02]. Scaled [CS00b, MZ06, TG02]. scales [Bro09]. Scaling [AL00, Eng04a, Lud09, ST06]. scan [CGNW01, CG02b, GZ06, SW00b]. scatter [Ger02]. scenarios [AL08]. scenery [FS09]. Scheffe [Sha03]. scheme [FT01a, FT01b, ST01a, TT08, WH07]. schemes [BCK06, FSX02, LWW09, Luc00, PB09]. Schur [But08]. Schur- [But08].

Schuster [Che00c]. Science [Ano01h]. score [Bru09, CÖ02, GZ06, LC05, Nic02, Sim07, TP01, WJ05]. score-type [GZ06]. scores [Buz09, Liu07b]. screening [EK04, WG03]. sd0971 [Ano01j, Ano01i]. SDE [AH06b]. SDEs [Bri00, He09, Hyn07, Kim01, NS06b, WX09]. Seasonal [DDN08, BLS04, DG07, Psa00, Rod00, SO00, Shi04a, UD09]. seasonality [Kuz05]. Seat [SD06]. Second [CB01, Cor04b, DS02b, ASG04, ADJ06, BR08, CW01, DP03, HRV08, Nic08c, RP02, Ris05, Sgi09]. Second-order [CB01, Cor04b, DS02b, DP03, Nic08c, Ris05]. seemingly [WV08]. seen [FS09, Lee04a]. segment [Huo05]. segregation [CP05a]. selected [Jen01, KK01, KMV09, Lia09c, VJ08]. Selecting [Rei08, Shi01]. selection [Ago02, AP02b, APS09b, BBR05, CCS06, FW08b, GP07, GMdRV02, Gua07, HH03a, HS07b, JFZ05, KF08, Kul01, Leh05, Lia02a, Mei08, Moj01a, MGB05, Omo07, Pre00, QF02, RM03, So03, WY08b, Yeo05, ZP02, ZK06, ZX09, dlHRB05]. selective [IR08]. selectively [Lop02]. selector [WCC07]. Self [Den07, Wan02c, AM02a, BM01b, CS00b, Den03, JH08, JR07, Koz05a, Koz05b, Kul06, LPS09, MC07a, ST06, Sal02, iS01b, VA07, Wan05b, YZ08, Zaj09]. self-adaptive [BM01b]. self-affine [Zaj09]. self-decomposability [CS00b, Koz05a, Koz05b, iS01b]. self-decomposable [AM02a, Koz05a, Koz05b]. self-exciting [Wan05b]. self-financing [Sal02]. Self-normalized [Den07, Wan02c, Den03, JH08, JR07, Kul06, MC07a]. self-similar [AM02a, LPS09, VA07]. self-weighted [YZ08]. selfdecomposability [Cho08a, Jur03]. selfdecomposable [MM07, Raj09, SS06a, Szu09]. selfsimilar [MiSW00]. Semi [APV06, Pen01b, Pon06, Cho08a, EFL02, FP07, HWL08, KKAK08, Lau06, Lim06, MiSW00, MM07, OL02, Raj09, ST06, SS06a]. Semi-functional [APV06]. semi-Markov [EFL02, FP07, Lim06, OL02, Pon06]. semi-Markovian [KKAK08]. Semi-parametric [Pen01b, Pon06, HWL08, Lau06]. semi-self-similar [ST06]. semi-selfdecomposable [MM07]. semi-selfsimilar [MiSW00]. semi-stability [Cho08a]. semicontinuous [CDMLDK01, RMLD07, Ter03]. semigroup [Dar03]. semimartingale [Pro09a]. semimartingales [Ara05, CO00, Kub08]. Semiparametric [SO00, SB08, YWS09, ZZ09, HWZ09b, HZ09c, KL08a, KPK02, Mai08, MRP04, PK07b, PK08, SS04, SZ00, SZ02b, Sun06a, WP08, YZ06, ZX09]. Sen [BB08b]. sense [Pen08]. sensitive [EG09, Lep03]. sensitivity [Orm00, Tia09]. Sensor [AFS^{+}07]. separability [Kow00]. Separable [LFdVA05, LZ05]. sequence [BPvZ04, Ber08, CL06b, CC08c, Ery08, FS02, GV02b, JD08c, Kol00b, Lin08c, Lou03, Mas02b, Ose09b, SS08b, SK06b, Shu09, zW06, Wan08b, Xie09]. sequences [Aur08a, CL08b, Coe08, CG02c, DZ02, DR05, sDXlyY05, Dud08, Ery07, EF08, GM00, HP00, HW05, Has01, HH02b, HX02, Hua09, JH08, JF00, Joh02, Kra00, LY03, Liu03, dL05b, MSW07, Nyr05, RS06b, Rao02c, Shi04b, SM06b, Sze08b, TP09, Vol06, WHSL08, WJ08, XYC09, YY08a, YAT09]. Sequential [Gom00, Mar01c, Pre00, UI00, AH04, BC00b, cICP09, Con05, FT04, HKS07, HY01, HK08, IF08, JD08a, JR08, KC03, KK00b, Lee03, RH02, Tia09]. sequentially [Cho08b, FI07, Lee04a]. Serial [Hus03, HWL08, PPRW08]. serially [DL06, Omo03]. series [APS09b, Aue04, BD05a, Bat05, BMR00, BR06c, COS03, Cai03, CCG09, CHD08, Che08b, CS08c, DGZ02, DvdAW08, DvZ05, FD05, FB02, GP05a, GP07, GP00b, Gir04, GLC09, GJP01, Gue07, HNRV00, Hua03b, HB01, Ioa04, JK07, Jur00b, KSO05, LH08a, MF05b, Nic08c, Omo03, Pen01b, PS06b, PPRW08, Pro08, RC03, Roz09b, RV08, Sco03, SN04, Sha06b, Shi04a, Sto03, SQMK06, TMN01, TR02, TP03, Tri08b, TN09, UD09, VZ03, Wan05a, Wei08, Zha06, ZB09c, ZB09d]. serology [KM08]. server [KS06]. Seshadri [GG02]. Set [Bal02, Yos09, AVV09, ASAK00, Ale01, Aur08b, BB08a, Bal01, BES01, BJR00, Bor00, Cas07, Had02, IS03, KTZ03, LGLDG00, MZ04, ÖW00, RM03, Wan01, WZ05]. Set-indexed [Bal02, Bal01, Bor00]. set-indexing [Ale01]. set-up [BB08a]. SETARMA [ANV06]. Sets [Ost01, BCAC01, Baí03, Ber01, BB05, CLM03, DKM02, DC01, ETP08, Erm00, Esh04, FT04, FM03a, HH00a, Har04, Joh02, LX08, OH07, RM08, Sch08, SS08c, TSJ09, Ter08, TB09, Vol06, Wen03, Zan08, dFF02]. setting [BB09b]. settings [HW01b]. Sevast’yanov [DK07]. several [BS00, Bro08b, Kak08, Li09a, NMO08, SL06, SGK09]. sex [Ibr08]. Shannon [Cha06b, ZN05]. Shape [BM03b, Che00b, FV09, Pai08, ZPK08]. shapes [KT08, Lac04]. sharing [LR08c]. Sharp [AJ05, DR03, DR04, Efr00, Gai07, Ilt00, Ima08, OR01, Ose09b, Roo01, PK04]. sheet [BJT03, DvZ05, Tot04]. shift [Mak08c]. shifted [BRR02]. Shiga [VA07]. Shiu [ZZG06]. shock [CH09b, GH05b, Li03a, LK07]. short [Aur08a, Gap04, HY04a, Hob07]. shortfall [Nak04]. Shot [SS07c, GvLM04]. Shot-noise [SS07c]. shoulder [Zha01b]. Shrinkage [AW04, Efr00, Hua02, IR08, MS09a, Nam03]. Sibuya [CS00b]. sided [BM04, CW07, HS00, KvD05, LGH03, LWW09, Lin08b, Log03, Wan02a, Wan07a, WY08a, XZJ08, ZJ09a]. Siegert [CLO09, DO07]. Sierpinski [LO09]. sieve [APR03, Psa03]. Sieves [MB06]. Sign [TKO03, ÖW00, PS06a, WZ05]. signal [BB09a, BH09, MN01a, NH07]. signatures [NRS05]. signed [DR05, Hel09]. significance [Kim03]. significant [IKK05]. signs [Dai04a, Lev00]. Simes [CS08a]. similar [AM02a, CS01, LPS09, Por02b, ST06, VA07, YK03]. Simmons [PR07]. Simons [Cha06a]. Simple [Van08, Ath09, BM05, CmHP02, CFR06, DR01a, DG08, El 08, FRG06, FvE01, FSC03, Hou05, Jam03, Jar03, Pan04, PB09, PD09, SS09a, Vol06, Vou06, Wan02b, Wan07b, Wit00, dUÁMM08]. simplex [SH09b]. simplices [Rom09]. Simplified [DW05, JNR09, LH08b, Per02]. simply [NW05]. simulated [Li00, OW03]. Simulating [Dev02, MS04b]. simulation [BCD09, Bha08, Cha06b, DS02a, Dev09, EZA04, GdL00, LS00b, Yao00]. simulation-based [GdL00, LS00b]. Simultaneous [HS00, MZ02, TT07b, Lu06, Sha03, Tri04]. simultaneously [Lee04a]. sine [NO00]. single [Aic09, Sun09a, YPD09]. single-covariate [Aic09]. single-index [YPD09]. singular [AL00, BP08, CD01b, He09, HNS08, LN03, PP06, Wan08c]. singularity [Kim01, KL03c, Ren01, RSY02]. sinusoids [KF08]. sites [CS00a]. situation [Gut00]. situations [Yue02]. Six [Yeh02]. size [AAC00, Asc09, BB08a, BC00b, BB07c, CM01, Coo02, CV04, EZA04, GH05a, GMM04, Hei01, HP05, Hug07, Jac08, KC03, LB02, Mak08c, Sch08, Tan09]. size-dependent [GMM04]. sizes [BLS08, DC01, Jia01a, MZG08, Mon06]. skew [AVdPS02, AB00, Ars08, Ars09, BJB06, FNA09, GHL01, HC06, HC07, JBB08, KM03a, Kim08, KM09a, Kim13, KN08, LL03, Lop01, Lop08, LRW09, Umb06, Umb08, WBG04, XWL09, YGK08]. skew-Cauchy [AB00, BJB06, HC07]. skew-distributions [AVdPS02]. skew-elliptical [Lop08]. skew-normal [FNA09, GHL01, JBB08, LL03, Lop01, LRW09, WBG04, XWL09, YGK08]. skew-slash [Ars08, Ars09]. skew-symmetric [Umb06, Umb08]. Skewed [NK03, BvdV08, BvdV09, DMRA07]. skewness [HV06, KMvE00, Mar07]. skipping [Pro08]. Skitovitch [KW00]. Skorohod [Cri04, Pro09a]. Skovgaard [FC08]. slash [Ars08, Ars09, GQT07, GV08]. slash-distributions [GQT07, GV08]. sliced [TL04]. slicing [Kul01]. SLLN [DW02, WX08]. slope [Dav05, Par01]. Slow [Bél00]. slowly [MS05]. Small [And00, Aur08b, DLL00, FL08, Roz09b, Shm06, Aur08a, BB08a, BB07c, FN07, Haf06, LZ06, LW03, Lon09, Naz09, Ryc09, Tsa01]. small-sample [Haf06]. Small-time [FL08]. smallest [DH09]. Smooth [BC06, Cat01, Che02a, CV04, QF03, QH09]. Smoothed [SBF00, BMR07, TP03]. smoothing [Gue07, Hua01, Kou03, Lee00b, Lee04b, LTY08, ZL07a]. smoothly [dL01]. smoothness [HZ04a, Lee04b]. soft [Dro06, Hua02]. software [BC07, Led08a, NBK08, Wan05b]. sojourn [BNO03]. Solution [BK02, HKV04, Leo09, Ser02]. solutions [AH06b, Bez08, Drt08, FH08, LR02, Liu08, LG09, LL09c, MN04, MS08c, Por02b, Roz04, ST01a, Sit02, Szk02, TT03, Wan08c, Xu08, Yin08]. solvability [GJZ06]. Some [Abu02, AJ00, AM07, Ahm04, AF09, AZB07, Ara05, AW06, BCD09, BM07, BV02, BK01b, Bos07a, CLO09, CN03, Cho08a, CO00, Dai04b, DS01c, Di 00, Din03, DZ04, EWI04, FC03, Gel09, GT08, Har08, Hat06, HX02, Jia09a, KZ08a, KP00, KN01b, KyS05, LLJ00, LLZ04, LH08a, Lia02a, Lia01, LZ07, Liu03, dL05b, MSZ08, MGB05, PMJ05, Pel02, QC04, Ren09b, Sia00, TO03, Umb06, Umb08, Var08, WO08, WJ08, XL08, Yan03b, YY07b, Zuo01, AG09a, AV01b, BRR02, Bez08, BO02, BhRH04, BI02, CD05, Con05, DO07, Dev09, Dey05, DMRA07, Dud03, Dun06, DF11, DF02, EKS05, FY00, GJ02, GDM06, JT08, KM02a, Kla09, KM05, Kuc04b, LN09, LS06a, LNS01, Mac08, Man05b, MSC08, Nad05, OMJ^{+}04, ÖK07, Ren01, RSY02, RB06, Sek08, Ste06]. some [Wan02b, Zah00]. Space [CM08b, DMP01b, AAV02, AM02b, CB07, DMP01a, Den03, Den07, FH08, FGMS08, FT01a, FT01b, FB02, GP00c, HCV01, Ilt00, Kim01, KKH08, KK08a, Kol00a, Lim06, LT08, Ma03, Mas06a, Mir01, Pak08b, PMM09, Pap06, PS00, Pro09b, Rao02b, RA08, ST06, Sen08b, Tóm05, TN09, WZL00, YY07b, Zha01c, ZB09d, dJP04]. space-scaling [ST06]. Space-time [CM08b, DMP01b, DMP01a, Ma03]. space-valued [AM02b, Kim01]. spaced [Omo03, VV00]. Spaces [Bos04b, Bos03b, CV01, ETP08, HCSV00, HP08, HNRV00, JPL09, Lan08, LFdVA05, Liu08, LL09c, MY09, QH09, Ros09a, Sit02, Tud08, Wan07a, Yan08]. spacings [BP02, BK08c, DR03, DR04, HW01d, JG04, Mir05, Tor06]. spanning [CBR00, GBQ03, KL03b]. spare [SS01a]. spares [SL06]. Sparre [LL09a, Ren09a, YH09]. sparse [Lie01, Lim05, SS08c]. Spatial [BRRZ03, LC04, LTY08, BL08, BPvZ04, Cer02, CP05a, CFCGM08, Faz03, Fin03a, Gao03, GLT08, Gri02, KTM08, KM03b, LB07, LK04, dL01, MT07b, NMO08, RC03, RL08, RV08, Sim00, ZS08]. spatial-temporal [LB07]. spatially [CHN04, SS00b]. spatially-correlated [CHN04]. spatio [PMZP07]. spatio-temporal [PMZP07]. Spearman [SS07b]. special [BM02]. Species [GDF01, HP00, LRM00, Pre09]. specific [CCDG02, SBF01]. specification [Gri02]. specified [Coo06]. Specifying [Bry00]. spectra [CQX08, FS00]. Spectral [BYB08, IKAT09, BM02, BR02b, CFCGM08, KTM08, LR08b, Sha04, Shi04a, SXLX07]. spectrum [BM01b, Kin08]. speed [Dev07, Kon09]. sphere [Pyc07]. spheres [Sch03]. spherical [AV01b, BM07, Del03, FWW01, Fot05, Ist06, IKIS09, JS02, STM01]. spherically [PR08]. spheroidal [LR08b]. Spiking [Pal08]. Spitzer [Lac08, Zha01a]. spline [Hua03a, Lee00b, Lee04b]. split [AZ04b, OMJ^{+}04, OK06, YZL07]. split-plot [AZ04b, OMJ^{+}04, OK06, YZL07]. splits [Shi01]. Splitting [HS07b, BH04, Clé01, Tri08a]. spread [BPR01, CHL01]. Spreij [Led08a]. square [Gra09b, KA07, KFW02, Lad07, Ose09a, ÖK07, QO09, Ten01, WBG04, Wit02, Yao01]. square-integrability [Yao01]. Squared [BBP09, KMvE00, BM01a, Bir08, CM04a, CL07, DGZ02, FS08b, GJ09, JNS02, LZD09, Man05b, ODO07, SZ08b, vE01]. squared-error [JNS02, ODO07]. squares [AT08b, BB07a, BE08, Blo07, GR01b, Gua07, Gue07, HLM06, HBK06, HKLB07, JKP01, Lad07, Lon09, Pal08, SN04, SKJ07, Shu04, TW06, Vou08a, YSkPkL02, ZB05]. Squeezing [MOM07]. Srivastava [Mar07]. SSAR [Hil01]. St. [CS07]. Stability [BHY09, CMNP07, Ete07, Gag02, JD08a, Xi04, Yan00b, AR03b, Bas00, Bib05, CF00, Cho08a, Ign09, Ker08, LT00b, SL09a, SX08, WD08]. stabilizing [Fuj00, Yu09a]. stable [Aur08b, BH09, BM04, BT02, Bro09, CC02, DDN08, Fer09, Gal01, Gon08, GK09, HKV04, Has06a, Jar08, JM08, Jur07, KP00, Koz05a, Koz05b, LS05b, MRK08, MC07a, MM09, Nag03, Pat09, Pin02, PTA03, SSA09, Woe03, YY08b, ZL07b]. stable-type [HKV04]. stage [DS00a, DS01b, DGZ02, KZS08, YZ07b]. stages [KZS08]. Stahel [DH09, Ger02]. Stam [GII08]. standard [BRSL08, Cho08b]. starting [Kra09]. Stat [Ano06o]. state [CQX08, Ery08, FGMS08, FB02, Ilt00, Lim06, LV07, Mir01, Pap06, SK06b, TN09, Vag03, ZB09d, dJP04]. state-space [FB02, Ilt00]. states [DHS00]. Stationarity [BB07b, HB04, Liu07a, AG09b, BL09b, EM08, LS05a, PS00, Psa03, Psa06]. Stationary [BR06a, Liu08, Ris06, ZJ09a, BPvZ04, Bou06, CG08a, CR08, CG02c, DZ02, DG07, sDXlyY05, Dud08, FS02, FP08, Fit06, GA00, Gne03, Gri03, JH08, Kab09, Kan07, KB01, Lim06, LL09c, Per09, PS00, PS07a, PTA03, PW06, Sas08, Sen01, Sen09, ST03b, Shu09, TP09, TN09, TP01]. Statist [BvdV09, DR15, Duc10, dE00b, FT01a, GV08, HV00, HWYZ09a, KM09a, Kim13, Lin00, Oeh01, Pal01b, PS09b, PK08, Rou01a, SF12, Yu09b, vEZ02]. statistic [Ali03, BH08, Ber08, CG02b, DWC04, FvE01, GL00, HK03b, Ing00, IKIS09, Jar03, Kla09, Lou03, Lu06, Pas06, SW00b, TO03, VvdMD08, YC04]. Statistical [BH09, CCC02, DOKN07, Iac01, Jan08, Kab08, Lop02, NH07, RS07, Wan05a, Bos07b, CC08a, CR08, CV08, De 08, GJP01, Jan03, KKRV01, LA07, LS00c, MP01a, Mil07, NOP01, PAJ05, TT03]. Statistics [Bro11, Gat09, Hal09, AB08a, AM09, AR09a, AS01, BC00a, BK02, BCK01, BCC02, BA03, BS08a, BB08b, BCD09, BLB08, Bar06, BEA09a, BB04, BB02a, BV02, BX00, BK07, BW08, Bra02a, CGNW01, CXH09, CKR02, Cra04, DS09, DR02, DR15, EB03, FMS04, For00, GO01a, GE07, GZ06, GV02b, HKM04, HZ05, HZ06, Hür04, Hus03, Hut00, IM05, JB09, JSV05, JGMGPM03, JYZ08, Jon02c, Kab08, KO01, KK00a, KK05, Kle02, Koc06, KG06, LB03, LNS01, Lop08, MR09, Man05b, Mav08, Nad08, Nas09, NP00, OR01, OAAA05, Pak08a, Pap01, Pyc07, Raq03, Ruk07a, Ruk07b, Ryc08, Ryc09, SS02, SS07b, Seg01, Sen07, Sta02, VT08, WN09a, WN09b, WN09c, Woz09, YJ01, Zha08, ZS08, dlCV01, dlCC05b, Joh07]. status [Rab00]. Steffensen [GO01a]. Stein [AVV09, AT08c, Kat09, Lan06, NO06, NS02, Röl08, YW09]. Stein-rule [NO06]. Stein-type [AT08c]. step [BKK09, BB02b, Gor07, GS08b, KSO05, KL00c, KW02, Lia07, Mon06, Ver07, ZZG06]. step-down [GS08b]. step-stress [BKK09]. step-up [Gor07, KL00c, KW02]. steps [WCS07]. Stepwise [TSGG02, Kul01]. Stieltjes [LG09, OS05, ST04b, Szu09]. Stochastic [Bar00, Bar01, BS01a, Cha01, GW08, HW01d, LT06, LS00d, MSH03, Mon08, RMAA01, Sha07, SS06b, WZ09a, YT08, ZB09b, AL08, AR03b, AT08c, AS02, AHS07, BHY09, Bar02b, BS06c, Bez08, CL06a, CD05, CLF06, CmHP02, CKS06b, DLS00, DZ04, DM05, DaY08, EGM06, FM03a, FT01a, FT01b, Fie03, Gne03, GM03b, GJZ06, HR08a, Hal09, HW01c, HZ05, HNS08, Ign09, Jia08, Jia05, Jur03, KS07, KKC08, Kub08, Kur08, Kyp09, Lan08, Lee03, LV03, Lel08, LT00a, LX05, LH08a, Lia09b, LNS01, LH01, Lin09, LT00b, LR02, LL09c, MM07, MR09, MN04, MS08c, Mon06, Nag09, Nic08c, NO09, Pet08, PR09b, Pro09b, Rad07, RH07, RB08b, SL09a, ST01a, Sek08, Sha08a, Sit02, Sor09, Spr03, WD08, zW06, Wan08c, WH09, WH08, Won09, WH07, Xu08, YY08a, YV06, Yu07b]. stochastic [Zha07a, Zha09b]. Stochastically [LL09c, Ban08, KW07a]. Stokes [BM03c]. stop [DV01, GMM02]. stop-loss [DV01, GMM02]. stopped [AE08]. stopped-sum [AE08]. stopping [Bis00, BZ06, CI09, GR06, HY01, Kla09, LS05b, Mül01, RB08b]. storage [LV05]. strata [KZS08]. strategies [Sal02]. strategy [FT04, KN05, YZ09]. stratification [Van08]. Stratified [KZS08, Bre08, FWW01, HMR08, KSH07]. strength [CFM01, DS00b]. stress [BKK09]. strictly [GA00, Pad01, ST03b]. strings [CD05]. Strong [Aue04, DFX05, FMS00, He09, IP03, JY02, KF08, Lab08, Led08b, MNOP03, Sun01, WHSL08, ZCLH01, AZB07, BC00a, Bal01, Ber03, BRRZ03, Blo07, CDMLDK01, DS09, GS02, HY00, HCW08a, LN04, Li07, LY03, Mas02c, OS06, QL06, QH09, RS03b, Sha08b, Shi04c, Sun06a, Wan08b, WJ08, YY08a, Yos09, Zha01a, CV09, Chr00, Kru08, Kuc05, SM06a, Ter03, zW06, YL00, YSY08, YJ01]. Strongly [Li07]. Structural [MR00, AGJ04, Coo06, JTQ09]. structurally [Udd02]. structure [CL09, Haa08, HB04, Mam02, Ost01, Pan04, PMZP07, SS06b, TV07]. structured [RL08]. structures [DMP01a, Had02, LQ07, Mar01a, YOH00]. Student [Oeh01, AT08c, IKIS09, Oeh00, Tar05]. Student- [AT08c]. Studentization [Fuj00]. Studentized [ADR05, DR01a, HMR08, Kab08]. Student’s [Wat01]. Student’s- [Wat01]. studies [Lec08]. Study [LZC05, BBP09, Che07b, Col08, Jan08]. studying [Man05a]. Sub [BGT04, GM08, HS03a, MS04b, Muk09]. Sub-fractional [BGT04]. sub-Gaussian [MS04b]. sub-sampled [HS03a]. sub-vector [Muk09]. sub-vectors [GM08]. Subadditive [HH00a]. subclass [JNS02]. subclasses [MM07]. subcritical [GMM00, GMdP04]. subdifferentials [Ter08]. subexponential [BLS08, CTY02, Gel09, hZbG08]. subfractional [Tud08]. subgaussian [AHV06]. Subgeometric [FM00]. subgraph [And00]. subject [Bro08b, DP03, HK09, IKK05, PM01, Sha03, SZ08a]. subject-deletion [HK09]. subjective [BL03]. submartingales [Bha05b, FI07]. subordinated [Koz05a, Koz05b, Wu08a]. Subordination [iS01b, Cho08a]. subordinator [Koz05a, Koz05b]. subpattern [HA03]. Subsampling [MP00]. subsequence [Den05]. subsequences [DK05, Sto08]. subset [CCS06, Jen01, MB08]. subspace [Pad01, Tri08a]. subsystems [Kle03]. subtrees [Mut08]. success [LB02]. successes [AB07]. successive [CH04a]. succinct [LS08b]. Sufficiency [Bos07b, Liu07b, NOP00]. Sufficient [CLF06, WY08b, CMNP07, CD01b, For00, HY04b, Jia09b, LWW09, MSW07, Mas06a, Pas06, sS00a, Wen07, Yoo09]. sum [AE08, BPC09, DMP01b, HS03b, Lad07, LCJ06, Mav07, Mir05, MC07b, Nas06, Pal08, Xia08, YT08]. summable [Deb07]. summands [Spo07]. Sums [YZ02, AAV02, AZB07, ARN08, BJ00, BK01b, CW05, Cha01, CNX06, CZ07, CG08b, Clé01, Den05, DV01, DK05, Dud08, FT08, GMM02, Gon08, GK09, HCSV00, HNRV00, HCV01, HRV08, JD08c, JR07, Khr01, KRV05, Kos09a, Kos09b, KH08b, LS04a, Leó01, LP03, LW08a, LQR09, Lia00a, Lin08a, LaYY03, LQ04, Pen02a, PQ03, Pin02, PS07b, Poz04, Qi03, QL06, RW05, Ruk07b, Sha07, Sha06a, SL08, Sun01, Sun07b, Sun09a, TSJZ01, Wan02c, WT04, zW06, WA01, hZbG08]. sumsets [Fai01]. sup [KK00b, Mim08]. sup-norm [KK00b, Mim08]. super [Eng04b, HL01, NT06a]. super-Brownian [Eng04b, HL01]. super-exponential [NT06a]. supercritical [Eng04b, SY04, SV00]. superdiffusions [Eng04a]. superficial [GP00c]. Superiority [ÖK07, WV08, WC03, ZWY05]. supermartingales [Cas07]. Supermodular [WH02]. Superposition [KS06, MY06]. superprocesses [LZ06]. supersaturated [But05, CL08a, GKM03, GSKP09, IV00, KS04, KM05, LL02, LLZ04, SLC09]. supersmooth [vEG08]. supplement [WZ06b]. supplementary [HN08]. supplemented [KM02a]. support [Eng04b, HP07, Llo05, PSS00, Pro03]. supported [SS01a, SL06]. supports [Pen02b]. suprema [RB06]. supremum [Ahm02, DZB04, Pat09, Sgi01a, Sgi01b, WCS07, vEG08]. Sure [GAW08, BW06a, CH07, CL06b, CL08b, Che02a, CG02c, Dud03, Dud08, Fah00, Gon08, HKM04, Hör06, Jam03, KKH08, KRV05, Lel08, LW08a, LS01, LT00b, Mat05, Mon06, NZ01, RAS08b, Shi04b, SM06b, Sta02, TP09, WZ06a, WL08a]. surface [But01, GM03a, Yue01]. surfaces [DP00, HF04b]. surplus [HWZ09a, YZ01]. surveillance [And05]. survey [SS05a]. Surveys [EG09, SVB07, VT00]. Survival [Mut08, SZ01, TMH02, AS07, Dub02, Fin03a, Han08, Kos02, KG06, KZ08d, LS00b, MZ02, MZ06, RCdCLN09, SDR01, sS00a, SH06, ZP09, ZZ06]. survivors [ZZ06]. susceptible [HYL04]. susceptibles [Kyp09]. sustainable [Ano02f]. Switching [OAAA05, BHY09, BhRH04, CC08b, Dav04, Gas09, Haa08, Liu07a, SH09a, Yao01]. symbiotic [DFX05]. Symmetric [Had02, CFUOV00, Dar04, DMRA07, Fan02a, FH07, FS08a, Haf06, HH02a, HC06, KS08a, KKRV01, MR09, MPV01, Mod03, Nad05, Ose09a, Ryc09, SV06, Umb06, Umb08]. symmetrical [CPG07, CV08, GPC05, QC04]. symmetrization [Gel04]. symmetry [Ber08, KA07, LS02b, Osm01, Sch02, Sim01, YJ01]. Syracuse [SM06b]. system [BA08, BS00, dCB00a, Dev07, KY00, Llo05, MW09, MNN09, Mi06, ST01a, SS01a, SL06, Tal08, VZ03]. systematic [LK06]. Systems [SS08a, BRR02, BS08b, BL02, Cui02, CKLX06, FF08, Hil03, Jar08, Ker08, KS00, Li04, LH08a, MS08a, MK08c, NRS05, Ser02, Spr03, Wu08a, Zah00].

tables [KA07, ZB09c]. Tail [BvdV08, BvdV09, JKP01, KL03b, NT06a, CFCL03, Fal05, Gan00, GP00b, GO03, HNRV00, KG06, LS02a, LP09, Liu06, NIK02, Sch01b, hZbG08]. tailed [CS06, Cho05b, HS03a, HK03b, Kle02, Lin08a, MY06, Pen01a, Psa09a, SD09, TSJZ01, Wan06, WCS07, ZJ09b]. tailedness [Ibr08]. tails [AKK^{+}02, BB07b, CNX06, CZ07, dE00a, dE00b, HS03b, Jur00a, Liu07c, Liu09, SL08, WT04]. Talagrand [Zan08]. Tamhane [KL00c]. Tanner [Lia09d]. TARCH [GLML09]. Taylor [GLML09, Haa09]. teams [Poz08]. technique [BM01b, Clé01, ET05, SS09a]. techniques [BD01b, Dar03]. telegrapher [Iac01]. Tempered [Fot04, Jur07]. temporal [LB07, PMZP07, RV08]. temporally [CC08b]. tensors [Rom02]. tent [Hai03]. term [LJ01, Mam02, Mir05, RCdCLN09, ZZ06]. terms [AB04, AE00, For00, Man05b]. Terry [SZ06]. tessellations [TM09]. Test [Bha05a, Jur00a, LS08a, LA07, Alt08, BD05a, BKK09, BL07, BBP09, Ban08, BX00, Ber08, BHHM04, Bra02a, Cac01, CK05a, Che04a, CC01b, CS08c, CKS06b, Con05, Coo02, Coo04, CFCGM08, CF06, Det00, DR05, Dur03, Fie03, GR01b, GL08, HL02, HS08a, HS00, HK03b, HT05, HWL08, Hu09, HB01, HcW08b, JG04, Jia01b, JL09b, Kab08, KKY09, KC03, KY04, KL00c, Lan08, LN02, LN04, Lee06, LS00b, LS04b, LWT08, LP09, LX08, Log03, LF07, LZ05, Mak08b, Mak08c, Nag09, NMO08, NZY09, Oht01, Oht02, Orm00, Osm01, ÖW00, Ozt07, Pal01a, PS06a, PLLJ06, PP01, Psa03, Rod00, Sch08, Sim00, SKPP03, Sim07, SGK09, SL00, SZ02a, SR04, Tar05, TKO03, Wan07b, WF00, YW07, YC04, YPD09, ZS08, ZLQW07]. Test-based [LA07]. Testing [BL08, BBP09, BHK04, Coo06, DV08, Dub02, Hol08, JB03, KD03, KD04, LGH03, Lim05, OS03, Rab00, Sch02, sS00a, SZ02a, SX08, WP00, YZ07a, YC05, AY02, AZS08, BD01b, BD05b, Bis08, CP08, CP05a, CS00c, CY02b, De 08, Fie03, GJ02, GVS00, GVGP08, Gor07, GS08b, Gor09, Jan08, Jar03, Kim00b, KKRV01, KS05, KY06, Li06, LPY04, LA07, Log03, Mei09, NP00, PP05, PLLJ06, PPRW08, Psa06, SH09a, Son09, Ste06, TMH02, Tor06, TT07b, Vou08b, Wan07a]. Tests [Cer02, AP02a, AH01, AW06, BPR01, Bur00, dBC06, CW07, CS01, Coo03, CV04, EHMM08, ED02, FCN02, FC08, Fig07, Hu06b, JTQ09, JL09b, JS02, Kla00, LG03, Lia04, Lie01, LC05, Llo08, Lo05a, MU04, MS07b, Miz00, Nic02, NG02, Psa00, Psa02, RL08, Sen01, Sen08a, Sen09, SO00, Voc06, WG03, WZ05, Wan06, WJ05, WN09b, WN09c, WX00, YZ07a, ZD04, ZR05]. tests-of-fit [Bur00]. th [AK09, CV01, Cha05, DR04, DLB05, Dem08, KR04, Kli06, LBMD05, NT06a]. th-order [CV01]. Theil [She09]. Theil-type [She09]. their [ARB09, AA08, Als01, AA07, CH01, CV08, Fot04, GHL01, HK03a, KM03a, KM09a, MSH03, ST06, Sch00, Sha04, Yeh02, vEZ01, vEZ02]. Theorem [Che08b, AKK^{+}02, BP06, BW06a, BR02a, Blo02, Cai00, Cha09, CM03, Cri04, DK07, DR02, DR15, DW05, Fah00, FH08, FI07, GS03a, GW05, Gon08, Gut06, HR08a, Hal09, HP05, HNRV00, Ing00, KW00, KO01, Lin08b, LR02, LaYY03, Ma09, MV01, Moj01b, Nyr05, PMJ08, PR07, Pro04, RB08b, Sap08, SW00a, Shu09, Ste01, Sto08, Tal08, VB08, Xia09, Yan00b, Yu09c, ZZ08, BHH08, Bla07, BdB09, CC02, CPD00, Dud03, FU08, GA00, GAW08, JY08, JH08, JR07, KB01, KK08a, LW08a, Mal01, Neu04, RW00, RS03b, Sou01, WL08a, Web06, vZ00]. Theorems [KMM00, Als01, AZB07, Bir08, BMR00, CG02c, DS06, Faz03, HH00a, HKM04, HS03b, HPR02, JJQ08, Jia05, Kag02, Kul06, Lau05, LZ06, LCZ09, Li09c, LY03, LW03, Mat05, MS05, NC09b, PMJ05, PMM09, Shi04c, Son00, Sta02, Sto09, zW06, Wan08b, Woz09, WJ08, WX09, YY08a, CL08b, Dud08, JH02, SW06]. theoretic [Ili00, Irl04]. theory [BEA09a, BS01a, BL02, CYJL01a, DLS00, FF08, Gut06, Hör06, HH00b, LLZ04, Nag09, NOP01]. there [Oht02, Sen07, Sen08a]. theta [Dev09, GAW08, Sza01]. theta-dependence [GAW08]. thick [MS02b]. Thin [MS02b]. Thinking [DOKN07]. thinning [TD08]. third [Pen02a]. thoughts [Kel06]. Three [D’A06, ASS02, GKM03, KK06b, KR09, LSSG08]. three-level [GKM03]. three-parameter [KR09, LSSG08]. Threshold [DFS08, HW01e, BE09, Coo03, Coo04, FR01, HB04, Ibr08, Liu06, Liu07a, PS06a, PBH09, ZCCH09]. thresholding [Che08c, Dro06]. tied [Eis09]. Tight [Ryc09]. Tightness [GJ01]. Tikhonov [DSBA09]. tilted [Cer07, FvE01]. tilting [KT03]. Time [Gue07, Shu03, APS09b, AH04, Aue04, BD05a, Bat05, BNO03, BO02, Bib05, BMR00, BR02b, COS03, Cai03, CCG09, CG08a, CHD08, CP06a, Cha05, CS06, CQX08, Che08b, CS08c, CY02a, CM08b, DMP01a, DGZ02, DMP01b, DY04, DD06, DN09, Dit07, Dow09, DvdAW08, ETC00, EsO08, FS01, FL08, FG04, FD05, FB02, GP05a, GP07, GP00b, Gir09, GL04, GJP01, GK04b, HH03b, Hoo02, HB01, IPP04, Ioa04, IKAT09, IS03, JKZ09, JK07, KKY09, Kin01, KZ08c, Kra09, Lab08, LS04b, LW08b, Ma03, Mac08, Mak08a, Mam02, MS05, MNX09, Mi06, MS03b, Min09, Nic08c, OMM09, Omo03, PP09, PMM09, PI09, Pen01b, PS06b, PPRW08, PTA03, Pro08, RC03, SVB07, SDR01, Sco03, Sen08b, SSB09, SN04, Sha06b, Shi04a, SV06, SLF09, SQMK06, SZ08a]. time [Tal08, TMN01, TR02, TP03, Tri08b, TN09, UD09, VA07, Wan01, Wan05a, WY08a, WH08, Wei08, WZ03, WL08b, XZJ08, YY07b, YZL08, YY08b, Zha01c, ZB09c, ZB09d]. time-delayed [GK04b]. time-dependent [Bib05]. Time-frequency [Shu03]. time-homogeneous [Dow09]. time-integrated [Mac08]. time-series [JK07, Sco03]. time-space [Sen08b, YY07b]. time-use [SVB07]. time-varying [FG04, Mam02, SZ08a]. times [Als01, BE09, BGT04, BC07, BN08, Che07a, CO00, FP07, FD09, HF04a, HZ08, Ker08, Kla09, Kok01, LC00, MS05, MY06, Nik08, Pal09, RB08b, SH09a, ST09, Sze08b, dUÁMM08]. Titterington [WZ06a]. Tobit [Wan07b]. tolerated [MMB07]. tomography [dlPGM08]. tool [Luc00]. tool-wear [Luc00]. Topchii [AR03a]. topologies [Ale01]. topology [MSW07]. Total [Baí03, HWZ09a, AAC00, BLS08, BK04, CBR00, DV01, IMS06, KKY09, KC03, LS04b, LL09a]. totally [Moj01a]. tournament [And00]. toxic [CS00a]. TP [GV02a]. trading [Ara01]. traffic [Din00c]. training [Had03]. trait [ZCL08]. transaction [Pou08]. transcription [YPD09]. transect [Zha01b]. Transfer [Kag02]. transform [Bar00, HH02a, LCZ09, Mon07, Roz09a, Sha07, SY07]. transformation [BN00, BE08, BGN08, Bro07, Bro11, CWTB09, CF01, Fuj00, GR01a, GL00, IWWJ04, KS05, QF03, Vou08a, Yeo05]. transformations [CYJL01a, D’A06, KKRV01, Yu09a]. transformed [CYJL01b, HB04, Liu06, Liu07a, Sar00]. transforming [YJ01]. transforms [Cai02, MS07b, MY09, MV01, NQMRLÚF01, Szu09, Yu08a]. transience [Tak04]. transient [Dai05, Leo09]. transition [CL08c, CV04, DS01b, Dow09, Pas06, Ros06, SY07, ZCLH01]. transitive [PRB99, Pal01b]. Transposition [Cho05b]. tray [KFW02]. treatment [AGG02, AJ00, ADJ06, ADJ07, Bru09, DWZ05, HC04, KG09, MLT00, PK07b, PK08, Ruk07b, YZ07a]. treatments [HT05]. tree [CP05b, Dai05, GBQ03, JL09c, KL03b, Kin08, MN09, MS02b, Mut08, SGK09, YL00, Yan03b]. tree-order [CP05b]. trees [BT04, Che07a, CM01, CBR00, DDP08, HS07b, Jon02a, KW07a, KKM^{+}08, Kon02, KO03, KO08, Kon09, Kou06, LW03, Pal09, Shi01, SLF06, TT08, dlPGM08]. trend [AP02a, AHS07, BHHM04, BH05, CK05a, LV03, Psa02, Sen01]. trend-break [Sen01]. trended [CCG09]. trial [KG09]. trials [Ery08, MPS06, Ruk03, SK06b, Tan09, YW07]. triangular [BW07, BDM08, CPD00, GX08, Has05, Has06b, Kan07]. trigonometric [DMB02]. Trimmed [KH08b, BD01a, BW08, GL00, Kim00a, LTC04, Oli01]. triplewise [Kan07]. triply [RL08]. trivariate [JB09]. Truncated [TP01, BM09, HcW08b, IP03, Lel08, LL09b, Pon06, Poz04, SZ00, SZ01, Sun06a, Tar01, Wan08a, ZR05]. truncation [HY00, Ima08, KL08a, LG03, MS08a, SB08]. truncation-by-death [Ima08]. truth [FL02]. Tukey [HR08b]. tuple [Lou03]. turn [And05]. turning [And05]. Two [Aic09, Als01, CP05c, Ozt07, Xia08, YZ07b, ASS02, ASG04, AM09, AZ04a, BCD09, BJT03, Bha05a, But04, But08, Cai00, CFM01, Cha01, Che00b, CG02b, CCC02, Che04b, CKS04, CH01, Dar01, Erm00, Esh04, EKM02, EK03a, EK04, EKS05, GM08, GQR06, Geo07, GX08, GL00, Han03, HC04, HY04a, Hei01, HH06, Jac08, JBB08, KG09, KK00a, KK04, KSO05, KvD05, KZS08, LZZ08, MW09, Ma09, MB08, MN01b, NZ01, Neu04, Ome08, Pui08, SS02, Sch08, SH06, SKPP03, TG02, TKO03, VZ03, Vin08, Wan02a, WY08a, WV08, WY08b, WC03, YS00, YSN01, YB07, YT08, YKB07, ZZG06, ZJ09a, Zuo01, vEZ01, vEZ02]. two-component [VZ03]. two-dimensional [CG02b]. two-factor [AZ04a, YB07]. Two-group [Aic09]. two-level [But04, But08, EK04, YB07]. two-marginal [CH01]. two-order [Hei01]. two-parameter [BJT03, Che00b, HH06, JBB08, MN01b, Pui08, YSN01]. two-part [Han03]. two-period [KG09]. Two-sample [Ozt07, GL00, KK00a, LZZ08, Neu04, SS02]. two-sided [KvD05, WY08a, ZJ09a]. Two-stage [YZ07b, KZS08]. two-step [KSO05, ZZG06]. two-type [Ma09]. two-variable [Geo07]. two-way [WC03]. Type [Has09, AAB04, AT08c, Arc00, Ars05, BCK01, CM08a, CGNW01, CM04b, CMS05, D’A06, DZ04, Dur03, Fer00, FM04, GO01a, GR06, GSTS07, GZ06, HKV04, Hen00, HCSV00, HWYZ09a, HWYZ09b, IKP08, IKIS09, JPL09, JKK05, JMP06, KO04, Kar02, KN01b, Kra02, KZ08d, Kur08, Li03a, LWT08, LQR09, Ma09, Nan08, NP00, Ome08, Ose09a, Pel00, PQ03, Rao02a, Rao02c, Raq03, Rou01a, Rou01b, She09, Shi02, SXL05, WA01, XNF07, YSkPkL02, YW04b, ZS08, dSF06a, dSF06b, dSF08, dlCC05a, BM03a, Fot05, LF07, Zhe01]. type-II [AAB04, BM03a]. types [Ome08, WY08b, Zuo01].

U [YW04b, ZJ09a]. U-type [YW04b]. UAR [RP02]. Uhlenbeck [Bis01, Bis06, Bis08, BN08, Dit07, EM08, GJ09, KS08b, Liu08, Lon09]. ultimate [YH09]. unbalanced [Gra09b, GSKP09, HS07a, Wul08]. unbiased [TW06]. Unbiasedness [Hu06b]. Unbounded [Zap08, BP06, RW00]. uncertain [MLT00]. uncertainty [D’A06]. unconditional [OH05]. uncorrelation [Ost01]. Understanding [Kin01]. unequal [GM08, LL08, Lo05a]. unequally [Omo03, VV00]. unexpected [GS08a].

Unfair [Bea07]. Unfolding [Pro09a]. unification [RCdCLN09]. unified [AGG02, EKM02, FSC03]. Uniform [Bas00, Ber03, hCW09, Har04, Kav08, LL05, Woz09, AR02, BR02a, Blo07, BPC09, Bur00, CV01, Che02a, DG09, Din00a, EWI04, FY00, FQ03, FvE01, FS08c, Gag02, GMZ07, GR00, HNRV00, HCV01, Jon02a, Jon02c, KV02, Khr01, Lac04, Lee04a, MNZ05, MK08b, MR08a, Mla09, Neu07, OS06, PW00, RP02, Seg01, Ter08, Wit05, YJ01, ZY08, hZbG08]. uniformity [Fig07, ZQ06]. Uniformly [Yan03a, Fit06, GO02, JY08, Jia09a, LYZ08, QH09, SH09b]. unimodal [KMvE00, Ryc09]. Unimodality [Ali03, BB02a, Cra04, ODO07]. Unimprovability [Gor07]. Unimprovable [Ser02]. union [HN08]. uniquely [CH01]. Uniqueness [Jon02a, STM01, BM03a, DFX05, He09, Jia09a, Krä07, Wan08c, YS00, Zha09b]. Unit [BRF06, Fuk05, Sen09, Vou08b, CK05a, Coo03, CV04, Coo04, ED02, HK03b, LX00b, Mak08c, MR00, Nag09, PS06a, Psa00, Psa02, SS01a, Sen07, Sen08a, SO00, SH09b, Vou06, Wan06]. Unit-root [Fuk05, Psa02, Vou06]. units [Li02, WK07]. univariate [BC03, GO01b, Kos02, Lia02b, SC05, YS00]. Universal [Bro07, Bro11, MT04]. universally [Udd02]. unknown [AAC00, Ahm04, LdC07, TSGG02]. unobservable [CKM03]. unobserved [RS06a]. unrelated [WV08]. unreplicated [ML03]. unspecified [SS08b]. unstable [BPvZ04]. unsupervised [KKM^{+}08]. until [Cha05, HA03]. upcrossings [Gri03, TS01]. updating [Jon08, So03]. upgrade [SS01a]. upon [LS08c]. Upper [GMM02, Nyr05, YH09, CK05b, CDMLDK01, DK08, HBA05, KL00a, RMLD07, Ryc02, Ter03, Wor01]. urn [CH09b, KMR00, MMB07, OM06]. Use [AT08c, ARB09, BP02, CP05a, Hug07, JB05, KM08, KK06b, Lia09b, RH02, SVB07]. used [AAK00, Din00b]. Useful [Mus00, MW00]. using [AE08, AM05, AR09a, And05, AR09c, BC06, BH09, BL07, BT04, BBZ08, Bha08, Bis01, Bou09, Bru09, Bur00, DMP01b, EK03b, FWW01, FNA09, FSX02, Gal01, GVMSN04, GP00c, HW07, Hug07, JK07, Kab08, KM03b, KSG08, Lee00b, LRM00, Liu07b, Mad02, Mid08, MP08b, MZ04, NRS05, Nav08, PK04, PNJ08, PS00, Psa09a, RM03, Rei08, SSB09, TS09a, TT03, Tom09, Tor06, TT07b, YZ07a, Yoh08, ZJ09b]. usual [Lou01, Sor09, Wan02a]. utility [Bro09, FW08b]. utilizing [FK08, KG09].

validation [Gua07, Lia09a, Son09, Wal02]. Validity [DS01b]. value [AP04, BV02, FR02, Fal05, Fit06, FF08, GN08, Haa09, Hür03, HLM06, Li09a, LC05, MR09, Pen08, Sch01b, TT07a, XW08, YZL08, dHC06]. valued [AAV02, AM02b, BS09, Den05, DD06, DvdAW08, DaY08, Ery07, EF08, FT01a, FT01b, FR06, GLX07, HS03a, HCV01, Kim01, LGLDG00, Rao02b, Shi04c, Tóm05, Wan08b, Yu08a, Zap03, ZB08]. values [AB09, AA07, Bis00, CS08a, Dah06, FS02, GJ08, GVMSN04, Hai03, HBA05, KPI09, KS07, KM02b, KL00c, Li06, dL01, MS03b, Moj01b, MC09, Mül01, NZ01, Sim06, SBH03, TT07b, Wan07a, WL08b, ZB09b, Hu09]. VaR [IKK05]. Variability [Kle02]. Variable [JFZ05, Kul01, Yeo05, ZX09, AL06, BBR05, D’A06, Ewa08, FD05, Geo07, GMdRV02, HS07b, Pro08, Roz09a, Sar00, Shi04b, Wan02b, WY08b, WHSL08, WCC07]. variables [AR09a, AA08, AZB07, AK09, AVG08, ARN08, Aur08a, BW07, BDM08, BPC09, Cac01, Cai02, Cha01, CZ07, CG08b, CM03, CH09a, Den05, DG09, Dev02, DR02, Dud03, Duk07, Eis08, EL03, Esh04, Ete07, FRG06, Fer03, FT08, FJ01, Fur07, Fur02, GP01, GW08, GX08, GMZ07, GRT07, HO08, Hol08, HY04b, HRV08, HWL08, HWYZ09a, HWZ09b, HWYZ09b, Hua03b, HZ06, IS01a, II08, JJQ08, JH02, JKK05, JD08c, Jur00b, KW00, KO04, Khr01, KKH08, KK08a, KRV05, Kra02, Kru08, Kuc05, Kuc07, Kuc09, LS04a, Lec08, Lee04a, Leh05, LP03, Li00, Li09b, LQR09, Lia00a, LY03, LaYY03, Liu03, Liu07c, LL09b, Liu09, Lyh05, MNZ05, MK08a, MB08, Mat05, Mav07, ML09b, MSZ08, MSH03, MPY03, Moo08, Muk09, Mül01, NO06, Nas06, NC09b, Nyr05, Oht02, Oli05]. variables [Ost01, PL07, Pel02, PS07b, Poz04, QL06, RW00, Roz09b, Ruk07b, Ruk09, Sch03, Sch08, Sha06a, SL08, Shi01, Sim01, Son09, Sun01, SHV05, Sun07b, Sun07c, Sun08b, TB09, Wan04, WT04, Wan08b, WZ09b, Wen03, Wit02, WJ08, WSM09, Xia08, YT08, YZ02, ZL07a, ZB09a, hZbG08, DR15]. Variance [Fuj00, TS09a, Yu09a, Alt08, And00, Ara05, AS06, BD01a, BCLBMR00, DW00, dE00a, dE00b, EGM06, GPC09, Gut00, HP09, JTQ09, KR04, KvD05, KGC09, Leó01, LL08, LBSM00, MP00, Mak08c, Mal01, Moo08, Nat01, NZY09, Oht01, Oht02, Pen01b, Pom01, QL06, Ros07, SS05a, Sen07, Sen09, SV02, VvdMD08, Wul08, YZ08, ZL09]. variance-optimal [Ara05]. variances [BCK01, CGK06, GMM02, Lo05a, OS03, Ryc08, SD06, SGK09, TSGG02, Wan02a, XY07]. variants [SC08]. variate [Fan02a, Tri08b]. variates [LX08, Pap01, SZ02a]. variation [Bre06, CH07, DV01, FV08, GH05b, Has06c, MP08b, Yos09, YYL08]. Variational [Com07]. Variations [BT02, GP00c, Ros09a]. variogram [GSGMFB03, PMZP07]. variograms [Bar05, SG06]. various [Hil02]. varying [BLS04, BO02, CNX06, CZ07, CF06, FG04, HWL08, HWZ09b, HZ09b, HZ09c, KO08, Liu07c, dL01, Mam02, MS05, RA08, SL08, SY04, SZ08a, YZ06, YP09, ZX09, ZY04]. varying-coefficient [HWL08, HWZ09b, HZ09b, HZ09c, YZ06, ZY04]. Vatutin [AR03a]. vector [BR06c, Cho05a, DD06, Duc04, Duc10, Gal09, Haf06, KMS08, LPS09, Muk09, OS02, PR09a, PS07a, UD09, Yan00a]. vectors [BC06, BP08, CP01, FM03a, FJ04, GM08, GHL01, Has06c, KPI09, KM03a, KM09a, Lop01, MW09, MS04b, Pel00, Pen02a, San05, Shi04c, Sou01, TKO03]. verifiable [Lel08]. version [Bar07, Dai04a, Dud08, Kra09, MS03a, PN09, Sza07, WZ08]. versions [BW06a, RS03b, Sta02]. versus [Bru09, JS09]. Vertex [Gup09, Dai03, Dai04b]. vertex-reinforced [Dai03, Dai04b]. via [Ago02, Aur08b, BBG05, BD01b, CFM01, Cha06b, cICP09, Clé01, Coo02, EZA04, GLX07, HH02a, JB09, JD08a, KKM^{+}08, LH01, MZ02, MN08, OR01, OAAA05, TMH02, TM09, YC06a, ZS08]. viability [Maz00, MR03]. view [Hua01, MNOP03, NOP00]. violation [Sch07, Wan05a]. Viot [dSF06a, dSF06b, dSF08]. virtual [Min09]. Vogelsang [HS08a]. volatility [DL06, Haa09, Sek08, TAP05, TPA08]. Volterra [AR03b, OE08a, WZ09a, Wan08c]. Volume [Ano01c, Ano01e, Ano01f, Ano01d]. volumes [Vit08]. voter [Pre09].

w.r.t [CB07]. Waiting [BE09, HH03b, Cha05, MS05, MS03b, Min09]. Wald [LWT08, dlPZ02]. walk [AKK^{+}02, Che07a, CFR06, Dai05, FBVW00, FS09, GLX07, KK04, KKAK08, Lac08, Let02, Sta00, Tan07, Vol06, WCS07, WL08b, Zha01a]. Walker [SS06c]. walks [Che06, CWZ08, Dai03, Dai04b, DO07, DDP08, Dev07, Hil02, Kin08, Kon02, Kon09, LC00, Lud09, MS05, MNX09, Pad01, PRB99, Pal01b, Pal09, QS05, TT08, dlPGM08]. waste [CS00a]. Watanabe [VA07]. Watson [Cai01, Dai05, FG06, Gho03, GMM00, Kin08, Ma09, MN09, MS02b, Mut08, SV00]. wave [AL00, Mac09]. Wavelet [PMA08, ZY04, AD01, ABG07, BBG05, Cai02, Efr00, Hua02, LVL09, LYZ08, LX01, OW03, OS03, TN09, Var08]. wavelets [PK04, Pen02b, WIL01]. way [Gut00, Hut03, Sia00, WC03]. ways [BL07]. Weak [BJR00, BJT03, JPL09, MN04, vEG08, BS06a, Bis06, CW08b, CPD00, DS06, DM05, Fuj08, GLS08, GJ01, HR08a, Hal09, HCSV00, HNRV00, Hua03a, Kra00, Küh01, LC04, Mas02a, Maz00, ML09b, Nas09, NOP01, Ose09a, Pak08b, SZ09b, SBH03, Sun06b, SHV05, Wu08a, YZL07, Zap08, ZW01, Gut06]. weakly [Che08a, CL08b, DHS00, DS09, JM08, KO04, LZ01, dL05b, San05]. wealth [LS04b]. wear [Luc00]. Wei [OM06]. Weibull [BK08a, FV09, JL07, KR09, Orm00, Zhe01]. Weierstrass [Szu02]. weighing [CGK06, GKM03, Nic08a]. weight [Bla08]. Weighted [Cai01, HK09, HLM06, SH09b, Wu04, Ago02, Ago06, AAV02, AZB07, BK08b, BH08, Bar09, Bur00, But01, Che04a, CNX06, CLR08, CBR00, DB09, DLL00, Ete07, Fan03, FNA09, FN07, GE07, GvLM04, HKS00, Hou05, HNRV00, HCV01, Hua03b, JB05, JMP06, JD08c, KG06, LS04a, LQR09, Lia00a, Mav08, ODO07, PQ03, Pin02, PS07b, Roz09b, Ruk07a, SS08a, SL08, sS09c, Sto03, Sun01, Sun07b, Sun09a, TW06, Web06, Yan03a, YZ07a, YZ08, Yeo07, YT08, YSkPkL02, hZbG08]. weighting [BS06c, Bla08, DW00, TT08]. weights [Aur08a, FS08b, HCV01, JY02, WC04]. well [CFCL03]. well-matching [CFCL03]. Wherefore [CS01]. which [BLV04, EsO08, Kan07, Kla09, YOH00]. white [MN01a, Wan05a]. Whittle [LS06a, Rao02b]. whose [QS05]. Wick [VB08]. wide [CG08a]. Width [MZ06, IF08, LB02]. Width-scaled [MZ06]. Wiener [BJR00, Bre06, CC01a, CB07, Gir09, MMV01, Ros09b, Shm06, WZ03]. Wigert [LG09]. Wilcoxon [DR05, Ome08]. Wilcoxon-signed [DR05]. Wilcoxon-type [Ome08]. Wild [YC06b]. Wilks [UWF06]. Wills [VW08]. Winsorized [TG02]. wise [Bar02a, BN01]. Wishart [RW05, SvPtK08]. withheld [CS07]. within [BS08b]. without [Fro03, HH02c, LLS06, NO06, Pen02a, Zha01b]. Wittmann [Den07]. wls [YOH00]. Woodruff [SW01]. Woyczyński [HNRV00].

XY [Nad05]. years [Joh07]. yield [AH02]. Yor [Hob07, IK04, Kou06]. Yule [SS06c].

Zabczyk [Sza07]. Zero [HWL08, CQ03, CL07, II08, Kok01, SS08c]. zeros [Mah04, ZZ09]. zeta [Lau05]. zeta-function [Lau05]. Zhou [KN05]. Zipf [Ano02f, Yeh02]. Zygmund [DS09, GS09, Kra02, RL01]. Zygmund-type [Kra02].

References

Abdelhamid:2001:MRC

[AA01] Hassairi Abdelhamid and Masmoudi Afif. The mean radius of curvature of an exponential family. Statistics & Probability Letters, 55(2):213–220, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001353.

Amini:2007:FIR

[AA07] Morteza Amini and J. Ahmadi. Fisher information in record values and their concomitants about dependence and correlation parameters. Statistics & Probability Letters, 77(10):964– 972, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000338.

Alamatsaz:2008:OCN

[AA08] Mohammad Hossein Alamatsaz and Somayyeh Abbasi. Ordering comparison of negative binomial random variables with their mixtures. Statistics & Probability Letters, 78 (14):2234–2239, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001168.

Alvarez-Andrade:2004:EQP

[AAB04] Sergio Alvarez-Andrade and Laurent Bordes. Empirical quantile process under type-II progressive censoring. Statistics & Probability Letters, 68(1):111–123, June 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000756.

Ahmad:2000:EPT

[AAC00] Manzoor Ahmad, Serge Alalouf, and Yogendra P. Chaubey. Estimation of the population total when the population size is unknown. Statistics & Probability Letters, 49(3): 211–216, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000407.

Al-Awadhi:2004:IAR

[AAHJ04] Fahimah Al-Awadhi, Merrilee Hurn, and Christopher Jennison. Improving the acceptance rate of reversible jump MCMC proposals. Statistics & Probability Letters, 69(2):189–198, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520400183X.

Abouammoh:2000:NRB

[AAK00] A. M. Abouammoh, R. Ahmad, and A. Khalique. On new renewal better than used classes of life distributions. Statistics & Probability Letters, 48(2):189–194, June 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299002047.

Ahmed:2002:RCC

[AAV02] S. Ejaz Ahmed, Rita Giuliano Antonini, and Andrei Volodin. On the rate of complete convergence for weighted sums of arrays of Banach space valued random elements with application to moving average processes. Statistics & Probability Letters, 58(2):185–194, June 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001268.

Arnold:2000:SCD

[AB00] Barry C. Arnold and Robert J. Beaver. The skew-Cauchy distribution. Statistics & Probability Letters, 49(3):285– 290, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000596.

Ahmadi:2004:CIQ

[AB04] J. Ahmadi and N. Balakrishnan. Confidence intervals for quantiles in terms of record range. Statistics & Probability Letters, 68(4):395–405, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001294.

Ahmadi:2005:DFC

[AB05] J. Ahmadi and N. Balakrishnan. Distribution-free confidence intervals for quantile intervals based on current records. Statistics & Probability Letters, 75(3):190–202, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002269.

Antzoulakos:2007:DMO

[AB07] Demetrios L. Antzoulakos and Michael V. Boutsikas. A direct method to obtain the joint distribution of successes, failures and patterns in enumeration problems. Statistics & Probability Letters, 77(1):32–39, January 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001921.

Ahmadi:2008:NCI

[AB08a] J. Ahmadi and N. Balakrishnan. Nonparametric confidence intervals for quantile intervals and quantile differences based on record statistics. Statistics & Probability Letters, 78(10):1236–1245, August 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003999.

Aknouche:2008:EHO

[AB08b] Abdelhakim Aknouche and Mohamed Bentarzi. On the existence of higher-order moments of periodic GARCH models. Statistics & Probability Letters, 78(18):3262–3268, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003118.

Ahmadi:2009:PCR

[AB09] Jafar Ahmadi and N. Balakrishnan. Pitman closeness of record values to population quantiles. Statistics & Probability Letters, 79(19):2037–2044, October 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002375.

Antoniadis:2007:PWM

[ABG07] Anestis Antoniadis, Jéremie Bigot, and Irène Gijbels. Penalized wavelet monotone regression. Statistics & Probability Letters, 77(16):1608–1621, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001368.

Abundo:2002:SCC

[Abu02] Mario Abundo. Some conditional crossing results of Brownian motion over a piecewise-linear boundary. Statistics & Probability Letters, 58(2):131–145, June 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001086.

Arnold:2000:MMD

[ACSGV00] Barry C. Arnold, Enrique Castillo, José María Sarabia, and Laureano González-Vega. Multiple modes in densities with normal conditionals. Statistics & Probability Letters, 49(4):355–363, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000687.

Amato:2001:CPM

[AD01] Umberto Amato and Daniela De Canditiis. Convergence in probability of the Mallows and GCV wavelet and Fourier regularization methods. Statistics & Probability Letters, 54(3):325–329, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001110.

Aggarwal:2006:BRT

[ADJ06] M. L. Aggarwal, Lih-Yuan Deng, and Mithilesh Kumar Jha. Balanced residual treatment effects designs of first and second order. Statistics & Probability Letters, 76(6):597–600, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003512.

Aggarwal:2007:BRT

[ADJ07] M. L. Aggarwal, Lih-Yuan Deng, and Mithilesh Kumar Jha. Balanced residual treatment effects designs of first order for correlated observations. Statistics & Probability Letters, 77(9):901–906, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000284.

Adamidis:2005:EEG

[ADL05] Konstantinos Adamidis, Theodora Dimitrakopoulou, and Sotirios Loukas. On an extension of the exponential-geometric distribution. Statistics & Probability Letters, 73(3):259–269, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001100.

Agho:2005:ESA

[ADR05] Kingsley Agho, Wen Dai, and John Robinson. Empirical saddlepoint approximations of the Studentized ratio and regression estimates for finite populations. Statistics & Probability Letters, 71(3):237–247, March 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204003050.

Asadi:2000:REC

[AE00] Majid Asadi and Nader Ebrahimi. Residual entropy and its characterizations in terms of hazard function and mean residual life function. Statistics & Probability Letters, 49(3):263–269, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000560.

Abd-Elfattah:2008:BSS

[AE08] Ehab F. Abd-Elfattah. Bivariate stopped-sum distributions using saddlepoint methods. Statistics & Probability Letters, 78(13):1857–1862, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000515.

Asadi:2005:DGI

[AES05] Majid Asadi, Nader Ebrahimi, and Ehsan S. Soofi. Dynamic generalized information measures. Statistics & Probability Letters, 71(1):85–98, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002871.

Ahmad:2001:OBK

[AF01] Ibrahim A. Ahmad and Yanqin Fan. Optimal bandwidths for kernel density estimators of functions of observations. Statistics & Probability Letters, 51(3):245–251, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520000136X.

Ahmadi:2009:SCO

[AF09] Jafar Ahmadi and M. Fashandi. Some characterization and ordering results based on entropies of current records. Statistics & Probability Letters, 79(19):2053–2059, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002399.

Alba-Fernandez:2005:BED

[AFMGJG05] V. Alba-Fernández, J. Muñoz-García, and M. D. Jiménez-Gamero. Bootstrap estimation of the distribution of Matusita distance in the mixed case. Statistics & Probability Letters, 73(3):277–285, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001203.

Anderson:2007:SAR

[AFS^{+}07] Dale N. Anderson, Deborah K. Fagan, Rey Suarez, Jim C. Hayes, and Justin I. McIntyre. Sensor analytics: radioactive gas concentration estimation and error propagation. Statistics & Probability Letters, 77(8):769–773, April 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600349X.

Apanasovich:2008:PEF

[AG08] Tatiyana V. Apanasovich and Edward Goldstein. On prediction error in functional linear regression. Statistics & Probability Letters, 78(13):1807–1810, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000436.

Aknouche:2009:SPP

[AG09a] Abdelhakim Aknouche and Hafida Guerbyenne. On some probabilistic properties of double periodic AR models. Statistics & Probability Letters, 79(3):407–413, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800429X.

Aknouche:2009:PSR

[AG09b] Abdelhakim Aknouche and Hafida Guerbyenne. Periodic stationarity of random coefficient periodic autoregressions. Statistics & Probability Letters, 79(7):990–996, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005658.

Abdous:2002:UTD

[AGG02] Belkacem Abdous, Stéphane Germain, and Nadia Ghazzali. A unified treatment of direct and indirect estimation of a probability density and its derivatives. Statistics & Probability Letters, 56(3):239–250, February 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100164X.

Antoch:2004:DSC

[AGJ04] Jaromír Antoch, Gérard Gregoire, and Daniela Jarusková. Detection of structural changes in generalized linear models. Statistics & Probability Letters, 69(3):315–332, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001828.

Agostinelli:2002:RMS

[Ago02] Claudio Agostinelli. Robust model selection in regression via weighted likelihood methodology. Statistics & Probability Letters, 56(3):289–300, February 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001936.

Agostinelli:2006:NPR

[Ago06] Claudio Agostinelli. Notes on Pearson residuals and weighted likelihood estimating equations. Statistics & Probability Letters, 76(17):1930–1934, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520600160X.

Antoch:2001:PTC

[AH01] Jaromír Antoch and Marie Husková. Permutation tests in change point analysis. Statistics & Probability Letters, 53 (1):37–46, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000098.

Aalen:2002:FMY

[AH02] Odd O. Aalen and Nils Lid Hjort. Frailty models that yield proportional hazards. Statistics & Probability Letters, 58(4):335–342, July 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000901.

Aue:2004:DTS

[AH04] Alexander Aue and Lajos Horváth. Delay time in sequential detection of change. Statistics & Probability Letters, 67(3):221–231, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000057.

Ai:2006:EMI

[AH06a] Mingyao Ai and Shuyuan He. An efficient method for identifying clear effects in blocked fractional factorial designs. Statistics & Probability Letters, 76(17):1889–1894, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001556.

Antonelli:2006:ESB

[AH06b] Fabio Antonelli and Said̈ Hamadène. Existence of the solutions of backward-forward SDE’s with continuous monotone coefficients. Statistics & Probability Letters, 76(14):1559–1569, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520600099X.

AL-Hussaini:2004:ADN

[AHEA04] Essam K. AL-Hussaini and Magdy E. El-Adll. Asymptotic distribution of normalized maximum under finite mixture models. Statistics & Probability Letters, 70(1):109–117, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204002329.

Ai:2008:OFP

[AHL08] Mingyao Ai, Fred J. Hickernell, and Dennis K. J. Lin. Optimal foldover plans for regular s-level fractional factorial designs. Statistics & Probability Letters, 78(7):896–903, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002970.

Ahmad:2002:MIS

[Ahm02] Ibrahim A. Ahmad. On moment inequalities of the supremum of empirical processes with applications to kernel estimation. Statistics & Probability Letters, 57(3):215–220, April 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202000299.

Ahmad:2004:SPC

[Ahm04] Ibrahim A. Ahmad. Some properties of classes of life distributions with unknown age. Statistics & Probability Letters, 69(3):333–342, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001841.

Aue:2007:RRA

[AHS07] Alexander Aue, Lajos Horváth, and Josef Steinebach. Rescaled range analysis in the presence of stochastic trend. Statistics & Probability Letters, 77(12):1165–1175, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000818.

Antonini:2006:CPS

[AHV06] Rita Giuliano Antonini, Tien-Chung Hu, and Andrei Volodin. On the concentration phenomenon for ϕ-subgaussian random elements. Statistics & Probability Letters, 76(5):465–469, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003263.

Aickin:2009:EDA

[Aic09] Mikel Aickin. Effect of design-adaptive allocation on inference for a regression parameter: Two-group, singlecovariate and double-covariate cases. Statistics & Probability Letters, 79(1):16–20, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002964.

Aggarwal:2000:SCB

[AJ00] M. L. Aggarwal and Mithilesh Kumar Jha. Some constructions for balanced n-ary residual treatment effects designs. Statistics & Probability Letters, 47(4):347–350, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001765.

Adell:2005:SEM

[AJ05] José A. Adell and Pedro Jodrá. Sharp estimates for the median of the Γ(n + 1, 1) distribution. Statistics & Probability Letters, 71(2):185–191, February 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002937.

Angelopoulos:2007:MEC

[AK07] P. Angelopoulos and C. Koukouvinos. Maximum estimation capacity projection designs from Hadamard matrices with 32, 36 and 40 runs. Statistics & Probability Letters, 77(2):220–229, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002276.

Archibald:2009:APM

[AK09] Margaret Archibald and Arnold Knopfmacher. The average position of the d th maximum in a sample of geometric random variables. Statistics & Probability Letters, 79(7):864–872, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005270.

Asmussen:2002:LLT

[AKK^{+}02] Søren Asmussen, Vladimir Kalashnikov, Dimitrios Konstantinides, Claudia Klüppelberg, and Gurami Tsitsiashvili. A local limit theorem for random walk maxima with heavy tails. Statistics & Probability Letters, 56(4):399–404, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202000330.

Anh:2000:SLF

[AL00] V. V. Anh and N. N. Leonenko. Scaling laws for fractional diffusion-wave equations with singular data. Statistics & Probability Letters, 48(3):239–252, July 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000031.

Adell:2006:ERV

[AL06] José A. Adell and Alberto Lekuona. Every random variable satisfies a certain nontrivial integrability condition. Statistics & Probability Letters, 76(15):1603–1606, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520600112X.

Anh:2008:LNL

[AL08] V. V. Anh and N. N. Leonenko. Log-normal, log-gamma and log-negative inverted gamma scenarios in multifractal products of stochastic processes. Statistics & Probability Letters, 78(11):1274–1282, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004051.

Albin:2008:CNB

[Alb08] J. M. P. Albin. A continuous non-Brownian motion martingale with Brownian motion marginal distributions. Statistics & Probability Letters, 78(6):682–686, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003124.

Aletti:2001:DTS

[Ale01] Giacomo Aletti. On different topologies for set-indexing collections. Statistics & Probability Letters, 54(1):67–73, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000621.

Aleskeviciene:2008:EML

[Ale08] A. K. Aleskeviciene. On the existence of moments of ladder heights. Statistics & Probability Letters, 78(3):282–285, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002398.

Aliev:2003:CUD

[Ali03] Fazil Alioglu Aliev. A comment on ‘unimodality of the distribution of record statistics’. Statistics & Probability Letters, 64(1):39–40, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001287.

Azzedine:2008:RNR

[ALOS08] Nadjia Azzedine, Ali Laksaci, and Elias Ould-Saïd. On robust nonparametric regression estimation for a functional regressor. Statistics & Probability Letters, 78(18):3216–3221, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003015.

Asci:2008:BHD

[ALP08] Claudio Asci, Gérard Letac, and Mauro Piccioni. Betahypergeometric distributions and random continued fractions. Statistics & Probability Letters, 78(13):1711–1721, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000308.

Alsmeyer:2001:TCT

[Als01] Gerold Alsmeyer. Two comparison theorems for nonlinear first passage times and their linear counterparts. Statistics & Probability Letters, 55(2):163–171, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001079.

Altman:2008:VCT

[Alt08] Rachel MacKay Altman. A variance component test for mixed hidden Markov models. Statistics & Probability Letters, 78(13):1885–1893, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000552.

Alzer:2006:CPM

[Alz06] Horst Alzer. A convexity property of the median of the gamma distribution. Statistics & Probability Letters, 76 (14):1510–1513, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000927.

Akita:2002:CSD

[AM02a] Koji Akita and Makoto Maejima. On certain self-decomposable self-similar processes with independent increments. Statistics & Probability Letters, 59(1):53–59, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001633.

Allam:2002:GAR

[AM02b] Abdelazziz Allam and Tahar Mourid. Geometric absolute regularity of Banach space-valued autoregressive processes. Statistics & Probability Letters, 60(3):241–252, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202002018.

Aggarwal:2005:OFF

[AM05] M. L. Aggarwal and Mukta Datta Mazumder. Optimal fractional factorial plans using minihypers. Statistics & Probability Letters, 75(4):291–297, December 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002439.

Agrawal:2007:SEE

[AM07] M. C. Agrawal and Chand K. Midha. Some efficient estimators of the domain parameters. Statistics & Probability Letters, 77(7):704–709, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206003397.

Ahmadi:2009:PIF

[AM09] J. Ahmadi and S. M. T. K. MirMostafaee. Prediction intervals for future records and order statistics coming from two parameter exponential distribution. Statistics & Probability Letters, 79(7):977–983, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005634.

Andersson:2000:SVS

[And00] Pontus Andersson. Small variance of subgraph counts in a random tournament. Statistics & Probability Letters, 49 (2):135–138, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000419.

Andersson:2005:LDT

[And05] E. Andersson. On-line detection of turning points using nonparametric surveillance: The effect of the growth after the turn. Statistics & Probability Letters, 73(4):433–439, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001471.

Annis:2007:NQL

[Ann07] David H. Annis. A note on quasi-likelihood for exponential families. Statistics & Probability Letters, 77(4):431– 437, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002628.

Anonymous:2000:Ia

[Ano00a] Anonymous. Index. Statistics & Probability Letters, 46 (4):421–422, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001972.

Anonymous:2000:Ib

[Ano00b] Anonymous. Index. Statistics & Probability Letters, 47(4): 421–422, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000328.

Anonymous:2000:Ic

[Ano00c] Anonymous. Index. Statistics & Probability Letters, 48(4): 421–422, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520000081X.

Anonymous:2000:Id

[Ano00d] Anonymous. Index. Statistics & Probability Letters, 49 (4):417–418, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001620.

Anonymous:2000:Ie

[Ano00e] Anonymous. Index. Statistics & Probability Letters, 50 (4):417–418, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001632.

Anonymous:2001:AA

[Ano01a] Anonymous. Author advert. Statistics & Probability Letters, 55(4):II, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201002103.

Anonymous:2001:AI

[Ano01b] Anonymous. Author index. Statistics & Probability Letters, 55(4):451–452, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201002073.

Anonymous:2001:AIVa

[Ano01c] Anonymous. Author index for volume 51 (2001). Statistics & Probability Letters, 51(4):433–434, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002467.

Anonymous:2001:AIVd

[Ano01d] Anonymous. Author index for volume 54 (2001). Statistics & Probability Letters, 54(4):451–452, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001547.

Anonymous:2001:AIVb

[Ano01e] Anonymous. Author index volume 52. Statistics & Probability Letters, 52(4):451–452, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000943.

Anonymous:2001:AIVc

[Ano01f] Anonymous. Author index volume 53 (2001). Statistics & Probability Letters, 53(4):449–450, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000967.

Anonymous:2001:CD

[Ano01g] Anonymous. Contents Direct. Statistics & Probability Letters, 54(1):EX2, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001006.

Anonymous:2001:SD

[Ano01h] Anonymous. Science Direct. Statistics & Probability Letters, 54(1):EX1, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000992.

Anonymous:2001:Sb

[Ano01i] Anonymous. SD0971. Statistics & Probability Letters, 55 (4):I, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000979.

Anonymous:2001:Sa

[Ano01j]  Anonymous.  sd0971+es0971.  Statistics & Probabil-

ity Letters, 55(2):I–II, November 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001705.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index. Statistics & Probability Letters, 56(4):447–448, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000688.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index. Statistics & Probability Letters, 57(4):409–410, May 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001670.

Anonymous:2002:AIc

[Ano02c] Anonymous. Author index. Statistics & Probability Letters, 58(4):435–436, July 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202002249.

Anonymous:2002:AId

[Ano02d] Anonymous. Author index. Statistics & Probability Letters, 59(4):437–438, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202002961.

Anonymous:2002:AIe

[Ano02e] Anonymous. Author index. Statistics & Probability Letters, 60(4):483–484, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003899.

Anonymous:2002:ZLM

[Ano02f] Anonymous. Zipf’s law and maximum sustainable growth. Statistics & Probability Letters, 56(2):??, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165188913000341.

Anonymous:2003:AIa

[Ano03a] Anonymous. Author index. Statistics & Probability Letters, 61(4):447–448, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000294.

Anonymous:2003:AIb

[Ano03b] Anonymous. Author index. Statistics & Probability Letters, 62(4):429–430, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001160.

Anonymous:2003:AIc

[Ano03c] Anonymous. Author index. Statistics & Probability Letters, 63(4):441–442, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001871.

Anonymous:2003:AId

[Ano03d] Anonymous. Author index. Statistics & Probability Letters, 64(4):443–444, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002451.

Anonymous:2003:AIe

[Ano03e] Anonymous. Author index. Statistics & Probability Letters, 65(4):467–468, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520300350X.

Anonymous:2003:EBa

[Ano03f] Anonymous. Editorial board. Statistics & Probability Letters, 62(1):ii, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203000725.

Anonymous:2003:EBb

[Ano03g] Anonymous. Editorial board. Statistics & Probability Letters, 62(2):ifc, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203000749.

Anonymous:2003:EBc

[Ano03h] Anonymous. Editorial board. Statistics & Probability Letters, 63(1):ii, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001378.

Anonymous:2003:EBd

[Ano03i] Anonymous. Editorial board. Statistics & Probability Letters, 64(1):ii, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001937.

Anonymous:2003:EBe

[Ano03j] Anonymous. Editorial board. Statistics & Probability Letters, 65(1):ii, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002761.

Anonymous:2003:IEBa

[Ano03k] Anonymous. IFC (editorial board). Statistics & Probability Letters, 63(2):ifc, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001573.

Anonymous:2003:IEBb

[Ano03l] Anonymous. IFC (editorial board). Statistics & Probability Letters, 63(3):ifc, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001639.

Anonymous:2003:IEBc

[Ano03m] Anonymous. IFC (editorial board). Statistics & Probability Letters, 63(4):ifc, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001846.

Anonymous:2003:IEBd

[Ano03n] Anonymous. IFC (editorial board). Statistics & Probability Letters, 64(2):ifc, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001998.

Anonymous:2003:IEBe

[Ano03o] Anonymous. IFC (editorial board). Statistics & Probability Letters, 64(3):ifc, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002323.

Anonymous:2003:IEBf

[Ano03p] Anonymous. IFC (editorial board). Statistics & Probability Letters, 64(4):ifc, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002426.

Anonymous:2003:IEBg

[Ano03q] Anonymous. IFC (editorial board). Statistics & Probability Letters, 65(2):ifc, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003134.

Anonymous:2003:IEBh

[Ano03r] Anonymous. IFC (editorial board). Statistics & Probability Letters, 65(3):ifc, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003213.

Anonymous:2003:IEBi

[Ano03s] Anonymous. IFC (editorial board). Statistics & Probability Letters, 65(4):ifc, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003432.

Anonymous:2003:IFC

[Ano03t] Anonymous. Inside front cover (editorial board). Statistics & Probability Letters, 62(4):ifc, May 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001135.

Anonymous:2003:PN

[Ano03u] Anonymous. Publisher’s note. Statistics & Probability Letters, 64(3):223, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002402.

Anonymous:2004:AIa

[Ano04a] Anonymous. Author index. Statistics & Probability Letters, 66(4):457–458, March 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000409.

Anonymous:2004:AIb

[Ano04b] Anonymous. Author index. Statistics & Probability Letters, 67(4):343–344, May 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000938.

Anonymous:2004:AIc

[Ano04c] Anonymous. Author index. Statistics & Probability Letters, 68(4):435–436, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520400152X.

Anonymous:2004:AId

[Ano04d] Anonymous. Author index. Statistics & Probability Letters, 69(4):477–478, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204002408.

Anonymous:2004:AIe

[Ano04e] Anonymous. Author index. Statistics & Probability Letters, 70(4):309–310, December 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003189.

Anonymous:2004:EBa

[Ano04f] Anonymous. Editorial board. Statistics & Probability Letters, 66(1):ii, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203003705.

Anonymous:2004:EBb

[Ano04g] Anonymous. Editorial board. Statistics & Probability Letters, 67(1):ii, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000471.

Anonymous:2004:EBc

[Ano04h] Anonymous. Editorial board. Statistics & Probability Letters, 68(1):ii, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001154.

Anonymous:2004:EBd

[Ano04i] Anonymous. Editorial board. Statistics & Probability Letters, 69(1):ii, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204002044.

Anonymous:2004:EBe

[Ano04j] Anonymous. Editorial board. Statistics & Probability Letters, 70(1):ii, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204002469.

Anonymous:2004:IEBa

[Ano04k] Anonymous. IFC (editorial board). Statistics & Probability Letters, 66(1):ifc, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003687.

Anonymous:2004:IEBb

[Ano04l] Anonymous. IFC (editorial board). Statistics & Probability Letters, 66(2):ifc, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003808.

Anonymous:2004:IEBc

[Ano04m] Anonymous. IFC (editorial board). Statistics & Probability Letters, 66(3):ifc, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000148.

Anonymous:2004:IEBd

[Ano04n] Anonymous. IFC (editorial board). Statistics & Probability Letters, 66(4):ifc, March 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000215.

Anonymous:2004:IEBe

[Ano04o] Anonymous. IFC (editorial board). Statistics & Probability Letters, 67(2):ifc, April 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000537.

Anonymous:2004:IEBf

[Ano04p] Anonymous. IFC (editorial board). Statistics & Probability Letters, 67(3):ifc, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000811.

Anonymous:2004:IEBg

[Ano04q] Anonymous. IFC (editorial board). Statistics & Probability Letters, 67(4):ifc, May 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000902.

Anonymous:2004:IEBh

[Ano04r] Anonymous. IFC (editorial board). Statistics & Probability Letters, 68(1):ifc, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001130.

Anonymous:2004:IEBi

[Ano04s] Anonymous. IFC (editorial board). Statistics & Probability Letters, 68(2):ifc, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520400135X.

Anonymous:2004:IEBj

[Ano04t] Anonymous. IFC (editorial board). Statistics & Probability Letters, 68(3):ifc, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001427.

Anonymous:2004:IEBk

[Ano04u] Anonymous. IFC (editorial board). Statistics & Probability Letters, 68(4):ifc, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520400149X.

Anonymous:2004:IEBl

[Ano04v] Anonymous. IFC (editorial board). Statistics & Probability Letters, 69(1):ifc, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204002020.

Anonymous:2004:IEBm

[Ano04w] Anonymous. IFC (editorial board). Statistics & Probability Letters, 69(2):ifc, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002111.

Anonymous:2004:IEBn

[Ano04x] Anonymous. IFC (editorial board). Statistics & Probability Letters, 69(3):ifc, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002184.

Anonymous:2004:IEBo

[Ano04y] Anonymous. IFC (editorial board). Statistics & Probability Letters, 69(4):ifc, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002378.

Anonymous:2004:IEBp

[Ano04z] Anonymous. IFC (editorial board). Statistics & Probability Letters, 70(1):ifc, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002445.

Anonymous:2004:IEBq

[Ano04-27] Anonymous. IFC (editorial board). Statistics & Probability Letters, 70(2):ifc, November 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002664.

Anonymous:2004:IEBr

[Ano04-28] Anonymous. IFC (editorial board). Statistics & Probability Letters, 70(3):CO2, December 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002962.

Anonymous:2004:IEBs

[Ano04-29] Anonymous. IFC (editorial board). Statistics & Probability Letters, 70(4):CO2, December 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003153.

Anonymous:2005:AIa

[Ano05a] Anonymous. Author index. Statistics & Probability Letters, 71(4):395–396, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520500060X.

Anonymous:2005:AIb

[Ano05b] Anonymous. Author index. Statistics & Probability Letters, 72(4):357–358, May 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001173.

Anonymous:2005:AIc

[Ano05c] Anonymous. Author index. Statistics & Probability Letters, 73(4):469–470, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002348.

Anonymous:2005:AId

[Ano05d] Anonymous. Author index. Statistics & Probability Letters, 74(4):379–380, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205003408.

Anonymous:2005:AIe

[Ano05e] Anonymous. Author index. Statistics & Probability Letters, 75(4):331–332, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004244.

Anonymous:2005:EBa

[Ano05f] Anonymous. Editorial board. Statistics & Probability Letters, 71(1):ii, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204003256.

Anonymous:2005:EBb

[Ano05g] Anonymous. Editorial board. Statistics & Probability Letters, 72(1):ii, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000672.

Anonymous:2005:EBc

[Ano05h] Anonymous. Editorial board. Statistics & Probability Letters, 73(1):ii, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001343.

Anonymous:2005:EBd

[Ano05i] Anonymous. Editorial board. Statistics & Probability Letters, 74(1):ii, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002555.

Anonymous:2005:EBe

[Ano05j] Anonymous. Editorial board. Statistics & Probability Letters, 75(1):ii, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205003561.

Anonymous:2005:EBf

[Ano05k] Anonymous. Editorial board. Statistics & Probability Letters, 75(4):CO2, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004219.

Anonymous:2005:IEBa

[Ano05l] Anonymous. IFC (editorial board). Statistics & Probability Letters, 71(1):CO2, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003232.

Anonymous:2005:IEBb

[Ano05m] Anonymous. IFC (editorial board). Statistics & Probability Letters, 71(2):CO2, February 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520500043X.

Anonymous:2005:IEBc

[Ano05n] Anonymous. IFC (editorial board). Statistics & Probability Letters, 71(3):CO2, March 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000507.

Anonymous:2005:IEBd

[Ano05o] Anonymous. IFC (editorial board). Statistics & Probability Letters, 71(4):CO2, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520500057X.

Anonymous:2005:IEBe

[Ano05p] Anonymous. IFC (editorial board). Statistics & Probability Letters, 72(1):CO2, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000659.

Anonymous:2005:IEBf

[Ano05q] Anonymous. IFC (editorial board). Statistics & Probability Letters, 72(2):CO2, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000878.

Anonymous:2005:IEBg

[Ano05r] Anonymous. IFC (editorial board). Statistics & Probability Letters, 72(3):CO2, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000933.

Anonymous:2005:IEBh

[Ano05s] Anonymous. IFC (editorial board). Statistics & Probability Letters, 72(4):CO2, May 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001148.

Anonymous:2005:IEBi

[Ano05t] Anonymous. IFC (editorial board). Statistics & Probability Letters, 73(1):CO2, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520500132X.

Anonymous:2005:IEBj

[Ano05u] Anonymous. IFC (editorial board). Statistics & Probability Letters, 73(2):CO2, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520500177X.

Anonymous:2005:IEBk

[Ano05v] Anonymous. IFC (editorial board). Statistics & Probability Letters, 73(3):CO2, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001999.

Anonymous:2005:IEBl

[Ano05w] Anonymous. IFC (editorial board). Statistics & Probability Letters, 73(4):CO2, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002312.

Anonymous:2005:IEBm

[Ano05x] Anonymous. IFC (editorial board). Statistics & Probability Letters, 74(2):CO2, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002762.

Anonymous:2005:IEBn

[Ano05y] Anonymous. IFC (editorial board). Statistics & Probability Letters, 74(3):CO2, October 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003196.

Anonymous:2005:IEBo

[Ano05z] Anonymous. IFC (editorial board). Statistics & Probability Letters, 74(4):CO2, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003378.

Anonymous:2005:IEBp

[Ano05-27] Anonymous. IFC (editorial board). Statistics & Probability Letters, 75(1):CO2, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003548.

Anonymous:2005:IEBq

[Ano05-28] Anonymous. IFC (editorial board). Statistics & Probability Letters, 75(2):CO2, November 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003718.

Anonymous:2005:IEBr

[Ano05-29] Anonymous. IFC (editorial board). Statistics & Probability Letters, 75(3):CO2, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003950.

Anonymous:2006:EBa

[Ano06a] Anonymous. Editorial board. Statistics & Probability Letters, 76(1):ii, January 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004323.

Anonymous:2006:EBb

[Ano06b] Anonymous. Editorial board. Statistics & Probability Letters, 76(2):CO2, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004591.

Anonymous:2006:EBc

[Ano06c] Anonymous. Editorial board. Statistics & Probability Letters, 76(4):CO2, February 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000034.

Anonymous:2006:EBd

[Ano06d] Anonymous. Editorial board. Statistics & Probability Letters, 76(5):CO2, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000125.

Anonymous:2006:EBe

[Ano06e] Anonymous. Editorial board. Statistics & Probability Letters, 76(8):CO2, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000630.

Anonymous:2006:EBf

[Ano06f] Anonymous. Editorial board. Statistics & Probability Letters, 76(9):CO2, May 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520600071X.

Anonymous:2006:EBg

[Ano06g] Anonymous. Editorial board. Statistics & Probability Letters, 76(10):CO2, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000873.

Anonymous:2006:EBh

[Ano06h] Anonymous. Editorial board. Statistics & Probability Letters, 76(11):CO2, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001040.

Anonymous:2006:EBi

[Ano06i] Anonymous. Editorial board. Statistics & Probability Letters, 76(12):CO2, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001325.

Anonymous:2006:EBj

[Ano06j] Anonymous. Editorial board. Statistics & Probability Letters, 76(13):CO2, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001684.

Anonymous:2006:EBk

[Ano06k] Anonymous. Editorial board. Statistics & Probability Letters, 76(15):CO2, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002057.

Anonymous:2006:EBl

[Ano06l] Anonymous. Editorial board. Statistics & Probability Letters, 76(16):CO2, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520600246X.

Anonymous:2006:EBm

[Ano06m] Anonymous. Editorial board. Statistics & Probability Letters, 76(17):CO2, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520600277X.

Anonymous:2006:EBn

[Ano06n] Anonymous. Editorial board. Statistics & Probability Letters, 76(18):CO2, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206003038.

Anonymous:2006:EPF

[Ano06o] Anonymous. Erratum to “Product formula and independence criterion for multiple Huang–Cambanis integrals” [Stat. Prob. Lett. 76 (2006) 1255–1260]. Statistics & Probability Letters, 76(18):2036, December 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002240. See [CA06a].

Anonymous:2006:IEBa

[Ano06p] Anonymous. IFC (editorial board). Statistics & Probability Letters, 76(3):CO2, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004815.

Anonymous:2006:IEBb

[Ano06q] Anonymous. IFC (editorial board). Statistics & Probability Letters, 76(6):CO2, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000319.

Anonymous:2007:C

[Ano07a] Anonymous. Contents. Statistics & Probability Letters, 77(12):xxi–xxii, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001757.

Anonymous:2007:EBa

[Ano07b] Anonymous. Editorial board. Statistics & Probability Letters, 77(1):ii, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206003129.

Anonymous:2007:EBb

[Ano07c] Anonymous. Editorial board. Statistics & Probability Letters, 77(2):CO2, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520600318X.

Anonymous:2007:EBc

[Ano07d] Anonymous. Editorial board. Statistics & Probability Letters, 77(3):CO2, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206003245.

Anonymous:2007:EBd

[Ano07e] Anonymous. Editorial board. Statistics & Probability Letters, 77(4):CO2, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206003300.

Anonymous:2007:EBe

[Ano07f] Anonymous. Editorial board. Statistics & Probability Letters, 77(5):CO2, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000090.

Anonymous:2007:EBf

[Ano07g] Anonymous. Editorial board. Statistics & Probability Letters, 77(6):CO2, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520700051X.

Anonymous:2007:EBh

[Ano07h] Anonymous. Editorial board. Statistics & Probability Letters, 77(7):ifc, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000570.

Anonymous:2007:EBi

[Ano07i] Anonymous. Editorial board. Statistics & Probability Letters, 77(8):ifc, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001095.

Anonymous:2007:EBj

[Ano07j] Anonymous. Editorial board. Statistics & Probability Letters, 77(10):ifc, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001381.

Anonymous:2007:EBk

[Ano07k] Anonymous. Editorial board. Statistics & Probability Letters, 77(11):ifc, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001691.

Anonymous:2007:EBl

[Ano07l] Anonymous. Editorial board. Statistics & Probability Letters, 77(12):ifc, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001770.

Anonymous:2007:EBm

[Ano07m] Anonymous. Editorial board. Statistics & Probability Letters, 77(13):ii, July 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002271.

Anonymous:2007:EBn

[Ano07n] Anonymous. Editorial board. Statistics & Probability Letters, 77(13):ifc, July 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002258.

Anonymous:2007:EBo

[Ano07o] Anonymous. Editorial board. Statistics & Probability Letters, 77(14):ifc, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002337.

Anonymous:2007:EBp

[Ano07p] Anonymous. Editorial board. Statistics & Probability Letters, 77(15):ifc, September 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002672.

Anonymous:2007:EBq

[Ano07q] Anonymous. Editorial board. Statistics & Probability Letters, 77(16):ifc, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520700274X.

Anonymous:2007:EBr

[Ano07r] Anonymous. Editorial board. Statistics & Probability Letters, 77(17):ifc, November 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520700315X.

Anonymous:2007:EBs

[Ano07s] Anonymous. Editorial board. Statistics & Probability Letters, 77(18):ifc, December 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003537.

Anonymous:2007:PA

[Ano07t] Anonymous. Publisher’s announcement. Statistics & Probability Letters, 77(13):v, July 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001745.

Anonymous:2007:RJB

[Ano07u] Anonymous. R. A. Johnson: a brief biography. Statistics & Probability Letters, 77(12):ix–xix, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001071.

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial board. Statistics & Probability Letters, 78(1):ii, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003732.

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial board. Statistics & Probability Letters, 78(1):CO2, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003719.

Anonymous:2008:EBc

[Ano08c] Anonymous. Editorial board. Statistics & Probability Letters, 78(2):CO2, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207004336.

Anonymous:2008:EBd

[Ano08d] Anonymous. Editorial board. Statistics & Probability Letters, 78(3):ifc, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207004397.

Anonymous:2008:EBe

[Ano08e] Anonymous. Editorial board. Statistics & Probability Letters, 78(4):ifc, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000928.

Anonymous:2008:EBf

[Ano08f] Anonymous. Editorial board. Statistics & Probability Letters, 78(5):ifc, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001004.

Anonymous:2008:EBg

[Ano08g] Anonymous. Editorial board. Statistics & Probability Letters, 78(6):ifc, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001612.

Anonymous:2008:EBh

[Ano08h] Anonymous. Editorial board. Statistics & Probability Letters, 78(7):ifc, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208002125.

Anonymous:2008:EBi

[Ano08i] Anonymous. Editorial board. Statistics & Probability Letters, 78(8):ifc, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208002241.

Anonymous:2008:EBj

[Ano08j] Anonymous. Editorial board. Statistics & Probability Letters, 78(9):ifc, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208002435.

Anonymous:2008:EBk

[Ano08k] Anonymous. Editorial board. Statistics & Probability Letters, 78(10):ifc, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208002903.

Anonymous:2008:EBl

[Ano08l] Anonymous. Editorial board. Statistics & Probability Letters, 78(11):ifc, August 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208003374.

Anonymous:2008:EBm

[Ano08m] Anonymous. Editorial board. Statistics & Probability Letters, 78(12):ifc, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208003763.

Anonymous:2008:EBn

[Ano08n] Anonymous. Editorial board. Statistics & Probability Letters, 78(13):ifc, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800391X.

Anonymous:2008:EBo

[Ano08o] Anonymous. Editorial board. Statistics & Probability Letters, 78(14):ifc, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004082.

Anonymous:2008:EBp

[Ano08p] Anonymous. Editorial board. Statistics & Probability Letters, 78(15):ifc, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004318.

Anonymous:2008:EBq

[Ano08q] Anonymous. Editorial board. Statistics & Probability Letters, 78(16):ifc, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004434.

Anonymous:2008:EBr

[Ano08r] Anonymous. Editorial board. Statistics & Probability Letters, 78(17):ifc, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004744.

Anonymous:2008:EBs

[Ano08s] Anonymous. Editorial board. Statistics & Probability Letters, 78(18):ifc, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004811.

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial board. Statistics & Probability Letters, 79(1):ii, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005385.

Anonymous:2009:EBb

[Ano09b] Anonymous. Editorial board. Statistics & Probability Letters, 79(1):ifc, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005361.

Anonymous:2009:EBc

[Ano09c] Anonymous. Editorial board. Statistics & Probability Letters, 79(2):ifc, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005555.

Anonymous:2009:EBd

[Ano09d] Anonymous. Editorial board. Statistics & Probability Letters, 79(3):ifc, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000145.

Anonymous:2009:EBe

[Ano09e] Anonymous. Editorial board. Statistics & Probability Letters, 79(4):ifc, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000273.

Anonymous:2009:EBf

[Ano09f] Anonymous. Editorial board. Statistics & Probability Letters, 79(5):ifc, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000479.

Anonymous:2009:EBg

[Ano09g] Anonymous. Editorial board. Statistics & Probability Letters, 79(6):ifc, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000613.

Anonymous:2009:EBh

[Ano09h] Anonymous. Editorial board. Statistics & Probability Letters, 79(7):ifc, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000807.

Anonymous:2009:EBi

[Ano09i] Anonymous. Editorial board. Statistics & Probability Letters, 79(8):ifc, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000881.

Anonymous:2009:EBj

[Ano09j] Anonymous. Editorial board. Statistics & Probability Letters, 79(9):ifc, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001308.

Anonymous:2009:EBk

[Ano09k] Anonymous. Editorial board. Statistics & Probability Letters, 79(10):ifc, May 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001424.

Anonymous:2009:EBl

[Ano09l] Anonymous. Editorial board. Statistics & Probability Letters, 79(11):ifc, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001680.

Anonymous:2009:EBm

[Ano09m] Anonymous. Editorial board. Statistics & Probability Letters, 79(12):ifc, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520900176X.

Anonymous:2009:EBn

[Ano09n] Anonymous. Editorial board. Statistics & Probability Letters, 79(13):ifc, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002053.

Anonymous:2009:EBo

[Ano09o] Anonymous. Editorial board. Statistics & Probability Letters, 79(14):ifc, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002193.

Anonymous:2009:EBp

[Ano09p] Anonymous. Editorial board. Statistics & Probability Letters, 79(15):ifc, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002260.

Anonymous:2009:EBq

[Ano09q] Anonymous. Editorial board. Statistics & Probability Letters, 79(16):ifc, August 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002557.

Anonymous:2009:EBr

[Ano09r] Anonymous. Editorial board. Statistics & Probability Letters, 79(17):ifc, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002788.

Anonymous:2009:EBs

[Ano09s] Anonymous. Editorial board. Statistics & Probability Letters, 79(18):ifc, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002958.

Anonymous:2009:EBt

[Ano09t] Anonymous. Editorial board. Statistics & Probability Letters, 79(19):ifc, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003137.

Anonymous:2009:EBu

[Ano09u] Anonymous. Editorial board. Statistics & Probability Letters, 79(20):ifc, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003368.

Anonymous:2009:EBv

[Ano09v] Anonymous. Editorial board. Statistics & Probability Letters, 79(21):ifc, November 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003605.

Anonymous:2009:EBw

[Ano09w] Anonymous. Editorial board. Statistics & Probability Letters, 79(22):ifc, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003745.

Anonymous:2009:EBx

[Ano09x] Anonymous. Editorial board. Statistics & Probability Letters, 79(23):ifc, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003824.

Anonymous:2009:EBy

[Ano09y] Anonymous. Editorial board. Statistics & Probability Letters, 79(24):ifc, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209004003.

Amendola:2006:MSM

[ANV06] Alessandra Amendola, Marcella Niglio, and Cosimo Vitale. The moments of SETARMA models. Statistics & Probability Letters, 76(6):625–633, March 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003615.

Alvo:2002:MNP

[AP02a] Mayer Alvo and Jincheol Park. Multivariate non-parametric tests of trend when the data are incomplete. Statistics & Probability Letters, 57(3):281–290, April 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000627.

Aneiros-Perez:2002:BSP

[AP02b] Germán Aneiros-Pérez. On bandwidth selection in partial linear regression models under dependence. Statistics & Probability Letters, 57(4):393–401, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000962.

Arakelian:2004:BAM

[AP04] V. Arakelian and V. Papathanasiou. On bounding the absolute mean value. Statistics & Probability Letters, 69(4): 447–450, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001981.

Alonso:2003:SBP

[APR03] Andrés M. Alonso, Daniel Peña, and Juan Romo. On sieve bootstrap prediction intervals. Statistics & Probability Letters, 65(1):13–20, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002141.

Abe:2009:PEC

[APS09a] Toshihiro Abe, Arthur Pewsey, and Kunio Shimizu. On Papakonstantinou’s extension of the cardioid distribution.

Statistics & Probability Letters, 79(20):2138–2147, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520900251X.

Antoniadis:2009:BSF

[APS09b] Anestis Antoniadis, Efstathios Paparoditis, and Theofanis Sapatinas. Bandwidth selection for functional time series prediction. Statistics & Probability Letters, 79(6):733–740, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005075.

Aneiros-Perez:2006:SFP

[APV06] Germán Aneiros-Pérez and Philippe Vieu. Semi-functional partial linear regression. Statistics & Probability Letters, 76 (11):1102–1110, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004530.

Appel:2002:CGU

[AR02] Martin J. B. Appel and Ralph P. Russo. The connectivity of a graph on uniform points on [0, 1]^{d}. Statistics & Probability Letters, 60(4):351–357, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200233X.

Alsmeyer:2003:BCT

[AR03a] Gerold Alsmeyer and Uwe Rösler. The best constant in the Topchii–Vatutin inequality for martingales. Statistics & Probability Letters, 65(3):199–206, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002529.

Appleby:2003:NES

[AR03b] John A. D. Appleby and David W. Reynolds. Non-exponential stability of scalar stochastic Volterra equations. Statistics & Probability Letters, 62(4):335–343, May 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300035X.

Antonelli:2005:PBR

[AR05] Sabrina Antonelli and Giuliana Regoli. On the Poissonbinomial relative error. Statistics & Probability Letters, 71 (3):249–256, March 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204003062.

Akbari:2009:OSU

[AR09a] M. GH. Akbari and A. H. Rezaei. Order statistics using fuzzy random variables. Statistics & Probability Letters, 79(8):1031–1037, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005725.

Antonelli:2009:ADD

[AR09b] Sabrina Antonelli and Giuliana Regoli. Asymptotic distribution of density ratios. Statistics & Probability Letters, 79 (3):289–294, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004008.

Asha:2009:CUG

[AR09c] G. Asha and C. John Rejeesh. Characterizations using the generalized reversed lack of memory property. Statistics & Probability Letters, 79(12):1480–1487, June 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001126.

Arai:2001:OMM

[Ara01] Takuji Arai. The p-optimal martingale measure in continuous trading models. Statistics & Probability Letters, 54(1):93–99, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000815.

Arai:2005:SPV

[Ara05] Takuji Arai. Some properties of the variance-optimal martingale measure for discontinuous semimartingales. Statistics & Probability Letters, 74(2):163–170, September 1, 2005.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001720.

Ahmadi:2009:CRT

[ARB09] J. Ahmadi, M. Razmkhah, and N. Balakrishnan. Current k-records and their use in distribution-free confidence intervals. Statistics & Probability Letters, 79(1):29–37, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003325.

Arcudi:2000:CPM

[Arc00] Ornella Arcudi. On convergences of probability measures in different Prohorov-type metrics. Statistics & Probability Letters, 46(4):365–369, February 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001236.

Arcones:2002:LMD

[Arc02] Miguel A. Arcones. Large and moderate deviations of empirical processes with nonstandard rates. Statistics & Probability Letters, 57(4):315–326, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000822.

Arevalillo:2003:ISA

[Are03] Jorge M. Arevalillo. Inverting a saddlepoint approximation. Statistics & Probability Letters, 61(4):421–428, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202004029.

Asmussen:2008:ASL

[ARN08] Søren Asmussen and Leonardo Rojas-Nandayapa. Asymptotics of sums of lognormal random variables with Gaussian copula. Statistics & Probability Letters, 78(16):2709–2714, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001892.

Arslan:2005:NCM

[Ars05] Olcay Arslan. A new class of multivariate distributions: Scale mixture of Kotz-type distributions. Statistics & Probability Letters, 75(1):18–28, November 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002063.

Arslan:2008:AMS

[Ars08] Olcay Arslan. An alternative multivariate skew-slash distribution. Statistics & Probability Letters, 78(16):2756–2761, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001971.

Arslan:2009:MLP

[Ars09] Olcay Arslan. Maximum likelihood parameter estimation for the multivariate skew-slash distribution. Statistics & Probability Letters, 79(20):2158–2165, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002533.

Athreya:2001:ADO

[AS01] J. S. Athreya and S. Sethuraman. On the asymptotics of discrete order statistics. Statistics & Probability Letters, 54(3):243–249, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000426.

Aue:2002:NEC

[AS02] Alexander Aue and Josef Steinebach. A note on estimating the change-point of a gradually changing stochastic process. Statistics & Probability Letters, 56(2):177–191, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001845.

Arnab:2006:NME

[AS06] Raghunath Arnab and Sarjinder Singh. A new method for estimating variance from data imputed with ratio method of imputation. Statistics & Probability Letters, 76(5):513–519, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003330.

Aletti:2007:RSA

[AS07] Giacomo Aletti and Diane Saada. Randomization in survival analysis. Statistics & Probability Letters, 77(5):497–502, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002707.

Al-Saleh:2000:DRS

[ASAK00] M. Fraiwan Al-Saleh and M. Ali Al-Kadiri. Double-ranked set sampling. Statistics & Probability Letters, 48(2):205–212, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299002060.

Asci:2009:ABA

[Asc09] Claudio Asci. Asymptotic behavior of an affine random recursion in Z_{p}^{k} defined by a matrix with an eigenvalue of size 1. Statistics & Probability Letters, 79(11):1421–1428, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001035.

Aggarwal:2004:OBD

[ASG04] M. L. Aggarwal, Poonam Singh, and Nidhi Gupta. Orthogonal block designs in two blocks for second degree Kmodel. Statistics & Probability Letters, 66(4):423–434, March 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003602.

Aggarwal:2002:OOD

[ASS02] M. L. Aggarwal, V. Sarin, and Poonam Singh. Optimal orthogonal designs in two blocks for Becker’s mixture models in three and four components. Statistics & Probability Letters, 59(4):385–396, October 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002377.

Astrauskas:2001:HLE

[Ast01] Arvydas Astrauskas. On high-level exceedances of i.i.d. random fields. Statistics & Probability Letters, 52(3):271–277, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520000211X.

Akkaya:2008:EPA

[AT08a] Aysen D. Akkaya and Moti L. Tiku. Estimating parameters in autoregressive models with asymmetric innovations. Statistics & Probability Letters, 78(17):2916, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002368.

Amano:2008:AEC

[AT08b] Tomoyuki Amano and Masanobu Taniguchi. Asymptotic efficiency of conditional least squares estimators for ARCH models. Statistics & Probability Letters, 78(2):179–185, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002118.

Arashi:2008:STI

[AT08c] M. Arashi and S. M. M. Tabatabaey. Stein-type improvement under stochastic constraints: Use of multivariate Student-t model in regression. Statistics & Probability Letters, 78(14):2142–2153, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000898.

Athreya:2008:GRP

[Ath08] K. B. Athreya. Growth rates for pure birth Markov chains. Statistics & Probability Letters, 78(12):1534–1540, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000084.

Atherton:2009:NSI

[Ath09] Juli Atherton. A new simple interpretation of an optimal design criterion. Statistics & Probability Letters, 79(6):707–710, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005014.

Aue:2004:SAR

[Aue04] Alexander Aue. Strong approximation for RCA(1) time series with applications. Statistics & Probability Letters, 68 (4):369–382, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001269.

Aurzada:2008:SNS

[Aur08a] Frank Aurzada. A short note on small deviations of sequences of i.i.d. random variables with exponentially decreasing weights. Statistics & Probability Letters, 78(15):2300–2307, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001247.

Aurzada:2008:SDS

[Aur08b] Frank Aurzada. Small deviations for stable processes via compactness properties of the parameter set. Statistics & Probability Letters, 78(6):577–581, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002982.

Adimari:2001:RIG

[AV01a] Gianfranco Adimari and Laura Ventura. Robust inference for generalized linear models with application to logistic regression. Statistics & Probability Letters, 55(4):413–419, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001572.

Arellano-Valle:2001:SCS

[AV01b] Reinaldo B. Arellano-Valle. On some characterizations of spherical distributions. Statistics & Probability Letters, 54 (3):227–232, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001917.

Arellano-Valle:2002:DPP

[AVdPS02] R. B. Arellano-Valle, G. del Pino, and E. San Martín. Definition and probabilistic properties of skew-distributions. Statistics & Probability Letters, 58(2):111–121, June 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000883.

Arellano-Valle:2008:EDM

[AVG08] Reinaldo B. Arellano-Valle and Marc G. Genton. On the exact distribution of the maximum of absolutely continuous dependent random variables. Statistics & Probability Letters, 78(1):27–35, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001678.

Ahmed:2009:HOA

[AVV09] S. Ejaz Ahmed, Andrei I. Volodin, and Igor N. Volodin. High order approximation for the coverage probability by a confident set centered at the positive-part James–Stein estimator. Statistics & Probability Letters, 79(17):1823–1828, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001874.

Amirdjanova:2004:SEC

[AW04] Anna Amirdjanova and Michael Woodroofe. Shrinkage estimation for convex polyhedral cones. Statistics & Probability Letters, 70(1):87–94, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002305.

Arcones:2006:SNT

[AW06] Miguel A. Arcones and Yishi Wang. Some new tests for normality based on U-processes. Statistics & Probability Letters, 76(1):69–82, January 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002683.

Abu-Youssef:2002:MID

[AY02] S. E. Abu-Youssef. A moment inequality for decreasing (increasing) mean residual life distributions with hypothesis testing application. Statistics & Probability Letters, 57(2):171– 177, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000457.

Ai:2004:DCC

[AZ04a] Mingyao Ai and Runchu Zhang. s^{n−m} Designs containing clear main effects or clear two-factor interactions. Statistics & Probability Letters, 69(2):151–160, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001701.

Ai:2004:MFF

[AZ04b] Mingyao Ai and Runchu Zhang. Multistratum fractional factorial split-plot designs with minimum aberration and maximum estimation capacity. Statistics & Probability Letters, 69(2):161–170, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001713.

Amini:2007:SSL

[AZB07] M. Amini, H. Zarei, and A. Bozorgnia. Some strong limit theorems of weighted sums for negatively dependent generalized Gaussian random variables. Statistics & Probability Letters, 77(11):1106–1110, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000491.

Al-Zahrani:2008:MID

[AZS08] Bander Al-Zahrani and Jordan Stoyanov. Moment inequalities for DVRL distributions, characterization and testing for exponentiality. Statistics & Probability Letters, 78(13):1792–1799, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000412.

Balakrishnan:2003:CLR

[BA03] N. Balakrishnan and I. S. Akhundov. A characterization by linearity of the regression function based on order statistics. Statistics & Probability Letters, 63(4):435–440, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001238.

Bairamov:2008:RLR

[BA08] Ismihan Bairamov and Barry C. Arnold. On the residual lifelengths of the remaining components in an n − k + 1 out of n system. Statistics & Probability Letters, 78(8):945–952, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700346X.

Baillo:2003:TEP

[Baí03] Amparo Baíllo. Total error in a plug-in estimator of level sets. Statistics & Probability Letters, 65(4):411–417, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002943.

Balan:2001:SMP

[Bal01] R. M. Balan. A strong Markov property for set-indexed processes. Statistics & Probability Letters, 53(2):219–226, June 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000918.

Balan:2002:SIP

[Bal02] R. M. Balan. Set-indexed processes with independent increments. Statistics & Probability Letters, 59(4):415–424, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002419.

Banerjee:2008:DFG

[Ban08] Shuvadeep Banerjee. A distribution free goodness of fit test for a stochastically ordered alternative. Statistics & Probability Letters, 78(17):2868–2875, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002228.

Bartoszewicz:2000:SOB

[Bar00] Jarosław Bartoszewicz. Stochastic orders based on the Laplace transform and infinitely divisible distributions. Statistics & Probability Letters, 50(2):121–129, November 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000894.

Bartoszewicz:2001:SCR

[Bar01] Jarosław Bartoszewicz. Stochastic comparisons of random minima and maxima from life distributions. Statistics & Probability Letters, 55(1):107–112, November 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001390.

Barakat:2002:CRB

[Bar02a] H. M. Barakat. On the confidence region for the bivariate coordinate-wise quantiles of the continuous bivariate distribution functions. Statistics & Probability Letters, 56(1):37–43, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001456.

Bartoszewicz:2002:MED

[Bar02b] Jarosław Bartoszewicz. Mixtures of exponential distributions and stochastic orders. Statistics & Probability Letters, 57(1):23–31, March 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202000342.

Barry:2005:LAV

[Bar05] Ronald Paul Barry. The local approximation of variograms in R^{∈} by covariance functions. Statistics & Probability Letters, 74(2):171–177, September 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001732.

Barakat:2006:CRC

[Bar06] H. M. Barakat. Comments on the rate of convergence to asymptotic independence between order statistics. Statistics & Probability Letters, 76(1):35–38, January 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002658.

Bartolucci:2007:PVE

[Bar07] Francesco Bartolucci. A penalized version of the empirical likelihood ratio for the population mean. Statistics & Probability Letters, 77(1):104–110, January 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002021.

Bartoszewicz:2009:RWD

[Bar09] Jarosław Bartoszewicz. On a representation of weighted distributions. Statistics & Probability Letters, 79(15):1690–1694, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001564.

Basu:2000:USP

[Bas00] Sanjib Basu. Uniform stability of posteriors. Statistics & Probability Letters, 46(1):53–58, January 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000863.

Battaglia:2005:OFA

[Bat05] Francesco Battaglia. Outliers in functional autoregressive time series. Statistics & Probability Letters, 72(4):323–332, May 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000738.

Basak:2002:UDR

[BB02a] Prasanta Basak and Indrani Basak. Unimodality of the distribution of record statistics. Statistics & Probability Letters, 56(4):395–398, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000287.

Bianco:2002:ABO

[BB02b] Ana Bianco and Graciela Boente. On the asymptotic behavior of one-step estimates in heteroscedastic regression models. Statistics & Probability Letters, 60(1):33–47, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202002444.

Bertoli-Barsotti:2003:OPP

[BB03] Lucio Bertoli-Barsotti. An order-preserving property of the maximum likelihood estimates for the Rasch model. Statistics & Probability Letters, 61(1):91–96, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003516.

Barbe:2004:BND

[BB04] Ph. Barbe and M. Broniatowski. Blowing number of a distribution for a statistics and loyal estimators. Statistics & Probability Letters, 69(4):465–475, October 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002007.

Bhandari:2005:GCI

[BB05] Subir K. Bhandari and Ayanendranath Basu. On Gaussian correlation inequalities for “periodic” sets. Statistics & Probability Letters, 73(3):315–320, July 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001252.

Bailey:2007:ODO

[BB07a] Ralph W. Bailey and Peter Burridge. Ordering the dispersion of ordinary least squares under near-integration. Statistics & Probability Letters, 77(6):594–597, March 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002902.

Bellini:2007:SDP

[BB07b] Fabio Bellini and Leonardo Bottolo. Stationarity domains for δ-power Garch process with heavy tails. Statistics & Probability Letters, 77(13):1418–1427, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000727.

Bose:2007:OME

[BB07c] Mausumi Bose and Sunanda Bagchi. Optimal main effect plans in blocks of small size. Statistics & Probability Letters, 77(2):142–147, January 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002136.

Bagchi:2008:OME

[BB08a] Sunanda Bagchi and Mausumi Bose. Optimal main effect plans in nested row-column set-up of small size. Statistics & Probability Letters, 78(16):2720–2724, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001910.

Balakrishnan:2008:RSI

[BB08b] N. Balakrishnan and K. Balasubramanian. Revisiting Sen’s inequalities on order statistics. Statistics & Probability Letters, 78(6):616–621, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003033.

Babenko:2009:PPC

[BB09a] Alexandra Babenko and Eduard Belitser. On the posterior pointwise convergence rate of a Gaussian signal under a conjugate prior. Statistics & Probability Letters, 79(5):670–675, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004963.

Bali:2009:PPE

[BB09b] Juan Lucas Bali and Graciela Boente. Principal points and elliptical distributions from the multivariate setting to the functional case. Statistics & Probability Letters, 79(17):1858–1865, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001928.

Bezandry:2005:CED

[BBG05] Paul H. Bezandry, George E. Bonney, and Ali Gannoun. Consistent estimation of the density and hazard rate functions for censored data via the wavelet method. Statistics & Probability Letters, 74(4):366–372, October 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001963.

Baltagi:2009:TFE

[BBP09] Badi H. Baltagi, Georges Bresson, and Alain Pirotte. Testing the fixed effects restrictions? a Monte Carlo study of Chamberlain’s minimum chi-squared test. Statistics & Probability Letters, 79(10):1358–1362, May 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000960.

Berlinet:2005:OBS

[BBR05] Alain Berlinet, Gérard Biau, and Laurent Rouvière. Optimal L_{1} bandwidth selection for variable kernel density estimates. Statistics & Probability Letters, 74(2):116–128, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001689.

Bassetti:2006:MKD

[BBR06] Federico Bassetti, Antonella Bodini, and Eugenio Regazzini. On minimum Kantorovich distance estimators. Statistics & Probability Letters, 76(12):1298–1302, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000381.

Bassetti:2007:CMD

[BBR07] Federico Bassetti, Antonella Bodini, and Eugenio Regazzini. Consistency of minimum divergence estimators based on grouped data. Statistics & Probability Letters, 77(10):937–941, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000053.

Basawa:2008:PEU

[BBZ08] I. V. Basawa, U. N. Bhat, and J. Zhou. Parameter estimation using partial information with applications to queueing and related models. Statistics & Probability Letters, 78(12):1375–1383, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520700418X.

Bai:2000:MSL

[BC00a] Z. D. Bai and Philip E. Cheng. Marcinkiewicz strong laws for linear statistics. Statistics & Probability Letters, 46 (2):105–112, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000930.

Bai:2000:NSE

[BC00b] Z. D. Bai and Mosuk Chow. A note on sequential estimation of the size of a population under a general loss function. Statistics & Probability Letters, 47(2):159–164, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001522.

Bose:2001:GBN

[BC01] Arup Bose and Snigdhansu Chatterjee. Generalised bootstrap in non-regular M-estimation problems. Statistics & Probability Letters, 55(3):319–328, December 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001614.

Biernacki:2003:DML

[BC03] Christophe Biernacki and Stéphane Chrétien. Degeneracy in the maximum likelihood estimation of univariate Gaussian mixtures with EM. Statistics & Probability Letters, 61(4):373–382, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003966.

Barroso:2005:BCH

[BC05] Lúcia P. Barroso and Gauss M. Cordeiro. Bartlett corrections in heteroskedastic t regression models. Statistics & Probability Letters, 75(2):86–96, November 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002142.

Babu:2006:SED

[BC06] G. Jogesh Babu and Yogendra P. Chaubey. Smooth estimation of a distribution and density function on a hypercube using Bernstein polynomials for dependent random vectors. Statistics & Probability Letters, 76(9):959–969, May 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004153.

Boland:2007:OTS

[BC07] Philip J. Boland and Nóra Ní Chuív. Optimal times for software release when repair is imperfect. Statistics & Probability Letters, 77(12):1176–1184, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700082X.

Baillo:2001:CRN

[BCAC01] Amparo Baíllo, Juan A. Cuesta-Albertos, and Antonio Cuevas. Convergence rates in nonparametric estimation of level sets. Statistics & Probability Letters, 53(1):27–35, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000062.

Balakrishnan:2002:ECM

[BCC02] N. Balakrishnan, A. Childs, and B. Chandrasekar. An efficient computational method for moments of order statistics under progressive censoring. Statistics & Probability Letters, 60(4):359–365, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202002675.

Balakrishnan:2009:SRO

[BCD09] N. Balakrishnan, E. Cramer, and K. Davies. Some results on order statistics generated by two simulation methods. Statistics & Probability Letters, 79(17):1847–1857, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001916.

Bravata:2001:MCP

[BCEO01] D. M. Bravata, R. W. Cottle, B. C. Eaves, and I. Olkin. Measuring conformability of probabilities. Statistics & Probability Letters, 52(3):321–327, April 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002376.

Berlinet:2001:ECM

[BCG01] Alain Berlinet, Benoît Cadre, and Ali Gannoun. Estimation of conditional L_{1}-median from dependent observations. Statistics & Probability Letters, 55(4):353–358, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000463.

Balakrishnan:2001:BMV

[BCK01] N. Balakrishnan, E. Cramer, and U. Kamps. Bounds for means and variances of progressive type II censored order statistics. Statistics & Probability Letters, 54(3):301–315, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001043.

Burkschat:2006:OSP

[BCK06] M. Burkschat, E. Cramer, and U. Kamps. On optimal schemes in progressive censoring. Statistics & Probability Letters, 76(10):1032–1036, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500444X.

Barranco-Chamorro:2000:LBV

[BCLBMR00] I. Barranco-Chamorro, F. López-Blázquez, and J. L. Moreno-Rebollo. Lower bounds for the variance in certain nonregular distributions. Statistics & Probability Letters, 50(1):75–81, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000845.

Berenhaut:2008:CDC

[BCT08] Kenneth S. Berenhaut, Donghui Chen, and Vy Tran. On the compatibility of Dyson’s conditions. Statistics & Probability Letters, 78(17):3110–3113, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002782.

Balakrishnan:2001:NVL

[BD01a] N. Balakrishnan and H. A. David. A note on the variance of a lightly trimmed mean when multiple outliers are present in the sample. Statistics & Probability Letters, 55(4):339–343, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002388.

Biedermann:2001:ODT

[BD01b] Stefanie Biedermann and Holger Dette. Optimal designs for testing the functional form of a regression via nonparametric estimation techniques. Statistics & Probability Letters, 52(2):215–224, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002443.

Bachmann:2005:NBR

[BD05a] Dirk Bachmann and Holger Dette. A note on the Bickel– Rosenblatt test in autoregressive time series. Statistics & Probability Letters, 74(3):221–234, October 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001215.

Birke:2005:NTC

[BD05b] Melanie Birke and Holger Dette. A note on testing the covariance matrix for large dimension. Statistics & Probability Letters, 74(3):281–289, October 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001872.

Boistard:2009:CLT

[BdB09] Hélène Boistard and Eustasio del Barrio. Central Limit Theorem for multiple integrals with respect to the empirical process. Statistics & Probability Letters, 79(2):188–195, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003672.

Bolfarine:2000:AMM

[BdC00] Heleno Bolfarine and Mário de Castro. ANOCOVA models with measurement errors. Statistics & Probability Letters, 50(3):257–263, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001000.

Barme-Delcroix:2007:LLM

[BDG07] Marie-Françoise Barme-Delcroix and Ursula Gather. Limit laws for multidimensional extremes. Statistics & Probability Letters, 77(18):1750–1755, December 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001617.

Bose:2008:MDT

[BDM08] Arup Bose, Amites Dasgupta, and Krishanu Maulik. Maxima of Dirichlet and triangular arrays of gamma variables. Statistics & Probability Letters, 78(16):2811–2820, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800206X.

Basu:2008:NTU

[BE08] Sanjib Basu and Nader Ebrahimi. A note on a transformation under censoring with application to partial least squares regression. Statistics & Probability Letters, 78(10):1161–1164, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207003458.

Bairamov:2009:WTE

[BE09] Ismihan Bairamov and Serkan Eryilmaz. Waiting times of exceedances in random threshold models. Statistics & Probability Letters, 79(5):676–683, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004975.

Beam:2007:UGP

[Bea07] John Beam. Unfair gambles in probability. Statistics & Probability Letters, 77(7):681–686, April 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003361.

Barakat:2009:ATE

[BEA09a] H. M. Barakat and Magdy E. El-Adll. Asymptotic theory of extreme dual generalized order statistics. Statistics & Probability Letters, 79(9):1252–1259, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000418.

Beare:2009:GHL

[Bea09b] Brendan K. Beare. A generalization of Hoeffding’s lemma, and a new class of covariance inequalities. Statistics & Probability Letters, 79(5):637–642, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004914.

Belisle:2000:SHR

[Bél00] Claude Bélisle. Slow hit-and-run sampling. Statistics & Probability Letters, 47(1):33–43, March 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001340.

Berkes:2001:LLN

[Ber01] István Berkes. The Law of Large Numbers with exceptional sets. Statistics & Probability Letters, 55(4):431–438, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001663.

Berrendero:2003:USC

[Ber03] José R. Berrendero. Uniform strong consistency of robust estimators. Statistics & Probability Letters, 64(2):159–168, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001482.

Berrahou:2008:LDP

[Ber08] Noureddine Berrahou. Large deviations probabilities for a symmetry test statistic based on delta-sequence density estimation. Statistics & Probability Letters, 78(3):238–248, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002192.

Barabesi:2001:ERS

[BES01] Lucio Barabesi and Abdel El-Sharaawi. The efficiency of ranked set sampling for parameter estimation. Statistics & Probability Letters, 53(2):189–199, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000785.

Betensky:2000:RAC

[Bet00] Rebecca A. Betensky. Redistribution algorithms for censored data. Statistics & Probability Letters, 46(4):385–389, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001273.

Bezandry:2006:LLN

[Bez06] Paul H. Bezandry. Laws of large numbers for D[0, 1]. Statistics & Probability Letters, 76(10):981–985, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004190.

Bezandry:2008:EAP

[Bez08] Paul H. Bezandry. Existence of almost periodic solutions to some functional integro-differential stochastic evolution equations. Statistics & Probability Letters, 78(17):2844–2849, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002186.

Boente:2000:KBF

[BF00]  Graciela Boente and Ricardo Fraiman.  Kernel-based

functional principal components. Statistics & Probability Letters, 48(4):335–345, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000146.

Bodnar:2009:IME

[BG09] Taras Bodnar and Arjun K. Gupta. An identity for multivariate elliptically contoured matrix distribution. Statistics & Probability Letters, 79(10):1327–1330, May 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000595.

Belzunce:2004:CHR

[BGHP04] Félix Belzunce, Xiaoli Gao, Taizhong Hu, and Franco Pellerey. Characterizations of the hazard rate order and IFR aging notion. Statistics & Probability Letters, 70(4):235–242, December 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002603.

Gao:2007:RPR

[bGhWhZhW07] Qi bing Gao, Yao hua Wu, Chun hua Zhu, and Guang hua Wei. Ruin problems in risk models with dependent rates of interest. Statistics & Probability Letters, 77(8):761–768, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206003488.

Bhansali:2007:CQF

[BGK07] R. J. Bhansali, L. Giraitis, and P. S. Kokoszka. Convergence of quadratic forms with nonvanishing diagonal. Statistics & Probability Letters, 77(7):726–734, April 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003427.

Basu:2003:EBP

[BGM03] Ruma Basu, J. K. Ghosh, and Rahul Mukerjee. Empirical Bayes prediction intervals in a normal regression model: higher order asymptotics. Statistics & Probability Letters, 63(2):197–203, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520300083X.

Bolance:2008:IBT

[BGN08] Catalina Bolancé, Montserrat Guillén, and Jens Perch Nielsen. Inverse beta transformation in kernel density estimation. Statistics & Probability Letters, 78(13):1757–1764, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000370.

Bojdecki:2004:SFB

[BGT04] Tomasz Bojdecki, Luis G. Gorostiza, and Anna Talarczyk. Sub-fractional Brownian motion and its relation to occupation times. Statistics & Probability Letters, 69(4):405–419, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001932.

Biswas:2002:NBB

[BH02] Atanu Biswas and Jing-Shiang Hwang. A new bivariate binomial distribution. Statistics & Probability Letters, 60(2): 231–240, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202003231.

Berkes:2003:RCQ

[BH03] István Berkes and Lajos Horváth. The rate of consistency of the quasi-maximum likelihood estimator. Statistics & Probability Letters, 61(2):133–143, January 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003425.

Barrera:2004:RSI

[BH04] Javiera Barrera and Thierry Huillet. On random splitting of the interval. Statistics & Probability Letters, 66(3):237– 250, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203003274.

Bischoff:2005:LBB

[BH05] Wolfgang Bischoff and Enkelejd Hashorva. A lower bound for boundary crossing probabilities of Brownian bridge/motion with trend. Statistics & Probability Letters, 74(3):265–271, October 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001653.

Belili:2006:AD

[BH06] Nacereddine Belili and Henri Heinich. Approximation of distributions. Statistics & Probability Letters, 76(3):298–303, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002981.

Baringhaus:2008:NWI

[BH08] L. Baringhaus and N. Henze. A new weighted integral goodness-of-fit statistic for exponentiality. Statistics & Probability Letters, 78(8):1006–1016, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003604.

Balakrishna:2009:SSE

[BH09] N. Balakrishna and G. Hareesh. Statistical signal extraction using stable processes. Statistics & Probability Letters, 79(7):851–856, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005257.

Bhandary:2005:TEA

[Bha05a] Madhusudan Bhandary. Test for the equality of autocorrelation coefficients for two populations in multivariate data when the errors are autocorrelated. Statistics & Probability Letters, 73(4):333–342, July 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001288.

Bhattacharya:2005:MIN

[Bha05b] P. K. Bhattacharya. A maximal inequality for nonnegative submartingales. Statistics & Probability Letters, 72(1):11–12, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002858.

Bhattacharya:2008:CEA

[Bha08] Sourabh Bhattacharya. Consistent estimation of the accuracy of importance sampling using regenerative simulation. Statistics & Probability Letters, 78(15):2522–2527, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001594.

Benhenni:2008:CRE

[BHGRV08] K. Benhenni, S. Hedli-Griche, M. Rachdi, and P. Vieu. Consistency of the regression estimator with functional data under long memory conditions. Statistics & Probability Letters, 78(8):1043–1049, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207004063.

Berkes:2008:FCL

[BHH08] István Berkes, Siegfried Hörmann, and Lajos Horváth. The functional Central Limit Theorem for a family of GARCH observations with applications. Statistics & Probability Letters, 78(16):2725–2730, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001922.

Bischoff:2004:PKT

[BHHM04] Wolfgang Bischoff, Enkelejd Hashorva, Jürg Hüsler, and Frank Miller. On the power of the Kolmogorov test to detect the trend of a Brownian bridge with applications to a change-point problem in regression models. Statistics & Probability Letters, 66(2):105–115, January 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002657.

Berkes:2004:TPC

[BHK04] István Berkes, Lajos Horváth, and Piotr Kokoszka. Testing for parameter constancy in GARCH(p, q) models. Statistics & Probability Letters, 70(4):263–273, December 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400272X.

Bibi:2004:PSB

[BhRH04] Abdelouahab Bibi and Moon ho Ringo Ho. Properties of some bilinear models with periodic regime switching. Statistics & Probability Letters, 69(3):221–231, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001610.

Bao:2009:SDN

[BHY09] Jianhai Bao, Zhenting Hou, and Chenggui Yuan. Stability in distribution of neutral stochastic differential delay equations with Markovian switching. Statistics & Probability Letters, 79(15):1663–1673, August 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001539.

Bansal:2000:NRF

[BHZ00] Naveen Bansal, G. G. Hamedani, and Hao Zhang. A note on random fields forming conditional bases. Statistics & Probability Letters, 50(4):397–400, December 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001292.

Bon:2002:NSN

[BI02] Jean-Louis Bon and Abbas Illayk. A note on some new renewal ageing notions. Statistics & Probability Letters, 57(2):151–155, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000366.

Biau:2002:OAQ

[Bia02] Gérard Biau. Optimal asymptotic quadratic errors of density estimators on random fields. Statistics & Probability Letters, 60(3):297–307, December 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200305X.

Bibi:2005:NSC

[Bib05] Abdelouahab Bibi. A note on the stability and causality of general time-dependent bilinear models. Statistics & Probability Letters, 73(2):131–138, June 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001021.

Balakrishnan:2009:PCS

[BIKM09] N. Balakrishnan, G. Iliopoulos, J. P. Keating, and R. L. Mason. Pitman closeness of sample median to population median. Statistics & Probability Letters, 79(16):1759–1766, August 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001667.

Birke:2008:CLT

[Bir08] Melanie Birke. Central limit theorems for the integrated squared error of derivative estimators. Statistics & Probability Letters, 78(13):1903–1913, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000576.

Bishop:2000:BVC

[Bis00] Jim Bishop. Bounds on p-values for a class of stopping rules. Statistics & Probability Letters, 48(4):385–392, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000201.

Bishwal:2001:ANA

[Bis01] J. P. N. Bishwal. Accuracy of normal approximation for the maximum likelihood estimator and Bayes estimators in the Ornstein–Uhlenbeck process using random normings. Statistics & Probability Letters, 52(4):427–439, May 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000268.

Biswas:2004:GCO

[Bis04] Atanu Biswas. Generating correlated ordinal categorical random samples. Statistics & Probability Letters, 70(1):25–35, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520400224X.

Bishwal:2006:RWC

[Bis06] Jaya P. N. Bishwal. Rates of weak convergence of approximate minimum contrast estimators for the discretely observed Ornstein–Uhlenbeck process. Statistics & Probability Letters, 76(13):1397–1409, July 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000472.

Bishwal:2008:LDT

[Bis08] Jaya P. N. Bishwal. Large deviations in testing fractional Ornstein–Uhlenbeck models. Statistics & Probability Letters, 78(8):953–962, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003471.

Bensaid:2000:LDI

[BJ00] Nadia Bensaïd and Pierre Jacob. Large deviations inequalities for partial sums of random measures. Statistics & Probability Letters, 49(2):105–112, August 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000304.

Behboodian:2006:NCS

[BJB06] J. Behboodian, A. Jamalizadeh, and N. Balakrishnan. A new class of skew-Cauchy distributions. Statistics & Probability Letters, 76(14):1488–1493, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000836.

Bardina:2000:WAW

[BJR00] Xavier Bardina, Maria Jolis, and Carles Rovira. Weak approximation of the Wiener process from a Poisson process: the multidimensional parameter set case. Statistics & Probability Letters, 50(3):245–255, November 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001139.

Bardina:2003:WCF

[BJT03] Xavier Bardina, Maria Jolis, and Ciprian A. Tudor. Weak convergence to the fractional Brownian sheet and other two-parameter Gaussian processes. Statistics & Probability Letters, 65(4):317–329, December 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002578.

Bhatt:2001:PCN

[BK01a] Abhay G. Bhatt and Rajeeva L. Karandikar. Path continuity of the nonlinear filter. Statistics & Probability Letters, 54(1):75–78, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000682.

Bon:2001:SEG

[BK01b] Jean-Louis Bon and Vladimir Kalashnikov. Some estimates of geometric sums. Statistics & Probability Letters, 55(1): 89–97, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001365.

Bai:2002:SDM

[BK02] Z. D. Bai and K. S. Kwong. Solution to Dalal and Mallows conjecture on monotone property of the joint distribution of order statistics. Statistics & Probability Letters, 59(1):29–35, August 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520200158X.

Bulinski:2004:NTP

[BK04] Alexander Bulinski and Andrei Khrennikov. Nonclassical total probability formula and quantum interference of probabilities. Statistics & Probability Letters, 70(1):49–58, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002275.

Beutner:2007:RCC

[BK07] Eric Beutner and Udo Kamps. Random convex combinations of order statistics. Statistics & Probability Letters, 77 (11):1133–1136, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520700065X.

Balakrishnan:2008:MLE

[BK08a] N. Balakrishnan and M. Kateri. On the maximum likelihood estimation of parameters of Weibull distribution based on complete and censored data. Statistics & Probability Letters, 78(17):2971–2975, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002514.

Balakrishnan:2008:CWP

[BK08b] N. Balakrishnan and Tomasz J. Kozubowski. A class of weighted Poisson processes. Statistics & Probability Letters, 78(15):2346–2352, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208001326.

Brilhante:2008:IDS

[BK08c] M. F. Brilhante and S. Kotz. Infinite divisibility of the spacings of a Kotz–Kozubowski–Podgórski generalized Laplace model. Statistics & Probability Letters, 78(15):2433–2436, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001454.

Balakrishnan:2009:MRU

[BKK09] N. Balakrishnan, U. Kamps, and M. Kateri. Minimal repair under a step-stress test. Statistics & Probability Letters, 79(13):1548–1558, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001199.

Breckling:2001:NMQ

[BKL01] Jens Breckling, Philip Kokic, and Oliver Lübke. A note on multivariate M-quantiles. Statistics & Probability Letters, 55(1):39–44, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001250.

Borkar:2004:FCK

[BKM04] V. S. Borkar, V. R. Konda, and S. K. Mitter. On de Finetti coherence and Kolmogorov probability. Statistics & Probability Letters, 66(4):417–421, March 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003572.

Braverman:2002:CSE

[BL02] Mark Braverman and Yan Lumelskii. Chebyshev systems and estimation theory for discrete distributions. Statistics & Probability Letters, 58(2):157–165, June 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001141.

Brown:2003:RBC

[BL03] Lawrence D. Brown and Yi Lin. Racetrack betting and consensus of subjective probabilities. Statistics & Probability Letters, 62(2):175–187, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000099.

Baltagi:2007:AWO

[BL07] Badi H. Baltagi and Long Liu. Alternative ways of obtaining Hausman’s test using artificial regressions. Statistics & Probability Letters, 77(13):1413–1417, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000703.

Baltagi:2008:TRE

[BL08] Badi H. Baltagi and Long Liu. Testing for random effects and spatial lag dependence in panel data models. Statistics & Probability Letters, 78(18):3304–3306, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003192.

Baltagi:2009:NAE

[BL09a] Badi H. Baltagi and Long Liu. A note on the application of EC2SLS and EC3SLS estimators in panel data models. Statistics & Probability Letters, 79(20):2189–2192, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209002673.

Bibi:2009:SMP

[BL09b] Abdelouahab Bibi and Radia Lessak. On stationarity and β-mixing of periodic bilinear processes. Statistics & Probability Letters, 79(1):79–87, January 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003489.

Blacher:2007:CLT

[Bla07] René Blacher. Central Limit Theorem by moments. Statistics & Probability Letters, 77(17):1647–1651, November 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001460.

Blazej:2008:PCL

[Bla08] Paweł Blazej. Preservation of classes of life distributions under weighting with a general weight function. Statistics & Probability Letters, 78(17):3056–3061, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002678.

Bar-Lev:2008:EFP

[BLB08] Shaul K. Bar-Lev and Daoud Bshouty. Exponential families are not preserved by the formation of order statistics. Statistics & Probability Letters, 78(16):2787–2792, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002022.

Bloznelis:2002:NML

[Blo02] M. Bloznelis. A note on the multivariate local limit theorem. Statistics & Probability Letters, 59(3):227–233, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001256.

Blondin:2007:RSU

[Blo07] David Blondin. Rates of strong uniform consistency for local least squares kernel regression estimators. Statistics & Probability Letters, 77(14):1526–1534, August 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001307.

Bar-Lev:2003:NNE

[BLP03] Shaul K. Bar-Lev and Denys Pommeret. A note on natural exponential families with cuts. Statistics & Probability Letters, 63(2):215–221, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000865.

Basawa:2004:FOS

[BLS04] I. V. Basawa, Robert Lund, and Qin Shao. First-order seasonal autoregressive processes with periodically varying parameters. Statistics & Probability Letters, 67(4):299–306, May 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000586.

Block:2005:MND

[BLS05] Henry W. Block, Yulin Li, and Thomas H. Savits. Mixtures of normal distributions: Modality and failure rate. Statistics & Probability Letters, 74(3):253–264, October 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001690.

Baltrunas:2008:PLD

[BLS08] Aleksandras Baltrunas, Remigijus Leipus, and Jonas Siaulys. Precise large deviation results for the total claim amount under subexponential claim sizes. Statistics & Probability Letters, 78(10):1206–1214, August 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700394X.

Bar-Lev:2004:NED

[BLV04] Shaul K. Bar-Lev and Frank A. Van der Duyn Schouten. A note on exponential dispersion models which are invariant under length-biased sampling. Statistics & Probability Letters, 70(4):275–284, December 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002743.

Besbeas:2001:ISE

[BM01a] Panagiotis Besbeas and Byron J. T. Morgan. Integrated squared error estimation of Cauchy parameters. Statistics & Probability Letters, 55(4):397–401, December 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001535.

Broniatowski:2001:SAT

[BM01b] Michel Broniatowski and Pascal Mignot. A self-adaptive technique for the estimation of the multifractal spectrum. Statistics & Probability Letters, 54(2):125–135, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200002108.

Bose:2002:LSD

[BM02] Arup Bose and Joydip Mitra. Limiting spectral distribution of a special circulant. Statistics & Probability Letters, 60(1):111–120, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202002894.

Balakrishnan:2003:EUM

[BM03a] N. Balakrishnan and Jie Mi. Existence and uniqueness of the MLEs for normal distribution based on general progressively Type-II censored samples. Statistics & Probability Letters, 64(4):407–414, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002062.

Bekker:2003:SCP

[BM03b] Leonid Bekker and Jie Mi. Shape and crossing properties of mean residual life functions. Statistics & Probability Letters, 64(3):225–234, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001470.

Belopolskaya:2003:AMN

[BM03c] Ya. Belopolskaya and G. N. Milstein. An approximation method for Navier–Stokes equations based on probabilistic approach. Statistics & Probability Letters, 64(2):201–211, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001524.

Bose:2003:OCD

[BM03d] Mausumi Bose and Bhramar Mukherjee. Optimal crossover designs under a general model. Statistics & Probability Letters, 62(4):413–418, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000506.

Besbeas:2004:ERE

[BM04] Panagiotis Besbeas and Byron J. T. Morgan. Efficient and robust estimation for the one-sided stable distribution of index 12. Statistics & Probability Letters, 66(3):251–257, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203003316.

Borovkov:2005:ABS

[BM05] K. Borovkov and A. Motyer. On the asymptotic behaviour of a simple growing point process model. Statistics & Probability Letters, 72(3):265–275, May 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000362.

Bischoff:2006:LFE

[BM06] Wolfgang Bischoff and Frank Miller. Lack-of-fit-efficiently optimal designs to estimate the highest coefficient of a polynomial with large degree. Statistics & Probability Letters, 76(15):1701–1704, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001246.

Baldi:2007:SCS

[BM07] Paolo Baldi and Domenico Marinucci. Some characterizations of the spherical harmonics coefficients for isotropic random fields. Statistics & Probability Letters, 77(5):490–496, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002690.

Belisle:2008:IAA

[BM08] Claude Bélisle and Vince Melfi. Independence after adaptive allocation. Statistics & Probability Letters, 78(3):214– 224, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002167.

Bebu:2009:CIL

[BM09] Ionut Bebu and Thomas Mathew. Confidence intervals for limited moments and truncated moments in normal and lognormal models. Statistics & Probability Letters, 79(3):375–380, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004227.

Blais:2000:LTR

[BMR00] Michel Blais, Brenda MacGibbon, and Roch Roy. Limit theorems for regression models of time series of counts. Statistics & Probability Letters, 46(2):161–168, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001017.

Bouezmarni:2007:RCS

[BMR07] T. Bouezmarni, M. Mesfioui, and J. M. Rolin. L_{1}- rate of convergence of smoothed histogram. Statistics & Probability Letters, 77(14):1497–1504, August 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001216.

Balaji:2006:DEP

[BMW06] Srinivasan Balaji, Hosam M. Mahmoud, and Osamu Watanabe. Distributions in the Ehrenfest process. Statistics & Probability Letters, 76(7):666–674, April 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003676.

Bagdonavicius:2000:GFL

[BN00] Vilijandas Bagdonavicius and Mikhail Nikulin. On goodnessof-fit for the linear transformation and frailty models. Statistics & Probability Letters, 47(2):177–188, April 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001558.

Borovkov:2001:PWD

[BN01] K. Borovkov and A. Novikov. On a piece-wise deterministic Markov process model. Statistics & Probability Letters, 53(4):421–428, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000748.

Borovkov:2008:ETL

[BN08] Konstantin Borovkov and Alexander Novikov. On exit times of Lévy-driven Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 78(12):1517–1525, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000060.

Beghin:2003:HST

[BNO03] L. Beghin, Y. Nikitin, and E. Orsingher. How the sojourn time distributions of Brownian motion are affected by different forms of conditioning. Statistics & Probability Letters, 65(4):291–302, December 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300230X.

Bibi:2002:NPS

[BO02] Abdelouahab Bibi and Alwell J. Oyet. A note on the properties of some time varying bilinear models. Statistics & Probability Letters, 58(4):399–411, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001530.

Bondesson:2003:MCP

[Bon03] Lennart Bondesson. On a minimum correlation problem. Statistics & Probability Letters, 62(4):361–370, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000403.

Borisov:2000:NPA

[Bor00] I. S. Borisov. A note on Poisson approximation of rescaled set-indexed empirical processes. Statistics & Probability Letters, 46(1):101–103, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299000942.

Boshnakov:2003:CCD

[Bos03a] Georgi N. Boshnakov. Confidence characteristics of distributions. Statistics & Probability Letters, 63(4):353–360, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001007.

Bosq:2003:BEI

[Bos03b] Denis Bosq. Berry–Esseen inequality for linear processes in Hilbert spaces. Statistics & Probability Letters, 63(3):243–247, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000889. See erra-

tum [Bos04b].

Bose:2004:RPD

[Bos04a] Sudip Bose. On the robustness of the predictive distribution for sampling from finite populations. Statistics & Probability Letters, 69(1):21–27, August 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001026.

Bosq:2004:ECB

[Bos04b] Denis Bosq. Erratum and complements to “Berry–Esseen inequality for linear processes in Hilbert Spaces”. Statistics & Probability Letters, 70(2):171–174, November 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002251. See [Bos03b].

Boshnakov:2007:SMA

[Bos07a] Georgi N. Boshnakov. Some measures for asymmetry of distributions. Statistics & Probability Letters, 77(11):1111–1116, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000624.

Bosq:2007:SES

[Bos07b] Denis Bosq. Sufficiency and efficiency in statistical prediction. Statistics & Probability Letters, 77(3):280–287, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002343.

Boshnakov:2009:AEP

[Bos09] Georgi N. Boshnakov. Analytic expressions for predictive distributions in mixture autoregressive models. Statistics & Probability Letters, 79(15):1704–1709, August 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001588.

Boudou:2006:APC

[Bou06] Alain Boudou. Approximation of the principal components analysis of a stationary function. Statistics & Probability Letters, 76(6):571–578, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003482.

Boucher:2009:HIM

[Bou09] Thomas R. Boucher. A Hoeffding inequality for Markov chains using a generalized inverse. Statistics & Probability Letters, 79(8):1105–1107, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005816.

Balakrishnan:2002:USE

[BP02] N. Balakrishnan and N. Papadatos. The use of spacings in the estimation of a scale parameter. Statistics & Probability Letters, 57(2):193–204, April 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000500.

Barczy:2006:PTU

[BP06] Mátyás Barczy and Gyula Pap. Portmanteau theorem for unbounded measures. Statistics & Probability Letters, 76(17):1831–1835, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001428.

Bura:2008:DLS

[BP08] E. Bura and R. Pfeiffer. On the distribution of the left singular vectors of a random matrix and its applications. Statistics & Probability Letters, 78(15):2275–2280, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001211.

Buonocore:2009:NSU

[BPC09] Aniello Buonocore, Enrica Pirozzi, and Luigia Caputo. A note on the sum of uniform random variables. Statistics & Probability Letters, 79(19):2092–2097, October 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002442.

Belzunce:2001:FTR

[BPR01] F. Belzunce, J. F. Pinar, and J. M. Ruiz. A family of tests for right spread order. Statistics & Probability Letters, 54(1):79–92, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520100075X.

Baran:2004:AIN

[BPvZ04] Sándor Baran, Gyula Pap, and Martien C. A. van Zuijlen. Asymptotic inference for a nearly unstable sequence of stationary spatial AR models. Statistics & Probability Letters, 69(1):53–61, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001580.

Baron:2001:PAC

[BR01] Michael Baron and Andrew L. Rukhin. Perpetuities and asymptotic change-point analysis. Statistics & Probability Letters, 55(1):29–38, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001225.

Berti:2002:ULT

[BR02a] Patrizia Berti and Pietro Rigo. A uniform limit theorem for predictive distributions. Statistics & Probability Letters, 56(2):113–120, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520100089X.

Boudou:2002:SRM

[BR02b] Alain Boudou and Yves Romain. On spectral and random measures associated to discrete and continuous-time processes. Statistics & Probability Letters, 59(2):145–157, September 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001426.

Biswas:2004:MRA

[BR04] Atanu Biswas and J. N. K. Rao. Missing responses in adaptive allocation design. Statistics & Probability Letters, 70(1):59–70, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002287.

Bebu:2006:SDA

[BR06a] Ionut Bebu and Andrew L. Rukhin. Stationary distributions in the atom-on-demand problem. Statistics & Probability Letters, 76(15):1607–1616, September 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001131.

Boente:2006:REH

[BR06b] Graciela Boente and Daniela Rodriguez. Robust estimators of high order derivatives of regression functions. Statistics & Probability Letters, 76(13):1335–1344, July 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000290.

Bouhaddioui:2006:DRC

[BR06c] Chafik Bouhaddioui and Roch Roy. On the distribution of the residual cross-correlations of infinite order vector autoregressive series and applications. Statistics & Probability Letters, 76(1):58–68, January 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002701.

Balabdaoui:2008:SMI

[BR08] Fadoua Balabdaoui and Kaspar Rufibach. A second Marshall inequality in convex estimation. Statistics & Probability Letters, 78(2):118–126, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002027.

Bravo:2002:BEC

[Bra02a] Francesco Bravo. Blockwise empirical Cressie–Read test statistics for α-mixing processes. Statistics & Probability Letters, 58(3):319–325, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001621.

Brazauskas:2002:FIM

[Bra02b] Vytaras Brazauskas. Fisher information matrix for the Feller– Pareto distribution. Statistics & Probability Letters, 59(2): 159–167, September 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001438.

Braverman:2005:CLP

[Bra05] Michael Braverman. On a class of Lévy processes. Statistics & Probability Letters, 75(3):179–189, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002245.

Breton:2006:CVJ

[Bre06] Jean-Christophe Breton. Convergence in variation of the joint laws of multiple Wiener–Itô integrals. Statistics & Probability Letters, 76(17):1904–1913, November 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600157X.

Breunig:2008:NDE

[Bre08] Robert Breunig. Nonparametric density estimation for stratified samples. Statistics & Probability Letters, 78(14):2194– 2200, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001107.

Bhattacharyya:2006:URP

[BRF06] B. B. Bhattacharyya, G. D. Richardson, and P. V. Flores. Unit roots: Periodogram ordinate. Statistics & Probability Letters, 76(6):641–651, March 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003652.

Brigo:2000:SML

[Bri00] Damiano Brigo. On SDEs with marginal laws evolving in finite-dimensional exponential families. Statistics & Probability Letters, 49(2):127–134, August 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000390.

Brockwell:2007:URM

[Bro07] A. E. Brockwell. Universal residuals: a multivariate transformation. Statistics & Probability Letters, 77(14):1473–1478, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000715. See ac-

knowledgement of priority [Bro11].

Brown:2008:HOM

[Bro08a] Geoffrey W. Brown. Higher order moments of renewal counting processes and Eulerian polynomials. Statistics & Probability Letters, 78(17):3008–3013, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002599.

Brown:2008:MPP

[Bro08b] Marlo Brown. Monitoring a Poisson process in several categories subject to changes in the arrival rates. Statistics & Probability Letters, 78(16):2637–2643, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001764.

Brown:2009:LSD

[Bro09] B. M. Brown. The logit-stable distributions and latent utility scales in categorical data applications. Statistics & Probability Letters, 79(20):2182–2188, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002636.

Brockwell:2011:APU

[Bro11] A. E. Brockwell. Acknowledgement of priority to: “Universal residuals: a multivariate transformation. Statistics & Probability Letters 27, 2007, 1473–1478”. Statistics & Probability Letters, 81(12):1822, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001477. See [Bro07].

Breyer:2001:NGE

[BRR01] Laird Breyer, Gareth O. Roberts, and Jeffrey S. Rosenthal. A note on geometric ergodicity and floating-point roundoff error. Statistics & Probability Letters, 53(2):123–127, June 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000542.

Belzunce:2002:PSS

[BRR02] Félix Belzunce, José M. Ruiz, and M. Carmen Ruiz. On preservation of some shifted and proportional orders by systems. Statistics & Probability Letters, 60(2):141–154, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003024.

Bhattacharyya:2003:SAM

[BRRZ03] B. B. Bhattacharyya, J.-J. Ren, G. D. Richardson, and J. Zhang. Spatial autoregression model: strong consistency. Statistics & Probability Letters, 65(2):71–77, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002190.

Belzunce:2008:MDO

[BRSL08] Félix Belzunce, José M. Ruiz, and Alfonso Suárez-Llorens. On multivariate dispersion orderings based on the standard construction. Statistics & Probability Letters, 78(3):271– 281, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002386.

Brumback:2009:NUE

[Bru09] Babette A. Brumback. A note on using the estimated versus the known propensity score to estimate the average treatment effect. Statistics & Probability Letters, 79(4):537–542, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800463X.

Bhattacharjee:2003:NOP

[BRVM03] Manish C. Bhattacharjee, S. Ravi, R. Vasudeva, and N. R. Mohan. New order preserving properties of geometric compounds. Statistics & Probability Letters, 64(2):113–120, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000397.

Bryc:2000:SBD

[Bry00] Włodzimierz Bryc. Specifying bivariate distributions by polynomial regressions. Statistics & Probability Letters, 47(4):391– 394, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001832.

Biswas:2000:ASM

[BS00] Atanu Biswas and Jyotirmoy Sarkar. Availability of a system maintained through several imperfect repairs before a replacement or a perfect repair. Statistics & Probability Letters, 50(2):105–114, November 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000870.

Belzunce:2001:SCM

[BS01a] Félix Belzunce and Moshe Shaked. Stochastic comparisons of mixtures of convexly ordered distributions with applications in reliability theory. Statistics & Probability Letters, 53(4):363– 372, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520100030X.

Bulinski:2001:NAQ

[BS01b] Alexander Bulinski and Charles Suquet. Normal approximation for quasi-associated random fields. Statistics & Probability Letters, 54(2):215–226, September 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001080.

Balakrishnan:2004:NPK

[BS04] N. Balakrishnan and A. Stepanov. A note on the paper of Khmaladze et al. Statistics & Probability Letters, 68(4): 415–419, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001312.

Bairamov:2006:NLD

[BS06a] Ismihan Bairamov and Alexei Stepanov. A note on large deviations for weak records. Statistics & Probability Letters, 76(14):1449–1453, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000782.

Balakrishnan:2006:FIR

[BS06b] N. Balakrishnan and A. Stepanov. On the Fisher information in record data. Statistics & Probability Letters, 76(5):537–545, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003354.

Bartoszewicz:2006:PCL

[BS06c] Jarosław Bartoszewicz and Magdalena Skolimowska. Preservation of classes of life distributions and stochastic orders under weighting. Statistics & Probability Letters, 76(6):587–596, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003500.

Bisgaard:2006:WDI

[BS06d] Torben Maack Bisgaard and Zoltán Sasvári. When does E(X^{k} · Y^{ℓ}) = E(X^{k}) · E(Y^{ℓ}) imply independence? Statistics & Probability Letters, 76(11):1111–1116, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004542.

Balakrishnan:2008:APR

[BS08a] N. Balakrishnan and A. Stepanov. Asymptotic properties of the ratio of order statistics. Statistics & Probability Letters, 78(3):301–310, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002428.

Bhattacharya:2008:OAC

[BS08b] Debasis Bhattacharya and Francisco J. Samaniego. On the optimal allocation of components within coherent systems. Statistics & Probability Letters, 78(7):938–943, May 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003446.

Biswas:2009:DVA

[BS09] Atanu Biswas and Peter X.-K. Song. Discrete-valued ARMA processes. Statistics & Probability Letters, 79(17):1884– 1889, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001977.

Barreto-Souza:2009:GEP

[BSCN09] Wagner Barreto-Souza and Francisco Cribari-Neto. A generalization of the exponential-Poisson distribution. Statistics & Probability Letters, 79(24):2493–2500, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003447.

Barone:2001:GRM

[BSS01] Piero Barone, Giovanni Sebastiani, and Julian Stander. General over-relaxation Markov chain Monte Carlo algorithms for Gaussian densities. Statistics & Probability Letters, 52(2):115– 124, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001656.

Blei:2002:VFM

[BT02] R. C. Blei and Nasser Towghi. Variations of Frechet measures of p-stable motions. Statistics & Probability Letters, 56(4):355–360, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001560.

Barcena:2004:FMR

[BT04] María Jesús Barcena and Fernando Tusell. Fitting multivariate responses using scalar trees. Statistics & Probability Letters, 69(3):253–259, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001671.

Bhattacharya:2001:NDI

[BTW01] Rabi Bhattacharya, Enrique Thomann, and Edward Waymire. A note on the distribution of integrals of geometric Brownian motion. Statistics & Probability Letters, 55(2):187– 192, November 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001171.

Burke:2000:MTF

[Bur00] Murray D. Burke. Multivariate tests-of-fit and uniform confidence bands using a weighted bootstrap. Statistics & Probability Letters, 46(1):13–20, January 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000826.

Burman:2002:EEH

[Bur02] Prabir Burman. Estimation of equifrequency histograms. Statistics & Probability Letters, 56(3):227–238, February 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000591.

Butucea:2000:ARC

[But00] Cristina Butucea. The adaptive rate of convergence in a problem of pointwise density estimation. Statistics & Probability Letters, 47(1):85–90, March 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001418.

Butler:2001:WDO

[But01] Neil A. Butler. On weighted design optimality criteria and response surface parameterizations. Statistics & Probability Letters, 51(1):41–46, January 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001401.

Butler:2004:MAP

[But04] Neil A. Butler. Minimum G_{2}-aberration properties of two-level foldover designs. Statistics & Probability Letters, 67(2):121– 132, April 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000070.

Butler:2005:MRS

[But05] Neil A. Butler. Minimax 16-run supersaturated designs. Statistics & Probability Letters, 73(2):139–145, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001033.

Butler:2008:SOT

[But08] Neil A. Butler. Schur- and E-optimal two-level factorial designs. Statistics & Probability Letters, 78(5):518–527, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002830.

Buzas:2009:NCS

[Buz09] Jeffrey S. Buzas. A note on corrected scores for logistic regression. Statistics & Probability Letters, 79(22):2351– 2358, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520900306X.

Beirlant:2002:SCE

[BV02] J. Beirlant and B. Vandewalle. Some comments on the estimation of a dependence index in bivariate extreme value statistics. Statistics & Probability Letters, 60(3):265–278, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520200281X.

Barrios:2007:BMA

[BV07] M. Pilar Barrios and Santiago Velilla. A bootstrap method for assessing the dimension of a general regression problem. Statistics & Probability Letters, 77(3):247–255, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002306.

Braekers:2008:CKG

[BV08] Roel Braekers and Noël Veraverbeke. A conditional Koziol– Green model under dependent censoring. Statistics & Probability Letters, 78(7):927–937, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003434.

Banachewicz:2008:TDS

[BvdV08] Konrad Banachewicz and Aad van der Vaart. Tail dependence of skewed grouped t-distributions. Statistics & Probability Letters, 78(15):2388–2399, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0167715208001387. See corrigendum [BvdV09].

Banachewicz:2009:CTD

[BvdV09] Konrad Banachewicz and Aad van der Vaart. Corrigendum to: “Tail dependence of skewed grouped t-distributions”

[Statist. Probab. Lett. 78 (2008) 2388–2399]. Statistics & Probability Letters, 79(15):1731, August 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001643. See [BvdV08].

Biau:2005:NMD

[BW05] Gérard Biau and Marten Wegkamp. A note on minimum distance estimation of copula densities. Statistics & Probability Letters, 73(2):105–114, June 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000854.

Berkes:2006:ASV

[BW06a] István Berkes and Michel Weber. Almost sure versions of the Darling-Erdős theorem. Statistics & Probability Letters, 76(3):280–290, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520500297X.

Bryc:2006:CBP

[BW06b] Włodzimierz Bryc and Jacek Wesołowski. The classical bi-Poisson process: An invertible quadratic harness. Statistics & Probability Letters, 76(15):1664–1674, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001192.

Berkes:2007:CCT

[BW07] István Berkes and Michel Weber. On complete convergence of triangular arrays of independent random variables. Statistics & Probability Letters, 77(10):952–963, June 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000272.

Borovskikh:2008:ADN

[BW08] Yuri V. Borovskikh and N. C. Weber. Asymptotic distributions of non-degenerate U-statistics on trimmed samples. Statistics & Probability Letters, 78(4):336–346, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002465.

Bentler:2000:CTS

[BX00] Peter M. Bentler and Jun Xie. Corrections to test statistics in principal Hessian directions. Statistics & Probability Letters, 47(4):381–389, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001820.

Bensaid:2008:SDE

[BYB08] Nadia Bensaïd and Ouafae Yazourh-Benrabah. Spectral density estimation for linear processes with dependent innovations. Statistics & Probability Letters, 78(12):1601– 1611, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000175.

Brown:2006:NOS

[BZ06] Marlo Brown and Shelemyahu Zacks. A note on optimal stopping for possible change in the intensity of an ordinary Poisson process. Statistics & Probability Letters, 76(13):1417–1425, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000502.

Claeskens:2000:LEE

[CA00] Gerda Claeskens and Marc Aerts. On local estimating equations in additive multiparameter models. Statistics & Probability Letters, 49(2):139–148, August 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000420.

Capar:2001:NCR

[ÇA01] Ulug Çapar and Hüseyin Aktuglu. A new construction of random Colombeau distributions. Statistics & Probability Letters, 54(3):291–299, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001018.

Chivoret:2006:PFIa

[CA06a] Sébastien Chivoret and Anna Amirdjanova. Product formula and independence criterion for multiple Huang–Cambanis integrals. Statistics & Probability Letters, 76(12):1255–1260, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000216. See erra-

tum [Ano06o].

Chivoret:2006:PFIb

[CA06b] Sébastien Chivoret and Anna Amirdjanova. Product formula and independence criterion for multiple Huang–Cambanis integrals. Statistics & Probability Letters, 76(18):2037–2042, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002252.

Cacoullos:2001:THC

[Cac01]  Theophilos Cacoullos.  The F -test of homoscedastic-

ity for correlated normal variables. Statistics & Probability Letters, 54(1):1–3, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001899.

Cai:2000:NAT

[Cai00] Haiyan Cai. A normal approximation theorem in comparing two binomial distributions. Statistics & Probability Letters, 48(1):83–89, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001923.

Cai:2001:WNW

[Cai01] Zongwu Cai. Weighted Nadaraya–Watson regression estimation. Statistics & Probability Letters, 51(3):307–318, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001723.

Cai:2002:LIL

[Cai02] Haiyan Cai. A law of iterated logarithm for the wavelet transforms of i.i.d. random variables. Statistics & Probability Letters, 60(1):121–127, November 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002900.

Cai:2003:NEE

[Cai03] Zongwu Cai. Nonparametric estimation equations for time series data. Statistics & Probability Letters, 62(4):379–390, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000427.

Claeskens:2003:QBM

[CAM03] Gerda Claeskens, Marc Aerts, and Geert Molenberghs. A quadratic bootstrap method and improved estimation in logistic regression. Statistics & Probability Letters, 61(4):383– 394, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202003978.

Carbone:2004:BAB

[Car04] Raffaella Carbone. Binomial approximation of Brownian motion and its maximum. Statistics & Probability Letters, 69(3):271–285, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001750.

Cassese:2007:DSI

[Cas07] Gianluca Cassese. Decomposition of supermartingales indexed by a linearly ordered set. Statistics & Probability Letters, 77(8):795–802, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000041.

Cator:2001:DAS

[Cat01] Eric A. Cator. Deconvolution with arbitrarily smooth kernels. Statistics & Probability Letters, 54(2):205–214, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000839.

Cordeiro:2001:SOB

[CB01] Gauss M. Cordeiro and Denise A. Botter. Second-order biases of maximum likelihood estimates in overdispersed generalized linear models. Statistics & Probability Letters, 55(3):269–280, December 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520100150X.

Chen:2007:EFL

[CB07] Xiong Chen and N. Balakrishnan. Extensions of functional LIL w.r.t. (r, p)-capacities on Wiener space. Statistics & Probability Letters, 77(4):468–473, February 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002847.

Cortina-Borja:2000:EAC

[CBR00] Mario Cortina-Borja and Tony Robinson. Estimating the asymptotic constants of the total length of Euclidean minimal spanning trees with power-weighted edges. Statistics & Probability Letters, 47(2):125–128, April 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001479.

Chen:2001:FMC

[CC01a] Bin Chen and Miklós Csörgő. A functional modulus of continuity for a Wiener process. Statistics & Probability Letters, 51(3):215–223, February 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001206.

Chen:2001:LSD

[CC01b] Hanfeng Chen and Jiahua Chen. Large sample distribution of the likelihood ratio test for normal mixtures. Statistics & Probability Letters, 52(2):125–133, April 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001711.

Chen:2001:REE

[cC01c] Chu chih Chen. Rank estimating equations for random effects models. Statistics & Probability Letters, 54(1):5–12, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002261.

Cheng:2002:SCC

[CC02] Tsung-Lin Cheng and Yuan-Shih Chow. On stable convergence in the Central Limit Theorem. Statistics & Probability Letters, 57(4):307–313, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000597.

Cerqueti:2008:ABR

[CC08a] Roy Cerqueti and Mauro Costantini. On the asymptotic behaviour of random matrices in a multivariate statistical model. Statistics & Probability Letters, 78(14):2039–2045, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000758.

Chan:2008:NAP

[CC08b] Wai-Sum Chan and Yin-Ting Chan. A note on the autocorrelation properties of temporally aggregated Markov switching Gaussian models. Statistics & Probability Letters, 78(6):728–735, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003240.

Chen:2008:ELI

[CC08c] Xia Chen and Hengjian Cui. Empirical likelihood inference for partial linear models under martingale difference sequence. Statistics & Probability Letters, 78(17):2895–2901, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002332.

Chen:2002:SIC

[CCC02] Guijing Chen, Jiahua Chen, and Yuming Chen. Statistical inference on comparing two distribution functions with a possible crossing point. Statistics & Probability Letters, 60(3):329–341, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003206.

Choi:2002:OOB

[CCDG02] Kuey Chung Choi, Kashinath Chatterjee, Ashish Das, and Sudhir Gupta. Optimality of orthogonally blocked diallels with specific combining abilities. Statistics & Probability Letters, 57(2):145–150, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000354.

Caliskan:2009:ERS

[CCG09] Derya Caliskan, Christophe Croux, and Sarah Gelper. Efficient and robust scale estimation for trended time series. Statistics & Probability Letters, 79(18):1900–1905, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001989.

Coolen:2006:NPS

[CCS06] F. P. A. Coolen and P. Coolen-Schrijner. Nonparametric predictive subset selection for proportions. Statistics & Probability Letters, 76(15):1675–1684, September 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001222.

Costa:2000:IPM

[CD00] O. L. V. Costa and F. Dufour. Invariant probability measures for a class of Feller Markov chains. Statistics & Probability Letters, 50(1):13–21, October 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000754.

Campos:2001:KDE

[CD01a] V. S. M. Campos and C. C. Y. Dorea. Kernel density estimation: the general case. Statistics & Probability Letters, 55(2):173–180, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001092.

Costa:2001:NSC

[CD01b] O. L. V. Costa and F. Dufour. Necessary and sufficient conditions for non-singular invariant probability measures for Feller Markov chains. Statistics & Probability Letters, 53(1):47–57, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000165.

Chaudhuri:2005:SSM

[CD05] Probal Chaudhuri and Amites Dasgupta. On some stochastic models for replication of character strings. Statistics & Probability Letters, 75(3):230–236, December 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002373.

Colubi:2001:MDS

[CDMLDK01] Ana Colubi, J. Santos Domínguez-Menchero, Miguel López-Díaz, and Ralf Körner. A method to derive strong laws of large numbers for random upper semicontinuous functions. Statistics & Probability Letters, 53(3):269–275, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000578.

Croux:2001:REC

[CDRV01] Christophe Croux, Catherine Dehon, Peter J. Rousseeuw, and Stefan Van Aelst. Robust estimation of the conditional median function at elliptical models. Statistics & Probability Letters, 51(4):361–368, February 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001760.

Cekanavicius:2002:CMB

[Cek02] V. Cekanavicius. On the convergence of Markov binomial to Poisson distribution. Statistics & Probability Letters, 58 (1):83–91, May 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001189.

Cerioli:2002:THS

[Cer02] Andrea Cerioli. Tests of homogeneity for spatial populations. Statistics & Probability Letters, 58(2):123–130, June 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001074.

Cerquetti:2007:NBN

[Cer07] Annalisa Cerquetti. A note on Bayesian nonparametric priors derived from exponentially tilted Poisson–Kingman models. Statistics & Probability Letters, 77(18):1705–1711, December 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207001551.

Cerquetti:2008:GCN

[Cer08] Annalisa Cerquetti. On a Gibbs characterization of normalized generalized gamma processes. Statistics & Probability Letters, 78(18):3123–3128, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002770.

Carnal:2000:SPP

[CF00] H. Carnal and G. M. Feldman. A stability property for probability measures on Abelian groups. Statistics & Probability Letters, 49(1):39–44, August 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000286.

Cheung:2001:LEA

[CF01] Ying-Kuen Cheung and Jason P. Fine. Likelihood estimation after nonparametric transformation. Statistics & Probability Letters, 55(1):1–7, November 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001067.

Cysneiros:2006:ILR

[CF06] Audrey H. M. A. Cysneiros and Silvia L. P. Ferrari. An improved likelihood ratio test for varying dispersion in exponential family nonlinear models. Statistics & Probability Letters, 76(3):255–265, February 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002944.

Crujeiras:2008:TCS

[CFCGM08] Rosa M. Crujeiras, Rubén Fernández-Casal, and Wenceslao González-Manteiga. An L_{2}-test for comparing spatial spectral densities. Statistics & Probability Letters, 78(15):2543–2551, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001685.

Chang:2003:TPL

[CFCL03] C. J. Chang, C. S. J. Fann, W. C. Chou, and I. B. Lian. On the tail probability of the longest well-matching run. Statistics & Probability Letters, 63(3):267–274, July 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000919.

Croux:2002:BBM

[CFH02] Christophe Croux, Cécile Flandre, and Gentiane Haesbroeck. The breakdown behavior of the maximum likelihood estimator in the logistic regression model. Statistics & Probability Letters, 60(4):377–386, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002924.

Chan:2001:COA

[CFM01] Ling-Yau Chan, Kai-Tai Fang, and Rahul Mukerjee. A characterization for orthogonal arrays of strength two via a regression model. Statistics & Probability Letters, 54(2): 189–192, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000517.

Csaki:2006:HPD

[CFR06] Endre Csáki, Antónia Földes, and Pál Révész. Heavy points of a d-dimensional simple random walk. Statistics & Probability Letters, 76(1):45–57, January 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002622.

Cordeiro:2000:CML

[CFUOV00] Gauss M. Cordeiro, Silvia L. P. Ferrari, Miguel A. Uribe-Opazo, and Klaus L. P. Vasconcellos. Corrected maximumlikelihood estimation in a class of symmetric nonlinear regression models. Statistics & Probability Letters, 46(4):317– 328, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001182.

Castro:2002:MEE

[CG02a] Glaysar Castro and Valerie Girardin. Maximum of entropy and extension of covariance matrices for periodically correlated and multivariate processes. Statistics & Probability Letters, 59(1):37–52, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001591.

Chen:2002:ACT

[CG02b] Jie Chen and Joseph Glaz. Approximations for a conditional two-dimensional scan statistic. Statistics & Probability Letters, 58(3):287–296, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001220.

Csaki:2002:ASL

[CG02c] Endre Csáki and Khurelbaatar Gonchigdanzan. Almost sure limit theorems for the maximum of stationary Gaussian sequences. Statistics & Probability Letters, 58(2):195–203, June 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001281.

Castro:2008:CPC

[CG08a] Glaysar Castro and Valerie Girardin. Characterization of periodically correlated and multivariate stationary discrete time wide Markov processes. Statistics & Probability Letters, 78(2):158–164, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002076.

Chen:2008:MMN

[CG08b] Pingyan Chen and Shixin Gan. On moments of the maximum of normed partial sums of ρ-mixing random variables. Statistics & Probability Letters, 78(10):1215–1221, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003951.

Ceranka:2006:OCB

[CGK06] Bronisław Ceranka, Małgorzata Graczyk, and Krystyna Katulska. A-optimal chemical balance weighing design with nonhomogeneity of variances of errors. Statistics & Probability Letters, 76(7):653–665, April 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003664.

Chen:2001:BTI

[CGNW01] Jie Chen, Joseph Glaz, Joseph Naus, and Sylvan Wallenstein. Bonferroni-type inequalities for conditional scan statistics. Statistics & Probability Letters, 53(1):67–77, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000347.

Courbage:2001:EEP

[CH01] M. Courbage and D. Hamdan. Examples of ergodic processes uniquely determined by their two-marginal laws. Statistics & Probability Letters, 52(4):341–345, May 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001905.

Chen:2004:CBS

[CH04a] John T. Chen and Fred M. Hoppe. A connection between successive comparisons and ranking procedures. Statistics & Probability Letters, 67(1):19–25, March 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300261X.

Chou:2004:CGG

[CH04b] Chao-Wei Chou and Wen-Jang Huang. On characterizations of the gamma and generalized inverse Gaussian distributions. Statistics & Probability Letters, 69(4):381–388, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000306.

Cao:2007:QSV

[CH07] Guilan Cao and Kai He. Quasi-sure p-variation of fractional Brownian motion. Statistics & Probability Letters, 77(5):543–548, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002860.

Christofides:2009:EID

[CH09a] Tasos C. Christofides and Milto Hadjikyriakou. Exponential inequalities for N-demimartingales and negatively associated random variables. Statistics & Probability Letters, 79(19):2060–2065, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002405.

Cirillo:2009:UAG

[CH09b] Pasquale Cirillo and Jürg Hüsler. An urn approach to generalized extreme shock models. Statistics & Probability Letters, 79(7):969–976, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005622.

Chang:2001:SOS

[Cha01]  Kuo-Hwa Chang.  Stochastic orders of the sums of

two exponential random variables. Statistics & Probability Letters, 51(4):389–396, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001796.

Chaabane:2002:IPL

[Cha02] Faouzi Chaabane. Invariance principles with logarithmic averaging for continuous local martingales. Statistics & Probability Letters, 59(2):209–217, September 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002079.

Chamayou:2004:PDG

[Cha04] Jean-François Chamayou. Products of double gamma, gamma and beta distributions. Statistics & Probability Letters, 68(2):199–208, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000872.

Chang:2005:DWT

[Cha05] Yung-Ming Chang. Distribution of waiting time until the r th occurrence of a compound pattern. Statistics & Probability Letters, 75(1):29–38, November 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002075.

Chandra:2006:PSJ

[Cha06a] Tapas Kumar Chandra. On a problem of Simons and Johnson. Statistics & Probability Letters, 76(11):1132–1136, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004578.

Chang:2006:CRS

[Cha06b] Hyeong Soo Chang. On convergence rate of the Shannon entropy rate of ergodic Markov chains via sample-path simulation. Statistics & Probability Letters, 76(12):1261–1264, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000228.

Chandra:2008:BCL

[Cha08] Tapas Kumar Chandra. The Borel–Cantelli lemma under dependence conditions. Statistics & Probability Letters, 78 (4):390–395, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002520.

Chareka:2009:CLT

[Cha09] Patrick Chareka. The central limit theorem for capacities. Statistics & Probability Letters, 79(12):1456–1462, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001096.

Chabot-Halle:2008:DCM

[CHD08] Dominique Chabot-Hallé and Pierre Duchesne. Diagnostic checking of multivariate nonlinear time series models with martingale difference errors. Statistics & Probability Letters, 78(8):997–1005, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003598.

Chen:2000:CLA

[Che00a] Xia Chen. Chung’s law for additive functionals of positive recurrent Markov chains. Statistics & Probability Letters, 47(3):253–264, April 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001637.

Chen:2000:NTP

[Che00b] Zhenmin Chen. A new two-parameter lifetime distribution with bathtub shape or increasing failure rate function. Statistics & Probability Letters, 49(2):155–161, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000444.

Chen:2000:BPS

[Che00c] Zhiqiang Chen. Breakdown point of Schuster–Narvarte’s location estimator. Statistics & Probability Letters, 48(3):283–286, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000080.

Cheng:2002:ASU

[Che02a] Cheng Cheng. Almost-sure uniform error bounds of general smooth estimators of quantile density functions. Statistics & Probability Letters, 59(2):183–194, September 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001608.

Cheng:2002:CED

[Che02b] Fuxia Cheng. Consistency of error density and distribution function estimators in nonparametric regression. Statistics & Probability Letters, 59(3):257–270, October 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001554.

Chebana:2004:OWB

[Che04a] Fateh Chebana. On the optimization of the weighted Bickel– Rosenblatt test. Statistics & Probability Letters, 68(4): 333–345, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001087.

Cheng:2004:APM

[Che04b] Yu Cheng. Asymptotic probabilities of misclassification of two discriminant functions in cases of high dimensional data. Statistics & Probability Letters, 67(1):9–17, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003857.

Chen:2005:ALR

[Che05] Yuguo Chen. Another look at rejection sampling through importance sampling. Statistics & Probability Letters, 72(4): 277–283, May 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000374.

Cheliotis:2006:NRR

[Che06] Dimitrios Cheliotis. A note on recurrent random walks. Statistics & Probability Letters, 76(10):1025–1031, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004438.

Chen:2007:GFH

[Che07a] Haiyan Chen. The generating functions of hitting times for random walk on trees. Statistics & Probability Letters, 77 (15):1574–1579, September 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001332.

Chesneau:2007:RRD

[Che07b] Christophe Chesneau. Regression with random design: a minimax study. Statistics & Probability Letters, 77(1):40–53, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001933.

Chen:2008:AKD

[Che08a] Jia Chen. Asymptotics of kernel density estimators on weakly associated random fields. Statistics & Probability Letters, 78(18):3230–3237, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003039.

Cheng:2008:EGC

[Che08b] Fuxia Cheng. Extended Glivenko–Cantelli theorem in ARCH(p)-time series. Statistics & Probability Letters, 78(12): 1434–1439, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207004269.

Chesneau:2008:MCB

[Che08c] Christophe Chesneau. On the maxiset comparison between hard and block thresholding methods. Statistics & Probability Letters, 78(6):675–681, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003112.

Cheng:2009:EAE

[Che09] Tsung-Lin Cheng. An efficient algorithm for estimating a change-point. Statistics & Probability Letters, 79(5):559–565, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004653.

Clark:2001:ESR

[CHL01] Jim Clark, Lajos Horváth, and Mark Lewis. On the estimation of spread rate for a biological population. Statistics & Probability Letters, 51(3):225–234, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001231.

Chen:2008:NCF

[CHL08] Xiaohong Chen, Yingyao Hu, and Arthur Lewbel. A note on the closed-form identification of regression models with a mismeasured binary regressor. Statistics & Probability Letters, 78(12):1473–1479, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004464.

Cui:2004:MEL

[CHN04] Hengjian Cui, Xuming He, and Kai W. Ng. M-estimation for linear models with spatially-correlated errors. Statistics & Probability Letters, 66(4):383–393, March 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003304.

Choi:2005:IBN

[Cho05a] In Choi. Inconsistency of bootstrap for nonstationary, vector autoregressive processes. Statistics & Probability Letters, 75(1):39–48, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002087.

Choulakian:2005:TIP

[Cho05b] Vartan Choulakian. Transposition invariant principal component analysis in L_{1} for long tailed data. Statistics & Probability Letters, 71(1):23–31, January 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002792.

Choi:2008:SRS

[Cho08a] Gyeong Suk Choi. Some results on subordination, selfdecomposability and operator semi-stability. Statistics & Probability Letters, 78(6):780–784, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700332X.

Chopin:2008:EBS

[Cho08b] N. Chopin. On the equivalence between standard and sequentially ordered hidden Markov models. Statistics & Probability Letters, 78(14):2171–2174, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001065.

Christofides:2000:MID

[Chr00] Tasos C. Christofides. Maximal inequalities for demimartingales and a Strong Law of Large Numbers. Statistics & Probability Letters, 50(4):357–363, December 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001164.

Chen:2009:LBM

[CHV09] Pingyan Chen, Tien-Chung Hu, and Andrei Volodin. Limiting behaviour of moving average processes under ϕ-mixing assumption. Statistics & Probability Letters, 79(1):105–111, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003519.

Coeurjolly:2001:CRB

[CI01] Jean-François Coeurjolly and Jacques Istas. Cramèr-Rao bounds for fractional Brownian motions. Statistics & Probability Letters, 53(4):435–447, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001978.

Christensen:2009:NPC

[CI09] Sören Christensen and Albrecht Irle. A note on pasting conditions for the American perpetual optimal stopping problem. Statistics & Probability Letters, 79(3):349–353, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004185.

Chang:2009:CBL

[cICP09] Yuan chin Ivan Chang and Eunsik Park. Constructing the best linear combination of diagnostic markers via sequential sampling. Statistics & Probability Letters, 79(18):1921–1927, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002016.

Ciuperca:2009:EMP

[Ciu09] Gabriela Ciuperca. The M-estimation in a multi-phase random nonlinear model. Statistics & Probability Letters, 79 (5):573–580, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004689.

Cha:2001:ABI

[CK01] Ji Hwan Cha and Jae Joo Kim. On availability of Bayesian imperfect repair model. Statistics & Probability Letters, 53(2):181–187, June 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000761.

Chaturvedi:2005:BUR

[CK05a] Anoop Chaturvedi and Jitendra Kumar. Bayesian unit root test for model with maintained trend. Statistics & Probability Letters, 74(2):109–115, September 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001525.

Chen:2005:IUL

[CK05b] Zengjing Chen and Reg Kulperger. Inequalities for upper and lower probabilities. Statistics & Probability Letters, 73(3):233– 241, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001070.

Cui:2006:DRS

[CKLX06] Lirong Cui, Way Kuo, Jinlin Li, and Min Xie. On the dual reliability systems of (n, f, k) and ⟨n, f, k⟩. Statistics & Probability Letters, 76(11):1081–1088, June 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004499.

Chang:2003:PMP

[CKM03] In Hong Chang, Byung Hwee Kim, and Rahul Mukerjee. Probability matching priors for predicting unobservable random effects with application to ANOVA models. Statistics & Probability Letters, 62(3):223–228, April 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300004X.

Cramer:2002:EMG

[CKR02] Erhard Cramer, Udo Kamps, and Tomasz Rychlik. On the existence of moments of generalized order statistics. Statistics & Probability Letters, 59(4):397–404, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002389.

Cohen:2004:FAP

[CKS04] A. Cohen, J. Kolassa, and H. B. Sackrowitz. A four action problem with ordered categorical data: are two distributions the same, ordered, or otherwise? Statistics & Probability Letters, 70(3):223–234, December 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002597.

Cano:2006:APD

[CKS06a] J. A. Cano, M. Kessler, and D. Salmerón. Approximation of the posterior density for diffusion processes. Statistics & Probability Letters, 76(1):39–44, January 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500266X.

Cohen:2006:NTS

[CKS06b] Arthur Cohen, John Kolassa, and H. B. Sackrowitz. A new test for stochastic order of k ≥ 3 ordered multinomial populations. Statistics & Probability Letters, 76(10):1017–1024, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205004426.

Chafai:2006:NML

[CL06a] Djalil Chafaï and Jean-Michel Loubes. On nonparametric maximum likelihood for a class of stochastic inverse problems. Statistics & Probability Letters, 76(12):1225–1237, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004748.

Chen:2006:ASM

[CL06b] Shouquan Chen and Zhengyan Lin. Almost sure maxlimits for nonstationary Gaussian sequence. Statistics & Probability Letters, 76(11):1175–1184, June 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004724.

Chou:2006:AOJ

[CL06c] Ching-Sung Chou and Hsien-Jen Lin. Asian options with jumps. Statistics & Probability Letters, 76(18):1983–1993, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001829.

Chou:2007:PMZ

[CL07] Ching-Sung Chou and Hsien-Jen Lin. Pricing model for zero coupon bonds driven by Bessel-squared interest processes with a jump. Statistics & Probability Letters, 77(5):475–482, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002677.

Chen:2008:OML

[CL08a] Jie Chen and Min-Qian Liu. Optimal mixed-level supersaturated design with general number of runs. Statistics & Probability Letters, 78(15):2496–2502, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001557.

Chen:2008:ASF

[CL08b] Shouquan Chen and Zhengyan Lin. Almost sure functional Central Limit Theorems for weakly dependent sequences. Statistics & Probability Letters, 78(13):1683–1693, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000266.

Chiquet:2008:MCT

[CL08c] Julien Chiquet and Nikolaos Limnios. A method to compute the transition function of a piecewise deterministic Markov process with application to reliability. Statistics & Probability Letters, 78(12):1397–1403, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700421X.

Chiragiev:2009:MFP

[CL09] Arthur Chiragiev and Zinoviy Landsman. Multivariate flexible Pareto model: Dependency structure, properties and characterizations. Statistics & Probability Letters, 79(16):1733–1743, August 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001618.

Clemencon:2001:MPI

[Clé01] S. J. M. Clémençon. Moment and probability inequalities for sums of bounded additive functionals of regular Markov chains via the Nummelin splitting technique. Statistics & Probability Letters, 55(3):227–238, December 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002364.

Chen:2006:SNC

[CLF06] Neiping Chen, Wenbin Liu, and Jianfeng Feng. Sufficient and necessary condition for the convergence of stochastic approximation algorithms. Statistics & Probability Letters, 76(2):203–210, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002890.

Collet:2003:ORM

[CLM03] Pierre Collet, F. Javier López, and Servet Martínez. Order relations of measures when avoiding decreasing sets. Statistics & Probability Letters, 65(3):165–175, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002232.

Cammarota:2009:SDS

[CLO09] V. Cammarota, A. Lachal, and E. Orsingher. Some Darling– Siegert relationships connected with random flights. Statistics & Probability Letters, 79(2):243–254, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003842.

Chiu:2002:BCA

[CLR02] Grace Chiu, Richard Lockhart, and Richard Routledge. Bentcable asymptotics when the bend is missing. Statistics & Probability Letters, 59(1):9–16, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001384.

Chen:2006:GHR

[CLR06] Anyue Chen, Junping Li, and N. I. Ramesh. General Harris regularity criterion for non-linear Markov branching processes. Statistics & Probability Letters, 76(5):446–452, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520500324X.

Chen:2008:PAW

[CLR08] Anyue Chen, Junping Li, and N. I. Ramesh. Probabilistic approach in weighted Markov branching processes. Statistics & Probability Letters, 78(6):771–779, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003318.

Christophi:2001:DSR

[CM01] Costas A. Christophi and Hosam M. Mahmoud. Distribution of the size of random hash trees, pebbled hash trees and N-trees. Statistics & Probability Letters, 53(3):277–282, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000712.

Christofides:2003:CLT

[CM03] Tasos C. Christofides and Petroula M. Mavrikiou. Central limit theorem for dependent multidimensionally indexed random variables. Statistics & Probability Letters, 63(1):67–78, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000543.

Chang:2004:ARF

[CM04a] In Hong Chang and Rahul Mukerjee. Asymptotic results on the frequentist mean squared error of generalized Bayes point predictors. Statistics & Probability Letters, 67(1):65–71, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000094.

Csiszar:2004:CMP

[CM04b] Villő Csiszár and Tamás F. Móri. The convexity method of proving moment-type inequalities. Statistics & Probability Letters, 66(3):303–313, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003547.

Chang:2006:PMP

[CM06] In Hong Chang and Rahul Mukerjee. Probability matching property of adjusted likelihoods. Statistics & Probability Letters, 76(8):838–842, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004001.

Chen:2007:NPH

[CM07] Shande Chen and Amita K. Manatunga. A note on proportional hazards and proportional odds models. Statistics & Probability Letters, 77(10):981–988, June 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000375.

Chang:2008:MPF

[CM08a] In Hong Chang and Rahul Mukerjee. Matching posterior and frequentist cumulative distribution functions with empiricaltype likelihoods in the multiparameter case. Statistics & Probability Letters, 78(16):2793–2797, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002034.

Comas:2008:STD

[CM08b] C. Comas and J. Mateu. Space-time dependence dynamics for birth-death point processes. Statistics & Probability Letters, 78(16):2715–2719, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001909.

Cline:2002:NSM

[CmHP02] Daren B. H. Cline and Huay min H. Pu. A note on a simple Markov bilinear stochastic process. Statistics & Probability Letters, 56(3):283–288, February 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001924.

Chacon:2007:SUP

[CMNP07] J. E. Chacón, J. Montanero, A. G. Nogales, and P. Pérez. Stability under products of sufficient, minimal sufficient and complete σ-fields in the Bayesian case. Statistics & Probability Letters, 77(7):710–716, April 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003403.

Csiszar:2005:CTI

[CMS05] Villó Csiszár, Tamás F. Móri, and Gábor J. Székely. Chebyshev-type inequalities for scale mixtures. Statistics & Probability Letters, 71(4):323–335, March 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000040.

Choi:2003:SBC

[CN03] Changsun Choi and Dougu Nam. Some boundary-crossing results for linear diffusion processes. Statistics & Probability Letters, 62(3):281–291, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000154.

Chen:2006:MRW

[CNX06] Yiqing Chen, Kai W. Ng, and Xiangsheng Xie. On the maximum of randomly weighted sums with regularly varying tails. Statistics & Probability Letters, 76(10):971–975, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205004177.

Coquet:2000:SIS

[CO00] F. Coquet and Y. Ouknine. Some identities on semimartingales local times. Statistics & Probability Letters, 49(2): 149–153, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000432.

Choi:2002:NCS

[CÖ02] Young Hun Choi and Ömer Öztürk. A new class of score generating functions for regression models. Statistics & Probability Letters, 57(2):205–214, April 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000615.

Chatterjee:2006:NEQ

[CO06] Samprit Chatterjee and Ingram Olkin. Nonparametric estimation for quadratic regression. Statistics & Probability Letters, 76(11):1156–1163, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004694.

Coeurjolly:2008:BRS

[Coe08] Jean-François Coeurjolly. Bahadur representation of sample quantiles for functional of Gaussian dependent sequences under a minimal assumption. Statistics & Probability Letters, 78(15):2485–2489, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208001533.

Colangelo:2008:SLR

[Col08] Antonio Colangelo. A study on LTD and RTI positive dependence orderings. Statistics & Probability Letters, 78(14): 2222–2229, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001144.

Comman:2007:VFL

[Com07] Henri Comman. Variational form of the large deviation functional. Statistics & Probability Letters, 77(9):931–936, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000363.

Conti:2005:NST

[Con05] Pier Luigi Conti. A nonparametric sequential test with power 1 for the ruin probability in some risk models. Statistics & Probability Letters, 72(4):333–343, May 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500074X.

Cook:2002:CSD

[Coo02] Steven Cook. Correcting size distortion of the Dickey– Fuller test via recursive mean adjustment. Statistics & Probability Letters, 60(1):75–79, November 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002808.

Cook:2003:MUR

[Coo03] Steven Cook. Modified unit root tests and momentum threshold autoregressive processes. Statistics & Probability Letters, 64(1):83–88, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001330.

Cook:2004:MTA

[Coo04] Steven Cook. A momentum-threshold autoregressive unit root test with increased power. Statistics & Probability Letters, 67(4):307–310, May 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000598.

Cook:2006:TCP

[Coo06] Steven Cook. Testing for cointegration in the presence of mis-specified structural change. Statistics & Probability Letters, 76(13):1380–1384, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000459.

Cordeiro:2004:PRG

[Cor04a] Gauss M. Cordeiro. On Pearson’s residuals in generalized linear models. Statistics & Probability Letters, 66(3):213– 219, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002566.

Cordeiro:2004:SOC

[Cor04b] Gauss M. Cordeiro. Second-order covariance matrix of maximum likelihood estimates in generalized linear models. Statistics & Probability Letters, 66(2):153–160, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003043.

Corcuera:2008:APP

[Cor08] José M. Corcuera. Approximate predictive pivots for autoregressive processes. Statistics & Probability Letters, 78 (16):2685–2691, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001855.

Cai:2003:LEN

[COS03] Zongwu Cai and Elias Ould-Saïd. Local M-estimator for nonparametric time series. Statistics & Probability Letters, 65(4):433–449, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002992.

Chen:2000:GDI

[CP00] Zengjing Chen and Shige Peng. A general downcrossing inequality for g-martingales. Statistics & Probability Letters, 46(2):169–175, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001029.

Coles:2001:ELL

[CP01] Stuart Coles and Francesco Pauli. Extremal limit laws for a class of bivariate Poisson vectors. Statistics & Probability Letters, 54(4):373–379, October 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100102X.

Ceyhan:2005:UDN

[CP05a] Elvan Ceyhan and Carey E. Priebe. The use of domination number of a random proximity catch digraph for testing spatial patterns of segregation and association. Statistics & Probability Letters, 73(1):37–50, June 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000805.

Chaudhuri:2005:BML

[CP05b] Sanjay Chaudhuri and Michael D. Perlman. Biases of the maximum likelihood and Cohen–Sackrowitz estimators for the tree-order model. Statistics & Probability Letters, 71(3):267–276, March 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003086.

Crimaldi:2005:TIC

[CP05c] Irene Crimaldi and Luca Pratelli. Two inequalities for conditional expectations and convergence results for filters. Statistics & Probability Letters, 74(2):151–162, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001719.

Chan:2006:CTM

[CP06a] Wenyaw Chan and Nan Fu Peng. Continuous time Markov chains observed on an alternating renewal process with exponentially distributed durations. Statistics & Probability Letters, 76(4):362–368, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500307X.

Chen:2006:OCP

[CP06b] Zhiqiang Chen and Eswar Phadia. An optimal completion of the product limit estimator. Statistics & Probability Letters, 76(9):913–919, May 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004086.

Cabilio:2008:MRB

[CP08] P. Cabilio and J. Peng. Multiple rank-based testing for ordered alternatives with incomplete data. Statistics & Probability Letters, 78(16):2609–2613, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003830.

Coulon-Prieur:2000:TCL

[CPD00] Clémentine Coulon-Prieur and Paul Doukhan. A triangular Central Limit Theorem under a new weak dependence condition. Statistics & Probability Letters, 47(1):61–68, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001388.

Cysneiros:2007:HSL

[CPG07] Francisco José A. Cysneiros, Gilberto A. Paula, and Manuel Galea. Heteroscedastic symmetrical linear models. Statistics & Probability Letters, 77(11):1084–1090, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000466.

Chen:2003:ELB

[CQ03] Song Xi Chen and Jing Qin. Empirical likelihood-based confidence intervals for data with possible zero observations. Statistics & Probability Letters, 65(1):29–37, October 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002165.

Chatterjee:2008:NLD

[CQ08] Kashinath Chatterjee and Hong Qin. A new look at discrete discrepancy. Statistics & Probability Letters, 78(17): 2988–2991, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520800254X.

Chen:2008:RBR

[CQX08] Yong Chen, Min-Ping Qian, and Jian-Sheng Xie. On the relation between reversibility and monotonicity of fluctuation spectra for discrete time finite state Markov chains. Statistics & Probability Letters, 78(14):2258–2264, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001193.

Charlot:2008:SPS

[CR08] F. Charlot and M. Rachdi. On the statistical properties of a stationary process sampled by a stationary point process. Statistics & Probability Letters, 78(4):456–462, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002611.

Cramer:2004:LUP

[Cra04] Erhard Cramer. Logconcavity and unimodality of progressively censored order statistics. Statistics & Probability Letters, 68(1):83–90, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000720.

Crimaldi:2004:BCE

[Cri04] Irene Crimaldi. On the behavior of the conditional expectations in Skorohod representation theorem. Statistics & Probability Letters, 67(2):141–148, April 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000197.

Chen:2000:BDR

[CS00a] Ling Chen and Jun Shao. A Bayesian decision rule for remediation actions at toxic waste sites. Statistics & Probability Letters, 50(1):83–88, October 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000857.

Christoph:2000:SSD

[CS00b] Gerd Christoph and Karina Schreiber. Scaled Sibuya distribution and discrete self-decomposability. Statistics & Probability Letters, 48(2):181–187, June 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299002035.

Cohen:2000:PBT

[CS00c] Arthur Cohen and H. B. Sackrowitz. Properties of Bayes testing procedures in order restricted inference. Statistics & Probability Letters, 49(2):205–209, August 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000493.

Cohen:2001:WST

[CS01] Arthur Cohen and Harold B. Sackrowitz. Wherefore similar tests? Statistics & Probability Letters, 54(3):283–290, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000694.

Carvalho:2005:EEM

[CS05a] Alexandre Carvalho and Georgios Skoulakis. Ergodicity and existence of moments for local mixtures of linear autoregressions. Statistics & Probability Letters, 71(4):313–322, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000027.

Clarke:2005:MLE

[CS05b] Paul S. Clarke and Peter W. F. Smith. On maximum likelihood estimation for log-linear models with non-ignorable non-response. Statistics & Probability Letters, 73(4):441–448, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001483.

Chen:2006:FTR

[CS06] Yu Chen and Chun Su. Finite time ruin probability with heavy-tailed insurance and financial risks. Statistics & Probability Letters, 76(16):1812–1820, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001477.

Csorgo:2007:SPG

[CS07] Sándor Csörgő and Gordon Simons. St. Petersburg games with the largest gains withheld. Statistics & Probability Letters, 77(12):1185–1189, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000806.

Cai:2008:MSC

[CS08a] Gengqian Cai and Sanat K. Sarkar. Modified Simes’ critical values under independence. Statistics & Probability Letters, 78(12):1362–1368, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004166.

Charpentier:2008:CAC

[CS08b] Arthur Charpentier and Johan Segers. Convergence of Archimedean copulas. Statistics & Probability Letters, 78 (4):412–419, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002568.

Cheng:2008:GFT

[CS08c] Fuxia Cheng and Shuxia Sun. A goodness-of-fit test of the errors in nonlinear autoregressive time series models. Statistics & Probability Letters, 78(1):50–59, January 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001940.

Csorgo:2003:RCC

[Csö03] Sándor Csörgő. Rates in the complete convergence of bootstrap means. Statistics & Probability Letters, 64(4):359–368, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001809.

Chen:2004:FSB

[CT04] Zhiqiang Chen and David E. Tyler. On the finite sample breakdown points of redescending M-estimates of location. Statistics & Probability Letters, 69(3):233–242, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001622.

Cheng:2002:AMD

[CTY02] Yebin Cheng, Qihe Tang, and Hailiang Yang. Approximations for moments of deficit at ruin with exponential and subexponential claims. Statistics & Probability Letters, 59(4):367–378, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002341.

Cui:2002:IPC

[Cui02] Lirong Cui. The IFR property for consecutive-k-out-of-n:f systems. Statistics & Probability Letters, 59(4):405–414, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002407.

Cabrera:2001:CCU

[CV01] Manuel Ordóñez Cabrera and Andrei I. Volodin. On conditional compactly uniform p th-order integrability of random elements in Banach spaces. Statistics & Probability Letters, 55(3):301–309, December 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001596.

Cook:2004:FSS

[CV04] Steven Cook and Dimitrios Vougas. On the finite-sample size distortion of smooth transition unit root tests. Statistics & Probability Letters, 70(3):175–182, December 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002524.

Cysneiros:2008:RTS

[CV08] Francisco José A. Cysneiros and Luis Hernando Vanegas. Residuals and their statistical properties in symmetrical nonlinear models. Statistics & Probability Letters, 78(18):3269–3273, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800312X.

Cabrera:2009:CRS

[CV09] Manuel Ordóñez Cabrera and Andrei Volodin. On Cantrell– Rosalsky’s Strong Laws of Large Numbers. Statistics & Probability Letters, 79(6):842–847, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005245.

Cheng:2001:ALI

[CW01] Bo Cheng and Xizhi Wu. Assessing local influence in PLS regression by the second order approach. Statistics & Probability Letters, 53(2):113–121, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000104.

Choy:2003:EEP

[CW03] S. T. Boris Choy and Stephen G. Walker. The extended exponential power distribution and Bayesian robustness. Statistics & Probability Letters, 65(3):227–232, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300258X.

Cai:2005:MAP

[CW05] Jun Cai and Gordon E. Willmot. Monotonicity and aging properties of random sums. Statistics & Probability Letters, 73(4):381–392, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001422.

Chongcharoen:2007:PSI

[CW07] Samruam Chongcharoen and F. T. Wright. Permutation and scale invariant one-sided approximate likelihood ratio tests. Statistics & Probability Letters, 77(8):774–781, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000028.

Carbonell:2008:GGC

[CW08a] F. Carbonell and K. J. Worsley. On the geometry of a generalized cross-correlation random field. Statistics & Probability Letters, 78(18):3129–3134, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002800.

Coppersmith:2008:CWE

[CW08b] Don Coppersmith and Chai Wah Wu. Conditions for weak ergodicity of inhomogeneous Markov chains. Statistics & Probability Letters, 78(17):3082–3085, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800271X.

Chen:2008:OMD

[CWL08] Ray-Bing Chen, Weng Kee Wong, and Kun-Yu Li. Optimal minimax designs over a prespecified interval in a heteroscedastic polynomial model. Statistics & Probability Letters, 78(13):1914–1921, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000588.

Carbonell:2009:FTC

[CWTB09] F. Carbonell, K. J. Worsley, and N. J. Trujillo-Barreto. On the Fisher’s Z transformation of correlation random fields. Statistics & Probability Letters, 79(6):780–788, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005154.

Chen:2008:NFC

[CWZ08] Dayue Chen, Bei Wei, and Fuxi Zhang. A note on the finite collision property of random walks. Statistics & Probability Letters, 78(13):1742–1747, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000357.

Chen:2009:LCG

[CXH09] Huaihou Chen, Hongmei Xie, and Taizhong Hu. Logconcavity of generalized order statistics. Statistics & Probability Letters, 79(3):396–399, February 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004264.

Chiu:2002:FET

[CY02a] Sung Nok Chiu and Chuan Cun Yin. The first exit time and ruin time for a risk process with reserve-dependent income. Statistics & Probability Letters, 60(4):417–424, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003115.

Cook:2002:ADT

[CY02b] R. Dennis Cook and Xiangrong Yin. Asymptotic distributions for testing dimensionality in q-based pHd. Statistics & Probability Letters, 58(3):233–243, July 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000925.

Cho:2001:ATB

[CYJL01a] Kwanho Cho, In-Kwon Yeo, Richard A. Johnson, and Wei-Yin Loh. Asymptotic theory for Box–Cox transformations in linear models. Statistics & Probability Letters, 51(4):337–343, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520000170X.

Cho:2001:PIE

[CYJL01b] Kwanho Cho, In-Kwon Yeo, Richard A. Johnson, and Wei-Yin Loh. Prediction interval estimation in transformed linear models. Statistics & Probability Letters, 51(4):345–350, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001693.

Chen:2007:LDR

[CZ07] Yu Chen and Weiping Zhang. Large deviations for random sums of negatively dependent random variables with consistently varying tails. Statistics & Probability Letters, 77(5):530–538, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002756.

DAngio:2006:TRV

[D’A06] Roberto D’Angiò. Three random variable transformations giving Heisenberg-type uncertainty relations. Statistics & Probability Letters, 76(4):423–430, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003160.

Dahl:2006:CPP

[Dah06] Fredrik Andreas Dahl. On the conservativeness of posterior predictive p-values. Statistics & Probability Letters, 76 (11):1170–1174, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004712.

Dai:2000:RRC

[Dai00] Jack Jie Dai. A result regarding convergence of random logistic maps. Statistics & Probability Letters, 47(1):11–14, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001315.

Dai:2003:NVR

[Dai03] Jack Jie Dai. A note on vertex-reinforced random walks. Statistics & Probability Letters, 62(3):275–280, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000142.

Dai:2004:CVR

[Dai04a] Jack Jie Dai. A computable version of the random signs problem and Kolmogorov complexity. Statistics & Probability Letters, 67(1):27–31, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003870.

Dai:2004:SRR

[Dai04b] Jack Jie Dai. Some results regarding vertex-reinforced random walks. Statistics & Probability Letters, 66(3):259–266, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520300333X.

Dai:2005:OER

[Dai05] Jack Jie Dai. A Once edge-reinforced random walk on a Galton–Watson tree is transient. Statistics & Probability Letters, 73(2):115–124, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000982.

Darwich:2001:AAC

[Dar01] A. R. Darwich. About the absolute continuity and orthogonality for two probability measures. Statistics & Probability Letters, 52(1):1–8, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001735.

Daria:2003:RST

[Dar03] Loukianova Daria. Remark on semigroup techniques and the maximum likelihood estimation. Statistics & Probability Letters, 62(2):111–115, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202004315.

Daras:2004:LDE

[Dar04] Tryfon Daras. Large deviations for the empirical process of a symmetric measure: a lower bound. Statistics & Probability Letters, 66(2):197–204, January 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003390.

Davidson:2004:FMS

[Dav04] James Davidson. Forecasting Markov-switching dynamic, conditionally heteroscedastic processes. Statistics & Probability Letters, 68(2):137–147, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000628.

Davidov:2005:ESM

[Dav05] Ori Davidov. Estimating the slope in measurement error models — a different perspective. Statistics & Probability Letters, 71(3):215–223, March 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003037.

Duan:2008:GMV

[DaY08] Jinqiao Duan and Jia an Yan. General matrix-valued inhomogeneous linear stochastic differential equations and applications. Statistics & Probability Letters, 78(15):2361–2365, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800134X.

Dai:2009:IPW

[DB09] Hongsheng Dai and Yanchun Bao. An inverse probability weighted estimator for the bivariate distribution function under right censoring. Statistics & Probability Letters, 79(16):1789–1797, August 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001825.

Castro:2006:PDT

[dBC06] Tomas del Barrio Castro. On the performance of the DHF tests against nonstationary alternatives. Statistics & Probability Letters, 76(3):291–297, February 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002993.

Daly:2007:PRM

[DBS07] David Daly, Peter Buchholz, and William H. Sanders. A preorder relation for Markov reward processes. Statistics & Probability Letters, 77(11):1148–1157, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000673.

Dykstra:2001:DSO

[DC01] Richard Dykstra and Yeh-Fong Chen. The distribution of sizes of ordered level sets. Statistics & Probability Letters, 53(4):339–346, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001802.

Bueno:2000:ARL

[dCB00a] Vanderlei da Costa Bueno. Asymptotic reliability of a linearly connected system with an infinite number of components. Statistics & Probability Letters, 50(1):55–62, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000821.

Bueno:2000:CIR

[dCB00b] Vanderlei da Costa Bueno. Component importance in a random environment. Statistics & Probability Letters, 48(2): 173–179, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299002011.

delCastillo:2009:EGP

[dCD09] Joan del Castillo and Jalila Daoudi. Estimation of the generalized Pareto distribution. Statistics & Probability Letters, 79(5):684–688, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004987.

Dharamadhikari:2006:ATV

[DD06] A. D. Dharamadhikari and Isha Dewan. Association in time of a vector valued process. Statistics & Probability Letters, 76(11):1147–1155, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004682.

Drees:2003:LDE

[DdHL03] Holger Drees, Laurens de Haan, and Deyuan Li. On large deviation for extremes. Statistics & Probability Letters, 64 (1):51–62, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001305.

Das:2006:AFC

[DDM06] Ashish Das, Aloke Dey, and Chand K. Midha. Allocating factors to the columns of an orthogonal array when certain interactions are important. Statistics & Probability Letters, 76(14):1570–1577, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001003.

Diongue:2008:SFA

[DDN08] Abdou Kâ Diongue, Aliou Diop, and Mor Ndongo. Seasonal fractional ARIMA with stable innovations. Statistics & Probability Letters, 78(12):1404–1411, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004221.

DelGrecoM:2008:NCE

[DDP08] Fabiola Del Greco M., Marco Di Marzio, and Agnese Panzera. A new class of excited random walks on trees. Statistics & Probability Letters, 78(13):1981–1989, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000680.

DeMartini:2008:RPE

[De 08] Daniele De Martini. Reproducibility probability estimation for testing statistical hypotheses. Statistics & Probability Letters, 78(9):1056–1061, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003653.

Edwardes:2000:VLIa

[dE00a] Michael D. deB. Edwardes. Is variance larger if and only if tails are larger? Statistics & Probability Letters, 47(2):141– 147, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001509. See correc-

tion [dE00b].

Edwardes:2000:VLIb

[dE00b] Michael D. deB. Edwardes. Is variance larger if and only if tails are larger? [Statist. Probab. Lett. 47 (2000) 141– 147]. Statistics & Probability Letters, 50(2):203, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001267. See [dE00a].

Davydov:2006:CEP

[DE06] Youri Davydov and Vladimir Egorov. On convergence of empirical point processes. Statistics & Probability Letters, 76(17):1836–1844, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001489.

Debowski:2007:PSP

[Deb07] Lukasz Debowski. On processes with summable partial autocorrelations. Statistics & Probability Letters, 77(7):752–759, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206003476.

Debowski:2009:GDC

[Deb09] Lukasz Debowski. A general definition of conditional information and its application to ergodic decomposition. Statistics & Probability Letters, 79(9):1260–1268, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900042X.

Degras:2008:ANE

[Deg08] David Degras. Asymptotics for the nonparametric estimation of the mean function of a random process. Statistics & Probability Letters, 78(17):2976–2980, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002526.

Deheuvels:2007:KLE

[Deh07] Paul Deheuvels. A Karhunen–Loève expansion for a mean-centered Brownian bridge. Statistics & Probability Letters, 77(12):1190–1200, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000831.

Delmas:2003:PSC

[Del03] Céline Delmas. Projections on spherical cones, maximum of Gaussian fields and Rice’s method. Statistics & Probability Letters, 64(3):263–270, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001706.

Dembinska:2008:RGD

[Dem08] Anna Dembińska. k th records from geometric distribution. Statistics & Probability Letters, 78(12):1662–1670, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800028X.

Deng:2003:SNB

[Den03] Dianliang Deng. On the self-normalized bounded laws of iterated logarithm in Banach space. Statistics & Probability Letters, 64(3):277–286, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300172X.

Deng:2005:CCS

[Den05] Dianliang Deng. The complete convergence of subsequence for sums of independent B-valued random variables. Statistics & Probability Letters, 74(3):235–244, October 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001276.

Deng:2007:SNW

[Den07] Dianliang Deng. Self-normalized Wittmann’s laws of iterated logarithm in Banach space. Statistics & Probability Letters, 77(6):632–643, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002975.

Dette:2000:NTL

[Det00] Holger Dette. On a nonparametric test for linear relationships. Statistics & Probability Letters, 46(3):307–316, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001303.

Devroye:2002:SBR

[Dev02] Luc Devroye. Simulating Bessel random variables. Statistics & Probability Letters, 57(3):249–257, April 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200055X.

Devulder:2007:SBS

[Dev07] Alexis Devulder. The speed of a branching system of random walks in random environment. Statistics & Probability Letters, 77(18):1712–1721, December 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001563.

Devroye:2009:ESA

[Dev09] Luc Devroye. On exact simulation algorithms for some distributions related to Jacobi theta functions. Statistics & Probability Letters, 79(21):2251–2259, November 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002867.

Dey:2005:PPS

[Dey05] Aloke Dey. Projection properties of some orthogonal arrays. Statistics & Probability Letters, 75(4):298–306, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002440.

Dzhaparidze:2002:FDA

[DF02] K. Dzhaparidze and J. A. Ferreira. A frequency domain approach to some results on fractional Brownian motion. Statistics & Probability Letters, 60(2):155–168, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202003073.

Duncan:2011:CPS

[DF11] Tyrone E. Duncan and Holger Fink. Corrigendum to “Prediction for some processes related to a fractional Brownian motion”. Statistics & Probability Letters, 81(8):1336–1337, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001192. See [Dun06].

daFonseca:2002:LBM

[dFF02] V. Grunert da Fonseca and C. M. Fonseca. A link between the multivariate cumulative distribution function and the hitting function for random closed sets. Statistics & Probability Letters, 57(2):179–182, April 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000469.

Dabye:2003:BEM

[DFK03] Ali S. Dabye, Christian Farinetto, and Yury A. Kutoyants. On Bayesian estimators in misspecified change-point problems for Poisson process. Statistics & Probability Letters, 61(1):17–30, January 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002006.

Durante:2008:TCP

[DFS08] Fabrizio Durante, Rachele Foschi, and Fabio Spizzichino. Threshold copulas and positive dependence. Statistics & Probability Letters, 78(17):2902–2909, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002344.

Dawson:2005:SUC

[DFX05] Donald A. Dawson, Klaus Fleischmann, and Jie Xiong. Strong uniqueness for cyclically symbiotic branching diffusions. Statistics & Probability Letters, 73(3):251–257, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001094.

Diebolt:2003:NAN

[DG03] Jean Diebolt and Stéphane Girard. A note on the asymptotic normality of the ET method for extreme quantile estimation. Statistics & Probability Letters, 62(4):397–405, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000452.

Daley:2006:MIM

[DG06a] D. J. Daley and Charles M. Goldie. The moment index of minima (II). Statistics & Probability Letters, 76(8):831–837, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003937.

Du:2006:PLH

[DG06b] Pang Du and Chong Gu. Penalized likelihood hazard estimation: Efficient approximation and Bayesian confidence intervals. Statistics & Probability Letters, 76(3):244–254, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002932.

Diongue:2007:SSH

[DG07] Abdou Kâ Diongue and Dominique Guégan. The stationary seasonal hyperbolic asymmetric power ARCH model. Statistics & Probability Letters, 77(11):1158–1164, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000697.

Davidov:2008:NCE

[DG08] Ori Davidov and Vladimir Griskin. A note on constrained estimation in the simple linear measurement error model. Statistics & Probability Letters, 78(5):508–517, April 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002829.

Dettmann:2009:PIU

[DG09] Carl P. Dettmann and Orestis Georgiou. Product of n independent uniform random variables. Statistics & Probability Letters, 79(24):2501–2503, December 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003459.

Donauer:2009:GRR

[DGJ09] Stefanie Donauer, Piet Groeneboom, and Geurt Jongbloed. Global rate results for the MLE in a class of deconvolution models. Statistics & Probability Letters, 79(4):519–524, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004604.

DiCrescenzo:2008:NBD

[DGNR08] A. Di Crescenzo, V. Giorno, A. G. Nobile, and L. M. Ricciardi. A note on birth-death processes with catastrophes. Statistics & Probability Letters, 78(14):2248–2257, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001181.

DeGooijer:2002:MSE

[DGZ02] Jan G. De Gooijer, Ali Gannoun, and Dawit Zerom. Mean squared error properties of the kernel-based multistage median predictor for time series. Statistics & Probability Letters, 56(1):51–56, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001699.

Delaigle:2006:OKC

[DH06] Aurore Delaigle and Peter Hall. On optimal kernel choice for deconvolution. Statistics & Probability Letters, 76(15):1594–1602, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001155.

Debruyne:2009:IFS

[DH09] M. Debruyne and M. Hubert. The influence function of the Stahel–Donoho covariance estimator of smallest outlyingness. Statistics & Probability Letters, 79(3):275–282, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003970.

deHaan:2006:CDF

[dHC06] Laurens de Haan and Luisa Canto e Castro. A class of distribution functions with less bias in extreme value estimation. Statistics & Probability Letters, 76(15):1617–1624, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001143.

Date:2000:NES

[DHS00] Akira Date, Chii-Ruey Hwang, and Shuenn-Jyi Sheu. On the number of equilibrium states in weakly coupled random networks. Statistics & Probability Letters, 49(3):291– 297, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000602.

DiCrescenzo:2000:SRP

[Di 00] Antonio Di Crescenzo. Some results on the proportional reversed hazards model. Statistics & Probability Letters, 50(4):313–321, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001279.

Dindar:2000:RFG

[Din00a] Zacharie Dindar. Random fractals generated by oscillations of the uniform empirical process. Statistics & Probability Letters, 50(3):285–291, November 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000105X.

Ding:2000:CCC

[Din00b] A. Adam Ding. Can coverage of confidence intervals be used to judge the goodness of point estimators in finite population double sampling? Statistics & Probability Letters, 49(3):251–256, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000547.

Dinwoodie:2000:CEN

[Din00c] I. H. Dinwoodie. Conditional expectations in network traffic estimation. Statistics & Probability Letters, 47(1):99–103, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001431.

Dindar:2003:SMR

[Din03] Zacharie Dindar. Some more results on increments of the partially observed empirical process. Statistics & Probability Letters, 64(1):25–37, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001275.

Ditlevsen:2007:RFP

[Dit07] Susanne Ditlevsen. A result on the first-passage time of an Ornstein–Uhlenbeck process. Statistics & Probability Letters, 77(18):1744–1749, December 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001605.

deJong:2004:AMS

[dJP04] Piet de Jong and Jeremy Penzer. The ARMA model in state space form. Statistics & Probability Letters, 70(1):119–125, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002330.

Dreier:2002:NCF

[DK02] I. Dreier and S. Kotz. A note on the characteristic function of the t-distribution. Statistics & Probability Letters, 57(3):221–224, April 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000329.

Divanji:2005:LIR

[DK05] Gooty Divanji and Tadewos Koroto. Limit infimum results for subsequences of partial sums and random sums. Statistics & Probability Letters, 71(2):123–129, February 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002615.

Denker:2007:ST

[DK07] M. Denker and N. Kan. On Sevast’yanov’s theorem. Statistics & Probability Letters, 77(3):272–279, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002331.

Debicki:2008:NUE

[DK08] Krzysztof Debicki and Paweł Kisowski. A note on upper estimates for Pickands constants. Statistics & Probability Letters, 78(14):2046–2051, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800076X.

Dharmadhikari:2002:AEM

[DKM02] Avinash Dharmadhikari, S. B. Kulathinal, and Vidyadhar Mandrekar. An algorithm for the estimation of minimal cut and path sets from field failure data. Statistics & Probability Letters, 58(1):1–11, May 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200086X.

Luna:2001:RNH

[dL01] Sara Sjöstedt de Luna. Resampling non-homogeneous spatial data with smoothly varying mean values. Statistics & Probability Letters, 53(4):373–379, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000402.

DiCrescenzo:2004:MDB

[DL04] Antonio Di Crescenzo and Maria Longobardi. A measure of discrimination between past lifetime distributions. Statistics & Probability Letters, 67(2):173–182, April 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400032X.

deLeon:2005:PLA

[dL05a] A. R. de Leon. Pairwise likelihood approach to grouped continuous model and its extension. Statistics & Probability Letters, 75(1):49–57, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002099.

Luna:2005:SPW

[dL05b] Sjöstedt de Luna. Some properties of weakly approaching sequences of distributions. Statistics & Probability Letters, 75(2):119–126, November 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002178.

Dahl:2006:NEV

[DL06] Christian M. Dahl and Michael Levine. Nonparametric estimation of volatility models with serially dependent innovations. Statistics & Probability Letters, 76(18):2007–2016, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001878.

Dembinska:2005:RDD

[DLB05] Anna Dembińska and Fernando López-Blázquez. k th records from discrete distributions. Statistics & Probability Letters, 71(3):203–214, March 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003025.

delaCal:2005:GOT

[dlCC05a] J. de la Cal and J. Cárcamo. A general Ostrowski-type inequality. Statistics & Probability Letters, 72(2):145–152, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000246.

delaCal:2005:IEE

[dlCC05b] J. de la Cal and J. Cárcamo. Inequalities for expected extreme order statistics. Statistics & Probability Letters, 73(3):219– 231, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001069.

delaCal:2001:MPC

[dlCV01] Jesús de la Cal and Ana M. Valle. Monotonicity properties of certain expected extreme order statistics. Statistics & Probability Letters, 52(3):313–319, April 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002327.

delaHorra:2005:BMS

[dlHRB05] Julián de la Horra and María Teresa Rodríguez-Bernal. Bayesian model selection: a predictive approach with losses based on distances L^{1} and L^{2}. Statistics & Probability Letters, 71(3):257–265, March 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204003074.

Dunker:2000:SBP

[DLL00] T. Dunker, W. V. Li, and W. Linde. Small ball probabilities for integrals of weighted Brownian motion. Statistics & Probability Letters, 46(3):211–216, February 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900098X.

delaPena:2008:IPR

[dlPGM08] Victor de la Pena, Henryk Gzyl, and Patrick McDonald. Inverse problems for random walks on trees: Network tomography. Statistics & Probability Letters, 78(18):3176–3183, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002885.

delaPena:2002:DDI

[dlPZ02] Victor H. de la Peña and Ingrid-Mona Zamfirescu. Decoupling and domination inequalities with application to Wald’s identity for martingales. Statistics & Probability Letters, 57(2):157–170, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000421.

Denuit:2000:THC

[DLS00] Michel Denuit, Claude Lefèvre, and Moshe Shaked. On the theory of high convexity stochastic orders. Statistics & Probability Letters, 47(3):287–293, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001662.

Drmota:2005:RWC

[DM05] Michael Drmota and Jean-François Marckert. Reinforced weak convergence of stochastic processes. Statistics & Probability Letters, 71(3):283–294, March 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204003116.

Dette:2002:FAD

[DMB02] Holger Dette, Viatcheslav B. Melas, and Stefanie Biedermann. A functional-algebraic determination of D-optimal designs for trigonometric regression models on a partial circle. Statistics & Probability Letters, 58(4):389–397, July 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001529.

Donati-Martin:2000:PNM

[DMMY00] Catherine Donati-Martin, Hiroyuki Matsumoto, and Marc Yor. On positive and negative moments of the integral of geometric Brownian motions. Statistics & Probability Letters, 49(1):45–52, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000298.

DeCesare:2001:EMS

[DMP01a] L. De Cesare, D. E. Myers, and D. Posa. Estimating and modeling space-time correlation structures. Statistics & Probability Letters, 51(1):9–14, January 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001310.

DeIaco:2001:STA

[DMP01b] S. De Iaco, D. E. Myers, and D. Posa. Space-time analysis using a general product-sum model. Statistics & Probability Letters, 52(1):21–28, March 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002005.

Dominguez-Molina:2007:IDS

[DMRA07] J. Armando Domínguez-Molina and Alfonso Rocha-Arteaga. On the infinite divisibility of some skewed symmetric distributions. Statistics & Probability Letters, 77(6):644–648, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002987.

Diedhiou:2009:CTB

[DN09] Alassane Diédhiou and Papa Ngom. Cutoff time based on generalized divergence measure. Statistics & Probability Letters, 79(10):1343–1350, May 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000704.

DeGregorio:2007:DSF

[DO07] A. De Gregorio and E. Orsingher. A Darling–Siegert formula relating some Bessel integrals and random walks. Statistics & Probability Letters, 77(7):667–680, April 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003014.

Doksum:2007:TOB

[DOKN07] Kjell Doksum, Akichika Ozeki, Jihoon Kim, and Elias Chaibub Neto. Thinking outside the box: Statistical inference based on Kullback–Leibler empirical projections. Statistics & Probability Letters, 77(12):1201–1213, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000843.

Downes:2009:BTD

[Dow09] A. N. Downes. Bounds for the transition density of time-homogeneous diffusion processes. Statistics & Probability Letters, 79(6):835–841, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005233.

Draper:2000:RAM

[DP00] Norman R. Draper and Friedrich Pukelsheim. Ridge analysis of mixture response surfaces. Statistics & Probability Letters, 48(2):131–140, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001959.

Draper:2003:CRS

[DP03] Norman R. Draper and Friedrich Pukelsheim. Canonical reduction of second-order fitted models subject to linear restrictions. Statistics & Probability Letters, 63(4):401–410, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001196.

Dewan:2000:EBL

[DR00] Isha Dewan and B. L. S. Prakasa Rao. Explicit bounds on Levy–Prohorov distance for a class of multidimensional distribution functions. Statistics & Probability Letters, 48(2): 105–119, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001790.

Dai:2001:ESA

[DR01a] Wen Dai and John Robinson. Empirical saddlepoint approximations of the Studentized mean under simple random sampling. Statistics & Probability Letters, 53(3):331–337, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520100092X.

Dohler:2001:ARN

[DR01b] Sebastian Döhler and Ludger Rüschendorf. An approximation result for nets in functional estimation. Statistics & Probability Letters, 52(4):373–380, May 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002248.

Dewan:2002:CLT

[DR02] Isha Dewan and B. L. S. Prakasa Rao. Central limit theorem for U-statistics of associated random variables. Statistics & Probability Letters, 57(1):9–15, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001948. See corrigendum [DR15].

Danielak:2003:SBE

[DR03] Katarzyna Danielak and Tomasz Rychlik. Sharp bounds for expectations of spacings from DDA and DFRA families. Statistics & Probability Letters, 65(4):303–316, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002554.

Danielak:2004:SBE

[DR04] Katarzyna Danielak and Mohammad Z. Raqab. Sharp bounds on expectations of k th record spacings from restricted families. Statistics & Probability Letters, 69(2):175–187, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001804.

Dewan:2005:WSR

[DR05] Isha Dewan and B. L. S. Prakasa Rao. Wilcoxon-signed rank test for associated sequences. Statistics & Probability Letters, 71(2):131–142, February 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002731.

Quintela-del-Rio:2006:NEM

[dR06] A. Quintela del Río. Nonparametric estimation of the maximum hazard under dependence conditions. Statistics & Probability Letters, 76(11):1117–1124, June 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004554.

Davydov:2008:NCI

[DR08] Youri Davydov and Vladimir Rotar. On a non-classical invariance principle. Statistics & Probability Letters, 78(14): 2031–2038, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000746.

Dewan:2015:CT

[DR15] Isha Dewan and B. L. S. Prakasa Rao. Corrigendum to “Central limit theorem for U-statistics of associated random variables” [Statist. Probab. Lett. 57 (1) (2002) 9–15]. Statistics & Probability Letters, (Letters):147–148, November 1– 300, 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002412. See [DR02].

Droge:2006:MRC

[Dro06] Bernd Droge. Minimax regret comparison of hard and soft thresholding for estimating a bounded normal mean. Statistics & Probability Letters, 76(1):83–92, January 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002713.

Drton:2008:MSL

[Drt08] Mathias Drton. Multiple solutions to the likelihood equations in the Behrens–Fisher problem. Statistics & Probability Letters, 78(18):3288–3293, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003167.

Datta:2000:EFS

[DS00a] Somnath Datta and Glen A. Satten. Estimating future stage entry and occupation probabilities in a multistage model based on randomly right-censored data. Statistics & Probability Letters, 50(1):89–95, October 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000869.

DeCock:2000:FMO

[DS00b] Dean DeCock and John Stufken. On finding mixed orthogonal arrays of strength 2 with many 2-level factors. Statistics & Probability Letters, 50(4):383–388, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001255.

DallAglio:2001:WLM

[DS01a] Marco Dall’Aglio and Marco Scarsini. When Lorenz met Lyapunov. Statistics & Probability Letters, 54(1):101–105, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000840.

Datta:2001:VAJ

[DS01b] Somnath Datta and Glen A. Satten. Validity of the Aalen–Johansen estimators of stage occupation probabilities and Nelson–Aalen estimators of integrated transition hazards for non-Markov models. Statistics & Probability Letters, 55(4):403–411, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001559.

DiConsiglio:2001:SRA

[DS01c] L. Di Consiglio and M. Scanu. Some results on asymptotics in adaptive cluster sampling. Statistics & Probability Letters, 52(2):189–197, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520000239X.

Devanarayan:2002:ESE

[DS02a] Viswanath Devanarayan and Leonard A. Stefanski. Empirical simulation extrapolation for measurement error models with replicate measurements. Statistics & Probability Letters, 59(3):219–225, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000986.

Ducinskas:2002:SOA

[DS02b] Kestutis Ducinskas and Jurate Saltyte. Second-order asymptotic expansion for the risk in classification of curved exponential populations. Statistics & Probability Letters, 59 (3):271–279, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001578.

Dembinska:2006:LTR

[DS06] A. Dembińska and A. Stepanov. Limit theorems for the ratio of weak records. Statistics & Probability Letters, 76 (14):1454–1464, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000794.

Dehling:2009:MZS

[DS09] Herold G. Dehling and Olimjon Sh. Sharipov. Marcinkiewicz– Zygmund strong laws for U-statistics of weakly dependent observations. Statistics & Probability Letters, 79(19):2028–2036, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209002363.

Dahmani:2009:CTR

[DSBA09] Abdelnasser Dahmani, Ahmed Ait Saidi, Fatah Bouhmila, and Mouloud Aissani. Consistency of the Tikhonov’s regularization in an ill-posed problem with random data. Statistics & Probability Letters, 79(6):722–727, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005051.

daSilva:2006:IMJ

[dSF06a] Telles T. da Silva and Marcelo D. Fragoso. Invariant measures for jump-type Fleming–Viot processes. Statistics & Probability Letters, 76(8):796–802, April 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003925.

daSilva:2006:NJT

[dSF06b] Telles T. da Silva and Marcelo D. Fragoso. A note on jump-type Fleming–Viot processes. Statistics & Probability Letters, 76(8):821–830, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003913.

daSilva:2008:SPJ

[dSF08] Telles Timóteo da Silva and Marcelo D. Fragoso. Sample paths of jump-type Fleming–Viot processes with bounded mutation operators. Statistics & Probability Letters, 78(13): 1784–1791, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000400.

deUna-Alvarez:2003:LSR

[dUÁ03] Jacobo de Uña-Álvarez. Large sample results under biased sampling when covariables are present. Statistics & Probability Letters, 63(3):287–293, July 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000932.

deUna-Alvarez:2008:SEB

[dUÁMM08] Jacobo de Uña-Álvarez and Luis F. Meira-Machado. A simple estimator of the bivariate distribution function for censored gap times. Statistics & Probability Letters, 78(15):2440–2445, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001478.

Dubnicka:2002:TAO

[Dub02] Suzanne R. Dubnicka. Testing against ordered alternatives for multivariate censored survival data. Statistics & Probability Letters, 56(2):207–216, January 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001894.

Dubischar:2003:LCP

[Dub03] Daniel Dubischar. Lipschitz continuous processes for given marginals. Statistics & Probability Letters, 62(1):45–48, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202004236.

Dubickas:2008:EPD

[Dub08] Arturas Dubickas. An estimate for the probability of dependent events. Statistics & Probability Letters, 78(17): 2839–2843, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208002101.

Duchesne:2004:MMD

[Duc04] Pierre Duchesne. On matricial measures of dependence in vector ARCH models with applications to diagnostic checking. Statistics & Probability Letters, 68(2):149–160, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000768. See corrigen-

dum [Duc10].

Ducinskas:2009:AEE

[Duc09] Kestutis Ducinskas. Approximation of the expected error rate in classification of the Gaussian random field observations. Statistics & Probability Letters, 79(2):138–144, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800357X.

Duchesne:2010:CMM

[Duc10] Pierre Duchesne. Corrigendum to: “On matricial measures of dependence in vector ARCH models with applications to diagnostic checking” [Statist. Probab. Lett. 68 (2004) 149– 160]. Statistics & Probability Letters, 80(9–10):910, May 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000477. See [Duc04].

Dudzinski:2003:NAS

[Dud03] Marcin Dudziński. A note on the almost sure Central Limit Theorem for some dependent random variables. Statistics & Probability Letters, 61(1):31–40, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002912.

Dudzinski:2008:ASC

[Dud08] Marcin Dudziński. The almost sure Central Limit Theorems in the joint version for the maxima and sums of certain stationary Gaussian sequences. Statistics & Probability Letters, 78(4):347–357, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207002477.

Dukhovny:2007:LPR

[Duk07] Alexander Dukhovny. Lattice polynomials of random variables. Statistics & Probability Letters, 77(10):989–994, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000387.

Duncan:2006:PSP

[Dun06] T. E. Duncan. Prediction for some processes related to a fractional Brownian motion. Statistics & Probability Letters, 76(2):128–134, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002506. See [DF11].

Dupuy:2006:ECP

[Dup06] Jean-François Dupuy. Estimation in a change-point hazard regression model. Statistics & Probability Letters, 76 (2):182–190, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002877.

Durot:2003:KTT

[Dur03] Cécile Durot. A Kolmogorov-type test for monotonicity of regression. Statistics & Probability Letters, 63(4):425–433, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001226.

Denuit:2001:SLT

[DV01] Michel Denuit and Sébastien Van Bellegem. On the stop-loss and total variation distances between random sums. Statistics & Probability Letters, 53(2):153–165, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000670.

Debbarh:2008:TAN

[DV08] Mohammed Debbarh and Vivian Viallon. Testing additivity in nonparametric regression under random censorship. Statistics & Probability Letters, 78(16):2584–2591, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003520.

Drost:2008:NIV

[DvdAW08] Feike C. Drost, Ramon van den Akker, and Bas J. M. Werker. Note on integer-valued bilinear time series models. Statistics & Probability Letters, 78(8):992–996, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003586.

Dzhaparidze:2005:OES

[DvZ05] Kacha Dzhaparidze and Harry van Zanten. Optimality of an explicit series expansion of the fractional Brownian sheet. Statistics & Probability Letters, 71(4):295–301, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204003128.

Du:2000:JWD

[DW00] Zhiyi Du and Douglas P. Wiens. Jackknifing, weighting, diagnostics and variance estimation in generalized M-estimation. Statistics & Probability Letters, 46(3):287–299, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001170.

DeVinney:2002:SOD

[DW02] Jason DeVinney and John C. Wierman. A SLLN for a one-dimensional class cover problem. Statistics & Probability Letters, 59(4):425–435, October 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002432.

Dippon:2005:SAP

[DW05] Jürgen Dippon and Harro Walk. Simplified analytical proof of Blackwell’s renewal theorem. Statistics & Probability Letters, 74(1):15–20, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001513.

DuBien:2004:MRC

[DWC04] Janice L. DuBien, William D. Warde, and Seong S. Chae. Moments of Rand’s c statistic in cluster analysis. Statistics & Probability Letters, 69(3):243–252, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001646.

Dette:2005:EOD

[DWZ05] Holger Dette, Weng Kee Wong, and Wei Zhu. On the equivalence of optimality design criteria for the placebo-treatment problem. Statistics & Probability Letters, 74(4):337–346, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001938.

DeLaPena:2004:BFP

[DY04] Victor H. De La Peña and Ming Yang. Bounding the first passage time on an average. Statistics & Probability Letters, 67(1):1–7, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003341.

Dong:2001:GEF

[DZ01] Jianping Dong and Chuang Zheng. Generalized edge frequency polygon for density estimation. Statistics & Probability Letters, 55(2):137–145, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000773.

Davydov:2002:CGL

[DZ02] Youri Davydov and Ricardas Zitikis. Convergence of generalized Lorenz curves based on stationary ergodic random sequences with deterministic noise. Statistics & Probability Letters, 59(4):329–340, October 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200202X.

Ding:2004:SSO

[DZ04] Ying Ding and Xinsheng Zhang. Some stochastic orders of Kotz-type distributions. Statistics & Probability Letters, 69(4):389–396, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001567.

Debicki:2004:SGP

[DZB04] Krzystof Debicki, Bert Zwart, and Sem Borst. The supremum of a Gaussian process over a random interval. Statistics & Probability Letters, 68(3):221–234, July 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003286.

Eryilmaz:2003:NSR

[EB03] S. Eryilmaz and I. G. Bairamov. On a new sample rank of an order statistics and its concomitant. Statistics & Probability Letters, 63(2):123–131, June 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000592.

Evans:2002:NPB

[ED02] Barry A. Evans and David A. Dickey. Normalizations for periodogram-based unit root tests. Statistics & Probability Letters, 60(4):343–350, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001542.

Eryilmaz:2008:RCV

[EF08] Serkan Eryilmaz and James C. Fu. Runs in continuous-valued sequences. Statistics & Probability Letters, 78(6):759–765, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700329X.

Epifani:2002:CMS

[EFL02] I. Epifani, S. Fortini, and L. Ladelli. A characterization for mixtures of semi-Markov processes. Statistics & Probability Letters, 60(4):445–457, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003243.

Efromovich:2000:SLB

[Efr00] Sam Efromovich. Sharp linear and block shrinkage wavelet estimation. Statistics & Probability Letters, 49(4):323–329, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520000064X.

Efromovich:2003:LEB

[Efr03] Sam Efromovich. On the limit in the equivalence between heteroscedastic regression and filtering model. Statistics & Probability Letters, 63(3):239–242, July 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000877.

Efromovich:2008:NRE

[Efr08] Sam Efromovich. Nonparametric regression estimation with assigned risk. Statistics & Probability Letters, 78(13):1748– 1756, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000369.

Eisenberg:2001:GMI

[EG01] Bennett Eisenberg and B. K. Ghosh. A generalization of Markov’s inequality. Statistics & Probability Letters, 53(1):59– 65, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000220.

Esponda:2009:SNQ

[EG09] Fernando Esponda and Victor M. Guerrero. Surveys with negative questions for sensitive items. Statistics & Probability Letters, 79(24):2456–2461, December 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003307.

Epifani:2006:SEL

[EGM06] I. Epifani, A. Guglielmi, and E. Melilli. A stochastic equation for the law of the random Dirichlet variance. Statistics & Probability Letters, 76(5):495–502, March 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003299.

Esteban:2008:DBT

[EHMM08] M. D. Esteban, T. Hobza, D. Morales, and Y. Marhuenda. Divergence-based tests for model diagnostic. Statistics & Probability Letters, 78(13):1702–1710, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000291.

Eisenberg:2008:EMI

[Eis08] Bennett Eisenberg. On the expectation of the maximum of IID geometric random variables. Statistics & Probability Letters, 78(2):135–143, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002040.

Eisenberg:2009:NPT

[Eis09] Bennett Eisenberg. The number of players tied for the record. Statistics & Probability Letters, 79(3):283–288, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003982.

Evangelaras:2003:ECB

[EK03a] H. Evangelaras and C. Koukouvinos. Effects confounded with blocks in factorial designs: a projective geometric approach with two blocks. Statistics & Probability Letters, 64(1):105–111, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001457.

Evangelaras:2003:MDE

[EK03b] H. Evangelaras and C. Koukouvinos. On maximizing a design estimation efficiency using hidden projection properties. Statistics & Probability Letters, 62(4):419–427, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000518.

Evangelaras:2004:GPT

[EK04] H. Evangelaras and C. Koukouvinos. On generalized projectivity of two-level screening designs. Statistics & Probability Letters, 68(4):429–434, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001336.

Evangelaras:2006:NIO

[EKK06] H. Evangelaras, E. Kolaiti, and C. Koukouvinos. Nonisomorphic orthogonal arrays obtained by juxtaposition. Statistics & Probability Letters, 76(3):274–279, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002968.

Evangelaras:2002:UAA

[EKM02] H. Evangelaras, C. Koukouvinos, and P. Mantas. A unified approach in addition or deletion of two level factorial designs. Statistics & Probability Letters, 59(2):125–133, September 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001116.

Evangelaras:2005:ESN

[EKS05] H. Evangelaras, C. Koukouvinos, and S. Stylianou. Evaluation of some non-orthogonal saturated designs with two levels. Statistics & Probability Letters, 74(4):322–329, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001914.

ElOtmani:2008:BDS

[El 08] Mohamed El Otmani. BSDE driven by a simple Lévy process with continuous coefficient. Statistics & Probability Letters, 78(11):1259–1265, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004038.

Engel:2003:BLE

[EL03] Hans-Andreas Engel and Christoph Leuenberger. Benford’s law for exponential random variables. Statistics & Probability Letters, 63(4):361–365, July 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001019.

Endo:2008:SMG

[EM08] Kotaro Endo and Muneya Matsui. The stationarity of multidimensional generalized Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 78(14):2265–2272, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800120X.

Eaton:2009:NPP

[EM09] Morris L. Eaton and Robb J. Muirhead. The “North Pole Problem” and random orthogonal matrices. Statistics & Probability Letters, 79(17):1878–1883, September 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001953.

El-Nouty:2003:FMF

[EN03] Charles El-Nouty. The fractional mixed fractional Brownian motion. Statistics & Probability Letters, 65(2):111– 120, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002281.

Englander:2004:ECC

[Eng04a] János Engländer. An example and a conjecture concerning scaling limits of superdiffusions. Statistics & Probability Letters, 66(3):363–368, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300364X.

Englander:2004:LDG

[Eng04b] János Engländer. Large deviations for the growth rate of the support of supercritical super-Brownian motion. Statistics & Probability Letters, 66(4):449–456, March 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003869.

Enikeeva:2006:AME

[Eni06] Farida Enikeeva. Adaptive minimax estimation of a fractional derivative. Statistics & Probability Letters, 76(14):1441–1448, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000770.

Estevez-Perez:2002:CRQ

[EP02] Graciela Estévez-Pérez. On convergence rates for quadratic errors in kernel hazard estimation. Statistics & Probability Letters, 57(3):231–241, April 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000524.

Ermakov:2000:DTN

[Erm00] M. S. Ermakov. On distinguishability of two nonparametric sets of hypothesis. Statistics & Probability Letters, 48(3):275– 282, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000079.

Eryilmaz:2007:ERC

[Ery07] Serkan Eryilmaz. Extension of runs to the continuousvalued sequences. Statistics & Probability Letters, 77(4): 383–388, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002549.

Eryilmaz:2008:DRS

[Ery08] Serkan Eryilmaz. Distribution of runs in a sequence of exchangeable multi-state trials. Statistics & Probability Letters, 78(12):1505–1513, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000047.

El-Shehawey:2009:GRP

[ES09] M. A. El-Shehawey. On the gambler’s ruin problem for a finite Markov chain. Statistics & Probability Letters, 79 (14):1590–1595, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001254.

Esaulova:2004:LDE

[Esa04] Veronika Esaulova. Large deviations of L_{1}-error of empirical measures on partitions. Statistics & Probability Letters, 69(4):439–445, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520400197X.

Eshima:2004:CEM

[Esh04] Nobuoki Eshima. Canonical exponential models for analysis of association between two sets of variables. Statistics & Probability Letters, 66(2):135–144, January 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002980.

Es-sebaiy:2008:HRC

[EsO08] Khalifa Es-sebaiy and Youssef Ouknine. How rich is the class of processes which are infinitely divisible with respect to time? Statistics & Probability Letters, 78(5):537–547, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002854.

Es-Sebaiy:2009:EDF

[ESOO09] Khalifa Es-Sebaiy, Idir Ouassou, and Youssef Ouknine. Estimation of the drift of fractional Brownian motion. Statistics & Probability Letters, 79(14):1647–1653, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001515.

Eberly:2005:LRP

[ET05] Lynn E. Eberly and Lisa M. Thackeray. On Lange and Ryan’s plotting technique for diagnosing non-normality of random effects. Statistics & Probability Letters, 75(2):77–85, November 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002129.

Eshima:2007:ECC

[ET07] Nobuoki Eshima and Minoru Tabata. Entropy correlation coefficient for measuring predictive power of generalized linear models. Statistics & Probability Letters, 77(6):588–593, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002732.

Elalaoui-Talibi:2000:NML

[ETC00] Hussain Elalaoui-Talibi and Kameswarrao S. Casukhela. A note on the multivariate local time intensity of exchangeable interval partitions. Statistics & Probability Letters, 48(3):269– 273, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000067.

Etemadi:2007:SWA

[Ete07] N. Etemadi. Stability of weighted averages of 2-exchangeable random variables. Statistics & Probability Letters, 77(4): 389–395, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002550.

Elalaoui-Talibi:2008:CDR

[ETP08] Hussain Elalaoui-Talibi and Lisa D. Peterson. Convergence in distribution of random compact sets in Polish spaces. Statistics & Probability Letters, 78(6):736–738, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003252.

Ewald:2008:NMD

[Ewa08] Christian-Oliver Ewald. A note on the Malliavin derivative operator under change of variable. Statistics & Probability Letters, 78(2):173–178, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002106.

Esteves:2004:SCU

[EWI04] L. G. Esteves, S. Wechsler, and P. L. Iglesias. Some characterizations of uniform models. Statistics & Probability Letters, 69(4):397–404, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001920.

Eickhoff:2004:SSC

[EZA04] Jens C. Eickhoff, Jun Zhu, and Yasuo Amemiya. On the simulation size and the convergence of the Monte Carlo EM algorithm via likelihood-based distances. Statistics & Probability Letters, 67(2):161–171, April 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000252.

Fahrner:2000:EAS

[Fah00] Ingo Fahrner. An extension of the almost sure max-limit theorem. Statistics & Probability Letters, 49(1):93–103, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000365.

Fainleib:2001:MSF

[Fai01] A. S. Fainleib. Multifold sumsets and fast decreasing of concentration functions. Statistics & Probability Letters, 51(4):415–421, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520000184X.

Falk:2005:GME

[Fal05] Michael Falk. On the generation of a multivariate extreme value distribution with prescribed tail dependence parameter matrix. Statistics & Probability Letters, 75(4):307–314, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002415.

Fang:2002:ELP

[Fan02a] B. Q. Fang. Estimation of the location parameter of the l_{1}-norm symmetric matrix variate distributions. Statistics & Probability Letters, 57(3):269–280, April 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000603.

Fang:2002:ODP

[Fan02b] Zhide Fang. D-optimal designs for polynomial regression models through origin. Statistics & Probability Letters, 57(4):343– 351, May 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000858.

Fang:2003:ODW

[Fan03] Zhide Fang. D-optimal designs for weighted polynomial regression. Statistics & Probability Letters, 63(2):205–213, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000841.

Fazekas:2003:LTE

[Faz03] István Fazekas. Limit theorems for the empirical distribution function in the spatial case. Statistics & Probability Letters, 62(3):251–262, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000105.

Fukasawa:2002:ECG

[FB02] T. Fukasawa and I. V. Basawa. Estimation for a class of generalized state-space time series models. Statistics & Probability Letters, 60(4):459–473, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003255.

Fisher:2000:ENR

[FBVW00] Evan Fisher, Matthew Berman, Natalie Vowels, and Christine Wilson. Excursions of a normal random walk above a boundary. Statistics & Probability Letters, 48(2):141–151, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001960.

Frey:2003:SRC

[FC03] Jesse Frey and Noel Cressie. Some results on constrained Bayes estimators. Statistics & Probability Letters, 65(4): 389–399, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002888.

Ferrari:2008:SAL

[FC08] Silvia L. P. Ferrari and Audrey H. M. A. Cysneiros. Skovgaard’s adjustment to likelihood ratio tests in exponential family nonlinear models. Statistics & Probability Letters, 78(17):3047–3055, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002666.

Ferrari:2002:CMP

[FCN02] Silvia L. P. Ferrari and Francisco Cribari-Neto. Corrected modified profile likelihood heteroskedasticity tests. Statistics & Probability Letters, 57(4):353–361, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000871.

Fried:2005:LVA

[FD05] Roland Fried and Vanessa Didelez. Latent variable analysis and partial correlation graphs for multivariate time series. Statistics & Probability Letters, 73(3):287–296, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001227.

Frostig:2009:DFT

[FD09] Esther Frostig and Michel Denuit. Dependence in failure times due to environmental factors. Statistics & Probability Letters, 79(4):487–495, February 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004550.

Finkelstein:2008:AFR

[FE08] Maxim Finkelstein and Veronica Esaulova. On asymptotic failure rates in bivariate frailty competing risks models. Statistics & Probability Letters, 78(10):1174–1180, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003896.

Fernandez:2000:BIT

[Fer00] Arturo J. Fernández. Bayesian inference from type II doubly censored Rayleigh data. Statistics & Probability Letters, 48(4):393–399, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000213.

Ferreira:2003:EAV

[Fer03] H. Ferreira. Extremes of associated variables. Statistics & Probability Letters, 63(4):333–338, July 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000968.

Fernandes:2004:BPD

[Fer04] Marcelo Fernandes. Bounds for the probability distribution function of the linear ACD process. Statistics & Probability Letters, 68(2):169–176, June 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000781.

Ferger:2009:ASM

[Fer09] Dietmar Ferger. Argmax-stable marked empirical processes. Statistics & Probability Letters, 79(9):1203–1206, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000248.

Freitas:2008:LBD

[FF08] Ana Cristina Moreira Freitas and Jorge Milhazes Freitas. On the link between dependence and independence in extreme value theory for dynamical systems. Statistics & Probability Letters, 78(9):1088–1093, July 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003781.

Ferrante:2009:PFA

[FF09] Marco Ferrante and Nadia Frigo. Particle filtering approximations for a Gaussian-generalized inverse Gaussian model. Statistics & Probability Letters, 79(4):442–449, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208004410.

Francq:2004:ETV

[FG04] Christian Francq and Antony Gautier. Estimation of timevarying ARMA models with Markovian changes in regime. Statistics & Probability Letters, 70(4):243–251, December 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204002639.

Figueiredo:2006:GFM

[FG06] Adelaide Figueiredo and Paolo Gomes. Goodness-of-fit methods for the bipolar Watson distribution defined on the hypersphere. Statistics & Probability Letters, 76(2):142–152, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002749.

Feigin:2008:FPR

[FGMS08] Paul D. Feigin, Phillip Gould, Gael M. Martin, and Ralph D. Snyder. Feasible parameter regions for alternative discrete state space models. Statistics & Probability Letters, 78(17):2963–2970, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002502.

Fuentes-Garcia:2009:NDP

[FGMW09] Ruth Fuentes-García, Ramsés H. Mena, and Stephen G. Walker. A nonparametric dependent process for Bayesian regression. Statistics & Probability Letters, 79(8):1112–1119, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209000030.

Fu:2006:APE

[FH06] Wenjiang J. Fu and Peter Hall. Asymptotic properties of estimators in age-period-cohort analysis. Statistics & Probability Letters, 76(17):1925–1929, November 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001593.

Farah:2007:CDD

[FH07] Mohamed Ben Farah and Abdelhamid Hassairi. Characterization of the Dirichlet distribution on symmetric matrices. Statistics & Probability Letters, 77(4):357–364, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002446.

Fan:2008:LTS

[FH08] Sheng-Jun Fan and Jian-Hua Hu. A limit theorem for solutions to BSDEs in the space of processes. Statistics & Probability Letters, 78(8):1024–1033, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003628.

Futschik:2006:CKD

[FI06] A. Futschik and E. Isogai. On the consistency of kernel density estimates under modality constraints. Statistics & Probability Letters, 76(4):431–437, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003159.

Fenoy:2007:OST

[FI07] M. Mar Fenoy and Pilar Ibarrola. The optional sampling theorem for submartingales in the sequentially planned context. Statistics & Probability Letters, 77(8):826–831, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000168.

Fierro:2003:TGF

[Fie03] Raúl Fierro. A test of goodness of fit testing for stochastic intensities associated to counting processes. Statistics & Probability Letters, 64(3):287–292, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001731.

Figueiredo:2007:CTU

[Fig07]  Adelaide Figueiredo.  Comparison of tests of unifor-

mity defined on the hypersphere. Statistics & Probability Letters, 77(3):329–334, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002392.

Finkelstein:2003:MSS

[Fin03a] M. S. Finkelstein. A model for spatial survival. Statistics & Probability Letters, 62(4):371–378, May 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000415.

Finkelstein:2003:OCB

[Fin03b] M. S. Finkelstein. On one class of bivariate distributions. Statistics & Probability Letters, 65(1):1–6, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002098.

Finkelstein:2006:ROM

[Fin06] Maxim Finkelstein. On relative ordering of mean residual lifetime functions. Statistics & Probability Letters, 76(9):939–944, May 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205004116.

Fitzgibbon:2006:SSA

[Fit06] L. J. Fitzgibbon. On sampling stationary autoregressive model parameters uniformly in r^{2} value. Statistics & Probability Letters, 76(4):349–352, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003056.

Fotopoulos:2001:MLE

[FJ01] Stergios Fotopoulos and Venkata Jandhyala. Maximum likelihood estimation of a change-point for exponentially distributed random variables. Statistics & Probability Letters, 51(4):423–429, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001851.

Fotopoulos:2004:BIA

[FJ04] Stergios B. Fotopoulos and Venkata K. Jandhyala. Bessel inequalities with applications to conditional log returns under GIG scale mixtures of normal vectors. Statistics & Probability Letters, 66(2):117–125, January 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002852.

Fotopoulos:2007:HEI

[FJ07] S. B. Fotopoulos and V. K. Jandhyala. On Hinkley’s estimator: Inference about the change point. Statistics & Probability Letters, 77(13):1449–1458, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000752.

Fakoor:2009:AEI

[FJA09] Vahid Fakoor, Sarah Jomhoori, and Hasanali Azarnoosh. Asymptotic expansion for ISE of kernel density estimators under censored dependent model. Statistics & Probability Letters, 79(17):1809–1817, September 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001849.

Frahm:2003:ECA

[FJS03] Gabriel Frahm, Markus Junker, and Alexander Szimayer. Elliptical copulas: applicability and limitations. Statistics & Probability Letters, 63(3):275–286, July 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000920.

Farchione:2008:CIN

[FK08] David Farchione and Paul Kabaila. Confidence intervals for the normal mean utilizing prior information. Statistics & Probability Letters, 78(9):1094–1100, July 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003793.

Fallis:2002:FRI

[FL02] Don Fallis and Gerrard Liddell. Further results on inquiry and truth possession. Statistics & Probability Letters, 60(2):169–182, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003085.

Figueroa-Lopez:2008:STM

[FL08]  José E. Figueroa-López.  Small-time moment asymp-

totics for Lévy processes. Statistics & Probability Letters, 78(18):3355–3365, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208003271.

Fang:2003:NOF

[FLQ03] Kai-Tai Fang, Dennis K. J. Lin, and Hong Qin. A note on optimal foldover design. Statistics & Probability Letters, 62(3):245–250, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000087.

Fort:2000:SEH

[FM00] Gersende Fort and Eric Moulines. V -subgeometric ergodicity for a Hastings–Metropolis algorithm. Statistics & Probability Letters, 49(4):401–410, October 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000742.

Fernandez:2003:SOR

[FM03a] Ignacio Cascos Fernández and Ilya Molchanov. A stochastic order for random vectors and random sets based on the Aumann expectation. Statistics & Probability Letters, 63(3):295– 305, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000944.

Filipiak:2003:ONB

[FM03b] K. Filipiak and A. Markiewicz. Optimality of neighbor balanced designs under mixed effects model. Statistics & Probability Letters, 61(3):225–234, February 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003395.

Filipiak:2004:OTO

[FM04] K. Filipiak and A. Markiewicz. Optimality of type I orthogonal arrays for general interference model with correlated observations. Statistics & Probability Letters, 68(3):259–265, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001002.

Freeland:2005:APC

[FM05] R. Keith Freeland and Brendan McCabe. Asymptotic properties of CLS estimators in the Poisson AR(1) model. Statistics & Probability Letters, 73(2):147–153, June 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001045.

Freeman:2006:IBC

[FM06] Jade Freeman and Reza Modarres. Inverse Box–Cox: The power-normal distribution. Statistics & Probability Letters, 76(8):764–772, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520500386X.

Frolov:2000:SLM

[FMS00] Andrei Frolov, Alexander Martikainen, and Josef Steinebach. Strong laws for the maximal gain over increasing runs. Statistics & Probability Letters, 50(3):305–312, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000119X.

Fay:2004:EEL

[FMS04] Gilles Faÿ, Eric Moulines, and Philippe Soulier. Edgeworth expansions for linear statistics of possibly longrange-dependent linear processes. Statistics & Probability Letters, 66(3):275–288, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003407.

Furrer:2007:PWM

[FN07] Reinhard Furrer and Philippe Naveau. Probability weighted moments properties for small samples. Statistics & Probability Letters, 77(2):190–195, January 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002215.

Flecher:2009:ECS

[FNA09] C. Flecher, P. Naveau, and D. Allard. Estimating the closed skew-normal distribution parameters using weighted moments. Statistics & Probability Letters, 79(19):1977–1984, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209002168.

Forchini:2000:DSS

[For00] G. Forchini. The density of the sufficient statistics for a Gaussian AR(1) model in terms of generalized functions. Statistics & Probability Letters, 50(3):237–243, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001115.

Fotopoulos:2004:TDT

[Fot04] Stergios B. Fotopoulos. Tempered distributions and their application in computing conditional moments for normal mixtures. Statistics & Probability Letters, 67(3):257–266, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000379.

Fotopoulos:2005:TSD

[Fot05] Stergios B. Fotopoulos. Type G and spherical distributions on R^{d}. Statistics & Probability Letters, 72(1):23–32, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000106.

Fotopoulos:2009:GCR

[Fot09] Stergios B. Fotopoulos. The geometric convergence rate of the classical change-point estimate. Statistics & Probability Letters, 79(2):131–137, January 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003568.

Ferreira:2007:CLC

[FP07] Fátima Ferreira and António Pacheco. Comparison of levelcrossing times for Markov and semi-Markov processes. Statistics & Probability Letters, 77(2):151–157, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600215X.

Ferreira:2008:HCE

[FP08] H. Ferreira and L. Pereira. How to compute the extremal index of stationary random fields. Statistics & Probability Letters, 78(11):1301–1304, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004099.

Fang:2003:NCN

[FQ03] Kai-Tai Fang and Hong Qin. A note on construction of nearly uniform designs with large number of runs. Statistics & Probability Letters, 61(2):215–224, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003577.

Falk:2001:ECD

[FR01] Michael Falk and Rolf-Dieter Reiss. Estimation of canonical dependence parameters in a class of bivariate peaksover-threshold models. Statistics & Probability Letters, 52 (3):233–242, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001942.

Falk:2002:CRC

[FR02] Michael Falk and Rolf Dieter Reiss. A characterization of the rate of convergence in bivariate extreme value models. Statistics & Probability Letters, 59(4):341–351, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202002092.

Fung:2006:CCI

[FR06] Thomas Fung and John Robinson. Continuity corrections for integer-valued saddlepoint approximations. Statistics & Probability Letters, 76(14):1465–1469, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000800.

Frahm:2006:EDC

[Fra06] Gabriel Frahm. On the extremal dependence coefficient of multivariate distributions. Statistics & Probability Letters, 76(14):1470–1481, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000812.

Fekri:2006:RES

[FRG06] M. Fekri and A. Ruiz-Gazen. Robust estimation in the simple errors-in-variables model. Statistics & Probability Letters, 76(16):1741–1747, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001301.

Frolov:2003:AMG

[Fro03] Andrei N. Frolov. On asymptotics of the maximal gain without losses. Statistics & Probability Letters, 63(1):13–23, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202004200.

Frolov:2005:CCR

[Fro05] Andrei N. Frolov. Converses to the Csörgő-Révész laws. Statistics & Probability Letters, 72(2):113–123, April 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000210.

Frostig:2006:RDM

[Fro06] Esther Frostig. On risk dependence and mrl ordering. Statistics & Probability Letters, 76(3):231–243, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002920.

Filipiak:2008:OCD

[FRSWT08] Katarzyna Filipiak, Rafal Rózański, Aneta Sawikowska, and Dominika Wojtera-Tyrakowska. On the E-optimality of complete designs under an interference model. Statistics & Probability Letters, 78(15):2470–2477, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800151X.

Fan:2000:MSC

[FS00] Ai Hua Fan and Narn-Rueih Shieh. Multifractal spectra of certain random Gibbs measures. Statistics & Probability Letters, 47(1):25–31, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001339.

Fang:2001:CNM

[FS01] Hong-Bin Fang and Jianguo Sun. Consistency of nonparametric maximum likelihood estimation of a distribution function based on doubly interval-censored failure time data. Statistics & Probability Letters, 55(3):311–318, December 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001602.

Ferreira:2002:ALH

[FS02] H. Ferreira and M. Scotto. On the asymptotic location of high values of a stationary sequence. Statistics & Probability Letters, 60(4):475–482, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003267.

Farrell:2006:NLC

[FS06] Patrick J. Farrell and Brajendra C. Sutradhar. A non-linear conditional probability model for generating correlated binary data. Statistics & Probability Letters, 76(4):353–361, February 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003068.

Farahmand:2008:APR

[FS08a] K. Farahmand and C. T. Stretch. Algebraic polynomials with random non-symmetric coefficients. Statistics & Probability Letters, 78(11):1305–1313, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004105.

Freels:2008:SGR

[FS08b] S. Freels and K. Sinha. R-squared for general regression models in the presence of sampling weights. Statistics & Probability Letters, 78(12):1671–1672, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000242.

Fung:2008:CSM

[FS08c] Thomas Fung and Eugene Seneta. A characterisation of scale mixtures of the uniform distribution. Statistics & Probability Letters, 78(17):2883–2888, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002307.

Franke:2009:ERS

[FS09] Brice Franke and Tatsuhiko Saigo. The extremes of a random scenery as seen by a random walk in a random environment. Statistics & Probability Letters, 79(8):1025–1030, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005713.

Fu:2003:UMC

[FSC03] James C. Fu, Galit Shmueli, and Y. M. Chang. A unified Markov chain approach for computing the run length distribution in control charts with simple or compound rules. Statistics & Probability Letters, 65(4):457–466, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003110.

Fu:2002:ARL

[FSX02] James C. Fu, Fred A. Spiring, and Hansheng Xie. On the average run lengths of quality control schemes using a Markov chain approach. Statistics & Probability Letters, 56(4):369–380, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001833.

Fierro:2001:CES

[FT01a] Raúl Fierro and Soledad Torres. Correction on “The Euler scheme for Hilbert space valued stochastic differential equations” [Statist. Probab. Lett. 51 (3) (2001) 207–213]. Statistics & Probability Letters, 54(1):113, August 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000177. See [FT01b].

Fierro:2001:ESH

[FT01b] Raúl Fierro and Soledad Torres. The Euler scheme for Hilbert space valued stochastic differential equations. Statistics & Probability Letters, 51(3):207–213, February 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001188. See [FT01a].

Ferguson:2004:OSS

[FT04] Thomas S. Ferguson and Curtis Tatsuoka. An optimal strategy for sequential classification on partially ordered sets. Statistics & Probability Letters, 68(2):161–168, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400077X.

Ford:2008:LLS

[FT08] Kevin Ford and Gérald Tenenbaum. Localized large sums of random variables. Statistics & Probability Letters, 78(1):84– 89, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207001988.

Fukuyama:2008:CLT

[FU08] Katusi Fukuyama and Yôhei Ueno. On the Central Limit Theorem and the law of the iterated logarithm. Statistics & Probability Letters, 78(12):1384–1387, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004191.

Fujisawa:2000:VST

[Fuj00] Hironori Fujisawa. Variance stabilizing transformation and Studentization for estimator of correlation coefficient. Statistics & Probability Letters, 47(3):213–217, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001583.

Fujii:2007:NAD

[Fuj07] Takayuki Fujii. A note on the asymptotic distribution of the maximum likelihood estimator in a non-regular case. Statistics & Probability Letters, 77(16):1622–1627, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001435.

Fujii:2008:WCL

[Fuj08] Takayuki Fujii. On weak convergence of the likelihood ratio process in multi-phase regression models. Statistics & Probability Letters, 78(14):2066–2074, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000795.

Fukuda:2005:URD

[Fuk05] Kosei Fukuda. Unit-root detection allowing for measurement error. Statistics & Probability Letters, 74(4):373–377, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001975.

Furrer:2002:EDR

[Fur02] Reinhard Furrer. M-estimation for dependent random variables. Statistics & Probability Letters, 57(4):337–341, May 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202000846.

Furman:2007:CNB

[Fur07] Edward Furman. On the convolution of the negative binomial random variables. Statistics & Probability Letters, 77 (2):169–172, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002185.

Furman:2008:MGD

[Fur08] Edward Furman. On a multivariate gamma distribution. Statistics & Probability Letters, 78(15):2353–2360, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001338. See erra-

tum [SF12].

Forkman:2008:DMA

[FV08] Johannes Forkman and Steve Verrill. The distribution of McKay’s approximation for the coefficient of variation. Statistics & Probability Letters, 78(1):10–14, January 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001630.

Franco:2009:CGM

[FV09] Manuel Franco and Juana-María Vivo. Constraints for generalized mixtures of Weibull distributions with a common shape parameter. Statistics & Probability Letters, 79(15):1724–1730, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001746.

Froda:2001:UBC

[FvE01] Sorana Froda and Constance van Eeden. A uniform bound on the characteristic function of the exponentially tilted nulldistribution of a simple linear rank statistic and its rate of convergence. Statistics & Probability Letters, 54(1):107–112, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000931.

Favaro:2008:GCD

[FW08a] S. Favaro and S. G. Walker. A generalized constructive definition for the Dirichlet process. Statistics & Probability Letters, 78(16):2836–2838, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002095.

Ferland:2008:FAU

[FW08b] René Ferland and François Watier. FBSDE approach to utility portfolio selection in a market with random parameters. Statistics & Probability Letters, 78(4):426–434, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002581.

Fang:2001:SSM

[FWW01] Kai-Tai Fang, Song-Gui Wang, and Gang Wei. A stratified sampling model in spherical feature inspection using coordinate measuring machines. Statistics & Probability Letters, 51(1):25–34, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001334.

Fang:2000:UDE

[FY00] Kai-Tai Fang and Zhen-Hai Yang. On uniform design of experiments with restricted mixtures and generation of uniform distribution on some domains. Statistics & Probability Letters, 46(2):113–120, January 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000954.

Gao:2000:CLT

[GA00] Jiti Gao and Vo Anh. A Central Limit Theorem for a random quadratic form of strictly stationary processes. Statistics & Probability Letters, 49(1):69–79, August 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000341.

Gagnon:2002:SOP

[Gag02] Gregory Gagnon. Stability of option prices under uniform ellipticity. Statistics & Probability Letters, 59(4):361–365, October 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002328.

Gaiffas:2007:SEN

[Gai07] Stéphane Gaif̈fas. Sharp estimation in sup norm with random design. Statistics & Probability Letters, 77(8):782–794, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700003X.

Gallagher:2001:MFS

[Gal01] Colin M. Gallagher. A method for fitting stable autoregressive models using the autocovariation function. Statistics & Probability Letters, 53(4):381–390, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000414.

Gallego:2009:ELF

[Gal09] Jose L. Gallego. The exact likelihood function of a vector autoregressive moving average process. Statistics & Probability Letters, 79(6):711–714, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005026.

Gantert:2000:NLT

[Gan00] Nina Gantert. A note on logarithmic tail asymptotics and mixing. Statistics & Probability Letters, 49(2):113–118, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000377.

Gao:2001:MDM

[Gao01] Fuqing Gao. Moderate deviations for the maximum likelihood estimator. Statistics & Probability Letters, 55(4):345–352, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000232.

Gao:2003:DDB

[Gao03] Yonghong Gao. Data depth based on spatial rank. Statistics & Probability Letters, 65(3):217–225, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002542.

Gapeev:2004:AMS

[Gap04] Pavel V. Gapeev. On arbitrage and Markovian short rates in fractional bond markets. Statistics & Probability Letters, 70(3):211–222, December 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002585.

Gapeev:2008:IOM

[Gap08] Pavel V. Gapeev. The integral option in a model with jumps. Statistics & Probability Letters, 78(16):2623–2631, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001697.

Gassem:2009:PES

[Gas09] Anis Gassem. On parameter estimation for switching diffusion process. Statistics & Probability Letters, 79(24):2484– 2492, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003435.

Gatto:2008:SAP

[Gat08] Riccardo Gatto. A saddlepoint approximation to the probability of ruin in the compound Poisson process with diffusion. Statistics & Probability Letters, 78(13):1948–1954, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000631. See erra-

tum [Gat09].

Gatto:2009:ESA

[Gat09] Riccardo Gatto. Erratum to: “A saddlepoint approximation to the probability of ruin in the compound Poisson process with diffusion” [Statistics & Probability Letters 78 (2008) 1948–1954]. Statistics & Probability Letters, 79(7):997–998, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800549X. See [Gat08].

Giuliano-Antonini:2008:TDC

[GAW08] R. Giuliano-Antonini and M. Weber. The theta-dependence coefficient and an Almost Sure Limit Theorem for random iterative models. Statistics & Probability Letters, 78(5):564–575, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700288X.

Gonzalez-Barrios:2003:CPB

[GBQ03] José María González-Barrios and Adolfo J. Quiroz. A clustering procedure based on the comparison between the k nearest neighbors graph and the minimal spanning tree. Statistics & Probability Letters, 62(1):23–34, March 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202004212.

Genon-Catalot:2003:NLE

[GC03] Valentine Genon-Catalot. A non-linear explicit filter. Statistics & Probability Letters, 61(2):145–154, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003449.

Ghosh:2004:ODC

[GD04] Himadri Ghosh and Ashish Das. Optimal diallel cross designs for the interval estimation of heredity. Statistics & Probability Letters, 67(1):47–55, March 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000045.

Gorostiza:2001:SAF

[GDF01] L. G. Gorostiza and E. Díaz-Francés. Species accumulation functions and pure birth processes. Statistics & Probability Letters, 55(2):221–226, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001481.

Genton:2000:RSB

[GdL00] Marc G. Genton and Xavier de Luna. Robust simulationbased estimation. Statistics & Probability Letters, 48(3):253– 259, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000043.

Ghosh:2006:SEP

[GDM06] Himadri Ghosh, Ashish Das, and C. K. Midha. On some efficient partial diallel cross designs. Statistics & Probability Letters, 76(16):1768–1774, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001404.

Gillen:2007:NCW

[GE07] Daniel L. Gillen and Scott S. Emerson. Nontransitivity in a class of weighted logrank statistics under nonproportional hazards. Statistics & Probability Letters, 77(2):123–130, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002112.

Geluk:2004:ASI

[Gel04] Jaap Geluk. Asymptotics in the symmetrization inequality. Statistics & Probability Letters, 69(1):63–68, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001592.

Geluk:2009:SCP

[Gel09] Jaap Geluk. Some closure properties for subexponential distributions. Statistics & Probability Letters, 79(8):1108–1111, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209000029.

Georgiou:2007:NTV

[Geo07] Stelios D. Georgiou. New two-variable full orthogonal designs and related experiments with linear regression models. Statistics & Probability Letters, 77(1):25–31, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520600191X.

Gervini:2002:IFS

[Ger02] Daniel Gervini. The influence function of the Stahel–Donoho estimator of multivariate location and scatter. Statistics & Probability Letters, 60(4):425–435, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003152.

Gerami:2008:IFE

[Ger08] A. Gerami. Incomplete factorial experiments in completely randomized and randomized complete block designs. Statistics & Probability Letters, 78(14):2058–2065, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000783.

Gupta:2000:CRM

[GG00] Ramesh C. Gupta and Pushpa L. Gupta. On the crossings of reliability measures. Statistics & Probability Letters, 46(3):301–305, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001248.

Ghomrasni:2002:ESI

[GG02] Raouf Ghomrasni and Svend Erik Graversen. An extension of Seshadri’s identities for Brownian motion. Statistics & Probability Letters, 59(4):379–384, October 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002353.

Garel:2000:RBP

[GH00] Bernard Garel and Marc Hallin. Rank-based partial autocorrelations are not asymptotically distribution-free. Statistics & Probability Letters, 47(3):219–227, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001595.

Ghorbel:2005:FSD

[GH05a] M. Ghorbel and T. Huillet. Fragment size distributions in random fragmentations with cutoff. Statistics & Probability Letters, 71(1):47–60, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002834.

Gut:2005:RVS

[GH05b] Allan Gut and Jürg Hüsler. Realistic variation of shock models. Statistics & Probability Letters, 74(2):187–204, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001756.

Ghazal:2000:RFE

[Gha00] G. A. Ghazal. Recurrence formula for expectations of products of bilinear forms and expectations of bilinear forms and random matrices. Statistics & Probability Letters, 48(1):1–9, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001807.

Genton:2001:MSN

[GHL01] Marc G. Genton, Li He, and Xiangwei Liu. Moments of skewnormal random vectors and their quadratic forms. Statistics & Probability Letters, 51(4):319–325, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001644.

Ghosh:2003:PEG

[Gho03] Yashowanto N. Ghosh. On the probability of extinction of a Galton–Watson process. Statistics & Probability Letters, 65(3):195–198, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002517.

Gibilisco:2008:SI

[GII08] Paolo Gibilisco, Daniele Imparato, and Tommaso Isola. Stam inequality on Z_{n}. Statistics & Probability Letters, 78(13):1851–1856, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000503.

Girard:2004:ANE

[Gir04] Stéphane Girard. On the asymptotic normality of the L_{1}-error for Haar series estimates of Poisson point processes boundaries. Statistics & Probability Letters, 66(1):81–90, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003109.

Giraudo:2009:AFF

[Gir09] Maria Teresa Giraudo. An approximate formula for the firstcrossing-time density of a Wiener process perturbed by random jumps. Statistics & Probability Letters, 79(13):1559–1567, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001205.

Gut:2001:TWC

[GJ01] Allan Gut and Svante Janson. Tightness and weak convergence for jump processes. Statistics & Probability Letters, 52(1):101–107, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002406.

Gatto:2002:SAT

[GJ02] Riccardo Gatto and S. Rao Jammalamadaka. A saddlepoint approximation for testing exponentiality against some increasing failure rate alternatives. Statistics & Probability Letters, 58(1):71–81, May 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001177.

Girard:2008:NEV

[GJ08] Stéphane Girard and Pierre Jacob. A note on extreme values and kernel estimators of sample boundaries. Statistics & Probability Letters, 78(12):1634–1638, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000199.

Gao:2009:MDS

[GJ09] Fuqing Gao and Hui Jiang. Moderate deviations for squared radial Ornstein–Uhlenbeck process. Statistics & Probability Letters, 79(11):1378–1386, June 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000765.

Guerrero:2001:DGB

[GJP01] Víctor M. Guerrero, Rodrigo Juárez, and Pilar Poncela. Data graduation based on statistical time series methods. Statistics & Probability Letters, 52(2):169–175, April 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002133.

Guo:2006:SIH

[GJZ06] Dongmei Guo, Shaolin Ji, and Huaizhong Zhao. On the solvability of infinite horizon forward-backward stochastic differential equations with absorption coefficients. Statistics & Probability Letters, 76(18):1954–1960, December 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001659.

Grigorescu:2004:PCM

[GK04a] Ilie Grigorescu and Min Kang. Path collapse for multidimensional Brownian motion with rebirth. Statistics & Probability Letters, 70(3):199–209, December 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002573.

Gushchin:2004:OGB

[GK04b] Alexander A. Gushchin and Uwe Küchler. On oscillations of the geometric Brownian motion with time-delayed drift. Statistics & Probability Letters, 70(1):19–24, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001944.

Gonchigdanzan:2009:FLP

[GK09] Khurelbaatar Gonchigdanzan and Kamil M. Kosiński. On the functional limits for partial sums under stable law. Statistics & Probability Letters, 79(17):1818–1822, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001850.

Georgiou:2003:CMT

[GKM03] S. Georgiou, C. Koukouvinos, and P. Mantas. Construction methods for three-level supersaturated designs based on weighing matrices. Statistics & Probability Letters, 63(4): 339–352, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203000993.

Guo:2000:ITT

[GL00] Jiin-Huarng Guo and Wei-Ming Luh. An invertible transformation two-sample trimmed t-statistic under heterogeneity and nonnormality. Statistics & Probability Letters, 49(1):1–7, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000225.

Geng:2003:FPP

[GL03] Zhi Geng and Kaican Li. Factorization of posteriors and partial imputation algorithm for graphical models with missing data. Statistics & Probability Letters, 64(4):369–379, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001810.

Gittenberger:2004:RBB

[GL04] Bernhard Gittenberger and Guy Louchard. Reflected Brownian bridge local time conditioned on its local time at the origin. Statistics & Probability Letters, 68(1):51–60, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000689.

Gilliland:2008:NCI

[GL08] Dennis Gilliland and Mingfei Li. A note on confidence intervals for the power of t-test. Statistics & Probability Letters, 78(5):488–489, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207002659.

Golnabi:2009:FPS

[GLC09] Saeed Golnabi, Martin S. Levy, and James J. Cochran. Finitizing power series distributions. Statistics & Probability Letters, 79(17):1829–1832, September 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001886.

Goncalves:2009:MAD

[GLML09] Esmeralda Gonçalves, Joana Leite, and Nazaré Mendes-Lopes. A mathematical approach to detect the Taylor property in TARCH processes. Statistics & Probability Letters, 79(5):602–610, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004720.

Gouet:2008:LLN

[GLS08] Raúl Gouet, F. Javier López, and Gerardo Sanz. Laws of large numbers for the number of weak records. Statistics & Probability Letters, 78(14):2010–2017, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000709.

Gao:2008:MIS

[GLT08] Jiti Gao, Zudi Lu, and Dag Tjøstheim. Moment inequalities for spatial processes. Statistics & Probability Letters, 78(6):687–697, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003136.

Gustafson:2000:PBA

[GLV00] Paul Gustafson, Nhu D. Le, and Marc Vallée. Parametric Bayesian analysis of case-control data with imprecise exposure measurements. Statistics & Probability Letters, 47(4):357–363, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001789.

Gilliland:2007:NAP

[GLX07] Dennis Gilliland, Shlomo Levental, and Yimin Xiao. A note on absorption probabilities in one-dimensional random walk via complex-valued martingales. Statistics & Probability Letters, 77(11):1098–1105, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700048X.

Guglielmi:2000:AFM

[GM00] Alessandra Guglielmi and Eugenio Melilli. Approximating de Finetti’s measures for partially exchangeable sequences. Statistics & Probability Letters, 48(3):309–315, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000134.

Gilmour:2003:BDC

[GM03a] Steven G. Gilmour and Roger Mead. A Bayesian design criterion for locating the optimum point on a response surface. Statistics & Probability Letters, 64(3):235–242, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001548.

Grigelionis:2003:FMS

[GM03b] Bronius Grigelionis and Vigirdas Mackevicius. The finiteness of moments of a stochastic exponential. Statistics & Probability Letters, 64(3):243–248, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520300155X.

Ghosh:2006:PPC

[GM06] Yashowanto N. Ghosh and Bhramar Mukherjee. On probabilistic properties of conditional medians and quantiles. Statistics & Probability Letters, 76(16):1775–1780, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001416.

Gamage:2008:IMS

[GM08] Jinadasa Gamage and Thomas Mathew. Inference on mean sub-vectors of two multivariate normal populations with unequal covariance matrices. Statistics & Probability Letters, 78(4):420–425, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520700257X.

Gonzalez:2004:LDS

[GMdP04] M. González, M. Molina, and I. del Puerto. Limiting distribution for subcritical controlled branching processes with random control function. Statistics & Probability Letters, 67(3):277–284, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000392.

Gonzalez-Manteiga:2002:NVS

[GMdRV02] Wenceslao González-Manteiga, Alejandro Quintela del Río, and Philippe Vieu. A note on variable selection in nonparametric regression with dependent data. Statistics & Probability Letters, 57(3):259–268, April 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000561.

Gonzalez:2000:LBS

[GMM00] M. González, M. Molina, and M. Mota. Limit behaviour for a subcritical bisexual Galton–Watson branching process with immigration. Statistics & Probability Letters, 49(1):19–24, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000262.

Genest:2002:USL

[GMM02] Christian Genest, Étienne Marceau, and Mhamed Mesfioui. Upper stop-loss bounds for sums of possibly dependent risks with given means and variances. Statistics & Probability Letters, 57(1):33–41, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000391.

Gonzalez:2004:MPS

[GMM04] Miguel González, Rodrigo Martínez, and Manuel Mota. Multitype population size-dependent branching processes with dependent offspring. Statistics & Probability Letters, 70(2): 145–154, November 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204002342.

Grzenda:2007:CPU

[GMZ07] Wioletta Grzenda, Dariusz Majerek, and Wiesław Zieb̧a. Comments to the paper “On uniform integrability of random variables”. Statistics & Probability Letters, 77(16):1644–1646, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207001459.

Gomes:2008:ACM

[GN08] M. Ivette Gomes and Cláudia Neves. Asymptotic comparison of the mixed moment and classical extreme value index estimators. Statistics & Probability Letters, 78(6):643–653, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207003070.

Gneri:2003:RES

[Gne03] Mario Antonio Gneri. Resistant estimators for stationary ergodic stochastic processes. Statistics & Probability Letters, 64(1):97–103, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001445.

Gajek:2001:IST

[GO01a] L. Gajek and A. Okolewski. Improved Steffensen type bounds on expectations of record statistics. Statistics & Probability Letters, 55(2):205–212, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001286.

Garcia:2001:MRR

[GO01b] Gloria García and Josep M. Oller. Minimum Riemannian risk equivariant estimator for the univariate normal model. Statistics & Probability Letters, 52(1):109–113, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001930.

Glynn:2002:HIU

[GO02] Peter W. Glynn and Dirk Ormoneit. Hoeffding’s inequality for uniformly ergodic Markov chains. Statistics & Probability Letters, 56(2):143–146, January 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001584.

Gomes:2003:CEP

[GO03] M. Ivette Gomes and Orlando Oliveira. Censoring estimators of a positive tail index. Statistics & Probability Letters, 65(3):147–159, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002116.

Goenner:2009:IKP

[Goe09] Lorenz Goenner. On an inequality by Kulikova and Prokhorov. Statistics & Probability Letters, 79(14):1630–1633, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001485.

Gombay:2000:SCP

[Gom00] Edit Gombay. Sequential change-point detection with likelihood ratios. Statistics & Probability Letters, 49(2):195–204, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000481.

Gonchigdanzan:2008:ASL

[Gon08] Khurelbaatar Gonchigdanzan. An almost sure limit theorem for the product of partial sums with stable distribution. Statistics & Probability Letters, 78(18):3170–3175, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002873.

Gomez:2009:EGB

[GOPB09] Héctor W. Gómez, Juan F. Olivares-Pacheco, and Heleno Bolfarine. An extension of the generalized Birnbaum–Saunders distribution. Statistics & Probability Letters, 79(3):331–338, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800415X.

Gorgens:2002:NCR

[Gør02] Tue Gørgens. Nonparametric comparison of regression curves by local linear fitting. Statistics & Probability Letters, 60(1):81–89, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202002833.

Gordon:2007:UBP

[Gor07] Alexander Y. Gordon. Unimprovability of the Bonferroni procedure in the class of general step-up multiple testing procedures. Statistics & Probability Letters, 77(2):117–122, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002100.

Gordon:2009:IBG

[Gor09] Alexander Y. Gordon. Inequalities between generalized familywise error rates of a multiple testing procedure. Statistics & Probability Letters, 79(19):1996–2004, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002326.

Garren:2000:ANM

[GP00a] Steven T. Garren and Shyamal D. Peddada. Asymptotic normality in multivariate nonlinear regression and multivariate generalized linear regression models under repeated measurements with missing data. Statistics & Probability Letters, 48(3):293–302, July 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000109.

Geluk:2000:AOE

[GP00b] J. L. Geluk and Liang Peng. An adaptive optimal estimate of the tail index for MA(l) time series. Statistics & Probability Letters, 46(3):217–227, February 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000991.

Guyon:2000:ISD

[GP00c] Xavier Guyon and Olivier Perrin. Identification of space deformation using linear and superficial quadratic variations. Statistics & Probability Letters, 47(3):307–316, April 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001716.

Garcia:2001:IMN

[GP01] Nancy Lopes Garcia and José Luis Palacios. On inverse moments of nonnegative random variables. Statistics & Probability Letters, 53(3):235–239, June 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000086.

Galeano:2005:NPI

[GP05a] Pedro Galeano and Daniel Peña. A note on prediction and interpolation errors in time series. Statistics & Probability Letters, 73(1):71–78, June 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001112.

Gimenez:2005:ECC

[GP05b] Patricia Giménez and María Laura Patat. Estimation in comparative calibration models with replicate measurement. Statistics & Probability Letters, 71(2):155–164, February 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204002901.

Galtchouk:2006:AEE

[GP06] L. Galtchouk and S. Pergamenshchikov. Asymptotically efficient estimates for nonparametric regression models. Statistics & Probability Letters, 76(8):852–860, April 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004025.

Galeano:2007:CBM

[GP07] Pedro Galeano and Daniel Peña. On the connection between model selection criteria and quadratic discrimination in ARMA time series models. Statistics & Probability Letters, 77(9):896–900, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000260.

Park:2002:EMM

[gPB02] Jeong gun Park and I. V. Basawa. Estimation for mixtures of Markov processes. Statistics & Probability Letters, 59(3):235–244, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520200127X.

Galea:2005:DSN

[GPC05] Manuel Galea, Gilberto A. Paula, and Francisco José A. Cysneiros. On diagnostics in symmetrical nonlinear models. Statistics & Probability Letters, 73(4):459–467, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001562.

Gomes:2009:NAV

[GPC09] M. Ivette Gomes, Dinis Pestana, and Frederico Caeiro. A note on the asymptotic variance at optimal levels of a bias-corrected Hill estimator. Statistics & Probability Letters, 79(3):295–303, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800401X.

Genest:2006:JAB

[GQR06] Christian Genest, Jean-François Quessy, and Bruno Rémillard. On the joint asymptotic behavior of two rank-based estimators of the association parameter in the gamma frailty model. Statistics & Probability Letters, 76(1):10–18, January 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002609.

Gomez:2007:NFS

[GQT07] Héctor W. Gómez, Fernando A. Quintana, and Francisco J. Torres. A new family of slash-distributions with elliptical contours. Statistics & Probability Letters, 77(7):717–725, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206003415. See erra-

tum [GV08].

Gupta:2000:REU

[GR00] Rameshwar D. Gupta and Donald St. P. Richards. Radix expansions and the uniform distribution. Statistics & Probability Letters, 46(3):263–270, February 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900111X.

Genest:2001:MPI

[GR01a] Christian Genest and Louis-Paul Rivest. On the multivariate probability integral transformation. Statistics & Probability Letters, 53(4):391–399, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000475.

Gijbels:2001:NLS

[GR01b] Irène Gijbels and Valentin Rousson. A nonparametric leastsquares test for checking a polynomial relationship. Statistics & Probability Letters, 51(3):253–261, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001528.

Gapeev:2006:OSP

[GR06] Pavel V. Gapeev and Markus Reiß. An optimal stopping problem in a diffusion-type model with delay. Statistics & Probability Letters, 76(6):601–608, March 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003524.

Gottard:2007:CGM

[GR07] Anna Gottard and Carla Rampichini. Chain graphs for multilevel models. Statistics & Probability Letters, 77(3):312–318, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002380.

Guo:2008:OBH

[GR08] Wenge Guo and M. Bhaskara Rao. On optimality of the Benjamini–Hochberg procedure for the false discovery rate. Statistics & Probability Letters, 78(14):2024–2030, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000734.

Grafstrom:2009:RPS

[Gra09a] Anton Grafström. Repeated Poisson sampling. Statistics & Probability Letters, 79(6):760–764, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005117.

Grau:2009:MUN

[Gra09b] Daniel Grau. Moments of the unbalanced non-central chisquare distribution. Statistics & Probability Letters, 79(3): 361–367, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004203.

Gregorio:2009:PEP

[Gre09] Alessandro De Gregorio. Parametric estimation for planar random flights. Statistics & Probability Letters, 79(20): 2193–2199, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002685.

Griffith:2002:SFS

[Gri02] Daniel A. Griffith. A spatial filtering specification for the auto-Poisson model. Statistics & Probability Letters, 58(3):245–251, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000998.

Grigelionis:2003:PMU

[Gri03] Bronius Grigelionis. On point measures of ϵ-upcrossings for stationary diffusions. Statistics & Probability Letters, 61 (4):403–410, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202004005.

Grigelionis:2008:PMM

[Gri08] Bronius Grigelionis. On Pólya mixtures of multivariate Gaussian distributions. Statistics & Probability Letters, 78(12):1459–1465, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207004300.

Gross:2003:RRE

[Gro03] Jürgen Groß. Restricted ridge estimation. Statistics & Probability Letters, 65(1):57–64, October 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002256.

Gu:2007:CRF

[GRT07] Wentao Gu, George G. Roussas, and Lanh T. Tran. On the convergence rate of fixed design regression estimators for negatively associated random variables. Statistics & Probability Letters, 77(12):1214–1224, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000867.

Gannoun:2002:NPS

[GS02] Ali Gannoun and Jérôme Saracco. A new proof of strong consistency of kernel estimation of density function and mode under random censorship. Statistics & Probability Letters, 59(1):61–66, August 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001669.

Galambos:2003:CRL

[GS03a] J. Galambos and I. Simonelli. Comments on a recent limit theorem of Quine. Statistics & Probability Letters, 63(1):89–95, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000567.

Grasshoff:2003:APO

[GS03b] Ulrike Graßhoff and Rainer Schwabe. On the analysis of paired observations. Statistics & Probability Letters, 65(1):7–12, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002128.

Gao:2007:APD

[GS07] Sujuan Gao and Jianzhao Shen. Asymptotic properties of a double penalized maximum likelihood estimator in logistic regression. Statistics & Probability Letters, 77(9):925–930, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000351.

Gani:2008:URA

[GS08a] J. Gani and R. J. Swift. An unexpected result in an approximate carrier-borne epidemic process. Statistics & Probability Letters, 78(14):2116–2120, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000850.

Gordon:2008:OHP

[GS08b] Alexander Y. Gordon and Peter Salzman. Optimality of the Holm procedure among general step-down multiple testing procedures. Statistics & Probability Letters, 78(13):1878–1884, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000540.

Gut:2009:AMZ

[GS09] Allan Gut and Ulrich Stadtmüller. An asymmetric Marcinkiewicz– Zygmund LLN for random fields. Statistics & Probability Letters, 79(8):1016–1020, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005695.

Garcia-Soidan:2003:LLR

[GSGMFB03] Pilar H. García-Soidán, Wenceslao González-Manteiga, and Manuel Febrero-Bande. Local linear regression estimation of the variogram. Statistics & Probability Letters, 64(2):169–179, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001494.

Gupta:2009:CEU

[GSKP09] V. K. Gupta, Poonam Singh, Basudev Kole, and Rajender Parsad. Construction of efficient unbalanced mixedlevel supersaturated designs. Statistics & Probability Letters, 79(22):2359–2366, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003071.

Gerardi:2000:EMM

[GST00] Anna Gerardi, Fabio Spizzichino, and Barbara Torti. Exchangeable mixture models for lifetimes: the role of “occupation numbers”. Statistics & Probability Letters, 49(4):365–375, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000699.

Ge:2007:HTP

[GSTS07] Yongchao Ge, Stuart C. Sealfon, Chi-Hong Tseng, and Terence P. Speed. A Holm-type procedure controlling the false discovery rate. Statistics & Probability Letters, 77 (18):1756–1762, December 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001642.

Goldenshluger:2004:EED

[GT04] A. Goldenshluger and A. Tsybakov. Estimating the endpoint of a distribution in the presence of additive observation errors. Statistics & Probability Letters, 68(1):39–49, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000677.

Gorostiza:2008:SLR

[GT08] Luis G. Gorostiza and Martha Takane. Some limits related to random iterations of a lamplighter group. Statistics & Probability Letters, 78(1):21–26, January 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001666.

Garibotti:2006:LLA

[GTH06] Gilda Garibotti, John V. Tsimikas, and Joseph Horowitz. Likelihood-look-ahead inference on the equilibrium distribution of Markov chains. Statistics & Probability Letters, 76 (10):991–1000, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004384.

Guan:2007:LSC

[Gua07] Yongtao Guan. A least-squares cross-validation bandwidth selection approach in pair correlation function estimations. Statistics & Probability Letters, 77(18):1722–1729, December 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001575.

Guerrero:2007:TSS

[Gue07] Victor M. Guerrero. Time series smoothing by penalized least squares. Statistics & Probability Letters, 77(12):1225–1234, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000855.

Guillas:2001:RCA

[Gui01] Serge Guillas. Rates of convergence of autocorrelation estimates for autoregressive Hilbertian processes. Statistics & Probability Letters, 55(3):281–291, December 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001511.

Guo:2008:NCP

[Guo08] Zhi Jun Guo. A note on the CIR process and the existence of equivalent martingale measures. Statistics & Probability Letters, 78(5):481–487, April 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002647.

Gupta:2009:VDR

[Gup09] Bhupendra Gupta. Vertex degree of random geometric graph on exponentially distributed points. Statistics & Probability Letters, 79(19):2020–2027, October 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002351.

Guttman:2000:CEM

[Gut00] Irwin Guttman. A contrast of EB, modal and EM-algorithm estimates arising in the one-way analysis of variance situation. Statistics & Probability Letters, 48(1):33–58, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001893.

Gut:2006:GRT

[Gut06] Allan Gut. Gnedenko–Raikov’s theorem, central limit theory, and the Weak Law of Large Numbers. Statistics & Probability Letters, 76(17):1935–1939, November 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001611.

Genest:2002:TOK

[GV02a] Christian Genest and François Verret. The TP_{2} ordering of Kimeldorf and Sampson has the normal-agreeing property. Statistics & Probability Letters, 57(4):387–391, May 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202000950.

Gyorfi:2002:ADG

[GV02b] L. Györfi and I. Vajda. Asymptotic distributions for goodnessof-fit statistics in a sequence of multinomial models. Statistics & Probability Letters, 56(1):57–67, January 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001729.

Giummole:2006:RPL

[GV06] F. Giummolè and L. Ventura. Robust prediction limits based on M-estimators. Statistics & Probability Letters, 76 (16):1735–1740, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001295.

Gomez:2008:ENF

[GV08] Héctor W. Gómez and Osvaldo Venegas. Erratum to: “A new family of slash-distributions with elliptical contours” [Statist. Probab. Lett. 77 (2007) 717–725]. Statistics & Probability Letters, 78(14):2273–2274, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001272. See [GQT07].

Gomez-Villegas:2008:MPN

[GVGP08] Miguel A. Gómez-Villegas and Beatriz González-Pérez. The multivariate point null testing problem: A Bayesian discussion. Statistics & Probability Letters, 78(17):3070–3074, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002691.

Gregori:2004:MFS

[GvLM04] P. Gregori, M. N. M. van Lieshout, and J. Mateu. Mixture formulae for shot noise weighted point processes. Statistics & Probability Letters, 67(4):311–320, May 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000604.

Gomez-Villegas:2004:ARB

[GVMSN04] Miguel A. Gómez-Villegas, Paloma Maín, Luis Sanz, and Hilario Navarro. Asymptotic relationships between posterior probabilities and p-values using the hazard rate. Statistics & Probability Letters, 66(1):59–66, January 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003080.

Gomez-Villegas:2000:CPT

[GVS00] Miguel A. Gómez-Villegas and Luis Sanz. ϵ-contaminated priors in testing point null hypothesis: a procedure to determine the prior probability. Statistics & Probability Letters, 47(1):53–60, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001376.

Gao:2005:RCF

[GW05] Fuqing Gao and Qinghua Wang. The rate of convergence in the functional limit theorem for increments of a Brownian motion. Statistics & Probability Letters, 73(2):165–177, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204003098.

Giovagnoli:2008:SOD

[GW08] A. Giovagnoli and H. P. Wynn. Stochastic orderings for discrete random variables. Statistics & Probability Letters, 78(16):2827–2835, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002083.

Gao:2009:RPL

[GW09] Qingwu Gao and Yuebao Wang. Ruin probability and local ruin probability in the random multi-delayed renewal risk model. Statistics & Probability Letters, 79(5):588–596, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208004707.

Glickman:2008:DPT

[GX08] Theodore S. Glickman and Feng Xu. The distribution of the product of two triangular random variables. Statistics & Probability Letters, 78(16):2821–2826, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002071.

Glaz:2006:MSS

[GZ06] Joseph Glaz and Zhenkui Zhang. Maximum scan score-type statistics. Statistics & Probability Letters, 76(13):1316–1322, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000253.

Hirano:2003:NOS

[HA03]  Katuomi Hirano and Sigeo Aki.  Number of occur-

rences of subpattern until the first appearance of a pattern and geometric distribution. Statistics & Probability Letters, 65(3):259–262, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002694.

Haas:2008:ASM

[Haa08] Markus Haas. The autocorrelation structure of the Markovswitching asymmetric power GARCH process. Statistics & Probability Letters, 78(12):1480–1489, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004476.

Haas:2009:PVC

[Haa09] Markus Haas. Persistence in volatility, conditional kurtosis, and the Taylor property in absolute value GARCH processes. Statistics & Probability Letters, 79(15):1674–1683, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001540.

Hadjicostas:2002:SIP

[Had02] Petros Hadjicostas. Symmetric and isomorphic properties of qualitative probability structures on a finite set. Statistics & Probability Letters, 56(3):309–319, February 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001997.

Hadjicostas:2003:CLR

[Had03] Petros Hadjicostas. Consistency of logistic regression coefficient estimates calculated from a training sample. Statistics & Probability Letters, 62(3):293–303, April 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000361.

Hafidi:2006:SSC

[Haf06] Bezza Hafidi. A small-sample criterion based on Kullback’s symmetric divergence for vector autoregressive modeling. Statistics & Probability Letters, 76(15):1647–1654, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001209.

Hagmark:2009:NCC

[Hag09] Per-Erik Hagmark. A new concept for count distributions. Statistics & Probability Letters, 79(8):1120–1124, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209000042.

Haiman:2003:EVT

[Hai03] George Haiman. Extreme values of the tent map process. Statistics & Probability Letters, 65(4):451–456, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520300302X.

Halidias:2009:RCE

[Hal09] Nikolaos Halidias. Remarks and corrections on “An existence theorem for stochastic functional differential equations with delays under weak assumptions”, Statistics & Probability Letters 78, 2008, by N. Halidias and Y. Ren. Statistics & Probability Letters, 79(20):2220–2222, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002739. See [HR08a].

Han:2003:NOD

[Han03] Cong Han. A note on optimal designs for a two-part model. Statistics & Probability Letters, 65(4):343–351, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002670.

Haner:2005:NE

[Han05] Matthew Haner. The nonorthogonal estimator. Statistics & Probability Letters, 73(4):347–356, July 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001306.

Hanagal:2008:MHB

[Han08] David D. Hanagal. Modelling heterogeneity for bivariate survival data by the log-normal distribution. Statistics & Probability Letters, 78(9):1101–1109, July 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700380X.

Hartigan:2004:UPC

[Har04] J. A. Hartigan. Uniform priors on convex sets improve risk. Statistics & Probability Letters, 67(4):285–288, May 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000276.

Harremoes:2008:SNM

[Har08] Peter Harremoës. Some new maximal inequalities. Statistics & Probability Letters, 78(16):2776–2780, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002009.

Hashorva:2001:ARF

[Has01] Enkelejd Hashorva. Asymptotic results for FGM random sequences. Statistics & Probability Letters, 54(4):417–425, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001201.

Hashorva:2002:RDM

[Has02] Enkelejd Hashorva. Remarks on domination of maxima. Statistics & Probability Letters, 60(1):101–109, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202002870.

Hashorva:2004:BMI

[Has04] Enkelejd Hashorva. Bivariate maximum insurance claim and related point processes. Statistics & Probability Letters, 69 (2):117–128, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204001634.

Hashorva:2005:ETA

[Has05] Enkelejd Hashorva. Elliptical triangular arrays in the maxdomain of attraction of Hüsler–Reiss distribution. Statistics & Probability Letters, 72(2):125–135, April 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000222.

Hashorva:2006:NCB

[Has06a] Enkelejd Hashorva. A novel class of bivariate max-stable distributions. Statistics & Probability Letters, 76(10):1047–1055, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205004451.

Hashorva:2006:MHR

[Has06b] Enkelejd Hashorva. On the multivariate Hüsler–Reiss distribution attracting the maxima of elliptical triangular arrays. Statistics & Probability Letters, 76(18):2027–2035, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001866.

Hashorva:2006:RVE

[Has06c] Enkelejd Hashorva. On the regular variation of elliptical random vectors. Statistics & Probability Letters, 76(14):

1427–1434, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000496.

Hashorva:2009:AKT

[Has09] Enkelejd Hashorva. Asymptotics for Kotz Type III elliptical distributions. Statistics & Probability Letters, 79(7):927–935, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005476.

Hattori:2006:SPM

[Hat06] Satoshi Hattori. Some properties of misspecified additive hazards models. Statistics & Probability Letters, 76(15): 1641–1646, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001180.

Hazelton:2000:MDE

[Haz00] Martin L. Hazelton. Marginal density estimation from incomplete bivariate data. Statistics & Probability Letters, 47(1):75–84, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001406.

Hwang:2001:NTS

[HB01] Sun Y. Hwang and I. V. Basawa. Nonlinear time series contiguous to AR(1) processes and a related efficient test for linearity. Statistics & Probability Letters, 52(4):381–390, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200002340.

Hwang:2004:SMS

[HB04] S. Y. Hwang and I. V. Basawa. Stationarity and moment structure for Box–Cox transformed threshold GARCH(1,1) processes. Statistics & Probability Letters, 68(3):209–220, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520300292X.

Hofmann:2005:CHF

[HBA05] Glenn Hofmann, N. Balakrishnan, and Jafar Ahmadi. Characterization of hazard function factorization by Fisher information in minima and upper record values. Statistics & Probability Letters, 72(1):51–57, April 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000155.

Hwang:2006:LSE

[HBK06] S. Y. Hwang, I. V. Basawa, and Tae Yoon Kim. Least squares estimation for critical random coefficient firstorder autoregressive processes. Statistics & Probability Letters, 76(3):310–317, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003020.

Huh:2002:ERF

[HC02] J. Huh and K. C. Carrière. Estimation of regression functions with a discontinuity in a derivative with local polynomial fits. Statistics & Probability Letters, 56(3):329–343, February 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000172.

Han:2004:NOD

[HC04] Cong Han and Kathryn Chaloner. A note on optimal designs for two or more treatment groups. Statistics & Probability Letters, 69(1):81–89, August 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001683.

Huang:2006:QFM

[HC06] Wen-Jang Huang and Yan-Hau Chen. Quadratic forms of multivariate skew normal-symmetric distributions. Statistics & Probability Letters, 76(9):871–879, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004037.

Huang:2007:GSC

[HC07] Wen-Jang Huang and Yan-Hau Chen. Generalized skew-Cauchy distribution. Statistics & Probability Letters, 77 (11):1137–1147, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000661.

Hwang:2009:MLS

[HC09] S. Y. Hwang and M. S. Choi. Modeling and large sample estimation for multi-casting autoregression. Statistics & Probability Letters, 79(18):1943–1950, September 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002119.

Hong:2000:WLR

[HCSV00] Dug Hun Hong, Manuel Ordóñez Cabrera, Soo Hak Sung, and Andrei I. Volodin. On the weak law for randomly indexed partial sums for arrays of random elements in martingale type p Banach spaces. Statistics & Probability Letters, 46(2):177–185, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001030.

Hu:2001:CRW

[HCV01] Tien-Chung Hu, Manuel Ordóñez Cabrera, and Andrei I. Volodin. Convergence of randomly weighted sums of Banach space valued random elements and uniform integrability concerning the random weights. Statistics & Probability Letters, 51(2):155–164, January 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001462.

Hu:2005:APB

[HCV05] Tien-Chung Hu, Manuel Ordóñez Cabrera, and Andrei Volodin. Asymptotic probability for the bootstrapped means deviations from the sample mean. Statistics & Probability Letters, 74(2):178–186, September 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001744.

Huang:2007:IAN

[HCW07] Yufen Huang, Ching-Ren Cheng, and Tai-Ho Wang. Influence analysis of non-Gaussianity by applying projection pursuit. Statistics & Probability Letters, 77(14):1515–1521, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520700123X.

Hu:2008:EBP

[HCW08a] Shuhe Hu, Guijing Chen, and Xuejun Wang. On extending the Brunk–Prokhorov strong law of large numbers for martingale differences. Statistics & Probability Letters, 78(18):3187–3194, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002976.

Hwang:2008:GFT

[HcW08b] Yi-Ting Hwang and Chun chao Wang. A goodness of fit test for left-truncated and right-censored data. Statistics & Probability Letters, 78(15):2420–2425, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001430.

Cai:2009:UBN

[hCW09] Guang hui Cai and Jian-Feng Wang. Uniform bounds in normal approximation under negatively associated random fields. Statistics & Probability Letters, 79(2):215–222, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003726.

He:2008:PEM

[He08] Xin He. A problem on extreme magnitudes of characteristic functions: The general case. Statistics & Probability Letters, 78(12):1514–1516, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000059.

He:2009:SUC

[He09] Hui He. Strong uniqueness for a class of singular SDEs for catalytic branching diffusions. Statistics & Probability Letters, 79(2):182–187, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208003660.

Heinrich:2001:SDT

[Hei01] P. Heinrich. The size of diagonal two-order Gaussian chaoses. Statistics & Probability Letters, 51(1):35–39, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001358.

Hellmund:2009:CRS

[Hel09] Gunnar Hellmund. Completely random signed measures. Statistics & Probability Letters, 79(7):894–898, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005324.

Hengjian:2000:PTD

[Hen00] Cui Hengjian. A projection type distribution function and quantile estimates in the presence of auxiliary information. Statistics & Probability Letters, 48(1):91–100, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001935.

Haiyan:2004:EHT

[HF04a] Chen Haiyan and Zhang Fuji. The expected hitting times for graphs with cutpoints. Statistics & Probability Letters, 66(1):9–17, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002876.

Huang:2004:NMB

[HF04b] Kuo-Nan Huang and Shu-Kai S. Fan. A note on minimum bias estimation in response surfaces. Statistics & Probability Letters, 70(1):71–85, October 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002299.

Hansen:2000:SET

[HH00a] Jennie C. Hansen and Paul Hulse. Subadditive ergodic theorems for random sets in infinite dimensions. Statistics & Probability Letters, 50(4):409–416, December 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001565.

Hu:2000:NBT

[HH00b] Feifang Hu and Jianhua Hu. A note on breakdown theory for bootstrap methods. Statistics & Probability Letters, 50(1):49–53, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000808.

Hashorva:2001:NPN

[HH01] Enkelejd Hashorva and Jürg Hüsler. On the number of points near the multivariate maxima. Statistics & Probability Letters, 55(2):113–124, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000281.

Hansen:2002:EGD

[HH02a] Niels Richard Hansen and Ernst Hansen. Establishing geometric drift via the Laplace transform of symmetric measures. Statistics & Probability Letters, 60(3):289–295, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002882.

Hashorva:2002:RCP

[HH02b] Enkelejd Hashorva and Jürg Hüsler. Remarks on compound Poisson approximation of Gaussian random sequences. Statistics & Probability Letters, 57(1):1–8, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001869.

Huwang:2002:PCI

[HH02c] Longcheen Huwang and J. T. Gene Hwang. Prediction and confidence intervals for nonlinear measurement error models without identifiability information. Statistics & Probability Letters, 58(4):355–362, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001414.

Hall:2003:IMB

[HH03a] Peter G. Hall and Rob J. Hyndman. Improved methods for bandwidth selection when estimating ROC curves. Statistics & Probability Letters, 64(2):181–189, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001500.

Han:2003:WTP

[HH03b] Qing Han and Katuomi Hirano. Waiting time problem for an almost perfect match. Statistics & Probability Letters, 65(1):39–49, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002177.

Hashorva:2005:MMM

[HH05] Enkelejd Hashorva and Jürg Hüsler. Multiple maxima in multivariate samples. Statistics & Probability Letters, 75 (1):11–17, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002051.

Huang:2006:EBE

[HH06] Wen-Tao Huang and Hui-Hsin Huang. Empirical Bayes estimation of the guarantee lifetime in a two-parameter exponential distribution. Statistics & Probability Letters, 76 (16):1821–1829, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001490.

He:2009:RBR

[HHC09] Kun He, Mingshang Hu, and Zengjing Chen. The relationship between risk measures and Choquet expectations in the framework of g-expectations. Statistics & Probability Letters, 79(4):508–512, February 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004586.

Hilgers:2000:ODB

[Hil00] Ralf-Dieter Hilgers. D-optimal design for Becker’s minimum polynomial. Statistics & Probability Letters, 49(2):175–179, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000468.

Hili:2001:HDE

[Hil01] Ouagnina Hili. Hellinger distance estimation of SSAR models. Statistics & Probability Letters, 53(3):305–314, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000864.

Hildebrand:2002:NVH

[Hil02] Martin Hildebrand. A note on various holding probabilities for random lazy random walks on finite groups. Statistics & Probability Letters, 56(2):199–206, January 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001882.

Hili:2003:HDE

[Hil03] Ouagnina Hili. Hellinger distance estimation of nonlinear dynamical systems. Statistics & Probability Letters, 63(2): 177–184, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203000804.

Huang:2006:CRR

[HJ06] Li-Fei Huang and Richard A. Johnson. Confidence regions for the ratio of percentiles. Statistics & Probability Letters, 76(4):384–392, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500310X.

Hamers:2003:BEM

[HK03a] Michael Hamers and Michael Kohler. A bound on the expected maximal deviation of averages from their means. Statistics & Probability Letters, 62(2):137–144, April 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000026.

Horvath:2003:BAU

[HK03b] Lajos Horváth and Piotr Kokoszka. A bootstrap approximation to a unit root test statistic for heavy-tailed observations. Statistics & Probability Letters, 62(2):163–173, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000075.

Hwang:2008:APO

[HK08] Leng-Cheng Hwang and Rohana J. Karunamuni. Asymptotically pointwise optimal allocation rules in Bayes sequential estimation. Statistics & Probability Letters, 78(15):2490–2495, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001545.

Hattori:2009:ASD

[HK09] Satoshi Hattori and Mai Kato. Approximate subjectdeletion influence diagnostics for Inverse Probability of Censoring Weighted (IPCW) method. Statistics & Probability Letters, 79(17):1833–1838, September 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001898.

Hwang:2007:GLS

[HKLB07] S. Y. Hwang, S. Kim, S. D. Lee, and I. V. Basawa. Generalized least squares estimation for explosive AR(1) processes with conditionally heteroscedastic errors. Statistics & Probability Letters, 77(13):1439–1448, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000740.

Holzmann:2004:ASL

[HKM04] Hajo Holzmann, Susanne Koch, and Aleksey Min. Almost sure limit theorems for U-statistics. Statistics & Probability Letters, 69(3):261–269, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001749.

Horvath:2000:AWB

[HKS00] Lajos Horváth, Piotr Kokoszka, and Josef Steinebach. Approximations for weighted bootstrap processes with an application. Statistics & Probability Letters, 48(1):59–70, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529900190X.

Horvath:2007:SDP

[HKS07] Lajos Horváth, Piotr Kokoszka, and Josef Steinebach. On sequential detection of parameter changes in linear regression. Statistics & Probability Letters, 77(9):885–895, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000223.

Hamedani:2004:SFE

[HKV04] G. G. Hamedani, Eric S. Key, and Hans Volkmer. Solution to a functional equation and its application to stable and stabletype distributions. Statistics & Probability Letters, 69(1):1–9, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000884.

Hong:2001:FSB

[HL01] Wenming Hong and Zenghu Li. Fluctuations of a super-Brownian motion with randomly controlled immigration. Statistics & Probability Letters, 51(3):285–291, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001619.

Ha:2002:CCT

[HL02] Jeongcheol Ha and Sangyeol Lee. Coefficient constancy test in AR–ARCH models. Statistics & Probability Letters, 57(1):65– 77, March 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000512.

Husler:2006:WLS

[HLM06] Jürg Hüsler, Deyuan Li, and Samuel Müller. Weighted least squares estimation of the extreme value index. Statistics & Probability Letters, 76(9):920–930, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004104.

Hall:2006:NEP

[HM06] Andreia Hall and Orlando Moreira. A note on the extremes of a particular moving average count data model. Statistics & Probability Letters, 76(2):135–141, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002671.

Hazelton:2009:LBK

[HM09] Martin L. Hazelton and Jonathan C. Marshall. Linear boundary kernels for bivariate density estimation. Statistics & Probability Letters, 79(8):999–1003, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800566X.

Hu:2008:ESA

[HMR08] Zhishui Hu, Chunsheng Ma, and John Robinson. Empirical saddlepoint approximations of the Studentized mean under stratified random sampling. Statistics & Probability Letters, 78(4):396–401, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207002532.

Hofmann:2002:RPR

[HN02] Glenn Hofmann and H. N. Nagaraja. A random power record model. Statistics & Probability Letters, 56(4):345–353, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520100147X.

Hoppe:2008:FOB

[HN08] Fred M. Hoppe and Mikhail Nediak. Fréchet optimal bounds on the probability of a union with supplementary information. Statistics & Probability Letters, 78(3):311–319, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700243X.

Hu:2000:ARN

[HNRV00] Tien-Chung Hu, Eunwoo Nam, Andrew Rosalsky, and Andrei I. Volodin. An application of the Ryll–Nardzewski– Woyczyński theorem to a uniform weak law for tail series of weighted sums of random elements in Banach spaces. Statistics & Probability Letters, 48(4):369–374, July 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000183.

Hu:2008:SSD

[HNS08] Yaozhong Hu, David Nualart, and Xiaoming Song. A singular stochastic differential equation driven by fractional Brownian motion. Statistics & Probability Letters, 78(14):2075–2085, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000801.

Henriques:2008:LDE

[HO08] Carla Henriques and Paulo Eduardo Oliveira. Large deviations for the empirical mean of associated random variables. Statistics & Probability Letters, 78(6):594–598, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003008.

Haupt:2009:ANN

[HO09] Harry Haupt and Walter Oberhofer. On asymptotic normality in nonlinear regression. Statistics & Probability Letters, 79(6):848–849, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005142.

Hobson:2007:SPI

[Hob07] David Hobson. A short proof of an identity for a Brownian Bridge due to Donati–Martin, Matsumoto and Yor. Statistics & Probability Letters, 77(2):148–150, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002148.

Holzmann:2008:TPM

[Hol08] Hajo Holzmann. Testing parametric models in the presence of instrumental variables. Statistics & Probability Letters, 78(6):629–636, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003057.

Hooghiemstra:2002:EOT

[Hoo02] Gerard Hooghiemstra. On explicit occupation time distributions for Brownian processes. Statistics & Probability Letters, 56(4):405–417, February 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000378.

Horfelt:2003:PIM

[Hör03] Per Hörfelt. A probabilistic interpretation of the θ-method. Statistics & Probability Letters, 62(2):117–122, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202004352.

Hory:2004:FEA

[Hor04] C. Hory. Formulation of the EM algorithm for a mixture of central and non-central χ^{2} distributions. Statistics & Probability Letters, 68(3):267–272, July 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001038.

Hormann:2006:EAS

[Hör06] Siegfried Hörmann. An extension of almost sure central limit theory. Statistics & Probability Letters, 76(2):191–202, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002889.

Hou:2005:SAD

[Hou05] Chia-Ding Hou. A simple approximation for the distribution of the weighted combination of non-independent or independent probabilities. Statistics & Probability Letters, 73(2):179–187, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520500009X.

Hansen:2000:PRE

[HP00] Ben Hansen and Jim Pitman. Prediction rules for exchangeable sequences related to species sampling. Statistics & Probability Letters, 46(3):251–256, February 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001091.

Hitczenko:2005:CLT

[HP05] Paweł Hitczenko and Robin Pemantle. Central limit theorem for the size of the range of a renewal process. Statistics & Probability Letters, 72(3):249–264, May 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000350.

Harman:2007:IRN

[HP07] Radoslav Harman and Luc Pronzato. Improvements on removing nonoptimal support points in D-optimum design algorithms. Statistics & Probability Letters, 77(1):90–94, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001994.

Houdre:2008:IRB

[HP08] Christian Houdré and Nicolas Privault. Isoperimetric and related bounds on configuration spaces. Statistics & Probability Letters, 78(14):2154–2164, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000904.

Hwang:2009:AIN

[HP09] Kyo-Shin Hwang and Tian-Xiao Pang. Asymptotic inference for nearly nonstationary AR(1) processes with possibly infinite variance. Statistics & Probability Letters, 79(22):2374–2379, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003204.

Hwang:2002:LTB

[HPR02] J. H. Hwang, B. U. Park, and W. Ryu. Limit theorems for boundary function estimators. Statistics & Probability Letters, 59(4):353–360, October 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002110.

Halidias:2008:ETS

[HR08a] Nikolaos Halidias and Yong Ren. An existence theorem for stochastic functional differential equations with delays under weak assumptions. Statistics & Probability Letters, 78(17):2864–2867, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002216. See remarks and correc-

tions [Hal09].

Hassairi:2008:TDC

[HR08b] Abdelhamid Hassairi and Ons Regaieg. On the Tukey depth of a continuous probability distribution. Statistics & Probability Letters, 78(15):2308–2313, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001259.

Harris:2009:MCE

[HR09] S. C. Harris and M. I. Roberts. Measure changes with extinction. Statistics & Probability Letters, 79(8):1129–1133, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209000066.

Hu:2008:CPS

[HRV08] Tien-Chung Hu, Andrew Rosalsky, and Andrei Volodin. On convergence properties of sums of dependent random variables under second moment and covariance restrictions. Statistics & Probability Letters, 78(14):1999–2005, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000679.

Hirotsu:2000:SCI

[HS00] Chihiro Hirotsu and Muni S. Srivastava. Simultaneous confidence intervals based on one-sided max t test. Statistics & Probability Letters, 49(1):25–37, August 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000274.

Hall:2003:ESS

[HS03a] A. Hall and M. G. Scotto. Extremes of sub-sampled integervalued moving average models with heavy-tailed innovations. Statistics & Probability Letters, 63(1):97–105, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000579.

Hu:2003:LTN

[HS03b] Zhishui Hu and Chun Su. Limit theorems for the number and sum of near-maxima for medium tails. Statistics & Probability Letters, 63(3):229–237, July 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000853.

Hush:2005:CHD

[HS05] Don Hush and Clint Scovel. Concentration of the hypergeometric distribution. Statistics & Probability Letters, 75 (2):127–132, November 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002191.

Holan:2007:MLE

[HS07a] Scott Holan and Christine Spinka. Maximum likelihood estimation for joint mean-covariance models from unbalanced repeated-measures data. Statistics & Probability Letters, 77(3):319–328, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002409.

Hsiao:2007:SVS

[HS07b] Wei-Cheng Hsiao and Yu-Shan Shih. Splitting variable selection for multivariate regression trees. Statistics & Probability Letters, 77(3):265–271, February 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600232X.

Haldrup:2008:NVT

[HS08a] Niels Haldrup and Andreu Sansó. A note on the Vogelsang test for additive outliers. Statistics & Probability Letters, 78(3):296–300, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002416.

Ho:2008:MMP

[HS08b] Lai Ping Ho and D. Stoyan. Modelling marked point patterns by intensity-marked Cox processes. Statistics & Probability Letters, 78(10):1194–1199, August 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003926.

Hu:1998:NCC

[HSV98] T.-C. Hu, D. Szynal, and A. I. Volodin. A note on complete convergence for arrays. Statistics & Probability Letters, 38(1):27–31, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715298001503. See [HV00].

Hanson:2005:CCD

[HSX05] Timothy Hanson, Jayaram Sethuraman, and Ling Xu. On choosing the centering distribution in Dirichlet process mixture models. Statistics & Probability Letters, 72(2):153–162, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000258.

Hsu:2005:OED

[HT05] Yung-Feng Hsu and Chao-Ping Ting. A-optimal and efficient diallel cross experiments for comparing test treatments with a control. Statistics & Probability Letters, 71(1):99–110, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002883.

Hernandez:2006:CRL

[HU06a] Víctor Hernández and Henar Urmeneta. Convergence rates for the Law of Large Numbers for arrays. Statistics & Probability Letters, 76(16):1714–1722, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001258.

Hu:2006:URT

[Hu06b] Xiaomi Hu. Unbiasedness on rays of the tests of locations of homogeneous means in isotonic regression. Statistics & Probability Letters, 76(10):976–980, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004189.

Hu:2008:LIL

[Hu08] Xiaoyu Hu. The law of the iterated logarithm for the Gaussian free field. Statistics & Probability Letters, 78(17):3023–3028, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002629.

Hu:2009:VTH

[Hu09] Xiaomi Hu. p-Values of a test on homogeneous means in a multivariate isotonic regression. Statistics & Probability Letters, 79(19):2005–2011, October 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002338.

Huang:2001:RPV

[Hua01] Li-Shan Huang. A roughness-penalty view of kernel smoothing. Statistics & Probability Letters, 52(1):85–89, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200002273.

Huang:2002:BAS

[Hua02] Su-Yun Huang. On a Bayesian aspect for soft wavelet shrinkage estimation under an asymmetric Linex loss. Statistics & Probability Letters, 56(2):171–175, January 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100181X.

Huang:2003:APS

[Hua03a] Jianhua Z. Huang. Asymptotics for polynomial spline regression under weak conditions. Statistics & Probability Letters, 65(3):207–216, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002530.

Huang:2003:LIL

[Hua03b] Wei Huang. A law of the iterated logarithm for geometrically weighted series of negatively associated random variables. Statistics & Probability Letters, 63(2):133–143, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000609.

Huang:2009:DSA

[Hua09] Huilin Huang. The degree sequences of an asymmetrical growing network. Statistics & Probability Letters, 79(4): 420–425, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004380.

Hueter:2003:GDI

[Hue03] Irene Hueter. Geometric decay of infection probabilities for the anisotropic contact process. Statistics & Probability Letters, 63(4):367–374, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001020.

Huggins:2001:NDA

[Hug01] Richard Huggins. A note on the difficulties associated with the analysis of capture-recapture experiments with heterogeneous capture probabilities. Statistics & Probability Letters, 54(2):147–152, September 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002339.

Huggins:2004:NNE

[Hug04] Richard Huggins. A note on nonparametric estimation for clustered data. Statistics & Probability Letters, 69(2):129–133, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001658.

Huggins:2007:ULM

[Hug07] Richard Huggins. On the use of linear models in the estimation of the size of a population using capture-recapture data. Statistics & Probability Letters, 77(6):649–653, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002999.

Huo:2005:MCB

[Huo05] Xiaoming Huo. Minimax correlation between a line segment and a beamlet. Statistics & Probability Letters, 72(1):71–81, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000180.

Hurlimann:2003:HLC

[Hür03] Werner Hürlimann. Hutchinson–Lai’s conjecture for bivariate extreme value copulas. Statistics & Probability Letters, 61(2):191–198, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003498.

Hurlimann:2004:RCA

[Hür04] Werner Hürlimann. On the rate of convergence to asymptotic independence between order statistics. Statistics & Probability Letters, 66(3):355–362, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003638.

Huskova:2003:SRS

[Hus03] M. Husková. Serial rank statistics for detection of changes. Statistics & Probability Letters, 61(2):199–213, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202003528.

Hutson:2000:ECB

[Hut00] Alan D. Hutson. Estimating the covariance of bivariate order statistics with applications. Statistics & Probability Letters, 48(2):195–203, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299002059.

Hutson:2002:QMR

[Hut02] Alan D. Hutson. Quasi-medians are robust and relatively efficient estimators of a common mean given multivariate normality. Statistics & Probability Letters, 57(4):403–408, May 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001153.

Hutson:2003:NEN

[Hut03] Alan D. Hutson. Nonparametric estimation of normal ranges given one-way ANOVA random effects assumptions. Statistics & Probability Letters, 64(4):415–424, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002074.

Hu:2000:ANC

[HV00]  Tien-Chung Hu and Andrei Volodin.  Addendum to

“A note on complete convergence for arrays”, Statist. Probab. Lett. 38 (1) (1998) 27–31. Statistics & Probability Letters, 47(2):209–211, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299002096. See [HSV98].

Hamedani:2006:IPC

[HV06] G. G. Hamedani and H. Volkmer. An inverse problem for a class of continuous distributions with skewness. Statistics & Probability Letters, 76(9):945–951, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500413X.

Hjort:2001:NKD

[HW01a] Nils Lid Hjort and Stephen G. Walker. A note on kernel density estimators with optimal bandwidths. Statistics & Probability Letters, 54(2):153–159, September 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100027X.

Hou:2001:DRS

[HW01b] X. Shirley Hou and C. F. J. Wu. On the determination of robust settings in parameter design experiments. Statistics & Probability Letters, 54(2):137–145, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002170.

Hu:2001:MSC

[HW01c] Taizhong Hu and Ying Wei. Multivariate stochastic comparisons of inspection and repair policies. Statistics & Probability Letters, 53(3):315–324, June 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000888.

Hu:2001:SCS

[HW01d] Taizhong Hu and Ying Wei. Stochastic comparisons of spacings from restricted families of distributions. Statistics & Probability Letters, 53(1):91–99, May 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100044X.

Hwang:2001:TAP

[HW01e] S. Y. Hwang and Mi-Ja Woo. Threshold ARCH(1) processes: asymptotic inference. Statistics & Probability Letters, 53(1):11–20, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002224.

Hariz:2005:RCC

[HW05] Samir Ben Hariz and Jonathan J. Wylie. Rates of convergence for the change-point estimator for long-range dependent sequences. Statistics & Probability Letters, 73(2):155–164, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001057.

Heller:2007:PDN

[HW07] B. Heller and M. Wang. Posterior distribution for negative binomial parameter p using a group invariant prior. Statistics & Probability Letters, 77(15):1542–1548, September 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001277.

Hu:2008:ZFO

[HWL08] Xuemei Hu, Zhizhong Wang, and Feng Liu. Zero finiteorder serial correlation test in a semi-parametric varyingcoefficient partially linear errors-in-variables model. Statistics & Probability Letters, 78(12):1560–1569, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000126.

Hu:2009:CPB

[HWLZ09] Shuhe Hu, Xuejun Wang, Xiaoqin Li, and Yuanyuan Zhang. Comments on the paper: A bilateral inequality on the Borel– Cantelli lemma. Statistics & Probability Letters, 79(7):889– 893, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005312.

Hu:2009:APC

[HWYZ09a] Shuhe Hu, Xuejun Wang, Wenzhi Yang, and Ting Zhao. Acknowledgement of priority concerning “The Hájek–Rènyitype inequality for associated random variables” [Statist. Probab. Lett. 79 (2009) 884–888]. Statistics & Probability Letters, 79(15):1732, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001606. See [HWYZ09b].

Hu:2009:HRT

[HWYZ09b] Shuhe Hu, Xuejun Wang, Wenzhi Yang, and Ting Zhao. The Hájek–Rényi-type inequality for associated random variables. Statistics & Probability Letters, 79(7):884–888, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005300. See acknowl-

edgement of priority [HWYZ09a].

He:2009:TDN

[HWZ09a] Jingmin He, Rong Wu, and Huayue Zhang. Total duration of negative surplus for the risk model with debit interest. Statistics & Probability Letters, 79(10):1320–1326, May 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209000583.

Hu:2009:ELS

[HWZ09b] Xuemei Hu, Zhizhong Wang, and Zhiwei Zhao. Empirical likelihood for semiparametric varying-coefficient partially linear errors-in-variables models. Statistics & Probability Letters, 79(8):1044–1052, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005749.

Huang:2002:SMI

[HX02] Wen-Tao Huang and Bing Xu. Some maximal inequalities and complete convergences of negatively associated random sequences. Statistics & Probability Letters, 57(2):183–191, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000494.

He:2000:SCP

[HY00] Shuyuan He and Grace L. Yang. On the strong convergence of the product-limit estimator and its integrals under censoring and random truncation. Statistics & Probability Letters, 49(3):235–244, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000523.

Hwang:2001:IEP

[HY01] Leng-Cheng Hwang and Chia-Chen Yang. An interval estimation procedure with deterministic stopping rule in Bayes sequential interval estimation. Statistics & Probability Letters, 52(3):243–248, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001966.

Hariya:2004:EDR

[HY04a] Yuu Hariya and Marc Yor. On an extension of Dufresne’s relation between exponential Brownian functionals from opposite drifts to two different drifts: a short proof. Statistics & Probability Letters, 67(4):331–341, May 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400063X.

Hu:2004:FDS

[HY04b] Taizhong Hu and Jianping Yang. Further developments on sufficient conditions for negative dependence of random variables. Statistics & Probability Letters, 66(3):369–381, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203003663.

Huggins:2004:NEI

[HYL04] Richard M. Huggins, Paul S. F. Yip, and Eric H. Y. Lau. A note on the estimation of the initial number of susceptible individuals in the general epidemic model. Statistics & Probability Letters, 67(4):321–330, May 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000616.

Hyndman:2007:FBS

[Hyn07] Cody Blaine Hyndman. Forward-backward SDEs and the CIR model. Statistics & Probability Letters, 77(17):1676–1682, November 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207001514.

Horvath:2003:AND

[HZ03] Lajos Horváth and Ricardas Zitikis. Asymptotics of the L_{p}- norms of density estimators in the first-order autoregressive models. Statistics & Probability Letters, 65(4):331–342, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002633.

Herrmann:2004:RCN

[HZ04a]  Eva Herrmann and Klaus Ziegler.  Rates of consis-

tency for nonparametric estimation of the mode in absence of smoothness assumptions. Statistics & Probability Letters, 68(4):359–368, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001245.

Horvath:2004:AND

[HZ04b] Lajos Horváth and Ricardas Zitikis. Asymptotics of the L_{p}- norms of density estimators in the first-order autoregressive models. Statistics & Probability Letters, 66(2):91–103, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002645.

Hu:2005:NSC

[HZ05] Taizhong Hu and Weiwei Zhuang. A note on stochastic comparisons of generalized order statistics. Statistics & Probability Letters, 72(2):163–170, April 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500026X.

Huang:2006:ANS

[HZ06] Wei Huang and Lin-Xi Zhang. Asymptotic normality for Ustatistics of negatively associated random variables. Statistics & Probability Letters, 76(11):1125–1131, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004566.

Husler:2008:FLR

[HZ08] Jürg Hüsler and Yueming Zhang. On first and last ruin times of Gaussian processes. Statistics & Probability Letters, 78(10):1230–1235, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003987.

Hu:2009:MRM

[HZ09a] Jianwei Hu and Runchu Zhang. Maximal rank minimum aberration and doubling. Statistics & Probability Letters, 79(7):915–919, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005452.

Huang:2009:EEA

[HZ09b] Zhensheng Huang and Riquan Zhang. Efficient estimation of adaptive varying-coefficient partially linear regression model. Statistics & Probability Letters, 79(7):943–952, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005506.

Huang:2009:ELN

[HZ09c] Zhensheng Huang and Riquan Zhang. Empirical likelihood for nonparametric parts in semiparametric varyingcoefficient partially linear models. Statistics & Probability Letters, 79(16):1798–1808, August 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001837.

Zhu:2008:UAT

[hZbG08] Chun hua Zhu and Qi bing Gao. The uniform approximation of the tail probability of the randomly weighted sums of subexponential random variables. Statistics & Probability Letters, 78(15):2552–2558, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001703.

Iacus:2001:SAI

[Iac01] Stefano Maria Iacus. Statistical analysis of the inhomogeneous telegrapher’s process. Statistics & Probability Letters, 55(1):83–88, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520100133X.

Iliopoulos:2009:CIB

[IB09] G. Iliopoulos and N. Balakrishnan. Conditional independence of blocked ordered data. Statistics & Probability Letters, 79(8):1008–1015, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005683.

Ibragimov:2008:HTT

[Ibr08] Rustam Ibragimov. Heavy-tailedness and threshold sex determination. Statistics & Probability Letters, 78(16):2804–2810, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002058.

Isogai:2008:CRS

[IF08] Eiichi Isogai and Andreas Futschik. On the convergence rate of sequential fixed-width confidence intervals for normal parameters. Statistics & Probability Letters, 78(13):1826–1834, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000473.

Ignatyev:2009:PAS

[Ign09] Oleksiy Ignatyev. Partial asymptotic stability in probability of stochastic differential equations. Statistics & Probability Letters, 79(5):597–601, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004719.

Ibragimov:2008:OCR

[II08] Marat Ibragimov and Rustam Ibragimov. Optimal constants in the Rosenthal inequality for random variables with zero odd moments. Statistics & Probability Letters, 78(2):186–189, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700212X.

Iksanov:2004:PYP

[IK04] Aleksander M. Iksanov and Che-Soong Kim. On a Pitman– Yor problem. Statistics & Probability Letters, 68(1):61–72, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000690.

Ishioka:2009:SAI

[IKAT09] Takahide Ishioka, Shunsuke Kawamura, Tomoyuki Amano, and Masanobu Taniguchi. Spectral analysis for intrinsic time processes. Statistics & Probability Letters, 79(23):2389–2396, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209003216.

Iwashita:2009:AED

[IKIS09] Toshiya Iwashita, Yoshihide Kakizawa, Tatsuki Inoue, and Takashi Seo. An asymptotic expansion of the distribution of Student’s t type statistic under spherical distributions. Statistics & Probability Letters, 79(18):1935–1942, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520900203X.

Inui:2005:VSS

[IKK05] Koji Inui, Masaaki Kijima, and Atsushi Kitano. VaR is subject to a significant positive bias. Statistics & Probability Letters, 72(4):299–311, May 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000635.

Inoue:2008:BIF

[IKP08] Akihiko Inoue, Yukio Kasahara, and Punam Phartyal. Baxter’s inequality for fractional Brownian motion-type processes with Hurst index less than 1/2. Statistics & Probability Letters, 78(17):2889–2894, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002320.

Iliopoulos:2000:NDT

[Ili00] George Iliopoulos. A note on decision theoretic estimation of ordered parameters. Statistics & Probability Letters, 50(1):33–38, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520000078X.

Iltis:2000:SAL

[Ilt00] Michael Iltis. Sharp asymptotics of large deviations for general state-space Markov-additive chains in R^{d}. Statistics & Probability Letters, 47(4):365–380, May 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001819.

Itoh:2005:ASM

[IM05] Yoshiaki Itoh and Hosam M. Mahmoud. Age statistics in the Moran population model. Statistics & Probability Letters, 74(1):21–30, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001537.

Iksanov:2008:NNR

[IM08] Alex Iksanov and Alex Marynych. A note on non-regular martingales. Statistics & Probability Letters, 78(17):3014–3017, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002605.

Imai:2008:SBC

[Ima08] Kosuke Imai. Sharp bounds on the causal effects in randomized experiments with “truncation-by-death”. Statistics & Probability Letters, 78(2):144–149, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002052.

Itoh:2006:PAM

[IMS06] Yoshiaki Itoh, Hosam Mahmoud, and Robert Smythe. Probabilistic analysis of maximal gap and total accumulated length in interval division. Statistics & Probability Letters, 76(13):1356–1363, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000423.

Inoue:2007:BMM

[INA07] Akihiko Inoue, Yumiharu Nakano, and Vo Anh. Binary market models with memory. Statistics & Probability Letters, 77(3):256–264, February 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002318.

Inglot:2000:LDT

[Ing00] Tadeusz Inglot. On large deviation theorem for data-driven Neyman’s statistic. Statistics & Probability Letters, 47(4):411– 419, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771529900187X.

Ioannides:2004:FDR

[Ioa04] D. A. Ioannides. Fixed design regression quantiles for time series. Statistics & Probability Letters, 68(3):235–245, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003882.

Iglesias-Perez:2003:SRC

[IP03] M. C. Iglesias-Pérez. Strong representation of a conditional quantile function estimator with truncated and censored data. Statistics & Probability Letters, 65(2):79–91, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002207.

Ibarrola:2004:LCC

[IPP04] P. Ibarrola and A. Pérez-Palomares. Linear completeness in a continuous time Gauss–Markov model. Statistics & Probability Letters, 69(2):143–149, August 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001695.

Ishwaran:2008:CGE

[IR08] Hemant Ishwaran and J. Sunil Rao. Clustering gene expression profile data by selective shrinkage. Statistics & Probability Letters, 78(12):1490–1497, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000023.

Irle:2004:MTA

[Irl04] A. Irle. A measure-theoretic approach to completeness of financial markets. Statistics & Probability Letters, 68(1):1–7, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520400029X.

Ibragimov:2001:BCR

[IS01a] R. Ibragimov and Sh. Sharakhmetov. The best constant in the Rosenthal inequality for nonnegative random variables. Statistics & Probability Letters, 55(4):367–376, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001341.

Sato:2001:SSD

[iS01b] Ken iti Sato. Subordination and self-decomposability. Statistics & Probability Letters, 54(3):317–324, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001109.

Ivanoff:2003:LTP

[IS03] B. Gail Ivanoff and P. Sawyer. Local time for processes indexed by a partially ordered set. Statistics & Probability Letters, 61(1):1–15, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200161X.

Istas:2006:KLE

[Ist06]  Jacques Istas.  Karhunen–Loève expansion of spheri-

cal fractional Brownian motions. Statistics & Probability Letters, 76(14):1578–1583, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001015.

Kelly:2000:APP

[IV00] Harrison W. Kelly III and Joseph O. Voelkel. Asymptoticpower problems in the analysis of supersaturated designs. Statistics & Probability Letters, 47(4):317–324, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001728.

Ivanoff:2004:PSP

[IW04] B. Gail Ivanoff and N. C. Weber. Predictable sampling for partially exchangeable arrays. Statistics & Probability Letters, 70(1):95–108, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002317.

Ip:2004:GRA

[IWWJ04] Wai Cheung Ip, Heung Wong, Song-Gui Wang, and Zhong-Zhen Jia. A GIC rule for assessing data transformation in regression. Statistics & Probability Letters, 68(1):105–110, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000744.

Jacroux:2008:NCO

[Jac08] Mike Jacroux. A note on the construction of optimal main effects plans in blocks of size two. Statistics & Probability Letters, 78(15):2366–2370, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001351.

James:2003:SPA

[Jam03] Lancelot F. James. A simple proof of the almost sure discreteness of a class of random measures. Statistics & Probability Letters, 65(4):363–368, December 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002839.

Janssen:2003:NCR

[Jan03] Arnold Janssen. A nonparametric Cramér-Rao inequality for estimators of statistical functionals. Statistics & Probability Letters, 64(4):347–358, October 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001792.

Janosky:2008:STA

[Jan08] Janine E. Janosky. Statistical testing alone and estimation plus testing: Reporting study outcomes in biomedical journals. Statistics & Probability Letters, 78(15):2327–2331, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001296.

Jaruskova:2003:ADS

[Jar03] Daniela Jarusková. Asymptotic distribution of a statistic testing a change in simple linear regression with equidistant design. Statistics & Probability Letters, 64(1):89–95, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001433.

Jaroszewska:2008:HCA

[Jar08] Joanna Jaroszewska. How to construct asymptotically stable iterated function systems. Statistics & Probability Letters, 78(12):1570–1576, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000138.

Jernigan:2003:TLM

[JB03] Robert W. Jernigan and Robert H. Baran. Testing lumpability in Markov chains. Statistics & Probability Letters, 64(1):17–23, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001263.

Johnson:2005:NUK

[JB05] Brent A. Johnson and Dennis D. Boos. A note on the use of kernel functions in weighted estimators. Statistics & Probability Letters, 72(4):345–355, May 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000775.

Jamalizadeh:2009:PTN

[JB09]  A. Jamalizadeh and N. Balakrishnan.  Prediction in

a trivariate normal distribution via a linear combination of order statistics. Statistics & Probability Letters, 79(21):2289–2296, November 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002910.

Jamalizadeh:2008:TPG

[JBB08] A. Jamalizadeh, J. Behboodian, and N. Balakrishnan. A two-parameter generalized skew-normal distribution. Statistics & Probability Letters, 78(13):1722–1726, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800031X.

Jasra:2008:SSM

[JD08a] Ajay Jasra and Arnaud Doucet. Stability of sequential Monte Carlo samplers via the Foster–Lyapunov condition. Statistics & Probability Letters, 78(17):3062–3069, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800268X.

Johansen:2008:NAP

[JD08b] Adam M. Johansen and Arnaud Doucet. A note on auxiliary particle filters. Statistics & Probability Letters, 78(12):1498–1504, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000035.

Jun:2008:CCW

[JD08c] An Jun and Yuan Demei. Complete convergence of weighted sums for ρ∗-mixing sequence of random variables. Statistics & Probability Letters, 78(12):1466–1472, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004312.

Jensen:2001:PSS

[Jen01] D. R. Jensen. Properties of selected subset diagnostics in regression. Statistics & Probability Letters, 51(4):377– 388, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001784.

Jensen:2004:CNE

[Jen04] D. R. Jensen. Condition numbers and D-efficiency. Statistics & Probability Letters, 66(3):267–274, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003389.

Johnson:2000:DIS

[JF00] Brad C. Johnson and James C. Fu. The distribution of increasing l-sequences in random permutations: a Markov chain approach. Statistics & Probability Letters, 49(4):337–344, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000663.

Jin:2005:VSG

[JFZ05] Man Jin, Yixin Fang, and Lincheng Zhao. Variable selection in generalized linear models with canonical link functions. Statistics & Probability Letters, 71(4):371–382, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000118.

Jammalamadaka:2004:TGF

[JG04] S. Rao Jammalamadaka and M. N. Goria. A test of goodness-of-fit based on Gini’s index of spacings. Statistics & Probability Letters, 68(2):177–187, June 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000793.

Jimenez-Gamero:2006:CRB

[JGMGdlV06] M. D. Jiménez-Gamero, J. Muñoz-García, and M. D. Cubiles de-la Vega. Consistency of the reduced bootstrap for sample means. Statistics & Probability Letters, 76(7):689–697, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003767.

Jimenez-Gamero:2003:BPE

[JGMGPM03] M. D. Jiménez-Gamero, J. Muñoz-García, and R. Pino-Mejías. Bootstrapping parameter estimated degenerate U and V statistics. Statistics & Probability Letters, 61(1):61–70, January 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003383.

Jiang:2002:ECL

[JH02] Xinxin Jiang and Marjorie G. Hahn. Empirical Central Limit Theorems for exchangeable random variables. Statistics & Probability Letters, 59(1):75–81, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002043.

Jiang:2008:SNC

[JH08] Xinxin Jiang and Marjorie Hahn. A self-normalized Central Limit Theorem for ρ-mixing stationary sequences. Statistics & Probability Letters, 78(12):1541–1547, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000096.

Jiang:2001:MEM

[Jia01a] Jiming Jiang. Mixed-effects models with random cluster sizes. Statistics & Probability Letters, 53(2):201–206, June 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000803.

Jiang:2001:NTA

[Jia01b] Jiming Jiang. A nonstandard χ^{2}-test with application to generalized linear model diagnostics. Statistics & Probability Letters, 53(1):101–109, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000499.

Jiang:2005:CCT

[Jia05] Long Jiang. Converse comparison theorems for backward stochastic differential equations. Statistics & Probability Letters, 71(2):173–183, February 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002925.

Jiang:2006:NEC

[Jia06] Long Jiang. A note on g-expectation with comonotonic additivity. Statistics & Probability Letters, 76(17):1895– 1903, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206001568.

Jia:2008:CBS

[Jia08] Guangyan Jia. A class of backward stochastic differential equations with discontinuous coefficients. Statistics & Probability Letters, 78(3):231–237, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002180.

Jia:2009:SUR

[Jia09a] Guangyan Jia. Some uniqueness results for one-dimensional BSDEs with uniformly continuous coefficients. Statistics & Probability Letters, 79(4):436–441, February 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004409.

Jiang:2009:NSC

[Jia09b] Long Jiang. A necessary and sufficient condition for probability measures dominated by g-expectation. Statistics & Probability Letters, 79(2):196–201, January 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003684.

Jahufer:2009:AGI

[JJ09] Aboobacker Jahufer and Chen Jianbao. Assessing global influential observations in modified ridge regression. Statistics & Probability Letters, 79(4):513–518, February 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004598.

James:2008:LTC

[JJQ08] Barry James, Kang James, and Yongcheng Qi. Limit theorems for correlated Bernoulli random variables. Statistics & Probability Letters, 78(15):2339–2345, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001314.

Jayakumar:2007:TSM

[JK07] K. Jayakumar and A. P. Kuttykrishnan. A time-series model using asymmetric Laplace distribution. Statistics & Probability Letters, 77(16):1636–1640, October 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003901.

Jedynak:2008:FNH

[JK08] Bruno Jedynak and Damianos Karakos. Finding a needle in a haystack: Conditions for reliable detection in the presence of clutter. Statistics & Probability Letters, 78(5):471–480, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002635.

Joo:2005:CTL

[JKK05] Sang Yeol Joo, Yun Kyung Kim, and Joong Sung Kwon. On Chung’s type Law of Large Numbers for fuzzy random variables. Statistics & Probability Letters, 74(1):67–75, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001604.

Jureckova:2001:TBL

[JKP01] Jana Jurecková, Roger Koenker, and Stephen Portnoy. Tail behavior of the least-squares estimator. Statistics & Probability Letters, 55(4):377–384, December 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001377.

Jackson:2009:RFH

[JKZ09] Ken Jackson, Alexander Kreinin, and Wanhe Zhang. Randomization in the first hitting time problem. Statistics & Probability Letters, 79(23):2422–2428, December 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003253.

Johnson:2007:PLD

[JL07] Richard A. Johnson and Wenqing Lu. Proof load designs for estimation of dependence in a bivariate Weibull model. Statistics & Probability Letters, 77(11):1061–1069, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000326.

Jin:2009:OLC

[JL09a] Hua Jin and Ying Lu. The optimal linear combination of multiple predictors under the generalized linear models. Statistics & Probability Letters, 79(22):2321–2327, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003010.

Jin:2009:PTN

[JL09b] Hua Jin and Ying Lu. Permutation test for non-inferiority of the linear to the optimal combination of multiple tests. Statistics & Probability Letters, 79(5):664–669, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004951.

Jin:2009:RRR

[JL09c] Hua Jin and Ying Lu. The ROC region of a regression tree. Statistics & Probability Letters, 79(7):936–942, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005488.

Jasiulis:2008:CBW

[JM08] B. H. Jasiulis and J. K. Misiewicz. On the connections between weakly stable and pseudo-isotropic distributions. Statistics & Probability Letters, 78(16):2751–2755, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800196X.

Jozani:2006:EWB

[JMP06] Mohammad Jafari Jozani, Éric Marchand, and Ahmad Parsian. On estimation with weighted balanced-type loss function. Statistics & Probability Letters, 76(8):773–780, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003858.

Jiang:2009:SAF

[JNR09] Jiming Jiang, Thuan Nguyen, and J. Sunil Rao. A simplified adaptive fence procedure. Statistics & Probability Letters, 79(5):625–629, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004896.

Jozani:2002:AME

[JNS02] Mohammad Jafari Jozani, Nader Nematollahi, and Khalil Shafie. An admissible minimax estimator of a bounded scale-parameter in a subclass of the exponential family under scale-invariant squared-error loss. Statistics & Probability Letters, 60(4):437–444, December 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003164.

Joarder:2008:MFB

[JO08] Anwar H. Joarder and M. H. Omar. A mass function based on correlation coefficient and its application. Statistics & Probability Letters, 78(18):3344–3349, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003258.

Johnson:2002:DIS

[Joh02] Brad C. Johnson. The distribution of increasing 2-sequences in random permutations of arbitrary multi-sets. Statistics & Probability Letters, 59(1):67–74, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001955.

Johnson:2007:MYF

[Joh07] Richard A. Johnson. My 25 years as Founding Editor of Statistics & Probability Letters. Statistics & Probability Letters, 77(12):vii–viii, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520700106X.

Jongbloed:2000:MLB

[Jon00] Geurt Jongbloed. Minimax lower bounds and moduli of continuity. Statistics & Probability Letters, 50(3):279–284, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001048.

Jonasson:2002:UUR

[Jon02a] Johan Jonasson. Uniqueness of uniform random colorings of regular trees. Statistics & Probability Letters, 57(3):243–248, April 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000548.

Jones:2002:DBD

[Jon02b] M. C. Jones. A dependent bivariate t distribution with marginals on different degrees of freedom. Statistics & Probability Letters, 56(2):163–170, January 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001808.

Jones:2002:FUO

[Jon02c] M. C. Jones. On fractional uniform order statistics. Statistics & Probability Letters, 58(1):93–96, May 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001190.

Jonasson:2008:DCC

[Jon08] Johan Jonasson. Dynamical circle covering with homogeneous Poisson updating. Statistics & Probability Letters, 78(15):2400–2403, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208001399.

Joutard:2007:ALD

[Jou07] Cyrille Joutard. Applications of large deviations to optimal experimental designs. Statistics & Probability Letters, 77(3):231–238, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002288.

Jowaheer:2006:MME

[Jow06] Vandna Jowaheer. Model misspecification effects in clustered count data analysis. Statistics & Probability Letters, 76(5):470–478, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205003275.

Jiao:2009:WTI

[JPL09] Yong Jiao, Lihua Peng, and Peide Liu. Weak type inequalities on Lorentz martingale spaces. Statistics & Probability Letters, 79(14):1581–1584, July 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001230.

Juodis:2007:CLT

[JR07] Mindaugas Juodis and Alfredas Rackauskas. A Central Limit Theorem for self-normalized sums of a linear process. Statistics & Probability Letters, 77(15):1535–1541, September 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001265.

Jokiel-Rokita:2008:SEP

[JR08] Alicja Jokiel-Rokita. A sequential estimation procedure for the parameter of an exponential distribution under asymmetric loss function. Statistics & Probability Letters, 78(17):3091–3095, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002733.

Jensen:2009:CRR

[JR09] D. R. Jensen and D. E. Ramirez. Concentration reversals in ridge regression. Statistics & Probability Letters, 79(21):2237–2241, November 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002764.

Johnson:2002:TIS

[JS02] Richard A. Johnson and Grace S. Shieh. On tests of independence for spherical data-invariance and centering. Statistics & Probability Letters, 57(4):327–335, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000834.

Jammalamadaka:2007:IED

[JS07] S. Rao Jammalamadaka and D. Sengupta. Inclusion and exclusion of data or parameters in the general linear model. Statistics & Probability Letters, 77(12):1235–1247, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000879.

Jowaheer:2009:GVG

[JS09] Vandna Jowaheer and Brajendra Sutradhar. GMM versus GQL inferences for panel count data. Statistics & Probability Letters, 79(18):1928–1934, September 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002028.

Janssen:2001:MBC

[JSV01] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Modified bootstrap consistency rates for U-quantiles. Statistics & Probability Letters, 54(3):261–268, October 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000554.

Janssen:2002:MBE

[JSV02] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. The modified bootstrap error process for Kaplan–Meier quantiles. Statistics & Probability Letters, 58(1):31–39, May 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001001.

Janssen:2005:BMG

[JSV05] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Bootstrapping modified goodness-of-fit statistics with estimated parameters. Statistics & Probability Letters, 71(2):111– 121, February 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204002561.

Johnson:2008:PSL

[JT08] Richard A. Johnson and James R. Taylor. Preservation of some life length classes for age distributions associated with age-dependent branching processes. Statistics & Probability Letters, 78(17):2981–2987, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002538.

Jin:2009:BTS

[JTQ09] Hao Jin, Zheng Tian, and Ruibing Qin. Bootstrap tests for structural change with infinite variance observations. Statistics & Probability Letters, 79(19):1985–1995, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900217X.

Jose:2008:APN

[JTS08] K. K. Jose, Lishamol Tomy, and J. Sreekumar. Autoregressive processes with normal-Laplace marginals. Statistics & Probability Letters, 78(15):2456–2462, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001491.

Jureckova:2000:TTB

[Jur00a] Jana Jurecková. Test of tails based on extreme regression quantiles. Statistics & Probability Letters, 49(1):53–61, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000316.

Jurek:2000:NGR

[Jur00b] Zbigniew J. Jurek. A note on gamma random variables and Dirichlet series. Statistics & Probability Letters, 49 (4):387–392, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000717.

Jurek:2003:GLS

[Jur03] Zbigniew J. Jurek. Generalized Lévy stochastic areas and selfdecomposability. Statistics & Probability Letters, 64(2): 213–222, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001536.

Jurek:2007:RIR

[Jur07] Zbigniew J. Jurek. Random integral representations for freeinfinitely divisible and tempered stable distributions. Statistics & Probability Letters, 77(4):417–425, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002604.

Jimenez:2002:SLE

[JY02] Raúl Jiménez and J. E. Yukich. Strong laws for Euclidean graphs with general edge weights. Statistics & Probability Letters, 56(3):251–259, February 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001754.

Jasra:2008:RPC

[JY08] Ajay Jasra and Chao Yang. A regeneration proof of the Central Limit Theorem for uniformly ergodic Markov chains. Statistics & Probability Letters, 78(12):1649–1655, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000230.

Jing:2008:ELN

[JYZ08] Bing-Yi Jing, Junqing Yuan, and Wang Zhou. Empirical likelihood for non-degenerate U-statistics. Statistics & Probability Letters, 78(6):599–607, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700301X.

Jones:2004:ANM

[JZ04] C. Matthew Jones and Anatoly A. Zhigljavsky. Approximating the negative moments of the Poisson distribution.

Statistics & Probability Letters, 66(2):171–181, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003328.

Kateri:2007:COQ

[KA07] Maria Kateri and Alan Agresti. A class of ordinal quasisymmetry models for square contingency tables. Statistics & Probability Letters, 77(6):598–603, March 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002914.

Kabaila:2001:BBC

[Kab01] Paul Kabaila. Better Buehler confidence limits. Statistics & Probability Letters, 52(2):145–154, April 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001991.

Kabaila:2008:SPE

[Kab08] Paul Kabaila. Statistical properties of exact confidence intervals from discrete data using Studentized test statistics. Statistics & Probability Letters, 78(6):720–727, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003239.

Kabluchko:2009:LDC

[Kab09] Zakhar Kabluchko. Limiting distribution of the continuity modulus for Gaussian processes with stationary increments. Statistics & Probability Letters, 79(7):953–956, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005518.

Kagan:2002:TTE

[Kag02] Abram Kagan. Transfer theorems in exponential families. Statistics & Probability Letters, 57(3):225–229, April 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000470.

Kakizawa:2008:MCS

[Kak08] Yoshihide Kakizawa. Multiple comparisons of several heteroscedastic multivariate populations. Statistics & Probability Letters, 78(11):1328–1338, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004142.

Kantorovitz:2007:EST

[Kan07] Miriam Ruth Kantorovitz. An example of a stationary, triplewise independent triangular array for which the CLT fails. Statistics & Probability Letters, 77(5):539–542, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002859.

Karlis:2002:ETA

[Kar02] Dimitris Karlis. An EM type algorithm for maximum likelihood estimation of the normal-inverse Gaussian distribution. Statistics & Probability Letters, 57(1):43–52, March 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202000408.

Koul:2007:P

[KAS07] Hira L. Koul, Michael Akritas, and Anton Schick. Preface. Statistics & Probability Letters, 77(12):v–vi, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001058.

Kattumannil:2009:SIA

[Kat09] Sudheesh Kumar Kattumannil. On Stein’s identity and its applications. Statistics & Probability Letters, 79(12):1444–1449, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001060.

Kavalieris:2008:UCA

[Kav08] Laimonis Kavalieris. Uniform convergence of autocovariances. Statistics & Probability Letters, 78(6):830–838, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003392.

Kim:2001:CLT

[KB01] Tae-Sung Kim and Jong-Il Baek. A Central Limit Theorem for stationary linear processes generated by linearly positively quadrant-dependent process. Statistics & Probability Letters, 51(3):299–305, February 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001681.

Khmaladze:2007:BPR

[KBH07] Estate Khmaladze, Ray Brownrigg, and John Haywood. Brittle power: On Roman Emperors and exponential lengths of rule. Statistics & Probability Letters, 77(12):1248–1257, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000880.

Kepner:2003:MTS

[KC03] James L. Kepner and Myron N. Chang. On the maximum total sample size of a group sequential test about binomial proportions. Statistics & Probability Letters, 62(1):87–92, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202004364.

Kounias:2008:EPL

[KC08] Stratis Kounias and Miltiadis Chalikias. Estimability of parameters in a linear model and related characterizations. Statistics & Probability Letters, 78(15):2437–2439, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001466.

Kirmani:2003:TRR

[KD03] Syed N. U. A. Kirmani and Jean-Yves Dauxois. Testing relative risk under random censoring. Statistics & Probability Letters, 62(1):1–7, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003553.

Kirmani:2004:TKG

[KD04] Syed N. U. A. Kirmani and Jean-Yves Dauxois. Testing the Koziol–Green model against monotone conditional odds for censoring. Statistics & Probability Letters, 66(3):327– 334, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203003596.

Ker:2005:EBN

[KE05] Alan P. Ker and A. T. Ergün. Empirical Bayes nonparametric kernel density estimation. Statistics & Probability Letters, 75(4):315–324, December 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002464.

Kella:2006:RT

[Kel06] Offer Kella. Reflecting thoughts. Statistics & Probability Letters, 76(16):1808–1811, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001465.

Kernane:2008:CSI

[Ker08] Tewfik Kernane. Conditions for stability and instability of retrial queueing systems with general retrial times. Statistics & Probability Letters, 78(18):3244–3248, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003052.

Kevei:2007:GPP

[Kev07] Péter Kevei. Generalized n-Paul paradox. Statistics & Probability Letters, 77(11):1043–1049, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001149.

Kliger:2008:SCF

[KF08] Mark Kliger and Joseph M. Francos. Strong consistency of a family of model order selection rules for estimating 2D sinusoids in noise. Statistics & Probability Letters, 78(17):3075–3081, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002708.

Konis:2009:SDR

[KF09] Kjell Konis and Konstantinos Fokianos. Safe density ratio modeling. Statistics & Probability Letters, 79(18):1915– 1920, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002004.

Kotz:2002:MST

[KFW02] Samuel Kotz, Hong-Bin Fang, and Gang Wei. Multiple square tray distributions. Statistics & Probability Letters, 59(2):195–207, September 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202002031.

Kosorok:2006:RTI

[KG06] Michael R. Kosorok and Ronald E. Gangnon. Resolving the tail instability in weighted log-rank statistics for clustered survival data. Statistics & Probability Letters, 76(3):304–309, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003007.

Kabaila:2009:LSC

[KG09] Paul Kabaila and Khageswor Giri. Large-sample confidence intervals for the treatment difference in a two-period crossover trial, utilizing prior information. Statistics & Probability Letters, 79(5):652–658, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004938.

Kyung:2009:CVI

[KGC09] Minjung Kyung, Jeff Gill, and George Casella. Characterizing the variance improvement in linear Dirichlet random effects models. Statistics & Probability Letters, 79(22):2343–2350, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003058.

Kim:2001:APL

[KH01] Jeankyung Kim and Jinsoo Hwang. Asymptotic properties of location estimators based on projection depth. Statistics & Probability Letters, 55(3):293–299, December 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001523.

Kocherginsky:2007:EMC

[KH07] Masha Kocherginsky and Xuming He. Extensions of the Markov chain marginal bootstrap. Statistics & Probability Letters, 77(12):1258–1268, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000892.

Kouame:2008:MDE

[KH08a] Euloge F. Kouamé and Ouagnina Hili. Minimum distance estimation of k-factors GARMA processes. Statistics & Probability Letters, 78(18):3254–3261, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003106.

Kulik:2008:TSL

[KH08b] Rafal Kulik and Mohamedou Ould Haye. Trimmed sums of long range dependent moving averages. Statistics & Probability Letters, 78(15):2536–2542, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001673.

Khrennikov:2001:LBS

[Khr01] Andrei Khrennikov. Limit behaviour of sums of independent random variables with respect to the uniform p-adic distribution. Statistics & Probability Letters, 51(3):269–276, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001589.

Kim:2000:RCD

[Kim00a] Jeankyung Kim. Rate of convergence of depth contours: with application to a multivariate metrically trimmed mean. Statistics & Probability Letters, 49(4):393–400, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000730.

Kim:2000:IPT

[Kim00b] Seong W. Kim. Intrinsic priors for testing exponential means. Statistics & Probability Letters, 46(2):195–201, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001066.

Kim:2001:GAS

[Kim01]  Yoon Tae Kim.  A geometric approach to singular-

ity for Hilbert space-valued SDEs. Statistics & Probability Letters, 52(1):35–45, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002042.

Kim:2003:APS

[Kim03] Ji-Hyun Kim. Assessing practical significance of the proportional odds assumption. Statistics & Probability Letters, 65(3):233–239, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002591.

Kim:2008:NSM

[Kim08] Hyoung-Moon Kim. A note on scale mixtures of skew normal distribution. Statistics & Probability Letters, 78(13):

1694–1701, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000278.

See corrigendum [Kim13].

Kim:2013:CNS

[Kim13] Hyoung-Moon Kim. Corrigendum to “A note on scale mixtures of skew normal distribution” [Statist. Probab. Lett. 78 (2008) 1694–1701]. Statistics & Probability Letters, 83 (8):1937, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001491.

See [Kim08].

Kinateder:2001:UAB

[Kin01] Kimberly K. J. Kinateder. Understanding average Brownian exit time. Statistics & Probability Letters, 51(1):1–8, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001218.

Kinnison:2008:MSH

[Kin08] Adam L. Kinnison. The multifractal spectrum of harmonic measure for forward moving random walks on a Galton–Watson tree. Statistics & Probability Letters, 78(17):3114–3121, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002794.

Krishnarani:2008:CAP

[KJ08] S. D. Krishnarani and K. Jayakumar. A class of autoregressive processes. Statistics & Probability Letters, 78(12): 1355–1361, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207004154.

Khaledi:2000:DOB

[KK00a] Baha-Eldin Khaledi and Subhash Kochar. On dispersive ordering between order statistics in one-sample and two-sample problems. Statistics & Probability Letters, 46(3):257–261, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001108.

Kim:2000:RCS

[KK00b] Jae-Chun Kim and Alexander Korostelev. Rates of convergence for the sup-norm risk in image models under sequential designs. Statistics & Probability Letters, 46(4):391–399, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001285.

Koo:2000:LDE

[KK00c] Ja-Yong Koo and Charles Kooperberg. Logspline density estimation for binned data. Statistics & Probability Letters, 46(2):133–147, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299000978.

Kumar:2001:EQS

[KK01] Somesh Kumar and Aditi Kar. Estimating quantiles of a selected exponential population. Statistics & Probability Letters, 52(1):9–19, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001747.

Khaniyev:2004:AEM

[KK04] Tahir Khaniyev and Zafer Kucuk. Asymptotic expansions for the moments of the Gaussian random walk with two barriers. Statistics & Probability Letters, 69(1):91–103, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001737.

Khaledi:2005:DOG

[KK05] Baha-Eldin Khaledi and Subhash Kochar. Dependence orderings for generalized order statistics. Statistics & Probability Letters, 73(4):357–367, July 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001409.

Kokonendji:2006:LMI

[KK06a] Célestin C. Kokonendji and Mohamed Khoudar. On Lévy measures for infinitely divisible natural exponential families. Statistics & Probability Letters, 76(13):1364–1368, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206000435.

Kolaiti:2006:UTL

[KK06b] E. Kolaiti and C. Koukouvinos. On the use of three level orthogonal arrays in robust parameter design. Statistics & Probability Letters, 76(3):266–273, February 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002956.

Kim:2008:CLT

[KK08a] Tae-Sung Kim and Mi-Hwa Ko. A Central Limit Theorem for the linear process generated by associated random variables in a Hilbert space. Statistics & Probability Letters, 78(14):2102–2109, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000837.

Kim:2008:CMC

[KK08b] Tae-Sung Kim and Mi-Hwa Ko. Complete moment convergence of moving average processes under dependence assumptions. Statistics & Probability Letters, 78(7):839–846, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002891.

Khaniyev:2008:AEM

[KKAK08] T. Khaniyev, T. Kesemen, R. Aliyev, and A. Kokangul. Asymptotic expansions for the moments of a semi-Markovian random walk with exponential distributed interference of chance. Statistics & Probability Letters, 78(6):785–793, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003331.

Kim:2008:IPL

[KKC08] Tae-Sung Kim, Mi-Hwa Ko, and Yong-Kab Choi. The invariance principle for linear multi-parameter stochastic processes generated by associated fields. Statistics & Probability Letters, 78(18):3298–3303, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003180.

Kim:2002:APP

[KKH02] Tae Soo Kim, Hae Kyung Kim, and Sun Hur. Asymptotic properties of a particular nonlinear regression quantile estimation. Statistics & Probability Letters, 60(4):387–394, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003036.

Kim:2008:ASC

[KKH08] Tae-Sung Kim, Mi-Hwa Ko, and Kwang-Hee Han. On the almost sure convergence for a linear process generated by negatively associated random variables in a Hilbert space. Statistics & Probability Letters, 78(14):2110–2115, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000849.

Karakos:2008:MCI

[KKM^{+}08] Damianos Karakos, Sanjeev Khudanpur, David J. Marchette, Adrian Papamarcou, and Carey E. Priebe. On the minimization of concave information functionals for unsupervised classification via decision trees. Statistics & Probability Letters, 78(8):975–984, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003501.

Kang:2000:KED

[KKP00] Kee-Hoon Kang, Ja-Yong Koo, and Cheol-Woo Park. Kernel estimation of discontinuous regression functions. Statistics & Probability Letters, 47(3):277–285, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001650.

Klebanov:2001:CDS

[KKRV01] L. B. Klebanov, T. J. Kozubowski, S. T. Rachev, and V. E. Volkovich. Characterization of distributions symmetric with respect to a group of transformations and testing of corresponding statistical hypothesis. Statistics & Probability Letters, 53(3):241–247, June 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000116.

Kawczak:2009:EPT

[KKY09] Janusz Kawczak, Reg Kulperger, and Hao Yu. Equivalent processes of total time on test, Lorenz and inverse Lorenz processes. Statistics & Probability Letters, 79(1):125–130, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003611.

Kabaila:2000:CCU

[KL00a] Paul Kabaila and Chris J. Lloyd. A computable confidence upper limit from discrete data with good coverage properties. Statistics & Probability Letters, 47(2):189–198, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529900156X.

Kabaila:2000:WDB

[KL00b] Paul Kabaila and Chris J. Lloyd. When do best confidence limits exist? Statistics & Probability Letters, 50 (2):115–120, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000882.

Kwong:2000:CCV

[KL00c] K. S. Kwong and W. Liu. Calculation of critical values for Dunnett and Tamhane’s step-up multiple test procedure. Statistics & Probability Letters, 49(4):411–416, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000766.

Kabaila:2003:EBC

[KL03a] Paul Kabaila and Chris J. Lloyd. The efficiency of Buehler confidence limits. Statistics & Probability Letters, 65(1):21–28, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002153.

Kim:2003:TBM

[KL03b] Jeong Han Kim and Sungchul Lee. Tail bound for the minimal spanning tree of a complete graph. Statistics & Probability Letters, 64(4):425–430, October 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002086.

Kim:2003:CRD

[KL03c] Yoon Tae Kim and Kee Won Lee. On a criterion of Riemannian distance for singularity and absolute continuity of probability measures. Statistics & Probability Letters, 63(2):185–195, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000828.

Kulikov:2006:LPD

[KL06] Vladimir N. Kulikov and Hendrik P. Lopuhaä. The limit process of the difference between the empirical distribution function and its concave majorant. Statistics & Probability Letters, 76(16):1781–1786, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600143X.

Karlsson:2008:SRE

[KL08a] Maria Karlsson and Thomas Laitila. A semiparametric regression estimator under left truncation and right censoring. Statistics & Probability Letters, 78(16):2567–2571, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800309X.

Kouritzin:2008:ECF

[KL08b] Michael A. Kouritzin and Hongwei Long. On extending classical filtering equations. Statistics & Probability Letters, 78(18):3195–3202, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002988.

Klar:2000:CTE

[Kla00] Bernhard Klar. A class of tests for exponentiality against HN-BUE alternatives. Statistics & Probability Letters, 47(2):199– 207, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001571.

Klass:2009:DSW

[Kla09] Michael J. Klass. A denormalized U-statistic which cannot be decoupled from some associated stopping times. Statistics & Probability Letters, 79(13):1509–1511, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001011.

Kleiber:2002:VOH

[Kle02] Christian Kleiber. Variability ordering of heavy-tailed distributions with applications to order statistics. Statistics & Probability Letters, 58(4):381–388, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001517.

Klein:2003:IMM

[Kle03] Thomas Klein. Information matrices of maximal parameter subsystems in linear models. Statistics & Probability Letters, 62(4):355–360, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000385.

Klimczak:2006:PR

[Kli06] Monika Klimczak. Prediction of k th records. Statistics & Probability Letters, 76(2):117–127, January 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002476.

Ko:2008:CIL

[KLL08] Kyungduk Ko, Jaechoul Lee, and Robert Lund. Confidence intervals for long memory regressions. Statistics & Probability Letters, 78(13):1894–1902, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000564.

Kim:2001:CDL

[KLP01] Choongrak Kim, Yonjoo Lee, and Byeong U. Park. Cook’s distance in local polynomial regression. Statistics & Probability Letters, 54(1):33–40, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000311.

Kachlicka:2002:ESS

[KM02a] Danuta Kachlicka and Iwona Mejza. On the efficiency of some supplemented (α_{1}, α_{2}, …, α_{R})-resolvable block designs. Statistics & Probability Letters, 57(3):291–299, April 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000640.

Kost:2002:CDV

[KM02b] James T. Kost and Michael P. McDermott. Combining dependent p-values. Statistics & Probability Letters, 60(2): 183–190, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202003103.

Kim:2003:MRV

[KM03a] Hyoung-Moon Kim and Bani K. Mallick. Moments of random vectors with skew t distribution and their quadratic forms. Statistics & Probability Letters, 63(4):417–423, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001214. See corrigen-

dum [KM09a].

Kim:2003:NBS

[KM03b] Hyoung-Moon Kim and Bani K. Mallick. A note on Bayesian spatial prediction using the elliptical distribution. Statistics & Probability Letters, 64(3):271–276, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001718.

Koukouvinos:2005:CSO

[KM05] C. Koukouvinos and P. Mantas. Construction of some E(f_{NOD}) optimal mixed-level supersaturated designs. Statistics & Probability Letters, 74(4):312–321, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001902.

Katona:2006:NCS

[KM06] Zsolt Katona and Tamás F. Móri. A new class of scale free random graphs. Statistics & Probability Letters, 76(15):1587–1593, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206001106.

Kedia:2008:CGN

[KM08] R. G. Kedia and B. L. Misra. On construction of generalized neighbor design of use in serology. Statistics & Probability Letters, 78(3):254–256, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002210.

Kim:2009:CMR

[KM09a] Hyoung-Moon Kim and Bani K. Mallick. Corrigendum to: “Moments of random vectors with skew t distribution and their quadratic forms” [Statist. Probab. Lett. 63 (2003) 417– 423]. Statistics & Probability Letters, 79(19):2098–2099, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209002478. See [KM03a].

Kozubowski:2009:BID

[KM09b] Tomasz J. Kozubowski and Mark M. Meerschaert. A bivariate infinitely divisible distribution with exponential and Mittag–Leffler marginals. Statistics & Probability Letters, 79(14):1596–1601, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001266.

Kundu:2000:CLT

[KMM00] Subrata Kundu, Suman Majumdar, and Kanchan Mukherjee. Central limit theorems revisited. Statistics & Probability Letters, 47(3):265–275, April 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001649.

Kotz:2000:GPU

[KMR00] Samuel Kotz, Hosam Mahmoud, and Philippe Robert. On generalized Pólya urn models. Statistics & Probability Letters, 49(2):163–173, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000456.

Klein:2008:AEF

[KMS08] André Klein, Guy Mélard, and Abdessamad Saidi. The asymptotic and exact Fisher information matrices of a vector ARMA process. Statistics & Probability Letters, 78(12):1430–1433, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207004257.

Kumar:2009:RES

[KMV09] Somesh Kumar, Ajaya Kumar Mahapatra, and P. Vellaisamy. Reliability estimation of the selected exponential populations. Statistics & Probability Letters, 79(11):1372–1377, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209000753.

Klaassen:2000:SSM

[KMvE00] Chris A. J. Klaassen, Philip J. Mokveld, and Bert van Es. Squared skewness minus kurtosis bounded by 186/125 for unimodal distributions. Statistics & Probability Letters, 50 (2):131–135, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000900.

Kuk:2000:PLA

[KN00] Anthony Y. C. Kuk and David J. Nott. A pairwise likelihood approach to analyzing correlated binary data. Statistics & Probability Letters, 47(4):329–335, May 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001741.

Kagan:2001:HMM

[KN01a] Abram Kagan and Sergei Nagaev. How many moments can be estimated from a large sample? Statistics & Probability Letters, 55(1):99–105, November 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001389.

Kotz:2001:SET

[KN01b] Samuel Kotz and Saralees Nadarajah. Some extremal type elliptical distributions. Statistics & Probability Letters, 54 (2):171–182, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000396.

Klass:2005:GZI

[KN05] Michael J. Klass and Krzysztof Nowicki. The Grossman and Zhou investment strategy is not always optimal. Statistics & Probability Letters, 74(3):245–252, October 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001641.

Kozubowski:2008:IDS

[KN08] Tomasz J. Kozubowski and John P. Nolan. Infinite divisibility of skew Gaussian and Laplace laws. Statistics & Probability Letters, 78(6):654–660, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003082.

Kowalczyk:2000:DKI

[KNB00] T. Kowalczyk and M. Niewiadomska-Bugaj. Decomposition of Kendall’s τ: implications for clustering. Statistics & Probability Letters, 48(4):375–383, July 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000195.

Kaluszka:2001:EEN

[KO01] M. Kaluszka and A. Okolewski. An extension of the Erdős– Neveu–Rényi theorem with applications to order statistics. Statistics & Probability Letters, 55(2):181–186, November 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001146.

Konsowa:2003:DRT

[KO03] Mokhtar H. Konsowa and Tamer F. Oraby. Dimensions of random trees. Statistics & Probability Letters, 62(1):49–60, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202004248.

Kaluszka:2004:FTL

[KO04] M. Kaluszka and A. Okolewski. On Fatou-type lemma for monotone moments of weakly convergent random variables. Statistics & Probability Letters, 66(1):45–50, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003055.

Konsowa:2008:RDB

[KO08] Mokhtar H. Konsowa and Tamer F. Oraby. Resistance dimensions of branching processes in varying environments trees. Statistics & Probability Letters, 78(1):67–74, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001964.

Kochar:2006:LOO

[Koc06] Subhash Kochar. Lorenz ordering of order statistics. Statistics & Probability Letters, 76(17):1855–1860, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001519.

Kokonendji:2001:FPT

[Kok01] Célestin C. Kokonendji. First passage times on zero and one and natural exponential families. Statistics & Probability Letters, 51(3):293–298, February 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001668.

Kolassa:2000:SAE

[Kol00a] John E. Kolassa. Saddlepoint approximation at the edges of a conditional sample space. Statistics & Probability Letters, 50(4):343–349, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001012.

Kolodynski:2000:NSE

[Kol00b] Slawomir Kolodynski. A note on the sequence of expected extremes. Statistics & Probability Letters, 47(3):295–300, April 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001698.

Konsowa:2002:RWT

[Kon02] Mokhtar H. Konsowa. Random walks on trees and the law of iterated logarithm. Statistics & Probability Letters, 56(2):193–197, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001857.

Kontorovich:2008:CPP

[Kon08] Leonid (Aryeh) Kontorovich. Constructing processes with prescribed mixing coefficients. Statistics & Probability Letters, 78(17):2910–2915, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208002356.

Konsowa:2009:SRW

[Kon09] Mokhtar H. Konsowa. On the speed of random walks on random trees. Statistics & Probability Letters, 79(20):2230–2236, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209002752.

Kosorok:2000:MCE

[Kos00] Michael R. Kosorok. Monte Carlo error estimation for multivariate Markov chains. Statistics & Probability Letters, 46(1):85–93, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000905.

Kosorok:2002:GCB

[Kos02] Michael R. Kosorok. On global consistency of a bivariate survival estimator under univariate censoring. Statistics & Probability Letters, 56(4):439–446, February 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000445.

Kosinski:2009:ASF

[Kos09a] Kamil Kosiński. Asymptotics for sums of a function of normalized independent sums. Statistics & Probability Letters, 79(4):415–419, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004276.

Kosinski:2009:FLS

[Kos09b] Kamil M. Kosiński. On the functional limits for sums of a function of partial sums. Statistics & Probability Letters, 79(13):1522–1527, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001151.

Kou:2003:EBM

[Kou03] S. C. Kou. Is C_{p} an empirical Bayes method for smoothing parameter choice? Statistics & Probability Letters, 65(2):139–146, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002621.

Koudou:2006:LBM

[Kou06] Angelo Efoévi Koudou. A link between the Matsumoto–Yor property and an independence property on trees. Statistics & Probability Letters, 76(11):1097–1101, June 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004529.

Kowalczyk:2000:LBG

[Kow00] Teresa Kowalczyk. Link between grade measures of dependence and of separability in pairs of conditional distributions. Statistics & Probability Letters, 46(4):371–379, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529900125X.

Kozubowski:2005:NSDa

[Koz05a] Tomasz J. Kozubowski. A note on self-decomposability of stable process subordinated to self-decomposable subordinator. Statistics & Probability Letters, 73(4):343–345, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520500129X.

Kozubowski:2005:NSDb

[Koz05b] Tomasz J. Kozubowski. A note on self-decomposability of stable process subordinated to self-decomposable subordinator. Statistics & Probability Letters, 74(1):89–91, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001628.

Koenker:2000:SPR

[KP00] Roger Koenker and Stephen Portnoy. Some pathological regression asymptotics under stable conditions. Statistics & Probability Letters, 50(3):219–228, November 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001073.

Kozek:2002:DFC

[KP02] Andrzej S. Kozek and Mirosław Pawlak. Distribution-free consistency of kernel non-parametric M-estimators. Statistics & Probability Letters, 58(4):343–353, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001062.

Kasahara:2009:DRV

[KPI09] Yukio Kasahara, Mohsen Pourahmadi, and Akihiko Inoue. Duals of random vectors and processes with applications to prediction problems with missing values. Statistics & Probability Letters, 79(14):1637–1646, July 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001503.

Kim:2002:IDS

[KPK02] Choongrak Kim, Byeong U. Park, and Woochul Kim. Influence diagnostics in semiparametric regression models. Statistics & Probability Letters, 60(1):49–58, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002687.

Klimczak:2004:MVK

[KR04] Monika Klimczak and Tomasz Rychlik. Maximum variance of k th records. Statistics & Probability Letters, 69(4):421–430, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001956.

Kundu:2009:ETP

[KR09] Debasis Kundu and Mohammad Z. Raqab. Estimation of R = P (Y < X) for three-parameter Weibull distribution. Statistics & Probability Letters, 79(17):1839–1846, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001904.

Krajka:2000:CWL

[Kra00] A. Krajka. Characterization of weak limits of randomly indexed sequences. Statistics & Probability Letters, 50(2): 155–163, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000948.

Krantsberg:2002:KMZ

[Kra02]  Alexander Krantsberg.  Khinchin and Marcinkiewicz–

Zygmund-type inequalities for quadratic forms of independent random variables. Statistics & Probability Letters, 60 (3):321–327, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202003176.

Kraus:2004:GFI

[Kra04] David Kraus. Goodness-of-fit inference for the Cox–Aalen additive-multiplicative regression model. Statistics & Probability Letters, 70(4):285–298, December 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002779.

Kratschmer:2007:UEE

[Krä07] Volker Krätschmer. The uniqueness of extremum estimation. Statistics & Probability Letters, 77(10):942–951, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000259.

Krasnosielska:2009:VEP

[Kra09] Anna Krasnosielska. A version of the Elfving problem with random starting time. Statistics & Probability Letters, 79(23):2429–2436, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003265.

Kruglov:2008:SLL

[Kru08] Victor M. Kruglov. A Strong Law of Large Numbers for pairwise independent identically distributed random variables with infinite means. Statistics & Probability Letters, 78(7):890–895, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002969.

Klesov:2005:ASG

[KRV05] Oleg Klesov, Andrew Rosalsky, and Andrei I. Volodin. On the almost sure growth rate of sums of lower negatively dependent nonnegative random variables. Statistics & Probability Letters, 71(2):193–202, February 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002949.

Kwiecinska:2000:RDS

[KS00] Anna A. Kwiecińska and Wojciech Slomczyński. Random dynamical systems arising from iterated function systems with place-dependent probabilities. Statistics & Probability Letters, 50(4):401–407, December 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001309.

Koukouvinos:2004:OML

[KS04] C. Koukouvinos and S. Stylianou. Optimal multi-level supersaturated designs constructed from linear and quadratic functions. Statistics & Probability Letters, 69(2):199–211, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001853.

Koul:2005:GFT

[KS05] Hira L. Koul and Lyudmila Sakhanenko. Goodness-of-fit testing in regression: a finite sample comparison of bootstrap methodology and Khmaladze transformation. Statistics & Probability Letters, 74(3):290–302, October 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001896.

Kella:2006:SRP

[KS06] Offer Kella and Wolfgang Stadje. Superposition of renewal processes and an application to multi-server queues. Statistics & Probability Letters, 76(17):1914–1924, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001581.

Khaledi:2007:SOB

[KS07] Baha-Eldin Khaledi and Roohollah Shojaei. On stochastic orderings between residual record values. Statistics & Probability Letters, 77(13):1467–1472, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001241.

Kaziska:2008:KMC

[KS08a] David Kaziska and Anuj Srivastava. The Karcher mean of a class of symmetric distributions on the circle. Statistics & Probability Letters, 78(11):1314–1316, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004117.

Kiselak:2008:EOD

[KS08b] Jozef Kiselák and Milan Stehlík. Equidistant and D-optimal designs for parameters of Ornstein–Uhlenbeck process. Statistics & Probability Letters, 78(12):1388–1396, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004208.

Kuroda:2008:AEE

[KSG08] Masahiro Kuroda, Michio Sakakihara, and Zhi Geng. Acceleration of the EM and ECM algorithms using the Aitken δ^{2} method for log-linear models with partially classified data. Statistics & Probability Letters, 78(15):2332–2338, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001302.

Kim:2007:CAE

[KSH07] Jong-Min Kim, Engin A. Sungur, and Tae-Young Heo. Calibration approach estimators in stratified sampling. Statistics & Probability Letters, 77(1):99–103, January 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600201X.

Kitano:2005:GCS

[KSO05] Masashi Kitano, Kunio Shimizu, and S. H. Ong. The generalized Charlier series distribution as a distribution with two-step recursion. Statistics & Probability Letters, 75(4):280–290, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002427.

Karunamuni:2006:ANA

[KSW06a] R. J. Karunamuni, T. N. Sriram, and J. Wu. Asymptotic normality of an adaptive kernel density estimator for finite mixture models. Statistics & Probability Letters, 76(2):211–220, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002907.

Karunamuni:2006:RCA

[KSW06b] R. J. Karunamuni, T. N. Sriram, and J. Wu. Rates of convergence of an adaptive kernel density estimator for finite mixture models. Statistics & Probability Letters, 76(3):221–230, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002919.

Kathman:2003:PAC

[KT03] Steven J. Kathman and George R. Terrell. Poisson approximation by constrained exponential tilting. Statistics & Probability Letters, 61(1):83–89, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003504.

Kuchler:2008:SBG

[KT08] Uwe Küchler and Stefan Tappe. On the shapes of bilateral Gamma densities. Statistics & Probability Letters, 78 (15):2478–2484, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001521.

Kanaan:2008:CSP

[KTM08] M. N. Kanaan, Paul C. Taylor, and M. A. Mugglestone. Crossspectral properties of a spatial point-lattice process. Statistics & Probability Letters, 78(18):3238–3243, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003040.

Koutras:2003:ERB

[KTZ03] M. V. Koutras, S. Tsitmidelis, and V. Zissimopoulos. Evaluation of reliability bounds by set covering models. Statistics & Probability Letters, 61(2):163–175, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003462.

Kubilius:2008:CSI

[Kub08] K. Kubilius. On the convergence of stochastic integrals with respect to p-semimartingales. Statistics & Probability Letters, 78(15):2528–2535, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001661.

Kuchler:2004:IRL

[Küc04a] Uwe Küchler. On integrals with respect to Lévy processes. Statistics & Probability Letters, 66(2):145–151, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003018.

Kuczmaszewska:2004:SCC

[Kuc04b] Anna Kuczmaszewska. On some conditions for complete convergence for arrays. Statistics & Probability Letters, 66(4): 399–405, March 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203003523.

Kuczmaszewska:2005:SLL

[Kuc05] Anna Kuczmaszewska. The Strong Law of Large Numbers for dependent random variables. Statistics & Probability Letters, 73(3):305–314, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001240.

Kuczmaszewska:2007:CCA

[Kuc07] Anna Kuczmaszewska. On complete convergence for arrays of rowwise dependent random variables. Statistics & Probability Letters, 77(11):1050–1060, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000235.

Kuczmaszewska:2009:CCA

[Kuc09] Anna Kuczmaszewska. On complete convergence for arrays of rowwise negatively associated random variables. Statistics & Probability Letters, 79(1):116–124, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003556.

Kuhn:2001:ENC

[Küh01] Christoph Kühn. An estimator of the number of change points based on a weak invariance principle. Statistics & Probability Letters, 51(2):189–196, January 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001553.

Kulasekera:2001:VSS

[Kul01] K. B. Kulasekera. Variable selection by stepwise slicing in nonparametric regression. Statistics & Probability Letters, 51(4):327–336, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520000167X.

Kulik:2006:LTS

[Kul06] Rafal Kulik. Limit theorems for self-normalized linear processes. Statistics & Probability Letters, 76(18):1947–1953, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001623.

Kumar:2002:EGH

[Kum02] C. Satheesh Kumar. Extended generalized hypergeometric probability distributions. Statistics & Probability Letters, 59(1):1–7, August 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001025.

Kurenok:2008:DSE

[Kur08]  V. P. Kurenok.  On degenerate stochastic equations

of Itô type with jumps. Statistics & Probability Letters, 78(17):2917–2925, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800237X.

Kuzin:2005:RDD

[Kuz05] Vladimir Kuzin. Recursive demeaning and deterministic seasonality. Statistics & Probability Letters, 72(3):195–204, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000283.

Karandikar:2002:RUC

[KV02] Rajeeva L. Karandikar and M. Vidyasagar. Rates of uniform convergence of empirical means with mixing processes. Statistics & Probability Letters, 58(3):297–307, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001244.

Kotz:2005:LBT

[KvD05] Samuel Kotz and J. René van Dorp. A link between twosided power and asymmetric Laplace distributions: with applications to mean and variance approximations. Statistics & Probability Letters, 71(4):383–394, March 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500012X.

Kruglov:2006:CCA

[KVH06] Victor M. Kruglov, Andrei I. Volodin, and Tien-Chung Hu. On complete convergence for arrays. Statistics & Probability Letters, 76(15):1631–1640, September 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001179.

Kagan:2000:EDS

[KW00] Abram Kagan and Jacek Wesołowski. An extension of the Darmois–Skitovitch theorem to a class of dependent random variables. Statistics & Probability Letters, 47(1):69–73, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529900139X.

Kwong:2002:MPS

[KW02] Koon-Shing Kwong and Ee-Hwee Wong. A more powerful step-up procedure for controlling the false discovery rate under independence. Statistics & Probability Letters, 56(2):217–225, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001961.

Karabatsos:2007:BNI

[KW07a] George Karabatsos and Stephen G. Walker. Bayesian nonparametric inference of stochastically ordered distributions, with Pólya trees and Bernstein polynomials. Statistics & Probability Letters, 77(9):907–913, May 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000296.

Kume:2007:DNC

[KW07b] A. Kume and Andrew T. A. Wood. On the derivatives of the normalising constant of the Bingham distribution. Statistics & Probability Letters, 77(8):832–837, April 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700017X.

Kuo:2000:BRM

[KY00] Lynn Kuo and Tae Young Yang. Bayesian reliability modeling for masked system lifetime data. Statistics & Probability Letters, 47(3):229–241, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001601.

Krishnamoorthy:2004:MNV

[KY04] K. Krishnamoorthy and Jianqi Yu. Modified Nel and Van der Merwe test for the multivariate Behrens–Fisher problem. Statistics & Probability Letters, 66(2):161–169, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003298.

Koul:2006:GFT

[KY06] Hira L. Koul and Tingting Yi. Goodness-of-fit testing in interval censoring case 1. Statistics & Probability Letters, 76(7):709–718, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003792.

Kagan:2009:GPS

[KY09] Abram M. Kagan and Tinghui Yu. A geometric property of the sample mean and residuals. Statistics & Probability Letters, 79(11):1409–1413, June 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900100X.

Kypraios:2009:NML

[Kyp09] Theodore Kypraios. A note on maximum likelihood estimation of the initial number of susceptibles in the general stochastic epidemic model. Statistics & Probability Letters, 79(18):1972–1976, September 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002156.

Kuhfeld:2005:SNO

[KyS05] Warren F. Kuhfeld and Chung yi Suen. Some new orthogonal arrays OA(4r, r^{1}2^{p}, 2). Statistics & Probability Letters, 75(3):169–178, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002233.

Karunamuni:2008:SIB

[KZ08a] R. J. Karunamuni and S. Zhang. Some improvements on a boundary corrected kernel density estimator. Statistics & Probability Letters, 78(5):499–507, April 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002817.

Kim:2008:ELL

[KZ08b] Mi-Ok Kim and Mai Zhou. Empirical likelihood for linear models in the presence of nuisance parameters. Statistics & Probability Letters, 78(12):1445–1451, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004282.

Kong:2008:FTR

[KZ08c] Fanchao Kong and Gaofeng Zong. The finite-time ruin probability for ND claims with constant interest force. Statistics & Probability Letters, 78(17):3103–3109, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002769.

Kulasekera:2008:PBT

[KZ08d] K. B. Kulasekera and Meng Zhao. A pointwise Bayes-type estimator of the survival probability with censored data. Statistics & Probability Letters, 78(16):2597–2603, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003659.

Kozak:2008:STS

[KZS08] Marcin Kozak, Andrzej Zieliński, and Sarjinder Singh. Stratified two-stage sampling in domains: Sample allocation between domains, strata, and sampling stages. Statistics & Probability Letters, 78(8):970–974, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003495.

Liao:2007:TBC

[LA07] Shu-Min Liao and Michael Akritas. Test-based classification: a linkage between classification and statistical testing. Statistics & Probability Letters, 77(12):1269–1281, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000909.

Labrador:2008:SPC

[Lab08] Boris Labrador. Strong pointwise consistency of the k_{T} - occupation time density estimator. Statistics & Probability Letters, 78(9):1128–1137, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003847.

Lacaze:2004:MSU

[Lac04] B. Lacaze. Matched shapes for uniform sampling. Statistics & Probability Letters, 70(2):127–135, November 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001579.

Lachal:2008:NSI

[Lac08] Aimé Lachal. A note on Spitzer identity for random walk. Statistics & Probability Letters, 78(2):97–108, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207002003.

Ladoucette:2007:ABM

[Lad07] Sophie A. Ladoucette. Asymptotic behavior of the moments of the ratio of the random sum of squares to the square of the random sum. Statistics & Probability Letters, 77(10):1021–1033, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000429.

Laitila:2001:PQU

[Lai01] Thomas Laitila. Properties of the QME under asymmetrically distributed disturbances. Statistics & Probability Letters, 52(4):347–352, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002029.

Laloe:2008:NNA

[Lal08] Thomas Laloë. A k-nearest neighbor approach for functional regression. Statistics & Probability Letters, 78(10):1189–1193, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207003914.

Landsman:2006:GSL

[Lan06] Zinoviy Landsman. On the generalization of Stein’s Lemma for elliptical class of distributions. Statistics & Probability Letters, 76(10):1012–1016, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004414.

Lanconelli:2008:NIU

[Lan08] Alberto Lanconelli. A note on the invariance under change of measure for stochastic test functions and distribution spaces. Statistics & Probability Letters, 78(18):3135–3138, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002812.

Laurincikas:2005:LTE

[Lau05] Antanas Laurincikas. Limit theorems for the Estermann zetafunction. I. Statistics & Probability Letters, 72(3):227–235, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000325.

Lau:2006:BSP

[Lau06] John W. Lau. Bayesian semi-parametric modeling for mixed proportional hazard models with right censoring. Statistics & Probability Letters, 76(7):719–728, April 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003810.

Liu:2003:LTP

[LaYY03] Wen Liu, Jia an Yan, and Weiguo Yang. A limit theorem for partial sums of random variables and its applications. Statistics & Probability Letters, 62(1):79–86, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202004285.

Liu:2002:SSD

[LB02] W. Liu and B. J. R. Bailey. Sample size determination for constructing a constant width confidence interval for a binomial success probability. Statistics & Probability Letters, 56(1):1–5, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000293.

Lien:2003:CCA

[LB03] Donald Lien and N. Balakrishnan. Conditional correlation analysis of order statistics from bivariate normal distribution with an application to evaluating inventory effects in futures market. Statistics & Probability Letters, 63(3):249–257, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000890.

Landagan:2007:EPS

[LB07] Ohmar Z. Landagan and Erniel B. Barrios. An estimation procedure for a spatial-temporal model. Statistics & Probability Letters, 77(4):401–406, February 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002574.

Lopez-Blazquez:2005:NDR

[LBMD05] F. López-Blázquez, B. Salamanca Miño, and A. Dembińska. A note on the distribution of k th records from discrete distributions. Statistics & Probability Letters, 75(4):325–330, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002518.

Lopez-Blazquez:2000:RMP

[LBSM00] Fernando López-Blázquez and Begoña Salamanca-Miño. A reverse martingale property that characterizes the natural exponential family with quadratic variance function. Statistics & Probability Letters, 49(1):63–68, August 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000033X.

Larsson-Cohn:2000:IPF

[LC00] Lars Larsson-Cohn. Invariance principles for the first passage times of perturbed random walks. Statistics & Probability Letters, 48(4):347–351, July 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000158.

Lu:2004:SKR

[LC04] Zudi Lu and Xing Chen. Spatial kernel regression estimation: weak consistency. Statistics & Probability Letters, 68(2):125–136, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002918.

Liquet:2005:CVM

[LC05] Benoit Liquet and Daniel Commenges. Computation of the p-value of the maximum of score tests in the generalized linear model; application to multiple coding. Statistics & Probability Letters, 71(1):33–38, January 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002809.

Li:2006:NAP

[LCJ06] Gang Li, Kan Cheng, and Xiaoyue Jiang. Negative ageing property of random sum. Statistics & Probability Letters, 76(7):737–742, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003834.

Li:2006:NMA

[LCLZ06] Peng-Fei Li, Bao-Jiang Chen, Min-Qian Liu, and Run-Chu Zhang. A note on minimum aberration and clear criteria. Statistics & Probability Letters, 76(10):1007–1011, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004402.

Lee:2004:LLD

[LCPY04] Y. K. Lee, H. Choi, B. U. Park, and K. S. Yu. Local likelihood density estimation on random fields. Statistics & Probability Letters, 68(4):347–357, July 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001233.

Li:2009:CRD

[LCZ09] Gaorong Li, Shuang Chen, and Junhua Zhang. A class of random deviation theorems and the approach of Laplace transform. Statistics & Probability Letters, 79(2):202–210, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003696.

Lian:2007:ADC

[LdC07] Zilong Lian and Enrique del Castillo. Adaptive deadband control of a drifting process with unknown parameters. Statistics & Probability Letters, 77(8):843–852, April 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000211.

LeGall:2003:DMM

[Le 03] Françoise Le Gall. Determination of the modes of a multinomial distribution. Statistics & Probability Letters, 62(4):325– 333, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202004303.

LeGall:2006:MNM

[Le 06] Françoise Le Gall. The modes of a negative multinomial distribution. Statistics & Probability Letters, 76(6):619–624, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003639.

Le:2008:BGP

[Le08] H. Le. On bounded Gaussian processes. Statistics & Probability Letters, 78(6):669–674, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003100.

Lechner:2008:NEC

[Lec08] Michael Lechner. A note on endogenous control variables in causal studies. Statistics & Probability Letters, 78(2):190–195, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002131.

Ledoux:2008:SRM

[Led08a] James Ledoux. On a software reliability model by Koch and Spreij. Statistics & Probability Letters, 78(7):904–914, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003409.

Ledoux:2008:SCC

[Led08b] James Ledoux. Strong convergence of a class of nonhomogeneous Markov arrival processes to a Poisson process. Statistics & Probability Letters, 78(4):445–455, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700260X.

Lee:2000:PPN

[Lee00a] Oesook Lee. On probabilistic properties of nonlinear ARMA(p, q) models. Statistics & Probability Letters, 46 (2):121–131, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000966.

Lee:2000:RSS

[Lee00b] Thomas C. M. Lee. Regression spline smoothing using the minimum description length principle. Statistics & Probability Letters, 48(1):71–82, May 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001911.

Lee:2003:SES

[Lee03] Sangyeol Lee. The sequential estimation in stochastic regression model with random coefficients. Statistics & Probability Letters, 61(1):71–81, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003401.

Lee:2004:OCP

[Lee04a] R. Lee. An optimal choice problem for uniform random variables seen simultaneously or sequentially with availability depending on the observation. Statistics & Probability Letters, 66(1):1–8, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002220.

Lee:2004:ISS

[Lee04b] Thomas C. M. Lee. Improved smoothing spline regression by combining estimates of different smoothness. Statistics & Probability Letters, 67(2):133–140, April 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000100.

Lee:2006:BRT

[Lee06] Sangyeol Lee. The Bickel–Rosenblatt test for diffusion processes. Statistics & Probability Letters, 76(14):1494–1502, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000848.

Lefebvre:2008:GRP

[Lef08a] Mario Lefebvre. The gambler’s ruin problem for a Markov chain related to the Bessel process. Statistics & Probability Letters, 78(15):2314–2320, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001260.

Lefebvre:2008:GFP

[Lef08b] Mario Lefebvre. Generalized filtered Poisson processes and application in hydrology. Statistics & Probability Letters, 78(18):3274–3276, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208003131.

Lehmann:2005:PCS

[Leh05] Nils Lehmann. Principal components selection given extensively many variables. Statistics & Probability Letters, 74(1):51–58, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001574.

Lelong:2008:ASC

[Lel08] Jérôme Lelong. Almost sure convergence of randomly truncated stochastic algorithms under verifiable conditions. Statistics & Probability Letters, 78(16):2632–2636, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001715.

Leon:2001:MAV

[Leó01] Carlos A. León. Maximum asymptotic variance of sums of finite Markov chains. Statistics & Probability Letters, 54 (4):413–415, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001183.

Leonenko:2009:NFT

[Leo09] G. M. Leonenko. A new formula for the transient solution of the Erlang queueing model. Statistics & Probability Letters, 79(3):400–406, February 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004288.

Lepeltier:2003:NRS

[Lep03] J. P. Lepeltier. Nonanticipative risk sensitive control: the martingale method. Statistics & Probability Letters, 64(2): 191–199, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001512.

Letac:2002:DWD

[Let02] Gérard Letac. Donkey walk and Dirichlet distributions. Statistics & Probability Letters, 57(1):17–22, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000275.

Levental:2000:PSB

[Lev00] Shlomo Levental. Permutations, signs and the Brownian bridge. Statistics & Probability Letters, 46(3):271–276, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001121.

Lou:2007:ETT

[LF07] W. Y. Wendy Lou and James C. Fu. On exact Type I and Type II errors of Cochran’s test. Statistics & Probability Letters, 77(12):1282–1287, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000910.

Leao:2005:SHM

[LFdVA05] D. Leão Jr., M. D. Fragoso, J. B. R. do Val, and M. G. Andrade. Separable Hausdorff measurable Radon spaces. Statistics & Probability Letters, 71(4):347–359, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000064.

Li:2003:ORE

[LG03] Jianjun Li and Shanti S. Gupta. Optimal rate of empirical Bayes tests for lower truncation parameters. Statistics & Probability Letters, 65(3):177–185, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002499.

Lopez-Garcia:2009:CSS

[LG09] Marcos López-García. Characterization of solutions to the Stieltjes–Wigert moment problem. Statistics & Probability Letters, 79(10):1337–1342, May 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000698.

Li:2003:TMO

[LGH03] Gang Li, Xiong Gao, and Minqiang Huang. Testing multivariate one-sided hypotheses. Statistics & Probability Letters, 64(1):63–68, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001317.

Lopez-Garcia:2000:IVQ

[LGLDG00] Hortensia López-García, Miguel López-Díaz, and María Angeles Gil. Interval-valued quantification of the inequality associated with a random set. Statistics & Probability Letters, 46(2):149–159, January 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001005.

Lin:2001:CDS

[LH01] Gwo Dong Lin and Chin-Yuan Hu. Characterizations of distributions via the stochastic ordering property of random linear forms. Statistics & Probability Letters, 51(1):93–99, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001607.

Li:2008:SNS

[LH08a] Xiaohu Li and Xiaoxiao Hu. Some new stochastic comparisons for redundancy allocations in series and parallel systems. Statistics & Probability Letters, 78(18):3388–3394, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800343X.

Lin:2008:SAI

[LH08b] Tsung I. Lin and Hsiu J. Ho. A simplified approach to inverting the autocovariance matrix of a general ARMA(p, q) process. Statistics & Probability Letters, 78(1):36–41, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001927.

Lin:2004:PPD

[LHLC04] Zhengyan Lin, Kyo-Shin Hwang, Sungchul Lee, and Yong-Kab Choi. Path properties of a d-dimensional Gaussian process. Statistics & Probability Letters, 68(4):383–393, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001282.

Li:2000:ENE

[Li00] Tong Li. Estimation of nonlinear errors-in-variables models: a simulated minimum distance estimator. Statistics & Probability Letters, 47(3):243–248, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001613.

Li:2002:LBQ

[Li02] Heng Li. Likelihood-based quadratic analyses for diallel cross and other experiments involving data collected on ordered pairs of sampling units. Statistics & Probability Letters, 58(1):61–69, May 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001165.

Li:2003:AMP

[Li03a] Haijun Li. Association of multivariate phase-type distributions, with applications to shock models. Statistics & Probability Letters, 64(4):381–392, October 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001822.

Li:2003:MIL

[Li03b] Yulin Li. A martingale inequality and large deviations. Statistics & Probability Letters, 62(3):317–321, April 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000464.

Li:2004:CNC

[Li04] Yulin Li. Closure of NBU(2) class under the formation of parallel systems. Statistics & Probability Letters, 67(1):57–63, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000082.

Li:2006:VHT

[Li06] Baibing Li. The p-values of the hypothesis testing about relative risks. Statistics & Probability Letters, 76(16):1731–1734, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001271.

Li:2007:SMF

[Li07] Yangrong Li. Strongly monotone q-functions and a note on strong ergodicity of monotone q-functions. Statistics & Probability Letters, 77(4):396–400, February 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002562.

Li:2009:GVA

[Li09a] Xinmin Li. A generalized p-value approach for comparing the means of several log-normal populations. Statistics & Probability Letters, 79(11):1404–1408, June 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000996.

Li:2009:CRL

[Li09b] Yun-Xia Li. Convergence rates in the law of the iterated logarithm for negatively associated random variables with multidimensional indices. Statistics & Probability Letters, 79(8):1038–1043, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005737.

Li:2009:LTG

[Li09c] Yuqiang Li. Limit theorems for generalized Jirina processes. Statistics & Probability Letters, 79(2):158–164, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800360X.

Liang:2000:CCW

[Lia00a] Han-Ying Liang. Complete convergence for weighted sums of negatively associated random variables. Statistics & Probability Letters, 48(4):317–325, July 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000002X.

Liao:2000:OPF

[Lia00b] C. T. Liao. Orthogonal parallel-flats designs for hierarchical models. Statistics & Probability Letters, 46(4):329–335, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001194.

Liao:2001:SCO

[Lia01] C. T. Liao. Some classes of orthogonal 2^{n_{1}} × 3^{n_{2}} mixed factorial designs of median-resolution. Statistics & Probability Letters, 52(4):401–411, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000074.

Liang:2002:SCB

[Lia02a] Faming Liang. Some connections between Bayesian and non-Bayesian methods for regression model selection. Statistics & Probability Letters, 57(1):53–63, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000482.

Liao:2002:ILC

[Lia02b] Jason J. Z. Liao. An insight into linear calibration: univariate case. Statistics & Probability Letters, 56(3):271–281, February 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001900.

Liang:2004:OCR

[Lia04] Ta Chen Liang. On optimal convergence rate of empirical Bayes tests. Statistics & Probability Letters, 68(2):189–198, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000860.

Lian:2007:CBE

[Lia07] Heng Lian. Consistency of Bayesian estimation of a step function. Statistics & Probability Letters, 77(1):19–24, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001908.

Lian:2009:CVC

[Lia09a] Heng Lian. Cross-validation for comparing multiple density estimation procedures. Statistics & Probability Letters, 79(1):112–115, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208003544.

Liang:2009:USA

[Lia09b] Faming Liang. On the use of stochastic approximation Monte Carlo for Monte Carlo integration. Statistics & Probability Letters, 79(5):581–587, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004690.

Liang:2009:CEP

[Lia09c] TaChen Liang. Comments on “Estimating the parameter of the population selected from discrete exponential family”. Statistics & Probability Letters, 79(20):2208–2211, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209002715.

Liang:2009:EBE

[Lia09d] TaChen Liang. Empirical Bayes estimation for Borel–Tanner distributions. Statistics & Probability Letters, 79(20):2212– 2219, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002727.

Lichtendahl:2009:RQD

[Lic09] Kenneth C. Lichtendahl, Jr. Random quantiles of the Dirichlet process. Statistics & Probability Letters, 79(4):501–507, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004574.

Liero:2001:TSM

[Lie01] Hannelore Liero. L_{2}-tests for sparse multinomials. Statistics & Probability Letters, 55(2):147–158, November 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000797.

Lillo:2004:CPC

[Lil04] Rosa E. Lillo. On the characterizing property of the convex conditional mean function. Statistics & Probability Letters, 66(1):19–24, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520300289X.

Lim:2005:TIS

[Lim05] Johan Lim. Testing independent sparse heterogeneous mixtures. Statistics & Probability Letters, 74(2):205–211, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001859.

Limnios:2006:ESD

[Lim06] N. Limnios. Estimation of the stationary distribution of semi-Markov processes with Borel state space. Statistics & Probability Letters, 76(14):1536–1542, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000964.

Lin:1997:MPO

[Lin97] Gwo Dong Lin. On the moment problems. Statistics & Probability Letters, 35(1):85–90, August 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002209. See erratum [Lin00].

Lin:2000:EMP

[Lin00] Gwo Dong Lin. Erratum: “On the moment problems” [Statist. Probab. Lett. 35 (1997) 85–90]. Statistics & Probability Letters, 50(2):205, November 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000153X. See [Lin97].

Lin:2004:NMB

[Lin04] Yi Lin. A note on margin-based loss functions in classification. Statistics & Probability Letters, 68(1):73–82, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000707.

Lin:2008:GPP

[Lin08a] Jianxi Lin. The general principle for precise large deviations of heavy-tailed random sums. Statistics & Probability Letters, 78(6):749–758, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003288.

Lin:2008:OSL

[Lin08b] Jianxi Lin. A one-sided large deviation local limit theorem. Statistics & Probability Letters, 78(16):2679–2684, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001843.

Ling:2008:BRS

[Lin08c] Nengxiang Ling. The Bahadur representation for sample quantiles under negatively associated sequence. Statistics & Probability Letters, 78(16):2660–2663, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001818.

Lin:2009:CBD

[Lin09] Qian Lin. A class of backward doubly stochastic differential equations with non-Lipschitz coefficients. Statistics & Probability Letters, 79(20):2223–2229, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002740.

Liu:2000:LPR

[Liu00] Wen Liu. A limit property of random conditional probabilities. Statistics & Probability Letters, 49(3):299–304, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000614.

Liu:2003:SLP

[Liu03] Wen Liu. Some limit properties of the multivariate function sequences of discrete random variables. Statistics & Probability Letters, 61(1):41–50, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003048.

Liu:2006:TBB

[Liu06] Ji-Chun Liu. On the tail behaviors of Box–Cox transformed threshold GARCH(1,1) process. Statistics & Probability Letters, 76(13):1323–1330, July 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000277.

Liu:2007:SMS

[Liu07a] Ji-Chun Liu. Stationarity for a Markov-switching Box– Cox transformed threshold GARCH process. Statistics & Probability Letters, 77(13):1428–1438, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000739.

Liu:2007:EGP

[Liu07b] Lian Liu. Estimation of generalized partially linear models with measurement error using sufficiency scores. Statistics & Probability Letters, 77(15):1580–1588, September 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001344.

Liu:2007:PLD

[Liu07c] Yan Liu. Precise large deviations for negatively associated random variables with consistently varying tails. Statistics & Probability Letters, 77(2):181–189, January 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002203.

Liu:2008:SSR

[Liu08] Kai Liu. Stationary solutions of retarded Ornstein–Uhlenbeck processes in Hilbert spaces. Statistics & Probability Letters, 78(13):1775–1783, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000394.

Liu:2009:PLD

[Liu09] Li Liu. Precise large deviations for dependent random variables with heavy tails. Statistics & Probability Letters, 79(9):1290–1298, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000546.

Lu:2001:GEM

[LJ01] Zudi Lu and Zhenyu Jiang. L_{1} geometric ergodicity of a multivariate nonlinear AR model with an ARCH term. Statistics & Probability Letters, 51(2):121–130, January 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001383.

Lekshmi:2006:APP

[LJ06] V. Seetha Lekshmi and K. K. Jose. Autoregressive processes with Pakes and geometric Pakes generalized Linnik marginals.

Statistics & Probability Letters, 76(3):318–326, February 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003019.

Low:2002:EMF

[LK02] Mark G. Low and Yung-Gyung Kang. Estimating monotone functions. Statistics & Probability Letters, 56(4):361– 367, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001791.

Louie:2004:CPC

[LK04] Mary M. Louie and Eric D. Kolaczyk. On the covariance properties of certain multiscale spatial processes. Statistics & Probability Letters, 66(4):407–416, March 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003560.

Leu:2006:MBC

[LK06] Ching-Ho Leu and Fei-Fei Kao. Modified balanced circular systematic sampling. Statistics & Probability Letters, 76 (4):373–383, February 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205003093.

Li:2007:LBS

[LK07] Zehui Li and Xinbing Kong. Life behavior of δ-shock model. Statistics & Probability Letters, 77(6):577–587, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002598.

Li:2002:DAS

[LL02] Runze Li and Dennis K. J. Lin. Data analysis in supersaturated designs. Statistics & Probability Letters, 59(2): 135–144, September 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202001402.

Liseo:2003:BIM

[LL03] Brunero Liseo and Nicola Loperfido. A Bayesian interpretation of the multivariate skew-normal distribution. Statistics & Probability Letters, 61(4):395–401, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200398X.

Li:2004:CEP

[LL04] Bo Li and Luqin Liu. The characterization of equilibrium potentials and last exit distributions for elliptic diffusion processes. Statistics & Probability Letters, 70(2):155–169, November 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002354.

Loukianova:2005:ULL

[LL05] D. Loukianova and O. Loukianov. Uniform Law of Large Numbers and consistency of estimators for Harris diffusions. Statistics & Probability Letters, 74(4):347–355, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520500194X.

Liu:2006:PAN

[LL06] Weidong Liu and Zhengyan Lin. Precise asymptotics for a new kind of complete moment convergence. Statistics & Probability Letters, 76(16):1787–1799, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001441.

Liu:2008:FRP

[LL08] Fuxiang Liu and Yanyan Liu. The failure rate properties of a bimodal mixture of normal distributions in an unequal variance case. Statistics & Probability Letters, 78(14):2006–2009, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000692.

Li:2009:DTD

[LL09a] Shuanming Li and Yi Lu. The distribution of total dividend payments in a Sparre Andersen model. Statistics & Probability Letters, 79(9):1246–1251, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000406.

Liu:2009:MBT

[LL09b] Guoqing Liu and Wenbo V. Li. Moment bounds for truncated random variables. Statistics & Probability Letters, 79(18):1951–1956, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002120.

Luo:2009:SBS

[LL09c] Jiaowan Luo and Guolie Lan. Stochastically bounded solutions and stationary solutions of stochastic differential equations in Hilbert spaces. Statistics & Probability Letters, 79(21):2260–2265, November 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002879.

Lim:2006:ILC

[LLC06] Johan Lim, Sungim Lee, and Hyungwon Choi. Information loss from censoring in rank-based procedures. Statistics & Probability Letters, 76(16):1705–1713, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001118.

Li:2000:SRA

[LLJ00] Xiaohu Li, Zehui Li, and Bing-Yi Jing. Some results about the NBUC class of life distributions. Statistics & Probability Letters, 46(3):229–237, February 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001042.

Lacruz:2000:DGM

[LLL00] Beatriz Lacruz, Pilar Lasala, and Alberto Lekuona. Dynamic graphical models and nonhomogeneous hidden Markov models. Statistics & Probability Letters, 49(4):377–385, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000705.

Lloyd:2002:ECR

[Llo02] Chris J. Lloyd. Estimation of a convex ROC curve. Statistics & Probability Letters, 59(1):99–111, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002122.

Lloyd:2005:MLS

[Llo05] Chris J. Lloyd. Monotonicity of likelihood support bounds for system failure rates. Statistics & Probability Letters, 73(2):91– 97, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000799.

Lloyd:2008:MPE

[Llo08] Chris J. Lloyd. More powerful exact tests of binary matched pairs. Statistics & Probability Letters, 78(16):2592–2596, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003532.

Laksaci:2009:GAK

[LLOS09] Ali Laksaci, Mohamed Lemdani, and Elias Ould-Saïd. A generalized L^{1}-approach for a kernel estimator of conditional quantile with functional regressors: Consistency and asymptotic normality. Statistics & Probability Letters, 79 (8):1065–1073, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005762.

Lin:2006:FAB

[LLS06] C. Devon Lin, Wilson W. Lu, and R. R. Sitter. Fast approximately balanced bootstrap without construction. Statistics & Probability Letters, 76(17):1861–1872, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001520.

Liu:2009:RCI

[LLW09] Yonghong Liu, Luoqing Li, and Chenggao Wan. The rate of convergence for increments of a Brownian motion in Hölder norm. Statistics & Probability Letters, 79(12):1463–1472, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001102.

Li:2004:STC

[LLZ04] Peng-Fei Li, Min-Qian Liu, and Run-Chu Zhang. Some theory and the construction of mixed-level supersaturated designs. Statistics & Probability Letters, 69(1):105–116, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001762.

Lee:2002:BRT

[LN02] Sangyeol Lee and Seongryong Na. On the Bickel–Rosenblatt test for first-order autoregressive models. Statistics & Probability Letters, 56(1):23–35, January 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001432.

Lalley:2003:IAC

[LN03] Steven P. Lalley and Andrew Nobel. Indistinguishability of absolutely continuous and singular distributions. Statistics & Probability Letters, 62(2):145–154, April 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000038.

Lee:2004:NTC

[LN04] Sangyeol Lee and Seongryong Na. A nonparametric test for the change of the density function in strong mixing processes. Statistics & Probability Letters, 66(1):25–34, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002906.

Lei:2009:DBB

[LN09] Pedro Lei and David Nualart. A decomposition of the bifractional Brownian motion and some applications. Statistics & Probability Letters, 79(5):619–624, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004884.

Lillo:2001:PSL

[LNS01] Rosa E. Lillo, Asok K. Nanda, and Moshe Shaked. Preservation of some likelihood ratio stochastic orders by order statistics. Statistics & Probability Letters, 51(2):111–119, January 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001371.

Lo:2005:LRT

[Lo05a] Yungtai Lo. Likelihood ratio tests of the number of components in a normal mixture with unequal variances. Statistics & Probability Letters, 71(3):225–235, March 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204003049.

Lockhart:2005:NMC

[LO05b] Richard Lockhart and Federico O’Reilly. A note on Moore’s conjecture. Statistics & Probability Letters, 74(2):212– 220, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001860.

Leorato:2009:BSG

[LO09] S. Leorato and E. Orsingher. Branching on a Sierpinski graph. Statistics & Probability Letters, 79(2):145–154, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003581.

Logan:2003:COM

[Log03] Brent R. Logan. A cone order monotone test for the one-sided multivariate testing problem. Statistics & Probability Letters, 63(3):315–323, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520300097X.

Long:2009:LSE

[Lon09] Hongwei Long. Least squares estimator for discretely observed Ornstein–Uhlenbeck processes with small Lévy noises. Statistics & Probability Letters, 79(19):2076–2085, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209002429.

Loperfido:2001:QFS

[Lop01] Nicola Loperfido. Quadratic forms of skew-normal random vectors. Statistics & Probability Letters, 54(4):381–387, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001031.

Loperfido:2002:SIS

[Lop02] Nicola Loperfido. Statistical implications of selectively reported inferential results. Statistics & Probability Letters, 56(1):13–22, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000529.

Loperfido:2008:NSE

[Lop08] Nicola Loperfido. A note on skew-elliptical distributions and linear functions of order statistics. Statistics & Probability Letters, 78(18):3184–3186, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002952.

Lemdani:2002:EAE

[LOS02] Mohamed Lemdani and Elias Ould-Saïd. Exact asymptotic L_{∞}-error of a kernel density estimator under censored data. Statistics & Probability Letters, 60(1):59–68, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520200278X.

Lemdani:2003:DKM

[LOS03] Mohamed Lemdani and Elias Ould-Saïd. L_{∞}-deficiency of the Kaplan–Meier estimator. Statistics & Probability Letters, 63(2):145–155, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000622.

Lou:2000:EDC

[Lou00]  W. Y. Wendy Lou.  The exact distribution of the

continuity of care measure NOP. Statistics & Probability Letters, 48(4):361–368, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000171.

Lou:2001:DUP

[Lou01] W. Y. Wendy Lou. The distribution of the usual provider continuity index under Markov dependence. Statistics & Probability Letters, 54(3):269–276, October 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100058X.

Lou:2003:EDT

[Lou03] W. Y. Wendy Lou. The exact distribution of the ktuple statistic for sequence homology. Statistics & Probability Letters, 61(1):51–59, January 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003371.

Louis:2005:ICP

[Lou05] P.-Y. Louis. Increasing coupling of probabilistic cellular automata. Statistics & Probability Letters, 74(1):1–13, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001501.

Leon:2003:EPS

[LP03] Carlos A. León and François Perron. Extremal properties of sums of Bernoulli random variables. Statistics & Probability Letters, 62(4):345–354, May 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000373.

Li:2009:GFT

[LP09] Deyuan Li and Liang Peng. Goodness-of-fit test for tail copulas modeled by elliptical copulas. Statistics & Probability Letters, 79(8):1097–1104, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005804.

Lavancier:2009:CFV

[LPS09] Frédéric Lavancier, Anne Philippe, and Donatas Surgailis. Covariance function of vector self-similar processes. Statistics & Probability Letters, 79(23):2415–2421, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003241.

Liang:2004:CBP

[LPY04] Jiajuan Liang, William S. Y. Pan, and Zhen-Hai Yang. Characterization-based Q–Q plots for testing multinormality. Statistics & Probability Letters, 70(3):183–190, December 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002536.

Lu:2004:NAD

[LQ04] Xuewen Lu and Yongcheng Qi. A note on asymptotic distribution of products of sums. Statistics & Probability Letters, 68(4):407–413, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001300.

Li:2007:HSA

[LQ07] Deli Li and Yongcheng Qi. Hierarchical structures associated with order functions. Statistics & Probability Letters, 77(5):525–529, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002744.

Li:2009:ILT

[LQR09] Deli Li, Yongcheng Qi, and Andrew Rosalsky. Iterated logarithm type behavior for weighted sums of i.i.d. random variables. Statistics & Probability Letters, 79(5):643–651, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208004926.

Liu:2002:CTS

[LR02] Jicheng Liu and Jiagang Ren. Comparison theorem for solutions of backward stochastic differential equations with continuous coefficient. Statistics & Probability Letters, 56(1):93–100, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520100178X.

Liu:2007:QPP

[LR07] Xu-Qing Liu and Jian-Ying Rong. Quadratic prediction problems in finite populations. Statistics & Probability Letters, 77(5):483–489, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002689.

Lebensztayn:2008:DPM

[LR08a] E. Lebensztayn and P. M. Rodríguez. The disk-percolation model on graphs. Statistics & Probability Letters, 78(14): 2130–2136, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000874.

Lii:2008:PSS

[LR08b] K. S. Lii and M. Rosenblatt. Prolate spheroidal spectral estimates. Statistics & Probability Letters, 78(11):1339–1348, August 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002563.

Ludkovski:2008:CPO

[LR08c] Michael Ludkovski and Ludger Rüschendorf. On comonotonicity of Pareto optimal risk sharing. Statistics & Probability Letters, 78(10):1181–1188, August 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003902.

Leite:2000:BAE

[LRM00] José G. Leite, Josemar Rodrigues, and Luis A. Milan. A Bayesian analysis for estimating the number of species in a population using nonhomogeneous Poisson process. Statistics & Probability Letters, 48(2):153–161, June 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001984.

Lysenko:2009:MEG

[LRW09] Natalia Lysenko, Parthanil Roy, and Rolf Waeber. Multivariate extremes of generalized skew-normal distributions. Statistics & Probability Letters, 79(4):525–533, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208004616.

Lee:2000:GEM

[LS00a] Oesook Lee and Dong Wan Shin. On geometric ergodicity of the MTAR process. Statistics & Probability Letters, 48(3):229– 237, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299002084.

Li:2000:SBG

[LS00b] Gang Li and Yanqing Sun. A simulation-based goodnessof-fit test for survival data. Statistics & Probability Letters, 47(4):403–410, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001868.

Liptser:2000:DPB

[LS00c] R. Liptser and V. Spokoiny. Deviation probability bound for martingales with applications to statistical estimation. Statistics & Probability Letters, 46(4):347–357, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001212.

Lopez:2000:SCG

[LS00d] F. Javier López and Gerardo Sanz. Stochastic comparisons for general probabilistic cellular automata. Statistics & Probability Letters, 46(4):401–410, February 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001297.

Lifshits:2001:LDP

[LS01] M. A. Lifshits and E. S. Stankevich. On the large deviation principle for the almost sure CLT. Statistics & Probability Letters, 51(3):263–267, February 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001541.

Lee:2002:GTE

[LS02a] Sungchul Lee and Zhonggen Su. Gaussian tail for empirical distributions of MST on random graphs. Statistics & Probability Letters, 58(4):363–368, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200144X.

Lee:2002:SMI

[LS02b] Sungchul Lee and Zhonggen Su. The symmetry in the martingale inequality. Statistics & Probability Letters, 56(1):83–91, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001778.

Lanzinger:2004:RBK

[LS04a] H. Lanzinger and U. Stadtmüller. Refined Baum–Katz laws for weighted sums of iid random variables. Statistics & Probability Letters, 69(3):357–368, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001877.

Li:2004:OTT

[LS04b] Xiaohu Li and Moshe Shaked. The observed total time on test and the observed excess wealth. Statistics & Probability Letters, 68(3):247–258, July 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000975.

Lee:2005:SMP

[LS05a] O. Lee and D. W. Shin. On stationarity and β-mixing property of certain nonlinear GARCH(p, q) models. Statistics & Probability Letters, 73(1):25–35, June 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000763.

Letac:2005:ESD

[LS05b] G. Letac and V. Seshadri. Exponential stopping and drifted stable processes. Statistics & Probability Letters, 72(2): 137–143, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000234.

Leonenko:2006:WES

[LS06a] N. N. Leonenko and L. M. Sakhno. On the Whittle estimators for some classes of continuous-parameter random processes and fields. Statistics & Probability Letters, 76(8):781–795, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003883.

Li:2006:CPA

[LS06b] Deli Li and Aurel Spataru. A converse to precise asymptotic results. Statistics & Probability Letters, 76(5):503–506, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003317.

Lee:2008:TPC

[LS08a] Sangyeol Lee and Junmo Song. Test for parameter change in ARMA models with GARCH innovations. Statistics & Probability Letters, 78(13):1990–1998, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000722.

Liu:2008:SBS

[LS08b] Rong Liu and Yong Shi. Sampling based succinct matrix approximation. Statistics & Probability Letters, 78(9):

1138–1147, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003859.

Loquin:2008:HDE

[LS08c] Kevin Loquin and Olivier Strauss. Histogram density estimators based upon a fuzzy partition. Statistics & Probability Letters, 78(13):1863–1868, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000527.

Leonenko:2000:ENA

[LSEB00] Nikolai N. Leonenko, Michail M. Sharapov, and Ahmed H. El-Bassiouny. On the exactness of normal approximation of LSE of regression coefficient of long-memory random fields. Statistics & Probability Letters, 48(2):121–130, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001947.

Liu:2000:NMC

[LSL00] Zhenjun Liu, Thanasis Stengos, and Qi Li. Nonparametric model check based on local polynomial fitting. Statistics & Probability Letters, 48(4):327–334, July 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000122.

Leiva:2008:NTP

[LSSG08] Víctor Leiva, Antonio Sanhueza, Andrés Silva, and Manuel Galea. A new three-parameter extension of the inverse Gaussian distribution. Statistics & Probability Letters, 78(11): 1266–1273, August 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520700404X.

Leon:2000:CDS

[LT00a] Jorge A. León and Constantin Tudor. Chaos decomposition of stochastic bilinear equations with drift in the first Poisson– Itô chaos. Statistics & Probability Letters, 48(1):11–22, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001856.

Liu:2000:NAS

[LT00b] Kai Liu and Aubrey Truman. A note on almost sure exponential stability for stochastic partial functional differential equations. Statistics & Probability Letters, 50(3):273–278, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001036.

Landsman:2006:SOB

[LT06] Zinoviy Landsman and Andreas Tsanakas. Stochastic ordering of bivariate elliptical distributions. Statistics & Probability Letters, 76(5):488–494, March 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003305.

Lin:2008:PBE

[LT08] Lu Lin and Lin Tan. Proper Bayesian estimating equation based on Hilbert space method. Statistics & Probability Letters, 78(9):1119–1127, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003823.

Lai:2004:GPG

[LTC04] Yi-Hsuan Lai, Peter Thompson, and Lin-An Chen. Generalized and pseudo-generalized trimmed means for the linear regression with AR(1) error model. Statistics & Probability Letters, 67(3):203–211, April 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002669.

Li:2001:ICM

[LTS01] Chin-Shang Li, Jeremy M. G. Taylor, and Judy P. Sy. Identifiability of cure models. Statistics & Probability Letters, 54(4):389–395, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001055.

Lu:2008:SSN

[LTY08] Zudi Lu, Dag Tjøstheim, and Qiwei Yao. Spatial smoothing, Nugget effect and infill asymptotics. Statistics & Probability Letters, 78(18):3145–3151, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002848.

Lu:2006:DSA

[Lu06] Xiaojing Lu. On the distribution of a statistic arising in exact simultaneous confidence bands. Statistics & Probability Letters, 76(7):743–747, April 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003809.

Luceno:2000:MCD

[Luc00] Alberto Luceño. Minimum cost dead band adjustment schemes under tool-wear effects and delayed dynamics. Statistics & Probability Letters, 50(2):165–178, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000095X.

Ludena:2009:CIS

[Lud09] Carenne Ludeña. Confidence intervals for the scaling function of multifractal random walks. Statistics & Probability Letters, 79(9):1186–1193, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000224.

Leipus:2003:LMS

[LV03] Remigijus Leipus and Marie-Claude Viano. Long memory and stochastic trend. Statistics & Probability Letters, 61 (2):177–190, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202003474.

Lowe:2005:SCH

[LV05] Matthias Löwe and Franck Vermet. The storage capacity of the Hopfield model and moderate deviations. Statistics & Probability Letters, 75(4):237–248, December 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002294.

Lowe:2007:CSP

[LV07] Matthias Löwe and Franck Vermet. The capacity of qstate Potts neural networks with parallel retrieval dynamics. Statistics & Probability Letters, 77(14):1505–1514, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001228.

Laukaitis:2009:EAO

[LVL09] Algirdas Laukaitis, Olegas Vasilecas, and Ricardas Laukaitis. Estimation of the autoregressive operator by wavelet packets. Statistics & Probability Letters, 79(1):38–43, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003337.

Liu:2003:MAR

[LW03] Wen Liu and Liying Wang. The Markov approximation of the random fields on Cayley trees and a class of small deviation theorems. Statistics & Probability Letters, 63(2):113–121, June 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000580.

Li:2008:ASC

[LW08a] Yun-Xia Li and Jian-Feng Wang. An almost sure Central Limit Theorem for products of sums under association. Statistics & Probability Letters, 78(4):367–375, March 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002490.

Liu:2008:EDP

[LW08b] Guoxin Liu and Ying Wang. On the expected discounted penalty function for the continuous-time compound binomial risk model. Statistics & Probability Letters, 78(15):2446–2455, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800148X.

Li:2007:DFP

[LWD07] Zhigang Li, Rong Wu, and Yonghong Du. The distribution of the first β point in the classical risk model with interest. Statistics & Probability Letters, 77(9):873–880, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000193.

Li:2008:BTC

[LWT08] Xiao Li, Yaohua Wu, and Dongsheng Tu. A Bartlett type correction for Wald test in Cox regression model. Statistics & Probability Letters, 78(16):2614–2622, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003866.

Li:2009:NSC

[LWW09] Zhonghua Li, Zhaojun Wang, and Zhang Wu. Necessary and sufficient conditions for non-interaction of a pair of onesided EWMA schemes with reflecting boundaries. Statistics & Probability Letters, 79(3):368–374, February 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004215.

Lai:2000:NFP

[LX00a] C. D. Lai and M. Xie. A new family of positive quadrant dependent bivariate distributions. Statistics & Probability Letters, 46(4):359–364, February 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001224.

Li:2000:BRU

[LX00b] Hongyi Li and Zhijie Xiao. On bootstrapping regressions with unit root processes. Statistics & Probability Letters, 48(3):261– 267, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000055.

Luan:2001:WIJ

[LX01] Yihui Luan and Zhongjie Xie. The wavelet identification for jump points of derivative in regression model. Statistics & Probability Letters, 53(2):167–180, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000700.

Lepeltier:2005:PMR

[LX05] J.-P. Lepeltier and M. Xu. Penalization method for reflected backward stochastic differential equations with one r.c.l.l. barrier. Statistics & Probability Letters, 75(1):58–66, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002105.

Lin:2008:HTI

[LX08] Zhengyan Lin and Yanbiao Xiang. A hypothesis test for independence of sets of variates in high dimensions. Statistics & Probability Letters, 78(17):2939–2946, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800240X.

Lu:2009:RPC

[LXZS09] Zhaoyang Lu, Wei Xu, Yan Zhang, and Yingling Sun. On the ruin probability for the Cox correlated risk model perturbed by diffusion. Statistics & Probability Letters, 79(3):381–389, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004239.

Liu:2003:CSL

[LY03] Wen Liu and Weiguo Yang. A class of strong limit theorems for the sequences of arbitrary random variables. Statistics & Probability Letters, 64(2):121–131, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001044.

Lee:2005:PBN

[LY05] Stephen M. S. Lee and G. Alastair Young. Parametric bootstrapping with nuisance parameters. Statistics & Probability Letters, 71(2):143–153, February 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002895.

Lyhagen:2005:ECM

[Lyh05] Johan Lyhagen. The exact covariance matrix of dynamic models with latent variables. Statistics & Probability Letters, 75(2):133–139, November 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500221X.

Li:2008:CUA

[LYZ08] Yongming Li, Shanchao Yang, and Yong Zhou. Consistency and uniformly asymptotic normality of wavelet estimator in regression model with associated samples. Statistics & Probability Letters, 78(17):2947–2956, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002411.

Lin:2001:BEE

[LZ01] Lu Lin and Runchu Zhang. Blockwise empirical Euclidean likelihood for weakly dependent processes. Statistics & Probability Letters, 53(2):143–152, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000669.

Li:2002:BNP

[LZ02a] Xuefeng Li and Linda H. Zhao. Bayesian nonparametric point estimation under a conjugate prior. Statistics & Probability Letters, 58(1):23–30, May 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000974.

Lin:2002:PQL

[LZ02b] Lu Lin and Runchu Zhang. Profile quasi-likelihood. Statistics & Probability Letters, 56(2):147–154, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001687.

Lu:2005:LRT

[LZ05] Nelson Lu and Dale L. Zimmerman. The likelihood ratio test for a separable covariance matrix. Statistics & Probability Letters, 73(4):449–457, July 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001495.

Li:2006:FLT

[LZ06] Zenghu Li and Mei Zhang. Fluctuation limit theorems of immigration superprocesses with small branching. Statistics & Probability Letters, 76(4):401–411, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003123.

Lin:2007:SPM

[LZ07] Zhengyan Lin and Jing Zheng. Some properties of a multifractional Brownian motion. Statistics & Probability Letters, 77(7):687–692, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206003373.

Li:2005:SRM

[LZC05] Zehui Li, Jinxia Zhu, and Feng Chen. Study of a risk model based on the entrance process. Statistics & Probability Letters, 72(1):1–10, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002627.

Li:2009:ENC

[LZD09] Qizhai Li, Junjian Zhang, and Shuai Dai. On estimating the non-centrality parameter of a chi-squared distribution. Statistics & Probability Letters, 79(1):98–104, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003507.

Liu:2008:ELT

[LZZ08] Yukun Liu, Changliang Zou, and Runchu Zhang. Empirical likelihood for the two-sample mean problem. Statistics & Probability Letters, 78(5):548–556, April 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002866.

Ma:2003:NCF

[Ma03] Chunsheng Ma. Nonstationary covariance functions that model space-time interactions. Statistics & Probability Letters, 61(4):411–419, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202004017.

Ma:2009:LTT

[Ma09] Chunhua Ma. A limit theorem of two-type Galton–Watson branching processes with immigration. Statistics & Probability Letters, 79(15):1710–1716, August 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900159X.

Mabrouk:2008:HOM

[Mab08] Anouar Ben Mabrouk. A higher order multifractal formalism. Statistics & Probability Letters, 78(12):1412–1421, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207004233.

Macci:2008:LDT

[Mac08] Claudio Macci. Large deviations for the time-integrated negative parts of some processes. Statistics & Probability Letters, 78(1):75–83, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001976.

Macci:2009:CLD

[Mac09] Claudio Macci. Convergence of large deviation rates based on a link between wave governed random motions and ruin processes. Statistics & Probability Letters, 79(2):255–263, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003854.

Madi:2002:IEE

[Mad02] Mohamed T. Madi. On the invariant estimation of an exponential scale using doubly censored data. Statistics & Probability Letters, 56(1):77–82, January 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001742.

Madi:2008:IEE

[Mad08] Mohamed T. Madi. Improved estimation of an exponential scale ratio based on records. Statistics & Probability Letters, 78(2):165–172, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002088.

Mahanti:2004:ENR

[Mah04] Mina Ketan Mahanti. Expected number of real zeros of random hyperbolic polynomial. Statistics & Probability Letters, 70(1):11–18, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001725.

Maity:2008:EEP

[Mai08] Arnab Maity. Efficient estimation of population quantiles in general semiparametric regression models. Statistics & Probability Letters, 78(16):2744–2750, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001958.

Majumdar:2005:CRL

[Maj05] Suman Majumdar. On convergence of random linear functionals. Statistics & Probability Letters, 75(4):249–255, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002385.

Makasu:2008:MET

[Mak08a] Cloud Makasu. On mean exit time from a curvilinear domain. Statistics & Probability Letters, 78(17):2859–2863, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002204.

Makhoukhi:2008:APF

[Mak08b] Mohammed Boukili Makhoukhi. An approximation for the power function of a non-parametric test of fit. Statistics & Probability Letters, 78(8):1034–1042, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700363X.

Maki:2008:SPN

[Mak08c] Daiki Maki. The size performance of a nonparametric unit root test under a variance shift. Statistics & Probability Letters, 78(6):743–748, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003276.

Makasu:2009:EPI

[Mak09] Cloud Makasu. Exit probability for an integrated geometric Brownian motion. Statistics & Probability Letters, 79 (11):1363–1365, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000728.

Maltz:2001:CLT

[Mal01] Alberto L. Maltz. A Central Limit Theorem for nonuniform ϕ-mixing random fields with infinite variance. Statistics & Probability Letters, 51(4):351–359, February 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001759.

Mamon:2002:TVM

[Mam02] Rogemar S. Mamon. A time-varying Markov chain model of term structure. Statistics & Probability Letters, 60(3):309–312, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003139.

Mandal:2005:ASA

[Man05a] Abhyuday Mandal. An approach for studying aliasing relations of mixed fractional factorials based on product arrays. Statistics & Probability Letters, 75(3):203–210, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002257.

Mansuy:2005:ISG

[Man05b] Roger Mansuy. An interpretation and some generalizations of the Anderson–Darling statistics in terms of squared Bessel bridges. Statistics & Probability Letters, 72(2):171–177, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000271.

Mancini:2008:LDP

[Man08] Cecilia Mancini. Large deviation principle for an estimator of the diffusion coefficient in a jump-diffusion process. Statistics & Probability Letters, 78(7):869–879, May 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002933.

Markiewicz:2001:DSP

[Mar01a] Augustyn Markiewicz. On dependence structures preserving optimality. Statistics & Probability Letters, 53(4):415–419, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000645.

Marsh:2001:EEG

[Mar01b] Patrick Marsh. Edgeworth expansions in Gaussian autoregression. Statistics & Probability Letters, 54(3):233–241, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002091.

Martinsek:2001:SEM

[Mar01c] Adam T. Martinsek. Sequential estimation of the mean in a random coefficient autoregressive model with beta marginals. Statistics & Probability Letters, 51(1):53–61, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001425.

Martinsek:2002:EMM

[Mar02] Adam T. Martinsek. Estimation of the maximum and minimum in a model for bounded, dependent data. Statistics & Probability Letters, 56(4):381–393, February 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001985.

Marichal:2006:CDF

[Mar06] Jean-Luc Marichal. Cumulative distribution functions and moments of lattice polynomials. Statistics & Probability Letters, 76(12):1273–1279, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000198.

Maruyama:2007:SMS

[Mar07] Yosihito Maruyama. On Srivastava’s multivariate sample skewness and kurtosis under non-normality. Statistics & Probability Letters, 77(3):335–342, February 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002410.

Mas:2002:RWC

[Mas02a] André Mas. Rates of weak convergence for images of measures by families of mappings. Statistics & Probability Letters, 56(1):7–12, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000384.

Mashayekhi:2002:CEM

[Mas02b] Mostafa Mashayekhi. Compound estimation of a monotone sequence. Statistics & Probability Letters, 60(1):7–15, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002390.

Masry:2002:MPD

[Mas02c] Elias Masry. Multivariate probability density estimation for associated processes: strong consistency and rates. Statistics & Probability Letters, 58(2):205–219, June 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001050.

Mas:2006:SCC

[Mas06a] André Mas. A sufficient condition for the CLT in the space of nuclear operators — application to covariance of random functions. Statistics & Probability Letters, 76(14):1503–1509, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520600085X.

Masmoudi:2006:CNE

[Mas06b] Afif Masmoudi. Conditional natural exponential families. Statistics & Probability Letters, 76(17):1882–1888, November 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001544.

Mashayekhi:2008:CCM

[Mas08] Mostafa Mashayekhi. On construction of consistent mixing distributions for compound decisions. Statistics & Probability Letters, 78(16):2644–2646, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001776.

Mathew:2003:LPI

[Mat03] George Mathew. Limit points of independent copies of sample maxima. Statistics & Probability Letters, 65(2):103– 109, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002244.

Matula:2005:ASL

[Mat05] Przemysław Matula. On almost sure limit theorems for positively dependent random variables. Statistics & Probability Letters, 74(1):59–66, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001598.

Mavrikiou:2007:KIP

[Mav07] Petroula M. Mavrikiou. Kolmogorov inequalities for the partial sum of independent Bernoulli random variables. Statistics & Probability Letters, 77(11):1117–1122, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000636.

Mavrikiou:2008:KIW

[Mav08] Petroula M. Mavrikiou. A Kolmogorov inequality for weighted U-statistics. Statistics & Probability Letters, 78(18):3294– 3297, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208003179.

Mazliak:2000:NWV

[Maz00] Laurent Mazliak. A note on weak viability for controlled diffusion. Statistics & Probability Letters, 49(4):331–336, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000651.

Mourid:2006:SEO

[MB06] Tahar Mourid and Nawel Bensmain. Sieves estimator of the operator of a functional autoregressive process. Statistics & Probability Letters, 76(1):93–108, January 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002737.

Marrelec:2008:CIB

[MB08] Guillaume Marrelec and Habib Benali. Conditional independence between two variables given any conditioning subset implies block diagonal covariance matrix for multivariate Gaussian distributions. Statistics & Probability Letters, 78(13):1922–1928, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800059X.

Maronna:2000:IBR

[MBY00] Ricardo A. Maronna, Matías Salibian Barrera, and Víctor J. Yohai. Improving bias-robustness of regression estimates through projections. Statistics & Probability Letters, 47(2): 149–158, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001510.

Melnykov:2007:CBS

[MC07a] Igor Melnykov and John T. Chen. A connection between self-normalized products and stable laws. Statistics & Probability Letters, 77(17):1662–1665, November 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001496.

Moon:2007:APP

[MC07b] Hee-Jin Moon and Yong-Kab Choi. Asymptotic properties for partial sum processes of a Gaussian random field. Statistics & Probability Letters, 77(1):9–18, January 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001891.

Mudholkar:2009:DV

[MC09] Govind S. Mudholkar and Yogendra P. Chaubey. On defining p-values. Statistics & Probability Letters, 79(18):1963– 1971, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209002144.

McGill:2006:AEB

[McG06] Paul McGill. An asymptotic estimate for Brownian motion with drift. Statistics & Probability Letters, 76(11):1164–1169, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205004700.

Molina:2007:CCB

[MdPR07] Manuel Molina, Inés M. del Puerto, and Alfonso Ramos. A class of controlled bisexual branching processes with mating depending on the number of progenitor couples. Statistics & Probability Letters, 77(18):1737–1743, December 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001599.

Meister:2007:OCR

[Mei07] Alexander Meister. Optimal convergence rates for density estimation from grouped data. Statistics & Probability Letters, 77(11):1091–1097, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000478.

Meinshausen:2008:NLG

[Mei08] Nicolai Meinshausen. A note on the Lasso for Gaussian graphical model selection. Statistics & Probability Letters, 78(7):880– 884, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002945.

Meister:2009:TLM

[Mei09]  Alexander Meister.  On testing for local monotonic-

ity in deconvolution problems. Statistics & Probability Letters, 79(3):312–319, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004033.

Merkle:2005:JIM

[Mer05] Milan Merkle. Jensen’s inequality for medians. Statistics & Probability Letters, 71(3):277–281, March 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204003104.

Martins:2005:MED

[MF05a] A. P. Martins and H. Ferreira. Measuring the extremal dependence. Statistics & Probability Letters, 73(2):99–103, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000830.

Medina:2005:CCR

[MF05b] Juan Miguel Medina and Bruno Cernuschi Frías. On the a.s. convergence of certain random series to a fractional random field in D′(R^{⌈}). Statistics & Probability Letters, 74(1):39–49, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001550.

Mukhopadhyay:2005:SBP

[MGB05] Nitai Mukhopadhyay, Jayanta K. Ghosh, and James O. Berger. Some Bayesian predictive approaches to model selection. Statistics & Probability Letters, 73(4):369–379, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001410.

Mandal:2008:CCB

[MGSB08] S. Mandal, D. K. Ghosh, R. K. Sharma, and S. C. Bagui. A complete class of balanced incomplete block designs (7, 35, 15, 3, 5) with repeated blocks. Statistics & Probability Letters, 78(18):3338–3343, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003246.

Metzler:2002:PMG

[MGW02] Dirk Metzler, Steffen Grossmann, and Anton Wakolbinger. A Poisson model for gapped local alignments. Statistics & Probability Letters, 60(1):91–100, November 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002845.

Mi:2006:LAS

[Mi06] Jie Mi. Limiting availability of system with non-identical lifetime distributions and non-identical repair time distributions. Statistics & Probability Letters, 76(7):729–736, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003822.

Middleton:2008:BRE

[Mid08] Joel A. Middleton. Bias of the regression estimator for experiments using clustered random assignment. Statistics & Probability Letters, 78(16):2654–2659, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001806.

Milionis:2007:ECM

[Mil07] Alexandros E. Milionis. Efficient capital markets: a statistical definition and comments. Statistics & Probability Letters, 77(6):607–613, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002938.

Mimoto:2008:CDS

[Mim08] Nao Mimoto. Convergence in distribution for the sup-norm of a kernel density estimator for GARCH innovations. Statistics & Probability Letters, 78(7):915–923, May 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003410.

Minkevicius:2009:AVW

[Min09] S. Minkevicius. Analysis of the virtual waiting time in multiphase queues. Statistics & Probability Letters, 79(2):173–176, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003635.

Mira:2001:EFS

[Mir01] Antonietta Mira. Efficiency of finite state space Monte Carlo Markov chains. Statistics & Probability Letters, 54(4):405–411, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001158.

Mirakhmedov:2005:LER

[Mir05] Sherzod A. Mirakhmedov. Lower estimation of the remainder term in the CLT for a sum of the functions of k-spacings. Statistics & Probability Letters, 73(4):411–424, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001458.

Mirakhmedov:2007:ANA

[Mir07] Sherzod M. Mirakhmedov. Asymptotic normality associated with generalized occupancy problems. Statistics & Probability Letters, 77(15):1549–1558, September 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001289.

Maejima:2000:DSS

[MiSW00] Makoto Maejima, Ken iti Sato, and Toshiro Watanabe. Distributions of selfsimilar and semi-selfsimilar processes with independent increments. Statistics & Probability Letters, 47(4):395–401, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001844.

Mizushima:2000:MTS

[Miz00] Takamasa Mizushima. Multisample tests for scale based on kernel density estimation. Statistics & Probability Letters, 49(1):81–91, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000353.

Manesh:2008:LRO

[MK08a] Sirous Fathi Manesh and Baha-Eldin Khaledi. On the likelihood ratio order for convolutions of independent generalized Rayleigh random variables. Statistics & Probability Letters, 78(18):3139–3144, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002824.

Marichal:2008:DFL

[MK08b] Jean-Luc Marichal and Ivan Kojadinovic. Distribution functions of linear combinations of lattice polynomials from the uniform distribution. Statistics & Probability Letters, 78(8): 985–991, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003513.

Milienos:2008:LBR

[MK08c] F. S. Milienos and M. V. Koutras. A lower bound for the reliability function of multiple failure mode systems. Statistics & Probability Letters, 78(12):1639–1648, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000229.

McGrath:2003:ALD

[ML03] Richard N. McGrath and Dennis K. J. Lin. Analyzing location and dispersion in unreplicated fractional factorials. Statistics & Probability Letters, 65(4):369–377, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002840.

Meng:2009:MIL

[ML09a] Yanjiao Meng and Zhengyan Lin. Maximal inequalities and laws of large numbers for L_{q}-mixingale arrays. Statistics & Probability Letters, 79(13):1539–1547, July 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001187.

Meng:2009:WLL

[ML09b] Yanjiao Meng and Zhengyan Lin. On the weak laws of large numbers for arrays of random variables. Statistics & Probability Letters, 79(23):2405–2414, December 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900323X.

Mladenovic:2009:MPS

[Mla09] Pavle Mladenović. Maximum of a partial sample in the uniform AR(1) processes. Statistics & Probability Letters, 79(11):1414–1420, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001072.

Madi:2000:BIT

[MLT00] Mohamed T. Madi, Thomas Leonard, and Kam-Wah Tsui. Bayes inference for treatment effects with uncertain order constraints. Statistics & Probability Letters, 49(3):277– 283, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200000584.

Mizera:2002:BPC

[MM02] Ivan Mizera and Christine H. Müller. Breakdown points of Cauchy regression-scale estimators. Statistics & Probability Letters, 57(1):79–89, March 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000573.

Maejima:2007:CSS

[MM07] Makoto Maejima and Manabu Miura. A characterization of subclasses of semi-selfdecomposable distributions by stochastic integral representations. Statistics & Probability Letters, 77(8):838–842, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000181.

Mohammadi:2009:BLP

[MM09] Mohammad Mohammadi and Adel Mohammadpour. Best linear prediction for α-stable random processes. Statistics & Probability Letters, 79(21):2266–2272, November 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002880.

Mezzetti:2007:ARU

[MMB07] Maura Mezzetti, Pietro Muliere, and Paolo Bulla. An application of reinforced urn processes to determining maximum tolerated dose. Statistics & Probability Letters, 77(7):740–747, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206003452.

Memin:2001:IMW

[MMV01] Jean Mémin, Yulia Mishura, and Esko Valkeila. Inequalities for the moments of Wiener integrals with respect to a fractional Brownian motion. Statistics & Probability Letters, 51(2):197–206, January 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001577.

Mitov:2009:CRI

[MMY09] Georgi K. Mitov, Kosto V. Mitov, and Nikolay M. Yanev. Critical randomly indexed branching processes. Statistics & Probability Letters, 79(13):1512–1521, July 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900114X.

Milstein:2001:MLE

[MN01a] G. N. Milstein and M. Nussbaum. Maximum likelihood estimation of a nonparametric signal in white noise by optimal control. Statistics & Probability Letters, 55(2):193– 203, November 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001237.

Moret:2001:EIT

[MN01b] Sílvia Moret and David Nualart. Exponential inequalities for two-parameter martingales. Statistics & Probability Letters, 54(1):13–19, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002455.

Mi:2003:IAS

[MN03] Jie Mi and Arlene Naranjo. Inferences about the scale parameter of the gamma distribution based on data mixed from censoring and grouping. Statistics & Probability Letters, 62(3):229–243, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000051.

Mishura:2004:WSS

[MN04] Yu. Mishura and D. Nualart. Weak solutions for stochastic differential equations with additive fractional noise. Statistics & Probability Letters, 70(4):253–261, December 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002640.

Morel:2008:REP

[MN08] Jorge G. Morel and Nagaraj K. Neerchal. Ratio estimation via Poisson regression and generalized estimating equations. Statistics & Probability Letters, 78(14):2188–2193, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208001090.

Magome:2009:ARG

[MN09] Masahisa Magome and Tetsuo Nakagawa. The asymptotic relative generation distribution of the closest common ancestor of the multitype Galton–Watson tree conditioned on non-extinction. Statistics & Probability Letters, 79(18):1957–1962, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002132.

Mnatsakanov:2008:HMPa

[Mna08a] Robert M. Mnatsakanov. Hausdorff moment problem: Reconstruction of distributions. Statistics & Probability Letters, 78(12):1612–1618, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000187.

Mnatsakanov:2008:HMPb

[Mna08b] Robert M. Mnatsakanov. Hausdorff moment problem: Reconstruction of probability density functions. Statistics & Probability Letters, 78(13):1869–1877, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000539.

Mandrekar:2009:IMS

[MNN09] V. Mandrekar and U. V. Naik-Nimbalkar. Identification of a Markovian system with observations corrupted by a fractional Brownian motion. Statistics & Probability Letters, 79(7):965– 968, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005531.

Montanero:2003:SIU

[MNOP03] Jesús Montanero, Agustín G. Nogales, José A. Oyola, and Paloma Pérez. Strong invariance under a Bayesian point of view. Statistics & Probability Letters, 61(2):155–162, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003450.

Meerschaert:2009:CCT

[MNX09] Mark M. Meerschaert, Erkan Nane, and Yimin Xiao. Correlated continuous time random walks. Statistics & Probability Letters, 79(9):1194–1202, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000236.

Majerek:2005:UIR

[MNZ05] Dariusz Majerek, Wioletta Nowak, and Wiesław Zieba. On uniform integrability of random variables. Statistics & Probability Letters, 74(3):272–280, October 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001847.

Mangalam:2008:CNM

[MNZ08] Vasudevan Mangalam, Gopalan M. Nair, and Yun Zhao. On computation of NPMLE for middle-censored data. Statistics & Probability Letters, 78(12):1452–1458, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004294.

Modarres:2003:EBS

[Mod03] Reza Modarres. Estimation of a bivariate symmetric distribution function. Statistics & Probability Letters, 63(1):25–34, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202004297.

Mojirsheibani:2000:KBC

[Moj00] Majid Mojirsheibani. A kernel-based combined classification rule. Statistics & Probability Letters, 48(4):411–419, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000249.

Mojirsheibani:2001:CST

[Moj01a]  Majid Mojirsheibani.  Classifier selection from a to-

tally bounded class of functions. Statistics & Probability Letters, 52(4):391–400, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000050.

Mojirsheibani:2001:GCT

[Moj01b] Majid Mojirsheibani. The Glivenko–Cantelli theorem based on data with randomly imputed missing values. Statistics & Probability Letters, 55(4):385–396, December 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001420.

Mehrotra:2007:SLD

[MOM07] Kishan G. Mehrotra, Necati E. Ozgencil, and Nancy Mc-Cracken. Squeezing the last drop: Cluster-based classification algorithm. Statistics & Probability Letters, 77(12):1288–1299, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000922.

Monnez:2006:ASC

[Mon06] Jean-Marie Monnez. Almost sure convergence of stochastic gradient processes with matrix step sizes. Statistics & Probability Letters, 76(5):531–536, March 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003433.

Monoyios:2007:MEM

[Mon07] Michael Monoyios. The minimal entropy measure and an Esscher transform in an incomplete market model. Statistics & Probability Letters, 77(11):1070–1076, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000430.

Monnez:2008:SAF

[Mon08] Jean-Marie Monnez. Stochastic approximation of the factors of a generalized canonical correlation analysis. Statistics & Probability Letters, 78(14):2210–2216, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001120.

Moon:2008:FCL

[Moo08] H. J. Moon. The functional CLT for linear processes generated by mixing random variables with infinite variance. Statistics & Probability Letters, 78(14):2095–2101, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000825.

Morris:2002:BBW

[Mor02] Jeffrey S. Morris. The BLUPs are not “best” when it comes to bootstrapping. Statistics & Probability Letters, 56 (4):425–430, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520200041X.

Mora:2005:CDF

[Mor05] Juan Mora. Comparing distribution functions of errors in linear models: a nonparametric approach. Statistics & Probability Letters, 73(4):425–432, July 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500146X.

Mori:2009:DDD

[Mór09]  Tamás F. Móri.  Deviation of discrete distributions-

positive and negative results. Statistics & Probability Letters, 79(8):1089–1096, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005798.

MacEachern:2000:SGS

[MP00] Steven N. MacEachern and Mario Peruggia. Subsampling the Gibbs sampler: variance reduction. Statistics & Probability Letters, 47(1):91–98, March 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900142X.

Macci:2001:BFN

[MP01a] Claudio Macci and Silvia Polettini. Bayes factor for non-dominated statistical models. Statistics & Probability Letters, 53(1):79–89, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000359.

Marinucci:2001:ALR

[MP01b] D. Marinucci and S. Poghosyan. Asymptotics for linear random fields. Statistics & Probability Letters, 51(2):131–141, January 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001395.

Marchand:2002:MEB

[MP02] Éric Marchand and François Perron. On the minimax estimator of a bounded normal mean. Statistics & Probability Letters, 58(4):327–333, July 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202000895.

Marchand:2006:MEB

[MP06] Éric Marchand and Ahmad Parsian. Minimax estimation of a bounded parameter of a discrete distribution. Statistics & Probability Letters, 76(6):547–554, March 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003445.

McMurry:2008:BCI

[MP08a] Timothy L. McMurry and Dimitris N. Politis. Bootstrap confidence intervals in nonparametric regression with built-in bias correction. Statistics & Probability Letters, 78(15):2463–2469, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001508.

Mladenovic:2008:EER

[MP08b] Pavle Mladenović and Vladimir Piterbarg. On estimation of the exponent of regular variation using a sample with missing observations. Statistics & Probability Letters, 78(4):327– 335, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002453.

Maejima:2007:CRM

[MPA07] Makoto Maejima and Víctor Pérez-Abreu. A class of random matrices with infinitely divisible determinants. Statistics & Probability Letters, 77(2):166–168, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002173.

Moler:2006:ADC

[MPS06] José A. Moler, Fernando Plo, and Miguel San Miguel. An adaptive design for clinical trials with non-dichotomous response and prognostic factors. Statistics & Probability Letters, 76(17):1940–1946, November 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001635.

Marzio:2009:LPR

[MPT09] Marco Di Marzio, Agnese Panzera, and Charles C. Taylor. Local polynomial regression for circular predictors. Statistics & Probability Letters, 79(19):2066–2075, October 1, 2009.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002417.

Menshikov:2001:MPS

[MPV01] M. V. Menshikov, S. Yu. Popov, and M. Vachkovskaia. Multiscale percolation on k-symmetric mosaic. Statistics & Probability Letters, 52(1):79–84, March 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000225X.

Mitov:2003:EGV

[MPY03] Kosto V. Mitov, Anthony G. Pakes, and George P. Yanev. Extremes of geometric variables with applications to branching processes. Statistics & Probability Letters, 65(4):379–388, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002864.

Montanes:2000:SBU

[MR00] Antonio Montañés and Marcelo Reyes. Structural breaks, unit roots and methods for removing the autocorrelation pattern. Statistics & Probability Letters, 48(4):401–409, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000237.

Mazliak:2003:EPV

[MR03] Laurent Mazliak and Catherine Rainer. Exact and possible viability for controlled diffusions. Statistics & Probability Letters, 62(2):155–161, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000063.

Mattner:2008:MPC

[MR08a] Lutz Mattner and Bero Roos. Maximal probabilities of convolution powers of discrete uniform distributions. Statistics & Probability Letters, 78(17):2992–2996, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002551.

Muller:2008:MDA

[MR08b] Samuel Müller and Kaspar Rufibach. On the max-domain of attraction of distributions with log-concave densities. Statistics & Probability Letters, 78(12):1440–1444, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004270.

Malinovsky:2009:SOA

[MR09] Yaakov Malinovsky and Yosef Rinott. On stochastic orders of absolute value of order statistics in symmetric distributions. Statistics & Probability Letters, 79(19):2086–2091, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209002430.

Mallick:2008:BPS

[MRK08] Madhuja Mallick, Nalini Ravishanker, and Nandini Kannan. Bivariate positive stable frailty models. Statistics & Probability Letters, 78(15):2371–2377, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001363.

Moral:2004:EMI

[MRP04] Ignacio Moral and Juan M. Rodriguez-Poo. An efficient marginal integration estimator of a semiparametric additive modelling. Statistics & Probability Letters, 69(4):451–463, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001993.

Mukherjea:2006:NMG

[MRS06] A. Mukherjea, M. Rao, and S. Suen. A note on moment generating functions. Statistics & Probability Letters, 76 (11):1185–1189, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520500475X.

Mira:2002:ARB

[MS02a] José Mira and María Jesús Sánchez. Analytical results for a Bayesian bivariate cokriging model. Statistics & Probability Letters, 58(1):97–109, May 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001207.

Morters:2002:TTP

[MS02b] Peter Mörters and Narn-Rueih Shieh. Thin and thick points for branching measure on a Galton–Watson tree. Statistics & Probability Letters, 58(1):13–22, May 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000937.

Maltz:2003:CGF

[MS03a] Alberto L. Maltz and Jorge D. Samur. On a CLT for Gibbs fields and its functional version. Statistics & Probability Letters, 64(3):323–333, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001779.

Minkevicius:2003:LIL

[MS03b] Saulius Minkevicius and Stasys Steisunas. A law of the iterated logarithm for global values of waiting time in multiphase queues. Statistics & Probability Letters, 61(4):359–371, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003930.

Miller:2004:CCR

[MS04a] A. Miller and R. R. Sitter. Choosing columns from the 12-run Plackett–Burman design. Statistics & Probability Letters, 67(2):193–201, April 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000367.

Mohammadpour:2004:SES

[MS04b] Adel Mohammadpour and A. Reza Soltani. On simulating exchangeable sub-Gaussian random vectors. Statistics & Probability Letters, 69(1):29–36, August 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001257.

Meerschaert:2005:LTC

[MS05] Mark M. Meerschaert and Hans-Peter Scheffler. Limit theorems for continuous time random walks with slowly varying waiting times. Statistics & Probability Letters, 71(1):15–22, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002780.

Markiewicz:2007:ODM

[MS07a] A. Markiewicz and A. Szczepańska. Optimal designs in multivariate linear models. Statistics & Probability Letters, 77(4):426–430, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002616.

Meintanis:2007:BGF

[MS07b] Simos Meintanis and Jan Swanepoel. Bootstrap goodness-of-fit tests with estimated parameters based on empirical transforms. Statistics & Probability Letters, 77(10):1004–1013, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000405.

Marchetti:2008:NDI

[MS08a] Giovanni M. Marchetti and Elena Stanghellini. A note on distortions induced by truncation with applications to linear regression systems. Statistics & Probability Letters, 78 (6):824–829, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003380.

Merzbach:2008:CMC

[MS08b] Ely Merzbach and Yair Y. Shaki. Characterizations of multiparameter Cox and Poisson processes by the renewal property. Statistics & Probability Letters, 78(6):637–642, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003069.

Mishura:2008:ASA

[MS08c] Yu. Mishura and G. Shevchenko. Approximate solutions to anticipative stochastic differential equations. Statistics & Probability Letters, 78(1):60–66, January 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001952.

Makalic:2009:MML

[MS09a] Enes Makalic and Daniel F. Schmidt. Minimum Message Length shrinkage estimation. Statistics & Probability Letters, 79(9):1155–1161, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520900008X.

Mazumder:2009:BRM

[MS09b] Satyaki Mazumder and Robert Serfling. Bahadur representations for the median absolute deviation and its modifications. Statistics & Probability Letters, 79(16):1774–1783, August 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001862.

Marella:2008:MNS

[MSC08] Daniela Marella, Mauro Scanu, and Pier Luigi Conti. On the matching noise of some nonparametric imputation procedures. Statistics & Probability Letters, 78(12):1593–1600, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000163.

Misra:2003:SCP

[MSH03] Neeraj Misra, Harshinder Singh, and E. James Harner. Stochastic comparisons of Poisson and binomial random variables with their mixtures. Statistics & Probability Letters, 65(4):279–290, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520300227X.

Muller:2001:IEC

[MSW01] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Improved estimators for constrained Markov chain models. Statistics & Probability Letters, 54(4):427–435, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001213.

Ma:2007:LDP

[MSW07] Yutao Ma, Qiongxia Song, and Liming Wu. Large deviation principles with respect to the τ-topology for exchangeable sequences: a necessary and sufficient condition. Statistics & Probability Letters, 77(3):239–246, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600229X.

Muller:2009:EED

[MSW09] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Estimating the error distribution function in nonparametric regression with multivariate covariates. Statistics & Probability Letters, 79(7):957–964, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800552X.

Mi:2008:SPC

[MSZ08] J. Mi, W. Shi, and Y. Y. Zhou. Some properties of convolutions of Pascal and Erlang random variables. Statistics & Probability Letters, 78(15):2378–2387, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001375.

Marsaglia:2004:BURa

[MT04] George Marsaglia and Wai Wan Tsang. The 64-bit universal RNG. Statistics & Probability Letters, 66(2): 183–187, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://

www.doornik.com/research/randomdouble.pdf;   http:
//www.sciencedirect.com/science/article/pii/S0167715203003353.

Molchanov:2007:OPQ

[MT07a] Ilya Molchanov and Nikolay Tontchev. Optimal Poisson quantisation. Statistics & Probability Letters, 77(11):1123–1132, June 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000648.

Moller:2007:PCF

[MT07b] Jesper Møller and Giovanni Luca Torrisi. The pair correlation function of spatial Hawkes processes. Statistics & Probability Letters, 77(10):995–1003, June 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000399.

Militino:2001:ACF

[MU01] Ana F. Militino and M. Dolores Ugarte. Assessing the covariance function in geostatistics. Statistics & Probability Letters, 52(2):199–206, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002418.

Meintanis:2004:BGF

[MU04] S. Meintanis and N. G. Ushakov. Binned goodness-of-fit tests based on the empirical characteristic function. Statistics & Probability Letters, 69(3):305–314, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001816.

Mukhopadhyay:2009:DRV

[Muk09] Nitis Mukhopadhyay. On p × 1 dependent random variables having each (p − 1) × 1 sub-vector made up of IID observations with examples. Statistics & Probability Letters, 79(14):1585–1589, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001242.

Muller:2001:BOS

[Mül01] Alfred Müller. Bounds for optimal stopping values of dependent random variables with given marginals. Statistics & Probability Letters, 52(1):73–78, March 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002200.

Muselli:2000:UIL

[Mus00] Marco Muselli. Useful inequalities for the longest run distribution. Statistics & Probability Letters, 46(3):239–249, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771529900108X.

Mutafchiev:2008:SPA

[Mut08] Ljuben R. Mutafchiev. Survival probabilities for N-ary subtrees on a Galton–Watson family tree. Statistics & Probability Letters, 78(14):2165–2170, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001053.

Mishura:2001:MTG

[MV01] Yuliya Mishura and Esko Valkeila. Martingale transforms and Girsanov theorem for long-memory Gaussian processes. Statistics & Probability Letters, 55(4):421–430, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001626.

Martinez:2000:ACU

[MW00] W. L. Martinez and E. J. Wegman. An alternative criterion useful for finding exact E-optimal designs. Statistics & Probability Letters, 47(4):325–328, May 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900173X.

Morvai:2005:LIF

[MW05] Gusztáv Morvai and Benjamin Weiss. Limitations on intermittent forecasting. Statistics & Probability Letters, 72(4): 285–290, May 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000386.

Muller:2006:AFM

[MW06] Hans-Georg Müller and Newton Wai. Asymptotic fluctuations of mutagrams. Statistics & Probability Letters, 76 (12):1201–1210, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004761.

Ma:2009:EPT

[MW09] Tiefeng Ma and Songgui Wang. Estimation of the parameters in a two linear regression equations system with identical parameter vectors. Statistics & Probability Letters, 79(9):1135–1140, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005038.

Masuda:2004:ADE

[MY04] Hiroki Masuda and Nakahiro Yoshida. An application of the double Edgeworth expansion to a filtering model with Gaussian limit. Statistics & Probability Letters, 70(1):37–48, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002263.

Mitov:2006:SRP

[MY06] Kosto V. Mitov and Nikolay M. Yanev. Superposition of renewal processes with heavy-tailed interarrival times. Statistics & Probability Letters, 76(6):555–561, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003469.

Miao:2008:LIL

[MY08] Yu Miao and Guangyu Yang. The law of the iterated logarithm for additive functionals of Markov chains. Statistics & Probability Letters, 78(3):265–270, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002234.

Meng:2009:MTB

[MY09] Weiwei Meng and Lin Yu. Martingale transforms between Q_{1} and Q_{ϕ} of martingale spaces. Statistics & Probability Letters, 79(22):2328–2333, November 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003022.

McKeague:2002:SCB

[MZ02] Ian W. McKeague and Yichuan Zhao. Simultaneous confidence bands for ratios of survival functions via empirical likelihood. Statistics & Probability Letters, 60(4):405–415, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003097.

Modarres:2004:MLE

[MZ04] Reza Modarres and Gang Zheng. Maximum likelihood estimation of dependence parameter using ranked set sampling. Statistics & Probability Letters, 68(3):315–323, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001221.

Manzan:2005:KEP

[MZ05] Sebastiano Manzan and Dawit Zerom. Kernel estimation of a partially linear additive model. Statistics & Probability Letters, 72(4):313–322, May 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000726.

McKeague:2006:WSC

[MZ06] Ian W. McKeague and Yichuan Zhao. Width-scaled confidence bands for survival functions. Statistics & Probability Letters, 76(4):327–339, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003032.

Meng:2008:RMD

[MZG08] Qingbin Meng, Xin Zhang, and Junyi Guo. On a risk model with dependence between claim sizes and claim intervals. Statistics & Probability Letters, 78(13):1727–1734, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000333.

Nadarajah:2005:PXS

[Nad05] Saralees Nadarajah. On the product XY for some elliptically symmetric distributions. Statistics & Probability Letters, 75(1):67–75, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002117.

Nadarajah:2008:EEM

[Nad08] Saralees Nadarajah. Explicit expressions for moments of order statistics. Statistics & Probability Letters, 78(2):196–205, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002143.

Nagaev:2003:APS

[Nag03] A. V. Nagaev. Asymptotic properties of stable densities and the asymmetric large deviation problems. Statistics & Probability Letters, 61(4):429–438, February 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202004030.

Nagakura:2009:ATE

[Nag09] Daisuke Nagakura. Asymptotic theory for explosive random coefficient autoregressive models and inconsistency of a unit root test against a stochastic unit root process. Statistics & Probability Letters, 79(24):2476–2483, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003344.

Naito:2001:CCN

[Nai01] Kanta Naito. On a certain class of nonparametric density estimators with reduced bias. Statistics & Probability Letters, 51(1):71–78, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001450.

Nakano:2004:MSR

[Nak04] Yumiharu Nakano. Minimization of shortfall risk in a jumpdiffusion model. Statistics & Probability Letters, 67(1):87–95, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000124.

Nakata:2008:CPO

[Nak08] Toshio Nakata. Collision probability for an occupancy problem. Statistics & Probability Letters, 78(13):1929–1932, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000606.

Namba:2003:PDP

[Nam03] Akio Namba. PMSE dominance of the positive-part shrinkage estimator in a regression model when relevant regressors are omitted. Statistics & Probability Letters, 63(4):375–385, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203001032.

Nane:2008:ITI

[Nan08] Erkan Nane. Isoperimetric-type inequalities for iterated Brownian motion in R^{n}. Statistics & Probability Letters, 78(1):90– 95, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700199X.

Nane:2009:LIL

[Nan09] Erkan Nane. Laws of the iterated logarithm for a class of iterated processes. Statistics & Probability Letters, 79(16): 1744–1751, August 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520900162X.

Nason:2006:SGR

[Nas06] Guy P. Nason. On the sum of t and Gaussian random variables. Statistics & Probability Letters, 76(12):1280–1286, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206000241.

Nasari:2009:WAS

[Nas09] Masoud M. Nasari. On weak approximations of U-statistics. Statistics & Probability Letters, 79(13):1528–1535, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001163.

Natarajan:2001:PMJ

[Nat01] Ranjini Natarajan. On the propriety of a modified Jeffreys’s prior for variance components in binary random effects models. Statistics & Probability Letters, 51(4):409–414, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001826.

Navarro:2008:CUB

[Nav08] Jorge Navarro. Characterizations using the bivariate failure rate function. Statistics & Probability Letters, 78(12): 1349–1354, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207004129.

Nazarov:2009:LLC

[Naz09] Alexander Nazarov. Log-level comparison principle for small ball probabilities. Statistics & Probability Letters, 79(4): 481–486, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004549.

Nanda:2006:PPM

[NBA06] Asok K. Nanda, Subarna Bhattacharjee, and S. S. Alam. Properties of proportional mean residual life model. Statistics & Probability Letters, 76(9):880–890, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004049.

Nanda:2007:PAI

[NBA07] Asok K. Nanda, Subarna Bhattacharjee, and S. S. Alam. Properties of aging intensity function. Statistics & Probability Letters, 77(4):365–373, February 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002525.

Nayak:2008:IEP

[NBK08] Tapan K. Nayak, Sudip Bose, and Subrata Kundu. On inconsistency of estimators of parameters of non-homogeneous Poisson process models for software reliability. Statistics & Probability Letters, 78(14):2217–2221, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001132.

Nieto-Barajas:2007:GAM

[NBW07] Luis E. Nieto-Barajas and Stephen G. Walker. Gibbs and autoregressive Markov processes. Statistics & Probability Letters, 77(14):1479–1485, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000764.

Ni:2007:RIR

[NC07] Liqiang Ni and R. Dennis Cook. A robust inverse regression estimator. Statistics & Probability Letters, 77(3):343–349, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002422.

Nguyen:2009:CBD

[NC09a] Truc T. Nguyen and John T. Chen. A connection between the double gamma model and Laplace sample mean. Statistics & Probability Letters, 79(10):1305–1310, May 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900056X.

Nidhin:2009:LTG

[NC09b] K. Nidhin and C. Chandran. Limit theorems of general functions of independent and identically distributed random variables. Statistics & Probability Letters, 79(23):2397–2404, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209003228.

Negri:2001:EEI

[Neg01] Ilia Negri. On efficient estimation of invariant density for ergodic diffusion processes. Statistics & Probability Letters, 51(1):79–85, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001474.

Neumeyer:2004:CLT

[Neu04] Natalie Neumeyer. A Central Limit Theorem for two-sample U-processes. Statistics & Probability Letters, 67(1):73–85, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000112.

Neumeyer:2007:NUC

[Neu07] Natalie Neumeyer. A note on uniform consistency of monotone function estimators. Statistics & Probability Letters, 77(7):693–703, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206003385.

Nott:2009:GLA

[NFL09] David J. Nott, Mark Fielding, and Daniela Leonte. On a generalization of the Laplace approximation. Statistics & Probability Letters, 79(11):1397–1403, June 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000984.

Nugroho:2002:NTR

[NG02] Sigit Nugroho and Z. Govindarajulu. Nonparametric tests for random effects in the balanced incomplete block design. Statistics & Probability Letters, 56(4):431–437, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202000433.

Ni:2007:SII

[NH07] Xuelei (Sherry) Ni and Xiaoming Huo. Statistical interpretation of the importance of phase information in signal and image reconstruction. Statistics & Probability Letters, 77(4):447–454, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002641.

Nadar:2003:ACM

[NHO03] M. Nadar, T. P. Hettmansperger, and H. Oja. The asymptotic covariance matrix of the Oja median. Statistics & Probability Letters, 64(4):431–442, October 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002104.

Niaparast:2009:ODP

[Nia09] Mehrdad Niaparast. On optimal design for a Poisson regression model with random intercept. Statistics & Probability Letters, 79(6):741–747, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005087.

Nicolaou:2002:CST

[Nic02] Anna Nicolaou. Conditional score tests in the exponential family. Statistics & Probability Letters, 60(1):69–74, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002791.

Nickerson:2008:CWE

[Nic08a] David M. Nickerson. Counting by weighing and its effect on comparing population proportions. Statistics & Probability Letters, 78(15):2321–2326, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001284.

Nicolaie:2008:GRC

[Nic08b] Alina Nicolaie. A generalization of a result concerning the asymptotic behavior of finite Markov chains. Statistics & Probability Letters, 78(18):3321–3329, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003222.

Nicolau:2008:MFT

[Nic08c] João Nicolau. Modeling financial time series through secondorder stochastic differential equations. Statistics & Probability Letters, 78(16):2700–2704, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001879.

Nieminen:2004:FBM

[Nie04] Ari Nieminen. Fractional Brownian motion and martingaledifferences. Statistics & Probability Letters, 70(1):1–10, October 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000665.

Nandi:2002:EFP

[NIK02] Swagata Nandi, Srikanth K. Iyer, and Debasis Kundu. Estimation of frequencies in presence of heavy tail errors. Statistics & Probability Letters, 58(3):265–282, July 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001098.

Nikeghbali:2008:HBB

[Nik08] Ashkan Nikeghbali. How badly are the Burkholder–Davis– Gundy inequalities affected by arbitrary random times? Statistics & Probability Letters, 78(6):766–770, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003306.

Ninomiya:2005:ICG

[Nin05] Yoshiyuki Ninomiya. Information criterion for Gaussian change-point model. Statistics & Probability Letters, 72(3): 237–247, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000349.

Nadarajah:2003:SDG

[NK03] Saralees Nadarajah and Samuel Kotz. Skewed distributions generated by the normal kernel. Statistics & Probability Letters, 65(3):269–277, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002712.

Nkiet:2005:EDL

[Nki05] Guy Martial Nkiet. On estimation of the dimensionality in linear canonical analysis. Statistics & Probability Letters, 75(2):103–112, November 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002154.

Nevalainen:2008:SRT

[NMO08] Jaakko Nevalainen, Jyrki Möttönen, and Hannu Oja. A spatial rank test and corresponding estimators for several samples. Statistics & Probability Letters, 78(6):661–668, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003094.

Nikitin:2000:IAS

[NO00] Yakov Nikitin and Enzo Orsingher. The intermediate arcsine law. Statistics & Probability Letters, 49(2):119–125, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000389.

Namba:2006:PPS

[NO06] Akio Namba and Kazuhiro Ohtani. PMSE performance of the Stein-rule and positive-part Stein-rule estimators in a regression model with or without proxy variables. Statistics & Probability Letters, 76(9):898–906, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004074.

Nzi:2009:BDS

[NO09] Modeste N’zi and Jean-Marc Owo. Backward doubly stochastic differential equations with discontinuous coefficients. Statistics & Probability Letters, 79(7):920–926, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005464.

Nogales:2000:CIR

[NOP00] Agustín G. Nogales, José A. Oyola, and Paloma Pérez. On conditional independence and the relationship between sufficiency and invariance under the Bayesian point of view. Statistics & Probability Letters, 46(1):75–84, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000899.

Nogales:2001:WNE

[NOP01] Agustín G. Nogales, José A. Oyola, and Paloma Pérez. A weak notion of equivalence in Bayesian statistical theory. Statistics & Probability Letters, 53(4):401–407, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000566.

Novak:2003:AMC

[Nov03] S. Y. Novak. On the accuracy of multivariate compound Poisson approximation. Statistics & Probability Letters, 62(1): 35–43, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202004224.

Novak:2004:GCC

[Nov04] S. Y. Novak. On Gebelein’s correlation coefficient. Statistics & Probability Letters, 69(3):299–303, September 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001798.

Novak:2007:NCN

[Nov07] S. Y. Novak. A new characterization of the normal law. Statistics & Probability Letters, 77(1):95–98, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002008.

Neumann:2000:BTS

[NP00] Michael H. Neumann and Efstathios Paparoditis. On bootstrapping L_{2}-type statistics in density testing. Statistics & Probability Letters, 50(2):137–147, November 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000912.

Nair:2008:MED

[NP08a] N. Unnikrishnan Nair and M. Preeth. Multivariate equilibrium distributions of order n. Statistics & Probability Letters, 78(18):3312–3320, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003210.

Neocleous:2008:MRQ

[NP08b] Tereza Neocleous and Stephen Portnoy. On monotonicity of regression quantile functions. Statistics & Probability Letters, 78(10):1226–1229, August 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003975.

Neammanee:2005:AID

[NPC05] K. Neammanee, S. Pianskool, and R. Chonchaiya. On the AIDS incubation distribution. Statistics & Probability Letters, 73(1):13–23, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000751.

Nelsen:2001:DFC

[NQMRLÚF01] Roger B. Nelsen, José Juan Quesada-Molina, José Antonio Rodríguez-Lallena, and Manuel Úbeda-Flores. Distribution functions of copulas: a class of bivariate probability integral transforms. Statistics & Probability Letters, 54(3):277–282, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000608.

Nelsen:2003:KDF

[NQMRLÚF03] Roger B. Nelsen, José Juan Quesada-Molina, José Antonio Rodríguez-Lallena, and Manuel Úbeda-Flores. Kendall distribution functions. Statistics & Probability Letters, 65(3): 263–268, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203002700.

Navarro:2005:NCA

[NRS05] Jorge Navarro, Jose M. Ruiz, and Carlos J. Sandoval. A note on comparisons among coherent systems with dependent components using signatures. Statistics & Probability Letters, 72(2):179–185, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000295.

Nicoleris:2002:RFP

[NS02] Theodoros Nicoleris and Anthony Sagris. Random function prediction and Stein’s identity. Statistics & Probability Letters, 59(3):293–305, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001979.

Nematollahi:2006:IIC

[NS06a] A. R. Nematollahi and A. R. Soltani. An inequality involving correlations. Statistics & Probability Letters, 76 (4):369–372, February 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205003081.

Nourdin:2006:ACO

[NS06b] Ivan Nourdin and Thomas Simon. On the absolute continuity of one-dimensional SDEs driven by a fractional Brownian motion. Statistics & Probability Letters, 76(9):907–912, May 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205004062.

Nguyen:2009:CLL

[NS09] Danh V. Nguyen and Damla Sentürk. A consistent local linear estimator of the covariate adjusted correlation coefficient. Statistics & Probability Letters, 79(15):1684–1689, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001552.

Nishikawa:2003:CIP

[NT03] Masako Nishikawa and Toshiro Tango. Counter-intuitive properties of the Kaplan–Meier estimator. Statistics & Probability Letters, 65(4):353–361, December 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002724.

Nagaev:2006:TAC

[NT06a] A. Nagaev and G. Tsitsiashvili. Tail asymptotics of the n th convolution of super-exponential distributions. Statistics & Probability Letters, 76(9):861–870, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003147.

Nowak:2006:LIP

[NT06b] Emmanuel Nowak and Emmanuel Thilly. A local invariance principle for Gibbsian fields. Statistics & Probability Letters, 76(18):1975–1982, December 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001817.

Nashimoto:2005:NMC

[NW05] Kane Nashimoto and F. T. Wright. A note on multiple comparison procedures for detecting differences in simply ordered means. Statistics & Probability Letters, 73(4):393–401, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001434.

Nyrhinen:2005:UBG

[Nyr05] Harri Nyrhinen. Upper bounds of the Gärtner-Ellis theorem for the sequences of random variables. Statistics & Probability Letters, 73(1):57–60, June 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000829.

Nayak:2000:FIC

[NZ00] S. S. Nayak and Madhusudhan Zalki. On the fluctuations of independent copies of moving maxima. Statistics & Probability Letters, 50(4):351–356, December 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001152.

Nayak:2001:ASL

[NZ01] S. S. Nayak and Madhusudhan Zalki. Almost sure limit points of record values from two independent populations. Statistics & Probability Letters, 51(2):181–187, January 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001516.

Ning:2009:MBT

[NZY09] Wei Ning, Sanguo Zhang, and Chang Yu. A momentbased test for the homogeneity in mixture natural exponential family with quadratic variance functions. Statistics & Probability Letters, 79(6):828–834, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005221.

Oncel:2005:SRO

[OAAA05] Sevgi Yurt Oncel, Mohammad Ahsanullah, Fazil A. Aliev, and Funda Aygun. Switching record and order statistics via random contractions. Statistics & Probability Letters, 73(3):207– 217, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001008.

Ongaro:2004:DRP

[OC04] Andrea Ongaro and Carla Cattaneo. Discrete random probability measures: a general framework for nonparametric Bayesian inference. Statistics & Probability Letters, 67(1):

33–45, March 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000033.

Ojeda:2007:BLR

[ODO07] Enrique Calderín Ojeda, Emilio Gómez Déniz, and Ignacio J. Cabrera Ortega. Bayesian local robustness under weighted squared-error loss function incorporating unimodality. Statistics & Probability Letters, 77(1):69–74, January 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001969.

Ouknine:2008:EVP

[OE08a] Youssef Ouknine and Mohamed Erraoui. Equivalence of Volterra processes: Degenerate case. Statistics & Probability Letters, 78(4):435–444, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207002593.

Ouknine:2008:NRI

[OE08b] Youssef Ouknine and Mohamed Erraoui. Noncanonical representation with an infinite-dimensional orthogonal complement. Statistics & Probability Letters, 78(10):1200–1205, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207003938.

Oehlert:2000:SNK

[Oeh00] Gary W. Oehlert. Student–Newman–Kuels controls the false discovery rate. Statistics & Probability Letters, 46(4):381– 383, February 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001261.

See correction [Oeh01].

Oehlert:2001:CSN

[Oeh01] Gary W. Oehlert. Correction on “Student–Newman–Kuels controls the false discovery rate” [Statist. Probab. Lett. 46 (2000) 381–383]. Statistics & Probability Letters, 51(4): 431, February 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001875.

See [Oeh00].

Olive:2003:RRH

[OH03] David J. Olive and Douglas M. Hawkins. Robust regression with high coverage. Statistics & Probability Letters, 63(3):259– 266, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000907.

Oberhofer:2005:ADU

[OH05] Walter Oberhofer and Harry Haupt. The asymptotic distribution of the unconditional quantile estimator under dependence. Statistics & Probability Letters, 73(3):243–250, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001082.

Olive:2007:BES

[OH07] David J. Olive and Douglas M. Hawkins. Behavior of elemental sets in regression. Statistics & Probability Letters, 77(6):621–624, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002951.

Ohtani:2001:MDP

[Oht01] Kazuhiro Ohtani. MSE dominance of the pre-test iterative variance estimator over the iterative variance estimator in regression. Statistics & Probability Letters, 54(3):331–340, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001316.

Ohtani:2002:EDP

[Oht02] Kazuhiro Ohtani. Exact distribution of a pre-test estimator for regression error variance when there are omitted variables. Statistics & Probability Letters, 60(2):129–140, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002821.

Ozawa:2006:ISP

[OK06] Kazuhiro Ozawa and Shinji Kuriki. Incomplete split-plot designs generated from α-resolvable designs. Statistics & Probability Letters, 76(12):1245–1254, July 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000204.

Ozkale:2007:SCE

[ÖK07] M. Revan Özkale and Selahattin Kaçiranlar. Superiority of the class estimator over some estimators by the mean square error matrix criterion. Statistics & Probability Letters, 77(4):438–446, February 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600263X.

Ouhbi:2002:ROF

[OL02] Brahim Ouhbi and Nikolaos Limnios. The rate of occurrence of failures for semi-Markov processes and estimation. Statistics & Probability Letters, 59(3):245–255, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001396.

Olkin:2003:BBD

[OL03] Ingram Olkin and Ruixue Liu. A bivariate beta distribution. Statistics & Probability Letters, 62(4):407–412, May 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000488.

Oleszkiewicz:2001:CCP

[Ole01] Krzysztof Oleszkiewicz. Concentration of capital — the product form of the Law of Large Numbers in L_{1}. Statistics & Probability Letters, 55(2):159–162, November 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000955.

Oleszkiewicz:2008:FBM

[Ole08] Krzysztof Oleszkiewicz. On fake Brownian motions. Statistics & Probability Letters, 78(11):1251–1254, August 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004014.

Olive:2001:HBA

[Oli01] David J. Olive. High breakdown analogs of the trimmed mean. Statistics & Probability Letters, 51(1):87–92, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001486.

Oliveira:2005:EIA

[Oli05] Paulo Eduardo Oliveira. An exponential inequality for associated variables. Statistics & Probability Letters, 73(2):189–197, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000337.

Oden:2006:EMP

[OM06] Neal L. Oden and Matthew J. McIntosh. Exact moments and probabilities for Wei’s urn randomization model. Statistics & Probability Letters, 76(15):1694–1700, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001283.

Omelka:2008:CTT

[Ome08] Marek Omelka. Comparison of two types of confidence intervals based on Wilcoxon-type R-estimators. Statistics & Probability Letters, 78(18):3366–3372, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003283.

Ozawa:2004:ISP

[OMJ^{+}04] Kazuhiro Ozawa, Stanisław Mejza, Masakazu Jimbo, Iwona Mejza, and Shinji Kuriki. Incomplete split-plot designs generated by some resolvable balanced designs. Statistics & Probability Letters, 68(1):9–15, June 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000331.

Omey:2009:NRR

[OMM09] Edward Omey, Georgi K. Mitov, and Kosto V. Mitov. On the number of renewals in random time. Statistics & Probability Letters, 79(21):2281–2288, November 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002909.

Omori:2003:EUS

[Omo03] Yasuhiro Omori. Estimation for unequally spaced time series of counts with serially correlated random effects. Statistics & Probability Letters, 63(1):1–12, May 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003437.

Omori:2007:EGS

[Omo07] Yasuhiro Omori. Efficient Gibbs sampler for Bayesian analysis of a sample selection model. Statistics & Probability Letters, 77(12):1300–1311, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000934.

Okolewski:2001:SDF

[OR01] Andrzej Okolewski and Tomasz Rychlik. Sharp distributionfree bounds on the bias in estimating quantiles via order statistics. Statistics & Probability Letters, 52(2):207–213, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520000242X.

Orme:2000:SOT

[Orm00] Chris D. Orme. On the sensitivity of the overdispersion test in a Weibull model. Statistics & Probability Letters, 46(1):95–100, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299000917.

Ouassou:2002:EPV

[OS02] Idir Ouassou and William E. Strawderman. Estimation of a parameter vector restricted to a cone. Statistics & Probability Letters, 56(2):121–129, January 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001444.

Oyet:2003:TVW

[OS03] Alwell J. Oyet and Brajendra Sutradhar. Testing variances in wavelet regression models. Statistics & Probability Letters, 61(1):97–109, January 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520200353X.

Ostrovska:2005:SCI

[OS05] Sofiya Ostrovska and Jordan Stoyanov. Stieltjes classes for M-indeterminate powers of inverse Gaussian distributions. Statistics & Probability Letters, 71(2):165–171, February 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204002913.

Ould-Said:2006:SUC

[OS06] Elias Ould-Said̈. A strong uniform convergence rate of kernel conditional quantile estimator under random censorship. Statistics & Probability Letters, 76(6):579–586, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003494.

Osekowski:2009:BCW

[Ose09a] Adam Osekowski. On the best constant in the weak type inequality for the square function of a conditionally symmetric martingale. Statistics & Probability Letters, 79(13):1536–1538, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001175.

Osekowski:2009:SNC

[Ose09b] Adam Osekowski. Sharp norm comparison of the maxima of a sequence and its predictable projection. Statistics & Probability Letters, 79(16):1784–1788, August 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001813.

Osmoukhina:2001:LDP

[Osm01] Anna V. Osmoukhina. Large deviations probabilities for a test of symmetry based on kernel density estimator. Statistics & Probability Letters, 54(4):363–371, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000360.

Ostrovska:2001:SRV

[Ost01] Sofiya Ostrovska. Sets of random variables with a given uncorrelation structure. Statistics & Probability Letters, 55 (4):359–366, December 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001122.

Otmani:2009:BDL

[Otm09] Mohamed El Otmani. BSDEs driven by Lévy process with enlarged filtration and applications in finance. Statistics & Probability Letters, 79(1):44–49, January 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003349.

Ozturk:2000:ARS

[ÖW00] Ömer Öztürk and Douglas A. Wolfe. Alternative ranked set sampling protocols for the sign test. Statistics & Probability Letters, 47(1):15–23, March 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001327.

Oyet:2003:EMW

[OW03] Alwell J. Oyet and Douglas P. Wiens. On exact minimax wavelet designs obtained by simulated annealing. Statistics & Probability Letters, 61(1):111–121, January 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003619.

Ozkale:2008:JRE

[Özk08] M. Revan Özkale. A jackknifed ridge estimator in the linear regression model with heteroscedastic or correlated errors. Statistics & Probability Letters, 78(18):3159–3169, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002861.

Ozturk:2007:TSM

[Ozt07] Omer Ozturk. Two-sample median test for order restricted randomized designs. Statistics & Probability Letters, 77 (2):131–141, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002124.

Padmanabhan:2001:RWS

[Pad01] A. S. Padmanabhan. Random walks strictly confined to a subspace. Statistics & Probability Letters, 53(3):299–303, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000852.

Paindaveine:2008:CDS

[Pai08] Davy Paindaveine. A canonical definition of shape. Statistics & Probability Letters, 78(14):2240–2247, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800117X.

Pundir:2005:BCR

[PAJ05] Sudesh Pundir, Sangeeta Arora, and Kanchan Jain. Bonferroni curve and the related statistical inference. Statistics & Probability Letters, 75(2):140–150, November 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002208.

Pakes:2008:OSR

[Pak08a] Anthony G. Pakes. Order statistics and renormalization. Statistics & Probability Letters, 78(5):463–470, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002623.

Pakshirajan:2008:NWC

[Pak08b] R. P. Pakshirajan. A note on the weak convergence of probability measures in the D[0, 1] space. Statistics & Probability Letters, 78(6):716–719, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003227.

Palacios:2000:NCM

[Pal00] José Luis Palacios. A note on circular Markov chains. Statistics & Probability Letters, 47(3):301–306, April 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001704.

Pal:2001:LBO

[Pal01a] Chandranath Pal. On a locally Bahadur optimal test for no contamination in mixtures from the one-parameter exponential family. Statistics & Probability Letters, 53(4):429–434, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000827.

Palacios:2001:CRW

[Pal01b] José Luis Palacios. Correction on “Random walks on edgetransitive graphs (II)” [Statist. Probab. Lett. 43 (1999) 25– 32]. Statistics & Probability Letters, 53(1):111–112, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001863. See [PRB99].

Pal:2008:SPM

[Pal08] Jayanta Kumar Pal. Spiking problem in monotone regression: Penalized residual sum of squares. Statistics & Probability Letters, 78(12):1548–1556, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000102.

Palacios:2009:HTR

[Pal09] José Luis Palacios. On hitting times of random walks on trees. Statistics & Probability Letters, 79(2):234–236, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520800374X.

Pan:2004:DOD

[Pan04] Jianxin Pan. Discordant outlier detection in the growth curve model with Rao’s simple covariance structure. Statistics & Probability Letters, 69(2):135–142, August 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400166X.

Papadatos:2001:DEB

[Pap01] Nickos Papadatos. Distribution and expectation bounds on order statistics from possibly dependent variates. Statistics & Probability Letters, 54(1):21–31, August 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000141.

Papanastassiou:2006:CCM

[Pap06] Demetrios Papanastassiou. Computing the covariance matrix of QML estimators for a state space model. Statistics & Probability Letters, 76(10):1001–1006, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004396.

Park:2001:ESI

[Par01] Byeong U. Park. On estimating the slope of increasing boundaries. Statistics & Probability Letters, 52(1):69–72, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200002194.

Paszkiewicz:2006:WTC

[Pas06] Adam Paszkiewicz. When transition count for a Markov chains is a complete sufficient statistic. Statistics & Probability Letters, 76(8):757–763, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003457.

Patie:2009:FRS

[Pat09] P. Patie. A few remarks on the supremum of stable processes. Statistics & Probability Letters, 79(8):1125–1128, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209000054.

Pawitan:2000:CEL

[Paw00] Yudi Pawitan. Computing empirical likelihood from the bootstrap. Statistics & Probability Letters, 47(4):337–345, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001753.

Palma:2003:EGF

[PB03] Wilfredo Palma and Pascal Bondon. On the eigenstructure of generalized fractional processes. Statistics & Probability Letters, 65(2):93–101, November 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002219.

Park:2009:SCF

[PB09] Sangun Park and N. Balakrishnan. On simple calculation of the Fisher information in hybrid censoring schemes. Statistics & Probability Letters, 79(10):1311–1319, May 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000571.

Park:2009:PTG

[PBH09] J. A. Park, J. S. Baek, and S. Y. Hwang. Persistentthreshold-GARCH processes: Model and application. Statistics & Probability Letters, 79(7):907–914, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005348.

Park:2008:FIH

[PBZ08] Sangun Park, N. Balakrishnan, and Gang Zheng. Fisher information in hybrid censored data. Statistics & Probability Letters, 78(16):2781–2786, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002010.

Poiraud-Casanova:2000:AMR

[PCTA00] Sandrine Poiraud-Casanova and Christine Thomas-Agnan. About monotone regression quantiles. Statistics & Probability Letters, 48(1):101–104, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771529900200X.

Pradhan:2009:IDC

[PD09] Biswabrata Pradhan and Anup Dewanji. On induced dependent censoring for quality adjusted lifetime (QAL) data in a simple illness-death model. Statistics & Probability Letters, 79(20):2170–2176, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002612.

Priebe:2001:DDN

[PDM01] Carey E. Priebe, Jason G. DeVinney, and David J. Marchette. On the distribution of the domination number for random class cover catch digraphs. Statistics & Probability Letters, 55(3):239–246, December 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001298.

Pellerey:2000:RVH

[Pel00] Franco Pellerey. Random vectors with HNBUE-type marginal distributions. Statistics & Probability Letters, 50(3):265– 271, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001024.

Peligrad:2002:SRC

[Pel02] Magda Peligrad. Some remarks on coupling of dependent random variables. Statistics & Probability Letters, 60(2): 201–209, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202003188.

Pelletier:2005:KDE

[Pel05] Bruno Pelletier. Kernel density estimation on Riemannian manifolds. Statistics & Probability Letters, 73(3):297–304, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001239.

Peng:2001:EMH

[Pen01a] Liang Peng. Estimating the mean of a heavy tailed distribution. Statistics & Probability Letters, 52(3):255–264, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200002030.

Peng:2001:SPE

[Pen01b] Liang Peng. Semi-parametric estimation of long-range dependence index in infinite variance time series. Statistics & Probability Letters, 51(2):101–109, January 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000122X.

Peng:2002:AED

[Pen02a] Liang Peng. Asymptotic expansions of densities of sums of random vectors without third moment. Statistics & Probability Letters, 58(2):167–174, June 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001219.

Pensky:2002:DDB

[Pen02b] Marianna Pensky. Density deconvolution based on wavelets with bounded supports. Statistics & Probability Letters, 56 (3):261–269, February 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001766.

Peng:2008:FNM

[Pen08] Chien-Yu Peng. The first negative moment in the sense of the Cauchy principal value. Statistics & Probability Letters, 78(13):1765–1774, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000382.

Perera:2002:MQE

[Per02] Subashan Perera. Maximum quasilikelihood estimation for a simplified NEAR(1) model. Statistics & Probability Letters, 58(2):147–155, June 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001128.

Pereira:2009:ALM

[Per09] L. Pereira. The asymptotic location of the maximum of a stationary random field. Statistics & Probability Letters, 79(20):2166–2169, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209002600.

Petrov:2002:NBC

[Pet02] Valentin V. Petrov. A note on the Borel–Cantelli lemma. Statistics & Probability Letters, 58(3):283–286, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200113X.

Petrov:2004:GBC

[Pet04] Valentin V. Petrov. A generalization of the Borel–Cantelli Lemma. Statistics & Probability Letters, 67(3):233–239, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204000264.

Petrovic:2008:MES

[Pet08] Ljiljana Petrović. Markovian extensions of a stochastic process. Statistics & Probability Letters, 78(6):810–814, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003367.

Pardo-Fernandez:2007:CED

[PF07] Juan Carlos Pardo-Fernández. Comparison of error distributions in nonparametric regression. Statistics & Probability Letters, 77(3):350–356, February 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002434.

Pacheco-Gonzalez:2009:EBM

[PG09] Carlos G. Pacheco-González. Ergodicity of a bounded Markov chain with attractiveness towards the centre. Statistics & Probability Letters, 79(20):2177–2181, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002624.

Papana:2009:GHF

[PI09] Ariadni Papana and Hemant Ishwaran. Gene hunting with forests for multigroup time course data. Statistics & Probability Letters, 79(9):1146–1154, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000078.

Pingyan:2002:LBW

[Pin02] Chen Pingyan. Limiting behavior of weighted sums with stable distributions. Statistics & Probability Letters, 60(4): 367–375, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715202002869.

Pizzinga:2009:FIR

[Piz09] Adrian Pizzinga. Further investigation into restricted Kalman filtering. Statistics & Probability Letters, 79(2):264–269, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003878.

Park:2004:ERF

[PK04] Cheol-Woo Park and Woo-Chul Kim. Estimation of a regression function with a sharp change point using boundary wavelets. Statistics & Probability Letters, 66(4):435–448, March 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203003651.

Park:2007:FIM

[PK07a] Jeong-Soo Park and Tae Yoon Kim. Fisher information matrix for a four-parameter kappa distribution. Statistics & Probability Letters, 77(13):1459–1466, July 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000788.

Przystalski:2007:CET

[PK07b] Marcin Przystalski and Paweł Krajewski. Constrained estimators of treatment parameters in semiparametric models. Statistics & Probability Letters, 77(9):914–919, May 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000302. See erratum [PK08].

Przystalski:2008:ECE

[PK08] Marcin Przystalski and Paweł Krajewski. Erratum to “Constrained estimators of treatment parameters in semiparametric models” [Statist. Probab. Lett. 77 (2007) 914–919]. Statistics & Probability Letters, 78(9):1158–1159, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003835. See [PK07b].

Parzen:2007:PMR

[PL07] Michael Parzen and Stuart R. Lipsitz. Perturbing the minimand resampling with Gamma(1,1) random variables as an extension of the Bayesian bootstrap. Statistics & Probability Letters, 77(6):654–657, March 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003002.

Patriota:2009:BCM

[PL09a] Alexandre G. Patriota and Artur J. Lemonte. Bias correction in a multivariate normal regression model with general parameterization. Statistics & Probability Letters, 79(15):1655–1662, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001527.

Peng:2009:CME

[PL09b] Jun Peng and Zaiming Liu. On a class of mathematical ecosystems with random jumps. Statistics & Probability Letters, 79(5):630–636, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004902.

Pateiro-Lopez:2006:MPL

[PLGM06] Beatriz Pateiro-López and Wenceslao González-Manteiga. Multivariate partially linear models. Statistics & Probability Letters, 76(14):1543–1549, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000976.

Peng:2006:OMC

[PLLJ06] J. Peng, C. I. C. Lee, L. Liu, and J. Jiang. Orthogonal multiple contrast test for testing monotone on the average. Statistics & Probability Letters, 76(2):161–168, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002853.

Pang:2003:NOA

[PLZ03] Shanqi Pang, Sanyang Liu, and Yingshan Zhang. A note on orthogonal arrays obtained by orthogonal decomposition of projection matrices. Statistics & Probability Letters, 63(4):411–416, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001202.

Pazman:2001:ODE

[PM01] Andrej Pázman and Werner G. Müller. Optimal design of experiments subject to correlated errors. Statistics & Probability Letters, 52(1):29–34, March 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002017.

Pal:2006:ODO

[PM06] Manisha Pal and Nripes K. Mandal. Optimum designs for optimum mixtures. Statistics & Probability Letters, 76(13): 1369–1379, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000447.

Pal:2008:MDO

[PM08] Manisha Pal and Nripes Kumar Mandal. Minimax designs for optimum mixtures. Statistics & Probability Letters, 78 (6):608–615, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003021.

Porto:2008:WRC

[PMA08] Rogério F. Porto, Pedro A. Morettin, and Elisete C. Q. Aubin. Wavelet regression with correlated errors on a piecewise Hölder class. Statistics & Probability Letters, 78(16): 2739–2743, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001946.

Pan:2005:SLT

[PMJ05] Guangming Pan, Boqi Miao, and Baisuo Jin. Some limiting theorems of some random quadratic forms. Statistics & Probability Letters, 75(3):151–157, December 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002695.

Pan:2008:CLT

[PMJ08] Guangming Pan, Baiqi Miao, and Baisuo Jin. Central limit theorem of random quadratics forms involving random matrices. Statistics & Probability Letters, 78(6):804–809, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003355.

Pancheva:2009:LTE

[PMM09] Elisaveta Pancheva, Ivan K. Mitov, and Kosto V. Mitov. Limit theorems for extremal processes generated by a point process with correlated time and space components. Statistics & Probability Letters, 79(3):390–395, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004240.

Paula:2009:IDL

[PMVL09] Gilberto A. Paula, Marcio Medeiros, and Filidor E. Vilca-Labra. Influence diagnostics for linear models with firstorder autoregressive elliptical errors. Statistics & Probability Letters, 79(3):339–346, February 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004161.

Porcu:2007:MST

[PMZP07] E. Porcu, J. Mateu, A. Zini, and R. Pini. Modelling spatiotemporal data: a new variogram and covariance structure proposal. Statistics & Probability Letters, 77(1):83–89, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001982.

Pascu:2009:DVL

[PN09] M. N. Pascu and A. Nicolaie. On a discrete version of the Laugesen–Morpurgo conjecture. Statistics & Probability Letters, 79(6):797–806, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800518X.

Park:2008:EUR

[PNJ08] Hyeonah Park, Seongryong Na, and Jongwoo Jeon. Estimation using response probability under callbacks. Statistics & Probability Letters, 78(13):1735–1741, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000345.

Pommeret:2001:PVQ

[Pom01] Denys Pommeret. Posterior variance for quadratic natural exponential families. Statistics & Probability Letters, 53(4):357– 362, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000244.

Pons:2006:SPE

[Pon06] Odile Pons. Semi-parametric estimation for a semi-Markov process with left-truncated and right-censored observations. Statistics & Probability Letters, 76(9):952–958, May 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004141.

Poon:2006:IIO

[Poo06] Wai-Yin Poon. Identifying influential observations in logistic discriminant analysis. Statistics & Probability Letters, 76(13):1348–1355, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000411.

Porosinski:2002:CMC

[Por02a] Zdzisław Porosiński. Characterization of the monotone case for a best choice problem with a random number of objects. Statistics & Probability Letters, 56(4):419–423, February 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520200038X.

Porosinski:2002:BCP

[Por02b] Zdzisław Porosiński. On best choice problems having similar solutions. Statistics & Probability Letters, 56(3):321–327, February 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201002012.

Porosinski:2003:OCO

[Por03] Zdzisław Porosiński. On optimal choosing of one of the k best objects. Statistics & Probability Letters, 65(4):419–432, December 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002967.

Poufinas:2008:NDG

[Pou08] Thomas Poufinas. On the number of deviations of Geometric Brownian Motion with drift from its extreme points with applications to transaction costs. Statistics & Probability Letters, 78(17):3040–3046, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002654.

Pozdnyakov:2004:FCP

[Poz04] Vladimir Pozdnyakov. On the functional CLT for partial sums of truncated bounded from below random variables. Statistics & Probability Letters, 70(2):137–144, November 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715204001919.

Pozdnyakov:2008:OPM

[Poz08] Vladimir Pozdnyakov. On occurrence of patterns in Markov chains: Method of gambling teams. Statistics & Probability Letters, 78(16):2762–2767, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001983.

Pena:2000:KCL

[PP00] Daniel Peña and Francisco J. Prieto. The kurtosis coefficient and the linear discriminant function. Statistics & Probability Letters, 49(3):257–261, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000559.

Poggi:2001:ALT

[PP01] Jean-Michel Poggi and Bruno Portier. Asymptotic local test for linearity in adaptive control. Statistics & Probability Letters, 55(1):9–17, November 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100116X.

Paparoditis:2005:BHT

[PP05] Efstathios Paparoditis and Dimitris N. Politis. Bootstrap hypothesis testing in regression models. Statistics & Probability Letters, 74(4):356–365, October 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001951.

Pazman:2006:IBL

[PP06] Andrej Pázman and Luc Pronzato. On the irregular behavior of LS estimators for asymptotically singular designs. Statistics & Probability Letters, 76(11):1089–1096, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004517.

Palmowski:2009:CAF

[PP09] Zbigniew Palmowski and Martijn Pistorius. Cramér asymptotics for finite time first passage probabilities of general Lévy processes. Statistics & Probability Letters, 79(16):1752–1758, August 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001631.

Perera:2008:ESM

[PPRW08] D. I. Perera, M. S. Peiris, J. Robinson, and N. C. Weber. The empirical saddlepoint method applied to testing for serial correlation in panel time series data. Statistics & Probability Letters, 78(17):2876–2882, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002290.

Penev:2004:EEF

[PPSW04] Spiridon Penev, Hanxiang Peng, Anton Schick, and Wolfgang Wefelmeyer. Efficient estimators for functionals of Markov chains with parametric marginals. Statistics & Probability Letters, 66(3):335–345, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003614.

Peng:2003:CTL

[PQ03] Liang Peng and Yongcheng Qi. Chover-type laws of the iterated logarithm for weighted sums. Statistics & Probability Letters, 65(4):401–410, December 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002931.

Perrin:2007:GST

[PR07] Olivier Perrin and Edmond Redside. Generalization of Simmons’ theorem. Statistics & Probability Letters, 77(6):604– 606, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002926.

Presnell:2008:MRL

[PR08] Brett Presnell and Pavlina Rumcheva. The mean resultant length of the spherically projected normal distribution. Statistics & Probability Letters, 78(5):557–563, April 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002878.

Pai:2009:MPC

[PR09a] Jeffrey Pai and Nalini Ravishanker. A multivariate preconditioned conjugate gradient approach for maximum likelihood estimation in vector long memory processes. Statistics & Probability Letters, 79(9):1282–1289, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000455.

Porzio:2009:SOF

[PR09b] Giovanni C. Porzio and Giancarlo Ragozini. On the stochastic ordering of folded binomials. Statistics & Probability Letters, 79(9):1299–1304, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000558.

Palacios:1999:RWE

[PRB99] José Luis Palacios, José Miguel Renom, and Pedro Berrizbeitia. Random walks on edge-transitive graphs (II). Statistics & Probability Letters, 43(1):25–32, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0167715298002417. See correction [Pal01b].

Preater:2000:SSB

[Pre00] J. Preater. Sequential selection with a better-than-average rule. Statistics & Probability Letters, 50(2):187–191, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000973.

Preater:2009:SVM

[Pre09] J. Preater. A species of voter model driven by immigration. Statistics & Probability Letters, 79(20):2131–2137, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209002508.

Privault:2001:ECI

[Pri01] Nicolas Privault. Extended covariance identities and inequalities. Statistics & Probability Letters, 55(3):247–255, December 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001407.

Pronzato:2003:RNO

[Pro03] Luc Pronzato. Removing non-optimal support points in D-optimum design algorithms. Statistics & Probability Letters, 63(3):223–228, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000816.

Pronzato:2004:MET

[Pro04] Luc Pronzato. A minimax equivalence theorem for optimum bounded design measures. Statistics & Probability Letters, 68(4):325–331, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001051.

Proietti:2008:MDT

[Pro08] Tommaso Proietti. Missing data in time series: a note on the equivalence of the dummy variable and the skipping approaches. Statistics & Probability Letters, 78(3):257– 264, February 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002222.

Prokaj:2009:USR

[Pro09a] Vilmos Prokaj. Unfolding the Skorohod reflection of a semimartingale. Statistics & Probability Letters, 79(4):534– 536, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004628.

Pronzato:2009:APN

[Pro09b] Luc Pronzato. Asymptotic properties of nonlinear estimates in stochastic models with finite design space. Statistics & Probability Letters, 79(21):2307–2313, November 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002934.

Perrin:2000:RNS

[PS00] Olivier Perrin and Rachid Senoussi. Reducing non-stationary random fields to stationarity and isotropy using a space deformation. Statistics & Probability Letters, 48(1):23–32, May 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001881.

Peng:2002:EEL

[PS02] Hanxiang Peng and Anton Schick. On efficient estimation of linear functionals of a bivariate distribution with known marginals. Statistics & Probability Letters, 59(1):83–91, August 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202002067.

Park:2006:STU

[PS06a] Soo Jung Park and Dong Wan Shin. A sign test for unit roots in a momentum threshold autoregressive process. Statistics & Probability Letters, 76(10):986–990, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004372.

Pereira:2006:NNF

[PS06b] I. Pereira and M. G. Scotto. On the non-negative firstorder exponential bilinear time series model. Statistics & Probability Letters, 76(9):931–938, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004098.

Perrin:2007:CMG

[PS07a] Olivier Perrin and Martin Schlather. Can any multivariate Gaussian vector be interpreted as a sample from a stationary random process? Statistics & Probability Letters, 77(9):881– 884, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520700020X.

Pingyan:2007:LBW

[PS07b] Chen Pingyan and Gan Shixin. Limiting behavior of weighted sums of i.i.d. random variables. Statistics & Probability Letters, 77(16):1589–1599, October 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001320.

Peng:2009:IEM

[PS09a] Hanxiang Peng and Anton Schick. Improving efficient marginal estimators in bivariate models with parametric marginals. Statistics & Probability Letters, 79(23):2437–2442, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209003289.

Petsa:2009:EMC

[PS09b] Athanasia Petsa and Theofanis Sapatinas. Erratum to: “Minimax convergence rates under the L^{p}-risk in the functional deconvolution model” [Statist. Probab. Lett. 79 (2009) 1568– 1576]. Statistics & Probability Letters, 79(17):1890, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209002247. See [PS09c].

Petsa:2009:MCR

[PS09c] Athanasia Petsa and Theofanis Sapatinas. Minimax convergence rates under the L^{p}-risk in the functional deconvolution model. Statistics & Probability Letters, 79(13):1568–1576, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001278. See erra-

tum [PS09b].

Psaradakis:2000:BTU

[Psa00] Zacharias Psaradakis. Bootstrap tests for unit roots in seasonal autoregressive models. Statistics & Probability Letters, 50(4):389–395, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001280.

Psaradakis:2002:ABU

[Psa02] Zacharias Psaradakis. On the asymptotic behaviour of unitroot tests in the presence of a Markov trend. Statistics & Probability Letters, 57(1):101–109, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000639.

Psaradakis:2003:SBT

[Psa03] Zacharias Psaradakis. A sieve bootstrap test for stationarity. Statistics & Probability Letters, 62(3):263–274, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000129.

Psaradakis:2006:BBT

[Psa06] Zacharias Psaradakis. Blockwise bootstrap testing for stationarity. Statistics & Probability Letters, 76(6):562–570, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003470.

Psarrakos:2009:ARH

[Psa09a] Georgios Psarrakos. Asymptotic results for heavy-tailed distributions using defective renewal equations. Statistics & Probability Letters, 79(6):774–779, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005130.

Psarrakos:2009:NCC

[Psa09b] Georgios Psarrakos. A note on convolutions of compound geometric distributions. Statistics & Probability Letters, 79(9):1231–1237, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000376.

Pal:2000:ESD

[PSS00] Nabendu Pal, Wei-Hsiung Shen, and Bimal K. Sinha. Estimation of the support of a discrete distribution. Statistics & Probability Letters, 48(3):287–292, July 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000092.

Pipiras:2003:CCT

[PTA03] Vladas Pipiras, Murad S. Taqqu, and Patrice Abry. Can continuous-time stationary stable processes have discrete linear representations? Statistics & Probability Letters, 64(2):147–157, August 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203001469.

Puig:2008:NHL

[Pui08] Pedro Puig. A note on the harmonic law: a two-parameter family of distributions for ratios. Statistics & Probability Letters, 78(3):320–326, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002441.

Phillips:2000:NUB

[PW00] M. J. Phillips and Graham V. Weinberg. Non-uniform bounds for geometric approximation. Statistics & Probability Letters, 49(3):305–311, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000626.

Pitt:2006:ECS

[PW06] Michael K. Pitt and Stephen G. Walker. Extended constructions of stationary autoregressive processes. Statistics & Probability Letters, 76(12):1219–1224, July 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004785.

Padoan:2008:MMB

[PW08] S. A. Padoan and M. P. Wand. Mixed model-based additive models for sample extremes. Statistics & Probability Letters, 78(17):2850–2858, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002198.

Pycke:2003:MEA

[Pyc03] Jean-Renaud Pycke. Multivariate extensions of the Anderson– Darling process. Statistics & Probability Letters, 63(4): 387–399, July 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001111.

Pycke:2007:SBG

[Pyc07] J.-R. Pycke. U-statistics based on the Green’s function of the Laplacian on the circle and the sphere. Statistics & Probability Letters, 77(9):863–872, May 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003440.

Peng:2008:IMC

[PZ08] Yingwei Peng and Jiajia Zhang. Identifiability of a mixture cure frailty model. Statistics & Probability Letters, 78 (16):2604–2608, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208003714.

Pang:2004:FRO

[PZL04a] Shanqi Pang, Yingshan Zhang, and Sanyang Liu. Further results on the orthogonal arrays obtained by generalized Hadamard product. Statistics & Probability Letters, 68(1):17– 25, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000641.

Pang:2004:NMD

[PZL04b] Shanqi Pang, Yingshan Zhang, and Sanyang Liu. Normal mixed difference matrix and the construction of orthogonal arrays. Statistics & Probability Letters, 69(4):431–437, October 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001968.

Qin:2009:REL

[QBZ09] Guoyou Qin, Yang Bai, and Zhongyi Zhu. Robust empirical likelihood inference for longitudinal data. Statistics & Probability Letters, 79(20):2101–2108, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002454.

Qin:2004:SRG

[QC04] Hong Qin and Yin-Bao Chen. Some results on generalized minimum aberration for symmetrical fractional factorial designs. Statistics & Probability Letters, 66(1):51–57, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003067.

Qian:2002:LIL

[QF02] Guoqi Qian and Chris Field. Law of iterated logarithm and consistent model selection criterion in logistic regression. Statistics & Probability Letters, 56(1):101–112, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001912.

Qin:2003:DKE

[QF03] Huai-Zhen Qin and Shi-Yong Feng. Deconvolution kernel estimator for mean transformation with ordinary smooth error. Statistics & Probability Letters, 61(4):337–346, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001992.

Quang:2009:SLL

[QH09] Nguyen Van Quang and Nguyen Van Huan. On the strong law of large numbers and L_{p}-convergence for double arrays of random elements in p-uniformly smooth Banach spaces. Statistics & Probability Letters, 79(18):1891–1899, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001965.

Qi:2003:LDP

[Qi03] Yongcheng Qi. Limit distributions for products of sums. Statistics & Probability Letters, 62(1):93–100, March 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202004388.

Qiu:2006:VPS

[QL06] Jin Qiu and Zhengyan Lin. The variance of partial sums of strong near-epoch dependent variables. Statistics & Probability Letters, 76(17):1845–1854, November 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001507.

Qin:2009:ELL

[QLL09] Yongsong Qin, Ling Li, and Qingzhu Lei. Empirical likelihood for linear regression models with missing responses. Statistics & Probability Letters, 79(11):1391–1396, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209000972.

Qu:2009:CSS

[QO09] Xianggui Qu and Theophilus Ogunyemi. On the connectedness of saturated square designs. Statistics & Probability Letters, 79(2):211–214, January 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003702.

Qiao:2005:RWW

[QS05] Zhihua Qiao and J. Michael Steele. Random walks whose concave majorants often have few faces. Statistics & Probability Letters, 75(2):97–102, November 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002130.

Qu:2001:ECM

[QSP01] Roger P. Qu, Jun Shao, and Mari Palta. Efficiency comparison of methods for estimation in longitudinal regression models. Statistics & Probability Letters, 55(2):125–135, November 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000487.

Qin:2007:ELI

[QZ07] Gengsheng Qin and Yichuan Zhao. Empirical likelihood inference for the mean residual life under random censorship. Statistics & Probability Letters, 77(5):549–557, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002872.

Rodrigues:2008:PPB

[RA08] Alexandre Rodrigues and Renato Assunção. Propriety of posterior in Bayesian space varying parameter models with normal data. Statistics & Probability Letters, 78(15):2408–2411, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001417.

Rabinowitz:2000:TCS

[Rab00] Daniel Rabinowitz. Testing current status data for dependent censoring. Statistics & Probability Letters, 48(3):213–216, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299002023.

Radchenko:2007:BRS

[Rad07] Vadym M. Radchenko. Besov regularity of stochastic measures. Statistics & Probability Letters, 77(8):822–825, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000156.

Rajba:2009:NCS

[Raj09] Teresa Rajba. Nested classes of C-semi-selfdecomposable distributions. Statistics & Probability Letters, 79(24):2469– 2475, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003332.

Rao:2002:AET

[Rao02a] B. L. S. Prakasa Rao. Another Esseen-type inequality for multivariate probability density functions. Statistics & Probability Letters, 60(2):191–199, November 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003127.

Rao:2002:AWI

[Rao02b] B. L. S. Prakasa Rao. Application of Whittle’s inequality for Banach space valued martingales. Statistics & Probability Letters, 57(2):133–137, April 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200024X.

Rao:2002:HRT

[Rao02c] B. L. S. Prakasa Rao. Hajek–Rényi-type inequality for associated sequences. Statistics & Probability Letters, 57(2):139– 143, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202000251.

Raqab:2003:NBM

[Raq03] Mohammad Z. Raqab. P -Norm bounds for moments of progressive type II censored order statistics. Statistics & Probability Letters, 64(4):393–402, October 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002049.

Rahimov:2008:LBG

[RAS08a] I. Rahimov and W. S. Al-Sabah. Limiting behavior of a generalized branching process with immigration. Statistics & Probability Letters, 78(3):225–230, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002179.

Rassoul-Agha:2008:ASI

[RAS08b] F. Rassoul-Agha and T. Seppäläinen. An almost sure invariance principle for additive functionals of Markov chains. Statistics & Probability Letters, 78(7):854–860, May 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700291X.

Reynaud-Bouret:2006:CEI

[RB06] P. Reynaud-Bouret. Compensator and exponential inequalities for some suprema of counting processes. Statistics & Probability Letters, 76(14):1514–1521, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000939.

Raqab:2008:PIF

[RB08a] Mohammad Z. Raqab and N. Balakrishnan. Prediction intervals for future records. Statistics & Probability Letters, 78(13):1955–1963, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000643.

Roussas:2008:HIR

[RB08b] George G. Roussas and Debasis Bhattacharya. Hájek–Inagaki representation theorem, under a general stochastic processes framework, based on stopping times. Statistics & Probability Letters, 78(15):2503–2510, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001569.

Requena:2000:CMP

[RC00] F. Requena and N. Martin Ciudad. Characterization of maximum probability points in the Multivariate Hypergeometric distribution. Statistics & Probability Letters, 50(1):39–47, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000791.

Reinsel:2003:AMR

[RC03] Gregory C. Reinsel and Wai-Kwong Cheang. Approximate ML and REML estimation for regression models with spatial or time series AR(1) noise. Statistics & Probability Letters, 62(2):123–135, April 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202004376.

Rodrigues:2009:ULT

[RCdCLN09] Josemar Rodrigues, Vicente G. Cancho, Mário de Castro, and Francisco Louzada-Neto. On the unification of long-term survival models. Statistics & Probability Letters, 79(6):753–759, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005105.

Reiter:2008:SNI

[Rei08] Jerome P. Reiter. Selecting the number of imputed datasets when using multiple imputation for missing data and disclosure limitation. Statistics & Probability Letters, 78(1):15–20, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207001654.

Rekkas:2009:AIM

[Rek09] M. Rekkas. Approximate inference for the multinomial logit model. Statistics & Probability Letters, 79(2):237–242, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003829.

Ren:2001:BSP

[Ren01] Yanxia Ren. Boundary singularity problem of some nonlinear elliptic equations. Statistics & Probability Letters, 51 (2):173–179, January 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001504.

Ren:2008:BDG

[Ren08] Yao-Feng Ren. On the Burkholder–Davis–Gundy inequalities for continuous martingales. Statistics & Probability Letters, 78(17):3034–3039, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002642.

Ren:2009:CBD

[Ren09a] Jiandong Ren. A connection between the discounted and nondiscounted expected penalty functions in the Sparre Andersen risk model. Statistics & Probability Letters, 79(3):324–330, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004148.

Ren:2009:SON

[Ren09b] Yanbo Ren. Some Orlicz-norm inequalities for martingales. Statistics & Probability Letters, 79(9):1238–1241, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000388.

Rosenberger:2002:UGL

[RH02] William F. Rosenberger and Mingxiu Hu. On the use of generalized linear models following a sequential design. Statistics & Probability Letters, 56(2):155–161, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001717.

Ren:2007:RBS

[RH07] Yong Ren and Lanying Hu. Reflected backward stochastic differential equations driven by Lévy processes. Statistics & Probability Letters, 77(15):1559–1566, September 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001290.

Ristic:2005:BGA

[Ris05] Miroslav M. Ristić. A Beta–Gamma autoregressive process of the second-order (BGAR(2)). Statistics & Probability Letters, 73(4):403–410, July 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001446.

Ristic:2006:SBM

[Ris06] Miroslav M. Ristić. Stationary bivariate minification processes. Statistics & Probability Letters, 76(5):439–445, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003172.

Rivoirard:2004:MLP

[Riv04] Vincent Rivoirard. Maxisets for linear procedures. Statistics & Probability Letters, 67(3):267–275, April 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000380.

Ren:2001:BCM

[RL01] Yao-Feng Ren and Han-Ying Liang. On the best constant in Marcinkiewicz–Zygmund inequality. Statistics & Probability Letters, 53(3):227–233, June 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000153.

Roy:2008:LRT

[RL08] Anuradha Roy and Ricardo Leiva. Likelihood ratio tests for triply multivariate data with structured correlation on spatial repeated measurements. Statistics & Probability Letters, 78(13):1971–1980, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000667.

Rodriguez-Lallena:2003:DFM

[RLÚF03] José A. Rodríguez-Lallena and Manuel Úbeda-Flores. Distribution functions of multivariate copulas. Statistics & Probability Letters, 64(1):41–50, August 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001299.

Rodriguez-Lallena:2004:NCB

[RLÚF04] José Antonio Rodríguez-Lallena and Manuel Úbeda-Flores. A new class of bivariate copulas. Statistics & Probability Letters, 66(3):315–325, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003584.

Rahimov:2003:EPM

[RM03] I. Rahimov and H. A. Muttlak. Estimation of the population mean using random selection in ranked set samples. Statistics & Probability Letters, 62(2):203–209, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000130.

Reyen:2005:MIL

[RM05] Salem S. Reyen and John J. Miller. The moment of inertia and the linear discriminant function. Statistics & Probability Letters, 71(1):39–46, January 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002822.

Ren:2008:NEL

[RM08] Qunshu Ren and Majid Mojirsheibani. Nonparametric estimation of level sets under minimal assumptions. Statistics & Probability Letters, 78(17):3029–3033, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002630.

Ruiz-Medina:2001:SFO

[RMAA01] M. D. Ruiz-Medina, V. V. Anh, and J. M. Angulo. Stochastic fractional-order differential models with fractal boundary conditions. Statistics & Probability Letters, 54(1):47–60, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000438.

Rodriguez-Muniz:2007:EIE

[RMLD07] Luis J. Rodríguez-Muñiz and Miguel López-Díaz. On the exchange of iterated expectations of random upper semicontinuous functions. Statistics & Probability Letters, 77 (16):1628–1635, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001447.

Ruiz:2000:NLC

[RN00] Edilberto Ruiz and Fabio H. Nieto. A note on linear combination of predictors. Statistics & Probability Letters, 47(4):351– 356, May 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001777.

Rodrigues:2000:NAD

[Rod00] Paulo M. M. Rodrigues. A note on the application of the DF test to seasonal data. Statistics & Probability Letters, 47(2):171–175, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001546.

Rollin:2008:NEC

[Röl08] Adrian Röllin. A note on the exchangeability condition in Stein’s method. Statistics & Probability Letters, 78(13):1800–1806, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000424.

Romain:2002:PFT

[Rom02] Yves Romain. Perturbation of functional tensors with applications to covariance operators. Statistics & Probability Letters, 58(3):253–264, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001049.

Romanazzi:2009:DDR

[Rom09] Mario Romanazzi. Data depth, random simplices and multivariate dispersion. Statistics & Probability Letters, 79(12): 1473–1479, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001114.

Roos:2001:SCP

[Roo01] Bero Roos. Sharp constants in the Poisson approximation. Statistics & Probability Letters, 52(2):155–168, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000208X.

Rosenblatt:2006:ETF

[Ros06] M. Rosenblatt. An example and transition function equicontinuity. Statistics & Probability Letters, 76(18):1961–1964, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001647.

Rosadi:2007:IMA

[Ros07] Dedi Rosadi. Identification of moving average process with infinite variance. Statistics & Probability Letters, 77(14): 1490–1496, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520700079X.

Rosenbaum:2009:FOV

[Ros09a] Mathieu Rosenbaum. First order p-variations and Besov spaces. Statistics & Probability Letters, 79(1):55–62, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003441.

Rosenblatt:2009:CCW

[Ros09b] M. Rosenblatt. A comment on a conjecture of N. Wiener. Statistics & Probability Letters, 79(3):347–348, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208004173.

Roussas:2000:ANK

[Rou00] George G. Roussas. Asymptotic normality of the kernel estimate of a probability density function under association. Statistics & Probability Letters, 50(1):1–12, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000729.

Roussas:2001:EET

[Rou01a] G. G. Roussas. Erratum: “An Esseen-type inequality for probability density functions, with an application” [Statist. Probab. Lett. 51 (2001) 397–408]. Statistics & Probability Letters, 54(4):449, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000980. See [Rou01b].

Roussas:2001:ETI

[Rou01b] George G. Roussas. An Esseen-type inequality for probability density functions, with an application. Statistics & Probability Letters, 51(4):397–408, February 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001814. See erratum [Rou01a].

Rousson:2007:GCR

[Rou07] Valentin Rousson. The gamma coefficient revisited. Statistics & Probability Letters, 77(17):1696–1704, November 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700154X.

Rozkosz:2004:ESB

[Roz04]  Andrzej Rozkosz.  On existence of solutions of BS-

DEs with continuous coefficient. Statistics & Probability Letters, 67(3):249–256, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000318.

Rozovsky:2009:RLB

[Roz09a] Leonid Rozovsky. Remarks on a link between the Laplace transform and distribution function of a nonnegative random variable. Statistics & Probability Letters, 79(13):1501–1508, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001229.

Rozovsky:2009:SDS

[Roz09b] Leonid Rozovsky. Small deviations of series of weighted i.i.d. non-negative random variables with a positive mass at the origin. Statistics & Probability Letters, 79(13):1495–1500, July 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209001023.

Ristic:2002:UAP

[RP02] Miroslav M. Ristić and Biljana C. Popović. The uniform autoregressive process of the second order (UAR(2)). Statistics & Probability Letters, 57(2):113–119, April 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000536.

Rinott:2001:RQN

[RR01] Yosef Rinott and Vladimir Rotar. A remark on quadrant normal probabilities in high dimensions. Statistics & Probability Letters, 51(1):47–51, January 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001413.

Rachdi:2000:CEM

[RS00] Mustapha Rachdi and Rachid Sabre. Consistent estimates of the mode of the probability density function in nonparametric deconvolution problems. Statistics & Probability Letters, 47(2):105–114, April 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001455.

Ramos:2003:DMD

[RS03a] Héctor M. Ramos and Miguel A. Sordo. Dispersion measures and dispersive orderings. Statistics & Probability Letters, 61(2):123–131, January 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003413.

Rychlik:2003:SVC

[RS03b] Zdzisław Rychlik and Konrad S. Szuster. On strong versions of the Central Limit Theorem. Statistics & Probability Letters, 61(4):347–357, February 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003620.

Radjavi:2006:FAE

[RS06a] H. Radjavi and C. C. A. Sastri. A function arising in the estimation of unobserved probability. Statistics & Probability Letters, 76(7):682–688, April 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520500369X.

Rafajlowicz:2006:HHS

[RS06b] Ewaryst Rafajlowicz and Rainer Schwabe. Halton and Hammersley sequences in multivariate nonparametric regression. Statistics & Probability Letters, 76(8):803–812, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003895.

Ramprasath:2007:SOC

[RS07] L. Ramprasath and Kesar Singh. Statistical options: Crash resistant financial contracts based on robust estimation. Statistics & Probability Letters, 77(2):196–203, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002227.

Russo:2008:CEM

[RS08] Ralph P. Russo and Nariankadu D. Shyamalkumar. On the convergence of the empirical mass function. Statistics & Probability Letters, 78(15):2293–2299, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001235.

Ren:2002:ISP

[RSY02] Yan-Xia Ren, Xiu-Yun Suo, and Xiang-Dong Yu. Interior singularity problem of some nonlinear elliptic equations. Statistics & Probability Letters, 59(3):317–328, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002055.

Rukhin:2003:CIM

[Ruk03] Andrew. L. Rukhin. Covariance identity for multinomial trials. Statistics & Probability Letters, 63(1):107–112, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000610.

Rukhin:2007:CCI

[Ruk07a] Andrew L. Rukhin. Conservative confidence intervals based on weighted means statistics. Statistics & Probability Letters, 77(8):853–861, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000247.

Rukhin:2007:NOS

[Ruk07b] Andrew L. Rukhin. Normal order statistics and sums of geometric random variables in treatment allocation problems. Statistics & Probability Letters, 77(12):1312–1321, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000946.

Rukhin:2009:INM

[Ruk09] Andrew L. Rukhin. Identities for negative moments of quadratic forms in normal variables. Statistics & Probability Letters, 79(8):1004–1007, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005671.

Rukhin:2008:CCS

[RV08] Andrew L. Rukhin and Ronny Vallejos. Codispersion coefficients for spatial and temporal series. Statistics & Probability Letters, 78(11):1290–1300, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004087.

Romano:2000:MGC

[RW00] Joseph P. Romano and Michael Wolf. A more general Central Limit Theorem for m-dependent random variables with unbounded m. Statistics & Probability Letters, 47(2):115–124, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001467.

Rempala:2005:API

[RW05] Grzegorz Rempala and Jacek Wesołowski. Asymptotics for products of independent sums with an application to Wishart determinants. Statistics & Probability Letters, 74(2):129– 138, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001665.

Rychlik:2002:BUQ

[Ryc02] Tomasz Rychlik. Best upper quantile evaluations for NWU distributions. Statistics & Probability Letters, 58(2):175–184, June 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001232.

Rychlik:2008:EVO

[Ryc08]  Tomasz Rychlik.  Extreme variances of order statis-

tics in dependent samples. Statistics & Probability Letters, 78(12):1577–1582, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800014X.

Rychlik:2009:TEE

[Ryc09] Tomasz Rychlik. Tight evaluations for expectations of small order statistics from symmetric and symmetric unimodal populations. Statistics & Probability Letters, 79(12):1488–1493, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001138.

Salopek:2002:NCN

[Sal02] D. M. Salopek. A new class of nearly self-financing strategies. Statistics & Probability Letters, 56(1):69–75, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001730.

Sancetta:2005:DBN

[San05] Alessio Sancetta. Distance between nonidentically weakly dependent random vectors and Gaussian random vectors under the bounded Lipschitz metric. Statistics & Probability Letters, 75(3):158–168, December 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002221.

Sapozhnikov:2008:CRL

[Sap08] Artëm Sapozhnikov. Convergence rates in the local renewal theorem. Statistics & Probability Letters, 78(14):2230–2233, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001156.

Sarkar:2000:AMB

[Sar00] Nityananda Sarkar. Arch model with Box–Cox transformed dependent variable. Statistics & Probability Letters, 50(4): 365–374, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200001176.

Sasvari:2008:GCF

[Sas08] Zoltán Sasvári. On generalized correlation functions of intrinsically stationary processes of order k. Statistics & Probability Letters, 78(18):3381–3387, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003301.

Sakov:2000:EEB

[SB00] Anat Sakov and Peter J. Bickel. An Edgeworth expansion for the m out of n bootstrapped median. Statistics & Probability Letters, 49(3):217–223, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000050X.

Salibian-Barrera:2006:BME

[SB06] Matias Salibian-Barrera. Bootstrapping MM-estimators for linear regression with fixed designs. Statistics & Probability Letters, 76(12):1287–1297, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000265.

Subramanian:2008:SLT

[SB08] Sundarraman Subramanian and Dipankar Bandyopadhyay. Semiparametric left truncation and right censorship models with missing censoring indicators. Statistics & Probability Letters, 78(16):2572–2577, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003477.

Skare:2000:SLP

[SBF00] Øivind Skare, Fred Espen Benth, and Arnoldo Frigessi. Smoothed Langevin proposals in Metropolis–Hastings algorithms. Statistics & Probability Letters, 49(4):345–354, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000675.

Sashegyi:2001:CBP

[SBF01] Andreas I. Sashegyi, K. Stephen Brown, and Patrick J. Farrell. On the correspondence between population-averaged models and a class of cluster-specific models for correlated binary data. Statistics & Probability Letters, 52(2):135–144, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001838.

Stepanov:2003:EDF

[SBH03] A. V. Stepanov, N. Balakrishnan, and Glenn Hofmann. Exact distribution and Fisher information of weak record values. Statistics & Probability Letters, 64(1):69–81, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001329.

Seneta:2005:GEU

[SC05] Eugene Seneta and John T. Chen. A generator for explicit univariate lower bounds. Statistics & Probability Letters, 75(4):256–266, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002397.

Shang:2008:ADB

[SC08] Junfeng Shang and Joseph E. Cavanaugh. An assumption for the development of bootstrap variants of the Akaike information criterion in mixed models. Statistics & Probability Letters, 78(12):1422–1429, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004245.

Scheike:2000:CNP

[Sch00] Thomas H. Scheike. Comparison of non-parametric regression functions through their cumulatives. Statistics & Probability Letters, 46(1):21–32, January 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000838.

Schick:2001:ADA

[Sch01a] Anton Schick. On asymptotic differentiability of averages. Statistics & Probability Letters, 51(1):15–23, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001322.

Schlather:2001:ECT

[Sch01b] Martin Schlather. Examples for the coefficient of tail dependence and the domain of attraction of a bivariate extreme value distribution. Statistics & Probability Letters, 53(3):325–329, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000906.

Schott:2002:TES

[Sch02] James R. Schott. Testing for elliptical symmetry in covariance-matrix-based analyses. Statistics & Probability Letters, 60(4):395–404, December 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202003061.

Schott:2003:DSR

[Sch03] James R. Schott. Distributions on spheres and random variables distributed on the interval (-1,1). Statistics & Probability Letters, 62(4):391–396, May 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000439.

Schwingenschlogl:2007:PMM

[Sch07] Udo Schwingenschlögl. Probabilities of majority and minority violation in proportional representation. Statistics & Probability Letters, 77(17):1690–1695, November 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001538.

Schott:2008:TIT

[Sch08] James R. Schott. A test for independence of two sets of variables when the number of variables is large relative to the sample size. Statistics & Probability Letters, 78(17):3096–3102, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002757.

Scotto:2003:NAD

[Sco03] M. G. Scotto. A note on the asymptotic distribution of the maxima in disaggregated time-series models. Statistics & Probability Letters, 65(2):127–137, November 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002487.

Sutradhar:2001:HOA

[SD01] Brajendra C. Sutradhar and Kalyan Das. A higher-order approximation to likelihood inference in the Poisson mixed model. Statistics & Probability Letters, 52(1):59–67, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200002157.

Schwingenschlogl:2006:SEV

[SD06] Udo Schwingenschlögl and Mathias Drton. Seat excess variances of apportionment methods for proportional representation. Statistics & Probability Letters, 76(16):1723–1730, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520600126X.

Svensson:2009:LDH

[SD09] Jens Svensson and Boualem Djehiche. Large deviations for heavy-tailed factor models. Statistics & Probability Letters, 79(3):304–311, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004021.

Satten:2001:EMS

[SDR01] Glen A. Satten, Somnath Datta, and James Robins. Estimating the marginal survival function in the presence of time dependent covariates. Statistics & Probability Letters, 54(4):397–403, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001134.

Dong:2005:MDP

[sDXlyY05] Zhi shan Dong, Tan Xi-li, and Xiao yun Yang. Moderate deviation principles for moving average processes of real stationary sequences. Statistics & Probability Letters, 74(2):139–150, September 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001707.

Segers:2001:NUL

[Seg01] Johan Segers. On the normal uniform local domain of attraction for intermediate order statistics. Statistics & Probability Letters, 53(4):409–413, July 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000633.

Segers:2005:ADC

[Seg05] Johan Segers. Approximate distributions of clusters of extremes. Statistics & Probability Letters, 74(4):330–336, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001926.

Sekine:2008:MDS

[Sek08] Jun Sekine. Marginal distribution of some path-dependent stochastic volatility model. Statistics & Probability Letters, 78(13):1846–1850, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000497.

Sen:2001:BDF

[Sen01] Amit Sen. Behaviour of Dickey–Fuller F -tests under the trend-break stationary alternative. Statistics & Probability Letters, 55(3):257–268, December 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001468.

Sen:2007:DDF

[Sen07] Amit Sen. On the distribution of Dickey–Fuller unit root statistics when there is a break in the innovation variance. Statistics & Probability Letters, 77(1):63–68, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001957.

Sen:2008:BDF

[Sen08a] Amit Sen. Behaviour of Dickey–Fuller tests when there is a break under the unit root null hypothesis. Statistics & Probability Letters, 78(6):622–628, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003045.

Sengupta:2008:MPT

[Sen08b] Arindam Sengupta. Markov processes, time-space harmonic functions and polynomials. Statistics & Probability Letters, 78(18):3277–3280, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003143.

Sen:2009:URT

[Sen09] Amit Sen. Unit root tests in the presence of an innovation variance break that has power against the mean break stationary alternative. Statistics & Probability Letters, 79(3):354–360, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004197.

Serdobolskii:2000:NMD

[Ser00] V. I. Serdobolskii. Normal model for distribution-free multivariate analysis. Statistics & Probability Letters, 48(4): 353–360, July 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520000016X.

Serdobolskii:2002:USS

[Ser02] A. V. Serdobolskii. Unimprovable solution to systems of empirical linear algebraic equations. Statistics & Probability Letters, 60(1):1–6, November 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001487.

Su:2012:EMG

[SF12] Jianxi Su and Edward Furman. Erratum to “On a multivariate Gamma distribution” by E. Furman [Statist. Probab. Lett. 78 (2008) 2353–2360]. Statistics & Probability Letters, 82(5):1040–1041, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715212000399. See [Fur08].

Sinha:2003:REN

[SFS03] Sanjoy K. Sinha, Christopher A. Field, and Bruce Smith. Robust estimation of nonlinear regression with autoregressive errors. Statistics & Probability Letters, 63(1):49–59, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520300052X.

Schlather:2006:LAV

[SG06] Martin Schlather and Tilmann Gneiting. Local approximation of variograms by covariance functions. Statistics & Probability Letters, 76(12):1303–1304, July 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000393.

Sgibnev:2001:EABa

[Sgi01a]  M. S. Sgibnev.  Exact asymptotic behaviour of the

distribution of the supremum. Statistics & Probability Letters, 52(3):301–311, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002297.

Sgibnev:2001:EABb

[Sgi01b] M. S. Sgibnev. On the exact asymptotic behaviour of the distribution of the supremum in the “critical” case. Statistics & Probability Letters, 54(4):357–362, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000189.

Sgibnev:2009:RFW

[Sgi09] M. S. Sgibnev. On a renewal function when the second moment is infinite. Statistics & Probability Letters, 79(9): 1242–1245, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520900039X.

Singh:2009:THS

[SGK09] Parminder Singh, Amar Nath Gill, and Narendra Kumar. A test of homogeneity of several variances against tree ordered alternative. Statistics & Probability Letters, 79(22):2315–2320, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209003009.

Shen:2006:ELD

[SH06] Junshan Shen and Shuyuan He. Empirical likelihood for the difference of two survival functions under right censorship. Statistics & Probability Letters, 76(2):169–181, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002865.

Sen:2009:NTR

[SH09a] Rituparna Sen and Fushing Hsieh. A note on testing regime switching assumption based on recurrence times. Statistics & Probability Letters, 79(24):2443–2450, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003277.

Soltani:2009:WAR

[SH09b] A. R. Soltani and H. Homei. Weighted averages with random proportions that are jointly uniformly distributed over the unit simplex. Statistics & Probability Letters, 79(9): 1215–1218, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000340.

Shao:2001:RCB

[Sha01] Yongzhao Shao. Rate of convergence of bootstrapped empirical measures. Statistics & Probability Letters, 53(3):293–298, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000736.

Shapiro:2003:SMC

[Sha03] Alexander Shapiro. Scheffe’s method for constructing simultaneous confidence intervals subject to cone constraints.

Statistics & Probability Letters, 64(4):403–406, October 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002050.

Shaw:2004:RFL

[Sha04] Jason T. Shaw. Random fields and the limit of their spectral densities: existence and bounds. Statistics & Probability Letters, 67(3):213–220, April 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003894.

Shanchao:2006:MIS

[Sha06a] Yang Shanchao. Moment inequalities for sums of products of independent random variables. Statistics & Probability Letters, 76(18):1994–2000, December 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001830.

Shao:2006:MPA

[Sha06b] Q. Shao. Mixture periodic autoregressive time series models. Statistics & Probability Letters, 76(6):609–618, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003627.

Shaked:2007:SCM

[Sha07] Moshe Shaked. Stochastic comparisons of multivariate random sums in the Laplace transform order, with applications. Statistics & Probability Letters, 77(12):1339–1344, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000971.

Shapiro:2008:AMS

[Sha08a] Alexander Shapiro. Asymptotics of minimax stochastic programs. Statistics & Probability Letters, 78(2):150–157, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002064.

Shashkin:2008:SIP

[Sha08b] Alexey Shashkin. A strong invariance principle for positively or negatively associated random fields. Statistics & Probability Letters, 78(14):2121–2129, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000862.

Shen:2008:AOM

[She08] Gang Shen. Asymptotics of Oja Median Estimate. Statistics & Probability Letters, 78(14):2137–2141, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000886.

Shen:2009:ATT

[She09] Gang Shen. Asymptotics of a Theil-type estimate in multiple linear regression. Statistics & Probability Letters, 79 (8):1053–1064, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005750.

Shih:2001:SBS

[Shi01] Yu-Shan Shih. Selecting the best splits for classification trees with categorical variables. Statistics & Probability Letters, 54(4):341–345, October 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001887.

Shikimi:2002:LDK

[Shi02] Takuhisa Shikimi. Large deviations for kernel-type empirical distributions. Statistics & Probability Letters, 59(1):23–28, August 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001499.

Shin:2004:ESD

[Shi04a] Dong Wan Shin. Estimation of spectral density for seasonal time series models. Statistics & Probability Letters, 67(2):149– 159, April 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000240.

Shixin:2004:ASC

[Shi04b] Gan Shixin. Almost sure convergence for ρ– d-mixing random variable sequences. Statistics & Probability Letters, 67(4):289– 298, May 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000355.

Shixin:2004:MIV

[Shi04c] Gan Shixin. Moment inequalities for B-valued random vectors with applications to the strong limit theorems. Statistics & Probability Letters, 67(2):111–119, April 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000021.

Shimizu:2009:NDE

[Shi09] Yasutaka Shimizu. Notes on drift estimation for certain nonrecurrent diffusion processes from sampled data. Statistics & Probability Letters, 79(20):2200–2207, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002697.

Shmileva:2006:SBP

[Shm06] Elena Shmileva. Small ball probabilities for jump Lévy processes from the Wiener domain of attraction. Statistics & Probability Letters, 76(17):1873–1881, November 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001532.

Shumway:2003:TFC

[Shu03] Robert H. Shumway. Time-frequency clustering and discriminant analysis. Statistics & Probability Letters, 63(3):307–314, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000956.

Shuhe:2004:CLS

[Shu04] Hu Shuhe. Consistency for the least squares estimator in nonlinear regression model. Statistics & Probability Letters, 67(2):183–192, April 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204000343.

Shu:2009:LLN

[Shu09] F. C. Shu. A law of large numbers and central limit theorem for the logarithm of an autoregressive process with a stationary driving sequence. Statistics & Probability Letters, 79(9):1141–1145, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005178.

Sung:2005:WLA

[SHV05] Soo Hak Sung, Tien-Chung Hu, and Andrei Volodin. On the weak laws for arrays of random variables. Statistics & Probability Letters, 72(4):291–298, May 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000416.

Siatkowski:2000:SPM

[Sia00] Idzi Siatkowski. Some properties of multi-way block designs. Statistics & Probability Letters, 46(3):277–282, February 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001133.

Sim:2000:TSC

[Sim00] Songyong Sim. A test for spatial correlation for binary data. Statistics & Probability Letters, 47(2):129–134, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001480.

Simonelli:2001:CSI

[Sim01] Italo Simonelli. Convergence and symmetry of infinite products of independent random variables. Statistics & Probability Letters, 55(1):45–52, November 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001262.

Simic:2006:EVE

[Sim06] Slavko Simic. Extreme values and entire functions of finite order. Statistics & Probability Letters, 76(7):703–708, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003780.

Sim:2007:TBQ

[Sim07] Songyong Sim. A test based on quantile score for csample randomized block design. Statistics & Probability Letters, 77(12):1345–1353, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000983.

Simic:2008:NMI

[Sim08] Slavko Simic. On a new moments inequality. Statistics & Probability Letters, 78(16):2671–2678, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001831.

Simic:2009:MIL

[Sim09] Slavko Simic. On a moment inequality for laws on (−∞, +∞). Statistics & Probability Letters, 79(22):2334–2337, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003034.

Singpurwalla:2007:BRL

[Sin07] Nozer D. Singpurwalla. Betting on residual life: The caveats of conditioning. Statistics & Probability Letters, 77(12): 1354–1361, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000995.

Situ:2002:SBS

[Sit02] Rong Situ. On solutions of backward stochastic differential equations with jumps and with non-Lipschitzian coefficients in Hilbert spaces and stochastic control. Statistics & Probability Letters, 60(3):279–288, December 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002857.

Sutradhar:2001:EEG

[SK01] Brajendra C. Sutradhar and Pranesh Kumar. On the efficiency of extended generalized estimating equation approaches. Statistics & Probability Letters, 55(1):53–61, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001274.

Sam:2006:NRU

[SK06a] Abdoul G. Sam and Alan P. Ker. Nonparametric regression under alternative data environments. Statistics & Probability Letters, 76(10):1037–1046, May 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004463.

Shinde:2006:JDR

[SK06b] R. L. Shinde and K. S. Kotwal. On the joint distribution of runs in the sequence of Markov-dependent multi-state trials. Statistics & Probability Letters, 76(10):1065–1074, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004505.

Shin:2007:AEO

[SKJ07] Dong Wan Shin, Han Joon Kim, and Won-Chul Jhee. Asymptotic efficiency of the ordinary least-squares estimator for sur models with integrated regressors. Statistics & Probability Letters, 77(1):75–82, January 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001970.

Sim:2003:TSQ

[SKPP03] Songyong Sim, Seongbum Kim, Gilju Park, and Jaesoon Park. A two sample quantile test with applications to randomized block designs. Statistics & Probability Letters, 61(4):439–445, February 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202004194.

Strzalkowska-Kominiak:2009:MRL

[SKS09] E. Strzalkowska-Kominiak and W. Stute. Martingale representations of the Lynden–Bell estimator with applications. Statistics & Probability Letters, 79(6):814–820, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005208.

Sivaganesan:2000:BFT

[SL00] S. Sivaganesan and Rama T. Lingham. Bayes factors for a test about the drift of the Brownian motion under noninformative priors. Statistics & Probability Letters, 48(2):163–171, June 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001996.

Sarkar:2006:LAA

[SL06] Jyotirmoy Sarkar and Fang Li. Limiting average availability of a system supported by several spares and several repair facilities. Statistics & Probability Letters, 76(18):1965–1974, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001660.

Shen:2008:PLD

[SL08] Xinmei Shen and Zhengyan Lin. Precise large deviations for randomly weighted sums of negatively dependent random variables with consistently varying tails. Statistics & Probability Letters, 78(18):3222–3229, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003027.

Sakthivel:2009:ASN

[SL09a] R. Sakthivel and J. Luo. Asymptotic stability of nonlinear impulsive stochastic differential equations. Statistics & Probability Letters, 79(9):1219–1223, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000352.

Stoica:2009:EGM

[SL09b] George Stoica and Deli Li. Expected gains in the MacQueen– Heyde model. Statistics & Probability Letters, 79(14):1634– 1636, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001497.

Sarkar:2009:PCM

[SLC09] Angshuman Sarkar, Dennis K. J. Lin, and Kashinath Chatterjee. Probability of correct model identification in supersaturated design. Statistics & Probability Letters, 79(9):1224–1230, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209000364.

Su:2006:NDR

[SLF06] Chun Su, Jie Liu, and Qunqiang Feng. A note on the distance in random recursive trees. Statistics & Probability Letters, 76(16):1748–1755, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001374.

Song:2009:FET

[SLF09] Lixin Song, Dawei Lu, and Jinghai Feng. The first exit time for a Bessel process from the minimum and maximum random domains. Statistics & Probability Letters, 79(20):2115–2123, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520900248X.

Shuhe:2006:GAR

[SM06a] Hu Shuhe and Hu Ming. A general approach rate to the Strong Law of Large Numbers. Statistics & Probability Letters, 76(8):843–851, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004013.

Slakmon:2006:ASC

[SM06b] Alain Slakmon and Luc Macot. On the almost sure convergence of Syracuse sequences. Statistics & Probability Letters, 76(15):1625–1630, September 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001167.

Serfling:2009:EPI

[SM09] Robert Serfling and Satyaki Mazumder. Exponential probability inequality and convergence results for the median absolute deviation and its modifications. Statistics & Probability Letters, 79(16):1767–1773, August 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001734.

Shao:2004:LSE

[SN04] Q. Shao and P. P. Ni. Least-squares estimation and ANOVA for periodic autoregressive time series. Statistics & Probability Letters, 69(3):287–297, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001786.

Senturk:2009:APC

[SN09] Damla Sentürk and Danh V. Nguyen. Asymptotic properties of covariate-adjusted regression with correlated errors. Statistics & Probability Letters, 79(9):1175–1180, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900011X.

Shin:2000:STS

[SO00] Dong Wan Shin and Man-Suk Oh. Semiparametric tests for seasonal unit roots based on a semiparametric feasible GLSE. Statistics & Probability Letters, 50(3):207–218, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001061.

So:2003:MLL

[So03] Beong Soo So. Maximized log-likelihood updating and model selection. Statistics & Probability Letters, 64(3):293– 303, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001743.

Song:2000:LTN

[Son00] Kai-Sheng Song. Limit theorems for nonparametric sample entropy estimators. Statistics & Probability Letters, 49(1):9–18, August 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000250.

Sonin:2008:GGI

[Son08] Isaac M. Sonin. A generalized Gittins index for a Markov chain and its recursive calculation. Statistics & Probability Letters, 78(12):1526–1533, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000072.

Song:2009:LFT

[Son09] Weixing Song. Lack-of-fit testing in errors-in-variables regression model with validation data. Statistics & Probability Letters, 79(6):765–773, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005129.

Sordo:2009:RLI

[Sor09] Miguel A. Sordo. On the relationship of location-independent riskier order to the usual stochastic order. Statistics & Probability Letters, 79(2):155–157, January 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003593.

Soulier:2001:MBC

[Sou01] Philippe Soulier. Moment bounds and Central Limit Theorem for functions of Gaussian vectors. Statistics & Probability Letters, 54(2):193–203, September 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100061X.

Stoyanov:2000:RMC

[SP00] Jordan Stoyanov and Christo Pirinsky. Random motions, classes of ergodic Markov chains and beta distributions. Statistics & Probability Letters, 50(3):293–304, November 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001140.

Spouge:2007:IOB

[Spo07] John L. Spouge. Inequalities on the overshoot beyond a boundary for independent summands with differing distributions. Statistics & Probability Letters, 77(14):1486–1489, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000776.

Spreij:2001:MPF

[Spr01] Peter Spreij. On the Markov property of a finite hidden Markov chain. Statistics & Probability Letters, 52(3):279–288, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002169.

Spreij:2003:HMC

[Spr03] Peter Spreij. On hidden Markov chains and finite stochastic systems. Statistics & Probability Letters, 62(2):189–201, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203000117.

Stute:2006:MCH

[SQMK06] W. Stute, M. Presedo Quindimil, W. González Manteiga, and H. L. Koul. Model checks of higher order time series. Statistics & Probability Letters, 76(13):1385–1396, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206000460.

Szekely:2004:MDT

[SR04] Gábor J. Székely and Maria L. Rizzo. Mean distance test of Poisson distribution. Statistics & Probability Letters, 67 (3):241–247, April 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715204000288.

Shen:2000:TSF

[sS00a] Pao sheng Shen. Testing for sufficient follow-up in survival data. Statistics & Probability Letters, 49(4):313–322, October 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000638.

Shin:2000:AEO

[SS00b] Dong Wan Shin and Seuck Heun Song. Asymptotic efficiency of the OLSE for polynomial regression models with spatially correlated errors. Statistics & Probability Letters, 47(1):1–10, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299000929.

Sarkar:2001:API

[SS01a] Jyotirmoy Sarkar and Sahadeb Sarkar. Availability of a periodically inspected system supported by a spare unit, under perfect repair or perfect upgrade. Statistics & Probability Letters, 53(2):207–217, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000876.

Schlittgen:2001:ADI

[SS01b] Rainer Schlittgen and Rainer Schwabe. An alternative definition of the influence function. Statistics & Probability Letters, 51(2):143–153, January 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001449.

Schmuland:2001:E

[SS01c] Byron Schmuland and Wei Sun. On the equation μ_{t+s} = μ_{s} ∗ T_{s}μ_{t}. Statistics & Probability Letters, 52(2):183–188, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002352.

Schechtman:2002:CTS

[SS02]  E. Schechtman and G. Schechtman.  On characteri-

zation of two-sample U-statistics. Statistics & Probability Letters, 58(1):53–59, May 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001104.

Srivastav:2003:OCC

[SS03] Sudesh K. Srivastav and Arti Shankar. On the E-optimality of certain class of block designs. Statistics & Probability Letters, 65(3):187–193, November 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002505.

Sneddon:2004:SFL

[SS04] Gary Sneddon and Brajendra C. Sutradhar. On semiparametric familial-longitudinal models. Statistics & Probability Letters, 69(3):369–379, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001907.

Saigo:2005:JVE

[SS05a] H. Saigo and R. R. Sitter. Jackknife variance estimator with reimputation for randomly imputed survey data. Statistics & Probability Letters, 73(3):321–331, July 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001264.

Sun:2005:ECD

[SS05b] Xiaoqian Sun and Dongchu Sun. Estimation of the Cholesky decomposition of the covariance matrix for a conditional independent normal model. Statistics & Probability Letters, 73(1):1–12, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205000179.

Satheesh:2006:MSS

[SS06a] S. Satheesh and E. Sandhya. Max-semi-selfdecomposable laws and related processes. Statistics & Probability Letters, 76 (14):1435–1440, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206000769.

Shykula:2006:SSA

[SS06b] Mykola Shykula and Oleg Seleznjev. Stochastic structure of asymptotic quantization errors. Statistics & Probability Letters, 76(5):453–464, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003251.

Silva:2006:ADY

[SS06c] Isabel Silva and M. Eduarda Silva. Asymptotic distribution of the Yule–Walker estimator for INAR(p) processes. Statistics & Probability Letters, 76(15):1655–1663, September 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001210.

Schechtman:2007:MEM

[SS07a] E. Schechtman and C. Spiegelman. Mitigating the effect of measurement errors in quantile estimation. Statistics & Probability Letters, 77(5):514–524, March 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002720.

Schmid:2007:MES

[SS07b] Friedrich Schmid and Rafael Schmidt. Multivariate extensions of Spearman’s rho and related statistics. Statistics & Probability Letters, 77(4):407–416, February 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002586.

Schmidt:2007:SNP

[SS07c] Thorsten Schmidt and Winfried Stute. Shot-noise processes and the minimal martingale measure. Statistics & Probability Letters, 77(12):1332–1338, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520700096X.

Srivastav:2007:CEC

[SS07d] Sudesh K. Srivastav and Arti Shankar. On the construction and existence of a certain class of complete diallel cross designs. Statistics & Probability Letters, 77(1):111–115, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002033.

Samaniego:2008:SWC

[SS08a] Francisco J. Samaniego and Moshe Shaked. Systems with weighted components. Statistics & Probability Letters, 78(6): 815–823, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003379.

Sarnowski:2008:LDP

[SS08b] Wojciech Sarnowski and Krzysztof Szajowski. On-line detection of a part of a sequence with unspecified distribution. Statistics & Probability Letters, 78(15):2511–2516, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001570.

Subbiah:2008:CSD

[SS08c] M. Subbiah and M. R. Srinivasan. Classification of 2 × 2 sparse data sets with zero cells. Statistics & Probability Letters, 78(18):3212–3215, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003003.

Shaffer:2009:STC

[SS09a] Jeffrey G. Shaffer and Sudesh K. Srivastav. A simple technique for constructing optimal complete diallel cross designs. Statistics & Probability Letters, 79(9):1181–1185, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000121.

Shen:2009:HRL

[sS09b] Pao sheng Shen. Hazards regression for length-biased and right-censored data. Statistics & Probability Letters, 79(4):

457–465, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004513.

Shen:2009:IPW

[sS09c] Pao sheng Shen. An inverse-probability-weighted approach to the estimation of distribution function with doubly censored data. Statistics & Probability Letters, 79(9):1269–1276, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209000431.

Soltani:2009:CDD

[SSA09] A. R. Soltani, A. Shirvani, and F. Alqallaf. A class of discrete distributions induced by stable laws. Statistics & Probability Letters, 79(14):1608–1614, July 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001382.

Sengupta:2009:EML

[SSB09] Samindra Nath Sengupta, Bikas K. Sinha, and Srijib B. Bagchi. Estimation of mean life and reliability for exponential life distribution using time censored sample data. Statistics & Probability Letters, 79(7):899–906, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005336.

Satheesh:2008:GEA

[SSR08] S. Satheesh, E. Sandhya, and K. E. Rajasekharan. A generalization and extension of an autoregressive model. Statistics & Probability Letters, 78(12):1369–1374, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207004178.

Shanqi:2002:SOA

[SSY02] Pang Shanqi, Liu Sanyang, and Zhang Yingshan. Satisfactory orthogonal array and its checking method. Statistics & Probability Letters, 59(1):17–22, August 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001463.

SanMartin:2001:ESS

[ST01a] Jaime San Martín and Soledad Torres. Euler scheme for solutions of a countable system of stochastic differential equations. Statistics & Probability Letters, 54(3):251–259, October 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000505.

Spizzichino:2001:MNA

[ST01b] F. Spizzichino and G. L. Torrisi. Multivariate negative aging in an exchangeable model of heterogeneity. Statistics & Probability Letters, 55(1):71–82, November 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001328.

Shalabh:2003:PML

[ST03a] H. Shalabh and H. Toutenburg. Pseudo-minimax linear and mixed regression estimation of regression coefficients when prior estimates are available. Statistics & Probability Letters, 63(1):35–39, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202004340.

Shimizu:2003:ENL

[ST03b] Kunio Shimizu and Minoru Tanaka. Expected number of level-crossings for a strictly stationary ellipsoidal process. Statistics & Probability Letters, 64(3):305–310, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203001755.

Shishebor:2004:GNB

[ST04a] Z. Shishebor and M. Towhidi. On the generalization of negative binomial distribution. Statistics & Probability Letters, 66(2):127–133, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002979.

Stoyanov:2004:NSC

[ST04b] Jordan Stoyanov and Leonid Tolmatz. New Stieltjes classes involving generalized gamma distributions. Statistics & Probability Letters, 69(2):213–219, August 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001889.

Saigo:2006:OSS

[ST06] Tatsuhiko Saigo and Yozo Tamura. Operator semi-self-similar processes and their space-scaling matrices. Statistics & Probability Letters, 76(7):675–681, April 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003688.

Sun:2009:ALD

[ST09] Liuquan Sun and Xingwei Tong. Analyzing longitudinal data with informative observation times under biased sampling. Statistics & Probability Letters, 79(9):1162–1168, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209000091.

Stadje:2000:IAP

[Sta00] Wolfgang Stadje. An iterative approximation procedure for the distribution of the maximum of a random walk. Statistics & Probability Letters, 50(4):375–381, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001243.

Stadtmuller:2002:ASV

[Sta02] U. Stadtmüller. Almost sure versions of distributional limit theorems for certain order statistics. Statistics & Probability Letters, 58(4):413–426, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001566.

Stadje:2005:EAM

[Sta05] Wolfgang Stadje. The evolution of aggregated Markov chains. Statistics & Probability Letters, 74(4):303–311, October 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205001884.

Stenflo:2001:NTK

[Ste01] Örjan Stenflo. A note on a theorem of Karlin. Statistics & Probability Letters, 54(2):183–187, September 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000451.

Steland:2005:DCM

[Ste05] Ansgar Steland. On the distribution of the clipping median under a mixture model. Statistics & Probability Letters, 71(1):1–13, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002767.

Stehlik:2006:ELR

[Ste06] Milan Stehlík. Exact likelihood ratio scale and homogeneity testing of some loss processes. Statistics & Probability Letters, 76(1):19–26, January 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520500249X.

Stockis:2008:NIC

[STKF08] Jean-Pierre Stockis, Joseph Tadjuidje-Kamgaing, and Jürgen Franke. A note on the identifiability of the conditional expectation for the mixtures of neural networks. Statistics & Probability Letters, 78(6):739–742, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003264.

Shin:2001:UCO

[STM01] Hwashin H. Shin, Glen K. Takahara, and Duncan J. Murdoch. Uniqueness, consistency and optimality in spherical regression experiments. Statistics & Probability Letters, 54(1):61–65, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000530.

Stoica:2003:FLL

[Sto03] George Stoica. Functional local law of the iterated logarithm for geometrically weighted random series. Statistics & Probability Letters, 62(1):71–77, March 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202004261.

Stoica:2008:BKT

[Sto08] George Stoica. The Baum–Katz theorem for bounded subsequences. Statistics & Probability Letters, 78(7):924–926, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003422.

Stoica:2009:NBK

[Sto09] George Stoica. Noncommutative Baum–Katz theorems. Statistics & Probability Letters, 79(3):320–323, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208004045.

Stummer:2000:NAP

[Stu00] Wolfgang Stummer. On Novikov and arbitrage properties of multidimensional diffusion processes with exploding drift. Statistics & Probability Letters, 46(1):43–51, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000851.

Su:2005:LIL

[Su05] Zhonggen Su. The law of the iterated logarithm for character ratios. Statistics & Probability Letters, 71(4):337–346, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000052.

Subramanian:2009:MIB

[Sub09] Sundarraman Subramanian. The multiple imputations based Kaplan–Meier estimator. Statistics & Probability Letters, 79(18):1906–1914, September 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209001990.

Suen:2003:CMO

[Sue03] Chung-Yi Suen. Construction of mixed orthogonal arrays by juxtaposition. Statistics & Probability Letters, 65(3): 161–163, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520300213X.

Sun:2000:NPC

[Sun00] Jianguo Sun. A note on principal component analysis for multi-dimensional data. Statistics & Probability Letters, 46(1):69–73, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299000887.

Sung:2001:SLW

[Sun01] Soo Hak Sung. Strong laws for weighted sums of i.i.d. random variables. Statistics & Probability Letters, 52(4):413–419, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000207.

Sun:2005:EDP

[Sun05] Li-Juan Sun. The expected discounted penalty at ruin in the Erlang (2) risk process. Statistics & Probability Letters, 72(3):205–217, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000301.

Sun:2006:SLU

[Sun06a] Liuquan Sun. The strong law under a semiparametric model for truncated and censored data. Statistics & Probability Letters, 76(14):1550–1558, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000988.

Sun:2006:BRS

[Sun06b] Shuxia Sun. The Bahadur representation for sample quantiles under weak dependence. Statistics & Probability Letters, 76(12):1238–1244, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004797.

Sun:2007:MRB

[Sun07a] Ping Sun. Moment representation of Bernoulli polynomial, Euler polynomial and Gegenbauer polynomials. Statistics & Probability Letters, 77(7):748–751, April 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206003464.

Sung:2007:CCW

[Sun07b] Soo Hak Sung. Complete convergence for weighted sums of random variables. Statistics & Probability Letters, 77 (3):303–311, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002379.

Sung:2007:NEI

[Sun07c] Soo Hak Sung. A note on the exponential inequality for associated random variables. Statistics & Probability Letters, 77(18):1730–1736, December 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001587.

Sung:2008:NCC

[Sun08a] Soo Hak Sung. A note on the complete convergence for arrays of rowwise independent random elements. Statistics & Probability Letters, 78(11):1283–1289, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004075.

Sung:2008:NHR

[Sun08b] Soo Hak Sung. A note on the Hajek–Rényi inequality for associated random variables. Statistics & Probability Letters, 78(7):885–889, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207002957.

Sung:2009:LSL

[Sun09a] Soo Hak Sung. A law of the single logarithm for weighted sums of i.i.d. random elements. Statistics & Probability Letters, 79(10):1351–1357, May 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000716.

Sung:2009:NCC

[Sun09b] Soo Hak Sung. A note on the complete convergence of moving average processes. Statistics & Probability Letters, 79 (11):1387–1390, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000959.

Sriram:2000:MHD

[SV00] T. N. Sriram and A. N. Vidyashankar. Minimum Hellinger distance estimation for supercritical Galton–Watson processes. Statistics & Probability Letters, 50(4):331–342, December 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001127.

Shevlyakov:2002:MVE

[SV02] Georgy L. Shevlyakov and Nikita O. Vilchevski. Minimax variance estimation of a correlation coefficient for ϵcontaminated bivariate normal distributions. Statistics & Probability Letters, 57(1):91–100, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000585.

Song:2006:MFR

[SV06] Renming Song and Zoran Vondracek. On the monotonicity of a function related to the local time of a symmetric Lévy process. Statistics & Probability Letters, 76(14):1522–1528, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206000940.

Samaniego:2007:CEP

[SVB07] F. J. Samaniego, E. M. Vestrup, and D. Bhattacharya. On constrained estimation problems in time-use surveys. Statistics & Probability Letters, 77(2):204–210, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002239.

Sung:2005:MCC

[SVH05] Soo Hak Sung, Andrei I. Volodin, and Tien-Chung Hu. More on complete convergence for arrays. Statistics & Probability Letters, 71(4):303–311, March 15, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400313X.

Steerneman:2008:RND

[SvPtK08] A. G. M. Steerneman and Frederieke van Perlo-ten Kleij. The rank of a normally distributed matrix and positive definiteness of a noncentral Wishart distributed matrix. Statistics & Probability Letters, 78(3):249–253, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002209.

Serfling:2000:LDT

[SW00a] Robert Serfling and Wenyang Wang. A large deviation theorem for U-processes. Statistics & Probability Letters, 49(2):181–193, August 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520000047X.

Su:2000:NAD

[SW00b] Xiaoping Su and Sylvan Wallenstein. New approximations for the distribution of the r-scan statistic. Statistics & Probability Letters, 46(4):411–419, February 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001352.

Sitter:2001:NWC

[SW01] Randy R. Sitter and Changbao Wu. A note on Woodruff confidence intervals for quantiles. Statistics & Probability Letters, 52(4):353–358, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002078.

Sun:2004:IAL

[SW04] Rui-Bo Sun and Bo-Cheng Wei. On influence assessment for LAD regression. Statistics & Probability Letters, 67(2):97–110, April 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002955.

Schick:2006:PCR

[SW06] Anton Schick and Wolfgang Wefelmeyer. Pointwise convergence rates and Central Limit Theorems for kernel density estimators in linear processes. Statistics & Probability Letters, 76(16):1756–1760, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001386.

Schick:2007:PIL

[SW07a] Anton Schick and Wolfgang Wefelmeyer. Prediction in invertible linear processes. Statistics & Probability Letters, 77(12):1322–1331, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000958.

Sun:2007:IEN

[SW07b] Y. Sun and A. C. M. Wong. Interval estimation for the normal correlation coefficient. Statistics & Probability Letters, 77(17):1652–1661, November 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001472.

Shi:2009:NCR

[SWM09] Xiaoping Shi, Yuehua Wu, and Baiqi Miao. A note on the convergence rate of the kernel density estimator of the mode.

Statistics & Probability Letters, 79(17):1866–1871, September 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520900193X.

Su:2008:TPS

[SX08] Liangjun Su and Zhijie Xiao. Testing for parameter stability in quantile regression models. Statistics & Probability Letters, 78(16):2768–2775, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001995.

Stoica:2005:NTP

[SXL05] Petre Stoica, Luzhou Xu, and Jian Li. A new type of parameter estimation algorithm for missing data problems. Statistics & Probability Letters, 75(3):219–229, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002270.

Stoica:2007:OCI

[SXLX07] Petre Stoica, Luzhou Xu, Jian Li, and Yao Xie. Optimal correction of an indefinite estimated MA spectral density matrix. Statistics & Probability Letters, 77(10):973–980, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520700034X.

Shieh:2004:DSB

[SY04] Narn-Rueih Shieh and Jinghu Yu. Dimensions of supercritical branching processes in varying environments. Statistics & Probability Letters, 70(4):299–308, December 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400255X.

Song:2007:FTD

[SY07] Ruili Song and Jiangang Ying. A formula for transition density function under Girsanov transform. Statistics & Probability Letters, 77(6):658–666, March 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600335X.

Sun:2000:SMT

[SZ00] Liuquan Sun and Lixing Zhu. A semiparametric model for truncated and censored data. Statistics & Probability Letters, 48(3):217–227, July 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299002072.

Sun:2001:SFD

[SZ01] Liuquan Sun and Xian Zhou. Survival function and density estimation for truncated dependent data. Statistics & Probability Letters, 52(1):47–57, March 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002145.

Stern:2002:TIP

[SZ02a] J. M. Stern and S. Zacks. Testing the independence of Poisson variates under the Holgate bivariate distribution: the power of a new evidence test. Statistics & Probability Letters, 60(3):313–320, December 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003140.

Sun:2002:RAS

[SZ02b] Liuquan Sun and Xian Zhou. Regression analysis for a semiparametric model with panel data. Statistics & Probability Letters, 58(3):309–317, July 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001475.

Slominski:2005:INI

[SZ05] Leszek Slomiński and Bartosz Ziemkiewicz. Inequalities for the L^{p} norms of integrals with respect to a fractional Brownian motion. Statistics & Probability Letters, 73(1):79–90, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000842.

Su:2006:CBB

[SZ06] Yuhua Su and Mai Zhou. On a connection between the Bradley–Terry model and the Cox proportional hazards model. Statistics & Probability Letters, 76(7):698–702, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003779.

Sun:2008:IAR

[SZ08a] Liuquan Sun and Xian Zhou. Inference in the additive risk model with time-varying covariates subject to measurement errors. Statistics & Probability Letters, 78(16):2559–2566, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003076.

Sun:2008:IME

[SZ08b] Xiaoqian Sun and Xian Zhou. Improved minimax estimation of the bivariate normal precision matrix under the squared loss. Statistics & Probability Letters, 78(2):127–134, February 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002039.

Schmidt:2009:CAD

[SZ09a] Karl Michael Schmidt and Anatoly Zhigljavsky. A characterization of the arcsine distribution. Statistics & Probability Letters, 79(24):2451–2455, December 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003290.

Slominski:2009:WAI

[SZ09b] Leszek Slomiński and Bartosz Ziemkiewicz. On weak approximations of integrals with respect to fractional Brownian motion. Statistics & Probability Letters, 79(4):543–552, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004641.

Szablowski:2001:DND

[Sza01] Paweł J. Szablowski. Discrete normal distribution and its relationship with Jacobi Theta functions. Statistics & Probability Letters, 52(3):289–299, April 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002236.

Szajowski:2007:GVC

[Sza07] Krzysztof Szajowski. A game version of the Cowan–Zabczyk– Bruss’ problem. Statistics & Probability Letters, 77(17): 1683–1689, November 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001526.

Szewczak:2008:EEO

[Sze08a] Zbigniew S. Szewczak. Edgeworth expansions in operator form. Statistics & Probability Letters, 78(12):1583– 1592, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000151.

Szewczak:2008:MRT

[Sze08b] Zbigniew S. Szewczak. On moments of recurrence times for positive recurrent renewal sequences. Statistics & Probability Letters, 78(17):3086–3090, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002721.

Szkutnik:2002:NQM

[Szk02] Zbigniew Szkutnik. A note on quasi-maximum likelihood solutions to an inverse problem for Poisson processes. Statistics & Probability Letters, 60(3):253–263, December 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002699.

Szulga:2002:HDW

[Szu02] Jerzy Szulga. Hausdorff dimension of Weierstrass–Mandelbrot process. Statistics & Probability Letters, 56(3):301–307, February 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520100195X.

Szulga:2009:SST

[Szu09] Jerzy Szulga. On selfdecomposable Stieltjes transforms. Statistics & Probability Letters, 79(6):748–752, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005099.

Takahashi:2004:RTM

[Tak04] Hiroshi Takahashi. Recurrence and transience of multidimensional diffusion processes in reflected Brownian environments. Statistics & Probability Letters, 69(2):171–174, August 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001774.

Talarczyk:2008:FET

[Tal08] Anna Talarczyk. A functional ergodic theorem for the occupation time process of a branching system. Statistics & Probability Letters, 78(7):847–853, May 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002908.

Tang:2007:ORW

[Tan07] Qihe Tang. The overshoot of a random walk with negative drift. Statistics & Probability Letters, 77(2):158–165, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002161.

Tango:2009:SSF

[Tan09] Toshiro Tango. Sample size formula for randomized controlled trials with counts of recurrent events. Statistics & Probability Letters, 79(4):466–472, February 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004525.

Thavaneswaran:2005:FV

[TAP05] A. Thavaneswaran, S. S. Appadoo, and S. Peiris. Forecasting volatility. Statistics & Probability Letters, 75(1):1–10, November 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520500204X.

Tardella:2001:NED

[Tar01]  Luca Tardella.  A note on estimating the diameter

of a truncated moment class. Statistics & Probability Letters, 54(2):115–124, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200002054.

Tarasinska:2005:CIP

[Tar05] Joanna Tarasińska. Confidence intervals for the power of Student’s t-test. Statistics & Probability Letters, 73(2):125–130, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520500101X.

Tokarev:2009:EMS

[TB09] Daniel V. Tokarev and Konstantin A. Borovkov. On the expectations of maxima of sets of independent random variables. Statistics & Probability Letters, 79(23):2381–2388, December 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003095.

Teke:2008:IRT

[TD08] S. P. Teke and S. R. Deshmukh. Inverse renewal thinning of Cox and renewal processes. Statistics & Probability Letters, 78(16):2705–2708, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001880.

Tenreiro:2001:ABI

[Ten01] Carlos Tenreiro. On the asymptotic behaviour of the integrated square error of kernel density estimators with datadependent bandwidth. Statistics & Probability Letters, 53 (3):283–292, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000724.

Tenreiro:2003:ANM

[Ten03] Carlos Tenreiro. On the asymptotic normality of multistage integrated density derivatives kernel estimators. Statistics & Probability Letters, 64(3):311–322, September 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001767.

Teran:2003:SLL

[Ter03] Pedro Terán. A Strong Law of Large Numbers for random upper semicontinuous functions under exchangeability conditions. Statistics & Probability Letters, 65(3):251–258, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715203002682.

Teran:2008:ULL

[Ter08] Pedro Terán. On a uniform Law of Large Numbers for random sets and subdifferentials of random functions. Statistics & Probability Letters, 78(1):42–49, January 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001939.

Tille:2005:OAB

[TF05] Yves Tillé and Anne-Catherine Favre. Optimal allocation in balanced sampling. Statistics & Probability Letters, 74 (1):31–37, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205001549.

Tableman:2002:SWE

[TG02] Mara Tableman and Alix I. Gitelman. The scaled α-Winsorized estimate of exponential scale for censored data: an analysis based on two influence functions. Statistics & Probability Letters, 59(2):169–181, September 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001451.

Theodosopoulos:2007:RCB

[The07] Ted Theodosopoulos. A reversion of the Chernoff bound. Statistics & Probability Letters, 77(5):558–565, March 1, 2007.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002884.

Tian:2009:CML

[Tia09] Yubin Tian. Consistency of the maximum likelihood estimator and Bayesian estimator based on sequential sensitivity experiments. Statistics & Probability Letters, 79(6):728–732, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208005063.

Thomas:2006:GDM

[TJ06] Seemon Thomas and Joy Jacob. A generalized Dirichlet model. Statistics & Probability Letters, 76(16):1761–1767, October 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001398.

Tomy:2009:GNL

[TJ09] Lishamol Tomy and K. K. Jose. Generalized normal-Laplace AR process. Statistics & Probability Letters, 79(14):1615– 1620, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001400.

Taskinen:2003:STI

[TKO03] Sara Taskinen, Annaliisa Kankainen, and Hannu Oja. Sign test of independence between two random vectors. Statistics & Probability Letters, 62(1):9–21, March 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202003991.

Tian:2004:QRM

[TL04] Maozai Tian and Guoying Li. A quasi-residuals method in sliced inverse regression. Statistics & Probability Letters, 66 (2):205–212, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203003419.

Tchetgen:2008:MEI

[TLRvdV08] Eric Tchetgen, Lingling Li, James Robins, and Aad van der Vaart. Minimax estimation of the integral of a power of a density. Statistics & Probability Letters, 78(18):3307–3311, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003209.

Trippa:2009:BNB

[TM09] Lorenzo Trippa and Pietro Muliere. Bayesian nonparametric binary regression via random tessellations. Statistics & Probability Letters, 79(21):2273–2280, November 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002892.

Taylor:2002:SET

[TMH02] Jeremy M. G. Taylor, Susan Murray, and Chiu-Hsieh Hsu. Survival estimation and testing via multiple imputation. Statistics & Probability Letters, 58(3):221–232, July 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000305.

Terpstra:2001:GEA

[TMN01] Jeffrey T. Terpstra, Joseph W. McKean, and Joshua D. Naranjo. GR-estimates for an autoregressive time series. Statistics & Probability Letters, 51(2):165–172, January 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200001498.

Triantafyllopoulos:2009:NSS

[TN09] K. Triantafyllopoulos and G. P. Nason. A note on state space representations of locally stationary wavelet time series. Statistics & Probability Letters, 79(1):50–54, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003428.

Thas:2003:SGA

[TO03] O. Thas and J. P. Ottoy. Some generalizations of the Anderson–Darling statistic. Statistics & Probability Letters, 64(3):255–261, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520300169X.

Tomacs:2005:CRL

[Tóm05] Tibor Tómács. Convergence rates in the Law of Large Numbers for arrays of Banach space valued random elements. Statistics & Probability Letters, 72(1):59–69, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000167.

Toma:2009:ORE

[Tom09] Aida Toma. Optimal robust M-estimators using divergences. Statistics & Probability Letters, 79(1):1–5, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002319.

Torabi:2006:NMH

[Tor06] Hamzeh Torabi. A new method for hypotheses testing using spacings. Statistics & Probability Letters, 76(13):1345–1347, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520600040X.

Toth:2004:AIP

[Tot04] Daniell Toth. Adding interior points to an existing Brownian sheet lattice. Statistics & Probability Letters, 66(3): 221–227, February 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203003006.

Tzavelas:2001:TQS

[TP01] George Tzavelas and Athanasios G. Paliatsos. Truncated quasi-score function in the 1-dependent and stationary case. Statistics & Probability Letters, 53(1):1–9, May 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002285.

Thavaneswaran:2003:GSE

[TP03] A. Thavaneswaran and Shelton Peiris. Generalized smoothed estimating functions for nonlinear time series. Statistics & Probability Letters, 65(1):51–56, October 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203002189.

Tan:2009:ASC

[TP09] Zhongquan Tan and Zuoxiang Peng. Almost sure convergence for non-stationary random sequences. Statistics & Probability Letters, 79(7):857–863, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005269.

Thavaneswaran:2008:RCV

[TPA08] A. Thavaneswaran, S. Peiris, and S. Appadoo. Random coefficient volatility models. Statistics & Probability Letters, 78(6):582–593, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207002994.

Terpstra:2002:ADM

[TR02] Jeffrey T. Terpstra and M. Bhaskara Rao. On the asymptotic distribution of a multivariate GR-estimate for a VAR(p) time series. Statistics & Probability Letters, 60(2):219–230, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202003218.

Tribouley:2004:ASC

[Tri04] Karine Tribouley. Adaptive simultaneous confidence intervals in non-parametric estimation. Statistics & Probability Letters, 69(1):37–51, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001270.

Triacca:2008:SCS

[Tri08a] Umberto Triacca. Is a subspace containing a splitting subspace a splitting subspace? Statistics & Probability Letters, 78(17):2997–2999, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002575.

Triantafyllopoulos:2008:MOA

[Tri08b] K. Triantafyllopoulos. Missing observation analysis for matrix-variate time series data. Statistics & Probability Letters, 78(16):2647–2653, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800179X.

Tanaka:2001:DCE

[TS01] Minoru Tanaka and Kunio Shimizu. Discrete and continuous expectation formulae for level-crossings, upcrossings and excursions of ellipsoidal processes. Statistics & Probability Letters, 52(3):225–232, April 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001929.

Tang:2009:VIU

[TS09a] Hsiu-Khuern Tang and Chuen-Teck See. Variance inequalities using first derivatives. Statistics & Probability Letters, 79(9):1277–1281, May 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000443.

Tiku:2009:MGE

[TS09b] Moti L. Tiku and Baris Sürücü. MMLEs are as good as M-estimators or better. Statistics & Probability Letters, 79(7):984–989, April 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005646.

Tsao:2001:SSC

[Tsa01] Min Tsao. A small sample calibration method for the empirical likelihood ratio. Statistics & Probability Letters, 54 (1):41–45, August 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201000372.

Tsao:2004:NMC

[Tsa04] Min Tsao. A new method of calibration for the empirical loglikelihood ratio. Statistics & Probability Letters, 68(3):305– 314, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016771520400121X.

Tao:2002:SPI

[TSGG02] Jian Tao, Ning-Zhong Shi, Jianhua Guo, and Wei Gao. Stepwise procedures for the identification of minimum effective dose with unknown variances. Statistics & Probability Letters, 57(2):121–131, April 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201001973.

Taheriyoun:2009:NHM

[TSJ09] Ali Reza Taheriyoun, Khalil Shafie, and Mohammad Jafari Jozani. A note on the higher moments of the Euler characteristic of the excursion sets of random fields. Statistics & Probability Letters, 79(8):1074–1082, April 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005774.

Tang:2001:LDH

[TSJZ01] Qihe Tang, Chun Su, Tao Jiang, and Jinsong Zhang. Large deviations for heavy-tailed random sums in compound renewal model. Statistics & Probability Letters, 52(1):91–100, March 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002315.

Tsou:2000:ECB

[Tso00] Tsung-Shan Tsou. Exchangeable cluster binary data correlation coefficient estimation with generalized estimating equations. Statistics & Probability Letters, 50(2):179–186, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000961.

Tian:2003:ESS

[TT03] Guo-Liang Tian and Ming Tan. Exact statistical solutions using the inverse Bayes formulae. Statistics & Probability Letters, 62(3):305–315, April 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000440.

Tang:2007:DPG

[TT07a] Shijie Tang and Kam-Wah Tsui. Distributional properties for the generalized p-value for the Behrens–Fisher problem. Statistics & Probability Letters, 77(1):1–8, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001854.

Tsui:2007:STM

[TT07b] Kam-Wah Tsui and Shijie Tang. Simultaneous testing of multiple hypotheses using generalized p-values. Statistics & Probability Letters, 77(12):1362–1370, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001009.

Takei:2008:PDO

[TT08] Masato Takei and Masaki Takeshima. Phase diagram for oncereinforced random walks on trees with exponential weighting scheme. Statistics & Probability Letters, 78(17):3000–3007, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002587.

Tudor:2008:IPS

[Tud08] Constantin Tudor. Inner product spaces of integrands associated to subfractional Brownian motion. Statistics & Probability Letters, 78(14):2201–2209, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001119.

Tang:2007:IRP

[TV07] Qihe Tang and Raluca Vernic. The impact on ruin probabilities of the association structure among financial risks.

Statistics & Probability Letters, 77(14):1522–1525, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207001253.

Tian:2006:EPO

[TW06] Yongge Tian and Douglas P. Wiens. On equality and proportionality of ordinary least squares, weighted least squares and best linear unbiased estimators in the general linear model. Statistics & Probability Letters, 76(12):1265–1272, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520600023X.

Tang:2000:IDN

[TWW00] Nian-Sheng Tang, Bo-Cheng Wei, and Xue-Ren Wang. Influence diagnostics in nonlinear reproductive dispersion models. Statistics & Probability Letters, 46(1):59–68, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000875.

Ursu:2009:MSM

[UD09] Eugen Ursu and Pierre Duchesne. On multiplicative seasonal modelling for vector time series. Statistics & Probability Letters, 79(19):2045–2052, October 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002387.

Uddin:2002:MCC

[Udd02] Nizam Uddin. A method of construction of a class of universally optimal structurally incomplete row-column designs. Statistics & Probability Letters, 59(1):93–98, August 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002109.

Uddin:2008:MOB

[Udd08] Nizam Uddin. MV-optimal block designs for correlated errors. Statistics & Probability Letters, 78(17):2926–2931, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208002381.

Uno:2000:SEL

[UI00] Chikara Uno and Eiichi Isogai. Sequential estimation of a linear combination of means. Statistics & Probability Letters, 47(1):45–52, March 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001364.

Umbach:2009:BAC

[UJ09] Dale Umbach and S. Rao Jammalamadaka. Building asymmetry into circular distributions. Statistics & Probability Letters, 79(5):659–663, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800494X.

Umbach:2006:SMR

[Umb06] Dale Umbach. Some moment relationships for skew-symmetric distributions. Statistics & Probability Letters, 76(5):507–512, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003329.

Umbach:2008:SMR

[Umb08] Dale Umbach. Some moment relationships for multivariate skew-symmetric distributions. Statistics & Probability Letters, 78(12):1619–1623, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000205.

Ulyanov:2006:BEB

[UWF06] Vladimir V. Ulyanov, Hirofumi Wakaki, and Yasunori Fujikoshi. Berry–Esseen bound for high dimensional asymptotic approximation of Wilks’ Lambda distribution. Statistics & Probability Letters, 76(12):1191–1200, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004736.

Vuolle-Apiala:2007:SWT

[VA07]  Juha Vuolle-Apiala.  Shiga–Watanabe’s time inversion

property for self-similar diffusion processes. Statistics & Probability Letters, 77(9):920–924, May 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000314.

Vaggelatou:2003:LLR

[Vag03] Eutichia Vaggelatou. On the length of the longest run in a multi-state Markov chain. Statistics & Probability Letters, 62(3):211–221, April 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202004327.

VanderWeele:2008:SRB

[Van08] Tyler J. VanderWeele. Simple relations between principal stratification and direct and indirect effects. Statistics & Probability Letters, 78(17):2957–2962, December 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002496.

Varron:2008:SAR

[Var08] Davit Varron. Some asymptotic results on density estimators by wavelet projections. Statistics & Probability Letters, 78(15):2517–2521, October 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001582.

Vignat:2008:EWT

[VB08] C. Vignat and S. Bhatnagar. An extension of Wick’s theorem. Statistics & Probability Letters, 78(15):2404–2407, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001405.

vanDoorn:2005:BDP

[vDZ05] Erik A. van Doorn and Alexander I. Zeifman. Birth-death processes with killing. Statistics & Probability Letters, 72(1):33– 42, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000131.

vanEeden:2000:MEL

[vE00] Constance van Eeden. Minimax estimation of a lowerbounded scale parameter of an F distribution. Statistics & Probability Letters, 46(3):283–286, February 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001145.

vanEs:2001:EMI

[vE01] Bert van Es. On the expansion of the mean integrated squared error of a kernel density estimator. Statistics & Probability Letters, 52(4):441–450, May 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000323.

vanEs:2008:WCS

[vEG08] Bert van Es and Shota Gugushvili. Weak convergence of the supremum distance for supersmooth kernel deconvolution. Statistics & Probability Letters, 78(17):2932–2938, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002393.

Verrill:2007:RCS

[Ver07] Steve Verrill. Rate of convergence of k-step Newton estimators to efficient likelihood estimators. Statistics & Probability Letters, 77(12):1371–1376, July 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001010.

vanEeden:2001:EOT

[vEZ01] Constance van Eeden and James V. Zidek. Estimating one of two normal means when their difference is bounded. Statistics & Probability Letters, 51(3):277–284, February 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001590. See correction [vEZ02].

vanEeden:2002:CEO

[vEZ02] Constance van Eeden and James V. Zidek. Correction on “Estimating one of two normal means when their difference is bounded”: [Statist. Probab. Lett. 51 (2001) 277– 284]. Statistics & Probability Letters, 57(1):111, March 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201002024. See [vEZ01].

Vinogradov:2008:ATC

[Vin08] Vladimir Vinogradov. On approximations for two classes of Poisson mixtures. Statistics & Probability Letters, 78(4):358– 366, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002489.

Vitale:2008:GRI

[Vit08] R. A. Vitale. On the Gaussian representation of intrinsic volumes. Statistics & Probability Letters, 78(10):1246–1249, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207004002.

Vellaisamy:2008:EPP

[VJ08] P. Vellaisamy and Sushmita Jain. Estimating the parameter of the population selected from discrete exponential family. Statistics & Probability Letters, 78(9):1076–1087, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003690.

Vandenhende:2003:IRB

[VL03] François Vandenhende and Philippe Lambert. Improved rankbased dependence measures for categorical data. Statistics & Probability Letters, 63(2):157–163, June 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000634.

Vock:2006:GCM

[Voc06] Michael Vock. Graphical comparison of multivariate nonparametric location tests for restricted alternatives. Statistics & Probability Letters, 76(14):1529–1535, August 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000952.

Volkov:2006:NSR

[Vol06] Stanislav Volkov. A note on the simple random walk on Z^{2}: Probability of exiting sequences of sets. Statistics & Probability Letters, 76(9):891–897, May 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004050.

Vougas:2006:RAD

[Vou06] Dimitrios V. Vougas. Remark on the asymptotic distribution of the OLS estimator in a simple Gaussian unit-root autoregression. Statistics & Probability Letters, 76(1):27–34, January 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002634.

Vougas:2008:GLS

[Vou08a] Dimitrios V. Vougas. Generalized least squares transformation and estimation with autoregressive error. Statistics & Probability Letters, 78(4):402–404, March 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002544.

Vougas:2008:URT

[Vou08b] Dimitrios V. Vougas. Unit root testing based on BLUS residuals. Statistics & Probability Letters, 78(13):1943– 1947, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520800062X.

Vasdekis:2000:NEI

[VT00] V. G. S. Vasdekis and A. Trichopoulou. Nonparametric estimation of individual food availability along with bootstrap confidence intervals in household budget surveys. Statistics & Probability Letters, 46(4):337–345, February 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001200.

Veena:2008:CBD

[VT08] T. G. Veena and P. Yageen Thomas. Characterizations of bivariate distributions by properties of concomitants of order statistics. Statistics & Probability Letters, 78(18):3350–3354, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800326X.

Vellaisamy:2007:NBD

[VU07] P. Vellaisamy and N. S. Upadhye. On the negative binomial distribution and its generalizations. Statistics & Probability Letters, 77(2):173–180, January 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206002197.

Vilar:2000:FSP

[VV00] José A. Vilar and Juan M. Vilar. Finite sample performance of density estimators from unequally spaced data. Statistics & Probability Letters, 50(1):63–73, October 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000833.

Vermaat:2008:ABV

[VvdMD08] M. B. Vermaat, F. H. van der Meulen, and R. J. M. M. Does. Asymptotic behavior of the variance of the EWMA statistic for autoregressive processes. Statistics & Probability Letters, 78(12):1673–1682, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000321.

vanWieringen:2005:ICL

[vW05] Wessel N. van Wieringen. On identifiability of certain latent class models. Statistics & Probability Letters, 75(3):211–218, December 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205002282.

Vitale:2008:WFP

[VW08] Richard A. Vitale and Yazhen Wang. The Wills functional for Poisson processes. Statistics & Probability Letters, 78(14):2181–2187, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208001089.

vanZanten:2000:MCL

[vZ00] Harry van Zanten. A multivariate Central Limit Theorem for continuous local martingales. Statistics & Probability Letters, 50(3):229–235, November 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001085.

Valdes:2003:OAA

[VZ03] José E. Valdés and Rómulo I. Zequeira. On the optimal allocation of an active redundancy in a two-component series system. Statistics & Probability Letters, 63(3):325–332, July 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203000981.

Wimmer:2001:NTP

[WA01] Gejza Wimmer and Gabriel Altmann. A new type of partialsums distributions. Statistics & Probability Letters, 52(4):359– 364, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002121.

Waagepetersen:2004:CPD

[Waa04] Rasmus Waagepetersen. Convergence of posteriors for discretized log Gaussian Cox processes. Statistics & Probability Letters, 66(3):229–235, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003079.

Walk:2002:CVK

[Wal02] Harro Walk. On cross-validation in kernel and partitioning regression estimation. Statistics & Probability Letters, 59(2):113–123, September 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715202001037.

Wang:2000:MPI

[Wan00a] Qi-Hua Wang. Moment and probability inequalities for the bivariate product-limit estimator. Statistics & Probability Letters, 46(1):1–12, January 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299000814.

Wang:2000:BPD

[Wan00b] Xikui Wang. A bandit process with delayed responses. Statistics & Probability Letters, 48(3):303–307, July 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000110.

Wang:2001:PIL

[Wan01] Wensheng Wang. On the properties for increments of a local time — a look through the set of limit points. Statistics & Probability Letters, 52(3):329–340, April 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002303.

Wang:2002:ICE

[Wan02a] Hsiuying Wang. Improved confidence estimators for the usual one-sided confidence intervals for the ratio of two normal variances. Statistics & Probability Letters, 59(3):307–315, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001980.

Wang:2002:SAM

[Wan02b] Liqun Wang. A simple adjustment for measurement errors in some limited dependent variable models. Statistics & Probability Letters, 58(4):427–433, July 15, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202001657.

Wang:2002:SNL

[Wan02c] Wensheng Wang. Self-normalized lag increments of partial sums. Statistics & Probability Letters, 58(1):41–51, May 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001013.

Wang:2002:APB

[Wan02d] Xikui Wang. Asymptotic properties of bandit processes with geometric responses. Statistics & Probability Letters, 60(2): 211–217, November 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520200319X.

Wang:2004:MIA

[Wan04] Jianfeng Wang. Maximal inequalities for associated random variables and demimartingales. Statistics & Probability Letters, 66(3):347–354, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003626.

Wang:2005:SDT

[Wan05a] Peijie Wang. Statistical distributions of time series in the frequency domain and the patterns of violation of white noise conditions. Statistics & Probability Letters, 74(1):103–108, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205001677.

Wang:2005:MSE

[Wan05b] R. Wang. A mixture and self-exciting model for software reliability. Statistics & Probability Letters, 72(3):187–194, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000088.

Wang:2006:NUR

[Wan06] Gaowen Wang. A note on unit root tests with heavytailed GARCH errors. Statistics & Probability Letters, 76 (10):1075–1079, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004487.

Wang:2007:MVO

[Wan07a] Hsiuying Wang. Modified p-values for one-sided testing in restricted parameter spaces. Statistics & Probability Letters, 77(6):625–631, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002963.

Wang:2007:SNT

[Wan07b] Lan Wang. A simple nonparametric test for diagnosing nonlinearity in Tobit median regression model. Statistics & Probability Letters, 77(10):1034–1042, June 1, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000442.

Wang:2008:NCC

[Wan08a] Antai Wang. A note on checking the Clayton model assumption based on left-truncated bivariate data. Statistics & Probability Letters, 78(10):1168–1173, August 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003884.

Wang:2008:CSD

[Wan08b] Xuewu Wang. A class of strong deviation theorems for the sequence of nonnegative integer valued random variables. Statistics & Probability Letters, 78(18):3281–3287, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003155.

Wang:2008:EUS

[Wan08c] Zhidong Wang. Existence and uniqueness of solutions to stochastic Volterra equations with singular kernels and non-Lipschitz coefficients. Statistics & Probability Letters, 78(9):1062–1071, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003665.

Wang:2009:MPE

[Wan09a] Lihong Wang. Memory parameter estimation for long range dependent random fields. Statistics & Probability Letters, 79(21):2297–2306, November 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002922.

Wang:2009:MIM

[Wan09b] Wei Wang. Maximal inequalities for g-martingales. Statistics & Probability Letters, 79(9):1169–1174, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209000108.

Watanabe:2001:SDF

[Wat01] Toshiaki Watanabe. On sampling the degree-of-freedom of Student’s-t disturbances. Statistics & Probability Letters, 52(2):177–181, April 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200002212.

Wong:2005:EPA

[WB05] Wing-Keung Wong and Guorui Bian. Estimating parameters in autoregressive models with asymmetric innovations. Statistics & Probability Letters, 71(1):61–70, January 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204002846.

Wang:2004:NEB

[WBG04] Jiuzhou Wang, Joseph Boyer, and Marc G. Genton. A note on an equivalence between chi-square and generalized skew-normal distributions. Statistics & Probability Letters, 66(4):395–398, March 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203003377.

Wei:2003:EBE

[WC03] Laisheng Wei and Jiahua Chen. Empirical Bayes estimation and its superiority for two-way classification model. Statistics & Probability Letters, 63(2):165–175, June 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000798.

Wong:2004:CCM

[WC04] Hubert Wong and Bertrand Clarke. A characterization of consistency of model weights given partial information in normal linear models. Statistics & Probability Letters, 68(1):27–37, June 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204000653.

Wu:2007:VBS

[WCC07] Tiee-Jian Wu, Ching-Fu Chen, and Huang-Yu Chen. A variable bandwidth selector in multivariate kernel density estimation. Statistics & Probability Letters, 77(4):462–467, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002665.

Wang:2007:SRW

[WCS07] Dingcheng Wang, Pingyan Chen, and Chun Su. The supremum of random walk with negatively associated and heavytailed steps. Statistics & Probability Letters, 77(13):1403– 1412, July 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207000685.

Wan:2008:ESN

[WD08] Li Wan and Jinqiao Duan. Exponential stability of nonautonomous stochastic partial differential equations with finite memory. Statistics & Probability Letters, 78(5):490–498, April 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002726.

Weber:2006:WCL

[Web06] Michel Weber. A weighted Central Limit Theorem. Statistics & Probability Letters, 76(14):1482–1487, August 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206000824.

Weiss:2008:CIP

[Wei08]  Christian H. Weiß.  The combined INAR(p) models

for time series of counts. Statistics & Probability Letters, 78(13):1817–1822, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520800045X.

Weiss:2009:JBA

[Wei09] Christian H. Weiß. Jumps in binomial AR(1) processes. Statistics & Probability Letters, 79(19):2012–2019, October 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520900234X.

Wendl:2003:CPB

[Wen03] Michael C. Wendl. Collision probability between sets of random variables. Statistics & Probability Letters, 64(3): 249–254, September 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715203001688.

Wen:2007:NSD

[Wen07] Xuerong Meggie Wen. A note on sufficient dimension reduction. Statistics & Probability Letters, 77(8):817–821, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000144.

Worsley:2000:TC

[WF00] K. J. Worsley and K. J. Friston. A test for a conjunction. Statistics & Probability Letters, 47(2):135–140, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001492.

Wang:2003:ODR

[WG03] Ming-Dauh Wang and Seymour Geisser. Optimal dichotomization for repeated screening tests. Statistics & Probability Letters, 62(1):61–70, March 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520200425X.

Wang:2001:FML

[WH01] C. Y. Wang and Yijian Huang. Functional methods for logistic regression on random-effect-coefficients for longitudinal measurements. Statistics & Probability Letters, 53(4):347–356, July 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201000128.

Wei:2002:SDO

[WH02] Gang Wei and Taizhong Hu. Supermodular dependence ordering on a class of multivariate copulas. Statistics & Probability Letters, 57(4):375–385, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000949.

Wang:2005:FIC

[WH05] Yanhua Wang and Shuyuan He. Fisher information in censored data. Statistics & Probability Letters, 73(2):199–206, June 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000404.

Wu:2007:AAE

[WH07] Shujin Wu and Dong Han. Algorithmic analysis of Euler scheme for a class of stochastic differential equations with jumps. Statistics & Probability Letters, 77(2):211–219, January 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206002264.

Wei:2008:RPD

[WH08] Xiao Wei and Yijun Hu. Ruin probabilities for discrete time risk models with stochastic rates of interest. Statistics & Probability Letters, 78(6):707–715, April 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003215.

Wang:2009:BSD

[WH09] Ying Wang and Zhen Huang. Backward stochastic differential equations with non-Lipschitz coefficients. Statistics & Probability Letters, 79(12):1438–1443, June 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001059.

Wang:2008:SLL

[WHSL08] Xuejun Wang, Shuhe Hu, Yan Shen, and Nengxiang Ling. Strong law of large numbers and growth rate for a class of random variable sequences. Statistics & Probability Letters, 78(18):3330–3337, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208003234.

Wong:2001:DJW

[WIL01] Heung Wong, Waicheung Ip, and Yuan Li. Detection of jumps by wavelets in a heteroscedastic autoregressive model. Statistics & Probability Letters, 52(4):365–372, May 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200002182.

Willink:2005:NMH

[Wil05] R. Willink. Normal moments and Hermite polynomials. Statistics & Probability Letters, 73(3):271–275, July 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205001124.

Withers:2000:SEM

[Wit00] C. S. Withers. A simple expression for the multivariate Hermite polynomials. Statistics & Probability Letters, 47(2):165– 169, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715299001534.

Witkovsky:2002:EDP

[Wit02] Viktor Witkovský. Exact distribution of positive linear combinations of inverted chi-square random variables with odd degrees of freedom. Statistics & Probability Letters, 56(1):45–50, January 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001651.

Wittich:2005:ELU

[Wit05] O. Wittich. An explicit local uniform large deviation bound for Brownian bridges. Statistics & Probability Letters, 73(1):51– 56, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000817.

Watamori:2005:ILR

[WJ05] Yoko Watamori and Peter E. Jupp. Improved likelihood ratio and score tests on concentration parameters of von Mises–Fisher distributions. Statistics & Probability Letters, 72(2):93–102, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204002810.

Wu:2008:SSL

[WJ08] Qunying Wu and Yuanying Jiang. Some strong limit theorems for ρ̃-mixing sequences of random variables. Statistics & Probability Letters, 78(8):1017–1023, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003616.

Weerahandi:2007:EHE

[WK07] Samaradasa Weerahandi and Martin A. Koschat. Experimentation on heterogeneous experimental units. Statistics & Probability Letters, 77(11):1077–1083, June 15, 2007. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000454.

Wang:2008:NAS

[WL08a] Jiang-Feng Wang and Han-Ying Liang. A note on the almost sure Central Limit Theorem for negatively associated fields. Statistics & Probability Letters, 78(13):1964–1970, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000655.

Wen:2008:APC

[WL08b] Jiwei Wen and Zhengyan Lin. Asymptotic properties for Cauchy’s principal values of Brownian and random walk local time. Statistics & Probability Letters, 78(11):1317–1327, August 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207004130.

Wang:2001:AMA

[WLG01] Qiying Wang, Yan-Xia Lin, and Chandra M. Gulati. Asymptotics for moving average processes with dependent innovations. Statistics & Probability Letters, 54(4):347–356, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200001954.

Walker:2003:BDP

[WM03] Stephen Walker and Pietro Muliere. A bivariate Dirichlet process. Statistics & Probability Letters, 64(1):1–7, August 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520300124X.

Withers:2006:END

[WM06] C. S. Withers and P. N. McGavin. Expressions for the normal distribution and repeated normal integrals. Statistics & Probability Letters, 76(5):479–487, March 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003287.

Wang:2009:COS

[WN09a] Ke Wang and H. N. Nagaraja. Concomitants of order statistics for dependent samples. Statistics & Probability Letters, 79(4):553–558, February 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004665.

Withers:2009:ATIa

[WN09b] Christopher S. Withers and Saralees Nadarajah. Accurate tests and intervals based on linear cusum statistics. Statistics & Probability Letters, 79(5):689–697, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004999.

Withers:2009:ATIb

[WN09c] Christopher S. Withers and Saralees Nadarajah. Accurate tests and intervals based on nonlinear cusum statistics. Statistics & Probability Letters, 79(21):2242–2250, November 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209002855.

Wang:2008:SPK

[WO08] Antai Wang and David Oakes. Some properties of the Kendall distribution in bivariate Archimedean copula models under censoring. Statistics & Probability Letters, 78(16):2578–2583, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003350.

Woerner:2003:LAN

[Woe03] Jeannette H. C. Woerner. Local asymptotic normality for the scale parameter of stable processes. Statistics & Probability Letters, 63(1):61–65, May 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000531.

Wong:2009:ECS

[Won09] Bernard Wong. Explicit construction of stochastic exponentials with arbitrary expectation k ∈ (0, 1). Statistics & Probability Letters, 79(7):880–883, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005294.

Worms:2001:LMD

[Wor01] Julien Worms. Large and moderate deviations upper bounds for the Gaussian autoregressive process. Statistics & Probability Letters, 51(3):235–243, February 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001346.

Wozabal:2009:ULT

[Woz09] Nancy Wozabal. Uniform limit theorems for functions of order statistics. Statistics & Probability Letters, 79(12):1450–1455, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001084.

Wu:2000:TLC

[WP00] Lang Wu and Michael D. Perlman. Testing lattice conditional independence models based on monotone missing data. Statistics & Probability Letters, 50(2):193–201, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715200000985.

Wang:2008:MES

[WP08] Xueqin Wang and Hanxiang Peng. Moment estimation in a semiparametric generalized linear model. Statistics & Probability Letters, 78(12):1624–1633, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000217.

Wright:2000:CDF

[Wri00] Peter A. Wright. The cumulative distribution function of process capability index C_{pm}. Statistics & Probability Letters, 47(3):249–251, April 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715299001625.

Weems:2004:RML

[WS04] K. S. Weems and P. J. Smith. On robustness of maximum likelihood estimates for Poisson-lognormal models. Statistics & Probability Letters, 66(2):189–196, January 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003365.

Wu:2009:AAI

[WSM09] Tiee-Jian Wu, Xiaoping Shi, and Baiqi Miao. Asymptotic approximation of inverse moments of nonnegative random variables. Statistics & Probability Letters, 79(11):1366–1371, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520900073X.

Wang:2004:MSR

[WT04] Dingcheng Wang and Qihe Tang. Maxima of sums and random sums for negatively associated random variables with heavy tails. Statistics & Probability Letters, 68(3):287–295, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001099.

Wu:2004:WEL

[Wu04] Changbao Wu. Weighted empirical likelihood inference. Statistics & Probability Letters, 66(1):67–79, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003092.

Wu:2008:WCS

[Wu08a] Biao Wu. On the weak convergence of subordinated systems. Statistics & Probability Letters, 78(18):3203–3211, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800299X.

Wu:2008:AMN

[Wu08b] Lang Wu. An approximate method for nonlinear mixedeffects models with nonignorably missing covariates. Statistics & Probability Letters, 78(4):384–389, March 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002519.

Wulff:2008:ERA

[Wul08] Shaun S. Wulff. The equality of REML and ANOVA estimators of variance components in unbalanced normal classification models. Statistics & Probability Letters, 78(4):405– 411, March 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207002556.

Wang:2008:MSB

[WV08] Lichun Wang and Noël Veraverbeke. MSE superiority of Bayes and empirical Bayes estimators in two generalized seemingly unrelated regressions. Statistics & Probability Letters, 78(2):109–117, February 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002015.

Wang:2001:NKA

[WW01] Suojin Wang and C. Y. Wang. A note on kernel assisted estimators in missing covariate regression. Statistics & Probability Letters, 55(4):439–449, December 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001675.

Wuyungaowa:2008:AEI

[WW08] Wuyungaowa and Tianming Wang. Asymptotic expansions for inverse moments of binomial and negative binomial. Statistics & Probability Letters, 78(17):3018–3022, December 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208002617.

Wu:2000:LLR

[WX00] Yanhong Wu and Yongxia Xu. Local likelihood ratio tests in the normal mixture model. Statistics & Probability Letters, 50(4):323–329, December 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001103.

Wierman:2008:GSO

[WX08] John C. Wierman and Pengfei Xiang. A general SLLN for the one-dimensional class cover problem. Statistics & Probability Letters, 78(9):1110–1118, July 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003811.

Wu:2009:CTF

[WX09] Zhen Wu and Mingyu Xu. Comparison theorems for forward backward SDEs. Statistics & Probability Letters, 79 (4):426–435, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004392.

Wang:2008:MFP

[WY08a] Huiqing Wang and Chuancun Yin. Moments of the first passage time of one-dimensional diffusion with two-sided barriers. Statistics & Probability Letters, 78(18):3373–3380, December 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003295.

Wang:2008:SDR

[WY08b] Qin Wang and Xiangrong Yin. Sufficient dimension reduction and variable selection for regression mean function with two types of predictors. Statistics & Probability Letters, 78(16):2798–2803, November 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002046.

Wen:2003:PAL

[WZ03] Ji-Wei Wen and Li-Xin Zhang. Precise asymptotics in laws of the iterated logarithm for Wiener local time. Statistics & Probability Letters, 64(2):133–145, August 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001421.

Wang:2005:OST

[WZ05] You-Gan Wang and Min Zhu. Optimal sign tests for data from ranked set samples. Statistics & Probability Letters, 72(1):13– 22, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000039.

Wang:2006:ASC

[WZ06a] Shaojun Wang and Yunxin Zhao. Almost sure convergence of Titterington’s recursive estimator for mixture models. Statistics & Probability Letters, 76(18):2001–2006, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206001842.

Weidong:2006:SCC

[WZ06b] Liu Weidong and Lin Zhengyan. A supplement to the complete convergence. Statistics & Probability Letters, 76(7):748–754, April 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003846.

Wei:2008:MVC

[WZ08] Zhengyuan Wei and Xinsheng Zhang. A matrix version of Chernoff inequality. Statistics & Probability Letters, 78(13):1823–1825, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000461.

Wan:2009:SLV

[WZ09a] Li Wan and Qinghua Zhou. Stochastic Lotka–Volterra model with infinite delay. Statistics & Probability Letters, 79(5): 698–706, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005002.

Wei:2009:CMI

[WZ09b] Zhengyuan Wei and Xinsheng Zhang. Covariance matrix inequalities for functions of Beta random variables. Statistics & Probability Letters, 79(7):873–879, April 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005282.

Wan:2000:MME

[WZL00] Alan T. K. Wan, Guohua Zou, and Andy H. Lee. Minimax and Γ-minimax estimation for the Poisson distribution under LINEX loss when the parameter space is restricted. Statistics & Probability Letters, 50(1):23–32, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000778.

Wu:2003:ECP

[WZW03] C. Q. Wu, L. C. Zhao, and Y. H. Wu. Estimation in change-point hazard function models. Statistics & Probability Letters, 63(1):41–48, May 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203000476.

Xi:2004:SRD

[Xi04] Fubao Xi. Stability of a random diffusion with nonlinear drift. Statistics & Probability Letters, 68(3):273–286, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520400104X.

Xia:2008:TRC

[Xia08] Ye Xia. Two refinements of the Chernoff bound for the sum of nonidentical Bernoulli random variables. Statistics & Probability Letters, 78(12):1557–1559, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000114.

Xiao:2009:PDT

[Xia09] Yimin Xiao. A packing dimension theorem for Gaussian random fields. Statistics & Probability Letters, 79(1):88–97, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003490.

Xie:2008:BIB

[Xie08] Yuquan Xie. A bilateral inequality on the Borel–Cantelli lemma. Statistics & Probability Letters, 78(14):2052–2057, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000771.

Xie:2009:BIN

[Xie09] Yuquan Xie. A bilateral inequality on a nonnegative bounded random sequence. Statistics & Probability Letters, 79(14): 1577–1580, July 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001217.

Xiao:2006:MDM

[XL06] Zhihong Xiao and Luqin Liu. Moderate deviations of maximum likelihood estimator for independent not identically distributed case. Statistics & Probability Letters, 76(10): 1056–1064, May 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205004475.

Xiong:2008:SRC

[XL08] Shifeng Xiong and Guoying Li. Some results on the convergence of conditional distributions. Statistics & Probability Letters, 78(18):3249–3253, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003088.

Xiao:2009:LIL

[XL09] Zhihong Xiao and Luqin Liu. Laws of iterated logarithm for MLE of generalized linear model randomly censored with incomplete information. Statistics & Probability Letters, 79(6):789–796, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208005166.

Li:2004:CMC

[xLxZ04] Yun xia Li and Li xin Zhang. Complete moment convergence of moving-average processes under dependence assumptions. Statistics & Probability Letters, 70(3):191–197, December 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204002548.

Xie:2007:OOT

[XNF07] Min-Yu Xie, Jian-Hui Ning, and Kai-Tai Fang. Orthogonality and D-optimality of the U-type design under general Fourier regression models. Statistics & Probability Letters, 77(12):1377–1384, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001022.

Xia:2006:CAN

[XTW06] Tian Xia, Nian-Sheng Tang, and Xue-Ren Wang. Consistency and asymptotic normality of the maximum likelihood estimates in reproductive dispersion linear models. Statistics & Probability Letters, 76(11):1137–1146, June 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205004670.

Xu:2008:ESM

[Xu08] Siyan Xu. Explicit solutions for multivalued stochastic differential equations. Statistics & Probability Letters, 78(15): 2281–2292, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001223.

Xu:2008:NGV

[XW08] Li-Wen Xu and Song-Gui Wang. A new generalized pvalue for ANOVA under heteroscedasticity. Statistics & Probability Letters, 78(8):963–969, June 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003483.

Xiang:2009:COD

[XW09] Pengfei Xiang and John C. Wierman. A CLT for a one-dimensional class cover problem. Statistics & Probability Letters, 79(2):223–233, January 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003738.

Xie:2009:HDS

[XWL09] Feng-Chang Xie, Bo-Cheng Wei, and Jin-Guan Lin. Homogeneity diagnostics for skew-normal nonlinear regression models. Statistics & Probability Letters, 79(6):821–827, March 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520800521X.

Xiong:2005:SAO

[XWS05] Jian Xiong, Augustine Wong, and Donna Salopek. Saddlepoint approximations to option price in a general equilibrium model. Statistics & Probability Letters, 71(4):361–369, March 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000076.

Xu:2007:DBE

[XY07] Qinfeng Xu and Jinhong You. Difference-based estimation for error variances in repeated measurement regression models. Statistics & Probability Letters, 77(8):811–816, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207000077.

Xing:2009:MMI

[XYC09] Guodong Xing, Shanchao Yang, and Aiwu Chen. A maximal moment inequality for α-mixing sequences and its applications. Statistics & Probability Letters, 79(12):1429–1437, June 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715209001047.

Xing:2008:TRD

[XZJ08] Xiaoyu Xing, Wei Zhang, and Yiming Jiang. On the time to ruin and the deficit at ruin in a risk model with double-sided jumps. Statistics & Probability Letters, 78(16):2692–2699, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001867.

Yanagimoto:2000:PEE

[Yan00a] Takemi Yanagimoto. A pair of estimating equations for a mean vector. Statistics & Probability Letters, 50(1):97–103, October 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200000924.

Yanushkevichius:2000:SCN

[Yan00b] Romanas Yanushkevichius. Stability of the characterization of normal distribution in the Laha–Lukacs theorem. Statistics & Probability Letters, 49(3):225–233, September 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000511.

Yang:2003:UAN

[Yan03a] Shanchao Yang. Uniformly asymptotic normality of the regression weighted estimator for negatively associated samples. Statistics & Probability Letters, 62(2):101–110, April 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202004273.

Yang:2003:SLP

[Yan03b] Weiguo Yang. Some limit properties for Markov chains indexed by a homogeneous tree. Statistics & Probability Letters, 65(3):241–250, November 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002608.

Yan:2004:MII

[Yan04] Litan Yan. Maximal inequalities for the iterated fractional integrals. Statistics & Probability Letters, 69(1):69–79, August 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001609.

Yan:2008:WDC

[Yan08] Yi Yan. When do cylinder σ-algebras equal Borel σalgebras in Polish spaces? Statistics & Probability Letters, 78(10):1222–1225, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207003963.

Yao:2000:CSO

[Yao00] J. Yao. On constrained simulation and optimization by Metropolis chains. Statistics & Probability Letters, 46(2): 187–193, January 15, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001054.

Yao:2001:SIA

[Yao01] J. Yao. On square-integrability of an AR process with Markov switching. Statistics & Probability Letters, 52(3):265–270, April 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715200002066.

Yatracos:2002:AQB

[Yat02] Yannis Yatracos. Assessing the quality of bootstrap samples and of the bootstrap estimates obtained with finite resampling. Statistics & Probability Letters, 59(3):281–292, October 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715202001967.

Yuan:2009:LLN

[YAT09] De Mei Yuan, Jun An, and Bao Tao. Laws of large numbers for residual Cesàro alpha-integrable sequences under dependence assumptions. Statistics & Probability Letters, 79(5):611–618, March 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004872.

Yang:2007:NTL

[YB07] Guijun Yang and Neil A. Butler. Nonregular two-level designs of resolution IV or more containing clear two-factor interactions. Statistics & Probability Letters, 77(5):566–575, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002896.

Yin:2004:ADT

[YC04] Xiangrong Yin and R. Dennis Cook. Asymptotic distribution of test statistic for the covariance dimension reduction methods in regression. Statistics & Probability Letters, 68(4):421–427, July 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715204001324.

You:2005:THP

[YC05] Jinhong You and Gemai Chen. Testing heteroscedasticity in partially linear regression models. Statistics & Probability Letters, 73(1):61–70, June 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000994.

Yin:2006:DRM

[YC06a] Xiangrong Yin and R. Dennis Cook. Dimension reduction via marginal high moments in regression. Statistics & Probability Letters, 76(4):393–400, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003111.

You:2006:WBE

[YC06b] Jinhong You and Gemai Chen. Wild bootstrap estimation in partially linear models with heteroscedasticity. Statistics & Probability Letters, 76(4):340–348, February 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205003044.

Yeh:2002:SMZ

[Yeh02] Hsiaw-Chan Yeh. Six multivariate Zipf distributions and their related properties. Statistics & Probability Letters, 56 (2):131–141, January 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715201001493.

Yeo:2005:VST

[Yeo05] In-Kwon Yeo. Variable selection and transformation in linear regression models. Statistics & Probability Letters, 72(3):219– 226, May 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205000313.

Yeo:2007:GWA

[Yeo07]  In-Kwon Yeo.  Generalized weighted additive models

based on distribution functions. Statistics & Probability Letters, 77(12):1394–1402, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001046.

Yadegari:2008:GBS

[YGK08] Iraj Yadegari, Abbas Gerami, and Majid Jafari Khaledi. A generalization of the Balakrishnan skew-normal distribution. Statistics & Probability Letters, 78(10):1165–1167, August 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003872.

Yuan:2008:ARC

[YH08] Haili Yuan and Yijun Hu. Absolute ruin in the compound Poisson risk model with constant dividend barrier. Statistics & Probability Letters, 78(14):2086–2094, October 1, 2008.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000813.

Yang:2009:UBU

[YH09] Wenquan Yang and Yijun Hu. Upper bounds for ultimate ruin probabilities in the Sparre Andersen risk model with interest and a nonlinear dividend barrier. Statistics & Probability Letters, 79(1):63–69, January 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003453.

Yi:2005:ADC

[Yi05] Seongbaek Yi. The asymptotic distribution of the constant behavior of the generalized partial autocorrelation function of an ARMA process. Statistics & Probability Letters, 75(2):113–118, November 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002166.

Yin:2008:SCI

[Yin08] Juliang Yin. On solutions of a class of infinite horizon FBS-DEs. Statistics & Probability Letters, 78(15):2412–2419, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001429.

Yeo:2001:USL

[YJ01] In-Kwon Yeo and Richard A. Johnson. A uniform Strong Law of Large Numbers for U-statistics with application to transforming to near symmetry. Statistics & Probability Letters, 51(1):63–69, January 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200001437.

Yang:2000:FEA

[YK00] Wei-Shih Yang and David Klein. An “FKG equality” with applications to random environments. Statistics & Probability Letters, 46(2):203–209, January 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001078.

Yang:2003:CSD

[YK03] Zhenhai Yang and Samuel Kotz. Center-similar distributions with applications in multivariate analysis. Statistics & Probability Letters, 64(4):335–345, October 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001780.

Yang:2005:RAC

[YK05] Bo Yang and John E. Kolassa. A refinement to approximate conditional inference. Statistics & Probability Letters, 72(2):103–112, April 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205000209.

Yu:2007:EPC

[YKB07] Jihnhee Yu, James L. Kepner, and Brian N. Bundy. Exact power calculations for detecting hypotheses involving two correlated binary outcomes. Statistics & Probability Letters, 77(3):288–294, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002355.

Yang:2000:SLL

[YL00] Weiguo Yang and Wen Liu. Strong Law of Large Numbers for Markov chains field on a Bethe tree. Statistics & Probability Letters, 49(3):245–250, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000535.

Yan:2005:IIR

[YL05] Litan Yan and Jingyun Ling. Iterated integrals with respect to Bessel processes. Statistics & Probability Letters, 74(1):93–102, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520500163X.

Yu:2001:BQR

[YM01] Keming Yu and Rana A. Moyeed. Bayesian quantile regression. Statistics & Probability Letters, 54(4):437–447, October 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201001249.

Young:2000:NDC

[YOH00] Dean M. Young, Patrick L. Odell, and William Hahn. Nonnegative-definite covariance structures for which the blu, wls, and ls estimators are equal. Statistics & Probability Letters, 49(3):271–276, September 15, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200000572.

Yohai:2008:ORE

[Yoh08] Victor J. Yohai. Optimal robust estimates using the Kullback–Leibler divergence. Statistics & Probability Letters, 78(13):1811–1816, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208000448.

Yoo:2009:PMB

[Yoo09] Jae Keun Yoo. Partial moment-based sufficient dimension reduction. Statistics & Probability Letters, 79(4):450– 456, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208004501.

Yosef:2009:SIS

[Yos09] Arthur Yosef. Set indexed strong martingales and path independent variation. Statistics & Probability Letters, 79 (8):1083–1088, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208005786.

Yu:2009:EBP

[YP09] Jinghu Yu and Jinli Pei. Extinction of branching processes in varying environments. Statistics & Probability Letters, 79(17):1872–1877, September 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001941.

Yoo:2009:OBP

[YPD09] Jae Keun Yoo, Becky S. Patterson, and Susmita Datta. An OLS-based predictor test for a single-index model for predicting transcription rate from histone acetylation level. Statistics & Probability Letters, 79(20):2109–2114, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002466.

Yamamoto:2000:UTP

[YS00] Wataru Yamamoto and Nobuo Shinozaki. On uniqueness of two principal points for univariate location mixtures. Statistics & Probability Letters, 46(1):33–42, January 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900084X.

You:2002:JTW

[YSkPkL02] Jinhong You, Xiaoqian Sun, Wan kai Pang, and Ping kei Leung. Jackknifing type weighted least squares estimators in partially linear regression models. Statistics & Probability Letters, 60(1):17–31, November 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202002420.

Yamato:2001:OST

[YSN01] Hajime Yamato, Masaaki Sibuya, and Toshifumi Nomachi. Ordered sample from two-parameter GEM distribution. Statistics & Probability Letters, 55(1):19–27, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201001195.

Yang:2008:GMS

[YSY08] Shanchao Yang, Chun Su, and Keming Yu. A general method to the Strong Law of Large Numbers and its applications. Statistics & Probability Letters, 78(6):794–803, April 15, 2008.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003343.

Yilmaz:2008:SIW

[YT08] Mehmet Yilmaz and Birol Topçu. Stochastic inequalities for weighted sum of two random variables independently and identically distributed as exponential. Statistics & Probability Letters, 78(18):3152–3158, December 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800285X.

Yu:2007:NPH

[Yu07a] Qiqing Yu. A note on the proportional hazards model with discontinuous data. Statistics & Probability Letters, 77(7):735– 739, April 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206003439. See erra-

tum [Yu09b].

Yu:2007:LME

[Yu07b] Sheng-Hua Yu. The linear minimax estimator of stochastic regression coefficients and parameters under quadratic loss function. Statistics & Probability Letters, 77(1):54–62, January 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715206001945.

Yu:2008:ONI

[Yu08a] Lin Yu. Orlicz norm inequalities for operator-valued martingale transforms. Statistics & Probability Letters, 78(16): 2664–2670, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520800182X.

Yu:2008:MCC

[Yu08b] Yaming Yu. On the maximal correlation coefficient. Statistics & Probability Letters, 78(9):1072–1075, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003689.

Yu:2009:VST

[Yu09a] Guan Yu. Variance stabilizing transformations of Poisson, binomial and negative binomial distributions. Statistics & Probability Letters, 79(14):1621–1629, July 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001473.

Yu:2009:ENP

[Yu09b] Qiqing Yu. Erratum to: “A note on the proportional hazards model with discontinuous data” [Statist. Probab. Lett. 77 (2007) 735–739]. Statistics & Probability Letters, 79(3):414, February 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208004252. See [Yu07a].

Yu:2009:CME

[Yu09c] Yaming Yu. Complete monotonicity of the entropy in the central limit theorem for gamma and inverse Gaussian distributions. Statistics & Probability Letters, 79(2):270–274, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208003994.

Yue:2001:CRQ

[Yue01] Rong-Xian Yue. A comparison of random and quasirandom points for nonparametric response surface design. Statistics & Probability Letters, 53(2):129–142, June 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201000657.

Yue:2002:MRD

[Yue02] Rong-Xian Yue. Model-robust designs in multiresponse situations. Statistics & Probability Letters, 58(4):369–379, July 15, 2002. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715202001505.

Youndje:2006:NQE

[YV06] É. Youndjé and P. Vieu. A note on quantile estimation for long-range dependent stochastic processes. Statistics & Probability Letters, 76(2):109–116, January 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002452.

Yuan:2004:DPL

[YW04a] Ming Yuan and Grace Wahba. Doubly penalized likelihood estimator in heteroscedastic regression. Statistics & Probability Letters, 69(1):11–20, August 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001014.

Yue:2004:UTF

[YW04b] Rong-Xian Yue and Jing-Wen Wu. U-type and factorial designs for nonparametric Bayesian regression. Statistics & Probability Letters, 69(3):343–356, September 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204001865.

Yi:2006:CO

[YW06] Zhang Yi and Chengguo Weng. On the correlation order. Statistics & Probability Letters, 76(13):1410–1416, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206000484.

Yi:2007:GFT

[YW07] Yanqing Yi and Xikui Wang. Goodness-of-fit test for response adaptive clinical trials. Statistics & Probability Letters, 77(10):1014–1020, June 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207000417.

Ye:2009:IEC

[YW09] Ren-Dao Ye and Song-Gui Wang. Improved estimation of the covariance matrix under Stein’s loss. Statistics & Probability Letters, 79(6):715–721, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520800504X.

Yu:2009:SER

[YWS09] Zhuoxi Yu, Dehui Wang, and Ningzhong Shi. Semiparametric estimation of regression functions in autoregressive models. Statistics & Probability Letters, 79(2):165–172, January 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208003623.

Yan:2008:FRG

[YWWY08] Yun-Zhi Yan, Han-Xing Wang, Jun Wang, and Xiang-Feng Yin. H(n)-factors in random graphs. Statistics & Probability Letters, 78(11):1255–1258, August 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207004026.

Yun-Xia:2006:PAC

[YX06] Li Yun-Xia. Precise asymptotics in complete moment convergence of moving-average processes. Statistics & Probability Letters, 76(13):1305–1315, July 15, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520600109X.

Yakovlev:2007:ARL

[YY07a] A. Yakovlev and N. Yanev. Age and residual lifetime distributions for branching processes. Statistics & Probability Letters, 77(5):503–513, March 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206002719.

Yang:2007:SRL

[YY07b] Xiangfeng Yang and Litan Yan. Some remarks on local time-space calculus. Statistics & Probability Letters, 77(16): 1600–1607, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001356.

Yang:2008:NSL

[YY08a] Weiguo Yang and Xue Yang. A note on strong limit theorems for arbitrary stochastic sequences. Statistics & Probability Letters, 78(14):2018–2023, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000710.

Yano:2008:RDL

[YY08b] Kouji Yano and Yuko Yano. Remarks on the density of the law of the occupation time for Bessel bridges and stable excursions. Statistics & Probability Letters, 78(14):2175–2180, October 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208001077.

Yan:2008:VIF

[YYL08] Litan Yan, Xiangfeng Yang, and Yunsheng Lu. p-Variation of an integral functional driven by fractional Brownian motion. Statistics & Probability Letters, 78(9):1148–1157, July 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715207003860.

Yang:2001:DSI

[YZ01] Hailiang Yang and Lihong Zhang. On the distribution of surplus immediately before ruin under interest force. Statistics & Probability Letters, 55(3):329–338, December 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001638.

Yu:2002:SDB

[YZ02] Chang Yu and Daniel Zelterman. Sums of dependent Bernoulli random variables and disease clustering. Statistics & Probability Letters, 57(4):363–373, May 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715202000913.

Yan:2004:RIB

[YZ04a] Litan Yan and Bei Zhu. A ratio inequality for Bessel processes. Statistics & Probability Letters, 66(1):35–44, January 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203003031.

Yor:2004:RAN

[YZ04b] M. Yor and L. Zambotti. A remark about the norm of a Brownian bridge. Statistics & Probability Letters, 68(3):297–304, July 1, 2004. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715204001208.

You:2005:LIL

[YZ05] Jinhong You and Xian Zhou. The law of iterated logarithm of estimators for partially linear panel data models. Statistics & Probability Letters, 75(4):267–279, December 15, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205002403.

You:2006:ELS

[YZ06] Jinhong You and Yong Zhou. Empirical likelihood for semiparametric varying-coefficient partially linear regression models. Statistics & Probability Letters, 76(4):412–422, February 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715205003135.

Yang:2007:TTE

[YZ07a] Song Yang and Yichuan Zhao. Testing treatment effect by combining weighted log-rank tests and using empirical likelihood. Statistics & Probability Letters, 77(12):1385–1393, July 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207001034.

You:2007:TSE

[YZ07b] Jinhong You and Haibo Zhou. Two-stage efficient estimation of longitudinal nonparametric additive models. Statistics & Probability Letters, 77(17):1666–1675, November 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207001502.

Yang:2008:NSW

[YZ08] Xiao Rong Yang and Li Xin Zhang. A note on self-weighted quantile estimation for infinite variance quantile autoregression models. Statistics & Probability Letters, 78(16):2731– 2738, November 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208001934.

Yang:2009:PCP

[YZ09] Hu Yang and Zhimin Zhang. The perturbed compound Poisson risk model with multi-layer dividend strategy. Statistics & Probability Letters, 79(1):70–78, January 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003465.

Yuen:2008:RMD

[YZG08] Kam-Chuen Yuen, Ming Zhou, and Junyi Guo. On a risk model with debit interest and dividend payments. Statistics & Probability Letters, 78(15):2426–2432, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001442.

Yang:2007:CFF

[YZL07] Jianfeng Yang, Runchu Zhang, and Minqian Liu. Construction of fractional factorial split-plot designs with weak minimum aberration. Statistics & Probability Letters, 77 (15):1567–1573, September 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001319.

Yang:2008:TVA

[YZL08] Hu Yang, Zhimin Zhang, and Chunmei Lan. On the time value of absolute ruin for a multi-layer compound Poisson model under interest force. Statistics & Probability Letters, 78(13):1835–1845, September 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208000485.

Yang:2009:RPD

[YZL09] Hu Yang, Zhimin Zhang, and Chunmei Lan. Ruin problems in a discrete Markov risk model. Statistics & Probability Letters, 79(1):21–28, January 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715208003313.

Zahran:2003:GEO

[ZAC03] Alyaa Zahran and C. M. Anderson-Cook. A general equation and optimal design for a 2-factor restricted region. Statistics & Probability Letters, 64(1):9–16, August 1, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203001251.

Zaharopol:2000:FMN

[Zah00] Radu Zaharopol. Fortet–Mourier norms associated with some iterated function systems. Statistics & Probability Letters, 50(2):149–154, November 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715200000936.

Zajkowski:2009:CMS

[Zaj09] Krzysztof Zajkowski. Covariance matrices of self-affine measures. Statistics & Probability Letters, 79(8):1021–1024, April 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208005701.

Zanger:2003:CMC

[Zan03] Daniel Z. Zanger. Concentration of measure and cluster analysis. Statistics & Probability Letters, 65(1):65–70, October 15, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715203002268.

Zanger:2008:TIM

[Zan08] Daniel Z. Zanger. Talagrand’s inductive method and isoperimetric inequalities involving random sets. Statistics & Probability Letters, 78(7):861–868, May 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207002921.

Zapala:2003:CRC

[Zap03] August M. Zapala. A criterion for right continuity of filtrations generated by group-valued additive processes. Statistics & Probability Letters, 63(1):79–87, May 15, 2003. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203000555.

Zapala:2008:UMW

[Zap08] August Michal Zapala. Unbounded mappings and weak convergence of measures. Statistics & Probability Letters, 78(6):698–706, April 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003203.

Zhou:2005:LSE

[ZB05] J. Zhou and I. V. Basawa. Least-squares estimation for bifurcating autoregressive processes. Statistics & Probability Letters, 74(1):77–88, August 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205001616.

Zheng:2008:FOO

[ZB08] Haitao Zheng and Ishwar V. Basawa. First-order observationdriven integer-valued autoregressive processes. Statistics & Probability Letters, 78(1):1–9, January 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715207001629.

Zhao:2009:LRO

[ZB09a] Peng Zhao and N. Balakrishnan. Likelihood ratio ordering of convolutions of heterogeneous exponential and geometric random variables. Statistics & Probability Letters, 79 (15):1717–1723, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209001655.

Zhao:2009:SCM

[ZB09b] Peng Zhao and N. Balakrishnan. Stochastic comparison and monotonicity of inactive record values. Statistics & Probability Letters, 79(5):566–572, March 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208004677.

Zhen:2009:CTS

[ZB09c] X. Zhen and I. V. Basawa. Categorical time series models for contingency tables. Statistics & Probability Letters, 79(10):1331–1336, May 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715209000686.

Zhen:2009:ODG

[ZB09d] X. Zhen and I. V. Basawa. Observation-driven generalized state space models for categorical time series. Statistics & Probability Letters, 79(24):2462–2468, December 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003319.

Zhang:2009:ICC

[ZBZ09] H. Y. Zhang, L. H. Bai, and A. M. Zhou. Insurance control for classical risk model with fractional Brownian motion perturbation. Statistics & Probability Letters, 79(4):473–480, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004537.

Zhang:2009:NCR

[ZCCH09] Li-Xin Zhang, Wai-Sum Chan, Siu-Hung Cheung, and King-Chi Hung. A note on the consistency of a robust estimator for threshold autoregressive processes. Statistics & Probability Letters, 79(6):807–813, March 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208005191.

Zhang:2008:ERL

[ZCL08] Hong Zhang, Hanfeng Chen, and Zhaohai Li. An explicit representation of the limit of the LRT for interval mapping of quantitative trait loci. Statistics & Probability Letters, 78(3):207–213, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002155.

Zhang:2001:SEM

[ZCLH01] Hanjun Zhang, Anyue Chen, Xiang Lin, and Zhenting Hou. Strong ergodicity of monotone transition functions. Statistics & Probability Letters, 55(1):63–69, November 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715201001304.

Zhang:2004:PCB

[ZD04] Chunming Zhang and Holger Dette. A power comparison between nonparametric regression tests. Statistics & Probability Letters, 66(3):289–301, February 15, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715203003535.

Zelterman:2006:ODA

[Zel06] Daniel Zelterman. An occupancy distribution arising in a limiting dilution assay for HIV-1. Statistics & Probability Letters, 76(1):1–9, January 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205002488.

Zhang:2006:MEL

[ZF06] Chunming Zhang and Haoda Fu. Masking effects on linear regression in multi-class classification. Statistics & Probability Letters, 76(16):1800–1807, October 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001453.

Zheng:2001:FIR

[ZG01] Gang Zheng and Joseph L. Gastwirth. On the Fisher information in randomly censored data. Statistics & Probability Letters, 52(4):421–426, May 1, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000256.

Zhang:2001:SAK

[Zha01a] Li-Xin Zhang. The strong approximation for the Kesten– Spitzer random walk. Statistics & Probability Letters, 53 (1):21–26, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715200002431.

Zhang:2001:IKE

[Zha01b] Shunpu Zhang. Improvements on the kernel estimation in line transect sampling without the shoulder condition. Statistics & Probability Letters, 53(3):249–258, June 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520100013X.

Zhang:2001:QER

[Zha01c] Xicheng Zhang. On the quasi-everywhere regularity of the local time of one-dimensional diffusion process in Besov space. Statistics & Probability Letters, 54(2):161–169, September 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715201000335.

Zhang:2006:ELS

[Zha06] Junjian Zhang. Empirical likelihood for NA series. Statistics & Probability Letters, 76(2):153–160, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205002841.

Zhang:2007:SOD

[Zha07a] Xinsheng Zhang. On stochastic ordering for diffusion with jumps and applications. Statistics & Probability Letters, 77(6):614–620, March 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520600294X.

Zhang:2007:NEM

[Zha07b] Zhengmin Zhang. A note on extreme magnitudes of characteristic functions. Statistics & Probability Letters, 77(16): 1641–1643, October 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001484.

Zhang:2008:LDS

[Zha08] Tonglin Zhang. Limiting distribution of the G statistics. Statistics & Probability Letters, 78(12):1656–1661, September 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208000254.

Zhao:2009:LDI

[Zha09a] Guoqing Zhao. Lenglart domination inequalities for gexpectations. Statistics & Probability Letters, 79(22):2338– 2342, November 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003046.

Zhao:2009:EUS

[Zha09b] Huiyan Zhao. On existence and uniqueness of stochastic evolution equation with Poisson jumps. Statistics & Probability Letters, 79(22):2367–2373, November 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003083.

Zheng:2001:CFH

[Zhe01] Gang Zheng. A characterization of the factorization of hazard function by the Fisher information under Type II censoring with application to the Weibull family. Statistics & Probability Letters, 52(3):249–253, April 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520000198X.

Zhou:2009:OII

[Zho09] Qing Zhou. Optimal investment for an insurer in the Lévy market: The martingale approach. Statistics & Probability Letters, 79(14):1602–1607, July 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900128X.

Zhang:2009:SDR

[ZJ09a] Lidong Zhang and Chunmei Jiang. Stationary distribution of reflected O–U process with two-sided barriers. Statistics & Probability Letters, 79(2):177–181, January 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208003647.

Zhu:2009:MHT

[ZJ09b] Rong Zhu and Harry Joe. Modelling heavy-tailed count data using a generalised Poisson-inverse Gaussian family. Statistics & Probability Letters, 79(15):1695–1703, August 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209001576.

Zhao:2006:CLM

[ZK06] Meng Zhao and K. B. Kulasekera. Consistent linear model selection. Statistics & Probability Letters, 76(5):520–530, March 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205003342.

Zhang:2007:AEF

[ZL07a] Riquan Zhang and Guoying Li. Averaged estimation of functional-coefficient regression models with different smoothing variables. Statistics & Probability Letters, 77(4):455– 461, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002653.

Zhang:2007:CLI

[ZL07b] Rongmao Zhang and Zhengyan Lin. Chung LIL for integrated α stable process. Statistics & Probability Letters, 77 (3):295–302, February 1, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715206002367.

Zhou:2009:IMC

[ZL09] Jie Zhou and San Y. Liu. Inference for mean changepoint in infinite variance AR(p) process. Statistics & Probability Letters, 79(1):6–15, January 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208002836.

Zou:2007:ELR

[ZLQW07] Changliang Zou, Yukun Liu, Peng Qin, and Zhaojun Wang. Empirical likelihood ratio test for the change-point problem. Statistics & Probability Letters, 77(4):374–382, February 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715206002537.

Zhang:2006:FMM

[ZM06] Y. Zhang and A. I. McLeod. Fitting MA(q) models in the closed invertible region. Statistics & Probability Letters, 76(13):1331–1334, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715206000289.

Zografos:2005:ERS

[ZN05] K. Zografos and S. Nadarajah. Expressions for Rényi and Shannon entropies for multivariate distributions. Statistics & Probability Letters, 71(1):71–84, January 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520400286X.

Zhou:2008:LDA

[ZNS08] Yong-Dao Zhou, Jian-Hui Ning, and Xie-Bing Song. Lee discrepancy and its applications in experimental designs. Statistics & Probability Letters, 78(13):1933–1942, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000618.

Zhao:2002:MSU

[ZP02] Lincheng Zhao and Limin Peng. Model selection under order restriction. Statistics & Probability Letters, 57(4):301–306, May 1, 2002. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715201002000.

Zhang:2009:CHF

[ZP09] Jiajia Zhang and Yingwei Peng. Crossing hazard functions in common survival models. Statistics & Probability Letters, 79(20):2124–2130, October 15, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002491.

Zaigraev:2008:ESP

[ZPK08] A. Zaigraev and A. Podraza-Karakulska. On estimation of the shape parameter of the gamma distribution. Statistics & Probability Letters, 78(3):286–295, February 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002404.

Zhang:2006:MPU

[ZQ06] Shangli Zhang and Hong Qin. Minimum projection uniformity criterion and its application. Statistics & Probability Letters, 76(6):634–640, March 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715205003640.

Zhang:2005:LRT

[ZR05] Yanqiong Zhang and William F. Rosenberger. On linear rank tests for truncated binomial randomization. Statistics & Probability Letters, 72(1):83–92, April 1, 2005. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715205000192.

Zhang:2006:NGM

[ZR06] Haimeng Zhang and M. Bhaskara Rao. A note on the generalized maximum likelihood estimator in partial Koziol–Green model. Statistics & Probability Letters, 76(8):813–820, April 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205004165.

Zou:2009:NLD

[ZRQ09] Na Zou, Ping Ren, and Hong Qin. A note on Lee discrepancy. Statistics & Probability Letters, 79(4):496–500, February 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715208004562.

Zhou:2008:GMR

[ZS08] Weihua Zhou and Robert Serfling. Generalized multivariate rank type test statistics via spatial U-quantiles. Statistics & Probability Letters, 78(4):376–383, March 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002507.

Zuo:2001:SQR

[Zuo01] Yijun Zuo. Some quantitative relationships between two types of finite sample breakdown point. Statistics & Probability Letters, 51(4):369–375, February 15, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001772.

Zwane:2003:IPB

[ZvdH03] E. N. Zwane and P. G. M. van der Heijden. Implementing the parametric bootstrap in capture-recapture models with continuous covariates. Statistics & Probability Letters, 65(2):121–125, November 1, 2003. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715203002293.

Zhang:2001:WCN

[ZW01] Li-Xin Zhang and Jiwei Wen. A weak convergence for negatively associated fields. Statistics & Probability Letters, 53(3):259–267, June 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167715201000219.

Wang:2006:CRD

[zW06] Zhong zhi Wang. A class of random deviation theorems for sums of nonnegative stochastic sequence and Strong Law of Large Numbers. Statistics & Probability Letters, 76(18):2017–2026, December 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016771520600188X.

Zou:2007:ERC

[ZWWC07] Guohua Zou, Alan T. K. Wan, Xiaoyong Wu, and Ti Chen. Estimation of regression coefficients of interest when other regression coefficients are of no interest: The case of nonnormal errors. Statistics & Probability Letters, 77(8):803–810, April 15, 2007. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715207000065.

Zhang:2005:SEB

[ZWY05] Weiping Zhang, Laisheng Wei, and Yaning Yang. The superiority of empirical Bayes estimator of parameters in linear model. Statistics & Probability Letters, 72(1):43–50, April 1, 2005. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-

2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715205000143.

Zhao:2009:VSS

[ZX09] Peixin Zhao and Liugen Xue. Variable selection for semiparametric varying coefficient partially linear models. Statistics & Probability Letters, 79(20):2148–2157, October 15, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209002521.

Zhou:2004:WEV

[ZY04] Xian Zhou and Jinhong You. Wavelet estimation in varyingcoefficient partially linear regression models. Statistics & Probability Letters, 68(1):91–104, June 1, 2004. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000732.

Zhang:2008:PAL

[ZY08] Yong Zhang and Xiao-Yun Yang. Precise asymptotics in the law of the iterated logarithm and the complete convergence for uniform empirical process. Statistics & Probability Letters, 78(9):1051–1055, July 15, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207003641.

Zhao:2006:PHM

[ZZ06] Xiaobing Zhao and Xian Zhou. Proportional hazards models for survival data with long-term survivors. Statistics & Probability Letters, 76(15):1685–1693, September 1, 2006. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715206001234.

Zheng:2008:GET

[ZZ08] Shiqiu Zheng and Shengwu Zhou. A generalized existence theorem of reflected BSDEs with double obstacles. Statistics & Probability Letters, 78(5):528–536, April 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207002842.

Zhao:2009:SMM

[ZZ09] Xiaobing Zhao and Xian Zhou. Semiparametric modeling of medical cost data containing zeros. Statistics & Probability Letters, 79(9):1207–1214, May 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900025X.

Zhang:2006:GSD

[ZZG06] H. Y. Zhang, M. Zhou, and J. Y. Guo. The Gerber– Shiu discounted penalty function for classical risk model with a two-step premium rate. Statistics & Probability Letters, 76(12):1211–1218, July 1, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

REFERENCES 597

URL http://www.sciencedirect.com/science/article/
pii/S0167715205004773.